A Complete Bibliography of ACM Transactions on Embedded Computing Systems (TECS)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: https://www.math.utah.edu/~beebe/

02 January 2025 Version 1.95

Title word cross-reference

(m, k) [NRZ^{+}24]. 2 [LMS^{+}22, VWG^{+}17]. 3 [CCY^{+}13, CLLC17, DSXS^{+}14, HH13, HL14, LQN^{+}13, LMS^{+}22, MSCS16, NODP24, PRB15, PP24, SSPP23, SP19b, SBK^{+}23, SVC^{+}23, WDM17]. 8 [LPO^{+}17, ZSH^{+}19]. ^{2} [EAAS22]. GF(2^{m}) [HJ19]. K [KB23]. μ [LN04, WPW^{+}04]. R^{3} [WLH^{+}18].

-512 [XHL^{+}24]. -Bit [LPO^{+}17, ZSH^{+}19]. -constrained [NRZ^{+}24]. -MRAM [ZBCM09]. -Periodic [KB23]. -Re-entrant [VWG^{+}17]. -tree [LCC^{+}19, WCB20].

/Divergence [BSV17].

1 [STLX22a, SCZ20a]. 12 [BLG^{+}15]. 16-bit [KG05]. 1s [Ano13, Ano14].

2 [CBH22b, STLX22b, SCZ20b]. 2.0 [CD19]. 2011 [SN10]. 2015 [EE16, FX17]. 2016 [EH18]. 2021 [DGG24]. 25000 [AKB^{+}25]. 256 [ZSH^{+}19]. 2s [Ano13, Ano14].

32-bit [SSA21]. 3PXNet [RLG20]. 3s [Ano14].

4.0 [Shu18b]. 4s [Ano14].

512 [XHL^{+}24]. 5G

[PMS^{+}24, SAS^{+}23, VKMP20]. 5s [Ano14].

6 [GIA11]. 61499 [YRS12]. 64-bit

[KKS^{+}23]. 653 [DLD^{+}19].

802.1AS [TGP25].

A9 [SOL^{+}16]. AADL [GGGK08]. abstract [HDR^{+}06, RRW05, WBF^{+}06]. Abstraction [BBMG24, CMS17, GGMS24, KB17, LP19, MSS23, SKKR11, VF17, WMRB17, YHL23, ADI06, PDBR08, RS07]. Abstraction-Based [MSS23]. Abstraction-Refinement [KB17]. Abstractions [SPP^{+}10]. Abstracts [Ano13, Ano14, TEC12]. ACAS [CMP23]. Accelerate [XDL^{+}18, LHM14]. Accelerated [CMMD24, MMD22, RRC22, ZAL22]. Accelerating [Alg24, CDX^{+}19, DZL^{+}22, HSK18, NODP24, STLX22a, STLX22b, YZZ^{+}23, ZLSQ17, ZHCY13]. Acceleration [CAN^{+}23, DLS^{+}24, GWC^{+}24, GIA11, GV21a, HZH^{+}18, KSFS24, KSA^{+}18, LMS^{+}19, LZNL23, MDWL23, PRL^{+}23, QRW^{+}24, SLK^{+}22, SWWW17, SLZ^{+}24, THG24, WZY^{+}23, ZEJ^{+}23, ZRZ^{+}19]. Accelerator [ASS^{+}23, AP20, ARZ^{+}23, AKI^{+}23, AV20, BTA^{+}19, CS22, FFA^{+}23, GKSB25, GZZ^{+}16, HHL^{+}23a, IVJ^{+}23, JAB^{+}22, MKOA25, MSR^{+}17, SXXS^{+}16a, SXXS^{+}16b, SXMX^{+}18, Sus20, TLSJ23, VKW^{+}17, WZS^{+}24, CCA^{+}13, TLLL09]. Accelerator-Rich [MKOA25]. Accelerators [BZY^{+}23, CGSH19, DKA^{+}19, HNY18, KMS^{+}23, LG21, LPP^{+}21, LCY^{+}22, NVB^{+}20, PCC17, PVSG22, PMDC17, RR17, RWL^{+}18, TAP23, WM24, XCZW23, YHL23, ZQ24, KJRG13, MSH^{+}14]. Acceptance [JJS^{+}25]. Access [BP19, CLLC17, GDN03, IFA^{+}16, KCWH14, KKCS16, LYH^{+}15, PP19, PP24, RB21, TGBT17, WLK^{+}19, WHL23, TVK08]. Access-Aware [LYH^{+}15]. Access-Execute [WLK^{+}19]. Accesses [RC17]. Accountability [KS18]. Accounting [GD19]. accrual [WRJL06]. Accumulative [MH19]. Accuracy [JBDD20, LVSVRFCG23, MKR13, OSA^{+}18, WCK^{+}19, AC08, ITO^{+}24, PSZ12b, SD08]. Accurate [AD24b, KCJ^{+}16, TKT15, VJD^{+}07, VDK^{+}08, WSMF22, LM13, LLC^{+}13]. ACDC [SRG^{+}15]. ACDSE [FFA^{+}23]. Achieving [GHZH14, JSZ^{+}19, LPFG13, WCK^{+}19, WHL23]. ACM [BLG^{+}15, DST19, CJL17, CGZ18, DST19, SCKD23, Mit21, Shu18c]. acoustic [PSZ12b]. acoustic-based [PSZ12b]. across [GKS^{+}22, JSZ^{+}19, LBM^{+}24]. Action [KMP15, LFC17]. Activation [HZW^{+}23]. Activations [BVM19, ZRZ^{+}19]. Active [CBS19, SWS23, WDY^{+}16, XWL25, YMHB19, SPK^{+}12]. Activities [ALZR19]. Activity [AMJ21, BTA^{+}19, DBX^{+}22, HZX15, HZW^{+}23, HB23, HB24, LX22, LMB^{+}22, SG24, TLL^{+}12, WLWS15, ZRF^{+}12, HXZ^{+}13, NRL13]. actor [FZK^{+}10, LLN09, ZL08, RBS^{+}10]. Actor-based [RBS^{+}10]. actor-oriented [LLN09]. Actors [DPNA16]. actuation [PCM12]. Acyclic [ADJM19]. ad [KDN^{+}07]. ad-hoc [KDN^{+}07]. AdaFT [XKK17]. ADAPT [LCP^{+}17]. Adaptable [ARZ^{+}23, LSL20, SMW^{+}17, LB04, MSL13]. Adaptation [HKL^{+}23, MSD17, WLHC18, WLC^{+}18, ASTPH10, WYJ^{+}14, ZC04b]. Adaptations [KRS^{+}16]. Adapting [SCM20]. Adaptive [CSW15, CLL16, CYH20, CZHK23, CCC^{+}20, CAPL11, CCAP12, DBX^{+}22, DAHM16, FFA^{+}23, FPGS22, GDDD17, GRRR24, HHC^{+}16a, IDO^{+}22, KBRD22, KBD24, LCP^{+}17, MYL^{+}22, MTWE20, MALM04, MSSP22, MKD13, MF12, PMPP14, PCGD21, RJM19, SA21, SKN17, SSH14, TXL^{+}12, WLH16, XKK17, YKKD23, YYKK18, ZZG24, ZTRC03, ZLL^{+}11, ZSJ12, BO13, CWKH12, CCH13, DEG11, GMOB13, LKW02, TSG10, VSSS13, HBSA04]. AdaTest [CZHK23]. adder [DBH14]. Adders [DVC21, DNT18, FMHS23]. Address [CCC^{+}20, SEB12, CKIR06, HABT11, JKJ^{+}10, ZP08]. Address-Code [SEB12]. Addressable [RSK17, YCT16].

Addressing [ESBK23, YZA13]. adjusting [Wu10]. Adoption [NVB^{+}20]. Advanced [BP19, LAZ^{+}16, PJWY12, SXH^{+}19, BCG^{+}07, ISTE08, SBF^{+}05]. Adversarial [SG24, WZS^{+}24, XYLC23]. Adversarially [BMP23]. Adversaries [Shu19a]. Adversary [KFY^{+}22]. Adversary-aware [KFY^{+}22]. Affecting [EVS^{+}17]. Affine [KDR23, NNS13]. Against [FLF^{+}23, HMR23, HDZL20, KKL^{+}16, KH23, ARJ11, CK23, Geb04, LLZ^{+}22, RLL^{+}23, Rru22, WZS^{+}24, XYLC23]. agent [GNR^{+}10, PCBW13]. aggregates [LXK10]. Aggregation [DB19, JGX^{+}18, HBSA04, LS12]. aggregators [WSK14]. Aggressive [GDC19]. Agile [KMS^{+}23, RH23, THG24]. Agnostic [GRRR24, Sus20, ZLL^{+}11]. AHA [KMS^{+}23]. Ahead [NEP23, WZH13]. AI [LZNL23, PAF22, PKIT23, PRL^{+}23, STLX22a, STLX22b, WZG^{+}23, YSC22]. AIDA [HBSA04]. AiDe [BKMG12]. Aided [CD12, LLL14]. Air [SHL^{+}17, WLH^{+}18]. Airborne [CMP23]. ALEXIA [KRR20]. Algebra [GOC^{+}22, WMRB17]. Algebraic [AGS^{+}16, MM16]. Algebraic-Wavelet [MM16]. Algorithm [Ahm13, ADH^{+}23, BRA^{+}16, CLS16, KAKSP15, KKD^{+}12, MM16, MSR^{+}17, PJWY12, QP15, SZG^{+}23, SMZ^{+}21, WYL^{+}19, YCK^{+}18, YYKK18, CCH13, DNNP14, GNS04, LHCK04, LCH^{+}08, NBGS09, PMPP14, PL10, TJ10]. Algorithm-Hardware [ADH^{+}23]. Algorithm/Architecture [KKD^{+}12, YCK^{+}18]. Algorithmic [ORA16]. Algorithms [AMKA17, CCP^{+}19, CYH20, DLD^{+}19, FLF^{+}23, GIB^{+}12, Li21, RN14, Rat25, SSS11, SGW^{+}16, TTB23, GNW05, HABT11, PBP09a, PBP09b, ZC08]. Aligned [AKB^{+}25, HKL^{+}23]. Alignment [GW15]. all-optical [KYHY14]. Alleria [BP19]. Alloc [WDM17]. Allocation [ADJM19, CKN^{+}20, HZGW18, JLW^{+}15, LOF20, NFL^{+}22, OMH^{+}23, PCGD21, SWX17, SHQX19, WLW^{+}24, XLY18, YWLW23, AF14, ABS02, DF14, ESAS14, KK05b, LXK10, LOXL13, NDB09, PAP^{+}12, UDB06, ZZZ^{+}12, ZLF13]. Allocator [YC16]. ALOHA [Rat25]. ALOHA-FP2I [Rat25]. alteration [SKPL10]. alternative [ZNS13]. ambulatory [WYP^{+}10]. AMP [LLW^{+}24]. Amplification [HWTL23]. AMS [WZH13]. Analog [AKI^{+}23, DLS^{+}24]. Analysing [FF25]. Analysis [ARJ08, ARP12, AKD^{+}18, ABH^{+}18, AKTM16, ABS^{+}19, BVM19, BKMG12, BE17, BAG^{+}20, BDG^{+}15, BGO17, BB13, CR14, DHKS15, DHL17, DJZ13, DVCC19, DNT18, EYG^{+}23, FZK^{+}10, FMHS23, FMSS15, GWM16, GZZ^{+}16, GCU^{+}23, GUC^{+}23, HKP18, HFA^{+}14, HFL^{+}19, JDW24, KB17, LS20, LL15, LCD18, LWCL25, McI13, MHT13, MKM^{+}23a, MWK^{+}24, MAGR15, NS16, NBM^{+}16, PC14, PSD21, RLMP23, SRNW16, SE17, SC17, SR12b, SMZ^{+}21, SISS24, SLE^{+}17, SFZX18, SD17, TP19, TBEP16, VA18, VM23, WMRB17, WM24, WCM^{+}16, XZK^{+}19, YZZG23, YGW^{+}12, YKD^{+}24, ZLLC15, ZLL^{+}19, ZSJ12, AF14, ADI06, AFL13, BAR13c, BGVZ11, BC07, CMV10, CCR^{+}14, Cul13, DNNP14, GW08, GT05, GLYY14, HHB^{+}12, LLLT08, LLTL09, MEP04, MMR^{+}10, SD08, SE10, SHME13, SAMR06, SE07, TM07, VAR13, ZSM13, ZB13]. Analytic [WW09]. Analytical [BJH^{+}24, FHK21, JLSP18, MAKO19, LM13, WMZY13]. Analytics [DLPK16, HRT^{+}22, PDHC23]. analyzability [NKP^{+}12]. Analyzable [CQV^{+}13, CD12]. Analyzing [BS13b, CD19, HKVI05, JZL^{+}15, MKD15, MKE18, PP12, YZ08, YGD^{+}17]. anchors [CTK^{+}13]. Android [CSCC17, ESM^{+}17, SZL^{+}17, SKK^{+}14, STY^{+}14, WWN23, YGD^{+}19]. Annotation [AMJ21]. Anomaly [CLJ^{+}19, LL18]. Anonymous [LMW^{+}17, SBR^{+}15]. Anti [SA18, WYL^{+}19]. Anti-collision [WYL^{+}19]. Anti-Lock [SA18]. Antlab [GMS17]. ANV [NBH23]. ANV-PUF [NBH23]. Aperiodic [DSB17]. App [KJKM16, LZS20]. AppAxO [USA^{+}22]. appear [TEC12]. appearance [KMB07]. appliances [Edi14, GLWM14]. Application [AHMT17, BBM15, BO13, CCKM16, CHS15, DGC^{+}20, DASS12, DSXS15, DSW^{+}09, DLS^{+}24, ESBK23, HPBL12, KAKSP15, KJK17a, LKW02, LMA19, LX16, MSCS16, MPFG19, PSZ12a, PÖG^{+}13, RC08, RWL^{+}18, SCRY16, TBFR17, TGP25, TKHZ22, USA^{+}22, WWG^{+}18, WP11, WMLM12, ZTZ^{+}19, BM13, yCBR05, HBSA04, JHPR13, LLR14, MMR^{+}10, NSL11, XWHC06, ZNS13]. Application-adaptive [LKW02]. Application-Aware [KJK17a, BO13]. Application-centric [ESBK23]. Application-Focused [HPBL12]. application-independent [HBSA04]. Application-Level [TGP25]. Application-Specific [DASS12, MPFG19, PSZ12a, TBFR17, TKHZ22, RC08, USA^{+}22, WP11, BM13, yCBR05, JHPR13, XWHC06]. Applications [BZG19, BTA^{+}19, BJCHA17, BYIG21, CBH22a, CBH22b, CAPL11, DNBL22, DVC21, ETAV16, ESBK23, FSB^{+}21, GTH^{+}22, HJ19, KKD^{+}12, KCJ^{+}16, KGT^{+}23, KMP15, LKZ^{+}23, LPBF24, MAS^{+}24, MLR^{+}17, MBCM22, MKD15, MSSP22, MASG15, MGP25, NZCS19, PX18, PMS^{+}24, PJL^{+}17, RPHA19, RDP17, SLB^{+}15, DFC^{+}19, SPB^{+}17, TDD^{+}16, TBG^{+}17, TP16, UBF^{+}16, VCM19, WZM17, WH17, WZY^{+}23, XDL^{+}18, ZDZ14, ZSH^{+}24, ZSJ12, AMCM06, ABC^{+}07, CMV10, CLK13, CD10, CCAP12, Dea06, DKAL05, FO03, GFC^{+}10, GHB13, HHB^{+}12, IHK04, KVN^{+}09, KBDV08, KZH^{+}06, LO13, MEP04, MEP08, MAG14, DWCM14, PCK^{+}08, QP03, RMM03, SGT^{+}13, SJC^{+}03, SPP^{+}10, UCK^{+}09, YG02, YCLV^{+}02, ZNS13, ZWY^{+}10, ZXS03]. Applied [BGRV15, LCQ^{+}13]. Applying [LZJ^{+}20]. Approach [APRC16, Alg24, ABF^{+}21, BJH^{+}24, CMMD24, DMPC23, ETAV16, HDZL20, HB24, JJS^{+}25, KMS^{+}23, KDB19, LYH^{+}15, LLW^{+}17, MSRM24, McI13, NBM^{+}16, PHG^{+}17, RSW21, SUK23, SP20, SWX17, TBAS17, WZ12, YF19, ZRF^{+}12, BvB13, CAP^{+}07, CRM14, FZHT13, GNR^{+}10, JHPR13, KKH^{+}12, LLL14, LM13, MSCJ12, MSS^{+}03, OMA^{+}13, PB14, ZCS^{+}05, ZKKC05]. Approaches [CZH^{+}24, FHB^{+}17, GWM16, KOM^{+}23, PKT23, HGL14, LSC14]. Approximate [Ato20, ASJ21, AZS^{+}23, CGSH19, DVC21, DNT18, FMHS23, GRR23, LN19, LPP^{+}21, LLW^{+}24, MDS^{+}21, MMH^{+}23, NBE18, RR17, RSK17, SUK23, TAP23, USA^{+}22, YEK17]. Approximation [ADH^{+}23, BDK^{+}23, KSS24, PC14, PLT23, SC20, NBGS09, ZX08]. Approximation-Discovery [SC20]. Approximator [MGP25]. Apps [HLM^{+}24]. AQuA [PDHC23]. Arbiter [CCKM16, NBH23, RPB^{+}19, ZTZ^{+}19]. arbitrary [LA11]. Arbitration [BJP24, MBJ^{+}23, TTA^{+}20, PL10]. architectural [SPGB24, VGG^{+}13]. Architecture [ABF^{+}21, AAR^{+}17, BDB^{+}17, CHK^{+}14a, CWX^{+}23, DASS12, HW17, JDW24, KKD^{+}12, KY17, KKCS16, LWZ^{+}24, MAS^{+}24, MCSW12, MSR^{+}12, MZG15, MLAD23, OBO^{+}23, OSA^{+}18, RKM24, SK13, SJOL22, SLK^{+}22, SVS21, SC20, SSS11, TKV^{+}18, TKHZ22, TKT15, TLSJ23, VKMP20, WCB20, WZS^{+}24, YCK^{+}18, AP09, AAPN14, BCLN13, Bec09, BO13, CIC^{+}08, CIC^{+}09, DSW^{+}09, GJ13, GDN03, GM03, GLWM14, HPLD09, ISE10, KVK^{+}03, KXL10, KYHY14, KGR12, KTT13, LS09, MMSN14, MMD04, PCK^{+}08, PBP09a, PBP09b, RDM06, RMD09, SKW^{+}07, TKG13, THON12, YFPJ14, ZCK13, ZVL04]. Architecture-Aware [MZG15]. Architecture-Independent [SC20].

Architectures [AMKA17, ADH^{+}23, ARDG16, BBB16, BMP23, BJCHA17, CHS15, CDH^{+}16, DSXS15, DLPK16, FSC^{+}16, GPT^{+}23, KAS^{+}20, KOM^{+}23, LCD18, MKOA25, MG15, MBCM22, MKD15, MKAA17, MKASJ18, NASM18, NODP24, OMMK23, RDP17, SB23, SVC^{+}23, SXMX^{+}18, THA^{+}12, VCM19, WM24, WSHC14, BP14, BvB13, BMP03, BCG10, CP13a, GMOB13, HG09, IBMK10, LOG^{+}14, LWK^{+}10, LXL13, MF13, NB04, PCM12, PDBR08, SBX08, SM13a, ZTD^{+}06]. Area [AZHC19, BKMG12, BTL^{+}12, GMVV17, KSK13, MCM^{+}17, SNTÖ25, TLL^{+}12, WH17, ZJZL20, CRM14]. Area-efficient [KSK13]. Areas [PKIT23, SBB19]. ARES [ZZA^{+}22]. Arguments [DHB^{+}23]. ARINC [DLD^{+}19]. ARINC-653 [DLD^{+}19]. Arithmetic [BMS24, LS17, CMGM24, OP06, RGdZS14, TSG10]. arithmetic-level [OP06]. ARM [CYH^{+}17, DVC^{+}07, SOL^{+}16, SSA21]. ARM/Thumb [CYH^{+}17]. ARMv8 [KKS^{+}23, WCB20]. Array [EZL^{+}17, FO03, YZZ^{+}23, ZRZ^{+}19, BDP^{+}13, WL09]. Arrays [TWTH18, WWHT21, YCK^{+}18, VSSS13]. Art [Shu15b, WGP04]. Article [BLG^{+}15]. Articles [Shu18c]. Artifact [Shu18c]. Artificial [HZYJ22, Shu18b]. Artistic [SRY13]. ARX [IYL^{+}23, SJLK18]. ARX-Based [SJLK18]. ASIC [AVF^{+}09, MCM^{+}17, MKAA17]. ASIC-Based [MCM^{+}17]. ASIP [SKW^{+}07]. aspect [DRL^{+}10]. aspect-oriented [DRL^{+}10]. Assessment [DHB^{+}23, HPBL12, KAS^{+}20]. Assignment [AR14, LBS15, MBP14, MF12, PLM^{+}15, RN14, SR12a, SEB12, CKIR06, HABT11, LO13, MEP08, OOAL06, PL10, QRB10, ZZZ^{+}12]. Assisted [BKS^{+}23, AAR^{+}17, KSY17, PX18, CCY^{+}13, HLD^{+}09, LOG^{+}14, WJ17]. Association [GZZ^{+}16, YCLV^{+}02]. Association-Rule [GZZ^{+}16]. associative [LPC^{+}07]. Assorted [MS23]. Assume [NLSV^{+}19, STH17]. Assume-Guarantee [STH17]. Assumptions [PMAB19, CJMB05]. Assurance [DHB^{+}23, SSK21, RPHA19]. ASTROLABE [NBM^{+}16]. asymmetric [ESAS14, GLWM14]. Asynchronized [GCU^{+}23]. Asynchronous [BHXP19, DS24, GHR15, SLZ^{+}24, KW10, KASD07, ZM07]. asynchrony [CW14]. ATCN [BT22]. Atom [UGS^{+}21]. Atomic [YCJW24]. Attack [BPSS24, CZK^{+}22, FXP^{+}17, IYL^{+}23, IPL16, LCLW17, PS08a, YGW^{+}12]. Attack-Resilient [IPL16, CZK^{+}22]. Attack-tolerant [PS08a]. Attacker [ZJZL20]. Attacks [ARP12, BCHL19, CBRZ19, CK23, DBFH14, HMR23, HDZL20, RLL^{+}23, SE23, SGZS21, WZS^{+}24, XYLC23, ZTZ^{+}19, ARJ11, Geb04, RCDB24, WGP04]. Attention [YZZ^{+}23]. Attestation [CZ23, MGC^{+}23]. Attitude [HCS18]. Audio [TSO22, XYLC23, TKG13]. Augmentation [KML^{+}22]. AuthCropper [KLK^{+}19]. Authenticated [DS11, KLK^{+}19]. Authentication [GMVV17, SZG^{+}23, SRK^{+}18, DLN13, LN04]. Auto [RB21, SSA^{+}24a]. Auto-Scaling [RB21]. Automata [JFK15, PG24, SFB23, SK19, SH15, BS13b]. Automated [CDD^{+}07, CFGM15, CI17, FC16, LSL20, NNS13, RMK17, TBAS17]. Automatic [BF17, BZY^{+}23, CMK12, DP19, GNP06, GGJ12, HVG13, LVSVRFCG23, LLC^{+}13, POCR24, SFZX18, TM15, TFL16, VNK^{+}03, YCK^{+}18, AFG08, BAR13c, IBMK10]. Automatically [BTD^{+}18]. Automating [SVC^{+}23]. Automation [CWZ^{+}20, SVZ13, LCQ^{+}13]. automaton [TLLL09]. automaton-matching [TLLL09]. Automotive [VA18]. Autonomous [CGZ18, HXH^{+}24, ICW^{+}21, Kha13, MM16, SH23, SAS^{+}23, WMLM12, YKKD23]. Auxiliary [DL12, ZCG^{+}22]. Availability [LAB^{+}23, FF09]. Available [KCJ^{+}16].

AVB [WLW^{+}24]. average [MLL08]. Avionics [LLP^{+}21, ABC^{+}07, NKP^{+}12]. Avoid [LJLT17]. Avoidance [CMP23]. AVR [LPO^{+}17]. AVX2 [XHL^{+}24]. AVX2/ [XHL^{+}24]. AVX2/-512 [XHL^{+}24]. AVX512 [LHP^{+}23]. Aware [AMJ21, BMAB16, BZG19, BLSM19, CSM24, CWH^{+}16, COC22, DAHM16, DHL17, FSG23, FS13, FMSS15, FC16, GQC^{+}17, HGW^{+}20, HDG^{+}14, HPP17, HB16, JRSR17, JLW^{+}15, JEP16, KKD^{+}12, KJK17a, KBS17, KJK18, KRS^{+}16, KNK^{+}25, LSC19, LJP17, LKW^{+}24, LYH^{+}15, LZJ^{+}19, MSR^{+}12, MZG15, NASM18, OBO^{+}23, OHCK24, PSZ12a, PP24, PVV24, RR17, RLL^{+}23, RDSS21, SOL^{+}16, SP19b, SXXS^{+}16a, SWX17, SLS^{+}19, SAS^{+}23, TBAS17, TBEP16, TLBM15, VA18, WLWS15, WHN^{+}17, WZD^{+}17, YC16, ZZG24, ZLX^{+}23, AHM19, ABF^{+}21, ACK^{+}13, AZHC19, AZS^{+}23, BCDD24, BMP03, BO13, CCSC23, DKV14, DGC^{+}20, DLRTB^{+}19, DJS16, ESM^{+}17, EYG^{+}23, FZJ08, GHB13, GGI13, GNR^{+}10, HSD22, HH13, IVJ^{+}23, JC03, JP14, KBDV08, KYL13, KFY^{+}22, LO13, LQN^{+}13, LCC^{+}23, MS21, MSS^{+}03, MALM04, MAG14, NODP24, OMH^{+}23, OMA^{+}13, SSK23, SSA^{+}24a, SRS03, SPT^{+}21, SPT^{+}23, SR19]. aware [XSP22, YW13, ZC04a, ZSEP21, Zhu10, SAS^{+}23, LDV12, SAMR06, BJP24]. Awareness [RSW21, ZO16]. AXI [RPB^{+}19]. Axiom [TNR17]. Axiomatization [BGGT23]. AxOTreeS [SUK23].

B [DAASP21, LCC^{+}19, SAS^{+}23, WCB20, WKC07]. B-AWARE [SAS^{+}23]. B-tree [WKC07]. Backbone [PAF22]. Bad [KNY^{+}17]. Balance [JNI15, YMKH23]. balanced [WLW^{+}24]. Balancing [CGSH19, CWJ17, FS13, THA^{+}12, Mus10]. BAND [BKMG12]. BAND-AiDe [BKMG12]. Bandwidth [BCS^{+}23, DF14, FBM16, HHL^{+}23a, WLW^{+}24].

Bandwidth-limited [HHL^{+}23a]. Bank [TGBT17, LXL13, SBX08, ZP06]. Banks [CI17, MF12]. BarbequeRTRM [BMF15]. Bare [BYIG21]. Bare-Metal [BYIG21]. Barrier [HCL^{+}17, ZQD^{+}23]. Baseband [VKMP20]. Based [ADH^{+}23, ARDG16, AYS15, BCD^{+}22, BCS16, BSA17, BE17, BMS24, BPSS24, BP12, BSJ15, BRL16, CSCC17, CGM^{+}25, CPC17, CCM17, CCC^{+}20, CDH^{+}16, CKB17, DWRR14, DJZ13, EVS^{+}17, FND^{+}16, GD24, GSC19, GMCC18, HPBL12, HSMS16, HZYJ22, HPO^{+}15, HWTL23, HLLL20, HHL^{+}23b, HPS13, HW17, JKH22, JZL^{+}15, KJLS20, KAKSP15, KY17, KKL^{+}16, KH23, KCC^{+}16, KSA^{+}18, Kwo16, LL15, LPFL16, LHP^{+}23, LX22, LZL15, LJ24, LZS^{+}18, LHL^{+}19, LPO^{+}17, LZZ^{+}19, LSL20, LPBF24, MSS23, MSCS16, MCS^{+}15, MCM^{+}17, MS13b, MKD15, MSSP22, MKAA17, NASM18, NBH23, NYH^{+}20, PYJL15, PJWY12, PGR16, PNRC17, RLL^{+}23, SA18, SLB^{+}15, SJLK18, SSA21, SXH^{+}19, SWL^{+}23, SPC^{+}16, SCRY16, SJOL22, SLK^{+}22, SIC19, TBFR17, TNR17, TMXS17, TAMS18, UM13, WWY13, WDY^{+}16, WXY^{+}18, WCK^{+}19, WZY^{+}23, WLC^{+}18, WZ12, XHK16, XWL25, XDL^{+}18, YJD^{+}17, YZZG23, YC12, YLW15, YCT16, YYKK18]. Based [ZRZ^{+}19, ZCG^{+}22, ASTPH10, AÖÖ23, AP20, ANARR^{+}19, AZHC19, ABS02, BGD14, BGK^{+}23, BD14, BZ13, BDD^{+}24, BFQ10, BONA22, BMMV21, BJH^{+}24, CCA^{+}13, CSM24, CYKH13, CC13a, CDX^{+}19, CCP^{+}19, CMPP23, CGV10, DJ23, DEG11, DLN13, DAASP21, FZHT13, FMHS23, FKS^{+}19, FFA^{+}23, FLF^{+}23, GW08, GFC^{+}10, GDDD17, GPH24, GD14, GDN03, HKP18, HTB^{+}25, HZX^{+}14, HPLD09, HB23, JKJ^{+}10, JMO14, KKO^{+}06, KPK^{+}19, KKH^{+}12, KGR12, KT14, KNK^{+}25, KKS^{+}23, LCQ^{+}13, LPC^{+}07, LS13, LLR14, LC17, LLG^{+}20, LKZ^{+}23, LLW^{+}24, LWZ^{+}24, LCY^{+}22, LHCK04, LLLGR13, LV09, MGC^{+}23, MWK^{+}24, Mus10, NSL11, OMA^{+}13, PBC22, PCK^{+}08, PS08b, PW13, PDBR08, PAS^{+}09, PCGD21, PSZ12b, PGR^{+}08, PLT23, QRW^{+}24, RS07, RCDB24, SSK21, SUK23, SGT^{+}13, SCF12, SKH^{+}12, SGZS21, SBLM13, SB08, SCB^{+}22, SXMX^{+}18, SVS21, SLZ^{+}24, SC05, TXL^{+}12, TJ23, THG24, TP20, TAP23]. based [USA^{+}22, VJD^{+}07, VDK^{+}08, WM24, WSK14, XQQ^{+}24, YZZ^{+}23, YV23, YRF10, YLTY21, ZKKC05, ZJZL20, ZQ24, ZLF13, ZK25, ETAV16, GZZ^{+}16, SBDK22, CLLC17, FS14, RBS^{+}10, RSB^{+}09, ZBCM09]. Based0 [MASG15]. Bases [HWC^{+}20]. Basic [HDZL20]. basis [RMH04a]. BASS [VVKG23]. Battery [AKTM16, CGZ18, FHK21, KCJ^{+}16, LOD18, SPT^{+}21, VA18, WXY^{+}18, WLHC18, YTL^{+}20, RV03, ZSM13]. Battery-Aware [VA18, SPT^{+}21]. Battery-Free [CGZ18, LOD18, WXY^{+}18, WLHC18]. Battery-Less [AKTM16]. Battery-powered [YTL^{+}20, RV03]. battery-supported [ZSM13]. Batteryless [CAN^{+}23, GTH^{+}22, HTR^{+}16]. Bayesian [ADH^{+}23, DHJ^{+}17]. BBB [HDZL20]. BBB-CFI [HDZL20]. Be [JSD23, Val17, GT05]. beamforming [TKG13]. Become [RH23, Shu18e]. BeepBeep [PSZ12b]. Behavior [JC12, NS17]. Behaviors [BTD^{+}18]. Belief [HLLL12]. Benchmark [LWK^{+}17, MCSW12]. Benchmarked [MKAA17]. Benchmarking [KS22]. BenchPrime [LWK^{+}17]. Benders [ETAV16]. benefits [BAR13b]. Berkeley [SVP05]. best [MAG14, SRM^{+}13]. best-effort [MAG14]. Between [DPNA16, NNH^{+}14]. Beyond [KMB07, SRNW16, SGZS21]. Bicriteria [MG15]. Big [APRC16, CJL17, KSA^{+}18, PNRC17, Shu15a, Shu16c]. big.LITTLE [HTC^{+}16]. Bilinear [YLW15]. Binaries [CYH^{+}17]. Binary [CL13, CYH^{+}17, HLF^{+}18, MBR15, PBC22, PWL^{+}19, WDM^{+}23, ZSH^{+}24, ZDL22, ZGH^{+}19, BCDH12, RMH04a]. Bio [BDB^{+}17]. Bio-signal [BDB^{+}17]. Biochips [CKB17, EZL^{+}17, SIC19]. Biological [BSM^{+}21]. biology [LHM14]. Bionode [PQA^{+}19]. biopotential [CNC13]. BISDU [MKS23]. Bit [JNI15, LPO^{+}17, MKS23, PYL^{+}23, SJK20, SISS24, ZJZL20, CMGM24, GJ13, KG05, KKS^{+}23, SSA21, YZZG23, ZSH^{+}19]. Bit-Density [JNI15]. Bit-flexible [SJK20]. Bit-Serial [MKS23, PYL^{+}23]. Bit-sliced [SISS24]. Bit-time-based [ZJZL20]. BitSET [PYL^{+}23]. bitwidth [ÖNG08]. Bitwise [JDW24]. Black [BTD^{+}18, GGMS24, SOL^{+}16, SWS23]. Black-Box [BTD^{+}18, SWS23, GGMS24]. BLE [HLL^{+}23]. Blech [GGMS24]. Blind [LCLW17]. blinding [KHHH14]. Block [FMHS23, FXP^{+}17, HDZL20, KNY^{+}17, LCLW17, PBC22, SJLK18, Shu16a, Shu17b, TFL16]. Block-based [FMHS23]. Blockage [SAS^{+}23]. Blockchain [WLC^{+}22, XQQ^{+}24]. Blockchain-based [XQQ^{+}24]. Blocking [HGW^{+}20, SE17, ZC04a, DW10]. Blocking-Aware [HGW^{+}20, ZC04a]. Blocks [BYIG21, JFM23, SWK19]. Blockwise [LYY^{+}17]. BlueIO [JAD19]. Bluetooth [KYDC20, LLL14]. BMS [KNY^{+}17]. Board [CPP^{+}17, CGV10]. Boards [JKH22]. Body [AZHC19, BKMG12, BTL^{+}12, GGJ12, PP12, TLL^{+}12, ZLL^{+}11, LHX^{+}14, QRB10, WYP^{+}10]. Bonsai [SZG^{+}23]. Boosting [CMV10]. BORPH [SB08]. Both [ZWK23]. Bottleneck [Ahm13]. bound [ZX08]. Bounded [AFMT17, FS24, KHB^{+}23, KDR23]. Bounded-Rate [AFMT17]. Bounding [WZ12]. bounds [LA11, NNS13, RM10]. Box [BTD^{+}18, SOL^{+}16, SWS23, GGMS24].

Brain [KOL^{+}22]. Brain-inspired [KOL^{+}22]. Brake [SA18]. Brake-by-Wire [SA18]. BRAM [PLT23]. Branch [QZXO14, DNNP14, PO05, ZA07]. branch-and-prune [DNNP14]. Branching [FKS^{+}19, KMP15]. breadcrumb [LHX^{+}14]. Break [BVM19]. Breast [PCC17, CCC^{+}14]. Brief [BLG^{+}15]. Bringing [MMA^{+}23]. Broadcast [ANARR^{+}19, GMVV17, PGR16, SXH^{+}19, DLN13, LN04]. Broadcast-Free [PGR16]. Broken [PL13]. BTMonitor [ZJZL20]. Budget [BE17]. Budgets [ABS^{+}19]. Buffer [CSW15, DP08, WBS10, CH10, LPC^{+}07, LWB13, LO13, PMPP14]. buffer-based [LPC^{+}07]. Buffering [HC16, KB23]. buffers [IHK04]. Building [AEF^{+}14, BJH^{+}24, LML^{+}23, LWK^{+}17, PL13, SWK19]. Built [MSHS19]. Built-in [MSHS19]. burstiness [MRY^{+}10]. Bus [HH23, RPB^{+}19, SGZS21, ZCG^{+}22, BD14, LLC^{+}13, PDBR08, PL10]. Bus-based [SGZS21, BD14, PDBR08]. buses [SAYN09]. Bypass [BJP24, JHPR13]. Bypassing [ASJ21]. Byte [JDW24, YCT16, YLDM19]. Byte-Addressable [YCT16].

C [BYIG21, EAAS22, Gar05, LPD^{+}20, LL15]. C-based [RSB^{+}09]. CABARRE [MBJ^{+}23]. Cache [AHM19, ANARR^{+}19, AB15, Ato20, BHD15, BPSS24, CHK^{+}14a, CR14, CBRZ19, Cul13, CMP17, DJ23, DLD^{+}19, FF25, GWZ16, JLSP18, JLW^{+}15, KR18, Kwo16, LPB06, MSHS19, MPT^{+}22, MGB^{+}21, MBJ^{+}23, NS16, NS17, NYH^{+}20, QZXO14, RP10, SRG^{+}15, SGZS21, SWL^{+}23, SP20, SJOL22, VGN18, WMGR12, WZJ^{+}18, WKP24, XSP22, YHL23, ZW17, BJP24, BGD14, BP05, BO13, GRVD12, GLYY14, HKVI05, KVK^{+}03, LKW02, RG13, SE07, VLX07, WAD14, ZVL04, ZVN05, ZKKC05, ZTRC03, UAK^{+}03]. Cache-Based [Kwo16]. Cache-Partitioned [GWZ16]. Cache-Related [CR14, FF25]. Caches [AK21, CR14, KJK^{+}17b, KRS^{+}16, MMK22, SMR15, TTA^{+}20, GRCV03, LM13, TM07, YZ08, YFPJ14]. Caching [AK21, SVS21, WLZ^{+}23, ZLZ^{+}24, GGI13, UAK^{+}03]. CaffePresso [HSK18]. Calculating [BCD^{+}22]. Calculation [BMMV21, ZHCY13]. calculations [VLX07]. Calculi [BGRV15]. Calculus [SCG15]. Calibration [WJ17, TXL^{+}12]. Call [SN10]. calls [KMB07, KASD07]. CAM [DEG11]. Camaroptera [DNBL22]. camera [BDP^{+}13, SCF12]. camera-based [SCF12]. cameras [DZR09, LWK^{+}10]. CAMsure [RSK17]. Can [YGD^{+}19, GMVV17, PS19, SKH^{+}12, XZK^{+}19, ZCG^{+}22]. Cancer [PCC17, CCC^{+}14]. CAPA’08 [PBP09a, PBP09b]. CAPA’09 [Pla12]. capable [PMM^{+}13]. Capacity [BJH^{+}24, HLLL20, WBS10]. CAP’NN [HSD22]. Capping [PHDL18]. Captured [CMP17]. car [SCF12]. Card [SCRY16]. Cardiac [AAM^{+}17]. Cards [BSJ15]. care [BDP^{+}13]. Career [Shu18a]. CARES [SPGB24]. CARIn [PVV24]. Carnegie [KCG^{+}05]. carrier [AAPN14]. Carry [AZS^{+}23, GWM16]. Carry-aware [AZS^{+}23]. CASCADE [WLK^{+}19, BPSS24]. Cascades [BBD23]. Case [AKI^{+}23, LKZ^{+}23, LOF20, MKE18, MFG17, NS16, SRR^{+}23, WZ12, BMMV21, DEG11, FKS^{+}19, KT14, LHM14, MSS^{+}03, PE23, SKW^{+}07, SPK^{+}12, VJD^{+}07, VDK^{+}08, WEE^{+}08, YF19, YZ08]. Cash [SBR^{+}15]. Catching [SXH^{+}19]. Causality [ZL08]. Cause [DVCC19, GCU^{+}23, GTvdBC25, KHB^{+}23]. Cause-Effect [DVCC19, KHB^{+}23, GCU^{+}23]. causing [LLP^{+}17]. CCATB [PDBR08]. CDMA [PGR16]. CDMA-Based [PGR16]. CEDR [MHK^{+}23]. Cell [JNI15, LZJ^{+}20, YTL^{+}20, PJL^{+}17, SPK^{+}12, HLLL12]. Cells [PRM21]. center [BDP^{+}13]. Centrality [GAD^{+}24].

Centric [HTC^{+}16, MDWL23, MAS^{+}24, SKS21, ESBK23, LLLT08, LLTL09, SZG^{+}23]. Certificateless [ZSY19]. Certificates [BP12, HCL^{+}17, ZQD^{+}23]. CéU [SIR^{+}17]. CFI [HDZL20]. CGPredict [WZM17]. CGRA [DMPC23, WLK^{+}19]. CGRAs [KKL^{+}16, PJS15]. Chains [DVCC19, GCU^{+}23, GTvdBC25, KHB^{+}23, SE17, Shu16a, Shu17b, SWL07]. Chains-Risks [Shu17b]. Challenge [Shu19b]. Challenges [FSG23, RRM16, SKL^{+}25, DPP14, HKP08, RRKH04]. Challenging [GLYY14]. Chambolle [BRA^{+}16]. Change [AMJ21, SDMK19]. Channel [AAT^{+}21, BPSS24, BS22, BTL^{+}12, GW15, GWM16, HMLZ21, JGCS24, MM16, PX18, SBK^{+}23, SLS^{+}19, WM24, ZLSQ17, CW14, IYL^{+}23, MWK^{+}24, RCDB24]. Channel-RAID5 [PX18]. Channels [GAG15]. Characteristics [JLSP18]. Characterization [CMMD24, FHK21, VGB19]. Characterizing [GG24, SBLM13]. Charge [WDM17]. Charge-Trapping [WDM17]. Chargers [LFHS18]. Charging [LZS^{+}18]. Cheap [LPE^{+}23]. Checker [KDR23]. Checking [AD24a, RJS19, SUS^{+}17, SWS23, WZ12, CJMB05, Sch10, ZS05]. Checkpointing [ABA^{+}20, ZWK23]. checks [BCS^{+}06]. Chimp [AZHC19]. Chip [ABF^{+}21, BCHB18, BS22, CPC17, CEC23, DLPK16, DJS16, FLF17, FPGS22, FC16, GIB^{+}12, GPT^{+}23, HMR23, HMLZ21, IB23, KS18, LLG^{+}20, MST^{+}16, OMMK23, PVSG22, PSZ12a, PRK15, PGR16, SGZS21, SIC19, SR19, VKDG19, WRKG16, AKB14, BP14, BGD14, BD14, BJT^{+}23, CP13a, CHK14b, CZHK23, LJ14, GMOB13, GNR^{+}10, HXZ^{+}13, HQB06, Hüb13, ISTE08, KYHY14, KGR12, LQN^{+}13, PL10, PS10, SRM^{+}13, SJRS^{+}13b, SJC^{+}03, SAYN09, TSBY13, VNK^{+}03, WYJ^{+}14, WMZY13, XWHC06, YFPJ14, YZA13, ZRZ^{+}19, SSS11]. Chip-Free [HMLZ21]. chip-multiprocessor [PS10]. chip-multiprocessors [BD14]. Chip-to-World [SIC19]. Chiplets [SPT^{+}23]. Chips [BGK^{+}23, LX12, PL13, VKDG19]. Choices [KNK^{+}25]. CIM [HB23]. Cipher [DLS^{+}24, FXP^{+}17]. Ciphers [IYL^{+}23, KPC^{+}16, LCLW17, SJLK18]. Circuit [DLS^{+}24, IYL^{+}23, MCSW12, LLL14, ZBCM09]. Circuits [ETBK19, LAMS24, LEPP13, SWK19, SBLM13]. Cities [XWL25]. Clairvoyance [ZZG24]. Class [HSD22, BCLN13, WBF^{+}06]. Class-aware [HSD22]. Classes [LLN09, MAKO19]. Classification [GKS^{+}22, ITO^{+}24, SVS21, SRA12, LCH^{+}08]. Classifier [BBD23, ZCG^{+}22, SM13a]. Classifiers [ORA16]. Classifying [TKD07]. Cleaning [WLTW24]. Clients [GAG15]. Climate [VA18]. Clinically [FSVG19]. Clinically-robust [FSVG19]. Clock [KH23]. Clocks [HTR^{+}16]. Closed [DEG11, NZCS19, PQA^{+}19]. Closed-Loop [NZCS19, PQA^{+}19, DEG11]. Closure [SBK^{+}23]. Cloud [ALV^{+}22, CDBB24, GQC^{+}17, GRRR24, LMW^{+}17, QRW^{+}24, SSK21, SKS21, SSK^{+}22]. Cloud-Edge-End [QRW^{+}24]. Cluster [HS24, NGL17]. Clustered [BSA17, DS11, SWX17, BvB13]. Clustering [GKS^{+}22, HWTL23, HB23, LYY^{+}17]. Clustering-based [HB23]. Clusters [XZK^{+}19]. CMP [FS14]. CNN [BONA22, CS22, CMMD24, FFA^{+}23, GHD^{+}25, GRWV22, LPP^{+}21, LCY^{+}22, MSSP22, MFG17, NODP24, SPT^{+}23, WZY^{+}23, WM24, XDL^{+}18, YV23, ZRZ^{+}19, ZQ24]. CNN-Based [MSSP22, WZY^{+}23, XDL^{+}18, YV23]. CNNs [BDP24, CDX^{+}19, GG24, GKSB25, WHL23, ZDTM19]. Co [ADH^{+}23, CWX^{+}23, IVJ^{+}23, LFHS18, MBR15, MGP25, SBDK22, TBG^{+}17, TLSJ23, YCK^{+}18, JBN^{+}13, SPGB24, ST05]. Co-Approximator [MGP25]. Co-Deployment [LFHS18]. Co-Design [ADH^{+}23, TLSJ23, ST05]. Co-Exploration [IVJ^{+}23, JBN^{+}13]. Co-optimisation [YCK^{+}18]. Co-Optimization [SBDK22, CWX^{+}23]. Co-Processor [MBR15]. Co-Scheduling [TBG^{+}17]. co-simulation [SPGB24]. Coalescing [SR12a, AP09, KG05, OOAL06]. Coarse [BZY^{+}23, CGM^{+}25, JSD23, KMS^{+}23, LCD18, VNK^{+}03]. Coarse-Grained [CGM^{+}25, JSD23, KMS^{+}23, LCD18, BZY^{+}23, VNK^{+}03]. Coarsely [LM24]. Coarsely-quantized [LM24]. COBRRA [BJP24]. Code [BMS24, BGGT23, CI17, EK12, HDZL20, HTB^{+}25, HYY^{+}15, KBS17, KD08, LFC17, LBS15, LZJ^{+}19, MS21, MS23, MBFT09, OSF19, POCR24, SEB12, TP19, WKJ20, ZXS03, BAR13c, BSB14, CKIR06, CLR05, ELS08, FRRJ07, GRVD12, LLPM07, LSK^{+}08, LCS03, NP04, TBG^{+}13, YW13, ZMB03]. Code-based [HTB^{+}25]. Code-Inherent [OSF19]. Code-Level [TP19]. Code-Reuse [HDZL20]. code-size [NP04, ZMB03]. Code-size-aware [MS21]. CODEBench [TLSJ23]. Coded [ANARR^{+}19]. Codes [MBR15, LJ14]. Codesigning [SVC^{+}23]. Coding [FS13, PJWY12, KJRG13]. Coevolution [YLTY21]. Coevolution-based [YLTY21]. Coexploration [KKD^{+}12, MMD04]. Cognition [KOL^{+}22]. Cognitive [HZGW18, XLY18]. Coherence [CMP17, LPB06, YFPJ14, MMK22]. Coherent [PRSV19, YHL23, HH23]. Collaboration [QRW^{+}24]. Collaborative [AMCM06, HB16, KCCW17, LLG^{+}20, CHTC07, ZHM^{+}14]. Collaborativeness [LZJ17]. Collection [CLL16, CBS19, GMN21, KSY17, LLW^{+}17, XWL25, CKL04, CW14, CSK^{+}02, DKAL05, SP10]. Collision [CMP23, WYL^{+}19]. ColLoc [ZHM^{+}14].

Colocated [MGP25]. Colored [YKD^{+}24]. coloring [LXK10]. Combating [DBFH14]. combination [CHK14b]. Combinatorial [PYJL15]. Combining [EAAS22, GRVD12, KSS24, Mos13, RBNM19, VGN18, ZS05]. Coming [SOG15]. Comment [BLG^{+}15]. commodity [WP11]. Communication [APRC16, AZHC19, BHAC15, BLSM19, CCM17, CGZ18, FJM^{+}25, FND^{+}16, GRWV22, HYY^{+}15, LAZ^{+}16, NGL17, PE23, RJM19, SGZS21, TNR17, TKHZ22, VM23, ZDZ14, GHZH14, ISE10, KASD07, PDBR08, QRB10, SRS03, TKD07]. Communication-Aware [BLSM19]. Communications [HLL^{+}23, LOD18, PS19, AMN^{+}14]. Compact [HKL^{+}23, LTQ^{+}24, RKM24, SJLK18, Seo18, TV19, ÖNG08, ZRZ^{+}19]. Compaction [SLZ^{+}24, DVC^{+}07]. Comparative [GHPP18]. comparison [AFL13, MLV09]. Comparisons [BBB16]. Compensation [JR20, SGT^{+}13]. Competing [LMK^{+}18]. Compilation [BPP23, CHS15, MRA^{+}17, WWHT21, JMO14, PGS^{+}13, VNK^{+}03]. Compile [AC08, CWX^{+}23, NDB09, UDB06]. Compile-time [AC08, UDB06]. compile-time-unknown [NDB09]. Compiled [RG14, RMD09]. Compiler [DJO12, HLD^{+}09, KKC^{+}05, KKK^{+}11, LPD^{+}20, LJLT17, MHK^{+}23, Sus20, ZP11, AMCM06, PBV07, UAK^{+}03, ZKKC05]. Compiler-assisted [HLD^{+}09]. compiler-based [ZKKC05]. Compiler-Directed [LJLT17, KKC^{+}05]. compiler-enabled [UAK^{+}03]. Compiler-integrated [MHK^{+}23]. Compiler-Supported [ZP11]. Compilers [KMS^{+}23, RMBS20, SCZ20a, SCZ20b, GM03, LP09a, Sch07]. Compiling [CJ20, FSB^{+}21, LOF20, NP04]. Complementary [WWY13, PB14]. Complete [BLG^{+}15, BCLS17, GLT^{+}13, XHSS10].

Completion [TJ23]. Complex [SE17, STH17, MG05, VHB^{+}13]. Complexity [BJT^{+}23, FMSS15, UGS^{+}21, DRL^{+}10]. Compliant [DLD^{+}19, MWF^{+}16]. Component [ASTPH10, CKB17, HWC^{+}20, PW13]. Component-based [ASTPH10, PW13]. Components [DHB^{+}23, GSC19]. Composable [VFS^{+}21]. Composing [BCC^{+}08]. Composite [CWH^{+}23, PW13]. Composition [PRSV19]. Compositional [CMS17, ETBK19, GCU^{+}23, SL08, SLFC19, AFL13, DF14, KKH^{+}12]. Compositionality [TBG^{+}13]. Comprehension [CLW^{+}20]. Comprehensive [JFM23, LLK^{+}25, SKL^{+}25, YDS^{+}22]. Compressed [CLW^{+}20, HW17]. Compressing [LCS03]. Compression [BLSM19, EAAS22, FFA^{+}23, HKL^{+}23, ICZ^{+}23, JCS^{+}17, KSS24, KCBM21, LMS^{+}22, LLW^{+}24, MCM^{+}17, MZG15, QRW^{+}24, ZRZ^{+}19, BCS^{+}06, KD08, PZ12, YDLC10a, YDLC10b]. Compressive [KCCW17]. Computation [BFL18, CL13, DB19, DP19, FGK^{+}23, GRWV22, HLLL20, Li21, PYL^{+}23, RJM19, SA21, VM23, CAP^{+}07, HMM04, HPLD09, WBS10]. Computational [HRT^{+}22, JFM23, TBCB15, WLH^{+}18]. Computationally [BJ23]. Computations [LNA^{+}15, RLP^{+}21, RMH04b]. Compute [AGG^{+}17, PRL^{+}23]. Computer [BONA22, CD12, GK22, HJ25, LM24, FF09]. Computer-Aided [CD12]. Computers [ANB^{+}20, LP09b, SB08]. Computing [ABA^{+}20, AH13, AAR^{+}17, BFW^{+}19, BDD^{+}24, BLG^{+}15, CDBB24, CZH^{+}24, DZL^{+}22, DNT18, GLMP18, GQC^{+}17, GKS^{+}22, GCS24, HLLL20, HWC22, JBI17, KM13, KLK24, KS22, Li21, LWZ^{+}24, LWS^{+}23, LAMS24, LMW^{+}17, LLK^{+}25, MCP17, MMH^{+}23, Mit21, MEK^{+}22, OHCK24, PL13, PJWY12, RB21, RHG^{+}12, SSA^{+}24b, Shu16c, Shu18a, Shu18e, Shu19b, Shu19c, Shu20a, SP12, TP16, WX17, WHL23, YS23, YEK17, ZLZ^{+}24, ZAL22, DKV14, FZJ08, JC03, JGD^{+}09, LS09, MMSN14, MSS^{+}03, MB10, PGS^{+}13, PBP09a, PBP09b, TSWL10]. Computing-Based [PJWY12]. Concentration [BCHB18, EHF24]. Concentration-Resilient [BCHB18]. Concepts [MBCM22]. Concern [Shu18e]. Concurrency [BBM15, CFGM15]. Concurrent [BVM19, GHR15, GKSB25, JZL^{+}15, LMBL21, SPB^{+}17, JM06]. Condensed [XYLC23]. Conditional [CLJ^{+}19]. Conditions [ARS16, RKC^{+}22]. Conduction [AAM^{+}17]. Conference [DST19]. Configurable [CVG^{+}13, JGCS24, LVSVRFCG23, LLP^{+}17, OP06, PW13, PBP09a, ZVN05, PBP09b]. Configuration [FC13, GPB^{+}17, SL16, SSS11, TR24, GRVD12]. Configurations [BCS16, JHPR13]. Configured [PMS^{+}24]. Configuring [BLG^{+}15, KS22, BHET04, GLT^{+}13, SPGB24, PBP09a, PBP09b]. conflict [ZCK13]. conflict-free [ZCK13]. Conflicts [LZS20, TGBT17]. Confluence [Shu18b]. Conformance [WLT12]. Congestion [DGC^{+}20, KYL13]. Congestion-aware [DGC^{+}20, KYL13]. Connected [RN18, XDL^{+}18, Bec09]. Connectivity [ZLZ^{+}24, GDN03, KDN^{+}07]. Connex [Sus20]. Connex-S [Sus20]. Conquer [CJL17, CWJ17]. Conservation [KNL12]. Conserving [MRY^{+}10]. considering [ZNS13]. Consistency [AbSZ^{+}19, LZZK25, LAB^{+}23, LLN^{+}14, SB23]. consortium [HKLH05]. Constrained [AV20, BSJ15, EHF24, GLMP18, GWC^{+}24, JGL21, KPS23, KKCS16, LWB18, MFG17, MPFG19, Bar13a, KAK05, LQN^{+}13, LCC^{+}19, NRZ^{+}24, TSG10, UCK^{+}09, WBS10, YRS12, ZBG20]. Constraint [COC22, PVV24, ZSH^{+}19, BvB13, HCQ^{+}14, RS07]. Constraint-Aware [PVV24]. Constraints [CCKM16, LN19, MBKF15, NZCS19, PSZ12a, SB23, CCB^{+}06, HLD^{+}09, KDN^{+}07, LSK^{+}08, MBFSV07, MEP08, NP04, PAP^{+}12, RMM03, SRM^{+}13, WRJL06, YRF10]. Construction [JAB^{+}22]. Constructive [CMPP23, SMR^{+}18]. consumer [RV07]. Consumption [ANB^{+}20, FLF17, MV16, MSRM24, OBSO16, YCT16, Mus03]. Contactless [QWY^{+}18]. Containerized [BCDD24]. Containers [SCA^{+}24]. Content [CWH^{+}16, DLD^{+}19, RSK17, TLLL09]. Content-Addressable [RSK17]. Contention [AD24b, KBRD22, LES14, LCL^{+}19, RDP17, SP20, ZLX^{+}23, DNNP14, BJP24]. COntention-aware [BJP24]. Contention-Detectable [LCL^{+}19]. Contention-free [LES14]. Contests [Shu19b, WXY^{+}17]. Context [AMJ21, LS20, WYS^{+}13]. Context-Aware [AMJ21]. Contextual [KH18, KP13]. Contextualized [YSC22]. Continual [LX22, LJ24]. Continuing [Shu17a]. Continuous [DLRTB^{+}19, ZQD^{+}23]. Continuum [CDBB24]. Contract [LPFL16, PRSV19]. Contract-Based [LPFL16]. Contractions [KRHC20]. Contracts [NLSV^{+}19]. Contrastive [SRB23]. Control [BMF15, BF17, BHL^{+}20, BYIG21, DSB17, DHL17, GDDD17, GDD20, KKCS16, LJP17, LML20, MBP14, MCG22, MMY^{+}19, MBLA16, PP19, PMP17, RJS19, RLMP23, SSD^{+}19, SE23, SUS^{+}17, SPK^{+}12, SLFC19, TR24, TBCB15, TCD^{+}19, TFL16, VA18, VM23, WZH13, YKD^{+}24, ZW13, BMM13, BJM13, CAP^{+}07, FC13, KKH^{+}12, KT14, LK10, MTL14, PCM12, RV07, SWT^{+}14, VAHC^{+}06, VGG^{+}13, ZTRC03]. Control-Flow [DHL17, PMP17, SUS^{+}17, BHL^{+}20, MCG22]. Control-theoretic [SPK^{+}12]. control-theoretical [MTL14]. control/data [VAHC^{+}06]. control/data-flow [VAHC^{+}06]. Controlled [BCS^{+}23, HFL^{+}19, JNI15, WMLM12, YDLC10a]. Controller [GAG15, GMVV17, HDG^{+}14, HPP17, MSS23, NZCS19, ZJZL20, LCQ^{+}13]. Controllers [ARDG16, BF17, BDG^{+}15, GHPP18, HKP18, ICW^{+}21, KML13, NPAG12, SVZ13, YF19, KASD07]. Converging [Gar05]. conversion [AC08]. convex [SJRS^{+}13a]. Convolution [AP20, AABG22, MDWL23]. Convolutional [AP20, HSK18, HY22, MPFG19, NHS20, PYL^{+}23, ZQ24, KSK13]. Cool [UAK^{+}03]. Cool-Cache [UAK^{+}03]. Cooperation [LOD18]. Cooperative [ANARR^{+}19, SHL^{+}17, YLTY21, ZZX^{+}15]. Coordination [LBM^{+}24, PMDC17]. Coprocessor [LRZ16, BZ13]. coprocessors [HMMA04]. copy [AP09]. Core [CLJ^{+}19, HSMS16, HH23, KR18, LKA^{+}18, MKD15, PGR16, RC17, RWL^{+}18, RJM19, SSPP23, SDBD18, SRR^{+}23, TKV^{+}18, TGTT17, VKDG19, VCM19, WHN^{+}17, ACK^{+}13, CCC^{+}14, CLLC17, CMP^{+}07, DPP14, DP19, JAD19, LKB14, LOG^{+}14, LLR14, LLLT08, LLTL09, LOF20, MG05, Mus10, PMM^{+}13, PHG^{+}17, RDP17, VKMP20, WBF^{+}06, XSP22, YFPJ14]. core-centric [LLLT08, LLTL09]. Core-Memory [SSPP23]. Cores [HA24, KRS^{+}16, TDD^{+}16, TMXS17, TAMS18, VSD^{+}17, yCBR05, SM13a]. CORIDOR [MMK22]. Coroutines [BHXP19]. Correct [ARDG16, LPFL16, PB14]. Correction [CGSH19, FND^{+}16]. Corrections [FHB^{+}17]. Correctness [ZWK23]. Correlation [GW15, SMZ^{+}21, TBEP16]. Correlation-Aware [TBEP16]. Correlations [HC16]. Cortex [SOL^{+}16]. Cortex-A9 [SOL^{+}16]. Corunner [AKD^{+}18]. Corunner-Dependent [AKD^{+}18]. Coscheduling [LK10]. cosimulation [OP06]. COSMOS [PMDC17]. Cost [ABC^{+}17, BLG^{+}15, CS22, GAS^{+}17, LLC^{+}22, LLZ^{+}17, LZZ^{+}19, MGLP19, ZO16, CCH13, CRM14, EHF24, GLT^{+}13, Mus10, SJRS^{+}13a, SM13b, YFPJ14, ZCK13, ZP09]. Cost-Effective [BLG^{+}15, GLT^{+}13, Mus10]. Costs [CGSH19]. cosupplied [MKD13]. cosynthesis [KBDV08]. COTS [FSB^{+}21, HH23, PSZ12b, PJT^{+}23]. COTS-Coherent [HH23]. Count [SIC19]. Counter [ARP12, KJLS20, MKASJ18, PMAB19]. Counter-Based [KJLS20]. Counter-Examples [PMAB19]. Counterexample [LP19, ZQD^{+}23]. countermeasure [Geb06]. Counters [CZ23]. Counting [PKIT23]. Coupled [WWHT21]. Coupling [CMGM24]. Course [Shu17a]. Coverage [HSR18, SHK^{+}19, YGHS08]. CPS [DCZB19, LML^{+}23, Rru22, SSB24, TBCGO23]. CPU [BBL09, ISE10, LWB18, OFA^{+}15, PHDL18, PDL21, PDL^{+}23, RC17, DFC^{+}19, SPB^{+}17]. CPU/GPU [OFA^{+}15]. CPUs [LSC19]. Crab [WCB20]. Crab-tree [WCB20]. Crash [WCB20]. Create [CD12]. Creating [HWC^{+}20, YMHB19]. Creation [SL16]. Crenel [LZL15]. Crenel-Interval-Based [LZL15]. CRIMP [HKL^{+}23]. Crisis [Shu20b]. Criteria [SHK^{+}19, CAA^{+}24]. Critical [BHL^{+}20, CKN^{+}20, HSR18, IPL16, KWKP23, LS20, RHG^{+}14, Shu15d, WKP24, ZYL^{+}17, ASTPH10, PJL^{+}14, SVN04]. Criticality [AKTM16, BCDD24, GE18, HPP17, HHC^{+}16a, LCP^{+}17, LH18, RC17, TSP15, TGTT17, ZZG24, ZGZ15, ABS^{+}19, FHB^{+}17, HGL14, KGT^{+}23, LDRM12, ZQGZ22]. Criticality-aware [BCDD24]. Cropper [KLK^{+}19]. Cross [BDG^{+}15, JCW^{+}16, SRNW16, WWN23, ZP09, KST^{+}12]. Cross-Layer [BDG^{+}15, JCW^{+}16, ZP09, KST^{+}12]. Cross-Platform [WWN23]. Cross-Section [SRNW16]. Crossbar [HKL^{+}23, JR20]. Crossbar-Aligned [HKL^{+}23]. Crossbars [BDD^{+}24]. Crosstalk [FC16, LPE^{+}23]. Crosstalk-Aware [FC16]. Crowd [DBFH14, PKIT23]. Crowd-Sourced [DBFH14]. Crowdsensing [XQQ^{+}24]. Cryptographic [AMKA17, ARH^{+}18, BCHL19, BSJ15, EKAK24, MKAA17, ZSY19, RMH04b]. Cryptography [BDR24, DZL^{+}22, HTB^{+}25, LHP^{+}23, LWHS17, LPO^{+}17, MWK^{+}24, NVB^{+}20, SOG15, Seo18, SAKH20, Geb04]. CRYSTALS [JGCS24]. CS [KSA^{+}18]. CS-Based [KSA^{+}18]. CSD [BJH^{+}24]. CSD-based [BJH^{+}24]. CSDF [KB23]. CSI [QWY^{+}18]. CSP [Gar05, McI13]. CUDA [DLV16, KS13, PGS^{+}13]. CURA [LKH16]. CURE [NGL17]. current [MG05]. curriculum [CSVA^{+}05, Sev05, SBF^{+}05]. Curve [DZL^{+}22, LWHS17]. curves [BSKB^{+}09, WPW^{+}04]. Custom [KAKSP15, LPD^{+}20, TKG13, HVG13, LSC14, ÖNG08]. Customizable [TKV^{+}18]. customization [CGV10, PO05, ZP09]. Customized [Rru22, YTL^{+}20]. Cutting [AR14]. CV [PRB15]. CxDNN [JR20]. Cyber [AFS^{+}13, BHAC15, BKMG12, CKGN14, DWRR14, DHJ^{+}17, DHF18, GCJD20, GSN21, HZX15, IPL16, KCC^{+}16, LAB^{+}23, LWZ^{+}16, LLN^{+}14, MBKF15, MKS^{+}17, NLSV^{+}19, PRS^{+}17, SHL^{+}17, Shu16d, Shu17b, Shu19b, Shu19d, SMR20, TGV12, TCD^{+}19, UGS^{+}21, WDY^{+}16, WZBP19, XKK17, ZYM16, ZYL^{+}17, ZJC^{+}17, BWS14, BJM13, DDG^{+}13, GMOB13, Hüb13, LDRM12, SPK^{+}12, TXL^{+}12, WLT12, YRS12, ZSM13]. Cyber-Physical [AFS^{+}13, BHAC15, BKMG12, CKGN14, DWRR14, DHJ^{+}17, DHF18, GCJD20, GSN21, HZX15, IPL16, KCC^{+}16, LAB^{+}23, LWZ^{+}16, LLN^{+}14, MBKF15, MKS^{+}17, NLSV^{+}19, PRS^{+}17, SHL^{+}17, Shu19d, TGV12, TCD^{+}19, WDY^{+}16, WZBP19, XKK17, ZJC^{+}17, SMR20, BWS14, DDG^{+}13, Hüb13, LDRM12, SPK^{+}12, TXL^{+}12, WLT12, YRS12, ZSM13]. Cyber-Physical-Social [ZYM16, ZYL^{+}17]. Cybersecurity [Shu15a]. Cycle [LS12, HHB^{+}12]. Cyclo [DHKS15, SLCS16]. Cyclo-Static [DHKS15, SLCS16].

D [HL14, CCY^{+}13, CLLC17, DSXS^{+}14, HH13, LQN^{+}13, LMS^{+}22, MSCS16, NODP24, PRB15, PP24, SSPP23, SP19b, SBK^{+}23, SVC^{+}23, SRK^{+}18, WDM17]. D-PUF [SRK^{+}18]. D-Stacked [SSPP23]. D/ [LMS^{+}22]. DaCapo [KPS23]. DAG [BGS^{+}18, CLJ^{+}19]. DAGs [CAA^{+}24]. DASS [MLAD23]. Data [APRC16, AMJ21, AbSZ^{+}19, ASD^{+}24, BGJ17, BDD^{+}24, CJL17, CZK^{+}22, CBS19, CMPP23, DBFH14, FSC^{+}16, GQC^{+}17, GSS^{+}18, HKC18, HRT^{+}22, HWC^{+}20, HB23, HB24, JRR16, JCS^{+}17, JLW^{+}15, KK05a, KSA^{+}18, LPD^{+}20, LLZ^{+}22, LCC^{+}23, LCJ13, LLN^{+}14, LLW^{+}17, LSL20, MM16, MMH^{+}23, MF12, OHCK24, PP24, PE23, PqBM^{+}15, PM19, PNRC17, RP03, SMW^{+}17, SRG^{+}15, SPC^{+}16, SPT^{+}23, SZL^{+}17, SFCW23, Shu15a, SWWY13, SWWW17, VLX07, WKJ20, WWTSM19, WLC^{+}22, WQGR22, WLK^{+}19, XWL25, XQQ^{+}24, YZZG23, YCK^{+}18, YHL23, ZZX^{+}15, ZW17, BS13a, CC13a, HBSA04, HKVI05, LXK10, SLZ^{+}24, SAYN09, TBG^{+}13, UAK^{+}03, ZKKC05, ZLF13]. Data-Adaptable [LSL20, SMW^{+}17]. Data-Cache [ZW17]. Data-Dependent [HKC18]. Data-Driven [BGJ17]. Data-flow [CMPP23, VAHC^{+}06]. Data-to-Memory [FSC^{+}16]. Databases [KCC^{+}16, LZZK25, CH10]. Dataflow [ABH^{+}18, ADJM19, BPP23, DKA^{+}19, DHKS15, DPNA16, ETBK19, FGK^{+}23, GTH^{+}22, KAKSP15, LWB18, MS21, MDWL23, MKD15, NODP24, SB23, DFC^{+}19, SCB^{+}22, SLCS16, WM24, YLTY21, FZHT13, Gei10]. Dataflow-aware [NODP24]. Dataflow-based [SCB^{+}22, WM24]. Dataflow/von [SB23]. datapath [HMMA04]. DC4CD [GLMP18]. DCA [KCCW17]. DCT [HPLD09]. Dead [TM15]. Deadline [COC22, HQE20, OHCK24, MEP08, SN10]. Deadline-Aware [OHCK24]. Deadlines [YKKD23]. Deadlock [BSV17, DGC^{+}20, HPS13, LX12, WZH13, ZW13, BSV17]. Deadlock- [BSV17, BSV17]. Deadlock-/Divergence [BSV17]. Deadlock-free [DGC^{+}20]. Dealing [RSF20]. debug [AKB14]. Debugger [MZG14]. Debugging [DHF18, FLF17, MBLA16, UM13]. Decade [SOG15]. Decentralized [BRR19]. Deception [Rru22]. Decision [BCD^{+}22, BGK^{+}23, CL13, CSH^{+}22]. Decisions [PWL^{+}19, SPGT19, UDB06]. Declarative [OSA^{+}18]. Decoded [GGI13]. Decoder [FS13, SHME13]. Decoding [RRC22, WZD^{+}17, LJ14, HE12]. Decomposition [ETAV16, KM09, WL09]. Decompositions [LPFL16]. decompression [CLR05]. Decoupled [WLK^{+}19]. Deduplication [CCSC23]. Deep [AABG22, ALV^{+}22, AD24b, BJ23, BLSM19, CWZ23, CK23, CMS08, DBX^{+}22, GVS^{+}20, GOC^{+}22, HY22, HZW^{+}23, JR20, JBDD20, JKH22, KML^{+}22, LLK^{+}25, MYL^{+}22, MTWE20, MPFG19, PVSG22, PBC22, RKM24, SJK20, SA21, SHK^{+}19, TP20, VVKG23, YMKH23, YLDM19]. DEEPEYE [CLW^{+}20]. Deeply [CLW^{+}20]. defect [LLR14, VSSS13]. defect-tolerant [LLR14, VSSS13]. defending [ARJ11]. Defense [WDY^{+}16, XYLC23]. Defensive [WZS^{+}24]. Deferred [DBM^{+}15]. defined [LJR12, VKMP20]. Deformable [DDC^{+}24]. deformation [MMSN14]. Degradation [GSC19, ZZG24, RGdZS14]. deinterleaver [KSK13]. Delay [CCKM16, CR14, FF25, KJK18, LLT^{+}17, CLK13, GNS04, KAK05].

Delay-Aware [KJK18]. delay-constrained [KAK05]. Delayed [IAS23]. Delays [CZK^{+}22, GRWV22, RDP17]. Deletion [LLC^{+}22, SZL^{+}17]. delivery [LHX^{+}14]. Demand [CCC^{+}20, KKK^{+}11, ANARR^{+}19, HRH^{+}22, WSK14]. Demand-Based [CCC^{+}20]. Demonstrated [CBS19]. Demonstration [LKZ^{+}23]. Demystifying [ANB^{+}20, SUS^{+}17]. Dense [LMS^{+}19, WZS^{+}24]. Density [JNI15, YCK^{+}18]. dependability [CMV10]. Dependable [BDP^{+}13, Zhu10]. dependence [SWL07]. Dependencies [CAP15, LCS03]. Dependency [SWWW17]. Dependent [AKD^{+}18, HKC18, ABS^{+}19]. Depletion [FHK21]. Depletion-time [FHK21]. Deploying [YLDM19]. Deployment [GG24, LFHS18, LCY^{+}22, MSS23, OBO^{+}23, RIMS21, RWL^{+}18, SRA12, CGV10, ZC04c]. Depth [LG21, KTT13, LYL13]. Deriving [WWTSM19]. Descent [XCZW23]. Description [Alg24, GI24, MMD04]. Descriptor [PRB15]. Design [ABL^{+}20, AHMT17, ADH^{+}23, ARDG16, AV20, BKMG12, BBM15, BJT^{+}23, BTL^{+}12, BHET04, BRL16, DCZB19, DGG24, DST19, DJJ^{+}19, DEG11, DJZ13, DLS^{+}24, DNT18, FMHS23, FFA^{+}23, FSVG19, GLP^{+}11, GK22, GHD^{+}25, Geb04, GCJD20, GV21b, GSN21, HJ25, HFA^{+}14, IT16, IAS23, JFM23, JBDD20, JEP16, JBCS16, KJRG13, KMS^{+}23, KNK^{+}25, KB17, KDB19, LS20, Leo18, LEPP13, LMW^{+}17, LPBF24, LV09, MSHS19, MFG16, MSCS16, MYL^{+}22, MPZS13, NYH^{+}20, NLSV^{+}19, OP06, OFA^{+}15, PSZ12a, PDL^{+}23, PRSV19, POCR24, PGR16, PHG^{+}17, RLMP23, SWK19, SIR^{+}17, SLB^{+}15, SFCW23, Shu15d, Shu18e, SPGT19, SBDK22, SMR20, SLFC19, SCS16, SR19, TP19, TSP15, TBAS17, TKHZ22, TKL^{+}15, TLSJ23, TBCGO23, UGS^{+}21, VA18, VP16, WWY13, WLZ^{+}23, WWG^{+}18, WMLA16, WZG^{+}23, YGHS08, YLTY21, ZYM16, ZYL^{+}17, ZBG20, ZO16, ARJ08, BGD14, BE10, BMP03]. design [BFQ10, CYKH13, CMS08, CCB^{+}06, DPP14, DDG^{+}13, DBH14, GLC07, GNR^{+}10, HQB07, HMMA04, HHB^{+}12, JM06, JBN^{+}13, KKO^{+}06, KM09, KKH^{+}12, LAN06, LSK^{+}08, LLN09, LM13, LHM14, MSCJ12, MBFSV07, PGR^{+}08, RP03, RSB^{+}09, RI04, RAK14, SVP05, ST05, STW13, SM13a, WCJ07, XWHC06, ZTRC03, CMP^{+}07, RRKH04]. Design-Level [TP19]. Design-space [MPZS13, BFQ10]. Design-Space-Exploration [GCJD20]. Design-Technology [SBDK22]. Designation [GAD^{+}24]. designed [ZWY^{+}10]. Designing [BRL16, DQ14, SRM^{+}13, USA^{+}22, VHB^{+}13]. Designs [CJL17, IYL^{+}23, JAB^{+}22, LN19, TAP23, HH13]. desynchronization [GNP06]. Detailed [DLV16, ZLL^{+}18]. Details [HKP18]. Detectable [LCL^{+}19]. Detecting [CCP^{+}19, CMP17, PMP17, HT06]. Detection [AMKA17, AMJ21, BPSS24, CLL21, CZHK23, EVS^{+}17, FGL^{+}19, GPH24, GLS^{+}23, HA24, HZYJ22, HMLZ21, HPS13, KJLS20, LX12, LMS^{+}22, LJ24, LHYQ18, LWCL25, LJLT17, LLP^{+}17, LL18, MYL^{+}22, MKM^{+}23b, MKAA17, MKASJ18, MMD22, MAGR15, PCC17, PKIT23, QWY^{+}18, RCS23, SXH^{+}19, SMZ^{+}21, TMXS17, WDY^{+}16, YZZG23, YHL23, YKK^{+}13, ZCG^{+}22, ZJZL20, CCC^{+}14, HLD^{+}09, KLC^{+}10, KTT13, LHCK04, MVS^{+}13]. Detector [LZS20, TP16]. Determinant [LZZ^{+}23]. Determinism [Lee21]. Deterministic [GDA13, LMBL21, LBM^{+}24, SC05]. Development [CWZ^{+}20, MKMGS18, Mos13, RH23, SH23, DSW^{+}09, PJL^{+}14]. Device [ALZR19, ALV^{+}22, CFXY17, JCW^{+}16, KPS23, LHYQ18, MM16, POCR24, SRK^{+}18, WXY^{+}18, WWT^{+}22, WGN23, WT15, YTL^{+}20, ZSH^{+}19, BMM13, NRL13, PJL^{+}14, RV07, RALR24, RBNM19, SKPL10, SC05]. Device-Free [LHYQ18, WXY^{+}18]. Device-to-Device [JCW^{+}16]. Devices [AV20, BKMG12, BRA^{+}16, BPSS24, CSCC17, CJL17, CLW^{+}20, CSC23, GLMP18, GRWV22, GAS^{+}17, GDB22, GSN21, GP23, GMCC18, HTR^{+}16, HY22, HTC^{+}16, HLLL20, HLL^{+}23, ICZ^{+}23, JGL21, JRSR17, KRHC20, KKCS16, KNY^{+}17, Kwo16, LMA19, LLP^{+}21, LSL^{+}23, LWHS17, LNA^{+}15, MFG16, MPT^{+}22, MV16, MMA^{+}23, MFG17, PVV24, RCS23, RSW21, Shu17c, TP19, TP20, WLH16, YJD^{+}17, CHCC13, CMS08, LCJ13, NNH^{+}14, PSZ12b, RC08, TSWL10]. Dew [ZLZ^{+}24]. DFA [WH17]. DFA-Resistant [WH17]. DFR [IAS23]. DFSynthesizer [SCB^{+}22]. DIAC [LZS20]. Diagnosability [GHKS15]. Diagnosis [GZZ^{+}16, KH23, AKB14]. Diagonal [CKB17]. Diagrams [BCD^{+}22, CL13, TFL16]. didactic [GT05]. Dies [MASG15]. Different [HCS18]. Differentiable [MLAD23]. Differential [FXP^{+}17, HVG13]. Digit [AJ18, RMH04a]. digit-serial [RMH04a]. Digital [AYS15, BCHB18, EZL^{+}17, HPO^{+}15, IAS23, LMB^{+}22, JMO14]. Dilithium [LTQ^{+}24, RCDB24, SISS24, XHL^{+}24]. Dimensional [WWTSM19, WL09]. Direct [ZP08, LP10, SPK^{+}12]. Directed [ADJM19, BGGT23, LJLT17, QZXO14, KKC^{+}05]. Directions [MBCM22, HKP08]. DirectNVM [ZAL22]. Disassembly [ITO^{+}24]. Disclosure [FLF^{+}23]. Discovery [LAZ^{+}16, SC20]. Discrete [KL13, NDZ13, BBL09, TSCC05]. discrete-time [TSCC05]. Discussion [FHB^{+}17]. DISE [CLR05]. Disease [LJ24]. Disjunctive [AGG^{+}17]. Disks [WLTW24, CCH13, CW14]. Disorder [HZYJ22]. Disparity [LKA^{+}18, TKT15]. Display [MH19, Dea06]. Displays [LKH16]. dissemination [KAK05]. Distance [CLS16]. Distill [MPFG19]. Distill-Net [MPFG19]. Distillation [MPFG19]. Distortions [HCS18]. distributable [CRAJ10]. Distributed [BHAC15, BWS14, BZG19, BLSM19, CJL17, CZH^{+}24, DVCC19, GLMP18, GCU^{+}23, HRT^{+}22, KSS16, Kha13, LAB^{+}23, LC17, LLW^{+}17, MSM21, REPL15, RDSS21, SE23, SLB^{+}15, SDBD18, Shu16a, SHQX19, TGV12, TGP25, TAMS18, YMHB19, BVGVEA10, CRAJ10, JGD^{+}09, LWK^{+}10, LN04, MSCJ12, PS08b, PEP05, SAHE04, YGHS08, YFPJ14, ZZZ^{+}12, ZLF13, ZC04c]. distribution [ZCG^{+}22]. Disturb [LLZ^{+}22]. Disturbance [YJD^{+}17]. Disurbance [MMK22]. Divergence [BSV17]. Diversity [RN18]. Divide [CJL17, CWJ17]. Divide-and-Conquer [CJL17, CWJ17]. DL [LCY^{+}22]. DL-RSIM [LCY^{+}22]. DLIC [GGI13]. DLSeF [PNRC17]. DLSpace [SHQX19]. DM [WLTW24]. DM-SMR [WLTW24]. dMazeRunner [DKA^{+}19]. DMR [LJLT17]. DMS [BSA17]. DMS-Based [BSA17]. dMTS [BSV17]. DNA [LLC^{+}22]. DNN [BJ23, GK22, GWC^{+}24, GRRR24, HHL^{+}23a, HKL^{+}23, IVJ^{+}23, JSZ^{+}19, PP24, PVV24, QRW^{+}24, SLK^{+}22, TTB23, WZS^{+}24]. DNNs [KFY^{+}22, PKL22]. Do [STH17]. Docker [LPBF24]. DOCTOR [LJ24]. Does [RH23, RKK15]. Domain [BGK^{+}23, BJT^{+}23, CDX^{+}19, KOM^{+}23, OMMK23, Shu19b, SXXS^{+}16a, WZY^{+}23, YCJW24]. Domain-Specific [BGK^{+}23, KOM^{+}23, OMMK23, SXXS^{+}16a, BJT^{+}23]. dominance [WYJ^{+}14]. door [SCF12]. dose [ZHCY13]. Dot [EZL^{+}17, MKS23]. Dot-Product [MKS23]. Downtimeless [SVZ13]. DPM [CHK14b]. DPU [GKSB25]. DRAM [CLLC17, GHPP18, HPP17, HKP18, KRHC20, LO13, PP24, PMPP14, SJOL22, SRK^{+}18, YWLW23]. DRAM/PRAM [LO13]. Drift [KH23]. Drive [SYC^{+}17]. Drive-Thru [SYC^{+}17].

Driven [BGJ17, CWZ^{+}20, CEC23, FKJM18, GLP^{+}11, GTH^{+}22, MKMGS18, Rru22, WQGR22, WCM^{+}16, WZG^{+}23, CHCC13, DRL^{+}10, FRRJ07, FKS^{+}19, HG09, LP10, PEP05, RSB^{+}09, WLH16, BE10]. driver [KXL10]. drivers [BMM13]. Drives [CCSC23, CCC^{+}20, HWTL23, ISOD21, YWLW23]. Driving [HXH^{+}24, YKKD23]. drone [SPGB24]. Dropping [LCP^{+}17]. DSE [BZY^{+}23, SPGT19]. DSP [FO03, Geb04, KMB07, KGR12, LWB13, LPP^{+}21, ZXS03]. DSP-embedded [Geb04]. DSSoC [MHK^{+}23]. DSTL [CCC^{+}20]. DTLS [TNR17]. DTLS-Based [TNR17]. DTRL [BGK^{+}23]. Dual [DCZB19, MF12, SLS^{+}19, GLWM14, LLPM07, ZP06, VAHC^{+}06]. dual-bank [ZP06]. Dual-Channel [SLS^{+}19]. Dual-Mode [DCZB19]. dual-processor [GLWM14]. DUE [LJLT17]. Durability [YCJW24]. Durable [CLL16]. During [SPGT19]. DVFS [ACK^{+}13, CHCC13, CHK14b, PDL21, YC12]. DVS [QH07, ZM07, ZC08]. dwell [MLL08]. DWM [KY17]. DWM-Based [KY17]. DWMAcc [CDX^{+}19]. DWT [PZ12]. DyCo [YSC22]. Dylog [DLH16]. Dynamic [ALV^{+}22, CGM^{+}25, CPC17, CLJ^{+}19, CRCR13, DJ23, DLH16, ELS08, EMVR23, GVS^{+}20, GE18, GPB^{+}17, HS24, HLF^{+}18, HNY18, IHK04, KAK05, KBS17, KLK24, KG05, KFY^{+}22, LZL15, LLN^{+}14, LLLGR13, MLL^{+}17, MSD17, MKE18, NYH^{+}20, NPAG12, NZCS19, OZ22, PqBM^{+}15, PNRC17, QWY^{+}18, QZXO14, SL16, SKPL10, SBK^{+}23, TBEP16, TR24, UDB06, WMGR12, YGW^{+}12, YSC22, YKKD23, YKD^{+}24, ZRF^{+}12, ZC04b, ZW17, ASTPH10, ACK^{+}13, BCDH12, CRJ10, CLR05, FZHT13, FZJ08, HPLD09, ISG03, KR14, MMR^{+}10, NNS13, NSL11, NB04, OMA^{+}13, PZ12, SJRS^{+}13b, WYJ^{+}14, ZTD^{+}06, Zhu10, ZC08]. Dynamic-Priority [GE18]. Dynamical [GD19]. Dynamically [ARDG16, MZG14, GD14, HMMA04, KK05b, LB04, MSL13, VHB^{+}13]. Dynamics [ANB^{+}20]. DynO [ALV^{+}22]. DyPO [GPB^{+}17]. DyVEDeep [GVS^{+}20].

E-Cash-Toward [SBR^{+}15]. EACAN [PS19]. Eager [CSCC17]. Earliest [HQE20]. Early [BJT^{+}23, BZY^{+}23, GD24, KBD24, NVB^{+}20, PYL^{+}23, RG14, Shu18a, NKP^{+}12]. Early-stage [BJT^{+}23]. Earthquake [SKS21]. EAST [GQC^{+}17]. EASYR [YKKD23]. EAVE [LDV12]. EC [DZL^{+}22]. EC-ECC [DZL^{+}22]. ECAx [CGSH19]. ECC [DZL^{+}22, FLF^{+}23, HJ19, ZSH^{+}19]. ECG [CNC13, GZZ^{+}16, MVS^{+}13]. ECO [RSW21]. economic [WCH^{+}23]. Ecosystem [YMHB19]. Ed25519 [TV19]. EDA [FSG23, LAN06]. EDF [CHTC07, CAA^{+}24, ZB13, ZM07]. EDF-scheduled [ZB13]. Edge [BKS^{+}23, ASD^{+}24, BT22, BPSS24, BDK^{+}23, BONA22, CAN^{+}23, CDBB24, CZH^{+}24, CS22, CMMD24, DZL^{+}22, GK22, GOC^{+}22, GRWV22, GRR23, GRRR24, GDB22, HRH^{+}22, HY22, HLLL20, LLP^{+}21, LMS^{+}22, MBCM22, MMA^{+}23, MSSP22, MMD22, OHCK24, PAF22, PRL^{+}23, QRW^{+}24, RCS23, RB21, RLG20, STLX22a, STLX22b, WDM^{+}23, WZG^{+}23, YS23, YV23, ZDL22, KTT13, SSK^{+}22, BKS^{+}23, BPSS24, PMM^{+}17]. Edge-AI-Driven [WZG^{+}23]. Edge-Assisted [BKS^{+}23]. Edge-CaSCADe [BPSS24]. Edge-Cloud [GRRR24]. Edge-SLAM [BKS^{+}23]. Edge-TM [PMM^{+}17]. EdgeWise [GRWV22]. Editorial [BBM15, BE10, Bur05, CS16, CJL17, CGZ18, DPP14, DST19, EE16, EH18, FGIS12, FX17, HKP08, IT16, LB04, Leo18, LP09a, MCP17, Mit21, NKS12, DWCM14, PS14, Pla12, RRM16, Shu14a, Shu14b, Shu15a, Shu15b, Shu15c, Shu15d, Shu16a, Shu16b, Shu16c, Shu16d, Shu17a, Shu17b, Shu17c, Shu18a, Shu18b, Shu18c, Shu18e, Shu18d, Shu19a, Shu19b, Shu19c, Shu19d, Shu20a, Shu20b, VP16, WX17, ZQC16, Gup04, JM06, PBP09a, PBP09b, Sch07, SL04, ST05, Wha07]. Editors [HM17]. education [KCG^{+}05, SVP05, SBF^{+}05]. EEG [CNC13, MM16]. Effect [DVCC19, GTvdBC25, KHB^{+}23, GCU^{+}23]. Effective [BMF15, BLG^{+}15, CJL17, LMK^{+}18, LWK^{+}17, VGN18, GLT^{+}13, Mus10]. Effectiveness [SUS^{+}17]. Effects [DJO12, MGB^{+}21, RPHA19]. Efficiency [CRCR13, CEC23, HZH^{+}18, LSL^{+}23, LM24, OSA^{+}18, PC14, PMM^{+}17, SWL^{+}23, THA^{+}12, YJD^{+}17, YDS^{+}22, KVK^{+}03, LPFG13, SWL07, SJRS^{+}13b, SKPL10, SM13b, TVK08]. Efficiency-Driven [CEC23]. Efficient [APRC16, ABA^{+}20, AABG22, AJ18, ARZ^{+}23, ADJM19, BJ23, BDP24, BRR19, BGS^{+}18, BCS^{+}23, CAN^{+}23, CHK^{+}14a, CTK^{+}13, CSH^{+}22, CS22, CI17, CGV10, DCZB19, DMPC23, DLPK16, EKAK24, FGL^{+}19, FLF17, GQC^{+}17, GK22, GRR23, GSS^{+}18, GE18, HRT^{+}22, HY22, HLL^{+}23, HB23, JGL21, JFM23, JDW24, JJS^{+}25, JAB^{+}22, KW10, KCC^{+}16, KASD07, LS12, LX22, LL17, LAMS24, LX16, LWHS17, LMW^{+}17, LFC17, LBS15, LZZ^{+}23, LLK^{+}25, MYL^{+}22, MSR^{+}12, MMH^{+}23, MGLP19, MEK^{+}22, MKASJ18, NSL11, NWA12, NODP24, PVSG22, PP19, PCM^{+}15, PGS^{+}13, PHDL18, PS19, PLM^{+}15, PMP17, PLT23, PNRC17, RR17, Rat25, RMH04a, SNTÖ25, SSD^{+}19, SLB^{+}15, SK19, SA21, SPC^{+}16, SP19a, SJOL22, SPB^{+}17, SIC19, SWX17, SHQX19, TLL^{+}12, TBDdD11, TKT15, VKW^{+}17, WCK^{+}19, WLZ^{+}23, WZG^{+}23, YZZ^{+}23, YKKD23, ZLL^{+}18, AÖÖ23, ABF^{+}21, BCLN13, BT22, BMP23]. efficient [CAP^{+}07, CMGM24, yCBR05, DLC^{+}14, ESAS14, FZK^{+}10, GRWV22, HE12, HQB06, JGD^{+}09, KSK13, LAN06, LK10, LWZ^{+}24, MPT^{+}22, PO05, QH07, RGSS04, RP10, RKC^{+}22, RMD09, SKW^{+}07, SJRS^{+}13a, SPT^{+}23, SP20, SAYN09, UAK^{+}03, WRJL06, WKC07, WCH^{+}23, YS23, ZMB03, ZTRC03, ZP08, ZP07, ZC08, KMB07, CH10]. Efficient-Grad [HY22]. Effort [CRCR13, GVS^{+}20, MAG14, SRM^{+}13]. eIDs [ASWZ24]. elder [BDP^{+}13]. elder-care [BDP^{+}13]. Electric [VA18]. Electrode [EZL^{+}17, YCK^{+}18]. Elements [LVSVRFCG23, SBDK22, UFA^{+}24, HVG13]. ELFs [ZGH^{+}19]. Eliminate [DJ23]. Eliminating [RRW05]. Elimination [FND^{+}16]. Elliptic [DZL^{+}22, LWHS17, WPW^{+}04]. Elon [DLC^{+}14]. ELSA [AV20]. Embedded [ALZR19, Akd21, AB15, AKB^{+}25, ADJM19, BVM19, BD14, BHXP19, BDR24, BLG^{+}15, BP12, BJCHA17, CLL21, CS16, CKGN14, CBH22a, CBH22b, CGM^{+}25, CJL17, CCC^{+}17, CLS16, CFXY17, CQB^{+}15, DAHM16, DDC^{+}24, DLH16, DZL^{+}22, DBFH14, DQ14, DJS16, GLP^{+}11, GDDD17, GV21a, Goe14, HKC18, HJ19, HKP18, HXH^{+}24, HSK18, HNY18, HHL^{+}23a, IPEP12, JGL21, JRR16, JLSP18, JLW^{+}15, JEP16, JAD19, KE15, KML13, KRHC20, KPS23, KSP^{+}12, KM13, KCBM21, KBRD22, LS20, LG21, LDV12, LS12, LMA19, LJP17, LWZ^{+}24, LWHS17, LLZ^{+}17, LSL20, LL18, LLK^{+}25, MTWE20, MCP17, MCG22, Mit21, MGLP19, NBE18, OBSO16, PXY^{+}17, PCM^{+}15, PKIT23, Pau14, PqBM^{+}15, QZXO14, RRM16, RHG^{+}14, SG24, SLB^{+}15, SOG15, SSA^{+}24b, SDBD18, SCZ20a, SCZ20b, SCKD23, Shu14a, Shu14b, Shu15a, Shu15d, Shu16a, Shu16c, Shu16d, Shu18a, Shu18b, Shu18e, Shu19b]. Embedded [Shu19c, Shu19d, Shu20a, SRR^{+}23, SPGT19, SLE^{+}17, SVZ13, TDD^{+}16, TSP15, TBAS17, TBDdD11, TKL^{+}15, USA^{+}22, VFS^{+}21, VVKG23, VP16, VKW^{+}17, WDY^{+}16, WZM17, WXY^{+}17, WLC^{+}18, WX17, YGD^{+}17, ZDZ14, ZDTM19, ZSH^{+}19, ZQC16, ZAL22, ARJ08, ARJ11, ASTPH10, ABS02, AEF^{+}14, BYD09, BCDH12, BP05, BE10, BMP03, BMM13, BCS^{+}06, BMS13, BFQ10, BS13b, CMV10, CSVA^{+}05, CKL04, CC13a, CSK^{+}02, yCBR05, CRJ10, CGV10, CVG^{+}13, DKV14, Dea06, DKAL05, DZR09, DRL^{+}10, ESAS14, FRRJ07, Geb04, Geb06, GGGK08, GNP06, GRCV03, GT05, GGI13, GM03, Gup04, GKW08, HCK^{+}08, HG09, HFG13, HTLC10, HLD^{+}09, HXZ^{+}13, HZX^{+}14, HQB06, HQB07, HKLH05, JC03, JGD^{+}09, JKH^{+}13, JHPR13, KVN^{+}09, KST^{+}12, KBCL13, KCG^{+}05, KASD07, KD08, KGR12, LB04, LSK^{+}08, LS13, LOG^{+}14, LP09a, LCJ13, LOXL13, LHM14, MBFSV07, MSB08, MSL13, NRL13, DWCM14, NDB09]. embedded [NPP13, PCM12, PLKH08, PK13, PAS^{+}09, PO05, PEP05, RP03, RP11, RV03, RRKH04, RP10, RSB^{+}09, SVP05, SWL07, Sch07, SAHE04, SMG04, SL04, Sev05, SJC^{+}03, ST05, STW13, SVN04, SGDP12, SBF^{+}05, TRJ05, TSWL10, TSG10, TVK08, TLLL09, UAK^{+}03, VAHC^{+}06, VS05, VHB^{+}13, WTSR13, WMT12, WPW^{+}04, WRJL06, Wu10, WMZY13, XQ07, YDLC10b, ZCS^{+}05, ZC04b, ZVL04, ZVN05, ZB13, ZMB03, ZP08, ZP09, ZM07, Zhu10, ZP06, ZP07, DEG11, HKP08, Shu18d, KAS^{+}20]. embedded-system [BE10]. embedded/multimedia [UAK^{+}03]. Embedding [HHB^{+}12, SWWY13]. emergency [KLC^{+}10, WYS^{+}13]. Emerging [ZQC16, SRY13]. EMG [WGPH13]. Emissions [ISOD21]. eMRTDs [ASWZ24]. EMS [WLZ^{+}23]. EMS-i [WLZ^{+}23]. Emulation [AAM^{+}17, MRA^{+}17]. Emulator [MZG14, WT15]. Enable [BGGT23, LLC^{+}22]. Enabled [DJJ^{+}19, SAS^{+}23, VKDG19, RC08, SLZ^{+}24, UAK^{+}03, CZH^{+}24]. Enabling [BCDH12, CCSC23, CCC^{+}20, DLC^{+}14, JRR16, LYC^{+}18, LCC^{+}19, MGP25, PM19, QWY^{+}18, SJRS^{+}13b, SDMK19, WDM^{+}23, YCJW24]. Encapsulation [AAT^{+}21, KNK^{+}25]. Enciphering [MKASJ18]. Encoded [BDP24]. EncoDeep [SJK20]. Encoder [FS13]. Encoder/Decoder [FS13]. Encoding [SJK20, SAYN09, THON12, LDV12]. Encryption [GCS24, MSR^{+}17, SXH^{+}19, VOG15, CMGM24, SKW^{+}07]. End [DVCC19, GRRR24, GTvdBC25, GCS24, QRW^{+}24, SSK^{+}22, CZ23]. End-edge-cloud [SSK^{+}22]. End-To-End [GTvdBC25, DVCC19, GRRR24, GCS24]. Endomicroscopy [CLS16]. Endurance [GMCC18]. Energy [ABL^{+}20, AHM19, ABD^{+}19, ANB^{+}20, ABA^{+}20, ABTS24, AJ18, ARZ^{+}23, AKTM16, ABC^{+}17, ASJ21, AV20, BCLN13, BFW^{+}19, BMAB16, BSA17, BMP03, BTA^{+}19, BGS^{+}18, CHK^{+}14a, CHK14b, CLL^{+}18, DKV14, DAHM16, DJO12, DLPK16, ESAS14, EYG^{+}23, FRRJ07, FS13, FPGS22, FBM16, GRWV22, GRR23, GTH^{+}22, GDC19, HPBL12, HDG^{+}14, HSR18, HB16, HNY18, HZGW18, HQB06, HZH^{+}18, HLL^{+}23, HB23, JBDD20, JRSR17, KE15, KY17, KBDV08, KYDC20, KNL12, LDV12, LS12, LOD18, LYC^{+}18, LWZ^{+}24, LAMS24, LM24, MLL^{+}17, MAS^{+}24, MPT^{+}22, MSRM24, MSR^{+}12, MMH^{+}23, MGLP19, MKE18, NRZ^{+}24, PC14, PCM^{+}15, PMM^{+}17, PJL^{+}17, PHDL18, PLM^{+}15, QH07, QFL^{+}24, RGSS04, RR17, RV03, RV07, RKC^{+}22, RSW21, SA21, SPC^{+}16, SPT^{+}23, SP19a, SP20, SJOL22, SKN17, SPB^{+}17, SVN04, SAYN09, TRJ05, TLL^{+}12, WDJ^{+}18, WRJL06, XLY18, YJD^{+}17, YCT16]. Energy [YS23, YKKD23, YW13, YDS^{+}22, ZZG24, ZQ24, ZO16, ZC08, BZ13, BBL09, CAP^{+}07, CSK^{+}02, CLK13, FZJ08, GNW05, Geb06, GGI13, GHZH14, HE12, HLD^{+}09, KKC^{+}05, KHZS07, KVK^{+}03, KAK05, KDN^{+}07, LSK^{+}08, LK10, LWB13, LPFG13, LOXL13, MRY^{+}10, MKD13, PAP^{+}12, RP10, RMM03, SRS03, SKPL10, SJC^{+}03, SM13b, SC05, TTAG14, TVK08, UAK^{+}03, YK03, ZVN05, ZKKC05, ZTD^{+}06, ZA07, ZX08, ZP08, Zhu10, CH10]. Energy- [FS13]. Energy-Accuracy [JBDD20]. Energy-Aware [BMAB16, DAHM16, HB16, JRSR17, ZZG24, BMP03, DKV14, EYG^{+}23, KBDV08, YW13, FZJ08, GGI13, SRS03]. Energy-Constrained [AV20]. Energy-Efficient [ABA^{+}20, AJ18, ARZ^{+}23, BGS^{+}18, CHK^{+}14a, DLPK16, GRR23, HLL^{+}23, HB23, LS12, MSR^{+}12, MMH^{+}23, PCM^{+}15, PHDL18, PLM^{+}15, RR17, SA21, SPC^{+}16, SP19a, SJOL22, SPB^{+}17, TLL^{+}12, YKKD23, BCLN13, ESAS14, GRWV22, HQB06, LWZ^{+}24, MPT^{+}22, RKC^{+}22, SPT^{+}23, SP20, SAYN09, WRJL06, YS23, ZC08, CAP^{+}07, HE12, LK10, RP10, ZP08, CH10]. Energy-Fidelity [HPBL12]. Energy-Free [LYC^{+}18]. Energy-Harvesting [ABC^{+}17, HSR18, HZGW18, GHZH14]. Energy-Neutral [WDJ^{+}18]. energy-optimal [SC05, YK03]. Energy-optimizing [FRRJ07]. Energy-Proportional [FPGS22]. energy-synchronized [GHZH14]. Energy/QoS [LDV12]. ENFFiS [PK13]. Enforcement [PRS^{+}17]. Enforcer [CD17]. Enforcing [WWY13]. engine [TSWL10]. Engineering [BCHL19, EKAK24, HMR23, PL13, YGW^{+}12, DRL^{+}10, Sev05, SBF^{+}05]. Enhanced [PDL21, PK13]. Enhancement [CCC^{+}17, DDC^{+}24, PRK15, CYKH13, HL14, KD08]. Enhancing [CLK13, IAS23, LM24, OFA^{+}15, KK05a]. Ensemble [HZW^{+}23]. Ensure [NZCS19]. Ensuring [LMA19, WLC^{+}18]. Entanglement [CD19]. Entire [JSD23]. entrant [VWG^{+}17]. Environment [AARJ12, BRL16, MKOA25, RHG^{+}12, SSB24, TSY^{+}16, CSK^{+}02, DDG^{+}13, HT06, SB08, WCJ07, WTSR13]. environmentally [LAHS06]. Environments [BCDD24, BE17, ZLZ^{+}24, WP11]. EnviroSuite [LAHS06]. Envision [LLK^{+}25]. epileptic [MVS^{+}13]. Equation [BGGT23, HVG13]. Equation-Directed [BGGT23]. Equations [LZZ^{+}23, WWTSM19]. Equilibrium [YMHB19]. Erased [PRM21]. Error [BDB^{+}17, CGSH19, DVC21, EZL^{+}17, FND^{+}16, KKD^{+}12, LJLT17, MLR^{+}17, MKASJ18, OSA^{+}18, PMM^{+}17, SK13, SUS^{+}17, TP16, WZD^{+}17, YZZG23, YC12, HLD^{+}09, MMK22, LDV12]. Error-Aware [KKD^{+}12, LDV12]. Error-Recovery [EZL^{+}17]. Errors [KKL^{+}16, KJK^{+}17b, LLP^{+}17, RJS19, WCM^{+}16, YZA13]. ES-IMS [KBCL13]. Escape [LS20, SLE^{+}17]. Escapers [SXH^{+}19]. ESL [GD24, SOL^{+}16, WMLA16]. Essential [ZW13]. Essentiality [ZTZ^{+}19]. Establishment [DL12]. ESTEREL [TdS05, SMR^{+}18, YKK^{+}13]. Estimating [OBSO16]. Estimation [ADH^{+}23, AVR22, CLL21, CYH20, FHK21, HRH^{+}22, LKA^{+}18, OBA^{+}17, PJWY12, RHG^{+}12, SOL^{+}16, SRR^{+}23, TKT15, WZM17, WSMF22, KKC^{+}05, KS13, LSC14, MSL13, ÖNG08]. Estimations [RSF20]. ESTIMedia’08 [BCEP12]. ESTIMedia’09 [PCB12]. ESTIMedia’10 [Edi13]. ESTIMedia’11 [CC14]. ESTIMedia’12 [CP13b]. ESTIMedia’13 [PS14]. ESWEEK [EE16, EH18]. ETAP [EYG^{+}23]. ETC [YKD^{+}24]. Evaluating [BCS^{+}23, HABT11, MMS06, SCS16, Shu20b]. Evaluation [BHET04, FHB^{+}17, HHC^{+}16b, JLSP18, LZS^{+}18, MCSW12, RG14, SMG04, SSS11, TGP25, WT15, YDS^{+}22, CLR05, HGL14, KJRG13, LSC14]. Evenly [LC17]. Event [HHC^{+}16b, JZL^{+}15, KL13, MV16, MBLA16, NDZ13, KW10, DAASP21]. Event-B [DAASP21]. Event-Based [JZL^{+}15]. Events [HSR18, ZH12a]. everyone [Shu14a]. everywhere [Shu14a].

Evolution [SVZ13]. evolutionary [HMM04]. Evolve [RRM16]. Exact [XZK^{+}19, YLW15]. Example [WZS^{+}24]. Examples [PMAB19]. exascale [DBH14]. Exchange [AAT^{+}21]. executables [DVC^{+}07]. Execute [WLK^{+}19]. Executing [DKA^{+}19]. Execution [AKD^{+}18, AARJ12, ABS^{+}19, BCD^{+}22, BCS^{+}23, BMMV21, CBRZ19, EVS^{+}17, FSB^{+}21, GKSB25, KHB^{+}23, LMA19, LWCL25, MCM^{+}17, MZG15, MBLA16, REPL15, RKK15, TB23, WWG^{+}18, WZ12, HG09, MEP04, WEE^{+}08, YZ08]. Execution-Time [EVS^{+}17, WEE^{+}08]. executions [LES14]. Exit [KBD24]. Exotasks [ABI^{+}09]. Expansion [CKB17, BYD09]. Experiences [RIMS21, WYS^{+}13, CLK13, CMP^{+}07]. Experiment [TSY^{+}16]. Experimental [BHET04, LKZ^{+}23]. Explanation [SRB23]. Explicit [SSD^{+}19, WAD14]. Exploit [ZQ24]. exploitation [KVN^{+}09]. Exploiting [CKN^{+}20, CFGM15, FS14, HE12, HC16, LPD^{+}20, LPE^{+}23, MWK^{+}24, NBE18, PXY^{+}17, PMM^{+}17, PE23, SK13, SGZS21, SDMK19, SWWW17, XDL^{+}18, ZM07, CLK13, GFC^{+}10, ZA07]. Exploration [ABL^{+}20, BCS16, BJT^{+}23, CDH^{+}16, DJJ^{+}19, FFA^{+}23, FSC^{+}16, FSVG19, GHD^{+}25, GCJD20, GSN21, IVJ^{+}23, JFM23, KAKSP15, KB23, MPT^{+}22, OFA^{+}15, PSZ12a, PWL^{+}19, SLB^{+}15, SXXS^{+}16b, WSHC14, YLTY21, ZEJ^{+}23, ZBG20, BFQ10, CIC^{+}08, CIC^{+}09, GDN03, JBN^{+}13, KGR12, LM13, MPZS13, OP06, PDBR08, SKW^{+}07, YCLV^{+}02]. Exploratory [KNK^{+}25]. Explore [BMP23, CAP15]. Exploring [DJO12, IFA^{+}16, WKJ20, WSK14]. Exposing [HKP18, SWL07]. EXPRESS [GKSB25]. Expression [WZG^{+}23]. Extended [AJ18, LDV12, WSHC14]. Extending [GMCC18, OFA^{+}15, YGD^{+}19]. Extensible [MHK^{+}23]. Extension [GI24, LWS^{+}23, PRSV19, MBFT09, RMH04a]. Extensions [KRR20, LTQ^{+}24, PJT^{+}23]. External [JGL21]. Extractor [XHK16]. Extreme [RKC^{+}22, YV23]. Extremely [CJL17].

F [MSHS19]. Fabrics [HMR23]. Face [APF24]. Facial [WZG^{+}23]. Facilitating [AMJ21]. Factored [JFK15]. Factors [Shu19c]. Fails [SZL^{+}17]. Failure [BV15, GPH24, SLS^{+}19, TMXS17]. Failure-based [GPH24]. failures [CRAJ10]. Fair [RPB^{+}19, RGSS04]. Fairness [CLLC17, GHKS15, RPB^{+}19, CJMB05]. Fall [LMS^{+}22]. Falsification [AFS^{+}13]. Family [MFG16]. FARSI [BJT^{+}23]. Fast [AP20, ABA^{+}20, AGG^{+}17, AD24b, CSCC17, CHS15, EKAK24, NS16, PDBR08, YMHB19, YCNCC11, BWS14, LM13, LHCK04, TLLL09, VJD^{+}07, VDK^{+}08, SAMR06]. Faster [HLM^{+}24, LHP^{+}23]. Fault [AMKA17, BVM19, BHD15, CPC17, DSB17, FXP^{+}17, GAS^{+}17, HA24, IPEP12, LCD18, LCLW17, LPE^{+}23, MKMGS18, MCP17, MKAA17, MAGR15, NDZ13, NRZ^{+}24, RCDB24, Rru22, SA18, SISS24, SSH14, TAP23, TMXS17, XKK17, YGD^{+}17, AFG08, BGD14, CMV10, JGD^{+}09, RMH04b, SHME13, ZC04b]. Fault-injection [RCDB24]. fault-tolerance [AFG08]. Fault-Tolerant [BHD15, CPC17, DSB17, IPEP12, MCP17, SA18, SSH14, TMXS17, NRZ^{+}24, TAP23, BGD14, JGD^{+}09, RMH04b]. Faults [EVS^{+}17, VM23, VS08]. Faulty [BVM19]. FD [WZY^{+}23]. FD-CNN [WZY^{+}23]. FDL [GV21b, DGG24]. FE [XHK16]. FE-SViT [XHK16]. Feasibility [SGW^{+}16, YRF10]. feasible [LA11, RM10]. Feature [TR24]. Federated [GP23, NFL^{+}22, TSY^{+}16, TSO22, YS23]. FedHIL [GP23]. Feedback [IAS23, KT14, ZM07]. Feedforward [YF19]. FELIX [SLK^{+}22]. Fence [Shu16b].

Fencing [FND^{+}16]. Ferroelectric [SLK^{+}22]. FET [SLK^{+}22]. Fetal [FSVG19]. FFConv [AP20]. Fiat [VS08]. Fidelity [HPBL12]. FIDES [ISTE08]. Field [NWA12, Shu16b, ITO^{+}24]. fields [RMH04a, RMH04b]. FIFO [GNW05, TBG^{+}17]. File [CCC^{+}17, KSP^{+}12, LCC^{+}23, OBSO16, CWKH12, LS13, PK13]. File-system-aware [LCC^{+}23]. file-system-oriented [CWKH12]. Filed [HCS18]. filling [BSKB^{+}09]. Filter [HZW^{+}23, CMS08]. Filtering [UM13, YYKK18, MSH^{+}14, TSG10]. filters [CC13b, FF09]. final [GGGK08]. Finding [VSD^{+}17]. Fine [BHL^{+}20, CAN^{+}23, KJLS20, MSRM24, DFC^{+}19]. Fine-Grained [KJLS20, DFC^{+}19, BHL^{+}20, CAN^{+}23]. Fine-Tuning [MSRM24]. FinFET [CSM24]. Fingerprinting [BS22, HMLZ21, ISOD21, PRM21]. Fingerprints [TM15]. Finite [CHS15, DQ14, NWA12, ZPZG17, RMH04b]. Finite-State-Machine [CHS15]. FINN [AGG^{+}23]. Fire [MMD22, PKIT23]. Firmness [BAG^{+}20]. Firmware [MKMGS18, McI13]. First [HQE20]. Fix [DLV16]. Fixed [BMS24, DBM^{+}15, DHL17, LH18, LJVD23, SCM20, SD17, WHN^{+}17, ZLL^{+}19, AC08, DF14, LA11, QH07, YK03]. Fixed-Point [LJVD23, BMS24, SCM20, AC08]. Fixed-Priority [DHL17, LH18, LA11, QH07, YK03]. fixed-priority-scheduled [DF14]. Flash [BDG^{+}15, CSCC17, GMCC18, GKS^{+}22, ISOD21, JGL21, JCS^{+}17, JNI15, KKK^{+}11, KSP^{+}12, KNY^{+}17, Kwo16, MAW22, OBSO16, PRM21, SWJ^{+}13, WDM17, WC16, WZD^{+}17, CH10, CKL04, CWKH12, CYKH13, LPC^{+}07, PCK^{+}08, PK13, WKC07, Wu10, JKJ^{+}10, MSHS19]. Flash-Based [CSCC17, PCK^{+}08]. flash-memory [CKL04, CWKH12, CYKH13, WKC07]. FlashKV [ZLSQ17]. FlashLight [KSP^{+}12]. Flaws [SZL^{+}17]. Flexibility [IAS23, WSK14]. Flexible [ABTS24, BHD15, CC13b, DMPC23, HWC22, KLK24, NVB^{+}20, PP19, PJL^{+}17, TV19, VWG^{+}17, MTL14, SJK20, ZW10]. FlexRay [SKH^{+}12, TBEP16]. FlexWAFE [DSW^{+}09]. Float [WHL23, WHL23]. Floating [LPP^{+}21, MLR^{+}17, Rat25, AC08, DBH14, LYL13]. floating- [AC08]. Floating-point [LPP^{+}21, MLR^{+}17]. FLoorplan [SBB19]. FLORA [SBB19]. Florets [SPT^{+}23]. Flow [CKB17, DHL17, GD24, LPBF24, PMP17, RJS19, SUS^{+}17, SPC^{+}16, SPT^{+}23, SIC19, VAHC^{+}06, WLW^{+}24, YZZG23, BHL^{+}20, CC13a, CMPP23, LMST04, MCG22, PGR^{+}08, RI04, TBG^{+}13, WYJ^{+}14]. Flow-aware [SPT^{+}23]. Flow-Based [CKB17, SIC19]. Flow-Layer [CKB17]. FlowPaP [YJD^{+}17]. FlowReR [YJD^{+}17]. Fluid [ARS16]. Fly [PM19, UM13]. FMM [HZH^{+}18]. FMSs [HPS13]. Focused [HPBL12, MSRM24]. Focusing [JDW24]. Fog [AAR^{+}17, Li21, SKS21]. Fog-Assisted [AAR^{+}17]. Footprint [LMB^{+}22, CDD^{+}07, HFG13, PLKH08]. Forecasting [ZK25]. Forensics [CFXY17]. ForEVeR [PB14]. Fork [SGW^{+}16]. Fork-Join [SGW^{+}16]. Formal [BGVZ11, CD17, CD10, DST19, DHF18, GDA13, GGMS24, Leo18, LML20, MS13b, Mos13, SVZ13, TBFR17, ZQD^{+}23, CCB^{+}06, HHB^{+}12, KST^{+}12, PB14, RBS^{+}10, BVM19]. formalism [Gar05]. Formalization [MHT13]. Formalized [LFC17]. Formally [CMP23, MS23]. Format [BDP24, CPP^{+}17]. Format-Independent [CPP^{+}17]. Formats [MMH^{+}23]. ForSyDe [UGS^{+}21]. ForSyDe-Atom [UGS^{+}21]. Forwarding [HSR18]. Foundations [BCHL19]. FP2I [Rat25]. FPGA [AMKA17, AP20, BSKB^{+}09, BFST19, BRA^{+}16, CCA^{+}13, CCC^{+}14, CHS15, CDH^{+}16, EKAK24, GZZ^{+}16, HJ19, HNY18, HPLD09, HW17, JSZ^{+}19, LSC14, LS17, LMS^{+}19, MDWL23, MMSN14, MCSW12, MCM^{+}17, RPB^{+}19, RMK17, SSK21, SNTÖ25, SUK23, DFC^{+}19, SB08, SM13a, TAP23, TV19, USA^{+}22, WL09, WZY^{+}23, YDS^{+}22, ZBCM09, ZHCY13]. FPGA- [MCM^{+}17]. FPGA-Based [GZZ^{+}16, HW17, AP20, CCA^{+}13, HPLD09, SSK21, SUK23, SB08, TAP23, USA^{+}22]. FPGAs [LZNL23, AZS^{+}23, DSW^{+}09, HVG13, JSD23, KJRG13, LZJ^{+}20, OFA^{+}15, PGS^{+}13, PAS^{+}09, PLT23, QFL^{+}24, RBNM19, SBB19, SWWW17, WGP04, YZZ^{+}23]. FRAM [JRSR17]. Frame [ESM^{+}17, PEP05, SGT^{+}13]. frame-based [SGT^{+}13]. frames [NPP13]. Framework [BCD^{+}22, BTD^{+}18, BRR19, BMF15, BJT^{+}23, BP19, BHL^{+}20, CWZ23, CJ20, CWX^{+}23, DJJ^{+}19, DSXS^{+}14, DAASP21, GLP^{+}11, GKSB25, HSD22, HFA^{+}14, IVJ^{+}23, JKH22, JLSP18, KCWH14, KPC^{+}16, KSA^{+}18, LJ24, LKH16, LCY^{+}22, LZZ^{+}23, MWF^{+}16, OBO^{+}23, OMH^{+}23, PRK15, PKL22, RMK17, SRB23, SL16, SFCW23, DFC^{+}19, SSH14, SYC^{+}17, SXXS^{+}16a, SXXS^{+}16b, SVS21, SC20, SCA^{+}24, SRA12, TLSJ23, VKDG19, WZG^{+}23, XHK16, XKK17, YDS^{+}22, ZDTM19, ARJ11, BWS14, CCR^{+}14, DZR09, FZJ08, KKO^{+}06, KGR12, LSK^{+}08, LAHS06, PO05, RDM06, SJRS^{+}13a, SL08, STY^{+}14, SGDP12, UAK^{+}03, ZW10]. Frameworks [TP20]. Franca [LMBL21, LML^{+}23]. Free [CGZ18, CLJ^{+}19, CQB^{+}15, HMLZ21, LOD18, LYC^{+}18, LHYQ18, PGR16, WXY^{+}18, WLHC18, CRJ10, DGC^{+}20, HHB^{+}05, LES14, OZ22, PRL^{+}23, RP10, ZCK13, MMY^{+}19, MSHS19, MFMA17]. Frequency [LOD18, PC14, PHDL18, SOL^{+}16, WZY^{+}23, YGW^{+}12, SAMR06].

Frequency-Aware [SOL^{+}16, SAMR06]. Frequency-Domain [WZY^{+}23]. Frequent [YG02]. Friendly [ORA16, GDB22, ZZA^{+}22]. Front [CAP15]. FSIMR [LCC^{+}23]. fsync [CSCC17]. FTL [CWKH12, JKJ^{+}10, PCK^{+}08]. Fuel [YTL^{+}20, SPK^{+}12]. Fuel-Cell [YTL^{+}20]. Fully [GCS24, XDL^{+}18, ZQ24, CMGM24, LPC^{+}07]. fully-associative [LPC^{+}07]. Function [KBS17, LZJ^{+}19, MGC^{+}23, PLT23, WWTSM19, SWWY13]. Function-Level [KBS17]. Functional [CPP^{+}17, Fra12, KM09, AKB14, GD14]. functionality [PB14]. Functions [HLM^{+}24, KLK24, MFMA17, ZWH^{+}16]. Fusing [HXH^{+}24, MS13a]. Fusion [APRC16, CWZ23, CZK^{+}22, DDC^{+}24, HCS18, IPL16, LLW^{+}17, TXL^{+}12]. fusion-based [TXL^{+}12]. Future [AYS15, LLK^{+}25, MBCM22, HKP08, SM13b]. Fuzz [LZJ^{+}19]. Fuzzy [MMY^{+}19, XHK16, LLLGR13]. fuzzy-logic-based [LLLGR13].

gains [BZ13]. gait [VAR13]. Game [CZH^{+}24, HLLL20, SR12b, WDY^{+}16, XWL25]. Game-Based [HLLL20]. Game-Theoretic [SR12b]. Game-Theory-Based [WDY^{+}16]. Games [CMP23, PHDL18]. GAN [ZCG^{+}22]. Gana [ZCK13]. Gaps [Akd21]. Garbage [CLL16, GMN21, KSY17, CKL04, CW14, CSK^{+}02, DKAL05, SP10]. garbage-collection [CW14]. Gas [EHF24]. Gateway [GAD^{+}24, XZK^{+}19, SKH^{+}12]. Gateway-Integrated [XZK^{+}19]. Gating [WHL23]. Gaussian [TJ23]. GDB [MZG14]. general [CMGM24, GKW08, HVG13, MTL14]. general-purpose [CMGM24, GKW08]. Generalized [PSD21]. Generating [BP12]. Generation [ASWZ24, BMS24, BZY^{+}23, BMMV21, CMK12, EK12, FKS^{+}19, HWC^{+}20, HYY^{+}15, LVSVRFCG23, LFC17, MCSW12, MZG14, POCR24, SMZ^{+}21, SSB24, SRK^{+}18, CMP23, ISTE08, ISE10, IBMK10, KOL^{+}22, KM09, LLC^{+}13, NNS13, SRY13, TBG^{+}13]. Generators [MFG16]. Generic [BGRV15, SFCW23]. GENESIS [DSXS15]. Genetic [Ahm13, SMZ^{+}21]. Genomic [MGLP19]. Get [SPGT19]. Getaway [SLE^{+}17]. GHOST [ASS^{+}23]. GIS [MBB^{+}15]. Givens [SPC^{+}16]. Global [DBM^{+}15, DHL17, PYJL15, ZLL^{+}19, BMM13]. Globally [YMHB19]. GMAI [CKN^{+}20]. GNN [QFL^{+}24]. Goal [SSB24, SGDP12]. goal-oriented [SGDP12]. Golden [HMLZ21]. Good [AR14, MASG15, Shu16a]. Goppa [MBR15]. Governing [HTC^{+}16]. Governor [PDL21]. GPGPUs [Ato20, ASJ21, WZJ^{+}18]. GPU [PDL21, SPB^{+}17, CKN^{+}20, CCC^{+}14, DZL^{+}22, LJ14, HHL^{+}23b, KWKP23, LSC19, LWB18, OBA^{+}17, OFA^{+}15, PHDL18, RC17, WZM17, YW13, YC16, ZXCH13]. GPU-like [LJ14]. GPU-optimized [ZXCH13]. GPUrpc [IFA^{+}16]. GPUs [BONA22, HA24, IFA^{+}16, LL17]. Graceful [ZZG24, RGdZS14]. Grad [HY22]. Gradient [HY22, XCZW23, LM24]. Grading [KBRD22]. graduate [CSVA^{+}05]. Grained [CGM^{+}25, JSD23, KJLS20, KMS^{+}23, LCD18, DFC^{+}19, WLTW24, BHL^{+}20, BZY^{+}23, CAN^{+}23, VNK^{+}03]. Grant [PMS^{+}24]. Granularity [HWTL23, MFG17]. Granularity-Based [HWTL23]. Graph [ASS^{+}23, AÖÖ23, DLPK16, HPS13, LHL^{+}19, OBO^{+}23, QFL^{+}24, SGW^{+}16, YS23]. Graph-Based [LHL^{+}19]. Graphical [GSSR^{+}24, LCQ^{+}13]. Graphs [ADJM19, DHKS15, DP19, KAKSP15, LWB18, MS21, MG15, MBLA16, RDSS21, SB23, FZHT13, LXK10, TBG^{+}13, WBS10]. Grazing [Shu16b]. GRec [DP19]. Greedy [CLL16, WMT12]. Green [XLY18]. Green-Energy-Powered [XLY18]. Grid [TSY^{+}16, SGDP12]. Group [ALZR19, TNR17]. Groups [BMF15]. GroupSense [ALZR19]. GSFAP [TSG10]. Guarantee [NLSV^{+}19, STH17]. Guaranteed [ABD^{+}19, LLT^{+}17, PJT^{+}23, TBCB15]. Guarantees [HQE20, IB23, KT14]. Guest [BBM15, CS16, CJL17, CGZ18, DST19, EE16, EH18, FX17, Gup04, IT16, JM06, Leo18, MCP17, PBP09a, PBP09b, RRM16, Sch07, SL04, ST05, Wha07, WX17, ZQC16, HM17]. Guided [FS24, GDD20, LP19, ZQD^{+}23, FKS^{+}19, OMH^{+}23]. Guidelines [CSVA^{+}05]. GUSTO [IBMK10].

H.264 [SHME13]. Hack [DLV16]. Hadamard [PBC22]. Half [SWJ^{+}13]. Half-Wits [SWJ^{+}13]. Handheld [YJD^{+}17, CHCC13]. handler [LP10]. Handling [Rru22, KW10]. Hard [CQB^{+}15, HFA^{+}14, LOF20, OSF19, PSD21, SP19a, SLCS16, SD17, UBF^{+}16, CRM14, HQE20, PMM^{+}13, SRM^{+}13, SC05, YK03, ZZZ^{+}12]. Hardness [SGW^{+}16]. Hardware [ADH^{+}23, Alg24, ARH^{+}18, BVM19, BPSS24, BMP23, BJCHA17, BRL16, CAN^{+}23, CAP15, CZHK23, CEC23, FSB^{+}21, GMN21, GI24, GIA11, GV21a, GDB22, GLS^{+}23, HT06, HWC^{+}20, HZH^{+}18, HE24, JR20, JGCS24, JAD19, KJLS20, KSFS24, KE15, KNK^{+}25, LX12, LMB^{+}22, LPP^{+}21, LLG^{+}20, LKW^{+}24, MWS15, MCS^{+}15, MKM^{+}23b, ORA16, PRK15, PCM^{+}15, PKT23, PMDC17, PMP17, PM19, RRC22, RKM24, Rat25, RH23, RPHA19, SSK21, SNTÖ25, SKKR11, SZG^{+}23, SMZ^{+}21, SCB^{+}22, TSY^{+}16, THG24, TKT15, TLSJ23, UM13, VGG^{+}13, WCJ07, WRB15, ZEJ^{+}23, ZZA^{+}22, ZAL22, ARJ08, CCB^{+}06, JM06, KTT13, LOG^{+}14, NSL11, OP06, PZ12, PBP09b, RP11, RI04, SMG04, SB08, SVN04, TTAG14, VS08, DEG11]. Hardware-accelerated [RRC22, ZAL22]. hardware-assisted [LOG^{+}14].

Hardware-Aware [LKW^{+}24]. Hardware-Based [BPSS24, UM13, NSL11]. Hardware-centric [SZG^{+}23]. Hardware-Efficient [TKT15, BMP23]. Hardware-Friendly [ORA16, GDB22, ZZA^{+}22]. Hardware-Software [LLG^{+}20, JR20]. Hardware/software [WCJ07, ARJ08, SB08, DEG11]. Hardware/Software-Embedded [DEG11]. harmful [YKK^{+}13]. Harmonic [HSMS16]. Harmonicity [WHN^{+}17]. Harmonicity-Aware [WHN^{+}17]. Harnessing [LKB14]. HARP [LKB14]. HARS [LOG^{+}14]. Harvest [CLL^{+}18]. Harvested [ZQ24]. Harvesting [ABD^{+}19, ABC^{+}17, BFW^{+}19, HSR18, HZGW18, KY17, LOD18, MLL^{+}17, PJL^{+}17, SKN17, GHZH14, KHZS07]. Hash [MKAA17, MKASJ18, THG24]. Hash-Based [MKAA17, THG24]. Hash-Counter-Hash [MKASJ18]. HDL [GI24]. HDLRuby [GI24]. Head [HS24]. Health [BTA^{+}19, HPBL12, HZYJ22, LMW^{+}17, JLSK13, KS10]. Healthcare [AAR^{+}17, CD10]. Heap [OMH^{+}23, BS13a, CH08, BVGVEA10]. heart [BJM13]. Hennessy [VRF15]. Hephaestus [SVC^{+}23]. Hercules [YCJW24]. HESSLE [MMY^{+}19]. Heterogeneity [GP23, AMN^{+}14]. Heterogeneous [AR14, BCS^{+}23, CGM^{+}25, COC22, CEC23, CDH^{+}16, CAA^{+}24, ETAV16, FMHS23, GQC^{+}17, GPB^{+}17, HGW^{+}20, KS18, KGT^{+}23, KSA^{+}18, LLW^{+}17, LLZ^{+}17, MKOA25, MG15, MMY^{+}19, OBO^{+}23, PRB15, PVV24, PqBM^{+}15, PLM^{+}15, QP15, RC17, RN14, RLP^{+}21, RDSS21, SXXS^{+}16b, THA^{+}12, VFS^{+}21, VKW^{+}17, VSD^{+}17, YHL23, ZDTM19, AP09, BCC^{+}08, FC13, KBDV08, NBGS09, PGR^{+}08, VHB^{+}13, WSK14]. Heuristic [BJ23, FKS^{+}19, KAKSP15, Li21, SEB12, VSSS13, YCNCC11].

Heuristic-guided [FKS^{+}19]. Heuristics [MG15, WLW^{+}24, OMA^{+}13]. HiCH [AAR^{+}17]. Hidden [GGJ12]. hiding [XHSS10]. Hierarchical [AAR^{+}17, CZH^{+}24, DAHM16, GNR^{+}10, ITO^{+}24, MCSW12, SCA^{+}24, TAMS18, ZEJ^{+}23, AFL13, TBG^{+}13]. Hierarchies [FF25, MDS^{+}21]. Hierarchy [GKS^{+}22, OMH^{+}23, TBG^{+}17]. Hierarchy-aware [OMH^{+}23]. High [AÖÖ23, BRL16, CCP^{+}19, DLPK16, FMHS23, FLF17, HTB^{+}25, HHL^{+}23b, HHL^{+}23a, HW17, HZH^{+}18, IYL^{+}23, ITO^{+}24, KCWH14, KPC^{+}16, LG21, LWB13, LN19, LCH^{+}08, LPO^{+}17, MSR^{+}17, NASM18, PCM^{+}15, PKT23, PMDC17, PGR16, QFL^{+}24, RPHA19, SRG^{+}15, SZG^{+}23, SPT^{+}23, SP12, WLK^{+}19, WKP24, YDLC10a, YCK^{+}18, ZDTM19, ZAL22, BCLN13, BAR13b, CCA^{+}13, FO03, KKC^{+}05, LLC^{+}13, PGS^{+}13, PSZ12b, THON12]. High-accuracy [ITO^{+}24, PSZ12b]. High-assurance [RPHA19]. High-Density [YCK^{+}18]. High-Level [BRL16, FLF17, KPC^{+}16, LN19, PKT23, PMDC17, IYL^{+}23, BAR13b, CCA^{+}13, FO03, KKC^{+}05]. High-Performance [DLPK16, HHL^{+}23b, KCWH14, LPO^{+}17, NASM18, PCM^{+}15, PGR16, SRG^{+}15, SPT^{+}23, SP12, HHL^{+}23a, LWB13, LCH^{+}08, SZG^{+}23, YDLC10a, ZAL22, BCLN13, PGS^{+}13]. High-resolution [LG21]. High-Speed [HW17, MSR^{+}17, LLC^{+}13]. High-Throughput [HTB^{+}25, AÖÖ23, THON12]. High-voltage [CCP^{+}19]. Highly [CHK^{+}14a, yCBR05, SPP^{+}10, TTAG14, VHB^{+}13, ZVN05]. Hijacking [FGL^{+}19]. Hint [WQGR22]. Hint-Driven [WQGR22]. History [Shu19d]. HLS [AÖÖ23]. HLS-based [AÖÖ23]. HMAC [GWM16]. HMAC-SHA256 [GWM16]. HMT [SZG^{+}23]. hoc [KDN^{+}07]. HOL [MHT13].

Holistic [NFL^{+}22, OSA^{+}18]. home [LCQ^{+}13]. Homogeneous [NASM18]. Homomorphic [GCS24, MSR^{+}17, CMGM24]. Honey [ZGH^{+}19]. Hop [GDD20]. Horizontal [AAT^{+}21, RB21]. Host [RG14]. Host-Compiled [RG14]. HotSpot [WKJ20]. HRT [CQB^{+}15]. HSPA [HTB^{+}25]. Human [AMJ21, BTA^{+}19, DBX^{+}22, HZW^{+}23, HB23, HB24, LX22, MAS^{+}24, Shu19c, WXY^{+}18]. Human-Centric [MAS^{+}24]. Humanoid [GPT^{+}23]. Humans [QWY^{+}18]. HW [ZDTM19]. HW/SW [ZDTM19]. Hybrid [AK21, Ahm13, BP14, BDP24, BF17, CSW15, CMS17, DSXS^{+}14, DB19, DAASP21, HCL^{+}17, JRSR17, JMO14, KCWH14, KDR23, LP19, LWB18, LWK^{+}17, MM16, MRA^{+}17, QP15, RMD09, SFB23, SRB23, SB23, SLN^{+}16, SZG^{+}23, TSY^{+}16, WZJ^{+}18, WRW^{+}21, WZS^{+}24, WQGR22, WMLM12, YTL^{+}20, YLW15, ADI06, HZX^{+}14, JKJ^{+}10, LMST04, LO13, MLL08, RP11, RS07]. Hybrid-compiled [RMD09]. Hybrid-Mapping [CSW15]. Hybridation [DAASP21]. HyDREA [MEK^{+}22]. Hyperdimensional [GKS^{+}22, LAMS24, MEK^{+}22]. hyperelliptic [WPW^{+}04]. Hyperproperties [BMNN23, WZBP19].

I/O [CWH^{+}16, CCB^{+}06, EAAS22, JAD19, LSL^{+}23, MRY^{+}10, SKPL10, SC05, WGN23]. I4.0 [EHF24]. iAware [ZLX^{+}23]. IBBE [SXH^{+}19]. ICE [SDBD18]. ICNN [NHS20]. ICS [LZJ^{+}19, HL14]. Ideal [LHP^{+}23, LPO^{+}17, BMP23]. ideality [HKL^{+}23]. Identification [ZJZL20]. Identifying [BTD^{+}18, ZWH^{+}16]. Identity [HHL^{+}23b, SXH^{+}19]. Identity-Based [HHL^{+}23b, SXH^{+}19]. IDK [BBD23]. Idle [LS12]. IEC [YRS12]. IEC61499 [Kha13]. IEEE [DST19, HHL^{+}23b, TGP25]. IGOR [SPGT19]. II [PBP09b, SSA21]. IIR [FF09]. illegal [HT06]. Image [CWH^{+}16, DNBL22, KLK^{+}19, KNL12, PDHC23, SVC^{+}23, WZY^{+}23, XYLC23, PZ12, SCF12, SRY13, SWWY13]. Image-Content-Aware [CWH^{+}16]. image-media [SWWY13]. Image-Processing [WZY^{+}23]. Images [CPP^{+}17, LM24]. imaging [CCC^{+}14]. Imitation [SSB24]. immersive [LAHS06]. Impact [BTL^{+}12, GM12, LC17, WKJ20, HHB^{+}05]. Implant [PQA^{+}19]. Implementation [AGG^{+}23, BDP24, BSM^{+}21, BSJ15, CD17, CCP^{+}19, FHB^{+}17, FLF^{+}23, GHD^{+}25, HJ19, HKP18, HHL^{+}23b, HGL14, JGCS24, KY17, KKS^{+}23, LS20, LHP^{+}23, MFG16, RIMS21, SNTÖ25, SIR^{+}17, Seo18, SPC^{+}16, SFCW23, TGTT17, TV19, XHL^{+}24, CSST08, CLR05, HQB06, KASD07, LV09, WL09, WKC07, ZXS03]. Implementations [NPAG12, SSD^{+}19, SJLK18, ZSH^{+}19, DP08, SM13a, WGP04, YCLV^{+}02]. Implementing [AFG08, VOG15, YRS12, ZPZG17]. Implicit [DASS12, CHTC07]. Important [SPGT19]. Imprecise [JFM23]. impressionist [SRY13]. Improve [BHD15, RKK15, SWL^{+}23, FS14, RP11]. Improved [SLCS16]. Improvements [BBB16, HHC^{+}16b]. Improving [AK21, AB15, CSC23, GMCC18, HLF^{+}18, HLL^{+}23, JJS^{+}25, KJKM16, LS17, LPP^{+}21, LLP^{+}21, LSL^{+}23, LWCL25, PE23, SHME13, SC17, TVK08, WZD^{+}17, XQ07, YJD^{+}17, AC08, CW14]. Imputation [HB23]. IMS [KBCL13]. In-Memory [ADH^{+}23, HKL^{+}23, SLK^{+}22, BDD^{+}24, DLS^{+}24, YEK17]. In-Vehicle [XZK^{+}19, SKH^{+}12]. In/Near [SSA^{+}24b]. inActive [LKB14]. inaugural [Wol02]. Inclusive [FF25]. Increase [PJT^{+}23]. increasing [JHK^{+}06, SWL07]. Incremental [CJL17, DHKS15, Ise17, POCR24, NKP^{+}12]. Independent [CPP^{+}17, HQE20, SC20, HBSA04]. Index [KCC^{+}16, LCC^{+}19, LLC^{+}22, MKM^{+}23a]. indexing [PCBW13]. Individual [YTL^{+}20]. Indoor [GP23, PKIT23, TSW^{+}17, TM15, TP20]. Indoors [LYC^{+}18]. Inductive [Alg24, Ise17]. Industrial [JGX^{+}18, MSM21, SE23, UBF^{+}16]. Industry [Akd21, BCDD24, SXH^{+}19, Shu18b]. Inequalities [FS24]. Inertial [FGL^{+}19, HCS18, HXH^{+}24, WJ17]. Inexact [BDB^{+}17, LEPP13, PL13]. Infer [AGS^{+}16, WRB15]. Inference [BLSM19, CAN^{+}23, CSH^{+}22, CMMD24, DNBL22, GRR23, GDC19, HSD22, HKL^{+}23, JKH22, JSZ^{+}19, KML^{+}22, KBD24, MTWE20, MFG17, PVV24, QRW^{+}24, RKC^{+}22, SSK^{+}22, SPT^{+}23, WLZ^{+}23, WM24, YMKH23, ZSH^{+}24, ZDL22, QRB10]. Inferencing [NODP24]. Information [CBRZ19, GD24, HDZL20, LMST04, GLWM14, KTT13, YZA13]. Infrastructure [BLG^{+}15, GLT^{+}13, JBN^{+}13]. Infrastructures [LLK^{+}25]. Inherent [OSF19, YZA13]. inheritance [LLN09]. Initiated [LCL^{+}19]. Injection [ARP12, BCS^{+}23, MKMGS18, SISS24, YGD^{+}17, CMV10, RCDB24]. Injections [LCLW17]. Innovative [VP16]. Input [LM24, RR17, SFB23]. Input-Aware [RR17]. Input/ [SFB23]. Inputs [DPNA16, RLP^{+}21]. Insertion [LLC^{+}22]. inspired [KOL^{+}22]. Installment [SYC^{+}17]. Instant [LX12]. Instantaneous [MG05]. Instantiating [STB^{+}24]. Instantiation [PLT23]. Instantly [LKZ^{+}23]. Instruction [AJ18, ARP12, AB15, BCLS17, Fra12, ITO^{+}24, KAKSP15, LTQ^{+}24, QZXO14, SWX17, WSHC14, AC08, BP05, GRCV03, KVK^{+}03, LSC14, LLPM07, LM13, LXL13, MBFT09, RDM06, RMD09, RAK14, SD13, YZ08]. Instruction-Cache [AB15]. instruction-level [SD13]. instruction-set [AC08, RDM06, RMD09]. Instructions [DASS12, LPD^{+}20, NYH^{+}20, GGI13, KG05, SBX08]. Instrumenting [MZG14]. Integer [Rat25]. Integrated [EK12, FSC^{+}16, GMN21, GDD20, LSC19, dFMAdN12, LL18, MSCS16, PDL21, SXXS^{+}16b, XZK^{+}19, BvB13, MHK^{+}23]. Integrating [GIB^{+}12, SPP^{+}10]. Integration [LWZ^{+}16, MHT13, SWL^{+}14, CCB^{+}06, Dea06, KASD07, NKP^{+}12, SD13, WCJ07]. Integrity [BHL^{+}20, DBFH14, MCG22, ZZA^{+}22]. Intel [CMP^{+}07]. Intellectual [BS22]. Intelligence [MBCM22, MFG17, Shu18b]. Intelligent [LHL^{+}19, Pau14, RMK17, ZLZ^{+}24, LKW02]. Intensive [MLR^{+}17, TDD^{+}16]. Inter [LZS20, PVSG22]. Inter-app [LZS20]. Inter-Node [PVSG22]. Interaction [ZLX^{+}23]. Interactive [BC07, Bro21, LL15, KBCL13, KZH^{+}06, LCQ^{+}13, PCBW13]. Interactivity [WWT^{+}22]. interconnect [JP14]. Interconnection [SXXS^{+}16a]. Interconnection-Aware [SXXS^{+}16a]. Interconnects [CFGM15, RPB^{+}19, WMZY13]. Interface [SH15, LCQ^{+}13]. interfaces [NNH^{+}14, ZL08]. Interfacing [SIC19]. Interference [NS16, SRR^{+}23, WZD^{+}17, XSP22, BHM17, RP10]. Interference-aware [XSP22]. interference-free [RP10]. Interlaced [LCC^{+}23]. Interleaved [WSMF22]. Interleaving [BB13, FSC^{+}16]. Intermediate [KPK^{+}19, MDWL23, RMBS20]. Intermediate-Centric [MDWL23]. Intermittent [ABA^{+}20, ABTS24, CAN^{+}23, EYG^{+}23, HLL^{+}23, KML^{+}22, KS22, MH19, YMKH23]. Intermittent-Powered [HLL^{+}23]. Intermittently [JRSR17, RN18]. Intermittently-Powered [JRSR17]. internal [CW14]. Internals [CKN^{+}20].

International [DST19]. Internet [BCHL19, BHXP19, BGJ17, KLK24, RRM16, SXH^{+}19, Shu15a, SYC^{+}17, ZSY19]. Internet-of-Things [BGJ17]. Interpolation [CLS16]. Interposer [SPT^{+}23]. Interpretable [PDL21]. interpretation [RRW05]. Interpreter [CMPP23]. interprocess [TKD07]. Interrupt [FND^{+}16, LP10, dFMAdN12, WCM^{+}16]. Interrupt-Driven [WCM^{+}16, LP10]. Interrupt-Triggered [FND^{+}16]. Interruptible [CZ23]. Interrupts [LMK^{+}18]. Intersection [LHL^{+}19]. Interval [LZL15, PLT23, LXK10]. Intrinsic [BFW^{+}19, MFMA17]. Intrinsically [SRK^{+}18]. Introducing [SBF^{+}05]. Introduction [BCHL19, BCEP12, BM13, CP13a, CDBB24, CKGN14, CC14, CBH22a, CBH22b, CP13b, DV13, DGG24, DSD12, Edi13, FM12, GV21b, Goe14, GP07, HCK^{+}08, HTLC10, Hüb13, JB02, JB03, JLSK13, KS10, KL13, KM13, MS05, OMMK23, PCB12, RHG^{+}14, SOG15, STLX22a, STLX22b, SSA^{+}24b, SCZ20a, SCZ20b, STW13, TFC24, Wol02, Sch07]. Intrusion [CLL21, WDY^{+}16, ZCG^{+}22, ZJZL20, LHCK04]. Intrusive [AARJ12]. Invariant [BP12, MS23, SC17]. Invariant-Based [BP12]. Invariants [AGS^{+}16, AGG^{+}17]. invasive [FSVG19]. inversion [IBMK10, KHHH14]. Investigation [LKW^{+}24, WRB15]. INVISIOS [AARJ12]. Invited [DSXS15]. IOSR [LSL^{+}23]. IoT [ABL^{+}20, AAR^{+}17, BZG19, BLSM19, CZ23, CDBB24, CBH22a, CBH22b, CCM17, GAS^{+}17, GSN21, JRSR17, LZS20, LKZ^{+}23, LWCL25, LZZ^{+}19, MFG17, MPFG19, PP19, PJL^{+}17, RSW21, SKS21, SJLK18, Shu17b, TTB23, TP19, TNR17, WX17]. IoT-Fog-Cloud [SKS21]. IoT-to-Edge-to-Cloud [CDBB24]. IP [CCB^{+}06, RBNM19, SM13a, TKL^{+}15, WCJ07]. IPs [BRL16]. IPSec [SKW^{+}07]. Irons [Shu16d]. ISA [CYH^{+}17]. islands [FZHT13]. ISO [AKB^{+}25]. Isolation [AHMT17, RWL^{+}18]. iSSD [SFCW23]. Issue [BBM15, BCHL19, BDR24, BDK^{+}23, CDBB24, CS16, CKGN14, CBH22a, CBH22b, CJL17, CGZ18, DGG24, DST19, DSXS15, EE16, EH18, FX17, GV21b, Goe14, IT16, KL13, Leo18, LZNL23, MCP17, OMMK23, RHG^{+}14, STLX22a, STLX22b, SSA^{+}24b, SCZ20a, SCZ20b, SCKD23, TEC12, TFC24, VP16, WX17, WSHC14, ZQC16, BM13, DPP14, GM03, Gup04, GP07, HCK^{+}08, HTLC10, JC03, KS10, KBCL13, LB04, MS05, DWCM14, PBP09a, Sch07, SL04, ST05, Wol02, PBP09b]. Issues [Shu15c, JB02, JB03]. iSupplemental1 [TEC12]. Iterational [XHSS10]. Iterative [NHS20, SAHE04, BWS14, KFY^{+}22, PS08a]. Itself [Shu16b]. ITUbee [FXP^{+}17]. IXP [LCH^{+}08].

Java [ABC^{+}07, BVGVEA10, CWZ^{+}20, CSK^{+}02, CH08, CRAJ10, GW08, HT06, HTLC10, JMO14, KW10, MS13a, PS10, SKKR11, SPP^{+}10, TKL^{+}15]. Java-based [GW08, JMO14]. Jetson [JKH22]. Join [SGW^{+}16]. Joint [HZGW18, HZX15, LMS^{+}22, LXL13, LYY^{+}17, PKL22, WC16, XLY18]. JOM [WC16]. JPEG [THON12]. JSCD [YC12]. Jump [PP12]. JVM [WKJ20].

Karatsuba [MSR^{+}17]. Keep [YMKH23]. Kernel [CSC23, LL17, WRB15, CDD^{+}07]. Kernel-Level [WRB15]. kernels [PGS^{+}13]. Key [AAT^{+}21, DL12, KNK^{+}25, MKM^{+}23a, PNRC17, Seo18, SAKH20, PS08b]. Key-Length-Based [PNRC17]. Key-value [MKM^{+}23a]. Keyword [GV21a]. Kit [JAB^{+}22]. knapsack [YCNCC11]. kNN [SM13a]. Knowledge [HWC^{+}20]. KNOWME [TLL^{+}12]. Kryptonite [SRR^{+}23]. Kubernetes [LPBF24]. KV [SLZ^{+}24, ZLSQ17]. Kyber [JGCS24, RCDB24].

L [EAAS22]. L24 [SM13b]. Lab [BCHB18]. Lab-on-Chip [BCHB18]. LaDy [CCSC23]. Lagrange [YF19]. LAMBDA [KAS^{+}20]. LanCeX [XYLC23]. Lane [KCBM21]. Language [CMPP23, LFC17, SIR^{+}17, MMD04]. Languages [DGG24, GV21b, SCZ20a, SCZ20b, WWN23, LP09a]. Large [CJL17, JGX^{+}18, LZZ^{+}23, MRA^{+}17, PE23, HHB^{+}05, PS08b]. Large-Scale [CJL17, JGX^{+}18, LZZ^{+}23, PS08b]. LARK [DS11]. Last [KRS^{+}16, MPT^{+}22, TTA^{+}20, WZJ^{+}18, WKP24]. Last-Level [KRS^{+}16, WZJ^{+}18, MPT^{+}22, WKP24]. Latency [AYS15, CSC23, DS24, GRRR24, GTvdBC25, HKP18, KSY17, KH23, LPE^{+}23, MV16, MSRM24, QFL^{+}24, WLTW24, ABI^{+}09, SRM^{+}13, XHSS10]. Latency-Aware [BZG19]. Latency-based [HKP18]. Latency-Optimized [AYS15]. Latent [VAR13]. Lattice [AYS15, BSJ15, DLS^{+}24, HPO^{+}15, LHP^{+}23, LPO^{+}17, MWK^{+}24, RCDB24, VF17]. Lattice-Based [AYS15, BSJ15, HPO^{+}15, LHP^{+}23, LPO^{+}17, MWK^{+}24, RCDB24]. Launch [KJKM16, CLK13]. Law [AKI^{+}23]. Layer [BDG^{+}15, CCC^{+}20, CKB17, JCW^{+}16, Kwo16, SKKR11, CYKH13, CCY^{+}13, KST^{+}12, KXL10, LPC^{+}07, PCK^{+}08, WKC07, Wu10, ZP09, JKJ^{+}10]. Layers [AP20, PBC22, UGS^{+}21, WWT^{+}22, XDL^{+}18, ZQ24]. Lazy [KGT^{+}23]. LCTES [FX17]. LCTES’05 [GP07]. LCTES’11 [DV13]. LDPC [LJ14, WZD^{+}17]. Leakage [CBRZ19, JDW24, SP19b, CNK04, ZKKC05, ZTD^{+}06, ZA07]. Leaks [DLV16]. LEAP [MSR^{+}12]. Learn [GKS^{+}22]. Learning [AHM19, Alg24, AZHC19, BJ23, BGK^{+}23, BLSM19, CZHK23, CK23, DHB^{+}23, EHF24, GTH^{+}22, GDB22, GLS^{+}23, GP23, HWC^{+}20, HZW^{+}23, JKH22, KSY17, KPS23, KCBM21, KNK^{+}25, KCCW17, LX22, LJ24, LLK^{+}25, MTWE20, MEK^{+}22, NBH23, NYH^{+}20, OBA^{+}17, ORA16, PVSG22, PDL21, Pau14, POCR24, RB21, RALR24, RLG20, SFB23, SSA^{+}24a, SR12b, SSK^{+}22, SWS23, SSB24, SKN17, Shu18b, STB^{+}24, TP20, TCD^{+}19, ZQD^{+}23, ZK25, KR14, SBF^{+}05]. Learning-Assisted [KSY17]. Learning-based [AZHC19, TP20]. Lebesgue [MHT13]. Ledgers [Shu16a]. Legacy [SWL^{+}14, CCAP12]. legaSCi [SWL^{+}14]. LegUp [CCA^{+}13]. Length [PNRC17, Sus20, BAR13b, KD08, PL10]. LEON [PDL^{+}23]. Less [AKTM16, KML^{+}22, BYD09, PLKH08]. Let [JSD23]. Level [BRL16, FLF17, KPC^{+}16, KBS17, KHB^{+}23, KRS^{+}16, LN19, LMK^{+}18, LYH^{+}15, LZJ^{+}20, MFMA17, MF12, NBM^{+}16, PKT23, PMDC17, SSA21, SDMK19, TP19, TWTH18, TGP25, TTA^{+}20, WZJ^{+}18, WRB15, ZRF^{+}12, ZYM16, ZYL^{+}17, AVR22, BAR13b, CCA^{+}13, FO03, IYL^{+}23, JBN^{+}13, KKC^{+}05, KVN^{+}09, MSCJ12, MPT^{+}22, MSS^{+}03, MSL13, OP06, RDSS21, SGT^{+}13, SD08, SD13, VJD^{+}07, VDK^{+}08, WKP24, YCLV^{+}02, ZEJ^{+}23, ZBG20]. Level-II [SSA21]. Leveling [HCS^{+}22, CCH13, PMPP14]. Levels [KS22]. Leveraging [HMLZ21, HRT^{+}22, MF13, MMY^{+}19]. LiBrA [GMVV17, JNI15]. Libraries [STB^{+}24, ZGH^{+}19, PLKH08]. Library [BCC^{+}17]. Licensing [RBNM19]. LiDAR [HXH^{+}24]. Lifetime [GM12, SHQX19, LO13]. Light [OHCK24, ARH^{+}18]. Lightweight [AMKA17, AARJ12, BDB^{+}17, BCHL19, BPSS24, CBS19, DS11, GMVV17, HDZL20, JCS^{+}17, KAS^{+}20, KSP^{+}12, KCBM21, KRR20, MFG16, OMH^{+}23, RCS23, SL16, SP19b, XYLC23, XQQ^{+}24, ZSY19, ZSH^{+}19, ZGH^{+}19, PS04]. like [BCS16, LJ14]. Limit [VSD^{+}17]. Limitations [MKE18]. Limited [HLLL20, BS13a, CH08, HHL^{+}23a, Wu10]. LIN [SKH^{+}12]. Linear [BF17, GD19, JSZ^{+}19, JDW24, LZZ^{+}23, KJRG13]. Lines [TR24]. Lingua [LMBL21, LML^{+}23]. Link [DVC^{+}07, JJS^{+}25, KXL10]. Link-time [DVC^{+}07]. Linked [PqBM^{+}15]. links [QRB10]. Linux [BMF15, CDD^{+}07, MZG14, TGP25]. LiSP [PS04]. Literature [RH23]. Live [FND^{+}16]. Live-Out [FND^{+}16]. Liveness [GZ12, WWY13]. Liveness-Enforcing [WWY13]. LL [QFL^{+}24]. LL-GNN [QFL^{+}24]. LLM [BS13a]. LMP [WSK14]. LMP-based [WSK14]. Load [CWJ17, JBI17, KGT^{+}23, UM13, WLW^{+}24, Mus10, ZP06]. Load-balanced [WLW^{+}24]. Load-Balancing [CWJ17, Mus10]. Load-Store [JBI17]. load/stores [ZP06]. loader [WBF^{+}06]. Local [DNBL22, KAKSP15, LBS15, BS13a]. Locality [CCSC23, GFC^{+}10, KK05a, YG02, MMK22]. Locality-aware [CCSC23]. Localization [BKS^{+}23, GP23, MMD22, SKL^{+}25, SHL^{+}17, TP20, YV23, ZH12a, BHET04, CTK^{+}13, HHB^{+}05, LLL14, PS08a, PSZ12b, ZH12b, ZC04c]. Location [LLT^{+}17, TM15, ZHM^{+}14]. locations [PS08a]. Lock [CRJ10, PCM^{+}15, SA18]. Lock-free [CRJ10]. Locked [SMR15]. Locking [AB15, DLD^{+}19, QZXO14, SWK19, ZW17, VLX07]. LOCUS [TKV^{+}18]. Log [LM24, SHQX19, LPC^{+}07, TSG10]. Log-gradient [LM24]. Logging [CSW15, CSCC17, DLH16, GSS^{+}18, MWF^{+}16]. Logic [AFS^{+}13, Alg24, KMP15, KDB19, LJMP23, MKS^{+}17, RLMP23, VRF15, WRW^{+}21, LLLGR13, ETAV16]. Logic-Based [ETAV16]. Logical [DGC^{+}20, KHB^{+}23, ZWK23]. Logics [BMNN23]. Long [DNBL22, GSS^{+}18, JC12, KSY17, WLTW24, DLC^{+}14]. Long-range [DNBL22]. Long-Tail [KSY17, WLTW24].

Long-Term [GSS^{+}18, JC12, DLC^{+}14]. Look [BCC^{+}17, KKS^{+}23, WZH13]. Look-Ahead [WZH13]. Look-up [KKS^{+}23]. Lookup [RR17]. Loop [MS23, NZCS19, PQA^{+}19, SFZX18, TWTH18, VGN18, WWHT21, DEG11, GGI13, KVN^{+}09, NNS13, TKD07, XHSS10]. Loop-Invariant [MS23]. loop-level [KVN^{+}09]. Loop-Oriented [SFZX18]. Loops [DKA^{+}19, EK12, TdS05, TBDdD11, SVN04, ZA07]. Loosely [AD24b, BBB16]. Loosely-Timed [AD24b]. LoRa [FJM^{+}25]. Lossless [EAAS22, KCBM21, LLW^{+}24, WCK^{+}19]. Lossy [EAAS22, WCK^{+}19]. Low [ABL^{+}20, ABC^{+}17, ABI^{+}09, BHD15, BDP24, BTA^{+}19, CZ23, CS22, EHF24, GAS^{+}17, HJ25, JRR16, JAB^{+}22, KYDC20, KSA^{+}18, LMK^{+}18, LZZ^{+}19, LPE^{+}23, MAS^{+}24, MSRM24, NBE18, PAF22, QFL^{+}24, RKM24, SWJ^{+}13, SJC^{+}03, SBK^{+}23, SCM20, SLK^{+}22, SR19, TKV^{+}18, YC12, ZRZ^{+}19, BDB^{+}17, CCH13, DBH14, Geb06, GJ13, GRCV03, GLWM14, IHK04, KYHY14, LWB13, NPP13, ÖNG08, PKIT23, RAK14, SJRS^{+}13a, TTB23, TTAG14, TVK08, ZCK13, ZVN05, ZP09, MSR^{+}12]. Low-Cost [ABC^{+}17, CS22, GAS^{+}17, LZZ^{+}19, EHF24, CCH13, SJRS^{+}13a, ZCK13, ZP09]. Low-end [CZ23]. Low-Energy [MAS^{+}24, SJC^{+}03, Geb06, LWB13]. Low-Latency [LPE^{+}23, ABI^{+}09]. Low-Level [LMK^{+}18]. Low-Memory [BDP24]. Low-Power [NBE18, TKV^{+}18, YC12, PAF22, RKM24, SCM20, SR19, GJ13, GLWM14, IHK04, KYHY14, NPP13, ÖNG08, RAK14, TVK08]. Low-Voltage [SWJ^{+}13]. Lower [ZX08]. LPWAN [RIMS21]. LRU [GLYY14]. LSM [SLZ^{+}24]. LSM-tree-based [SLZ^{+}24]. LSTM [AV20]. LTE [AAPN14, VKMP20]. LTE/5G [VKMP20]. LTL [BRR19]. Lustre [BGGT23, GSSR^{+}24, TSCC05].

LWE [NVB^{+}20].

M2M [Pau14, RRM16]. MAC [BTL^{+}12, CHTC07, GDA13, LCL^{+}19, ZWY^{+}10]. Machine [APRC16, AHM19, CHS15, DHB^{+}23, EHF24, GTH^{+}22, GDB22, GLS^{+}23, KKCS16, KCBM21, KBRD22, LAZ^{+}16, MEK^{+}22, MFG17, NBH23, NYH^{+}20, OBA^{+}17, PDL21, PG24, RLG20, SSA^{+}24a, Shu18b, STB^{+}24, ABC^{+}07, CGV10]. machine-based [CGV10]. Machine-Learning-Resilient [NBH23]. Machine-to-Machine [APRC16, KKCS16, LAZ^{+}16]. Machines [BPP23, CMPP23, DQ14, KCWH14, ZPZG17, CH08]. macromodeling [LBP07, TRJ05]. made [PDL^{+}23]. MaGNAS [OBO^{+}23]. Magnetic [CPP^{+}17, CCSC23, HCS18, ISOD21, LCC^{+}19, LCC^{+}23]. MAGNETO [ISOD21]. Main [AVR22, HCS^{+}22, PXY^{+}17, SJOL22, WLWS15, WZJ^{+}18, HXZ^{+}13, PMPP14]. Maintaining [LLR14, KDN^{+}07]. Majority [BDD^{+}24, NASM18]. Majority-Based [NASM18, BDD^{+}24]. Making [LPE^{+}23, WCH^{+}23]. Malware [KJLS20, KAS^{+}20, LWCL25, RCS23, Rru22]. manage [CRM14]. Managed [HCS^{+}22, LBS15]. Management [ABD^{+}19, BMF15, CSW15, DAHM16, DSXS15, ESM^{+}17, ESBK23, FBM16, HB16, HNY18, HZX^{+}14, HHC^{+}16a, IDO^{+}22, KNY^{+}17, KBS17, KJK18, KR18, dFMAdN12, LZL15, LL17, LCC^{+}23, LHL^{+}19, MLL^{+}17, MMY^{+}19, MBJ^{+}23, NEP23, NRZ^{+}24, OMMK23, OZ22, PVSG22, PYJL15, Pau14, RC17, RALR24, RJM19, SPT^{+}21, SSPP23, SKN17, SP19b, SBK^{+}23, TDD^{+}16, TMXS17, TAMS18, VGN18, VCM19, WLWS15, WDM17, WZJ^{+}18, WWT^{+}22, WLC^{+}22, WQGR22, ZP11, AMCM06, ACK^{+}13, BDP^{+}13, BBL09, CCY^{+}13, CH08, ELS08, FZJ08, ISG03, JKH^{+}13, KHZS07, KR14, KXL10, MPZS13, RV03, SGT^{+}13, SRS03, WYS^{+}13, YCNCC11, ZC04b, Zhu10]. Manager [DAHM16, MDS^{+}21, CH10]. Managers [REPL15]. Managing [CRCR13, DRL^{+}10, MLR^{+}17, BS13a]. manner [SRY13]. MANTIS [MLV09]. Manual [LL15]. Manufacturing [GM12, VWG^{+}17]. Many [APF24, CCC^{+}14, CLLC17, JAD19, LKA^{+}18, MKD15, RWL^{+}18, RJM19, SDBD18, SXXS^{+}16a, SXXS^{+}16b, SXMX^{+}18, TDD^{+}16, TKV^{+}18, TMXS17, TAMS18, VCM19, VKMP20, ACK^{+}13, DPP14, LKB14, LOG^{+}14, LLR14, YFPJ14]. Many-Accelerator [SXXS^{+}16a, SXXS^{+}16b, SXMX^{+}18]. Many-Core [LKA^{+}18, MKD15, RWL^{+}18, RJM19, SDBD18, TKV^{+}18, VCM19, CCC^{+}14, CLLC17, JAD19, VKMP20, ACK^{+}13, DPP14, LKB14, LOG^{+}14, LLR14, YFPJ14]. Many-Cores [TDD^{+}16, TMXS17, TAMS18]. Many-to-Many [APF24]. Manycore [DJJ^{+}19, LLG^{+}20, NODP24, KYL13]. Map [TKT15]. MaPHeA [OMH^{+}23]. Mapping [BKS^{+}23, ABF^{+}21, BJ23, BRA^{+}16, CSW15, CLL16, CPC17, CCC^{+}20, DMPC23, ETAV16, FSC^{+}16, FC16, GIB^{+}12, GAG15, HC16, JRSR17, JSD23, LX16, MSCS16, NASM18, OBO^{+}23, PP24, PJWY12, QP15, RH23, RLP^{+}21, SB23, SKL^{+}25, SPB^{+}17, TWTH18, WWG^{+}18, YLTY21, ZNS13, DKV14, HH13, LWB13, MEP08, MAG14, OMA^{+}13, WW09]. Mapping-Aware [OBO^{+}23]. March [SN10]. Market [ZLF13]. Market-based [ZLF13]. Markov [GGJ12]. Marriage [RPHA19]. mask [Geb06]. Masked [WH17]. massive [Edi14, Mus10, ZXCH13]. Massively [GLP^{+}11, TWTH18]. Matching [CYH20, PMP17, UFA^{+}24, LHCK04, TLLL09]. MATLAB [LPD^{+}20]. MATLAB-to-C [LPD^{+}20]. Matrix [AGG^{+}23, FKJM18, GOC^{+}22, LZZ^{+}23, IBMK10]. Maximal [VRF15, HCQ^{+}14]. Maximally [WZH13]. Maximisation [DCZB19]. Maximising [IDO^{+}22]. maximization [HCQ^{+}14]. Maximizing [MASG15, PP24, RMM03]. MC [LCP^{+}17]. MC-ADAPT [LCP^{+}17]. McEliece [MBR15, VOG15]. MCU [GHD^{+}25]. MCUs [ABL^{+}20, JRSR17]. MDP [SWS23]. MDPC [VOG15]. Me [SPGT19]. Measure [MHT13]. Measurement [BYIG21, FGL^{+}19, ZO16, LYL13]. Measurements [ITO^{+}24]. Measures [FKJM18]. Measuring [DW10, YGD^{+}19]. Mechanism [CAPL11, FFA^{+}23, LCL^{+}19, WLZ^{+}23, WC16, YZZ^{+}23, CWKH12, RAK14]. Mechanisms [AbSZ^{+}19, CJL17, KNK^{+}25]. Mechanized [RPHA19]. media [HE12, SWWY13]. Median [LLW^{+}24]. Median-based [LLW^{+}24]. Medical [LJ24, MS13b, PJL^{+}14, KLC^{+}10]. medicine [WYS^{+}13]. MEDiSN [KLC^{+}10]. Medium [KKCS16]. meet [SRM^{+}13]. Meeting [LZZK25]. meets [BSKB^{+}09]. Mellon [KCG^{+}05]. MemFHE [GCS24]. MEMMU [BYD09]. MEMOCODE [DST19]. Memories [CDX^{+}19, KRHC20, KOL^{+}22, PqBM^{+}15, PRL^{+}23, SP19b, SDMK19, WLWS15, WCH^{+}23, BMP03, HXZ^{+}13]. Memory [AD24a, ADH^{+}23, AD24b, AVR22, BDP24, BLSM19, BCS^{+}06, BP19, BCS^{+}23, CBH22a, CBH22b, CI17, DPNA16, DKAL05, EAAS22, FLF17, FSC^{+}16, FLF^{+}23, FMSS15, GIB^{+}12, GAG15, GAS^{+}17, GGMS24, GCS24, HCS^{+}22, HKP18, HKL^{+}23, JGL21, JRSR17, JLW^{+}15, KPS23, KKK^{+}11, KS13, KJKM16, KNY^{+}17, KBS17, KRR20, LSL^{+}23, LWZ^{+}24, LYH^{+}15, LWB18, LBS15, LOF20, MDS^{+}21, MBKF15, MF12, NYH^{+}20, NDB09, OMH^{+}23, OZ22, PXY^{+}17, PP19, PP24, PMM^{+}17, PMDC17, PRM21, RC17, RRC22, RKC^{+}22, RSK17, SSK23, SWJ^{+}13, SSD^{+}19, SSPP23, SSA^{+}24b, SJOL22, SBK^{+}23, SLK^{+}22, SR19, Sus20, TDD^{+}16, TBG^{+}17, TGBT17, VCM19, VKW^{+}17, WDM17, WZJ^{+}18, WCB20, WWT^{+}22, WLZ^{+}23, WQGR22, WSMF22, WC16, WHL23, YCJW24, YYKK18, ZDZ14, ZQGZ22, ZZA^{+}22, ACK^{+}13, ABS02, BCLN13, BS13a, BCDH12, Bar13a, BAR13c, BCBJ24, BDD^{+}24, CH10, CDD^{+}07, CKL04, CWKH12]. memory [CYKH13, CC13a, CSK^{+}02, CH08, CVG^{+}13, DLS^{+}24, ELS08, GDN03, HFG13, HH13, HZX^{+}14, HL14, JB02, JB03, JKH^{+}13, KYL13, KGR12, LKW02, LO13, LXK10, LXL13, LPB06, MMD04, PLKH08, PK13, PMPP14, RP03, SGT^{+}13, SE10, SBX08, SJC^{+}03, UDB06, UCK^{+}09, WAD14, WKC07, XHSS10, YDLC10a, YDLC10b, YEK17, ZP08, ZP06, BYD09]. Memory- [BLSM19]. memory-based [CC13a, HZX^{+}14]. Memory-Constrained [JGL21, KPS23, LWB18, Bar13a]. Memory-Efficient [SSD^{+}19]. Memory-Intensive [TDD^{+}16]. memory-limited [CH08]. Memory-Model-Aware [FMSS15]. Memristive [YEK17]. Memristor [MCS^{+}15, ZK25]. Memristor-Based [MCS^{+}15, ZK25]. Mental [HZYJ22, WGPH13]. Merged [BZY^{+}23]. Merging [PRSV19]. Merkle [SZG^{+}23]. Mesh [AKI^{+}23, MSCS16, BP14, Bec09, SJRS^{+}13a]. Mesh-Based [MSCS16]. mesh-connected [Bec09]. Message [HM17, JJS^{+}25, KHHH14, LZJ17, XZK^{+}19, LBP07]. Message-Processing [XZK^{+}19]. Messages [ZSEP21]. Metadata [ZZA^{+}22]. Metal [BYIG21]. METEOR [BP14]. meters [Edi14]. methanol [SPK^{+}12]. Method [AGS^{+}16, AGG^{+}17, EVS^{+}17, FGL^{+}19, FFA^{+}23, GW15, HWTL23, HB23, ITO^{+}24, KCBM21, SXH^{+}19, WLTW24, XYLC23, YZZG23, CCB^{+}06, KHHH14, LWB13, LO13]. Methodologies [IT16, ST05]. Methodology [FSC^{+}16, GDDD17, HJ25, JKH22, NYH^{+}20, OBSO16, PSZ12a, SK19, TSW^{+}17, TGV12, TAP23, WWG^{+}18, DEG11, KST^{+}12, LAN06, Shu14b, XWHC06]. Methods [BBMG24, DST19, HHC^{+}16b, JR20, KCCW17, Leo18, Mos13, Pau14, TB23, VP16, AC08, SHME13, WEE^{+}08]. Metric [GZ12, PDHC23]. metroII [DDG^{+}13]. MHDeep [HZYJ22]. Micro [EZL^{+}17, JC12, MB10]. Micro-Electrode-Dot-Array [EZL^{+}17]. micro-satellite [MB10]. Micro-Solar [JC12]. Microarchitect [KJK^{+}17b]. Microarchitectural [DJO12, MPT^{+}22, SGZS21]. microarchitecture [NB04]. MicroBlaze [LV09]. MicroBlaze-based [LV09]. Microcontroller [PRM21]. Microcontrollers [CZ23, CI17, DBX^{+}22, JRR16, LPO^{+}17, MKS23, SWJ^{+}13, YLDM19, Sch10]. Microfluidic [BCHB18, CKB17, EZL^{+}17, SIC19]. Microprocessor [KE15]. microprocessors [RAK14]. Microsearch [TSWL10]. Microserver [MBB^{+}15]. microthreaded [YFPJ14]. microwave [CCC^{+}14]. migration [LP10]. Milner [VRF15]. Miniaturized [MVS^{+}13]. Minimal [CL13, GAD^{+}24, MKM^{+}23b, SBX08, Edi14, GNW05]. Minimal-Overlap [GAD^{+}24]. Minimally [AARJ12]. Minimising [TGBT17]. Minimization [HZX15, SSK21, SIC19, PAP^{+}12, ZX08]. Minimize [YCT16]. Minimizing [BBL09, GRRR24, LLZ^{+}17, SPDLK^{+}17, ZDZ14, ZQGZ22, ZW17, GNS04]. Minimum [ABD^{+}19, KAK05]. minimum-energy [KAK05]. Mining [BMNN23, GZZ^{+}16, KDB19, NCJF18, PMAB19, SC17]. MIPS [LCS03]. Mirroring [PX18]. Mirroring-Assisted [PX18]. Miss [NS17, MEP08]. Misses [ZLL^{+}18]. Missing [PMAB19]. Mitigate [KSY17, MMK22, SE23]. Mitigation [SUS^{+}17]. Mixed [AKTM16, ABS^{+}19, CYH^{+}17, FHB^{+}17, GE18, HPP17, HHC^{+}16a, KGT^{+}23, LCP^{+}17, LH18, LJVD23, SSD^{+}19, SLK^{+}22, TSP15, TGTT17, ZZG24, ZGZ15, ZQGZ22, ZDL22, HGL14, LDRM12]. Mixed-Criticality [AKTM16, GE18, HHC^{+}16a, LCP^{+}17, LH18, TSP15, TGTT17, ZZG24, ZGZ15, ABS^{+}19, FHB^{+}17, KGT^{+}23, ZQGZ22, HGL14, LDRM12]. Mixed-Precision [SSD^{+}19, ZDL22]. Mixed-Signal [SLK^{+}22]. Mixture [BCHB18, TJ23]. ML [TTB23]. MLC [CYKH13, NBE18]. MLC-based [CYKH13]. MLC-PCM [NBE18]. MLOps [MMA^{+}23]. MMU [BYD09, ELS08, PLKH08]. MMU-less [BYD09, PLKH08]. Mobile [CWH^{+}16, CHJ22, CZH^{+}24, CSC23, EMVR23, GQC^{+}17, GP23, HTC^{+}16, HLLL20, IDO^{+}22, JBDD20, JCS^{+}17, KCJ^{+}16, KJK17a, KJK18, KNL12, LDV12, Li21, LSL^{+}23, LKH16, LMW^{+}17, LNA^{+}15, MV16, PX18, PHDL18, SBR^{+}15, SJOL22, Shu17c, TP20, WTSR13, WLH16, WQGR22, WZG^{+}23, XDL^{+}18, YTL^{+}20, YDS^{+}22, ZLX^{+}23, BO13, CTK^{+}13, CLK13, FZJ08, ISTE08, ISE10, KSK13, KST^{+}12, LLL14, LCJ13, NNH^{+}14, PK13, RC08, VAR13, WRJL06, WYP^{+}10]. Mobiles [HLM^{+}24]. MobiSense [WYP^{+}10]. Modal [BV15, SH15, WBS10]. Mode [ABS^{+}19, DCZB19, JRR16, Rat25, yCBR05, SR19, YLTY21, ZTRC03]. Mode-dependent [ABS^{+}19]. Model [AD24a, ARS16, ARDG16, AAM^{+}17, AKB^{+}25, AAS18, BLSM19, BRL16, BJH^{+}24, CWZ^{+}20, FKS^{+}19, FSB^{+}21, FGK^{+}23, FMSS15, GLP^{+}11, GGJ12, HE24, IAS23, IVJ^{+}23, JFM23, KML^{+}22, KSS24, KH23, KFY^{+}22, KDR23, KBRD22, LC17, LAZ^{+}16, LSL20, MTWE20, MV16, MMA^{+}23, PDL21, PNRC17, SSD^{+}19, Sch10, SWL^{+}14, SOL^{+}16, SP20, SSB24, TJ23, TBFR17, TBCB15, WRW^{+}21, WZ12, CJMB05, DRL^{+}10, KKH^{+}12, OMA^{+}13, PJL^{+}14, RSB^{+}09, SL08, WMZY13, ZS05, BE10]. Model-Based [ARDG16, BRL16, KH23, LSL20, TBFR17, BJH^{+}24, FKS^{+}19, TJ23, KKH^{+}12, OMA^{+}13]. Model-Driven [CWZ^{+}20, GLP^{+}11, DRL^{+}10, RSB^{+}09, BE10]. Model-Predictive [TBCB15]. Modeling [BONA22, Bro21, Fra12, GFC^{+}10, HMM04, KSS16, KE15, KL13, KYDC20, LLLT08, LLTL09, LHL^{+}19, McI13, MKD15, MD04, MAGR15, NKP^{+}12, NDZ13, NBM^{+}16, PJL^{+}17, RHG^{+}12, SRSM21, TKHZ22, YKD^{+}24, ZYM16, ZTZ^{+}19, ASTPH10, MG05, SD08, SPK^{+}12, VJD^{+}07, VDK^{+}08, WW09, VAHC^{+}06]. Modelling [DAASP21]. Models [ABH^{+}18, AD24b, BTD^{+}18, BHM17, CD12, CD19, DST19, DVC21, GSSR^{+}24, HYY^{+}15, IT16, JBDD20, Leo18, LMS^{+}22, LZJ17, MAKO19, PRSV19, PMP17, SRB23, SBLM13, SGJ17, SGW^{+}16, TB23, YSC22, CC13a, DP08, HDR^{+}06, HVG13, LLC^{+}13, ST05, ZMB03]. Modern [BMB16, HA24, DFC^{+}19]. Modes [PXY^{+}17]. Modular [IAS23, MRA^{+}17, TBG^{+}13, CMGM24]. Modularization [LPFL16]. Module [BCS16, ARJ08, PAS^{+}09]. module-based [PAS^{+}09]. modulo [SWWY13]. Molen [PBV07]. Momentum [BFW^{+}19]. Monads [RPHA19]. Monitor [PDL^{+}23, MVS^{+}13]. Monitoring [BCDD24, BRR19, BFST19, Edi14, HHC^{+}16b, MBKF15, SKS21, TLL^{+}12, TBCGO23, VFS^{+}21, GJ13, GNR^{+}10, WYP^{+}10]. Monitors [BMMV21]. Montgomery [SLN^{+}16, SAKH20]. Moore [AKI^{+}23]. MOOS [DJJ^{+}19]. MORPHEUS [VHB^{+}13]. Motion [HPBL12, LHYQ18, MS23, PJWY12, WXY^{+}18, YW13]. Motion-Based [HPBL12]. mountain [VS05]. Moving [QWY^{+}18]. MP [JBN^{+}13, YCNCC11]. MP-SoC [JBN^{+}13, YCNCC11]. MPPT [BFW^{+}19]. MPSoC [CFGM15, CMP17, IDO^{+}22, LYH^{+}15, LBP07, MBLA16, OBO^{+}23, PAP^{+}12, PGR^{+}08, TBAS17]. MPSoCs [BLG^{+}15, AMN^{+}14, BMB16, CAP^{+}07, CMMD24, DJ23, GLT^{+}13, HHB^{+}12, KBDV08, KGT^{+}23, LPB06, MAG14, MASG15, OMA^{+}13, QP15, SPB^{+}17, TBG^{+}17, WWG^{+}18, GPB^{+}17]. MPSoCSim [WRKG16]. MRAM [ZBCM09, AVR22, LKZ^{+}23, LWZ^{+}24, MPT^{+}22, YJD^{+}17]. MRU [GLYY14]. MSP430X [Seo18, SAKH20]. MTSS [MSB08]. Multi [ALZR19, ABH^{+}18, BGK^{+}23, CH10, CLJ^{+}19, CAA^{+}24, DJJ^{+}19, DP19, GAD^{+}24, GMS17, GDD20, GIB^{+}12, HSMS16, HH23, HWC^{+}20, JSZ^{+}19, JDW24, KR18, KRS^{+}16, LMS^{+}19, LJ24, LLW^{+}17, LOF20, MKM^{+}23a, MMA^{+}23, PMS^{+}24, PVV24, PWL^{+}19, PGR16, PHG^{+}17, QRW^{+}24, RC17, Rat25, RB21, RDP17, SSK^{+}22, SHL^{+}17, SRR^{+}23, SLFC19, TWTH18, TGBT17, TGTT17, VKDG19, VCM19, VF17, WHN^{+}17, XSP22, YZZG23, YLTY21, ZEJ^{+}23, ZPZG17, yCBR05, ISE10, Mus10, NPP13, PMM^{+}13, PÖG^{+}13, ZHCY13]. Multi-Access [RB21]. Multi-Bank [TGBT17]. Multi-bit [YZZG23]. Multi-buffer [CH10]. Multi-Byte [JDW24]. Multi-Compression [QRW^{+}24]. Multi-Core [CLJ^{+}19, HSMS16, HH23, KR18, PGR16, RC17, SRR^{+}23, TGTT17, VKDG19, WHN^{+}17, DP19, LOF20, PHG^{+}17, RDP17, XSP22, Mus10, PMM^{+}13]. Multi-Cores [KRS^{+}16]. multi-CPU [ISE10]. Multi-criteria [CAA^{+}24]. Multi-Device [ALZR19]. Multi-Disease [LJ24]. Multi-DNN [PVV24]. Multi-FPGA [JSZ^{+}19, LMS^{+}19, ZHCY13]. multi-frames [NPP13]. Multi-Gateway [GAD^{+}24]. Multi-Hop [GDD20]. Multi-Level [TWTH18, ZEJ^{+}23]. Multi-Mode [Rat25, YLTY21, yCBR05]. Multi-Objective [BGK^{+}23, DJJ^{+}19, PWL^{+}19]. multi-processor [PÖG^{+}13]. Multi-Processors [GIB^{+}12]. Multi-Quadcopter [SHL^{+}17]. Multi-Rate [ZPZG17]. Multi-Representative [LLW^{+}17]. Multi-Robot [GMS17, SLFC19]. Multi-Scale [ABH^{+}18]. Multi-task [HWC^{+}20, LOF20]. Multi-tenant [MKM^{+}23a, MMA^{+}23]. Multi-Threaded [VCM19]. Multi-Traffic [PMS^{+}24]. Multi-user [SSK^{+}22]. Multi-valued [VF17]. multiapplication [HT06]. Multicast [JCW^{+}16]. Multichannel [CLL16, GAG15, HC16, ZO16, CCH13, CW14]. multichoice [YCNCC11]. multicluster [PEP05]. Multicopy [CHK^{+}14a]. Multicore [AbSZ^{+}19, BZ13, CPC17, CQB^{+}15, ETAV16, EVS^{+}17, HGW^{+}20, HDG^{+}14, HPP17, LLZ^{+}17, MS21, MSD17, OSA^{+}18, PCM^{+}15, PCGD21, PLM^{+}15, RG14, RLP^{+}21, SMR15, SP19a, THA^{+}12, TFL16, UBF^{+}16, WZ12, YKKD23, BP14, BS13a, CCR^{+}14, HG09, HZX^{+}14, HL14, TKG13]. Multicore-Based [EVS^{+}17, BZ13]. Multicores [CSM24, LBS15, PM19, RKK15, SP20, JP14, MPZS13]. multidimension [YCNCC11]. Multidimensional [APRC16, JBN^{+}13]. multifrequency [ZWY^{+}10]. MultiLayered [LNA^{+}15]. Multilevel [CR14, LN04, ZLLC15]. Multimedia [CAPL11, Kwo16, YC12, CLK13, CCAP12, HQB07, IHK04, KBDV08, KBCL13, PK13, PCBW13, QP03, RC08, SRY13, UAK^{+}03, WMZY13]. multimedia-enabled [RC08]. Multimodal [RKM24, TLL^{+}12, AF14]. Multimode [AFMT17, JEP16]. MultiNets [NNH^{+}14]. Multiple [HLLL20, LZJ^{+}20, LBM^{+}24, MAKO19, NEP23, PXY^{+}17, TBDdD11, WSHC14, HQB06, ISG03, LCS03, MMSN14, NRL13, NNH^{+}14, PL10]. multiple-FPGA [MMSN14]. Multiple-Issue [WSHC14]. multiple-QoS [PL10]. Multiple-Step-Ahead [NEP23].

Multiplication

[DLS^{+}24, HTB^{+}25, MWK^{+}24, SAKH20]. multiplicative [KHHH14]. Multiplier [NWA12]. Multipliers [AZS^{+}23, LPP^{+}21, YF19, RMH04a]. Multiprocessor [BGO17, CDH^{+}16, DBM^{+}15, LX16, MG15, WRKG16, ZQGZ22, CHK14b, DZR09, ESAS14, HQB06, Hüb13, ISTE08, JM06, KKO^{+}06, LWB13, LES14, LQN^{+}13, OP06, PS10, SE10, TSBY13, VSSS13]. Multiprocessors [AR14, DSB17, PSZ12a, PRK15, PYJL15, RN14, BGD14, Bar13a, BD14, HFG13, HXZ^{+}13]. Multirate [TFL16]. multisequence [ZH12b]. Multitask [CMMD24, CSST08, DP08, MSB08]. multitasked [ZP09]. Multitasking [NB04, PCGD21, RP10, TM07, WAD14]. Multithreaded [HYY^{+}15, KE15, SPDLK^{+}17, ZP11, LCH^{+}08, LP09b]. Multithreading [LRZ16, PJS15, DFC^{+}19]. Multiunit [LX12]. Multiversion [KCC^{+}16]. muscle [WGPH13]. Must [Shu18e]. Mutation [FKS^{+}19]. Mutation-driven [FKS^{+}19]. MVLevelDB [LZZK25]. MxU [PP19]. My [BVM19]. Myriad2 [LLP^{+}21].

n [GKS^{+}22]. NAND

[BDG^{+}15, GMCC18, JNI15, MSHS19,

MAW22, PCK^{+}08, PK13, WC16, WZD^{+}17]. NAND-Flash [MSHS19]. Native [WWN23]. NAVIDRO [SPGB24]. Near [BCS16, FPGS22, ITO^{+}24, LFHS18, LWZ^{+}24, SWT^{+}14, SSA^{+}24b, SFCW23, SLZ^{+}24]. Near-Data [SFCW23, SLZ^{+}24]. Near-field [ITO^{+}24]. Near-Optimal [LFHS18]. Near-Static [BCS16]. Necessary [ARS16]. Need [Shu18c, STH17, TTB23]. Negative [CLS16]. Nested [DKA^{+}19, WYL^{+}19, KMB07, NNS13, TKD07]. nested-loop [NNS13, TKD07]. Net [DJZ13, LZJ17, YKD^{+}24, MPFG19, RKM24, YLDM19]. Net-V2 [RKM24]. NetBench [MMS06]. Nets [ACR17, BSM^{+}21, BB13, BB15, CL13, DLRTB^{+}19, JFK15, NDZ13, WZH13, ZW13, VAHC^{+}06]. Network [ASS^{+}23, ANARR^{+}19, ABF^{+}21, AD24b, BS22, CPC17, CWX^{+}23, CLW^{+}20, DLPK16, ESBK23, FJM^{+}25, FS24, HMR23, HSD22, HFL^{+}19, ICZ^{+}23, ICW^{+}21, JAB^{+}22, KSS24, KJK18, KFY^{+}22, LLG^{+}20, LMS^{+}22, MST^{+}16, NHS20, PGR16, SPT^{+}23, TLL^{+}12, VKDG19, WCK^{+}19, WDM^{+}23, WRB15, WZG^{+}23, XCZW23, YF19, ZRF^{+}12, ZZX^{+}15, ZSH^{+}24, ZBG20, ZLZ^{+}24, ZJZL20, ZDL22, ZP11, BP14, BFQ10, CP13a, CMS08, GMOB13, HVG13, KJRG13, KYL13, LLLT08, LLTL09, LHCK04, LCH^{+}08, LLLGR13, LS09, NNH^{+}14, PCM12, TKD07, WYP^{+}10, WYJ^{+}14, WW09, YCLV^{+}02, YZA13, ZWY^{+}10, MMS06, SSS11]. network-flow [WYJ^{+}14]. Network-Level [ZRF^{+}12, ZBG20]. Network-on-Chip [ABF^{+}21, BS22, DLPK16, HMR23, LLG^{+}20, MST^{+}16, VKDG19, BP14, GMOB13, YZA13, SSS11]. Network-on-Chip-Based [CPC17]. Network-on-Interposer [SPT^{+}23]. Networked [DLH16, WLC^{+}18, BWS14, BFQ10, FC13, Gup04, KKH^{+}12, NKP^{+}12]. Networking [JJS^{+}25, LYC^{+}18, WGN23, ZSEP21, DGC^{+}20]. Networks [AP20, AABG22, ANARR^{+}19, ALV^{+}22, ARZ^{+}23, AZHC19, ABC^{+}17, BKMG12, BMS24, BSM^{+}21, BBMG24, BTL^{+}12, CWZ23, DBX^{+}22, DS11, FPGS22, FBM16, FC16, GAD^{+}24, GVS^{+}20, GM12, GOC^{+}22, GDD20, GMVV17, GGJ12, HSR18, HZYJ22, HSK18, HY22, HZGW18, IB23, JR20, JBDD20, JGX^{+}18, JSD23, KBD24, LMB^{+}22, LFHS18, LAZ^{+}16, LJVD23, LPE^{+}23, MYL^{+}22, MSM21, MPFG19, MLAD23, MAGR15, OHCK24, PBC22, PYL^{+}23, QFL^{+}24, RN18, RKM24, RLG20, SJK20, SA21, SSK^{+}22, SCB^{+}22, SHK^{+}19, TSW^{+}17, XLY18, YLDM19, ZBG20, ZLL^{+}11, AKB14, CTK^{+}13, DLN13, DLC^{+}14, FZK^{+}10, GHZH14, HBSA04, HHB^{+}05, KHZS07, KAK05, KXL10, KLC^{+}10, KYHY14, KDN^{+}07, LN04, LAHS06, MLV09, NNS13, PS04, PS08a, PS08b, SRM^{+}13, SKH^{+}12, SGDP12, WYJ^{+}14, XWHC06, YGHS08, ZL08, ZLF13, ZC04c]. Networks-of-Systems [ZBG20]. Networks-on-Chip [FPGS22, FC16, IB23, AKB14, KYHY14, SRM^{+}13, WYJ^{+}14, XWHC06]. Neumann [SB23]. Neural [ASS^{+}23, AP20, AABG22, ALV^{+}22, ARZ^{+}23, AD24b, BMS24, BMP23, BBMG24, CWZ23, CWX^{+}23, CLW^{+}20, DBX^{+}22, FS24, GVS^{+}20, GOC^{+}22, HZYJ22, HSD22, HY22, HFL^{+}19, ICZ^{+}23, ICW^{+}21, JR20, JAB^{+}22, JSD23, KSS24, KBD24, LMB^{+}22, LJVD23, MSS23, MYL^{+}22, MPFG19, MLAD23, NHS20, OBO^{+}23, PBC22, PYL^{+}23, QFL^{+}24, RKM24, SJK20, SA21, SCB^{+}22, SHK^{+}19, TLSJ23, WCK^{+}19, WDM^{+}23, XCZW23, YF19, YLDM19, ZSH^{+}24, ZQD^{+}23, ZDL22, LLL14, TTB23]. Neural-Network [HFL^{+}19, WCK^{+}19]. Neuromodulation [PQA^{+}19]. Neuromorphic [CEC23, LMB^{+}22, SCB^{+}22, SBDK22]. Neuron [CCP^{+}19]. Neurosymbolic [SSA^{+}24a]. NeuroTAP [PP24]. Neutral [WDJ^{+}18, BFW^{+}19]. Next [ASWZ24, CMP23, KOL^{+}22, ISTE08, ISE10]. Next-generation [CMP23, KOL^{+}22, ISE10]. NIR [MAS^{+}24]. NIR-sighted [MAS^{+}24]. NIST [SSA21, ZSH^{+}19]. no [KHHH14, BVGVEA10]. No-Heap [BVGVEA10]. NoC [BLG^{+}15, BGD14, CCY^{+}13, CLLC17, DNNP14, DJ23, DJJ^{+}19, GLT^{+}13, LLR14, MSCS16, MKD15, MASG15, NASM18, NODP24, OMA^{+}13, PB14, PCGD21, TKHZ22, TMXS17, TAMS18, ZCK13]. NoC-based [CLLC17, MKD15, TAMS18, BGD14, LLR14, OMA^{+}13, PCGD21]. NoC-Based0 [MASG15]. NoCs [MAKO19]. Node [McI13, PVSG22, SKN17, ZH12a, ZH12b]. Nodes [GSS^{+}18, SLS^{+}19, ZO16, SGDP12]. noise [SBLM13]. Noisy [ICZ^{+}23]. NOMA [CZH^{+}24]. NOMA-Enabled [CZH^{+}24]. Non [BHM17, BMP23, FF25, FSVG19, HCS^{+}22, HKL^{+}23, KFY^{+}22, LZJ^{+}20, TTB23, WLWS15, XSP22, YHL23, ZZA^{+}22, HXZ^{+}13]. Non-coherent [YHL23]. Non-ideal [BMP23]. Non-ideality [HKL^{+}23]. Non-Inclusive [FF25]. Non-interference [BHM17]. Non-invasive [FSVG19]. Non-iterative [KFY^{+}22]. Non-neural [TTB23]. Non-preemptive [XSP22]. Non-Volatile [HCS^{+}22, ZZA^{+}22, LZJ^{+}20, WLWS15, HXZ^{+}13]. Nonblocking [SP10]. noncontact [CNC13]. Nonconverging [BTD^{+}18]. Noninclusive [CR14]. nonintrusive [NSL11]. Nonlinear [CMS17, LLL14]. nonparametric [GKW08]. nonrenewable [MKD13]. Nonutilization [LA11]. Nonvolatile [LKZ^{+}23, MLL^{+}17, PXY^{+}17, RKC^{+}22, SLS^{+}19, HZX^{+}14]. NOR [PRM21, SWJ^{+}13]. normal [RMH04a]. Novel [AAM^{+}17, CLS16, EVS^{+}17, HB23, HB24, MKOA25, MCS^{+}15, SP20, DZR09, NPP13, ZCK13]. NQA [WYL^{+}19]. NUCA [FS14]. NUCA-based [FS14]. Nucleus [VSD^{+}17]. Number [Ano13, Ano14, LCLW17, MFG16, MASG15, SSA21, SISS24, SRK^{+}18, Edi14]. numbers [ZXCH13]. Numerical [AGG^{+}17, ADJM19]. NVM [NBH23, SBDK22, WCK^{+}19]. NVM-Based [SBDK22, NBH23, WCK^{+}19]. NVMe [ZAL22]. NWSLite [GKW08].

O [CWH^{+}16, CCB^{+}06, EAAS22, JAD19, LSL^{+}23, MRY^{+}10, SKPL10, SC05, WGN23]. OA [MM16]. Object [GPH24, GMCC18, SRSM21, KTT13, MMSN14, NPP13]. Object-Based [GMCC18]. Object-oriented [SRSM21]. Objective [BGK^{+}23, DJJ^{+}19, PWL^{+}19]. Objects [BVGVEA10]. Oblivious [ZSH^{+}24]. ObNoCs [HMR23]. Observations [KH23]. Ocarina [GGGK08]. Off [KB23, LKZ^{+}23, ZRF^{+}12, CP13a, CRCR13, CLK13, CGV10, HFG13, OSA^{+}18, SD08, SJC^{+}03, SAYN09]. off-board [CGV10]. off-chip [CP13a, SJC^{+}03, SAYN09]. OffloaD [GPH24]. Offloading [CZH^{+}24, GPH24, HLLL20, Li21, OHCK24]. Offs [IPEP12, MCM^{+}17, GFC^{+}10, LDV12, SM13b]. Offset [OOAL06, SR12a, SEB12, HABT11]. Oh [Shu15b]. OLED [LKH16]. omega [MRT13]. omega-regular [MRT13]. omnidirectional [SCF12]. On-Accelerator [VKW^{+}17]. On-Board [CPP^{+}17]. on-body [QRB10]. On-Chip [FLF17, PVSG22, PSZ12a, SGZS21, CZHK23, KGR12, YFPJ14, ZRZ^{+}19, CP13a, LJ14, PL10, SJRS^{+}13b, WMZY13]. On-demand [ANARR^{+}19]. On-Device [KPS23, POCR24, RALR24]. On-the-Fly [PM19, UM13]. On/Off [LKZ^{+}23]. Onboard [FGL^{+}19, BCG10]. One [FLF^{+}23, WZH13]. One-shot [FLF^{+}23]. One-Step [WZH13]. Online [ASD^{+}24, Ano13, Ano14, EVS^{+}17, ISG03, KR14, REPL15, RALR24, SFB23, SE23, SSK^{+}22, VWG^{+}17, WXY^{+}17, YDLC10b, MSL13, TTAG14, YDLC10a]. Onloading [ALV^{+}22]. Only [GW15, BS13a, GDC19]. OnNetwork [KJK18]. OnSRAM [PVSG22]. onto [CC13a, DSXS15, OFA^{+}15]. Open [CWH^{+}23, LZS20, ZLSQ17, CCA^{+}13]. Open-Channel [ZLSQ17]. open-source [CCA^{+}13]. OpenCL [SPB^{+}17, SXMX^{+}18]. OpenCL-based [SXMX^{+}18]. operand [LCS03]. Operating [RKC^{+}22, WDJ^{+}18, AMCM06, BMM13, FRRJ07, TRJ05, WP11, YDLC10a]. Operation [BHD15, WC16, FC13]. Operational [AKB^{+}25, SGJ17]. Operations [GSC19, VF17, BAR13b, SWWY13]. Operator [CWZ23, GK22]. Operators [PRSV19, SUK23, USA^{+}22]. Opinions [Akd21]. OPLE [KAKSP15]. OPPC [LZS^{+}18]. Opportunistic [JCW^{+}16]. Opportunities [FSG23, Shu17b]. OPS [ZCS^{+}05]. Optical [FC16, LLG^{+}20, TKHZ22, VVKG23, KYHY14]. Optimal [ABD^{+}19, AZS^{+}23, BBD23, GTH^{+}22, GAG15, GPB^{+}17, KK05b, LFHS18, LZS^{+}18, MBP14, ZWK23, ABS02, CHK14b, GJ13, GNS04, PL10, RV07, SWT^{+}14, SC05, TAP23, YK03]. optimisation [YCK^{+}18]. Optimization [AHM19, CWH^{+}16, CCP^{+}19, CYH20, CAA^{+}24, DHJ^{+}17, DJJ^{+}19, DVC21, DASS12, DHL17, FBM16, GIB^{+}12, HJ25, HZGW18, IPEP12, JBDD20, JKH22, LKA^{+}18, LYH^{+}15, LX16, LSL20, MWS15, MFG17, PYJL15, PMS^{+}24, PWL^{+}19, PLM^{+}15, PMDC17, SR12a, SEB12, SP12, SBDK22, SR19, TSP15, TJ23, WH17, YDS^{+}22, ZYL^{+}17, ZPZG17, ZZZ^{+}12, BWS14, BMP03, CHK14b, CWX^{+}23, DVC^{+}07, DP08, HZX^{+}14, IBMK10, JMO14, KKC^{+}05, LXL13, LLLGR13, RP03, SAHE04, SKK^{+}14, YGHS08]. Optimizations [BSA17, BDG^{+}15, DJO12, HY22, HYY^{+}15, KKK^{+}11, MS23]. Optimize [FLF17]. Optimized [ARH^{+}18, AYS15, AV20, BRA^{+}16, BGGT23, MBR15, ZDL22, JHPR13, ZXCH13]. Optimizer [FS24, SBB19]. Optimizing [BP05, BCG10, KRHC20, KLK24, MDS^{+}21, MTWE20, PDHC23, RMBS20, SHQX19, XHL^{+}24, FRRJ07, HMM04]. Optimum [SPGT19]. Optimus [CWZ23]. Optode [FSVG19]. Orchestration [BCDD24, SSK^{+}22, SCA^{+}24]. Order [ACR17, BHM17, JLSP18, JBI17, KH23, LWCL25, LLC^{+}13]. Organized [TMXS17]. Oriented [BKMG12, SFZX18, CWKH12, DRL^{+}10, KK05a, LLN09, SRSM21, SGDP12]. Oscillators [SCM20]. Out-of-distribution [ZCG^{+}22]. Out-of-Order [JLSP18, JBI17, KH23, LLC^{+}13]. Output [KPK^{+}19, SFB23]. Output-based [KPK^{+}19]. Outputs [DPNA16]. Outsourcing [LZZ^{+}23]. Over-the-Air [WLH^{+}18]. Overbooking [DWRR14]. Overcoming [TP20]. Overflow [PRL^{+}23, BCS^{+}06, RRW05]. Overflow-free [PRL^{+}23]. Overhead [KSA^{+}18]. Overlap [GAD^{+}24]. overlapping [CTK^{+}13]. Overlay [CHS15, DFC^{+}19]. Overload [LDRM12]. overview [SVP05, WEE^{+}08]. Oximetry [FSVG19].

P [CAA^{+}24, KNY^{+}17, WDM17, WPW^{+}04, ZSH^{+}19]. P-256 [ZSH^{+}19]. P-Alloc [WDM17]. P-BMS [KNY^{+}17]. P-EDF [CAA^{+}24]. P1363 [HHL^{+}23b]. PA [JGX^{+}18]. Pacemaker [BJM13]. Packed [RLG20]. Packet [JGX^{+}18, SVS21, CMS08, LCH^{+}08, LS09, Mus10, RGSS04]. packet-switched [LS09]. packing [PEP05]. pad [ABS02, NDB09, UDB06]. Page [CLL16, CSC23, HC16, Rru22, VKW^{+}17, WLWS15]. Page-Mapping [CLL16, HC16]. Paging [KKK^{+}11]. pair [RV07]. Pairwise [DL12]. PAIS [JP14]. PALP [SDMK19]. PANDORA [SC20]. Papers [TEC12, SN10]. Parallel [CS16, CD19, DSXS15, GLP^{+}11, Goe14, LKA^{+}18, LZJ17, LYY^{+}17, LFC17, NFL^{+}22, PRB15, PJWY12, PÖG^{+}13, RDP17, SWL^{+}14, SM13a, TWTH18, WMLA16, GNR^{+}10, MMSN14, THON12, WW09]. Parallelism [AMN^{+}14, HLF^{+}18, JP14, LPD^{+}20, SDMK19, SM13b, ZEJ^{+}23, CW14, KVN^{+}09, MB10, SD13]. Parallelism-aware [JP14]. Parallelization [CGM^{+}25, HLLL12, LL15, TFL16]. Parallelized [KNY^{+}17]. Parallelizing [BRA^{+}16, MKR13, SC20, TTB23, UBF^{+}16, ZP06, MSH^{+}14]. Parameter [CAP15, GG24, LYY^{+}17]. Parameterizable [BRL16]. Parameterization [BSM^{+}21].

Parameterized [DVC21, Ise17, CMA05]. Parameters [DJO12, HKC18, MPT^{+}22]. Parametric [ACR17, MMR^{+}10, VGB19]. Parametrised [SH15]. Pareto [CAP15, GPB^{+}17, TAP23]. Pareto-Optimal [GPB^{+}17, TAP23]. parMERASA [UBF^{+}16]. Parsimonious [LEPP13]. Part [CBH22b, STLX22a, STLX22b, SCZ20a, SCZ20b]. Partial [ACR17, BHM17, DLV16, KLK24, SBK^{+}23, HPLD09, LPFG13]. Partial-Order [ACR17]. Partially [PRM21, TJ23, CIC^{+}08, CIC^{+}09, RI04]. participatory [WTSR13]. Partition [LYY^{+}17, SDMK19]. Partition-Level [SDMK19]. Partitioned [AK21, BDP24, DBM^{+}15, GWZ16, HGW^{+}20, KVK^{+}03, TGTT17, ZQGZ22, CJMB05, LXL13, SBX08]. Partitioning [AbSZ^{+}19, Bar13a, CWH^{+}23, CI17, GRRR24, GTH^{+}22, HSMS16, KAKSP15, SMR15, SPB^{+}17, VGN18, WHN^{+}17, XSP22, KP13, LXL13, RP10, SVN04, TJ10, XHSS10]. Partitions [LC17, SJRS^{+}13a]. PArtNNer [GRRR24]. party [RBNM19]. Pass [KKK^{+}11]. Passing [LZJ17, LBP07]. Passive [BSB14, SFB23]. passivity [KKH^{+}12]. Path [FLF17, LZS^{+}18, GNW05, MSH^{+}14]. Paths [NS17]. Pattern [GDDD17, GDD20, PP24, UFA^{+}24, GDN03, KPK^{+}19]. Pattern-Aware [PP24]. pattern-based [GDN03]. patterns [GLC07, PÖG^{+}13]. Pay [RBNM19]. Pay-per-device [RBNM19]. Payment [SYC^{+}17]. Payments [SBR^{+}15]. PCA [KCCW17]. PCA/DCA [KCCW17]. PCM [KCC^{+}16, NBE18, SJOL22, YCT16, YYKK18]. PCM-Based [KCC^{+}16, SJOL22, YCT16, YYKK18]. PCMOS [SBLM13]. PDA [GW08]. PDE [AKI^{+}23]. Pedestrian [TM15]. Perceivable [CSC23]. Perceived [KJKM16]. Perfect [SLE^{+}17]. Perfectly [DKA^{+}19]. perform [AAPN14].

Performance

[AK21, AB15, AVR22, BFW^{+}19, BRA^{+}16, BDG^{+}15, BONA22, CZ23, DCZB19, DJO12, DLPK16, FMHS23, Fra12, GSC19, GMCC18, HHL^{+}23b, IPEP12, JLSP18, KJLS20, KCWH14, LKA^{+}18, LLP^{+}21, LCC^{+}19, LYH^{+}15, LPO^{+}17, MAKO19, MLR^{+}17, MGP25, NASM18, NZCS19, NBM^{+}16, OBSO16, OBA^{+}17, PP24, PCM^{+}15, PGR16, PP12, PJT^{+}23, RG14, RKK15, SRG^{+}15, SPT^{+}23, SP12, WZM17, WT15, WZD^{+}17, WKP24, YJD^{+}17, ZRF^{+}12, ZLSQ17, ZSJ12, BCLN13, BP05, BZ13, CMP^{+}07, HLD^{+}09, HHL^{+}23a, HHB^{+}12, JHK^{+}06, KS13, KD08, LWB13, LCH^{+}08, MSL13, PGS^{+}13, SE10, SZG^{+}23, SM13b, YDLC10a, ZAL22]. Performance-Maximisation [DCZB19]. Performance-Optimized [BRA^{+}16]. Performance-Power-Programmability [LLP^{+}21]. Performance/Error [MLR^{+}17]. Performance/Power [RKK15]. performances [FS14]. period [LK10]. Periodic [ARS16, KB23, LZL15, PSD21, SD17, HCQ^{+}14, SL08, XQ07]. Periodically [WMLM12]. Permissive [WZH13]. Permutations [ARH^{+}18]. Permuted [RLG20]. Perpetuu [MBB^{+}15]. Persistence [YCJW24, Cul13]. Persistent [HTR^{+}16, SXH^{+}19, WCB20, WCH^{+}23, YCJW24]. Persistently [ZZA^{+}22]. personal [RC08]. Personalized [HSD22, SKK^{+}14, YS23]. Perspective [DJS16, KJK^{+}17b, SG24, SUS^{+}17, BJM13, WBF^{+}06]. pervasive [CD10, TSWL10]. Petri [ACR17, BSM^{+}21, BB13, BB15, CL13, DLRTB^{+}19, DJZ13, JFK15, LZJ17, NDZ13, WZH13, YKD^{+}24, ZW13]. Phase [GW15, LH18, MSD17, SDMK19, ZLLC15]. Phase-Only [GW15]. Phased [TB23]. PHAT [YV23]. PhiNets [PAF22]. PHMMRGB [SNTÖ25]. phone [LLL14]. Photonic [AKI^{+}23, PGR16, XCZW23, BP14]. Photonics [ASS^{+}23]. Photovoltaic [BCS16]. Physical [AFS^{+}13, ALZR19, BHAC15, BKMG12, CKGN14, DWRR14, DHJ^{+}17, DHF18, GCJD20, GSN21, HZX15, IPL16, KCC^{+}16, LAB^{+}23, LWZ^{+}16, LLN^{+}14, MBKF15, MKS^{+}17, NRL13, NLSV^{+}19, PRS^{+}17, SHL^{+}17, Shu19d, TGV12, TLL^{+}12, TCD^{+}19, UGS^{+}21, WDY^{+}16, WZBP19, XKK17, ZYM16, ZYL^{+}17, ZJC^{+}17, BWS14, BJM13, DDG^{+}13, GMOB13, HVG13, Hüb13, LDRM12, SPK^{+}12, SMR20, TXL^{+}12, WLT12, YRS12, ZSM13]. Physically [MGC^{+}23]. Physics [PL13, QFL^{+}24, Rru22]. Physics-Driven [Rru22]. PICA [LS12]. Piecewise [SGJ17]. Piecewise-Smooth [SGJ17]. PIM [GWC^{+}24, NODP24]. Pin [SIC19]. Pin-Count [SIC19]. Pipeline [AÖÖ23, HZH^{+}18, MD04]. Pipelined [HH23, TBDdD11, BAR13b, CAP^{+}07, HG09, LLC^{+}13, THON12, ZXS03]. PISCOT [HH23]. Placement [CKB17, DSXS15, MSCS16, NS16, PqBM^{+}15, TP16, WJ17, BSKB^{+}09, JGD^{+}09, SBX08]. Places [WWY13]. Plan [SRB23]. Planes [AR14]. Planetary [LMS^{+}19]. Planning [BJH^{+}24, JFK15, LZS^{+}18]. Platform [CPC17, GLMP18, GRRR24, LMA19, MFMA17, PGR^{+}08, SSA^{+}24a, SLB^{+}15, SXXS^{+}16b, TGTT17, WWN23, YGD^{+}19, ARJ08, AKB14, ISTE08, VHB^{+}13, ZM07]. Platform-Agnostic [GRRR24]. Platform-aware [SSA^{+}24a]. Platform-Based [SLB^{+}15, PGR^{+}08]. Platform-Specific [LMA19, MFMA17]. Platforms [CAA^{+}24, CMP17, DDC^{+}24, DP19, ETAV16, HGW^{+}20, IDO^{+}22, JBDD20, LSC19, LKA^{+}18, LWB18, MPFG19, MBLA16, PRB15, PCGD21, SL16, SOG15, SJOL22, TFL16, VFS^{+}21, WHN^{+}17, ZQGZ22, ACK^{+}13, CCR^{+}14, JMO14, LLLGR13, MLV09, OP06, SM13b, TSBY13]. PLTL [CJMB05]. PLTL-partitioned [CJMB05]. Plug&Chip [DSXS^{+}14].

Plugging [CSW15]. PMC [HPP17]. Point [AABG22, LJVD23, Rat25, AC08, BMS24, DBH14, LPP^{+}21, MLR^{+}17, SCM20]. Pointer [SFZX18]. pointers [BAR13c]. Points [ZLL^{+}19]. Poisoning [CK23]. Polar [LZJ^{+}19]. Policies [GD24, KR14, KBDV08, LA11, RG13]. policing [DW10]. Policy [HPS13, PYJL15, Mus10]. Pollution [DJ23, SHL^{+}17]. Poly [PZ12]. Poly-DWT [PZ12]. PolyARBerNN [FS24]. Polyhedral [LP19, NNS13]. polymorphic [AVF^{+}09, PZ12]. Polynomial [DLS^{+}24, FS24, HTB^{+}25, MWK^{+}24, TJ10]. Poor [ZLZ^{+}24]. Portability [CHS15]. Portable [MGLP19, YC12, ABI^{+}09, ELS08]. position [QRB10]. Positioning [HXH^{+}24]. Post [AKB14, AKI^{+}23, AAT^{+}21, BDR24, HTB^{+}25, KSFS24, KKK^{+}11, MKAA17, MWK^{+}24, NVB^{+}20, RCDB24]. Post-Moore [AKI^{+}23]. Post-Pass [KKK^{+}11]. Post-Quantum [AAT^{+}21, BDR24, HTB^{+}25, KSFS24, MKAA17, MWK^{+}24, NVB^{+}20, RCDB24]. Post-silicon [AKB14]. postfabrication [CIC^{+}08, CIC^{+}09]. postural [QRB10]. Power [ABD^{+}19, ABF^{+}21, ARP12, ACK^{+}13, AZHC19, AVR22, BFW^{+}19, BDB^{+}17, Bec09, BGO17, FHK21, HJ25, HRT^{+}22, HZGW18, IYL^{+}23, JRR16, JC12, JEP16, JDW24, JAB^{+}22, KHZS07, LKA^{+}18, LPP^{+}21, LLP^{+}21, LZL15, LYH^{+}15, LKH16, LBP07, MLL^{+}17, MV16, MSR^{+}12, MCM^{+}17, NBE18, OBSO16, OMA^{+}13, PYJL15, Pau14, RKK15, RSW21, SLB^{+}15, SRS03, SOL^{+}16, SSPP23, SKN17, SBK^{+}23, SLK^{+}22, SR19, SLS^{+}19, TKV^{+}18, WHL23, XLY18, YGW^{+}12, YC12, ZRF^{+}12, ZRZ^{+}19, ZP07, AMCM06, BO13, CMP^{+}07, DBH14, Geb04, GJ13, GRCV03, GLWM14, IHK04, ISG03, JC03, JHK^{+}06, KR14, KXL10, KYHY14, MSS^{+}03, MALM04, MSL13, Mus03, NPP13, ÖNG08, PAF22, PKIT23, QRB10, RKM24, RAK14, SWL07, SJRS^{+}13b, SCM20, TTB23, TVK08, VJD^{+}07, VDK^{+}08, ZC04b, ZTRC03]. power-attacks [Geb04]. Power-Aware [JEP16, ACK^{+}13, AZHC19, OMA^{+}13, JC03, MSS^{+}03, MALM04]. Power-Efficient [HRT^{+}22, SLB^{+}15, ABF^{+}21, ZP07]. Power-mode-aware [SR19]. Power-neutral [BFW^{+}19]. Power-Performance [ZRF^{+}12]. power-saving [ISG03]. power-sensitive [BO13]. Power-Temperature [BGO17]. Powered [HLL^{+}23, JRSR17, TSW^{+}17, XLY18, ANB^{+}20, MBB^{+}15, RV03, YTL^{+}20]. Powerful [SGZS21]. Practical [BCLS17, BHL^{+}20, HPO^{+}15, LC17, PWL^{+}19, RIMS21]. Practice [FSB^{+}21, BSKB^{+}09, Cul13]. Practitioners [Akd21]. PRAM [LO13, PMPP14]. Pre [CIC^{+}09]. Pre- [CIC^{+}09]. Preaveraging [GWM16]. Precedence [SE17, MBFSV07]. Precise [MGB^{+}21, NS16, ZLL^{+}18]. Precision [SSD^{+}19, SE07, ZDL22]. Precomputation [HKC18]. Preconditioning [LLW^{+}24]. PredATW [DS24]. Predicate [ADI06]. Predictability [TSBY13, GLYY14]. Predictable [BCS^{+}23, FSB^{+}21, GHPP18, KR18, KWKP23, PP19, PW13, SSK23, SRG^{+}15, TBG^{+}17, VKMP20, WWG^{+}18, WKP24, AEF^{+}14, WAD14]. Predicting [DS24, DJO12, EHF24, JC12]. Prediction [GKSB25, KCJ^{+}16, MGP25, NS17, NEP23, QZXO14, SKS21, TKHZ22, GKW08, HE12]. Prediction-Directed [QZXO14]. Predictive [RN18, SSD^{+}19, TBCB15]. PredictNcool [SP19b]. Predictor [SP19b, WGPH13, ZA07]. Preemption [CR14, DBM^{+}15, FF25, GWZ16, TB23, ZGZ15, ZLL^{+}19, ZP09]. preemptions [RM10]. Preemptive [DSB17, TM07, WAD14, XSP22]. Prefabrication [CIC^{+}08]. Preface [AL05]. PReFeR [MGC^{+}23]. prefetching [YZ08, ZP07]. Preorders [BSV17].

Preparation [BCHB18]. Presence

[TBDdD11, LHX^{+}14, VS08]. PRESENT [WH17]. Preservation [HSR18]. Preserving [ACR17, KLK^{+}19, LLT^{+}17, CSST08]. Pretrained [JBDD20]. Prevention [ZW13]. pricing [WSK14]. Primary [Shu18e]. Primitive [MCS^{+}15]. Primitives [BSJ15, LBP07]. Principled [PHG^{+}17]. Prioritizing [SPGT19]. Priority [DBM^{+}15, DHL17, GE18, LH18, MBP14, MAKO19, SD17, WHN^{+}17, DF14, LA11, MEP08, QH07, YK03, ZZZ^{+}12]. Privacy [KLK^{+}19, KCCW17, LLT^{+}17]. Proactive [SWL^{+}23]. PROARTIS [CQV^{+}13]. Probabilistic [AFS^{+}13, CLL21, COC22, GUC^{+}23, HQB07, HCL^{+}17, KM13, LP19, LEPP13, MHT13, SWJ^{+}13, SCG15, SWS23, TBEP16, WHN^{+}17]. Probabilistically [CQV^{+}13]. Probability [MKM^{+}23b]. Problem [SEB12, WEE^{+}08, Ahm13]. Problems [KOM^{+}23, TJ10]. Procedural [GGMS24]. procedure [KMB07, KASD07]. Process [BGRV15, GM12, MZG14, MAG14, MASG15, WDM17, NNS13, TKD07]. Process-Variation [WDM17]. Process-variation-aware [MAG14]. Processes [LZJ17, PBP09a, PBP09b]. Processing [AÖÖ23, ASD^{+}24, BT22, BDB^{+}17, BCBJ24, DVC21, HRH^{+}22, HKL^{+}23, LVSVRFCG23, MKM^{+}23a, MGLP19, MKE18, SFCW23, SBDK22, SWWW17, SLZ^{+}24, VKMP20, WZY^{+}23, XZK^{+}19, AMN^{+}14, BCG^{+}07, BCG10, DSW^{+}09, GHB13, GJ13, HVG13, PÖG^{+}13, SCF12, VGG^{+}13, ZH12b, ZLF13, MSR^{+}12]. Processing-enabled [SLZ^{+}24]. Processing-in-memory [BCBJ24]. Processor [AKI^{+}23, BVM19, GOC^{+}22, KRR20, LWS^{+}23, MLL^{+}17, MBR15, MSD17, MMD04, PHG^{+}17, SK13, SOL^{+}16, SK19, SCS16, TWTH18, TKL^{+}15, WWHT21, ZZA^{+}22, CCA^{+}13, GLWM14, HL14, KGR12, KT14, LK10, LHCK04, LCH^{+}08, LV09, MG05, PMM^{+}13, PÖG^{+}13, ZC04a, LS12]. processor-based [KGR12, LHCK04]. Processor-memory [MMD04]. Processor-transparent [ZZA^{+}22]. processor/accelerator [CCA^{+}13]. Processors [AJ18, GIB^{+}12, HLLL12, HTC^{+}16, JLSP18, KKS^{+}23, PDL^{+}23, PCGD21, PJT^{+}23, RC17, SJLK18, SSA21, SCM20, SWX17, TTB23, TBDdD11, WZ12, YKKD23, YC16, ZP11, BS13a, BO13, BM13, CIC^{+}08, CIC^{+}09, CC13a, DPP14, Geb04, GGI13, HZX^{+}14, JHPR13, KD08, KK05b, LLPM07, LS13, LLLT08, LLTL09, Mus10, ÖNG08, PBV07, PO05, RP11, TLLL09, UAK^{+}03, WW09, YW13, ZMB03, ZP06, ZP07, LKB14, MMS06]. producer [RV07]. Product [MKS23, TR24]. Profile [OMH^{+}23, WKJ20, WLH16, BAR13c]. Profile-guided [OMH^{+}23]. Profiling [BP19, CGM^{+}25, FLF17, MGB^{+}21, MSRM24, MSL13, ZLL^{+}18, LLLGR13, NSL11, STY^{+}14]. Program [AAS18, BVM19, FHK21, KPK^{+}19, OSA^{+}18, RLP^{+}21, SRR^{+}23, WZD^{+}17, AFG08, MF13]. Programmability [LLP^{+}21, THA^{+}12]. Programmable [GOC^{+}22, LWS^{+}23, MAS^{+}24]. Programmatic [Bro21]. Programming [BHXP19, GWC^{+}24, WCK^{+}19, WWN23, ABI^{+}09, BWS14, BvB13, BMM13, Gar05, LP09b, LAHS06, PÖG^{+}13, SGDP12]. Programs [AGG^{+}17, CJ20, EYG^{+}23, GHR15, KH18, LL15, LLP^{+}17, LML20, MKR13, SPDLK^{+}17, TWTH18, WMRB17, WCM^{+}16, AFG08, BSB14, CSST08, CC13b, GNP06, KS13, NNS13, TKD07]. Progress [BHAC15, HLL^{+}23]. Promising [KOM^{+}23]. Proof [DAASP21, MS13b]. Proof-Based [MS13b]. Propagate [GWM16]. Propagation [HLLL12, RS07]. propagation-based [RS07]. Properties [BFST19, BBDR12, GZ12, CMA05]. Property [BS22, KM09]. Proportional [FPGS22]. Protecting [BS22, HMR23, KJK^{+}17b, LMW^{+}17]. Protection [JGCS24, RLL^{+}23, YC12, BCS^{+}06]. Protocol [AZHC19, CCM17, CBS19, FJM^{+}25, GDA13, KYDC20, LZJ^{+}19, MGC^{+}23, ZSY19, CHTC07, KASD07, PS04, YFPJ14]. Protocols [AAT^{+}21, EZL^{+}17]. prototype [GGGK08]. Prototypes [GD24]. Prototyping [CS16, DSXS^{+}14, Goe14, KPC^{+}16, SXMX^{+}18]. Provably [AR14]. Proven [PG24]. Providing [DLN13, KS18, LHX^{+}14]. provisioning [LDRM12]. Proximity [LNA^{+}15]. prune [DNNP14]. Pruned [RLG20]. Pruned-Permuted-Packed [RLG20]. Pruning [KFY^{+}22, PKL22, SC05]. Pruning-based [SC05]. Pseudorandom [MFG16]. Psi [BGRV15]. Psi-Calculi [BGRV15]. PTAT [ZLL^{+}18]. Public [Seo18, SAKH20, Shu16a]. Public-Key [Seo18]. PUF [CCKM16, CCM17, NBH23, RBNM19, SRK^{+}18]. PUF-Based [CCM17]. PUFs [LZZ^{+}19, ZTZ^{+}19]. Purchasing [CZH^{+}24]. purpose [CMGM24, GKW08]. PV [PJL^{+}17]. PV-cell [PJL^{+}17]. pWCET [RSF20]. Python [LHM14].

QC [VOG15]. QEMU [MZG14]. qLUT [RR17]. QoE [IDO^{+}22]. QoS [HLLL20, LDV12, PL10, RJM19, SSK23, ZLL^{+}11]. QoS-aware [SSK23]. QR [WL09]. QRD [SPC^{+}16]. Quadcopter [SHL^{+}17]. Quadratic [AGS^{+}16, AGG^{+}17]. Quadtrees [WCH^{+}23]. Quality [AKB^{+}25, BZG19, CLL^{+}18, CYH20, CRCR13, DDC^{+}24, LKH16, MST^{+}16, PDHC23, RDSS21, WKJ20]. Quality-level [RDSS21]. Quality-of-Service [MST^{+}16]. Quality-Retaining [LKH16]. Quality/Latency [BZG19]. Quality/Latency-Aware [BZG19]. Quantifying [CBRZ19]. Quantitative [SD08, SR12b]. Quantization [GG24, IVJ^{+}23, LJVD23, PKL22]. Quantization-aware [IVJ^{+}23]. Quantized [DBX^{+}22, PKL22, RR17, LM24]. Quantum [ASWZ24, AAT^{+}21, BDR24, HTB^{+}25, KSFS24, MKAA17, SWK19, MWK^{+}24, NVB^{+}20, RCDB24]. Quantum-Safe [ASWZ24]. QUAREM [IDO^{+}22]. quasi [FZHT13]. quasi-static [FZHT13]. Quasistatic [PLKH08]. Queries [LZZK25]. Query [BMMV21, MKM^{+}23a, WTSR13]. Query-based [BMMV21]. Queue [JBI17]. Queues [GHR15]. queuing [RGSS04]. QUIDAM [IVJ^{+}23].

Race [YHL23]. Racetrack [KRHC20, KOL^{+}22]. Racetracks [BCBJ24]. radar [BCG^{+}07]. Radiation [MGB^{+}21, ZHCY13]. Radio [HZGW18, LOD18, SRA12, XLY18, ZLL^{+}11, KXL10, LJR12]. Radio-Agnostic [ZLL^{+}11]. radiotherapy [ZXCH13]. RAID [GIA11, SWL^{+}23]. RAID5 [PX18]. RAIDs [BD14]. RAIDs-on-chip [BD14]. Rainbow [KKS^{+}23]. RAM [MMK22]. Random [SRK^{+}18, KJRG13]. Randomization [SE23]. Randomized [ARP12]. Range [HHB^{+}05, DNBL22]. Range-free [HHB^{+}05]. ranging [PSZ12b]. Rapid [DSXS^{+}14, HSR18, KPC^{+}16, LSC14, LP10, ZP09]. RapidIO [BCG^{+}07, BCG10]. Rapidly [STB^{+}24]. RapidRadio [SRA12]. Rare [HSR18]. Rasterization [OBA^{+}17]. Rate [AFMT17, ESM^{+}17, JJS^{+}25, SLS^{+}19, ZPZG17, BJM13, GNP06, SWT^{+}14]. Rates [WSMF22]. ratio [MEP08]. ray [ZXCH13]. rays [ZXCH13]. RCML [RHG^{+}12]. RDF [FGK^{+}23]. Re [LLW^{+}17, Shu20b, VWG^{+}17]. Re-evaluating [Shu20b]. Re-Fusion [LLW^{+}17]. Reach [KDR23]. Reachability [BF17, BB13, FKJM18, HFL^{+}19, JBCS16, MG15, ADI06]. Reachable [DB19, GD19]. ReachNN [HFL^{+}19]. Reaction [GUC^{+}23]. Reactive [JZL^{+}15, Mos13, BCC^{+}08, CJMB05, GNP06]. Read [HCS^{+}22, LLZ^{+}22, MMK22, YJD^{+}17, YCK^{+}18, YWLW23]. Read-Out [YCK^{+}18]. Read-Related [YWLW23]. Reads [PM19]. READY [DFC^{+}19]. Real [ARS16, AbSZ^{+}19, AYS15, BMAB16, BZG19, BFST19, BE17, BAG^{+}20, BGS^{+}18, CDBB24, CQV^{+}13, CSM24, CKL04, CKGN14, CWZ^{+}20, CSH^{+}22, CHJ22, CLS16, CQB^{+}15, CAA^{+}24, DLRTB^{+}19, DHL17, DJZ13, ESBK23, FJM^{+}25, FSB^{+}21, FBM16, GAD^{+}24, GG24, GAG15, GZZ^{+}16, GE18, HQE20, HGW^{+}20, HSMS16, HH23, HFA^{+}14, HHC^{+}16b, HE24, IB23, JSZ^{+}19, JAD19, JGX^{+}18, JBCS16, KSS16, KR18, KH23, KB17, LG21, LN19, dFMAdN12, LZL15, LX16, LOF20, MM16, MZG15, MCG22, MAW22, MSM21, NFL^{+}22, NRZ^{+}24, OSF19, PMS^{+}24, Pau14, PSD21, PJT^{+}23, PNRC17, RG14, RMK17, SSK23, SNTÖ25, SCG15, SMR15, SE10, SP19a, SP20, SLCS16, SCS16, SLE^{+}17, SCA^{+}24, SGW^{+}16, SD17, TSP15, TGP25, TKT15, UBF^{+}16, WDJ^{+}18, WMGR12, WHN^{+}17, WGN23, XSP22, XQQ^{+}24, ZDZ14, ZWK23, ZPZG17, ZJC^{+}17, ZSJ12, AMCM06, AF14]. real [AFL13, ABC^{+}07, ABI^{+}09, AFG08, BVGVEA10, BBL09, CMV10, CHK14b, CRJ10, CRM14, CHTC07, CCAP12, CRAJ10, DF14, DSW^{+}09, DW10, GNW05, HT06, HTLC10, HHB^{+}12, HCQ^{+}14, KBDV08, KW10, KTT13, LSK^{+}08, LES14, LQN^{+}13, LLR14, LHX^{+}14, MMSN14, MEP08, MRY^{+}10, MVS^{+}13, MALM04, MAG14, MKD13, DWCM14, NNH^{+}14, PMM^{+}13, PAP^{+}12, PL10, PS10, QH07, RMM03, SP10, SKPL10, SL08, SE07, SC05, TM07, WMT12, WP11, WAD14, YK03, ZC04a, ZC04b, ZB13, ZX08, Zhu10, ZZZ^{+}12, TTA^{+}20]. Real-Time [ARS16, AbSZ^{+}19, AYS15, BMAB16, BGS^{+}18, CDBB24, CQV^{+}13, CSM24, CKGN14, CWZ^{+}20, CSH^{+}22, CHJ22, CLS16, CQB^{+}15, DHL17, DJZ13, FJM^{+}25, FBM16, GAD^{+}24, GG24, GAG15, GZZ^{+}16, GE18, HGW^{+}20, HSMS16, HH23, HFA^{+}14, HHC^{+}16b, IB23, JSZ^{+}19, JAD19, JGX^{+}18, JBCS16, KSS16, KR18, KH23, KB17, LN19, dFMAdN12, LZL15, LX16, MM16, MZG15, MAW22, OSF19, PMS^{+}24, Pau14, PSD21, PJT^{+}23, PNRC17, RG14, RMK17, SNTÖ25, SCG15, SMR15, SP19a, SLCS16, SCS16, SLE^{+}17, SGW^{+}16, SD17, TSP15, TGP25, TKT15, UBF^{+}16, WDJ^{+}18, WMGR12, WHN^{+}17, WGN23, XQQ^{+}24, ZDZ14, ZPZG17, ZJC^{+}17, ZSJ12, BZG19, BFST19, BAG^{+}20, CKL04, CAA^{+}24, DLRTB^{+}19, ESBK23, HQE20, LG21, LOF20, MCG22, MSM21, NFL^{+}22, NRZ^{+}24, SSK23, SE10, SP20, SCA^{+}24, XSP22, ZWK23, AMCM06, AF14, AFL13, ABC^{+}07, ABI^{+}09]. real-time [AFG08, BVGVEA10, BBL09, CMV10, CHK14b, CRJ10, CRM14, CHTC07, CCAP12, CRAJ10, DF14, DSW^{+}09, DW10, GNW05, HT06, HTLC10, HHB^{+}12, HCQ^{+}14, KBDV08, KW10, KTT13, LSK^{+}08, LES14, LQN^{+}13, LLR14, LHX^{+}14, MMSN14, MEP08, MRY^{+}10, MVS^{+}13, MALM04, MAG14, MKD13, DWCM14, NNH^{+}14, PMM^{+}13, PAP^{+}12, PL10, PS10, QH07, RMM03, SP10, SKPL10, SL08, SE07, SC05, TM07, WMT12, WP11, WAD14, YK03, ZC04a, ZC04b, ZB13, ZX08, Zhu10, ZZZ^{+}12]. Realization [CSH^{+}22]. Realize [TB23]. Realizing [SJK20]. Reallocation [LLZ^{+}22]. Really [RPB^{+}19]. REC [ZQ24]. Receiver [LCL^{+}19, KSK13]. Receiver-Initiated [LCL^{+}19]. Rechargeable [LFHS18]. Reckoning [TM15]. Reclamation [CSC23, KJKM16]. Recoding [CD12]. Recognition [BTA^{+}19, BJCHA17, DBX^{+}22, GGJ12, HZW^{+}23, HB23, HB24, LX22, RRC22, SG24, TSO22, WZG^{+}23, XYLC23, ZRF^{+}12, KP13, NRL13, NPP13]. Recognizing [ALZR19]. Recommendation [WLZ^{+}23]. Reconfigurable [ARDG16, DSXS15, FGK^{+}23, GPT^{+}23, HHL^{+}23a, KMS^{+}23, LCD18, LZZ^{+}19, PJWY12, RHG^{+}12, SBB19, SFCW23, SP12, SVC^{+}23, SSS11, STY^{+}14, SRK^{+}18, WRKG16, YZZ^{+}23, AVF^{+}09, Bec09, CIC^{+}08, CIC^{+}09, CMS08, CRM14, GD14, HMMA04, LPFG13, LS09, LP09b, NBGS09, NB04, PBV07, PCK^{+}08, RI04, SB08, SGDP12, VNK^{+}03, VHB^{+}13, YMKH23]. Reconfiguration [AHM19, CWH^{+}23, DP19, FF09, KLK24, SA18, WMGR12, YKKD23, GNS04, HMM04, HKVI05, HPLD09, LJR12, LPFG13, PAS^{+}09, ZBCM09]. Reconfiguration-Based [SA18]. Reconfigurations [Kha13, KML13, ZSJ12, CRM14]. ReconOS [LP09b]. Reconstruction [HW17, SWL^{+}23]. Recording [CCSC23, LCC^{+}19, LCC^{+}23]. Records [LMW^{+}17]. Recoverable [WCB20]. Recovering [CRAJ10]. Recovery [BDB^{+}17, EZL^{+}17, HPS13, HB24, LCD18, LJLT17, SSK21, TMXS17, ZZA^{+}22, FO03]. Rectifying [CSC23]. Recurrent [ARZ^{+}23]. Recursive [SCM20]. REDEFINE [AVF^{+}09]. Redirection [MST^{+}16]. Reduce [WLTW24, CRM14, LOXL13, Mus03, YFPJ14]. Reduced [HLM^{+}24, RRC22]. Reducing [ASJ21, BB13, CW14, CKIR06, HWTL23, JHK^{+}06, LLC^{+}22, MV16, UCK^{+}09, ZKKC05, ZTD^{+}06, ZA07, ZLX^{+}23, CSK^{+}02]. Reduction [GDC19, LCLW17, LLW^{+}24, PYL^{+}23, SLN^{+}16, TBDdD11, WRW^{+}21, WHL23, YCK^{+}18, ZZX^{+}15, CDD^{+}07, HXZ^{+}13, LS13, PLKH08, ZXS03]. Redundancy [BB13, TTAG14, YZA13]. Redundant [AJ18, LPE^{+}23, NWA12, SSA21, MB10]. Redundant-Digit [AJ18]. references [HT06]. Refinement [DJZ13, DAASP21, KB17, LP19, MS13b, HDR^{+}06, RS07]. Refinement-based [DAASP21]. Refining [NS17]. Reflections [Shu19d]. Reflexes [SPP^{+}10]. Refresh [LLZ^{+}22]. Reg [MSRM24]. Reg-Tune [MSRM24]. Region [ZWH^{+}16]. Region-Wide [ZWH^{+}16]. regions [LA11]. Register [AP09, FND^{+}16, FLF^{+}23, LOXL13, SWX17, TBDdD11, YC16, CKIR06, HABT11, LS13]. Register-based [FLF^{+}23]. Register-to-Register [FND^{+}16]. Registers [NGL17, LOXL13]. Registration [SVC^{+}23]. RegKey [FLF^{+}23]. Regression [ASD^{+}24, MSRM24, RLP^{+}21, BMS13]. Regression-Focused [MSRM24]. Regular [CWH^{+}23, NCJF18, Shu15c, CMA05, MRT13]. Regularity [LC17]. Regularity-based [LC17]. Regulation [SSPP23, YFPJ14]. Reimagining [Mit21]. Reinforcement [BJ23, BGK^{+}23, CZHK23, KSY17, RB21, SKN17, TCD^{+}19, PCBW13]. Reinforcing [WXY^{+}17]. ReKeying [DS11]. Related [CR14, FF25, MGC^{+}23, Shu15c, YWLW23]. relation [VAHC^{+}06]. Relational [CMS17]. Relations [SE17]. Relative [LZZK25]. Relaying [WLHC18]. Reliability [BHD15, BDG^{+}15, DHB^{+}23, KRS^{+}16, LCY^{+}22, MB10, NASM18, PRK15, SRNW16, WDM17, WLC^{+}18, ZSEP21, Zhu10, CYKH13, RP11]. Reliability-Aware [KRS^{+}16, NASM18, ZSEP21, Zhu10]. Reliable [AMKA17, AVR22, GCJD20, GMOB13, HKL^{+}23, MKASJ18, PS19, RKC^{+}22, TBCGO23, WLH^{+}18, DKV14, LHX^{+}14, SKH^{+}12, TTAG14]. Remanence [SZL^{+}17]. remapping [RP03]. Remote [BVGVEA10, CZ23, DNBL22, IFA^{+}16, MGC^{+}23, LCQ^{+}13]. Removal [MM16]. renewable [MKD13]. ReNoC [SSS11]. rental [JKH^{+}13]. Reordering [HLM^{+}24, GRVD12]. reorganization [LCJ13]. Repair [AAS18]. REPAIRS [TJ23]. replacement [RG13]. replay [RAK14]. replication [FS14]. Reporting [MWF^{+}16]. Representation [ADJM19, CAP15, KPK^{+}19, NWA12, RMBS20, YLW15, TKD07]. Representative [LLW^{+}17]. reprogrammable [PO05]. Reprogramming [WLH^{+}18, DLC^{+}14]. Request [BJP24, BCS^{+}23, MBJ^{+}23, SSK23, TTA^{+}20]. Request-Response [BJP24]. Requirement [DHF18, HPP17, LPFL16, LLN^{+}14]. Requirement-Aware [HPP17]. Requirements [LZZK25, GFC^{+}10, UCK^{+}09]. requiring [KHHH14]. ReRAM [BDD^{+}24, LCY^{+}22, ZQ24]. ReRAM-based [LCY^{+}22, ZQ24]. rerouting [SJRS^{+}13b]. Research [BMB16, KOM^{+}23, Shu15a]. Researchers [Shu18a]. reservation [WAD14]. Reservoir [IAS23, ZK25]. ReSG [BDD^{+}24]. Reshaping [LSL^{+}23]. Resilience [OSA^{+}18, SK13, SHME13]. Resilient [BCHB18, GP23, HFA^{+}14, IPL16, MST^{+}16, NBH23, OSA^{+}18, CZK^{+}22, RAK14]. Resistance [YGW^{+}12, DLN13]. Resistant [IYL^{+}23, WH17]. Resistive [JR20]. resolution [GJ13, LG21, PO05]. Resonance [CPP^{+}17]. Resource [ADJM19, BT22, BMF15, CKN^{+}20, CWH^{+}23, CZH^{+}24, DCZB19, DWRR14, GWC^{+}24, HRH^{+}22, HZH^{+}18, IDO^{+}22, KKCS16, LX12, LX22, LVSVRFCG23, LC17, LZJ17, MMY^{+}19, MFG17, MPFG19, NFL^{+}22, NEP23, PMS^{+}24, PS19, REPL15, SPT^{+}21, SCA^{+}24, TLBM15, TMXS17, TAMS18, ZGZ15, ZBG20, ZLX^{+}23, ZSH^{+}19, AF14, BMM13, CHCC13, FF09, GFC^{+}10, HE12, MPZS13, TSG10, UCK^{+}09, WRJL06, Wu10, ZB13, ZMB03, ZLF13]. Resource-Aware [TLBM15]. Resource-Constrained [KKCS16, MFG17, MPFG19, ZBG20, TSG10, UCK^{+}09]. Resource-Constraint [ZSH^{+}19]. Resource-demand [HRH^{+}22]. resource-driven [CHCC13].

Resource-Efficient [DCZB19, LX22, PS19, BT22]. resource-limited [Wu10]. Resources [JSD23, RJM19, SP12, NBGS09]. Response [BJP24, BE17, MBJ^{+}23, SE17, ZLL^{+}19, FF09]. Response-Time [SE17]. Responsive [PVV24, SPP^{+}10]. Resprinting [TBCB15]. Restoring [RPB^{+}19]. restricted [LYL13]. Results [RCDB24, GT05]. Retaining [LKH16]. Retargetable [LPD^{+}20, RDM06]. Retargeting [MFMA17]. Retention [JRR16]. Rethinking [Shu20b, WWT^{+}22]. REtime [ZQ24]. retiming [XHSS10]. Retraining [KSS24]. Retransmissions [RN18]. Retrieval [KNL12]. Reuse [DPNA16, HDZL20, SA21, BCS^{+}06, HKVI05]. Revealing [WM24]. Reverse [HMR23]. Reverse-Engineering [HMR23]. Review [BBMG24, RH23]. Revisited [BBDR12]. rewards [RMM03]. ReWire [PHG^{+}17]. Rewiring [KFY^{+}22]. RF4CE [LCQ^{+}13]. RF4CE-based [LCQ^{+}13]. RFID [LHYQ18, WH17, WXY^{+}18, WYL^{+}19]. RFIDs [CBS19, WLH^{+}18]. Rich [GSS^{+}18, MKOA25]. Rider [MFMA17]. Rigorous [JKH^{+}13, NBM^{+}16, STW13]. Ring [NVB^{+}20, BP14, CTK^{+}13]. Ring-LWE [NVB^{+}20]. ring-mesh [BP14]. RISC [KSFS24, LWS^{+}23]. RISC-V [KSFS24]. RISE [RMK17]. Risk [COC22, LJMP23, RHG^{+}14]. Risk-Aware [COC22]. Risks [Shu17b]. RLUTs [RBNM19]. RMW [MSHS19]. RMW-F [MSHS19]. RMW-Free [MSHS19]. Robot [GMS17, LWZ^{+}16, SLFC19]. Robotic [LPBF24]. Robotics [Shu18b]. Robots [EMVR23, GPT^{+}23]. Robust [BBD23, BMP23, CK23, CQB^{+}15, CHTC07, CAPL11, GD19, GPT^{+}23, GP23, HXH^{+}24, KKL^{+}16, KHB^{+}23, KDB19, LAMS24, MKMGS18, MEK^{+}22, PCM12, PSZ12a, SSD^{+}19, SCF12, SMR20, YV23, FSVG19, ISE10, MRT13].

Robustness [KH23, LWCL25, LM24, RLMP23, Shu19a, ZSM13]. ROS [LPBF24, SLFC19]. ROS-based [LPBF24]. ROSES [WCJ07]. Rotating [SBK^{+}23]. Rotation [SPC^{+}16]. Rotation-Based [SPC^{+}16]. Round [CLLC17]. Round-trip [CLLC17]. Rounding [KNK^{+}25, Rat25]. Rounding-based [KNK^{+}25]. Router [FPGS22, YZA13]. Routerless [IB23]. Routing [DGC^{+}20, GDD20, LLT^{+}17, WLW^{+}24, ZSEP21, CCY^{+}13, JGD^{+}09, PS08b, SJRS^{+}13a]. Rovers [LMS^{+}19]. RQNoC [MST^{+}16]. RSA [KHHH14]. RSIM [LCY^{+}22]. RSU [SAS^{+}23]. RT [DSB17, WLC^{+}18]. RT-WiFi-Based [WLC^{+}18]. RTL [AGG^{+}23, CMK12, PMP17]. RTOS [DHL17, DLD^{+}19, HDR^{+}06, TBFR17]. RTOS-Aware [DHL17]. RTPL [FJM^{+}25]. RTSJ [ZW10]. Ruby [GI24]. Rule [GZZ^{+}16, FZHT13]. rule-based [FZHT13]. Rules [STH17]. Run [MSSP22, OMMK23, SPB^{+}17, YGD^{+}19, BCS^{+}06, GNS04, HMM04]. Run-Time [MSSP22, OMMK23, SPB^{+}17, WWG^{+}18, BCS^{+}06, GNS04, HMM04]. Runs [ACR17]. RunStream [KPC^{+}16]. Runtime [BGK^{+}23, BMF15, CLL21, DAHM16, DSXS15, GSC19, HKC18, HHC^{+}16b, KML13, LKA^{+}18, LL18, MHK^{+}23, MKOA25, MDS^{+}21, MWS15, MBKF15, MMY^{+}19, NEP23, PRS^{+}17, TDD^{+}16, TAMS18, WWN23, WCM^{+}16, YMKH23, ZJC^{+}17, CCY^{+}13, LOG^{+}14, LPFG13, MPZS13, MF13, PB14, SB08, STY^{+}14, YCNCC11, ZBCM09, AVF^{+}09]. Runtime-reconfigurable [YMKH23]. RVSDG [RMBS20].

S [Sus20]. S3PR [WWY13]. SA [GQC^{+}17]. Safe [ASWZ24, RB21, VVKG23, ZCS^{+}05]. SAFE-OPS [ZCS^{+}05]. Safely [SWL07]. Safety [BHAC15, BGO17, DHB^{+}23, ESBK23, GZ12, HCL^{+}17, IPL16, ICW^{+}21, KWKP23, KRR20, LS20, PJL^{+}14, RS07, TCD^{+}19, WKP24, YLW15, ASTPH10, CMA05, DKAL05]. Safety-Critical [IPL16, KWKP23, LS20, WKP24, PJL^{+}14, ASTPH10]. Salesman [Ahm13]. Sample [LYY^{+}17, ZTZ^{+}19]. Samples [PE23]. Sampling [CZHK23]. SAT [AAS18, KDR23]. SAT-Reach [KDR23]. satellite [MB10]. Satisfaction [YF19]. Saving [LKH16, ISG03]. savings [SVN04]. Saying [RH23]. Scabbard [KNK^{+}25]. Scaffolded [GK22]. Scalability [HPBL12, WMRB17, Bec09]. Scalable [AGS^{+}16, ABH^{+}18, HPLD09, JAD19, MBR15, PAF22, PYJL15, SE07, ZSH^{+}24, KYHY14, LCJ13, RGdZS14, SAHE04, TLLL09]. Scale [ABH^{+}18, CJL17, JGX^{+}18, MRA^{+}17, QRW^{+}24, HHB^{+}05, LZZ^{+}23, PS08b]. scaled [LAMS24]. Scaling [BFW^{+}19, CRCR13, GG24, JRR16, LKW^{+}24, RB21, YGW^{+}12, MMR^{+}10]. SCCharts [GSSR^{+}24, SRSM21]. Scenario [CBS19, MSSP22]. scenarios [Gei10]. SCEst [SMR^{+}18]. Schedulability [ARS16, AFMT17, AKD^{+}18, GE18, LZS^{+}18, MEP04, PSD21, PEP05, SD17, ZB13, AF14, AFL13, BC07]. Schedulability-driven [PEP05]. Schedule [RLL^{+}23, SE23, WLC^{+}18, QH07, SAHE04]. Schedule-Based [RLL^{+}23]. scheduled [DF14, ZB13]. Scheduler [GPH24, SSK23, VGB19]. schedulers [SMG04]. schedules [KMB07, SKPL10]. Scheduling [ARS16, AKTM16, ABS^{+}19, BMAB16, BZG19, BGK^{+}23, BE17, BGS^{+}18, CSM24, CPC17, CC13a, CLJ^{+}19, COC22, CHJ22, CAPL11, DBM^{+}15, DLRTB^{+}19, DSB17, EMVR23, FHB^{+}17, GWC^{+}24, GDDD17, GDD20, GWZ16, GE18, HQE20, HGW^{+}20, HSMS16, HDR^{+}06, HTC^{+}16, IPEP12, JCW^{+}16, JZL^{+}15, JGX^{+}18, KGT^{+}23, KB23, LCP^{+}17, LSC19, LJP17, LH18, LSL^{+}23, LLZ^{+}22, LWB18, LHL^{+}19, LLN^{+}14, LX16, LLZ^{+}17, MS21, MG15, MSM21, NFL^{+}22, PMS^{+}24, PCGD21, RLL^{+}23, RDP17, RDSS21, SMW^{+}17, SP19a, SP20, SLCS16, SWX17, SD17, SAS^{+}23, TGV12, TBG^{+}17, TLBM15, TGTT17, VWG^{+}17, WHN^{+}17, WZJ^{+}18, ZZG24, ZGZ15, ZQGZ22, ZLX^{+}23, ZLL^{+}19, ZSEP21, BvB13, CCAP12, DKV14, FZHT13, GNW05, HGL14, IHK04, JP14, KBDV08, LP10, LES14, LQN^{+}13, MTL14, MBFSV07, MALM04, MKD13, NBGS09, NB04, PW13, RGSS04, SL08, SC05, TTAG14, WRJL06, XQ07, XHSS10, YK03, ZW10, ZC04a]. scheduling [ZM07, ZC08]. Scheme [DS11, HHL^{+}23b, KPS23, KJKM16, KNY^{+}17, KCC^{+}16, LX12, LCC^{+}19, LZS^{+}18, LLC^{+}22, LLT^{+}17, PC14, PJS15, RBNM19, SLZ^{+}24, TAMS18, WZY^{+}23, WZD^{+}17, XWL25, YCT16, ABS02, BS13a, CHCC13, CTK^{+}13, JKJ^{+}10, VS08, WSK14]. Schemes [BSJ15, HPO^{+}15, MKASJ18, RCDB24, HL14, SKPL10]. Schizoid [Shu15d]. schizophrenic [YKK^{+}13]. Science [Shu16c]. SCOPES’09 [FM12]. ScorePlus [TSY^{+}16]. SCPS’09 [DSD12]. scratch [ABS02, NDB09, UDB06]. scratch-pad [NDB09, UDB06]. scratch-pad-based [ABS02]. Scratchpad [JLW^{+}15, KBS17, LXK10, PVSG22, Sus20, VCM19, WSMF22, BCDH12, CC13a, ELS08, HZX^{+}14]. Scratchpad-Memory [VCM19]. screening [GJ13]. Scriptable [MWF^{+}16]. SDC [LJLT17, LLP^{+}17, YZZG23]. SDC-causing [LLP^{+}17]. SDF [TBG^{+}13]. SDmesh [DGC^{+}20]. SDRAM [SJC^{+}03, TVK08]. Sea [LYL13]. Seamless [WJ17, ISE10]. SEAMS [MDS^{+}21]. Search [BJ23, FKS^{+}19, MKM^{+}23a, MLAD23, OBO^{+}23, RSK17, SUK23, YS23, PCBW13, TSWL10, VSSS13]. second [NPP13]. Secondary [MKM^{+}23a]. secret [CNK04]. Section [BCEP12, FGIS12, FM12, KM13, NKS12, PS14, Pla12, SRNW16, CP13a, CC14, CP13b, DV13, DSD12, Edi13, Hüb13, JLSK13, PCB12, STW13]. sector [LPC^{+}07]. Secure [ABL^{+}20, AARJ12, CCM17, CBS19, GCJD20, GSN21, JEP16, LMA19, LJR12, LMW^{+}17, LZZ^{+}23, MCP17, MKAA17, PP19, PS08b, PHG^{+}17, RSK17, SYC^{+}17, TNR17, XQQ^{+}24, YGD^{+}17, ZZA^{+}22, Geb04, Geb06, ISTE08]. Securely [WXY^{+}17]. Security [AYS15, BCHL19, CPP^{+}17, CFXY17, FSG23, GQC^{+}17, GSC19, KS22, LJP17, LZZ^{+}19, MCS^{+}15, PKT23, PNRC17, RRKH04, RLL^{+}23, SCKD23, Shu15b, Shu16b, Shu16d, Shu17b, Shu17c, Shu18b, Shu18e, Shu19b, TP19, TBAS17, TP20, VKDG19, WGP04, ZYL^{+}17, CVG^{+}13, PS04, SL04, VS08, XQ07, ZCS^{+}05]. Security-Aware [FSG23, GQC^{+}17, LJP17, RLL^{+}23, TBAS17]. Security-Critical [ZYL^{+}17]. See [WXY^{+}18]. See-through-Wall [WXY^{+}18]. Segment [HSMS16, TBEP16]. Segment-Based [HSMS16]. Segmentation [GGJ12, VAR13]. Segmented [FPGS22]. seizures [MVS^{+}13]. Selection [AbSZ^{+}19, AABG22, BCLS17, DLD^{+}19, GPB^{+}17, HS24, KAKSP15, KBRD22, MTWE20, ZRF^{+}12, BMS13, LSC14, LXL13, SWT^{+}14, SBX08]. Selective [CSCC17, KKL^{+}16, LLPM07, Gar05]. Self [BLG^{+}15, BHET04, CLL^{+}18, DJS16, LYC^{+}18, MDS^{+}21, RJM19, TMXS17, TSO22, TBCGO23, WHL23, YYKK18, DEG11, GLT^{+}13, GNR^{+}10, WYJ^{+}14, Wu10, ZVL04]. Self-Adaptive [RJM19, YYKK18, DEG11]. self-adjusting [Wu10]. Self-aware [DJS16, GNR^{+}10]. Self-Configuring [BLG^{+}15, BHET04, GLT^{+}13]. Self-Gating [WHL23]. Self-Optimizing [MDS^{+}21]. Self-Organized [TMXS17]. Self-Sustained [CLL^{+}18, TBCGO23]. Self-Sustaining [LYC^{+}18]. Self-Testing [BLG^{+}15, GLT^{+}13]. Self-training [TSO22]. self-tuning [WYJ^{+}14, ZVL04]. Semantic [LWZ^{+}16, RALR24]. Semantics [BB13, BGGT23, BV15, CSST08, CMPP23, GGMS24]. Semantics-preserving [CSST08]. Semi [HSMS16, TSO22, ZZG24]. Semi-Clairvoyance [ZZG24]. Semi-Partitioning [HSMS16]. Semi-supervised [TSO22]. semiring [YRF10]. semiring-based [YRF10]. Sense [RSW21]. Sensing [ALZR19, CGZ18, CLL^{+}18, DNBL22, HTR^{+}16, HZGW18, LYC^{+}18, LLG^{+}20, LLW^{+}17, LNA^{+}15, MSR^{+}12, VVKG23, WXY^{+}18, WTSR13, YGHS08]. Sensitive [JJS^{+}25, BO13, ZSEP21]. Sensitivity [RG13, YGD^{+}17]. SensiX [MMA^{+}23]. Sensor [ABC^{+}17, CZK^{+}22, DNBL22, DS11, GM12, GSS^{+}18, GGJ12, HSR18, HCS18, HB16, IPL16, JGX^{+}18, LZZK25, LX22, LFHS18, McI13, MAGR15, PE23, RN18, SG24, SKN17, SLS^{+}19, TSW^{+}17, WWTSM19, ZRF^{+}12, ZZX^{+}15, ZH12a, ZLL^{+}11, ZO16, ZC04c, BS13b, CTK^{+}13, DLN13, DLC^{+}14, GHZH14, HBSA04, HHB^{+}05, KHZS07, KAK05, KXL10, KLC^{+}10, LN04, LLLGR13, LAHS06, MLV09, PS04, PS08a, PS08b, SM13b, SGDP12, VGG^{+}13, WYP^{+}10, YGHS08, ZH12b, ZWY^{+}10, ZLF13]. Sensor-Based [LX22]. SensorGAN [HB24]. Sensors [ABTS24, DL12, EHF24, GSS^{+}18, HZYJ22, HXH^{+}24, HZW^{+}23, LJ24, PP12, WJ17, CNC13, LYL13, NRL13]. Sensory [MMA^{+}23]. Sentries [Shu16b]. Sequence [LL18, ZW13, ZK25]. Sequential [GHB13, LCC^{+}19, MKR13]. Sequential-write-constrained [LCC^{+}19]. Sequentialization [WCM^{+}16]. Sequentially [SMR^{+}18]. Serial [LS17, MKS23, PYL^{+}23, RMH04a]. Series [BT22, ZK25]. Server [ABS^{+}19, BE17, GMS17, MALM04]. Serverless [CBS19]. Servers [AHMT17]. Service [LAZ^{+}16, MST^{+}16, WSMF22, BDP^{+}13, LCJ13, WP11]. Services [JCW^{+}16, KBCL13, PCBW13, SRY13, WTSR13]. Serving [MMA^{+}23]. Set [AJ18, DB19, Fra12, GD19, LTQ^{+}24, AC08, LLPM07, MBFT09, RDM06, RMD09]. Sets [BB15]. SFA [PC14]. SG [WHL23]. SG-Float [WHL23]. SHA256 [GWM16]. Shader [YC16, YW13]. Shading [BCS16]. shadowing [LHX^{+}14]. Shamir [VS08]. shapers [WMT12]. Shaping [OSF19, RC08]. Shared [AD24a, CH08, GGMS24, JSD23, KR18, KRS^{+}16, MBJ^{+}23, NS16, SP12, TGBT17, VGN18, WZJ^{+}18, WKP24, ZGH^{+}19, LPB06, PLKH08, SE10]. shared-memory [LPB06]. Sharing [KSS24, LZJ17, RKK15, SDBD18, VKW^{+}17, VSD^{+}17, BZ13, MSB08, PS08b, ZB13]. SHARP [ARZ^{+}23]. ShaVe [SDBD18]. ShaVe-ICE [SDBD18]. Shift [CDX^{+}19]. Shift-based [CDX^{+}19]. Shingled [CCSC23, CCC^{+}20, LCC^{+}19]. Shortest [GNW05]. Shortest-path [GNW05]. shot [FLF^{+}23]. should [GT05]. Shrunk [ZGH^{+}19]. Side [AAT^{+}21, BPSS24, GW15, GWM16, HMLZ21, IYL^{+}23, JGCS24, MWK^{+}24, RCDB24, WM24]. Side-Channel [AAT^{+}21, BPSS24, GW15, GWM16, HMLZ21, JGCS24, WM24, IYL^{+}23, MWK^{+}24, RCDB24]. sifting [AP09]. sighted [MAS^{+}24]. Signal [DVC21, HW17, SLK^{+}22, SRA12, ZO16, AMN^{+}14, BDB^{+}17, GJ13]. Signals [CCP^{+}19]. Signature [FLF^{+}23, HPO^{+}15, HHL^{+}23b, KKS^{+}23, ZSY19, DLN13]. signature-based [DLN13]. Signatures [ABC^{+}17, AYS15, KSFS24, MKAA17, THG24]. SIKE [EKAK24, SSA21]. Silicon [ASS^{+}23, MSCS16, THA^{+}12, AKB14]. SIMD [FSC^{+}16, HLF^{+}18, SFZX18]. similarity [HE12, LLR14]. Simon [AMKA17]. Simple [SEB12]. Simplex [JBCS16]. Simulating [WRKG16]. Simulation [AVR22, BSM^{+}21, CD19, FKJM18, GD14, LCY^{+}22, MRA^{+}17, RG14, SXMX^{+}18, VF17, WMLA16, ZJC^{+}17, MMSN14, RDM06, RMD09, SPGB24]. Simulation-based [GD14]. Simulation-Driven [FKJM18]. Simulator [BCBJ24, CMP^{+}07]. Simulators [Fra12, SWL^{+}14]. Simulink [BCC^{+}17, DP08, HYY^{+}15, TSCC05]. Simultaneous [BKS^{+}23, LRZ16, SKL^{+}25, TTAG14, OOAL06]. simultaneously [LOXL13]. Single [LMS^{+}19, MM16, PC14, PVV24, WZM17, ZW17, KMB07]. Single- [LMS^{+}19, PVV24]. single-appearance [KMB07]. Single-Threaded [WZM17]. Sink [LFHS18, LLT^{+}17]. Sink-Location [LLT^{+}17]. Sinusoidal [SCM20]. SIPF [SYC^{+}17]. Siphons [CL13, ZW13]. Size [HLM^{+}24, CKIR06, LSK^{+}08, MS21, NP04, NDB09, ZMB03, ZXS03]. Skewed [CSC23]. Skiing [VS05]. Skills [Akd21]. SLAM [BKS^{+}23, GPT^{+}23]. SLAQA [RDSS21]. Sleep [JRR16]. Sleep-Mode [JRR16]. SLEXNet [KBD24]. sliced [SISS24]. Slicing [YKD^{+}24]. Sliding [GW15]. Slimmable [KBD24]. SLISCP [ARH^{+}18]. SLISCP-light [ARH^{+}18]. Slotless [PAS^{+}09]. SM2 [ZSH^{+}19]. Small [HJ19, SRG^{+}15, Shu16c, STB^{+}24]. Smart [BCS16, BSJ15, CYH20, DJS16, HDG^{+}14, LLT^{+}17, MFG16, SCRY16, TSY^{+}16, VP16, XWL25, CHCC13, DEG11, DZR09, Edi14, LWK^{+}10, MSCJ12, SCF12]. SmartLMK [KJKM16]. Smartphone [GW15]. Smartphones [SPT^{+}21, SKK^{+}14, ESM^{+}17]. Smartwatches [AMJ21]. Smooth [SGJ17]. SMR [CCC^{+}20, MSHS19, WLTW24]. SMT [PMM^{+}13]. Snake [BCS16]. Snake-like [BCS16]. Snapshot [LLN^{+}14]. SoC [DJS16, CGM^{+}25, GSC19, JM06, JBN^{+}13, KKO^{+}06, LLP^{+}21, LHM14, MKMGS18, PÖG^{+}13, RH23, TKL^{+}15, YCNCC11, ZDTM19]. SoC-Based [CGM^{+}25]. Social [ZYM16, ZYL^{+}17]. Society [Shu20a]. Socioecological [LAZ^{+}16]. SoCs [BCS^{+}23, COC22, DSXS^{+}14, GD24, HSK18, ISE10, RPB^{+}19, RJM19, VKW^{+}17, XDL^{+}18]. Soft [FND^{+}16, KKL^{+}16, KJK^{+}17b, LJLT17, OSA^{+}18, RJS19, SSK23, SUS^{+}17, TP16, WMGR12, HLD^{+}09, MMSN14, MEP08, SM13a]. Soft-Error [OSA^{+}18, SUS^{+}17]. soft-object [MMSN14]. Softcore [AZS^{+}23]. SoftRM [TMXS17]. Software [Akd21, AKB^{+}25, BVM19, CAP15, CMP17, Dea06, DBFH14, EMVR23, GLC07, GDC19, HJ25, HCS^{+}22, JLSP18, JNI15, KE15, KKL^{+}16, LS13, LLG^{+}20, LMK^{+}18, LBS15, LPBF24, MBLA16, OBA^{+}17, PJS15, RKM24, SWJ^{+}13, Seo18, SISS24, SCM20, SD13, SLFC19, STB^{+}24, TSY^{+}16, TBDdD11, VGN18, VKMP20, WQGR22, YMHB19, YGD^{+}17, ZPZG17, ZQC16, ARJ08, ARJ11, BCLN13, BS13a, BMS13, CMV10, CSVA^{+}05, DZR09, FRRJ07, FZJ08, HG09, HFG13, HQB06, HKLH05, JR20, JM06, KMB07, KASD07, LOG^{+}14, LJR12, LWK^{+}10, MRT13, MLV09, OP06, PGR^{+}08, RP11, Sch07, Sch10, SMG04, SB08, SE07, SVN04, SBF^{+}05, WCJ07, ZCS^{+}05, ZXS03]. Software-Based [KKL^{+}16, LS13]. Software-Controlled [JNI15]. Software-defined [VKMP20, LJR12]. Software-Embedded [DEG11]. Software-Hardware [TSY^{+}16]. Software-Managed [HCS^{+}22]. Software-only [GDC19, BS13a]. software-pipelined [ZXS03]. Solar [ABD^{+}19, JC12, MBB^{+}15, SKN17, SLS^{+}19, TSW^{+}17]. Solar-Powered [TSW^{+}17, MBB^{+}15]. Solid [HWTL23, YWLW23, CCH13, CW14]. Solid-State [HWTL23, YWLW23, CCH13, CW14]. solution [MTL14, ZHCY13]. Solutions [BCHL19, SEB12, SSH14]. Solver [CWJ17, FS24]. Solving [AAS18]. Sorting [JGL21]. SOS [YLW15]. Sound [LJVD23, YV23]. Source [MFMA17, MF12, YV23, CCA^{+}13, FRRJ07]. Source-Level [MFMA17, MF12]. Sourced [DBFH14]. sources [MKD13]. Space [ABL^{+}20, BJT^{+}23, CAP15, CCC^{+}17, DJJ^{+}19, FFA^{+}23, FSVG19, GHD^{+}25, GCJD20, GSN21, JFM23, KCC^{+}16, LLP^{+}21, OFA^{+}15, PSZ12a, SLB^{+}15, SHQX19, WCH^{+}23, WLTW24, YLTY21, ZBG20, BSKB^{+}09, BFQ10, BCG^{+}07, JBN^{+}13, KK05a, KASD07, LM13, MPZS13, OP06, RP03, VAR13]. Space-economic [WCH^{+}23]. Space-Efficient [KCC^{+}16, KASD07]. space-filling [BSKB^{+}09]. Space-Grained [WLTW24]. space-oriented [KK05a]. Spacecraft [LML20]. Spaces [RIMS21]. Sparing [NRZ^{+}24, TTAG14]. Sparse [BDP24, HTB^{+}25, HE24, LYY^{+}17, MLAD23, PKL22, WZS^{+}24]. Sparse-dense [WZS^{+}24]. Sparsity [XDL^{+}18]. Spatial [JBI17, RWL^{+}18, BvB13, GFC^{+}10]. Spatio [DDC^{+}24, LMS^{+}22, SRNW16]. Spatio-Temporal [DDC^{+}24, SRNW16, LMS^{+}22]. Speaker [BJCHA17]. Special [BBM15, BCHL19, BDR24, BDK^{+}23, CDBB24, CS16, CKGN14, CBH22a, CBH22b, CJL17, CGZ18, DPP14, DGG24, DST19, DSXS15, EE16, EH18, FGIS12, FX17, GV21b, GM03, IT16, JC03, KBCL13, KM13, Leo18, LZNL23, MCP17, NKS12, DWCM14, OMMK23, PS14, Pla12, RHG^{+}14, STLX22a, STLX22b, SSA^{+}24b, SCZ20a, SCZ20b, SCKD23, Shu15c, TFC24, VP16, WX17, ZQC16, BM13, CP13a, CC14, CP13b, DV13, DSD12, Edi13, GP07, HCK^{+}08, HTLC10, Hüb13, JB02, JB03, JLSK13, KS10, MS05, PCB12, Sch07, STW13, BCEP12, FM12, Goe14, Gup04, KL13, LB04, SL04, ST05, TEC12]. Specialized [WLZ^{+}23]. specially [ZWY^{+}10]. Specific [BGK^{+}23, DASS12, KOM^{+}23, LMA19, LWZ^{+}16, MFMA17, MPFG19, OMMK23, PSZ12a, SXXS^{+}16a, TBFR17, TKHZ22, ARH^{+}18, BJT^{+}23, BM13, yCBR05, GDB22, JHPR13, RC08, USA^{+}22, WP11, XWHC06].

Specification [DGG24, KDB19, MS13b, BGVZ11, CD10, GHB13, KW10, GV21b]. Specifications [CMK12, LJMP23, NCJF18, OFA^{+}15, RLMP23, WRW^{+}21, YF19, Gar05, MD04, ZNS13]. Specified [TJ23]. Speck [AMKA17]. SPECTRUM [VKMP20]. Speculating [Mus03]. Speculation [PCM^{+}15]. speculative [MF13]. Speech [RRC22, KP13]. Speed [HW17, MSR^{+}17, BBL09, KT14, LLC^{+}13, RV07, SD08]. speed/accuracy [SD08]. Speeds [QWY^{+}18]. Speedup [JSZ^{+}19]. speedups [SVN04]. SpikeHard [CEC23]. Spiking [LMB^{+}22, SCB^{+}22]. Spilling [TBDdD11]. SPIMulator [BCBJ24]. Spin [ZBCM09]. SpinBayes [ADH^{+}23]. Spindles [CCP^{+}19]. Spintronic [ADH^{+}23, BCBJ24]. Spintronic-Based [ADH^{+}23]. Split [HH23, SBR^{+}15, Geb06]. split-mask [Geb06]. Split-Transaction [HH23]. Splitting [KWKP23, PLT23]. SPM [TDD^{+}16]. SPMPool [TDD^{+}16]. SPMs [SDBD18]. SPN [LCLW17]. Sponge [ARH^{+}18]. Sponge-specific [ARH^{+}18]. Sporadic [BE17, DVCC19, FHB^{+}17, Bar13a, HGL14]. Spotting [GV21a]. Squared [CLS16]. SRAM [JRR16, JRSR17]. SRP [YV23]. SRP-PHAT [YV23]. SSD [KSY17, MKM^{+}23a, PX18, SWL^{+}23, SHQX19]. SSD-Based [SWL^{+}23]. SSDs [YWLW23, CSW15, CLL16, HC16, LLZ^{+}22, ZLSQ17, ZAL22]. SSI [BBDR12]. SSP [WBF^{+}06]. Stability [BGO17, ORA16, REPL15]. Stable [CLL16, SWWW17]. Stack [KY17, ZDZ14, ZQGZ22, MSB08, RRW05]. Stacked [SSPP23, HL14]. STADIA [XCZW23]. stage [BJT^{+}23, DBH14]. Stakeholders [YMHB19]. Standard [ABC^{+}17, BCC^{+}17, MWF^{+}16]. Standard-Compliant [MWF^{+}16]. Standby [NRZ^{+}24, TTAG14]. Standby-Sparing [NRZ^{+}24, TTAG14].

Start [HLM^{+}24]. Start-Up [HLM^{+}24].

Stash [ABTS24]. State

[ABA^{+}20, BPP23, CMPP23, CHS15, DQ14, HWTL23, PG24, WRB15, YWLW23, ZPZG17, CCH13, CW14, WGP04]. State-based [CMPP23]. State-of-the-art [WGP04]. statecharts [MS13a]. Stateful [THG24]. Stateless [MKAA17]. statements [YKK^{+}13]. States [SBK^{+}23, ISG03]. Static [BCS16, CYH^{+}17, DHKS15, GWC^{+}24, SMR15, SWL07, SC17, SAMR06, SLCS16, TBFR17, WCM^{+}16, ZMB03, FZHT13, SHME13, ZTRC03]. static-power-efficient [ZTRC03]. Stations [LFHS18]. Statistical [Fra12, MKR13, RLP^{+}21, WZBP19, SGT^{+}13]. Statistics [CNK04]. STDF [DDC^{+}24]. Stealing [LNA^{+}15]. STEAM [HDG^{+}14]. Step [NEP23, WZH13]. Stereo [CYH20, HLLL12, LMS^{+}19]. Stigmergy [GSC19]. Stigmergy-Based [GSC19]. STL [YF19]. STM [CQB^{+}15]. Stochastic [AH13, BSM^{+}21, DMPC23, GRWV22, HCL^{+}17, KDB19, LJMP23, NLSV^{+}19, VM23, WRW^{+}21, XCZW23, MEP04, ÖNG08]. Stop [Val17]. Storage [ABTS24, BJH^{+}24, CBH22a, CBH22b, CCC^{+}17, HRT^{+}22, ICZ^{+}23, JCS^{+}17, KCWH14, KNY^{+}17, Kwo16, LCC^{+}19, LLC^{+}22, MSHS19, MAW22, SWJ^{+}13, SR12a, SCRY16, SSA^{+}24b, WT15, YCT16, YYKK18, BCLN13, CKL04, CWKH12, CYKH13, MRY^{+}10, WKC07]. Store [GKS^{+}22, JBI17]. Store-n-Learn [GKS^{+}22]. Stores [SLZ^{+}24, ZP06]. strands [SWL07]. Strategies [DB19, GDC19, LS17, RWL^{+}18, ISG03]. Strategy [CSCC17, DCZB19, LCY^{+}22, SHQX19, YWLW23, ZWK23]. Stream [BFST19, CJ20, KPC^{+}16, LZZK25, MG15, MV16, MCM^{+}17, PNRC17, SWWW17, CC13b, DSW^{+}09, GHB13, HE12]. Stream-Monitoring [BFST19]. Streaming [BZG19, MAS^{+}24, MASG15, TBG^{+}17, WLK^{+}19, ZSJ12, HFG13, HHB^{+}12, LO13, MAG14, ZNS13]. Streams [HHC^{+}16b]. stress [WGPH13]. string [LHCK04]. string-matching [LHCK04]. Stripe [SWL^{+}23]. Structural [SHK^{+}19]. Structure [BDD^{+}24, SGZS21, ZO16]. Structured [CD12, GDDD17]. Structures [PqBM^{+}15]. STT [ZBCM09, AVR22, LKZ^{+}23, LWZ^{+}24, MMK22, MPT^{+}22, YJD^{+}17]. STT-MRAM-Based [YJD^{+}17, LKZ^{+}23, LWZ^{+}24]. Stubborn [BB15, Val17]. Studies [LKZ^{+}23]. Study [GHPP18, KNK^{+}25, MSD17, RH23, SKL^{+}25, SZL^{+}17, DEG11, LHM14, MSS^{+}03, MSH^{+}14, SKW^{+}07, SPK^{+}12, VJD^{+}07, VDK^{+}08]. Studying [MGB^{+}21]. Style [LWS^{+}23, SPGB24]. Sub [DGC^{+}20, YLDM19]. Sub-Byte [YLDM19]. Sub-networking [DGC^{+}20]. Subgraph [PMP17]. Subject [PSZ12a, VM23]. Subset [PG24]. Subspace [LYY^{+}17]. Subsystem [SR19, KYL13]. Sufficient [ARS16]. Suite [LWK^{+}17, GGGK08]. Suites [SPDLK^{+}17]. Super [JSZ^{+}19]. Super-Linear [JSZ^{+}19]. Superblock [JKJ^{+}10]. superblock-based [JKJ^{+}10]. SuperCISC [JHK^{+}06]. superperfect [LXK10]. supervised [TSO22]. supervisor [ZS05]. Supervisors [WWY13]. Supervisory [DSB17, TR24]. Supplemental [TEC12]. Supplements [Ano13, Ano14]. Support [ZJC^{+}17, HT06, NB04, PZ12, SJRS^{+}13a, VGG^{+}13]. Supported [ZP11, ZSM13]. Supporting [DSXS^{+}14, LDV12, SSH14]. Surrounding [LNA^{+}15]. Surveillance [KLK^{+}19, RMK17, MSCJ12]. Survey [AH13, ASWZ24, BMAB16, BHXP19, BJCHA17, GV21a, LLK^{+}25, MCG22, RCDB24, SP19a, WLC^{+}22, BMP03, WEE^{+}08]. Sustained [CLL^{+}18, TBCGO23]. Sustaining [LYC^{+}18]. SViT [XHK16]. SViT-Based [XHK16]. SVM [CWJ17]. SW [ZDTM19]. Swapping [APF24, KJK17a, LSL^{+}23]. SWARAM [MGLP19]. Switchable [CI17]. Switched [AGS^{+}16, LS09]. switches [SMG04]. Switching [BF17, MSSP22, NNH^{+}14]. Sybil [DBFH14]. Symbolic [BFL18, CBRZ19, TWTH18, WWHT21]. Synaptic [LMB^{+}22]. Synching [CSCC17]. Synchronised [AD24a]. Synchronization [BGJ17, PE23, WXY^{+}17, ZGZ15, AAPN14, CRJ10]. Synchronized [GGMS24, TGP25, GHZH14]. Synchronous [BMM13, BCC^{+}17, BPP23, CMPP23, DHKS15, Gei10, HE24, MS21, SIR^{+}17, WMRB17, ZPZG17, BSB14, CSST08, CC13a, QP03, TBG^{+}13, ZM07]. Synergistic [PHDL18]. Synergy [ZDTM19]. Synterface [SIC19]. Synthesis [BBD23, BF17, BRL16, CWZ^{+}20, yCBR05, CFGM15, CDH^{+}16, EZL^{+}17, FLF17, IYL^{+}23, KMP15, LPFL16, LN19, MSS23, NVB^{+}20, PMDC17, SCB^{+}22, SXXS^{+}16a, TBFR17, VRF15, WWTSM19, ZQD^{+}23, BAR13b, BAR13c, CCA^{+}13, FZK^{+}10, GM03, HG09, HFG13, HVG13, KMB07, MRT13, QP03, SPK^{+}12, ZS05]. Synthesizable [AÖÖ23, GI24]. Synthesizing [LEPP13, SUK23]. Sysfier [RBS^{+}10]. System [AAM^{+}17, AVR22, AKTM16, BTD^{+}18, BBM15, BGK^{+}23, BFQ10, BJT^{+}23, CD12, CGM^{+}25, CLL^{+}18, DST19, DHB^{+}23, DJS16, GIB^{+}12, GPT^{+}23, HZYJ22, HXH^{+}24, HB16, HHL^{+}23a, HWC22, IT16, JC12, JAD19, KSP^{+}12, KHB^{+}23, LX12, Leo18, LKZ^{+}23, LAMS24, LWK^{+}10, LYH^{+}15, MSCJ12, MYL^{+}22, MWS15, MSM21, MGLP19, MEK^{+}22, NCJF18, NBM^{+}16, NLSV^{+}19, OMMK23, PKIT23, PRSV19, QP03, RG14, SA18, SGT^{+}13, SRB23, SSA21, SCRY16, SHL^{+}17, SCKD23, SH23, SR19, SLS^{+}19, SVZ13, UGS^{+}21, VFS^{+}21, WXY^{+}18, WLZ^{+}23, WT15, YKKD23, YCLV^{+}02, YYKK18, YKD^{+}24, ZYM16, ZYL^{+}17, ZX08, ZK25, AMCM06, BE10, BDP^{+}13, BJM13, CWKH12, CSK^{+}02, CHK14b, Dea06, FRRJ07, LJ14, GGGK08, HQB06, HVG13, Hüb13, JBN^{+}13, KCG^{+}05, KZH^{+}06, KGR12, LCQ^{+}13, LKW02, LCC^{+}23, LHCK04, MSS^{+}03, MSL13, NPP13, NNH^{+}14, PK13, PSZ12b, SVP05, Sev05, SPK^{+}12, STY^{+}14]. system [TTAG14, TSBY13, VJD^{+}07, VDK^{+}08, VNK^{+}03, WAD14, YDLC10a, ZHM^{+}14]. system-driven [FRRJ07]. System-Level [KHB^{+}23, LYH^{+}15, NBM^{+}16, ZYM16, ZYL^{+}17, AVR22, MSCJ12, SGT^{+}13, YCLV^{+}02, JBN^{+}13, MSS^{+}03, MSL13, VJD^{+}07, VDK^{+}08]. system-on-a-chip [VNK^{+}03]. System-on-Chip [DJS16, GIB^{+}12, GPT^{+}23, OMMK23, SR19, BJT^{+}23, CHK14b, HQB06, Hüb13, TSBY13]. System-on-Chips [BGK^{+}23, LX12]. System-wide [ZX08]. System/network [BFQ10]. Systematic [CMMD24, JHPR13]. SystemC [BSM^{+}21, CMK12, CD19, FZK^{+}10, GD24, MWF^{+}16, RBS^{+}10, RSB^{+}09, SL16, SWL^{+}14, WMLA16]. SystemC-Based [GD24]. SystemC/C [RSB^{+}09]. SystemC/C-based [RSB^{+}09]. SystemJ [MSCJ12]. Systems [AFS^{+}13, ABD^{+}19, AbSZ^{+}19, AGS^{+}16, AFMT17, AB15, BHAC15, BFW^{+}19, BMAB16, BHXP19, BF17, BGJ17, BDR24, BGO17, BLG^{+}15, BP12, BHL^{+}20, BV15, BJH^{+}24, CKN^{+}20, CLL21, CS16, CQV^{+}13, CKGN14, CBH22a, CBH22b, CMS17, CLLC17, CCC^{+}17, CLJ^{+}19, CHJ22, CWH^{+}23, CS22, CEC23, CQB^{+}15, DAHM16, DWRR14, DHJ^{+}17, DJJ^{+}19, DHL17, DJZ13, DS24, DHF18, DLH16, DBFH14, DB19, DQ14, DAASP21, DVCC19, DJS16, EVS^{+}17, GLP^{+}11, GG24, GD19, GCJD20, GRR23, GZ12, Goe14, GTH^{+}22, GSN21, GE18, HJ25, HKC18, HSMS16, HPP17, HH23, HFA^{+}14, HNY18, HHC^{+}16a, HLLL20, HZX15, HCL^{+}17, HFL^{+}19, Ise17, IPL16, ICW^{+}21, IPEP12, JR20, JLW^{+}15, JZL^{+}15, JEP16, JAD19, KS18, KSS16, Kha13, KPS23, KY17, KSP^{+}12, KJK17a, KJK18, KLK^{+}19, KWKP23, KCBM21, KH23, KCC^{+}16, KNL12, KDR23, KB17, KBRD22, LP19, LS20, LDV12]. Systems [LS12, LAB^{+}23, LMB^{+}22, dFMAdN12, LZL15, LWZ^{+}16, LH18, LZS20, LLG^{+}20, LWZ^{+}24, LJMP23, LLN^{+}14, LX16, LLZ^{+}17, LSL20, LMBL21, LL18, LLK^{+}25, LOF20, MLL^{+}17, MS21, MRA^{+}17, MTWE20, MBKF15, MKS^{+}17, MH19, MS13b, MCG22, Mit21, MMY^{+}19, Mos13, NDZ13, NRZ^{+}24, NBE18, OSF19, OBSO16, PXY^{+}17, PCM^{+}15, PDHC23, PqBM^{+}15, PLM^{+}15, PRS^{+}17, PJT^{+}23, QZXO14, REPL15, RHG^{+}12, RALR24, RRM16, RLP^{+}21, RHG^{+}14, RDSS21, SSK23, SG24, SE23, SMW^{+}17, SCG15, SMR15, SKL^{+}25, SR12b, SWL^{+}23, SP19a, SSPP23, SSA^{+}24b, SDBD18, SCZ20a, SCZ20b, SZL^{+}17, Shu15a, Shu15d, Shu16a, Shu16c, Shu16d, Shu18d, Shu19d, SRR^{+}23, SPGT19, SGJ17, SMR20, SXXS^{+}16b, SLFC19, SCS16, SLE^{+}17, TSP15, TBAS17, TGV12, TJ23, TGP25, TCD^{+}19, TFL16, USA^{+}22, VVKG23, VWG^{+}17, VP16, VM23, WDJ^{+}18, WMGR12, WDY^{+}16, WCK^{+}19, WYL^{+}19, WZBP19, WRW^{+}21]. Systems [WRKG16, WLC^{+}18, WLC^{+}22, WQGR22, WSMF22, WMLM12, WKP24, XSP22, XKK17, YC12, YLW15, YCT16, YHL23, ZYM16, ZYL^{+}17, ZWK23, ZBG20, ZQD^{+}23, ZJC^{+}17, ZLX^{+}23, ZQC16, ARJ08, ARJ11, ASTPH10, AF14, ADI06, AFL13, ABS02, AEF^{+}14, BYD09, BCDH12, BWS14, BP05, Bar13a, BCC^{+}08, BMM13, BBL09, BCS^{+}06, BFQ10, BCG^{+}07, BHET04, CMA05, CCA^{+}13, CSVA^{+}05, CKL04, CWKH12, CYKH13, CCY^{+}13, yCBR05, CRJ10, CJMB05, CRM14, CGV10, CVG^{+}13, CHTC07, DKV14, DDG^{+}13, DF14, DEG11, DW10, DRL^{+}10, ELS08, ESAS14, FZJ08, FS14, FC13, Geb06, GJ13, GMOB13, GD14, GRCV03, GT05, GM03, GNR^{+}10, Gup04, GKW08, HCK^{+}08, HKP08, HTLC10, HLD^{+}09, HQB07, HCQ^{+}14, Hüb13, ISG03, JLSK13, JKH^{+}13, KST^{+}12, KBCL13, KKH^{+}12, LB04, LDRM12, LMST04, LSK^{+}08, LK10, LWB13, LP09a, LLR14, LPFG13, LOXL13, LHX^{+}14, LHM14, MBFSV07, MRY^{+}10]. systems [MSB08, MLL08, MKD13, MSL13, NKP^{+}12, NDB09, PLKH08, PEP05, QH07, RP03, RV03, RS07, RRKH04, RSB^{+}09, SWT^{+}14, Sch07, SE10, SAHE04, SRS03, SL04, SJC^{+}03, ST05, Shu14b, STW13, SVN04, SC05, SBF^{+}05, TRJ05, TM07, TXL^{+}12, TKG13, TSG10, TVK08, VAHC^{+}06, VS05, VHB^{+}13, VGG^{+}13, WMT12, WP11, WLT12, WRJL06, WKC07, Wu10, WMZY13, XQ07, YDLC10b, YRS12, YK03, ZC04b, ZVL04, ZVN05, ZSM13, ZB13, ZP08, ZP09, Zhu10, ZZZ^{+}12, ZC08, KL13]. Systems-on-Chip [CEC23, KS18, WRKG16, GNR^{+}10]. Systolic [YZZ^{+}23, ZRZ^{+}19, WL09].

TAB [ZDL22]. Table [KKS^{+}23, PLT23, RR17, VKW^{+}17, WLWS15, YCLV^{+}02]. Table-based [KKS^{+}23, PLT23]. Tableau [BRR19]. Tail [KSY17, WLTW24, LJLT17]. Tail-DMR [LJLT17]. Tailor [PDL^{+}23]. Tailor-made [PDL^{+}23]. Tailoring [ZGH^{+}19]. Taiwan [HKLH05]. TAMA [ABF^{+}21]. Tame [BJT^{+}23]. Taming [UGS^{+}21]. target [ZC04c]. Task [AR14, BGK^{+}23, CSM24, CGM^{+}25, CPC17, CZH^{+}24, GMS17, HLLL20, LCP^{+}17, dFMAdN12, MTL14, MEP08, NASM18, OHCK24, PCGD21, QP15, RN14, RDSS21, SMW^{+}17, SMR15, SE17, SLS^{+}19, SGW^{+}16, TLBM15, WHN^{+}17, XSP22, ZW17, ZLX^{+}23, Bar13a, DKV14, ESAS14, HWC^{+}20, LK10, LQN^{+}13, LOF20, MEP04, TTAG14, WBS10, ZP09, ZZZ^{+}12, ZC08, TBG^{+}17]. Task-FIFO [TBG^{+}17]. Tasks [ARS16, AKD^{+}18, BAG^{+}20, BGS^{+}18, CLJ^{+}19, FHB^{+}17, HQE20, LJP17, LLZ^{+}17, MBP14, NFL^{+}22, PSD21, SSK21, SPT^{+}23, SD17, WHN^{+}17, XZK^{+}19, ZLL^{+}19, GNW05, HGL14, LP10, MALM04, SPP^{+}10, XQ07, ZC04a, ZX08]. taught [GT05]. Taxicab [ZWH^{+}16]. TBES [CDH^{+}16]. TCAM [SVS21]. TCAM-based [SVS21]. TCX [LWS^{+}23]. TDES [DSB17]. Team [HB16]. Technique [ASD^{+}24, BRR19, DJ23, HPS13, LX16, SFB23, YCK^{+}18, BMS13, JGD^{+}09, ÖNG08, RP11, RMD09, ZXS03]. Techniques [ABS^{+}19, JEP16, KKK^{+}11, KKL^{+}16, KDN^{+}07, LEPP13, LBS15, MCG22, OMMK23, SWJ^{+}13, AP09, AFL13, BMP03, ESAS14, KM09, KK05b, SAYN09]. Technologies [ZQC16, BMP03, HTLC10, WP11]. Technology [CCSC23, SBDK22, DWCM14, SBF^{+}05]. TECS [DST19, Mit21, TEC12, CJL17, CGZ18, SCKD23, Shu20b]. TEFLON [NODP24]. telecom [YCLV^{+}02]. Telomere [MAW22]. Temperature [BGO17, CSM24, HDG^{+}14, JLW^{+}15, NZCS19, SP19b, HCQ^{+}14, KT14, LOXL13, TSBY13]. Temperature-Aware [CSM24, JLW^{+}15]. temperature-based [KT14]. Template [AÖÖ23, CDH^{+}16]. Template-Based [CDH^{+}16]. Temporal [AFS^{+}13, BMNN23, BTL^{+}12, DDC^{+}24, KDB19, LZZK25, LC17, LJMP23, LLN^{+}14, MKS^{+}17, RLMP23, SRNW16, WRW^{+}21, BvB13, LMS^{+}22, MMK22]. Ten [PL13]. tenant [MKM^{+}23a, MMA^{+}23]. Tensor [CLW^{+}20, HA24, HRH^{+}22, KRHC20, LMS^{+}22, LWS^{+}23]. Tensor-Compressed [CLW^{+}20]. TensorRT [JKH22]. TensorRT-Based [JKH22]. Term [GSS^{+}18, JC12, DLC^{+}14]. Terminal [CLW^{+}20]. terminals [ISTE08, ISE10]. Termination [PYL^{+}23]. Ternary [ZDL22]. TESLA [LN04]. Test [CMK12, FKS^{+}19, GE18, KPK^{+}19, MKMGS18, MKM^{+}23b, SPDLK^{+}17, SMZ^{+}21, SHK^{+}19, TSW^{+}17, BMS13, KM09]. Test-case [FKS^{+}19]. Test-Driven [MKMGS18]. Test-pattern [KPK^{+}19].

Testing [BLG^{+}15, BSV17, DHJ^{+}17, DHF18, FMSS15, KH18, LZJ^{+}19, MKS^{+}17, VKDG19, GLT^{+}13, WLT12, BSV17]. Tests [MKR13]. TF [YLDM19]. TF-Net [YLDM19]. TH [SFCW23]. TH-iSSD [SFCW23]. Theoretic [SR12b, SISS24, CAP^{+}07, SPK^{+}12]. Theoretical [CZH^{+}24, MTL14]. Theory [CCKM16, Cul13, HB16, KMP15, KB17, MHT13, SCZ20a, SCZ20b, WDY^{+}16, MRT13, BSKB^{+}09]. Thermal [ARS16, AHMT17, DAHM16, DLRTB^{+}19, FS13, HFA^{+}14, HH13, LSC19, LQN^{+}13, LLG^{+}20, PP24, SSPP23, SP19b, SBK^{+}23, CCY^{+}13]. Thermal-Aware [FS13, LSC19, DLRTB^{+}19, HH13, LQN^{+}13]. Thermal-Resilient [HFA^{+}14]. Thermally [NODP24]. Things [KLK24, BCHL19, BHXP19, BGJ17, RRM16, SXH^{+}19, Shu15a, ZSY19]. Thou [Shu15b]. Thread [MFG17, PLM^{+}15, SPB^{+}17, ZP11, CRAJ10, Dea06, KASD07, SD13]. Threaded [VCM19, WZM17]. thReads [LKB14]. Threat [CLL21, Geb04]. Threshold [GWZ16]. Thresholds [ZGZ15]. Through-Silicon [MSCS16]. Throughput [AV20, HG09, HFG13, HTB^{+}25, HCQ^{+}14, KB23, LS17, LX16, MCM^{+}17, WLK^{+}19, ZDTM19, AÖÖ23, THON12, WBS10]. Throughput-Buffering [KB23]. throughput-constrained [WBS10]. Throughput-driven [HG09]. Throughput-memory [HFG13]. Throughput-Optimized [AV20]. Thru [SYC^{+}17]. Thumb [CYH^{+}17]. Tiered [MBB^{+}15, GJ13]. tight [VLX07]. Tightening [RM10, RDP17]. Tightly [WWHT21]. tile [Mus10]. tile-based [Mus10]. Tiling [VGN18, KK05a]. Time [ARS16, AbSZ^{+}19, ACR17, AYS15, BT22, BMAB16, BBB16, BE17, BGS^{+}18, BB13, BB15, BYIG21, BMMV21, CDBB24, CQV^{+}13, CSM24, CKGN14, CWZ^{+}20, CSH^{+}22, CHJ22, CLS16, CQB^{+}15, DHL17, DJZ13, DS24, EVS^{+}17, FJM^{+}25, FBM16, GAD^{+}24, GG24, GKSB25, GAG15, GZZ^{+}16, GE18, GUC^{+}23, HGW^{+}20, HSMS16, HH23, HFA^{+}14, HHC^{+}16b, IB23, JSZ^{+}19, JAD19, JJS^{+}25, JGX^{+}18, JBCS16, KSS16, KCJ^{+}16, KJKM16, KR18, KMP15, KH23, KHB^{+}23, KB17, LCD18, LN19, dFMAdN12, LZL15, LX16, LL18, MM16, MZG15, MSSP22, MAW22, NPAG12, OSF19, OMMK23, PMS^{+}24, Pau14, PSD21, PJT^{+}23, PNRC17, REPL15, RG14, RMK17, SNTÖ25, SCG15, SMR15, SE17, SP19a, Shu20b, SPB^{+}17, SLCS16, SCS16, SLE^{+}17, SGW^{+}16, SD17, TSP15, TGP25, TKT15, UBF^{+}16, WDJ^{+}18, WMGR12, WHN^{+}17, WWG^{+}18, WGN23, WZ12, XLY18, XQQ^{+}24, YGD^{+}19, ZDZ14]. Time [ZPZG17, ZJC^{+}17, ZLL^{+}19, ZSEP21, ZSJ12, ZK25, AC08, AMCM06, AF14, AFL13, ABC^{+}07, ABI^{+}09, AFG08, BZG19, BVGVEA10, BFST19, BAG^{+}20, BBL09, BCS^{+}06, CMV10, CKL04, CHK14b, CRJ10, CRM14, CHTC07, CCAP12, CAA^{+}24, CRAJ10, DVC^{+}07, DLRTB^{+}19, DF14, DSW^{+}09, DW10, ESBK23, FHK21, GNW05, GHB13, GNS04, HQE20, HMM04, HT06, HTLC10, HHB^{+}12, HCQ^{+}14, KBDV08, KW10, KASD07, KTT13, LG21, LSK^{+}08, LES14, LQN^{+}13, LLR14, LHX^{+}14, LOF20, MMSN14, MEP08, MRY^{+}10, MVS^{+}13, MALM04, MAG14, MCG22, MLL08, MSM21, MKD13, DWCM14, NDB09, NFL^{+}22, NNH^{+}14, NRZ^{+}24, PMM^{+}13, PAP^{+}12, PL10, PS10, QH07, RMM03, SSK23, SE10, SP10, SKPL10, SP20, SL08, SE07, SCA^{+}24, SC05, TM07, TTAG14, TSCC05, UDB06, WMT12, WP11, WAD14, WEE^{+}08, XSP22, YZ08, YK03, ZC04a, ZC04b]. time [ZB13, ZWK23, ZX08, ZJZL20, Zhu10, ZZZ^{+}12]. time- [KASD07]. time-aware [GHB13]. time-portable [ABI^{+}09]. Time-Sensitive [JJS^{+}25, ZSEP21]. Time-Series [ZK25]. Time-Triggered [BBB16, NPAG12]. Time/Run [WWG^{+}18]. Time/Run-Time [WWG^{+}18]. Timed [AD24b, DLRTB^{+}19, Ise17, NCJF18, PG24, UFA^{+}24, BS13b]. Timelines [LBM^{+}24]. Timeliness [YGD^{+}19]. Timely [HLL^{+}23, SHL^{+}17]. timeout [KR14]. Times [AKD^{+}18, PE23, DW10, MEP04]. Timestamp [MKS^{+}17]. Timing [BS22, CD17, CLJ^{+}19, DVCC19, EYG^{+}23, GCU^{+}23, MBKF15, MKS^{+}17, SE23, SK13, TM07, TBEP16, WMRB17, ZWK23, AEF^{+}14, CCB^{+}06, LLR14, MMR^{+}10, TSBY13, VLX07, YRF10, SAMR06]. Timing-Anomaly [CLJ^{+}19]. Tiny [EHF24, GRCV03, SSA^{+}24a]. TinyM [RKM24]. TinyML [APF24, ASD^{+}24, POCR24, RALR24, LM24, TFC24]. TinyNS [SSA^{+}24a]. TinyOS [GLC07, McI13, MLV09]. TIOA [KSS16]. Tissue [VVKG23]. TLB [ZLL^{+}18]. TLC [Kwo16]. TLC-Based [Kwo16]. TLM [BFQ10, CMK12, CD19, LLC^{+}13]. TLM-2.0 [CD19]. TM [PMM^{+}17]. Tolerance [GAS^{+}17, LPE^{+}23, MAGR15, PMM^{+}17, XKK17, AFG08, ZC04b]. Tolerant [BHD15, CPC17, DSB17, IPEP12, MCP17, SA18, SSH14, TMXS17, WDM17, BGD14, JGD^{+}09, LLR14, NRZ^{+}24, PS08a, PAP^{+}12, RMH04b, TAP23, VSSS13]. Tomahawk [AMN^{+}14]. Tool [BKMG12, BGRV15, BMB16, MFMA17, ZLL^{+}18, CCA^{+}13, GGGK08, IBMK10, LAN06, PJL^{+}14]. Toolbox [POCR24]. Tools [SCZ20a, SCZ20b, LP09a, WEE^{+}08]. Toolset [LL15]. Topologies [BCS16]. torque [ZBCM09]. Total [LLW^{+}24]. Trace [LL15, MZG14, UM13]. Trace-Based [LL15]. Traces [CMP17, MZG15, NCJF18, SFB23]. Tracing [PM19, SK19, ZLL^{+}18, ZXCH13]. tracking [ZHM^{+}14]. Tractable [AF14]. Trade [CRCR13, IPEP12, KB23, LDV12, MCM^{+}17, ZRF^{+}12, CLK13, GFC^{+}10, HFG13, SD08, SM13b]. Trade-Off [KB23, ZRF^{+}12, CRCR13, CLK13, HFG13, SD08].

Trade-Offs [IPEP12, MCM^{+}17, LDV12, GFC^{+}10, SM13b]. Tradeoff [JBDD20, MLR^{+}17]. tradeoffs [LPB06]. Trades [OSA^{+}18]. Trading [XQQ^{+}24]. Traffic [MAKO19, OSF19, OHCK24, PMS^{+}24, WRB15, YFPJ14]. Trainable [PKL22]. Training [GK22, HY22, HWC^{+}20, PKL22, SA21, WCK^{+}19, WDM^{+}23, TSO22]. Trajectories [ZWH^{+}16]. Trajectory [LHYQ18]. Transaction [HH23, SD08]. Transactional [PMM^{+}17]. Transactions [BLG^{+}15, Mit21, Shu18c]. Transfer [ANARR^{+}19, ZBCM09, WLH16]. Transfer-based [ANARR^{+}19]. Transferability [SG24]. Transform [PBC22, SISS24]. Transform-based [PBC22]. Transformation [CWZ^{+}20, MFMA17, SPC^{+}16, LLPM07, MBFT09]. transformational [WBF^{+}06]. transformations [AFG08, FRRJ07, FO03]. Transformer [LKW^{+}24, RCS23]. Transient [GSS^{+}18, HA24, YCJW24, VS08, YZA13]. Transiently [ANB^{+}20]. Transition [BV15, GZ12, HPS13, MKM^{+}23b, SMW^{+}17]. Transition-Based [HPS13]. Translating [TSCC05]. Translation [CYH^{+}17, CCC^{+}20, GI24, HLF^{+}18, JKJ^{+}10, KPK^{+}19, Kwo16, PWL^{+}19, PG24, BCDH12, CYKH13, LPC^{+}07, PJL^{+}14, PCK^{+}08, Wu10, ZP08]. Transmission [GQC^{+}17, QRB10, RN18, WLHC18]. Transparency [IPEP12]. Transparency/Performance [IPEP12]. Transparent [IFA^{+}16, ZZA^{+}22]. Transport [AAPN14, CCY^{+}13]. Transport-layer-assisted [CCY^{+}13]. Transposed [MDWL23]. Trapezius [WGPH13]. Trapping [WDM17]. Traveling [Ahm13]. TREAFET [CSM24]. Treble [YMHB19]. Tree [BGK^{+}23, SUK23, SZG^{+}23, LCC^{+}19, SLZ^{+}24, WCB20, WCH^{+}23, WCB20, WKC07]. Tree-based [BGK^{+}23]. Trees [CSH^{+}22, MG15].

Trigger [HMLZ21]. Triggered [BBB16, FND^{+}16, NPAG12, AAPN14]. Trinity [Shu15a, LYC^{+}18]. trip [CLLC17]. Trireme [ZEJ^{+}23]. Trivial [ASJ21]. Trojan [CZHK23, GLS^{+}23, HMLZ21, MKM^{+}23b, SMZ^{+}21]. Trojans [PMP17, SSK21]. truly [WL09]. Trust [RHG^{+}14, Shu18e, XWL25]. Trusted [DQ14, ARJ08]. TrustFlow [BHL^{+}20]. TrustFlow-X [BHL^{+}20]. Trustworthy [SH23]. Truthful [XQQ^{+}24]. TSCH [GAD^{+}24]. TSN [PE23, WLW^{+}24]. TTL [MKS^{+}17]. Tune [MSRM24]. Tuning [CSK^{+}02, MSRM24, KST^{+}12, KZH^{+}06, WYJ^{+}14, ZVL04]. Turn [ABF^{+}21]. Turn-aware [ABF^{+}21]. Tutorial [CK23, GV21a, GLS^{+}23, GTvdBC25, MKOA25, PKT23]. Tutorials [SCKD23]. TV [JMO14, KSK13, RIMS21]. Tweakable [MKASJ18]. Two [AR14, LH18, RBNM19, JB02, JB03, WL09]. two-dimensional [WL09]. Two-party [RBNM19]. Two-Phase [LH18]. Two-Type [AR14]. TX2 [DZL^{+}22]. TyBox [POCR24]. Tÿcho [CJ20]. Type [AR14]. Types [TBDdD11].

UAV [FGL^{+}19]. UBAR [SPT^{+}21]. ubiquitous [BDP^{+}13]. Ultra [ABL^{+}20, BHD15, BDB^{+}17, BTA^{+}19, DBH14, GJ13, JRR16, PKIT23, TTB23]. Ultra-Low [BTA^{+}19, JRR16, BDB^{+}17]. Ultra-Low-Energy [ABL^{+}20]. Ultra-low-power [DBH14, PKIT23, TTB23]. UML [KKO^{+}06, PG24]. UML-based [KKO^{+}06]. Uncertain [CZK^{+}22]. Uncertainties [GD19]. Uncertainty [ADH^{+}23, AKTM16, KDB19, RSF20]. Uncontrolled [ZH12a]. Undergraduate [KCG^{+}05]. Underminer [BTD^{+}18]. Understanding [ALZR19, CKN^{+}20]. Unequal [YC12]. Unfoldings [KH18, SPDLK^{+}17]. Unified [CCR^{+}14, FMSS15, GOC^{+}22, KKCS16, PKL22, SSPP23, TGV12, VKDG19, YDS^{+}22, ZDL22, KXL10, OMA^{+}13, SB08]. Uniform [HGW^{+}20]. Unintentional [ISOD21]. Uniprocessor [MBFSV07]. Unit [AGG^{+}23, FGL^{+}19, MKS23]. Units [HRH^{+}22, RKK15, DBH14, RGdZS14]. Universal [BCLS17, SCRY16]. unknown [NDB09]. unnecessary [Mus03]. Unreliable [BHAC15]. Unsupervised [ASD^{+}24]. Update [HZX15, TM15, YCT16, WLH16]. Updates [EMVR23, YMHB19]. Updating [KLK24, MH19]. Upgrade [LCC^{+}19]. upon [Bar13a]. UPP2SF [PJL^{+}14]. UPPAAL [KSS16, BGVZ11, BS13b]. Upsampling [LG21]. Urban [ZWH^{+}16]. Usage [LWZ^{+}16, ZDZ14]. Usage-Specific [LWZ^{+}16]. USB [ISOD21, WGN23]. Use [GTH^{+}22, NRZ^{+}24, SWL^{+}23, Shu18d, GNW05, WMT12]. used [TSWL10]. User [CSC23, ESM^{+}17, GDB22, HTC^{+}16, KJKM16, SPT^{+}21, CLK13, LCQ^{+}13, SSK^{+}22, WLH16]. User- [SPT^{+}21]. User-aware [ESM^{+}17]. User-Centric [HTC^{+}16]. User-Perceivable [CSC23]. User-Perceived [KJKM16]. User-Profile-Driven [WLH16]. User-specific [GDB22]. Users [Li21, YTL^{+}20]. Using [AK21, AHM19, ADH^{+}23, AR14, BHD15, BDP24, BRR19, BMF15, BHXP19, BAR13c, BS22, CL13, CRCR13, CMP17, DLRTB^{+}19, DVC21, DL12, FKJM18, FGL^{+}19, FLF17, GD24, GSS^{+}18, GZZ^{+}16, GLS^{+}23, GGJ12, HDZL20, HB16, HPS13, HCL^{+}17, HZW^{+}23, IYL^{+}23, JLSP18, KKK^{+}11, KBD24, LHP^{+}23, LPP^{+}21, LM24, MSHS19, MM16, MMK22, MV16, MSD17, NS17, NWA12, NGL17, NDZ13, OHCK24, Pau14, PRM21, PP12, PLT23, QWY^{+}18, RC17, RB21, SOL^{+}16, SK19, SR12b, SMZ^{+}21, SKN17, SP19b, SIC19, SLE^{+}17, VF17, WWTSM19, WRB15, WZH13, WHL23, YF19, ZWH^{+}16, ZK25, ALZR19, ASS^{+}23, BJ23, BSKB^{+}09, BMNN23, BAR13b, BGVZ11, BCS^{+}06, CLR05, DNNP14, GGGK08, HMM04, HPLD09, KBDV08, KMB07, KM09, KASD07, KTT13, LPC^{+}07, LML^{+}23, MSCJ12, MMS06, MSR^{+}17, MMD04, MKM^{+}23b, MSH^{+}14, NKP^{+}12, NRL13, OBA^{+}17, OMA^{+}13, OOAL06, OP06]. using [PJL^{+}14, PSZ12b, SHME13, SSB24, SB08, SWWY13, TSG10, UDB06, WCJ07, WMRB17, ZKKC05]. Utility [DWRR14, MWS15, GKW08, WRJL06]. Utility-Based [DWRR14]. Utilization [ABD^{+}19, BCS^{+}23, CCC^{+}17, JJS^{+}25]. Utilizing [MEK^{+}22, ACK^{+}13, CTK^{+}13]. UWB [CCC^{+}14].

V [KSFS24]. v11 [TEC12]. V2 [RKM24]. V2X [ESBK23]. VADF [MMH^{+}23]. Validation [BGGT23, GDA13, GD24, KKL^{+}16, SMR20, MF13, MD04]. Value [UM13, MKM^{+}23a, YG02]. valued [VF17]. Variability [PSZ12a, PAP^{+}12, BJM13, SGT^{+}13]. Variability-Aware [PSZ12a]. Variability-tolerant [PAP^{+}12]. Variable [CD19, GVS^{+}20, MWS15, SR12a, BAR13b, KD08, KK05b, LXL13, OOAL06]. variable-length [BAR13b, KD08]. Variant [WCB20]. Variation [BTL^{+}12, LLW^{+}24, MASG15, WDM17, MAG14]. Variations [GM12, REPL15]. Varying [KH23, GFC^{+}10]. VBN [LLP^{+}21]. VCPSS’09 [FGIS12]. VecRA [YC16]. Vector [AGG^{+}23, LRZ16, PJT^{+}23, Sus20, YC16, BZ13]. Vector-Aware [YC16]. Vector-Length [Sus20]. Vectorization [LWB18, SFZX18]. Vectors [MKM^{+}23b]. Vehicle [WMLM12, XZK^{+}19, SKH^{+}12]. Vehicles [CYH20, LHL^{+}19, SAS^{+}23, VA18]. Vehicular [ANARR^{+}19, ASD^{+}24, LLN^{+}14, OHCK24, XQQ^{+}24, ZLZ^{+}24]. Verifiable [LML^{+}23]. Verification [BDD^{+}24, CMA05, CD17, DJZ13, DHF18, GHR15, GZ12, HCL^{+}17, Ise17, KL13, LMK^{+}18, LHL^{+}19, LML20, PNRC17, SSB24, SVZ13, TCD^{+}19, WZBP19, WMLM12, YLW15, ZJC^{+}17, ZZA^{+}22, ASTPH10, GD14, PB14, PS08a, RS07, RBS^{+}10]. Verified [BPP23, CMP23, JBCS16, MS23, Shu18c]. Verifying [BBMG24, ICW^{+}21, MLL08, WRW^{+}21]. Verilog [GI24]. Versatile [MMH^{+}23, XYLC23]. Versus [CSW15, Shu19d]. vertical [STY^{+}14]. Via [FKS^{+}19, LSL^{+}23, MSCS16, AMJ21, AAS18, CZ23, CBRZ19, FS13, Gar05, HLF^{+}18, HKL^{+}23, HZW^{+}23, ISOD21, JBDD20, LPFL16, LXK10, LLW^{+}24, LHYQ18, LJLT17, LYY^{+}17, MBLA16, PS08a, SC17, SSPP23, SWS23, SHQX19, TBCB15, VKW^{+}17, WRW^{+}21, WM24, WLK^{+}19, WCM^{+}16, ZQD^{+}23]. Victim [AK21]. Video [CLW^{+}20, DDC^{+}24, FS13, GDC19, JCW^{+}16, KKD^{+}12, LDV12, MYL^{+}22, PJWY12, PCBW13, PDHC23, RMK17, LCJ13, PÖG^{+}13]. viewpoint [MTL14]. Violations [CMP17]. ViPES [CS16, Goe14]. Virtual [CMP17, DSXS^{+}14, Goe14, GD24, KCWH14, MBLA16, SL16, SLB^{+}15, SSB24, SXXS^{+}16b, SXMX^{+}18, VKW^{+}17, ABC^{+}07, CH08, CGV10, NKP^{+}12, ZP08, CS16]. Virtualization [BE17, CWH^{+}16, JAD19, KR18, LRZ16, PDL^{+}23, SJRS^{+}13a]. Virtualized [SDBD18]. Virtualizing [AKI^{+}23]. VirtualSoC [BMB16]. Vision [BONA22, GK22, HJ25, HLLL12, KBRD22, LMS^{+}19, LKW^{+}24, LM24, MAS^{+}24, RCS23, WJ17]. Vision-assisted [WJ17]. Vision-based [BONA22]. VISTA [KZH^{+}06]. Visual [BKS^{+}23, GPT^{+}23, SKL^{+}25]. Visualizing [PP12]. ViT4Mal [RCS23]. Viterbi [RRC22]. VLIW [CIC^{+}08, CIC^{+}09, MG05, SCS16, SWX17, TBDdD11, ZTD^{+}06]. Volatile [HCS^{+}22, ZZA^{+}22, HXZ^{+}13, LZJ^{+}20, WLWS15]. Voltage [BHD15, JRR16, LAMS24, SWJ^{+}13, YGW^{+}12, CCP^{+}19, IHK04, KK05b, LK10, MMR^{+}10, SAHE04, YK03, ZC04a]. Voltage-scaled [LAMS24]. voltages [HQB06]. Volume [Ano13, Ano14, ZXCH13]. Voting [Shu18d]. VPO [KZH^{+}06]. VR [DS24]. vs [CRCR13, LAB^{+}23]. Vulnerabilities [AAT^{+}21, TP20]. Vulnerability [AHM19, YZZG23, LS13]. Vulnerability-aware [AHM19].

Wait [CQB^{+}15, OZ22]. Wait-Free [CQB^{+}15, OZ22]. Waiting [GHR15]. Walking [VKW^{+}17]. Wall [CDX^{+}19, WXY^{+}18]. Walsh [PBC22]. Walsh-Hadamard [PBC22]. Warbler [MFG16]. WARM [WCH^{+}23]. WARM-tree [WCH^{+}23]. warning [PL10]. warning-zone-length [PL10]. Warp [DS24, LV09]. WasmAndroid [WWN23]. Water [CLL^{+}18]. Wavefront [KWKP23]. Wavelet [CCP^{+}19, MM16, GFC^{+}10, PZ12]. Wavelet-based [CCP^{+}19, GFC^{+}10]. WCET [BCD^{+}22, BFL18, CCR^{+}14, GLYY14, HZX15, KBS17, SWX17, ZW17]. WCET-Aware [KBS17, SWX17]. WCRT [SSK21, WMRB17, XZK^{+}19]. Weak [GHKS15]. Weakly [HQE20, PSD21, SD17]. Weakly-hard [HQE20]. Wear [HCS^{+}22, JNI15, CCH13, PMPP14]. Wear-Leveling [HCS^{+}22, CCH13, PMPP14]. Wearable [BTA^{+}19, HZYJ22, HZW^{+}23, HB24, LJ24, MM16, PJL^{+}17, VVKG23, WJ17, ZRF^{+}12, GJ13, NRL13]. Wearables [TKV^{+}18]. week [HCK^{+}08]. Weight [GWC^{+}24, KSS24, SWWY13]. weight-function [SWWY13]. wfspan [OZ22]. Where [Shu15b]. While [RDP17, JHK^{+}06]. White [RIMS21]. WHS’09 [NKS12]. WIA [JGX^{+}18]. Wide [ZWH^{+}16, ZX08]. WiFi [WLC^{+}18]. Wildfire [TBCGO23]. Will [BVM19].

Window [GW15, RLL^{+}23]. Winograd [AABG22]. Wire [SA18]. Wireless [AZHC19, ABC^{+}17, BTL^{+}12, CLL^{+}18, CNC13, DS11, FBM16, HPBL12, HSR18, JGX^{+}18, KKD^{+}12, LOD18, MSM21, MAGR15, PP12, TSW^{+}17, VKDG19, YCK^{+}18, ZZX^{+}15, ZH12b, ZO16, CTK^{+}13, CD10, CHTC07, DLC^{+}14, GW08, GHZH14, HBSA04, JLSK13, KS10, KAK05, KXL10, KDN^{+}07, MVS^{+}13, MLV09, PCM12, PS04, QRB10, RGSS04, SWT^{+}14, YGHS08, ZWY^{+}10, ZLF13]. WirelessHART [ZHM^{+}14]. Within [ZGZ15, GHD^{+}25]. without [DKAL05]. Witness [BMMV21]. Wits [SWJ^{+}13]. Work [AÖÖ23, LNA^{+}15]. Work-efficient [AÖÖ23]. Workbench [BGRV15]. working [NPP13]. Workload [FS13, HHC^{+}16a, NEP23, PAP^{+}12, RC08, WW09]. Workloads [PVV24, LQN^{+}13]. World [SIC19]. Worst [BMMV21, KT14, LOF20, MKE18, NS16, PE23, SRR^{+}23, WZ12, YF19, WEE^{+}08, YZ08]. Worst-Case [MKE18, NS16, SRR^{+}23, WZ12, BMMV21, KT14, PE23, YF19, WEE^{+}08, YZ08]. Write [GMCC18, HCS^{+}22, HWTL23, HXZ^{+}13, HC16, PXY^{+}17, WLWS15, WCH^{+}23, YWLW23, LCC^{+}19]. Write-Activity-Aware [WLWS15]. Write-efficient [WCH^{+}23]. Write-Related [YWLW23]. Writing [LLC^{+}22]. WRSNs [LZS^{+}18]. WSANs [FC13]. WSN [DL12, HS24]. WSNs [BSA17, LYC^{+}18, LLT^{+}17].

X [BHL^{+}20, CMP23]. X25519 [TV19].

Xbars [BMP23]. Xilinx

[AZS^{+}23, GKSB25]. XimSwap [APF24]. XIP [PLKH08]. XML [MSH^{+}14]. XNOR [RLG20]. XOR [LZZ^{+}19]. XOR-Based [LZZ^{+}19]. XploreNAS [BMP23]. XR [THON12]. XScale [CMP^{+}07, VJD^{+}07, VDK^{+}08]. XScale-based [VJD^{+}07, VDK^{+}08]. XTREM [CMP^{+}07]. xTune [KST^{+}12].

Years [PL13]. Yield [HL14, PRK15]. Yield-enhancement [HL14].

Zeroconf [BGVZ11]. ZigBee [MLV09]. ZIP [GHD^{+}25]. ZIP-CNN [GHD^{+}25]. zone [PL10]. ZPP [DJ23].

References

Alam:2022:WCD

[AABG22] Syed Asad Alam, Andrew Anderson, Barbara Barabasz, and David Gregg. Winograd convolution for deep neural networks: Efficient point selection. ACM Transactions on Embedded Computing Systems, 21(6): 80:1–80:??, November 2022. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3524069.

Andalam:2017:NEM

[AAM^{+}17] Sidharta Andalam, Nathan Allen, Avinash Malik, Partha S. Roop, and Mark Trew. A novel emulation model of the cardiac conduction system. ACM Transactions on Embedded Computing Systems, 16(5s):157:1–157:??, October 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Anjum:2014:TTA

[AAPN14] Omer Anjum, Mubashir Ali, Teemu Pitkänen, and Jari Nurmi. Transport triggered architecture to perform carrier synchronization for LTE. ACM Transactions on Embedded Computing Systems, 13(4):89:1–89:??, February 2014. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Azimi:2017:HHF

[AAR^{+}17] Iman Azimi, Arman Anzanpour, Amir M. Rahmani, Tapio Pahikkala, Marco Levorato, Pasi Liljeberg, and Nikil Dutt. HiCH: Hierarchical fog-assisted computing architecture for healthcare IoT. ACM Transactions on Embedded Computing Systems, 16(5s):174:1– 174:??, October 2017. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Arora:2012:ILM

[AARJ12] Divya Arora, Najwa Aaraj, Anand Raghunathan, and Niraj K. Jha. INVISIOS: a lightweight, minimally intrusive secure execution environment. ACM Transactions on Embedded Computing Systems, 11(3):60:1– 60:??, September 2012. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Attie:2018:MPR

[AAS18] Paul C. Attie, Kinan Dak Al Bab, and Mouhammad Sakr. Model and program repair via SAT solving. ACM Transactions on Embedded Computing Systems, 17(2):32:1–32:??, April 2018. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Aydin:2021:HSC

[AAT^{+}21] Furkan Aydin, Aydin Aysu, Mohit Tiwari, Andreas Gerstlauer, and Michael Orshansky. Horizontal sidechannel vulnerabilities of post-quantum key exchange and encapsulation protocols. ACM Transactions on Embedded Computing Systems, 20(6):110:1–110:22, November 2021. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL

https://dl.acm.org/doi/
10.1145/3476799.

Anand:2015:ICL

[AB15] Kapil Anand and Rajeev Barua. Instruction-cache locking for improving embedded systems performance. ACM Transactions on Embedded Computing Systems, 14(3):53:1–53:??, April 2015. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Ahmed:2020:FEE

[ABA^{+}20] Saad Ahmed, Naveed Anwar Bhatti, Muhammad Hamad Alizai, Junaid Haroon Siddiqui, and Luca Mottola. Fast and energy-efficient state checkpointing for intermittent computing. ACM Transactions on Embedded Computing Systems, 19(6): 45:1–45:27, November 2020. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3391903.

Armbruster:2007:RTJ

[ABC^{+}07] Austin Armbruster, Jason Baker, Antonio Cunei, Chapman Flack, David Holmes, Filip Pizlo, Edward Pla, Marek Prochazka, and Jan Vitek. A real-time Java virtual machine with applications in avionics. ACM Transactions on Embedded Computing Systems, 7(1): 5:1–5:49, December 2007. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Ateniese:2017:LCS

[ABC^{+}17] Giuseppe Ateniese, Giuseppe Bianchi, Angelo T. Capossele, Chiara Petrioli, and Dora Spenza. Low-cost standard signatures for energyharvesting wireless sensor networks. ACM Transactions on Embedded Computing Systems, 16(3):64:1– 64:??, July 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Ahmed:2019:OPM

[ABD^{+}19] Rehan Ahmed, Bernhard Buchli, Stefan Draskovic, Lukas Sigrist, Pratyush Kumar, and Lothar Thiele. Optimal power management with guaranteed minimum energy utilization for solar energy harvesting systems. ACM Transactions on Embedded Computing Systems, 18(4):30:1–30:??, August 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/ft_gateway.
cfm?id=3317679.

Aligholipour:2021:TTA

[ABF^{+}21] Rashid Aligholipour, Mohammad Baharloo, Behnam Farzaneh, Meisam Abdollahi, and Ahmad Khonsari. TAMA: Turn-aware mapping and architecture — a power-efficient network-on-chip approach. ACM Transactions on Embedded Computing Systems, 20(5):44:1– 44:24, July 2021. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3462700.

Ara:2018:SAM

[ABH^{+}18] Hadi Alizadeh Ara, Amir Behrouzian, Martijn Hendriks, Marc Geilen, Dip Goswami, and Twan Basten. Scalable analysis for multiscale dataflow models. ACM Transactions on Embedded Computing Systems, 17(4): 80:1–80:??, August 2018. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Auerbach:2009:LLT

[ABI^{+}09] Joshua Auerbach, David F. Bacon, Daniel Iercan, Christoph M. Kirsch, V. T. Rajan, Harald Röck, and Rainer Trummer. Low-latency time-portable real-time programming with Exotasks. ACM Transactions on Embedded Computing Systems, 8(2):15:1– 15:??, January 2009. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Aerabi:2020:DSE

[ABL^{+}20] Ehsan Aerabi, Milad Bohlouli, Mohammad Hasan Ahmadi Livany, Mahdi Fazeli, Athanasios Papadimitriou, and David Hely. Design space exploration for ultralow-energy and secure IoT MCUs. ACM Transactions on Embedded Computing Systems, 19(3):19:1– 19:34, July 2020. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3384446.

Avissar:2002:OMA

[ABS02] Oren Avissar, Rajeev Barua, and Dave Stewart. An optimal memory allocation scheme for scratchpad-based embedded systems. ACM Transactions on Embedded Computing Systems, 1(1):6–26, November 2002. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Awan:2019:TAM

[ABS^{+}19] Muhammad Ali Awan, Konstantinos Bletsas, Pedro F. Souto, Benny Akesson, and Eduardo Tovar. Techniques and analysis for mixedcriticality scheduling with mode-dependent server execution budgets. ACM Transactions on Embedded Computing Systems, 18(5s): 109:1–109:??, October 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/ft_gateway.
cfm?id=3358234.

Al-bayati:2019:PSD

[AbSZ^{+}19] Zaid Al-bayati, Youcheng Sun, Haibo Zeng, Marco Di Natale, Qi Zhu, and Brett H. Meyer. Partitioning and selection of data consistency mechanisms for multicore real-time systems. ACM Transactions on Embedded Computing Systems, 18(4): 35:1–35:??, August 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/ft_gateway.
cfm?id=3320271.

Alsubhi:2024:SFE

[ABTS24] Arwa Alsubhi, Simeon Babatunde, Nicole Tobias, and Jacob Sorber. Stash: Flexible energy storage for intermittent sensors. ACM Transactions on Embedded Computing Systems, 23(2): 18:1–18:??, March 2024. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3641511.

Aamodt:2008:CTI

[AC08] Tor M. Aamodt and Paul Chow. Compile-time and instruction-set methods for improving floating- to fixedpoint conversion accuracy. ACM Transactions on Embedded Computing Systems, 7(3):26:1–26:??, April 2008. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Anagnostopoulos:2013:PAD

[ACK^{+}13]  Iraklis  Anagnostopoulos,

Jean-Michel Chabloz, Ioannis Koutras, Alexandros Bartzas, Ahmed Hemani, and Dimitrios Soudris. Poweraware dynamic memory management on many-core platforms utilizing DVFS. ACM Transactions on Embedded Computing Systems, 13(1s):40:1–40:??, November 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Andre:2017:PPO

[ACR17] Étienne André, Thomas Chatain, and César Rodríguez. Preserving partial-order runs in parametric time Petri nets. ACM Transactions on Embedded Computing Systems, 16(2):43:1–43:??, April 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Aguado:2024:SSM

[AD24a] Joaquín Aguado and Alejandra Duenas. Synchronised shared memory and model checking. ACM Transactions on Embedded Computing Systems, 23(5):70:1– 70:??, September 2024. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3626188.

Arasteh:2024:FLT

[AD24b] Emad M. Arasteh and Rainer Dömer. Fast looselytimed deep neural network models with accurate memory contention. ACM Transactions on Embedded Computing Systems, 23(5):75:1– 75:??, September 2024. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3607548.

Ahmed:2023:SAH

[ADH^{+}23] Soyed Tuhin Ahmed, Kamal Danouchi, Michael Hefenbrock, Guillaume Prenat, Lorena Anghel, and Mehdi B. Tahoori. Spin-Bayes: Algorithm-hardware co-design for uncertainty estimation using Bayesian in-memory approximation on spintronic-based architectures. ACM Transactions on Embedded Computing Systems, 22(5s):131:1– 131:??, October 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3609116.

Alur:2006:PAR

[ADI06] Rajeev Alur, Thao Dang, and Franjo Ivančić. Predicate abstraction for reachability analysis of hybrid systems. ACM Transactions on Embedded Computing Systems, 5(1):152– 199, February 2006. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Arrestier:2019:NRD

[ADJM19] Florian Arrestier, Karol Desnos, Eduardo Juarez, and Daniel Menard. Numerical representation of directed acyclic graphs for efficient dataflow embedded resource allocation. ACM Transactions on Embedded Computing Systems, 18(5s):101:1– 101:??, October 2019. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3358225.

Axer:2014:BTP

[AEF^{+}14] Philip Axer, Rolf Ernst, Heiko Falk, Alain Girault, Daniel Grund, Nan Guan, Bengt Jonsson, Peter Marwedel, Jan Reineke, Christine Rochange, Maurice Sebastian, Reinhard Von Hanxleden, Reinhard Wilhelm, and Wang Yi. Building timing predictable embedded systems. ACM Transactions on Embedded Computing Systems, 13(4):82:1– 82:??, February 2014. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Ahmed:2014:TSA

[AF14] Masud Ahmed and Nathan Fisher. Tractable schedulability analysis and resource allocation for real-time multimodal systems. ACM Transactions on Embedded Computing Systems, 13(2s): 65:1–65:??, January 2014. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Ayav:2008:IFT

[AFG08] Tolga Ayav, Pascal Fradet, and Alain Girault. Implementing fault-tolerance in real-time programs by automatic program transformations. ACM Transactions on Embedded Computing Systems, 7(4):45:1–45:??, July 2008. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Anand:2013:CCS

[AFL13] Madhukar Anand, Sebastian Fischmeister, and Insup Lee. A comparison of compositional schedulability analysis techniques for hierarchical real-time systems. ACM Transactions on Embedded Computing Systems, 13(1):2:1–2:??, August 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Alur:2017:SBR

[AFMT17] Rajeev Alur, Vojtech Forejt, Salar Moarref, and Ashutosh Trivedi. Schedulability of bounded-rate multimode systems. ACM Transactions on Embedded Computing Systems, 16(3):85:1– 85:??, July 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Abbas:2013:PTL

[AFS^{+}13] Houssam Abbas, Georgios Fainekos, Sriram Sankaranarayanan, Franjo Ivancić, and Aarti Gupta. Probabilistic temporal logic falsification of cyber-physical systems. ACM Transactions on Embedded Computing Systems, 12(2s):95:1–95:??, May 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Allamigeon:2017:FMC

[AGG^{+}17] Xavier Allamigeon, Stéphane Gaubert, Eric Goubault, Sylvie Putot, and Nikolas Stott. A fast method to compute disjunctive quadratic invariants of numerical programs. ACM Transactions on Embedded Computing Systems, 16(5s):166:1– 166:??, October 2017. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Alam:2023:RIF

[AGG^{+}23] Syed Asad Alam, David Gregg, Giulio Gambardella, Thomas Preusser, and Michaela Blott. On the RTL implementation of FINN matrix vector unit. ACM Transactions on Embedded Computing Systems, 22(6):94:1– 94:??, November 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3547141.

Allamigeon:2016:SAM

[AGS^{+}16] Xavier Allamigeon, Stéphane Gaubert, Nikolas Stott, Éric Goubault, and Sylvie Putot. A scalable algebraic method to infer quadratic invariants of switched systems. ACM Transactions on Embedded Computing Systems, 15(4): 69:1–69:??, August 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Alaghi:2013:SSC

[AH13] Armin Alaghi and John P. Hayes. Survey of stochastic computing. ACM Transactions on Embedded Computing Systems, 12(2s):92:1– 92:??, May 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Ahmed:2013:HGA

[Ahm13] Zakir Hussain Ahmed. A hybrid genetic algorithm for the Bottleneck Traveling Salesman Problem. ACM Transactions on Embedded Computing Systems, 12(1): 9:1–9:??, January 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Ahmed:2019:CRU

[AHM19] Alif Ahmed, Yuanwen Huang, and Prabhat Mishra. Cache reconfiguration using machine learning for vulnerabilityaware energy optimization. ACM Transactions on Embedded Computing Systems, 18(2):15:1–15:??, April 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/ft_gateway.
cfm?id=3309762.

Ahmed:2017:DAT

[AHMT17] Rehan Ahmed, Pengcheng Huang, Max Millen, and Lothar Thiele. On the design and application of thermal isolation servers. ACM Transactions on Embedded Computing Systems, 16(5s): 165:1–165:??, October 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Amanollahi:2018:ERD

[AJ18] Saba Amanollahi and Ghassem Jaberipur. Extended redundant-digit instruction set for energy-efficient processors. ACM Transactions on Embedded Computing Systems, 17(3):70:1– 70:??, June 2018. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Agarwal:2021:IPH

[AK21] Sukarn Agarwal and Hemangee K. Kapoor. Improving the performance of hybrid caches using partitioned victim caching. ACM Transactions on Embedded Computing Systems, 20(1):5:1– 5:27, January 2021. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3411368.

Abdel-Khalek:2014:PSP

[AKB14] Rawan Abdel-Khalek and Valeria Bertacco. Postsilicon platform for the functional diagnosis and debug of networks-on-chip. ACM Transactions on Embedded Computing Systems, 13(3s): 112:1–112:??, March 2014. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Argotti:2025:OQM

[AKB^{+}25] Yann Argotti, Yasmine Kenfaoui, Claude Baron, Alain Abran, and Philippe Esteban. An operational quality model of embedded software aligned with ISO 25000. ACM Transactions on Embedded Computing Systems, 24(1):9:1– 9:??, January 2025. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3691642.

Andersson:2018:SAT

[AKD^{+}18] BJörn Andersson, Hyoseung Kim, Dionisio De Niz, Mark Klein, Ragunathan (Raj) Rajkumar, and John Lehoczky. Schedulability analysis of tasks with corunner-dependent execution times. ACM Transactions on Embedded Computing Systems, 17(3):71:1– 71:??, June 2018. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Akdur:2021:SGI

[Akd21] Deniz Akdur. Skills gaps in the industry: Opinions of embedded software practitioners. ACM Transactions on Embedded Computing Systems, 20(5):43:1– 43:39, July 2021. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3463340.

Anderson:2023:VPM

[AKI^{+}23] Jeff Anderson, Engin Kayraklioglu, Hamid Reza Imani, Chen Shen, Mario Miscuglio, Volker J. Sorger, and Tarek El-Ghazawi. Virtualizing a post-Moore’s law analog mesh processor: The case of a photonic PDE accelerator. ACM Transactions on Embedded Computing Systems, 22(2):38:1– 38:??, March 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3544971.

Asyaban:2016:ASB

[AKTM16] Sedigheh Asyaban, Mehdi Kargahi, Lothar Thiele, and Morteza Mohaqeqi. Analysis and scheduling of a batteryless mixed-criticality system with energy uncertainty.

ACM Transactions on Embedded Computing Systems, 16(1):23:1–23:??, November 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Alur:2005:P

[AL05] Rajeev Alur and Insup Lee. Preface. ACM Transactions on Embedded Computing Systems, 4(4):707, November 2005. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Algahtani:2024:HAA

[Alg24] Eyad Algahtani. A hardware approach for accelerating inductive learning in description logic. ACM Transactions on Embedded Computing Systems, 23(4):62:1– 62:??, July 2024. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3665277.

Almeida:2022:DDO

[ALV^{+}22] Mario Almeida, Stefanos Laskaridis, Stylianos I. Venieris, Ilias Leontiadis, and Nicholas D. Lane. DynO: Dynamic onloading of deep neural networks from cloud to device. ACM Transactions on Embedded Computing Systems, 21(6):71:1– 71:??, November 2022. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3510831.

Abkenar:2019:GRU

[ALZR19] Amin B. Abkenar, Seng W.

Loke, Arkady Zaslavsky, and Wenny Rahayu. GroupSense: Recognizing and understanding group physical activities using multi-device embedded sensing. ACM Transactions on Embedded Computing Systems, 17(6): 98:1–98:??, January 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

AbouGhazaleh:2006:COS

[AMCM06] Nevine AbouGhazaleh, Daniel Mossé, Bruce R. Childers, and Rami Melhem. Collaborative operating system and compiler power management for real-time applications. ACM Transactions on Embedded Computing Systems, 5(1):82–115, February 2006. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Akbari:2021:FHA

[AMJ21] Ali Akbari, Jonathan Martinez, and Roozbeh Jafari. Facilitating human activity data annotation via contextaware change detection on smartwatches. ACM Transactions on Embedded Computing Systems, 20(2):15:1– 15:20, March 2021. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3431503.

Ahir:2017:LAR

[AMKA17] Prashant Ahir, Mehran Mozaffari-Kermani, and Reza Azarderakhsh. Lightweight architectures for reliable and fault detection Simon and Speck cryptographic algorithms on FPGA. ACM Transactions on Embedded Computing Systems, 16(4): 109:1–109:??, August 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Arnold:2014:TPH

[AMN^{+}14] Oliver Arnold, Emil Matus, Benedikt Noethen, Markus Winter, Torsten Limberg, and Gerhard Fettweis. Tomahawk: Parallelism and heterogeneity in communications signal processing MP-SoCs. ACM Transactions on Embedded Computing Systems, 13(3s):107:1–107:??, March 2014. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Ali:2019:CCT

[ANARR^{+}19] G. G. Md. Nawaz Ali, Md. Noor-A-Rahim, Md. Ashiqur Rahman, Beshah Ayalew, Peter H. J. Chong, and Yong Liang Guan. Cooperative cache transferbased on-demand network coded broadcast in vehicular networks. ACM Transactions on Embedded Computing Systems, 18(4):38:1– 38:??, August 2019. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3329865.

Ahmed:2020:DEC

[ANB^{+}20] Saad Ahmed, Muhammad Nawaz, Abu Bakar, Naveed Anwar Bhatti, Muhammad Hamad Alizai, Junaid Haroon Siddiqui, and Luca Mottola. Demystifying energy consumption dynamics in transiently powered computers. ACM Transactions on Embedded Computing Systems, 19(6):47:1– 47:25, November 2020. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3391893.

Anonymous:2013:AOS

[Ano13] Anonymous. Abstracts: Online supplements volume 12, number 1s, volume 12, number 2s. ACM Transactions on Embedded Computing Systems, 12(4):112:1– 112:??, June 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Anonymous:2014:AOS

[Ano14] Anonymous. Abstracts: Online supplements volume 13, number 1s volume 13, number 2s volume 13, number 3s volume 13, number 4s volume 13, number 5s. ACM Transactions on Embedded Computing Systems, 13(4):99:1–99:??, November 2014. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Ahangari:2023:HBH

[AÖÖ23] Hamzeh Ahangari, Muhammet Mustafa Özdal, and Özcan Öztürk. HLSbased high-throughput and work-efficient synthesizable graph processing template pipeline. ACM Transactions on Embedded Computing Systems, 22(2):34:1– 34:??, March 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3529256.

Ahn:2009:RCT

[AP09] Minwook Ahn and Yunheung Paek. Register coalescing techniques for heterogeneous register architecture with copy sifting. ACM Transactions on Embedded Computing Systems, 8(2): 16:1–16:??, January 2009. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Ahmad:2020:FFB

[AP20] Afzal Ahmad and Muhammad Adeel Pasha. FF-Conv: an FPGA-based accelerator for fast convolution layers in convolutional neural networks. ACM Transactions on Embedded Computing Systems, 19(2):15:1– 15:24, March 2020. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3380548.

Ancilotto:2024:XMM

[APF24] Alberto Ancilotto, Francesco Paissan, and Elisabetta Farella. XimSwap: Manyto-many face swapping for TinyML. ACM Transactions on Embedded Computing Systems, 23(3):49:1– 49:??, May 2024. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3603173.

Ahmad:2016:EMB

[APRC16] Awais Ahmad, Anand Paul, Mazhar Rathore, and Hangbae Chang. An efficient multidimensional big data fusion approach in machineto-machine communication. ACM Transactions on Embedded Computing Systems, 15(2):39:1–39:??, May 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Andersson:2014:PGT

[AR14] Björn Andersson and Gurulingesh Raravi. Provably good task assignment for two-type heterogeneous multiprocessors using cutting planes. ACM Transactions on Embedded Computing Systems, 13(5s):160:1– 160:??, September 2014. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

An:2016:MBD

[ARDG16] Xin An, Eric Rutten, Jean-Philippe Diguet, and Abdoulaye Gamatié. Modelbased design of correct controllers for dynamically reconfigurable architectures. ACM Transactions on Embedded Computing Systems, 15(3):51:1–51:??, July 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Altawy:2018:SLT

[ARH^{+}18] Riham Altawy, Raghvendra Rohit, Morgan He, Kalikinkar Mandal, Gangqiang Yang, and Guang Gong. SLISCP-light: Towards hardware optimized spongespecific cryptographic permutations. ACM Transactions on Embedded Computing Systems, 17(4):81:1– 81:??, August 2018. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Aaraj:2008:ADH

[ARJ08] Najwa Aaraj, Anand Raghunathan, and Niraj K. Jha. Analysis and design of a hardware/software trusted platform module for embedded systems. ACM Transactions on Embedded Computing Systems, 8(1):8:1– 8:??, December 2008. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Aaraj:2011:FDE

[ARJ11] Najwa Aaraj, Anand Raghunathan, and Niraj K. Jha. A framework for defending embedded systems against software attacks. ACM Transactions on Embedded Computing Systems, 10(3):33:1– 33:??, April 2011. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Ambrose:2012:RII

[ARP12] Jude A. Ambrose, Roshan G.

Ragel, and Sri Parameswaran. Randomized instruction injection to counter power analysis attacks. ACM Transactions on Embedded Computing Systems, 11(3): 69:1–69:??, September 2012. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Ahmed:2016:NSC

[ARS16] Rehan Ahmed, Parameswaran Ramanathan, and Kewal K. Saluja. Necessary and sufficient conditions for thermal schedulability of periodic real-time tasks under fluid scheduling model. ACM Transactions on Embedded Computing Systems, 15(3): 49:1–49:??, July 2016. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Aminabadi:2023:SAE

[ARZ^{+}23] Reza Yazdani Aminabadi, Olatunji Ruwase, Minjia Zhang, Yuxiong He, Jose-Maria Arnau, and Antonio Gonazalez. SHARP: an adaptable, energy-efficient accelerator for recurrent neural networks. ACM Transactions on Embedded Computing Systems, 22(2):30:1– 30:??, March 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3552513.

Andrade:2024:OPV

[ASD^{+}24] Pedro Andrade, Ivanovitch Silva, Marianne Diniz, Thommas Flores, Daniel G. Costa, and Eduardo Soares. Online processing of vehicular data on the edge through an unsupervised TinyML regression technique. ACM Transactions on Embedded Computing Systems, 23(3): 50:1–50:??, May 2024. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3591356.

Atoofian:2021:REG

[ASJ21] Ehsan Atoofian, Zayan Shaikh, and Ali Jannesari. Reducing energy in GPG-PUs through approximate trivial bypassing. ACM Transactions on Embedded Computing Systems, 20(2): 13:1–13:27, March 2021. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3429440.

Afifi:2023:GGN

[ASS^{+}23] Salma Afifi, Febin Sunny, Amin Shafiee, Mahdi Nikdast, and Sudeep Pasricha. GHOST: a graph neural network accelerator using silicon photonics. ACM Transactions on Embedded Computing Systems, 22(5s):102:1– 102:??, October 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3609097.

Adler:2010:CBM

[ASTPH10] Rasmus Adler, Ina Schaefer, Mario Trapp, and Arnd Poetzsch-Heffter. Componentbased modeling and verification of dynamic adaptation in safety-critical embedded systems. ACM Transactions on Embedded Computing Systems, 10(2):20:1– 20:??, December 2010. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Alnahawi:2024:TNG

[ASWZ24] Nouri Alnahawi, Nicolai Schmitt, Alexander Wiesmaier, and Chiara-Marie Zok. Toward next generation quantum-safe eIDs and eMRTDs: a survey. ACM Transactions on Embedded Computing Systems, 23(2): 34:1–34:??, March 2024. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3585517.

Atoofian:2020:ACG

[Ato20] Ehsan Atoofian. Approximate cache in GPGPUs. ACM Transactions on Embedded Computing Systems, 19(5):40:1–40:22, November 2020. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3407904.

Azari:2020:ETO

[AV20] Elham Azari and Sarma Vrudhula. ELSA: a throughputoptimized design of an LSTM accelerator for energyconstrained devices. ACM Transactions on Embedded Computing Systems, 19(1): 3:1–3:21, February 2020. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3366634.

Alle:2009:RRR

[AVF^{+}09] Mythri Alle, Keshavan Varadarajan, Alexander Fell, Ramesh Reddy C., Nimmy Joseph, Saptarsi Das, Prasenjit Biswas, Jugantor Chetia, Adarsh Rao, S. K. Nandy, and Ranjani Narayan. RE-DEFINE: Runtime reconfigurable polymorphic ASIC. ACM Transactions on Embedded Computing Systems, 9(2):11:1–11:??, October 2009. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Asifuzzaman:2022:PPE

[AVR22] Kazi Asifuzzaman, Rommel Sánchez Verdejo, and Petar Radojković. Performance and power estimation of STT–MRAM main memory with reliable system-level simulation. ACM Transactions on Embedded Computing Systems, 21(1):6:1– 6:25, January 2022. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3476838.

Aysu:2015:FRT

[AYS15] Aydin Aysu, Bilgiday Yuce, and Patrick Schaumont. The future of real-time security: Latency-optimized lattice-based digital signatures. ACM Transactions on Embedded Computing Systems, 14(3):43:1–43:??, May 2015. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Arghavani:2019:CLB

[AZHC19] Abbas Arghavani, Haibo Zhang, Zhiyi Huang, and Yawen Chen. Chimp: a learning-based power-aware communication protocol for wireless body area networks. ACM Transactions on Embedded Computing Systems, 18(2):18:1–18:??, April 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

                                        /dl.acm.org/ft_gateway.
                                        cfm?id=3309763.

Awais:2023:TOS

[AZS^{+}23] Muhammad Awais, Ali Zahir, Syed Ayaz Ali Shah, Pedro Reviriego, Anees Ullah, Nasim Ullah, Adam Khan, and Hazrat Ali. Toward optimal softcore carryaware approximate multipliers on Xilinx FPGAs. ACM Transactions on Embedded Computing Systems, 22(4): 76:1–76:??, July 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3564243.

Behrouzian:2020:FAR

[BAG^{+}20] Amir Behrouzian, Hadi Alizadeh Ara, Marc Geilen, Dip Goswami, and Twan Basten. Firmness analysis of real-time tasks. ACM Transactions on Embedded Computing Systems, 19(4): 26:1–26:24, July 2020. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3398328.

Baruah:2013:PST

[Bar13a] Sanjoy Baruah. Partitioning sporadic task systems upon memory-constrained multiprocessors. ACM Transactions on Embedded Computing Systems, 12(3):78:1– 78:??, March 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Ben-Asher:2013:BUV

[BAR13b] Yosi Ben-Asher and Nadav Rotem. The benefits of using variable-length pipelined operations in highlevel synthesis. ACM Transactions on Embedded Computing Systems, 13(3):52:1– 52:??, December 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Ben-Asher:2013:UMP

[BAR13c] Yosi Ben-Asher and Nadav Rotem. Using memory profile analysis for automatic synthesis of pointers code. ACM Transactions on Embedded Computing Systems, 12(3):68:1– 68:??, March 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Boucheneb:2013:RIS

[BB13] Hanifa Boucheneb and Kamel Barkaoui. Reducing interleaving semantics redundancy in reachability analysis of time Petri nets. ACM Transactions on Embedded Computing Systems, 12(1): 7:1–7:??, January 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Boucheneb:2015:SST

[BB15] Hanifa Boucheneb and Kamel Barkaoui. Stubborn sets for time Petri nets. ACM Transactions on Embedded Computing Systems, 14(1):11:1– 11:??, January 2015. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Baudart:2016:LTT

[BBB16] Guillaume Baudart, Albert Benveniste, and Timothy Bourke. Loosely timetriggered architectures: Improvements and comparisons. ACM Transactions on Embedded Computing Systems, 15(4):71:1–71:??, August 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Baruah:2023:OSR

[BBD23] Sanjoy Baruah, Alan Burns, and Robert Ian Davis. Optimal synthesis of robust IDK classifier cascades. ACM Transactions on Embedded Computing Systems, 22(5s): 150:1–150:??, October 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3609129.

Boissinot:2012:SPR

[BBDR12] Benoit Boissinot, Philip Brisk, Alain Darte, and Fabrice Rastello. SSI properties revisited. ACM Transactions on Embedded Computing Systems, 11S(1):21:1–21:??, 2012. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Bini:2009:MCE

[BBL09] Enrico Bini, Giorgio Buttazzo, and Giuseppe Lipari. Minimizing CPU energy in real-time systems with discrete speed management. ACM Transactions on Embedded Computing Systems, 8(4):31:1–31:??, July 2009. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Barkaoui:2015:GES

[BBM15] Kamel Barkaoui, Luca Bernardinello, and Andrey Mokhov. Guest editorial for special issue application of concurrency to system design. ACM Transactions on Embedded Computing Systems, 14(4):62:1–62:??, December 2015. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Boudardara:2024:RAM

[BBMG24] Fateh Boudardara, Abderraouf Boussif, Pierre-Jean Meyer, and Mohamed Ghazel. A review of abstraction methods toward verifying neural networks. ACM Transactions on Embedded Computing Systems, 23(4):58:1–58:??, July 2024. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3617508.

Bordoloi:2007:ISA

[BC07] Unmesh D. Bordoloi and Samarjit Chakraborty. Interactive schedulability analysis. ACM Transactions on Embedded Computing Systems, 7(1):7:1–7:27, December 2007. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Bera:2024:SSP

[BCBJ24] Pavia Bera, Stephen Cahoon, Sanjukta Bhanja, and Alex Jones. SPIMulator: a spintronic processing-in-memory simulator for racetracks. ACM Transactions on Embedded Computing Systems, 23(6):94:1– 94:??, November 2024. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3645112.

Benveniste:2008:CHR

[BCC^{+}08] Albert Benveniste, Benoît Caillaud, Luca P. Carloni, Paul Caspi, and Alberto L. Sangiovanni-Vincentelli. Composing heterogeneous reactive systems. ACM Transactions on Embedded Computing Systems, 7(4):43:1–43:??, July 2008. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Bourke:2017:SLS

[BCC^{+}17] Timothy Bourke, Francois Carcenac, Jean-Louis Colaço, Bruno Pagano, Cédric Pasteur, and Marc Pouzet. A synchronous look at the Simulink standard library. ACM Transactions on Embedded Computing Systems, 16(5s):176:1– 176:??, October 2017. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Bai:2022:FCW

[BCD^{+}22] Zhenyu Bai, Hugues Cassé, Marianne De Michiel, Thomas Carle, and Christine Rochange. A framework for calculating WCET based on execution decision diagrams. ACM Transactions on Embedded Computing Systems, 21(3):26:1–26:26, May 2022. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3476879.

Barletta:2024:CAM

[BCDD24] Marco Barletta, Marcello Cinque, Luigi De Simone, and Raffaele Della Corte. Criticality-aware monitoring and orchestration for containerized industry 4.0 environments. ACM Transactions on Embedded Computing Systems, 23(1):5:1– 5:??, January 2024. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3604567.

Baiocchi:2012:EDB

[BCDH12] José A. Baiocchi, Bruce R. Childers, Jack W. Davidson, and Jason D. Hiser. Enabling dynamic binary translation in embedded systems with scratchpad memory. ACM Transactions on Embedded Computing Systems, 11(4): 89:1–89:??, December 2012. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Berekovic:2012:ISS

[BCEP12] Mladen Berekovic, Samarjit Chakraborty, Petru Eles, and Andy D. Pimentel. Introduction to the Special Section on ESTIMedia’08. ACM Transactions on Embedded Computing Systems, 11S(1):11:1–11:??, 2012. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Bueno:2007:RRP

[BCG^{+}07] David Bueno, Chris Conger, Alan D. George, Ian Troxel, and Adam Leko. RapidIO for radar processing in advanced space systems. ACM Transactions on Embedded Computing Systems, 7(1): 1:1–1:38, December 2007. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Bueno:2010:ORA

[BCG10] David Bueno, Chris Conger, and Alan D. George. Optimizing rapidIO architectures for onboard processing. ACM Transactions on Embedded Computing Systems, 9(3):18:1–18:??, February 2010. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Bhattacharjee:2018:CRM

[BCHB18] Sukanta Bhattacharjee, Yi-Ling Chen, Juinn-Dar Huang, and Bhargab B. Bhattacharya. Concentrationresilient mixture preparation with digital microfluidic lab-on-chip. ACM Transactions on Embedded Computing Systems, 17(2):49:1– 49:??, April 2018. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Batina:2019:ISI

[BCHL19] Lejla Batina, Sherman S. M. Chow, Gerhard Hancke, and Zhe Liu. Introduction to the special issue on cryptographic engineering for Internet of Things: Security foundations, lightweight solutions, and attacks. ACM Transactions on Embedded Computing Systems, 18(3): 22:1–22:??, June 2019. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3322641.

Baek:2013:EEH

[BCLN13] Seungjae Baek, Jongmoo Choi, Donghee Lee, and Sam H. Noh. Energy-efficient and high-performance software architecture for storage class memory. ACM Transactions on Embedded Computing Systems, 12(3):81:1– 81:??, March 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Blindell:2017:CPU

[BCLS17] Gabriel Hjort Blindell, Mats Carlsson, Roberto Castañeda Lozano, and Christian Schulte. Complete and practical universal instruction selection. ACM Transactions on Embedded Computing Systems, 16(5s):119:1–119:??, October 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Biswas:2006:MOP

[BCS^{+}06] Surupa Biswas, Thomas Carley, Matthew Simpson, Bhuvan Middha, and Rajeev Barua. Memory overflow protection for embedded systems using run-time checks, reuse, and compression. ACM Transactions on Embedded Computing Systems, 5(4):719– 752, November 2006. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Baka:2016:NSS

[BCS16] Maria-Iro Baka, Francky Catthoor, and Dimitrios Soudris. Near-static shading exploration for smart photovoltaic module topologies based on snake-like configurations. ACM Transactions on Embedded Computing Systems, 15(2):27:1– 27:??, May 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Brilli:2023:ECM

[BCS^{+}23] Gianluca Brilli, Roberto Cavicchioli, Marco Solieri, Paolo Valente, and Andrea Marongiu. Evaluating controlled memory request injection for efficient bandwidth utilization and predictable execution in heterogeneous SoCs. ACM Transactions on Embedded Computing Systems, 22(1):7:1– 7:??, January 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3548773.

Bathen:2014:ERC

[BD14] Luis Angel D. Bathen and Nikil D. Dutt. Embedded RAIDs-on-chip for bus-based chip-multiprocessors. ACM Transactions on Embedded Computing Systems, 13(4): 83:1–83:??, February 2014. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Basu:2017:IUL

[BDB^{+}17] Soumya Basu, Loris Duch, Rubén Braojos, Giovanni Ansaloni, Laura Pozzi, and David Atienza. An inexact ultra-low power bio-signal processing architecture with lightweight error recovery. ACM Transactions on Embedded Computing Systems, 16(5s):159:1–159:??, October 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Bhunia:2024:RDS

[BDD^{+}24] Kousik Bhunia, Arighna Deb, Kamalika Datta, Muhammad Hassan, Saeideh Shirinzadeh, and Rolf Drechsler. ReSG: a data structure for verification of majoritybased in-memory computing on ReRAM crossbars. ACM Transactions on Embedded Computing Systems, 23(6):90:1–90:??, November 2024. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3615358.

Bertozzi:2015:PRA

[BDG^{+}15] Davide Bertozzi, Stefano Di Carlo, Salvatore Galfano, Marco Indaco, Piero Olivo, Paolo Prinetto, and Cristian Zambelli. Performance and reliability analysis of crosslayer optimizations of NAND flash controllers. ACM Transactions on Embedded Computing Systems, 14(1): 7:1–7:??, January 2015. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Bosio:2023:SIA

[BDK^{+}23] Alberto Bosio, Lara Dolecek, Alexandra Kourfali, Sri Parameswaran, and Alessandro Savino. Special issue: “Approximation at the Edge”. ACM Transactions on Embedded Computing Systems, 22(4):72:1– 72:??, July 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3605757.

Bharanitharan:2013:DMS

[BDP^{+}13] K. Bharanitharan, Jiun-Ren Ding, Anand Paul, Kuen-Ming Lee, and Ting-Wei Hou. Dependable management system for ubiquitous camera array service in an elder-care center. ACM Transactions on Embedded Computing Systems, 12(2): 29:1–29:??, February 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Basak:2024:ELM

[BDP24] Barnali Basak, Pallab Dasgupta, and Arpan Pal. Efficient low-memory implementation of sparse CNNs using encoded partitioned hybrid sparse format. ACM Transactions on Embedded Computing Systems, 23(6):103:1– 103:??, November 2024. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3687239.

Bhasin:2024:SIP

[BDR24] Shivam Bhasin, Fabrizio De Santis, and Francesco Regazzoni. Special issue on postquantum cryptography for embedded systems. ACM Transactions on Embedded Computing Systems, 23(2): 26:1–26:??, March 2024. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3641852.

Basten:2010:EMD

[BE10] Twan Basten and Rolf Ernst.

Editorial: Model-driven embedded-system design. ACM Transactions on Embedded Computing Systems, 10(2):15:1–15:??, December 2010. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Beckert:2017:RTA

[BE17] Matthias Beckert and Rolf Ernst. Response time analysis for sporadic server based budget scheduling in real time virtualization environments. ACM Transactions on Embedded Computing Systems, 16(5s):161:1– 161:??, October 2017. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Beckett:2009:PSM

[Bec09] Paul Beckett. Power scalability in a mesh-connected reconfigurable architecture. ACM Transactions on Embedded Computing Systems, 9(2):13:1–13:??, October 2009. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Benerecetti:2017:ASS

[BF17] Massimo Benerecetti and Marco Faella. Automatic synthesis of switching controllers for linear hybrid systems: Reachability control. ACM Transactions on Embedded Computing Systems, 16(4):104:1–104:??, August 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Ballabriga:2018:SWC

[BFL18] Clément Ballabriga, Julien Forget, and Giuseppe Lipari. Symbolic WCET computation. ACM Transactions on Embedded Computing Systems, 17(2):39:1–39:??, April 2018. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Bombieri:2010:SND

[BFQ10] Nicola Bombieri, Franco Fummi, and Davide Quaglia. System/network design-space exploration based on TLM for networked embedded systems. ACM Transactions on Embedded Computing Systems, 9(4):37:1–37:??, March 2010. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Baumeister:2019:FSM

[BFST19] Jan Baumeister, Bernd Finkbeiner, Maximilian Schwenger, and Hazem Torfah. FPGA stream-monitoring of realtime properties. ACM Transactions on Embedded Computing Systems, 18(5s): 88:1–88:??, October 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/ft_gateway.
cfm?id=3358220.

Balsamo:2019:MPN

[BFW^{+}19] Domenico Balsamo, Benjamin J. Fletcher, Alex S. Weddell, Giorgos Karatziolas, Bashir M. Al-Hashimi, and Geoff V. Merrett. Momentum: Power-neutral performance scaling with intrinsic MPPT for energy harvesting computing systems. ACM Transactions on Embedded Computing Systems, 17(6):93:1–93:??, January 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Banaiyanmofrad:2014:NBF

[BGD14] Abbas Banaiyanmofrad, Gustavo Girão, and Nikil Dutt. NoC-based fault-tolerant cache design in chip multiprocessors. ACM Transactions on Embedded Computing Systems, 13(3s):115:1– 115:??, March 2014. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Brun:2023:EDA

[BGGT23] Lélio Brun, Christophe Garion, Pierre-Loïc Garoche, and Xavier Thirioux. Equationdirected axiomatization of Lustre semantics to enable optimized code validation. ACM Transactions on Embedded Computing Systems, 22(5s):151:1–151:??, October 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3609393.

Bennett:2017:DDS

[BGJ17] Terrell R. Bennett, Nicholas Gans, and Roozbeh Jafari. Data-driven synchronization for Internet-of-Things systems. ACM Transactions on Embedded Computing Systems, 16(3):69:1–69:??, July 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Basaklar:2023:DDT

[BGK^{+}23] Toygun Basaklar, A. Alper Goksoy, Anish Krishnakumar, Suat Gumussoy, and Umit Y. Ogras. DTRL: Decision tree-based multiobjective reinforcement learning for runtime task scheduling in domain-specific systemon-chips. ACM Transactions on Embedded Computing Systems, 22(5s):113:1– 113:??, October 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3609108.

Bhat:2017:PTS

[BGO17] Ganapati Bhat, Suat Gumussoy, and Umit Y. Ogras. Power-temperature stability and safety analysis for multiprocessor systems. ACM Transactions on Embedded Computing Systems, 16(5s): 145:1–145:??, October 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Borgstrom:2015:PCW

[BGRV15] Johannes Borgstrom, Ramunas Gutkovas, Ioana Rodhe, and Björn Victor. The psi-calculi workbench: a generic tool for applied process calculi. ACM Transactions on Embedded Computing Systems, 14(1):9:1– 9:??, January 2015. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Bhuiyan:2018:EER

[BGS^{+}18] Ashikahmed Bhuiyan, Zhishan Guo, Abusayeed Saifullah, Nan Guan, and Haoyi Xiong. Energy-efficient realtime scheduling of DAG tasks. ACM Transactions on Embedded Computing Systems, 17(5):84:1–84:??, November 2018. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

                                        tronic). URL https://dl.
                                        acm.org/ft_gateway.cfm?
                                        id=3241049.

Berendsen:2011:FSA

[BGVZ11] Jasper Berendsen, Biniam Gebremichael, Frits W. Vaandrager, and Miaomiao Zhang. Formal specification and analysis of Zeroconf using Uppaal. ACM Transactions on Embedded Computing Systems, 10(3): 34:1–34:32, April 2011. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Bak:2015:SPD

[BHAC15] Stanley Bak, Zhenqi Huang, Fardin Abdi Taghi Abad, and Marco Caccamo. Safety and progress for distributed cyber-physical systems with unreliable communication. ACM Transactions on Embedded Computing Systems, 14(4):76:1–76:??, December 2015. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Banaiyanmofrad:2015:UFF

[BHD15] Abbas Banaiyanmofrad, Houman Homayoun, and Nikil Dutt.

Using a flexible faulttolerant cache to improve reliability for ultra low voltage operation. ACM Transactions on Embedded Computing Systems, 14(2):32:1– 32:??, March 2015. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Bulusu:2004:SCL

[BHET04] Nirupama Bulusu, John Heidemann, Deborah Estrin, and Tommy Tran. Selfconfiguring localization systems: Design and experimental evaluation. ACM Transactions on Embedded Computing Systems, 3(1): 24–60, February 2004. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Bresch:2020:TXP

[BHL^{+}20] Cyril Bresch, David Hély, Roman Lysecky, Stéphanie Chollet, and Ioannis Parissis. TrustFlow-X: a practical framework for fine-grained control-flow integrity in critical systems. ACM Transactions on Embedded Computing Systems, 19(5):36:1– 36:26, November 2020. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3398327.

Berard:2017:NIP

[BHM17] Béatrice Bérard, Loïc Hélouët, and John Mullins. Noninterference in partial order models. ACM Transactions on Embedded Computing Systems, 16(2):44:1– 44:??, April 2017. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Belson:2019:SAP

[BHXP19] Bruce Belson, Jason Holdsworth, Wei Xiang, and Bronson Philippa. A survey of asynchronous programming using coroutines in the Internet of Things and embedded systems. ACM Transactions on Embedded Computing Systems, 18(3):21:1–21:??, June 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

                                        /dl.acm.org/ft_gateway.
                                        cfm?id=3319618.

Bakshi:2023:CED

[BJ23] Suyash Bakshi and Lennart Johnsson. Computationally efficient DNN mapping search heuristic using deep reinforcement learning. ACM Transactions on Embedded Computing Systems, 22(5s):115:1–115:??, October 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3609110.

Bouraoui:2017:HAE

[BJCHA17] Hasna Bouraoui, Chadlia Jerad, Anupam Chattopadhyay, and Nejib Ben Hadj-Alouane. Hardware architectures for embedded speaker recognition applications: a survey. ACM Transactions on Embedded Computing Systems, 16(3): 78:1–78:??, July 2017. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Byun:2024:AMB

[BJH^{+}24] Hongsu Byun, Safdar Jamil, Jungwook Han, Sungyong Park, Myungcheol Lee, Changsoo Kim, Beongjun Choi, and Youngjae Kim. An analytical model-based capacity planning approach for building CSD-based storage systems. ACM Transactions on Embedded Computing Systems, 23(6):92:1– 92:??, November 2024. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3623677.

Bogdan:2013:PCH

[BJM13] Paul Bogdan, Siddharth Jain, and Radu Marculescu. Pacemaker control of heart rate variability: a cyber physical system perspective. ACM Transactions on Embedded Computing Systems, 12(1s):50:1–50:??, March 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Bagchi:2024:CCA

[BJP24] Aritra Bagchi, Dinesh Joshi, and Preeti Ranjan Panda. COBRRA: COntention-aware cache bypass with requestresponse arbitration. ACM Transactions on Embedded Computing Systems, 23(1): 12:1–12:??, January 2024. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3632748.

Boroujerdian:2023:FES

[BJT^{+}23] Behzad Boroujerdian, Ying Jing, Devashree Tripathy, Amit Kumar, Lavanya Subramanian, Luke Yen, Vincent Lee, Vivek Venkatesan, Amit Jindal, Robert Shearer, and Vijay Janapa Reddi. FARSI: an earlystage design space exploration framework to tame the domain-specific systemon-chip complexity. ACM Transactions on Embedded Computing Systems, 22(2): 31:1–31:??, March 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3544016.

Banerjee:2012:BAT

[BKMG12] Ayan Banerjee, Sailesh Kandula, Tridib Mukherjee, and Sandeep K. S. Gupta. BAND-AiDe: a tool for cyber-physical oriented analysis and design of body area networks and devices. ACM Transactions on Embedded Computing Systems, 11(S2):49:1–49:??, 2012. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Ali:2023:ESE

[BKS^{+}23] Ali J. Ben Ali, Marziye Kouroshli, Sofiya Semenova, Zakieh Sadat Hashemifar, Steven Y. Ko, and Karthik Dantu. Edge-SLAM: Edgeassisted visual simultaneous localization and mapping. ACM Transactions on Embedded Computing Systems, 22(1):18:1–18:??, January 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3561972.

Bishnoi:2015:BCC

[BLG^{+}15] Rimpy Bishnoi, Vijay Laxmi, Manoj Singh Gaur, José Flich, and Francisco Triviño. A brief comment on “A Complete Self-Testing and Self-Configuring NoC Infrastructure for Cost-Effective MPSoCs” [ACM Transactions on Embedded Computing Systems 12 (2013) Article 106]. ACM Transactions on Embedded Computing Systems, 14(1):2:1– 2:??, January 2015. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). See [GLT^{+}13].

Bhardwaj:2019:MCA

[BLSM19] Kartikeya Bhardwaj, Ching-Yi Lin, Anderson Sartor, and Radu Marculescu. Memoryand communication-aware model compression for distributed deep learning inference on IoT. ACM Transactions on Embedded Computing Systems, 18(5s):82:1– 82:??, October 2019. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

                                        tronic). URL https://dl.
                                        acm.org/ft_gateway.cfm?
                                        id=3358205.

Brisk:2013:ISI

[BM13] Philip Brisk and Tulika Mitra. Introduction to the special issue on applicationspecific processors. ACM Transactions on Embedded Computing Systems, 13(2): 15:1–15:??, September 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Bambagini:2016:EAS

[BMAB16] Mario Bambagini, Mauro Marinoni, Hakan Aydin, and Giorgio Buttazzo. Energyaware scheduling for realtime systems: a survey. ACM Transactions on Embedded Computing Systems, 15(1):7:1–7:??, February 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Bortolotti:2016:VRT

[BMB16] Daniele Bortolotti, Andrea Marongiu, and Luca Benini. VirtualSoC: a research tool for modern MPSoCs. ACM Transactions on Embedded Computing Systems, 16(1): 3:1–3:??, November 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Bellasi:2015:ERR

[BMF15] Patrick Bellasi, Giuseppe Massari, and William Fornaciari. Effective runtime resource management using Linux control groups with the BarbequeRTRM framework. ACM Transactions on Embedded Computing Systems, 14(2):39:1– 39:??, March 2015. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Berthier:2013:SPD

[BMM13] Nicolas Berthier, Florence Maraninchi, and Laurent Mounier. Synchronous programming of device drivers for global resource control in embedded operating systems. ACM Transactions on Embedded Computing Systems, 12(1s):39:1– 39:??, March 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Bur:2021:WCE

[BMMV21] Márton Búr, Kristóf Marussy, Brett H. Meyer, and Dániel Varró. Worst-case execution time calculation for querybased monitors by witness generation. ACM Transactions on Embedded Computing Systems, 20(6):107:1– 107:36, November 2021. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3471904.

Bartocci:2023:MHU

[BMNN23] Ezio Bartocci, Cristinel Mateis, Eleonora Nesterini, and Dejan Ničković. Mining hyperproperties using temporal logics. ACM Transactions on Embedded Computing Systems, 22(5s):156:1– 156:??, October 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3609394.

Benini:2003:EAD

[BMP03] Luca Benini, Alberto Macii, and Massimo Poncino. Energyaware design of embedded memories: a survey of technologies, architectures, and optimization techniques. ACM Transactions on Embedded Computing Systems, 2(1):5–32, February 2003. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Bhattacharjee:2023:XEA

[BMP23]  Abhiroop  Bhattacharjee,

Abhishek Moitra, and Priyadarshini Panda. XploreNAS: Explore adversarially robust and hardware-efficient neural architectures for nonideal xbars. ACM Transactions on Embedded Computing Systems, 22(4):62:1– 62:??, July 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3593045.

Biswas:2013:RTS

[BMS13] Swarnendu Biswas, Rajib Mall, and Manoranjan Satpathy. A regression test selection technique for embedded software. ACM Transactions on Embedded Computing Systems, 13(3): 47:1–47:??, December 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Benmaghnia:2024:CGN

[BMS24] Hanane Benmaghnia, Matthieu Martel, and Yassamine Seladji. Code generation for neural networks based on fixed-point arithmetic. ACM Transactions on Embedded Computing Systems, 23(5):68:1–68:??, September 2024. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3563945.

Bournoutian:2013:AAA

[BO13] Garo Bournoutian and Alex Orailoglu. Applicationaware adaptive cache architecture for power-sensitive mobile processors. ACM Transactions on Embedded Computing Systems, 13(3): 41:1–41:??, December 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Bouzidi:2022:PMC

[BONA22] Halima Bouzidi, Hamza Ouarnoughi, Smail Niar, and Abdessamad Ait El Cadi. Performance modeling of computer visionbased CNN on edge GPUs. ACM Transactions on Embedded Computing Systems, 21(5):64:1–64:??, September 2022. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3527169.

Bartolini:2005:OIC

[BP05] S. Bartolini and C. A. Prete. Optimizing instruction cache performance of embedded systems. ACM Transactions on Embedded Computing Systems, 4(4):934– 965, November 2005. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Blech:2012:GIB

[BP12] Jan Olaf Blech and Michaël Périn. Generating invariantbased certificates for embedded systems. ACM Transactions on Embedded Computing Systems, 11(2):34:1– 34:??, July 2012. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Bahirat:2014:MHP

[BP14] Shirish Bahirat and Sudeep Pasricha. METEOR: Hybrid photonic ring-mesh networkon-chip for multicore architectures. ACM Transactions on Embedded Computing Systems, 13(3s):116:1– 116:??, March 2014. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Brais:2019:AAM

[BP19] Hadi Brais and Preeti Ranjan Panda. Alleria: an advanced memory access profiling framework. ACM Transactions on Embedded Computing Systems, 18(5s):81:1– 81:??, October 2019. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3358193.

Bourke:2023:VCS

[BPP23] Timothy Bourke, Basile Pesin, and Marc Pouzet. Verified compilation of synchronous dataflow with state machines. ACM Transactions on Embedded Computing Systems, 22(5s):137:1– 137:??, October 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3608102.

Bhade:2024:LHB

[BPSS24] Pavitra Bhade, Joseph Paturel, Olivier Sentieys, and Sharad Sinha. Lightweight hardware-based cache sidechannel attack detection for edge devices (Edge-CaSCADe). ACM Transactions on Embedded Computing Systems, 23(4):56:1– 56:??, July 2024. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3663673.

Beretta:2016:PCA

[BRA^{+}16] Ivan Beretta, Vincenzo Rana, Abdulkadir Akin, Alessandro Antonio Nacci, Donatella Sciuto, and David Atienza. Parallelizing the chambolle algorithm for performance-optimized mapping on FPGA devices. ACM Transactions on Embedded Computing Systems, 15(3):

44:1–44:??, July 2016. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Butt:2016:DPH

[BRL16] Shahzad Ahmad Butt, Mehdi Roozmeh, and Luciano Lavagno. Designing parameterizable hardware IPs in a model-based design environment for high-level synthesis. ACM Transactions on Embedded Computing Systems, 15(2):32:1–32:??, May 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Broman:2021:IPM

[Bro21] David Broman. Interactive programmatic modeling. ACM Transactions on Embedded Computing Systems, 20(4):33:1–33:26, June 2021. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3431387.

Bataineh:2019:EDL

[BRR19] Omar Bataineh, David S.

Rosenblum, and Mark Reynolds. Efficient decentralized LTL monitoring framework using tableau technique. ACM Transactions on Embedded Computing Systems, 18(5s): 87:1–87:??, October 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

                                        /dl.acm.org/ft_gateway.
                                        cfm?id=3358219.

Bai:2013:SOS

[BS13a] Ke Bai and Aviral Shrivastava. A software-only scheme for managing heap data on limited local memory (LLM) multicore processors. ACM Transactions on Embedded Computing Systems, 13(1):5:1–5:??, August 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Bourke:2013:AES

[BS13b] Timothy Bourke and Arcot Sowmya. Analyzing an embedded sensor with timed automata in Uppaal. ACM Transactions on Embedded Computing Systems, 13(3):44:1–44:??, December 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Biswas:2022:PNC

[BS22] Arnab Kumar Biswas and Biplab Sikdar. Protecting network-on-chip intellectual property using timing channel fingerprinting. ACM Transactions on Embedded Computing Systems, 21(2): 17:1–17:21, March 2022. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3495565.

Bandari:2017:DBE

[BSA17] Maryam Bandari, Robert Simon, and Hakan Aydin. DMS-based energy optimizations for clustered WSNs. ACM Transactions on Embedded Computing Systems, 16(3):86:1–86:??, July 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Brandt:2014:PCS

[BSB14] Jens Brandt, Klaus Schneider, and Yu Bai. Passive code in synchronous programs. ACM Transactions on Embedded Computing Systems, 13(2s):67:1– 67:??, January 2014. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Boorghany:2015:CIL

[BSJ15] Ahmad Boorghany, Siavash Bayat Sarmadi, and Rasool Jalili. On constrained implementation of lattice-based cryptographic primitives and schemes on smart cards. ACM Transactions on Embedded Computing Systems, 14(3):42:1–42:??, April 2015. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Banerjee:2009:FPU

[BSKB^{+}09] Pritha Banerjee, Susmita Sur-Kolay, Arijit Bishnu, Sandip Das, Subhas C.

Nandy, and Subhasis Bhattacharjee. FPGA placement using space-filling curves: Theory meets practice. ACM Transactions on Embedded Computing Systems, 9(2): 12:1–12:??, October 2009. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Bombieri:2021:SIS

[BSM^{+}21] Nicola Bombieri, Silvia Scaffeo, Antonio Mastrandrea, Simone Caligola, Tommaso Carlucci, Franco Fummi, Carlo Laudanna, Gabriela Constantin, and Rosalba Giugno. SystemC implementation of stochastic Petri nets for simulation and parameterization of biological networks. ACM Transactions on Embedded Computing Systems, 20(4):31:1– 31:20, June 2021. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3427091.

Bujtor:2017:TPD

[BSV17] Ferenc Bujtor, Lev Sorokin, and Walter Vogler. Testing preorders for dMTS: Deadlock- and the new Deadlock-/Divergence Testing. ACM Transactions on Embedded Computing Systems, 16(2):41:1–41:??, April 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Baharani:2022:ARE

[BT22] Mohammadreza Baharani and Hamed Tabkhi. ATCN: Resource-efficient processing of time series on edge. ACM Transactions on Embedded Computing Systems, 21(5):61:1–61:??, September 2022. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3524070.

Bhat:2019:ULE

[BTA^{+}19] Ganapati Bhat, Yigit Tuncel, Sizhe An, Hyung Gyu Lee, and Umit Y. Ogras. An ultra-low energy human activity recognition accelerator for wearable health applications. ACM Transactions on Embedded Computing Systems, 18(5s):49:1– 49:??, October 2019. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3358175.

Balkan:2018:UFA

[BTD^{+}18] Ayca Balkan, Paulo Tabuada, Jyotirmoy V. Deshmukh, Xiaoqing Jin, and James Kapinski. Underminer: a framework for automatically identifying nonconverging behaviors in black-box system models. ACM Transactions on Embedded Computing Systems, 17(1):20:1– 20:??, January 2018. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Boulis:2012:IWC

[BTL^{+}12] Athanassios Boulis, Yuriy Tselishchev, Lavy Libman, David Smith, and Leif Hanlen. Impact of wireless channel temporal variation on MAC design for body area networks. ACM Transactions on Embedded Computing Systems, 11(S2): 51:1–51:??, 2012. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Burns:2005:E

[Bur05] Alan Burns. Editorial. ACM Transactions on Embedded Computing Systems, 4(3): 469–471, August 2005. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Bujtor:2015:FSM

[BV15] Ferenc Bujtor and Walter Vogler. Failure semantics for modal transition systems. ACM Transactions on Embedded Computing Systems, 14(4):67:1–67:??, December 2015. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Beg:2013:CPA

[BvB13] Mirza Beg and Peter van Beek. A constraint programming approach for integrated spatial and temporal scheduling for clustered architectures. ACM Transactions on Embedded Computing Systems, 13(1):14:1– 14:??, August 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Basanta-Val:2010:NHR

[BVGVEA10] Pablo Basanta-Val, Marisol García-Valls, and Iria Estévez-Ayres. No-Heap Remote Objects for distributed realtime Java. ACM Transactions on Embedded Computing Systems, 10(1):7:1– 7:??, August 2010. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Bajczi:2019:WMP

[BVM19] Levente Bajczi, András Vörös, and Vince Molnár. Will my program break on this faulty processor?: Formal analysis of hardware fault activations in concurrent embedded software. ACM Transactions on Embedded Computing Systems, 18(5s):89:1–89:??, October 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/ft_gateway.
cfm?id=3358238.

Balani:2014:DPF

[BWS14] Rahul Balani, Lucas F. Wanner, and Mani B. Srivastava. Distributed programming framework for fast iterative optimization in networked cyber-physical systems. ACM Transactions on Embedded Computing Systems, 13(2s):66:1–66:??, January 2014. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Bai:2009:MME

[BYD09] Lan S. Bai, Lei Yang, and Robert P. Dick. MEMMU: Memory expansion for MMUless embedded systems. ACM Transactions on Embedded Computing Systems, 8(3):23:1–23:??, April 2009. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Bruns:2021:TMC

[BYIG21] Friederike Bruns, Irune Yarza, Philipp Ittershagen, and Kim Grüttner. Time measurement and control blocks for bare-metal C++ applications. ACM Transactions on Embedded Computing Systems, 20(4):34:1– 34:26, June 2021. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3434401.

Beldianu:2013:MBV

[BZ13] Spiridon F. Beldianu and Sotirios G. Ziavras. Multicorebased vector coprocessor sharing for performance and energy gains. ACM Transactions on Embedded Computing Systems, 13(2):17:1– 17:??, September 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Barijough:2019:QLA

[BZG19] Kamyar Mirzazad Barijough, Zhuoran Zhao, and Andreas Gerstlauer. Quality/latency-aware real-time scheduling of distributed streaming IoT applications. ACM Transactions on Embedded Computing Systems, 18(5s):83:1–83:??, October 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/ft_gateway.
cfm?id=3358209.

Brumar:2023:EDA

[BZY^{+}23] Iulian Brumar, Georgios Zacharopoulos, Yuan Yao, Saketh Rama, David Brooks, and Gu-Yeon Wei. Early DSE and automatic generation of coarse-grained merged accelerators. ACM Transactions on Embedded Computing Systems, 22(2): 32:1–32:??, March 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3546070.

Cucinotta:2024:MCO

[CAA^{+}24] Tommaso Cucinotta, Alexandre Amory, Gabriele Ara, Francesco Paladino, and Marco Di Natale. Multicriteria optimization of realtime DAGs on heterogeneous platforms under P-EDF. ACM Transactions on Embedded Computing Systems, 23(1):3:1–3:??, January 2024. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL

https://dl.acm.org/doi/
10.1145/3592609.

Caronti:2023:FGH

[CAN^{+}23] Luca Caronti, Khakim Akhunov, Matteo Nardello, Kasim Sinan Yildirim, and Davide Brunelli. Fine-grained hardware acceleration for efficient batteryless intermittent inference on the edge. ACM Transactions on Embedded Computing Systems, 22(5): 82:1–82:??, September 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3608475.

Carta:2007:CTA

[CAP^{+}07] Salvatore Carta, Andrea Alimonda, Alessandro Pisano, Andrea Acquaviva, and Luca Benini. A control theoretic approach to energyefficient pipelined computation in MPSoCs. ACM Transactions on Embedded Computing Systems, 6(4): 27:1–27:??, September 2007. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Catania:2015:PSR

[CAP15] Vincenzo Catania, Andrea Araldo, and Davide Patti. Parameter space representation of Pareto front to explore hardware–software dependencies. ACM Transactions on Embedded Computing Systems, 14(4):77:1– 77:??, December 2015. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Cucinotta:2011:RMA

[CAPL11] Tommaso Cucinotta, Luca Abeni, Luigi Palopoli, and Giuseppe Lipari. A robust mechanism for adaptive scheduling of multimedia applications. ACM Transactions on Embedded Computing Systems, 10(4): 46:1–46:??, November 2011. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Chang:2022:ISIa

[CBH22a] Yuan-Hao Chang, Jalil Boukhobza, and Song Han. Introduction to the special issue on memory and storage systems for embedded and IoT applications. ACM Transactions on Embedded Computing Systems, 21(1): 1:1–1:4, January 2022. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3505283.

Chang:2022:ISIb

[CBH22b] Yuan-Hao Chang, Jalil Boukhobza, and Song Han. Introduction to the special issue on memory and storage systems for embedded and IoT applications: Part 2. ACM Transactions on Embedded Computing Systems, 21(3):21:1–21:2, May 2022. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3531707.

Chattopadhyay:2019:QIL

[CBRZ19] Sudipta Chattopadhyay, Moritz Beck, Ahmed Rezine, and Andreas Zeller. Quantifying the information leakage in cache attacks via symbolic execution. ACM Transactions on Embedded Computing Systems, 18(1):7:1– 7:??, February 2019. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3288758.

Cherif:2019:LSD

[CBS19] Amina Cherif, Malika Belkadi, and Damien Sauveron. A lightweight and secure data collection serverless protocol demonstrated in an active RFIDs scenario. ACM Transactions on Embedded Computing Systems, 18(3): 27:1–27:??, June 2019. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3274667.

Che:2013:SSD

[CC13a] Weijia Che and Karam S.

Chatha. Scheduling of synchronous data flow models onto scratchpad memorybased embedded processors. ACM Transactions on Embedded Computing Systems, 13(1s):30:1–30:??, November 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Collins:2013:FFS

[CC13b] Rebecca L. Collins and Luca P. Carloni. Flexible filters in stream programs. ACM Transactions on Embedded Computing Systems, 13(3):45:1–45:??, December 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Chang:2014:ISS

[CC14] Naehyuck Chang and Jian-Jia Chen. Introduction to the special section on ES-TIMedia’11. ACM Transactions on Embedded Computing Systems, 13(2s):58:1– 58:??, January 2014. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Canis:2013:LOS

[CCA^{+}13] Andrew Canis, Jongsok Choi, Mark Aldham, Victor Zhang, Ahmed Kammoona, Tomasz Czajkowski, Stephen D. Brown, and Jason H. Anderson. LegUp: an open-source high-level synthesis tool for FPGAbased processor/accelerator systems. ACM Transactions on Embedded Computing Systems, 13(2):24:1– 24:??, September 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Cucinotta:2012:ART

[CCAP12] Tommaso Cucinotta, Fabio Checconi, Luca Abeni, and Luigi Palopoli. Adaptive real-time scheduling for legacy multimedia applications. ACM Transactions on Embedded Computing Systems, 11(4):86:1– 86:??, December 2012. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Coussy:2006:FMH

[CCB^{+}06] Philippe Coussy, Emmanuel Casseau, Pierre Bomel, Adel Baganne, and Eric Martin. A formal method for hardware IP design and integration under I/O and timing constraints. ACM Transactions on Embedded Computing Systems, 5(1):29–53, February 2006. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Casu:2014:UMI

[CCC^{+}14] Mario R. Casu, Francesco Colonna, Marco Crepaldi, Danilo Demarchi, Mariagrazia Graziano, and Maurizio Zamboni. UWB microwave imaging for breast cancer detection: Manycore, GPU, or FPGA? ACM Transactions on Embedded Computing Systems, 13(3s): 109:1–109:??, March 2014. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Chen:2017:SUE

[CCC^{+}17] Tseng-Yi Chen, Yuan-Hao Chang, Shuo-Han Chen, Nien-I Hsu, Hsin-Wen Wei, and Wei-Kuan Shih. On space utilization enhancement of file systems for embedded storage systems. ACM Transactions on Embedded Computing Systems, 16(3):83:1–83:??, July 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Chuang:2020:DDB

[CCC^{+}20] Yi-Jing Chuang, Shuo-Han Chen, Yuan-Hao Chang, Yu-Pei Liang, Hsin-Wen Wei, and Wei-Kuan Shih. DSTL: a demand-based shingled translation layer for enabling adaptive address mapping on SMR drives. ACM Transactions on Embedded Computing Systems, 19(4):25:1– 25:21, July 2020. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3391892.

Chang:2013:ALC

[CCH13] Li-Pin Chang, Tung-Yang Chou, and Li-Chun Huang. An adaptive, low-cost wearleveling algorithm for multichannel solid-state disks. ACM Transactions on Embedded Computing Systems, 13(3):55:1–55:??, December 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Chatterjee:2016:TAD

[CCKM16] Urbi Chatterjee, Rajat Subhra Chakraborty, Hitesh Kapoor, and Debdeep Mukhopadhyay. Theory and application of delay constraints in arbiter PUF. ACM Transactions on Embedded Computing Systems, 15(1):10:1– 10:??, February 2016. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Chatterjee:2017:PBS

[CCM17] Urbi Chatterjee, Rajat Subhra Chakraborty, and Debdeep Mukhopadhyay. A PUFbased secure communication protocol for IoT. ACM Transactions on Embedded Computing Systems, 16(3): 67:1–67:??, July 2017. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Chen:2019:OIW

[CCP^{+}19] Yu-Chieh Chen, Ching-Chih Chang, Ramesh Perumal, Shih-Rung Yeh, Yen-Chung Chang, and Hsin Chen. Optimization and implementation of wavelet-based algorithms for detecting highvoltage spindles in neuron signals. ACM Transactions on Embedded Computing Systems, 18(5):39:1– 39:??, October 2019. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3329864.

Chattopadhyay:2014:UWA

[CCR^{+}14] Sudipta Chattopadhyay, Lee Kee Chong, Abhik Roychoudhury, Timon Kelter, Peter Marwedel, and Heiko Falk. A unified WCET analysis framework for multicore platforms. ACM Transactions on Embedded Computing Systems, 13(4s):124:1– 124:??, April 2014. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Chang:2023:LEL

[CCSC23] Jung-Hsiu Chang, Tzu-Yu Chang, Yi-Chao Shih, and Tseng-Yi Chen. LaDy: Enabling locality-aware deduplication technology on shingled magnetic recording drives. ACM Transactions on Embedded Computing Systems, 22(5s):127:1– 127:??, October 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3607921.

Chao:2013:TLA

[CCY^{+}13] Chih-Hao Chao, Kun-Chih Chen, Tsu-Chu Yin, Shu-Yen Lin, and An-Yeu (Andy) Wu. Transport-layer-assisted routing for runtime thermal management of 3D NoC systems. ACM Transactions on Embedded Computing Systems, 13(1):11:1– 11:??, August 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Coronato:2010:FSW

[CD10] Antonio Coronato and Giuseppe De Pietro. Formal specification of wireless and pervasive healthcare applications. ACM Transactions on Embedded Computing Systems, 10(1):12:1–12:??, August 2010. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Chandraiah:2012:CAR

[CD12] Pramod Chandraiah and Rainer Dömer. Computeraided recoding to create structured and analyzable system models. ACM Transactions on Embedded Computing Systems, 11S(1):23:1– 23:??, 2012. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Chaki:2017:FVT

[CD17] Sagar Chaki and Dionisio De Niz. Formal verification of a timing enforcer implementation. ACM Transactions on Embedded Computing Systems, 16(5s):168:1– 168:??, October 2017. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Cheng:2019:AVE

[CD19] Zhongqi Cheng and Rainer Dömer. Analyzing variable entanglement for parallel simulation of SystemC TLM-2.0 models. ACM Transactions on Embedded Computing Systems, 18(5s):79:1– 79:??, October 2019. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3358194.

Casini:2024:ISI

[CDBB24] Daniel Casini, Dakshina Dasari, Matthias Becker, and Giorgio Buttazzo. Introduction to the special issue on real-time computing in the IoT-to-edge-to-cloud continuum. ACM Transactions on Embedded Computing Systems, 23(1):1:1– 1:??, January 2024. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3605180.

Chanet:2007:ARM

[CDD^{+}07] Dominique Chanet, Bjorn De Sutter, Bruno De Bus, Ludo Van Put, and Koen De Bosschere. Automated reduction of the memory footprint of the Linux kernel. ACM Transactions on Embedded Computing Systems, 6(4):23:1–23:??, September 2007. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Corre:2016:TTB

[CDH^{+}16] Youenn Corre, Jean-Philippe Diguet, Dominique Heller, Dominique Blouin, and Loïc Lagadec. TBES: Templatebased exploration and synthesis of heterogeneous multiprocessor architectures on FPGA. ACM Transactions on Embedded Computing Systems, 15(1):9:1– 9:??, February 2016. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Chen:2019:DAS

[CDX^{+}19] Zhengguo Chen, Quan Deng, Nong Xiao, Kirk Pruhs, and Youtao Zhang. DW-MAcc: Accelerating shiftbased CNNs with domain wall memories. ACM Transactions on Embedded Computing Systems, 18(5s):69:1– 69:??, October 2019. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3358199.

Clair:2023:SED

[CEC23] Judicael Clair, Guy Eichler, and Luca P. Carloni. Spike-Hard: Efficiency-driven neuromorphic hardware for heterogeneous systems-on-chip. ACM Transactions on Embedded Computing Systems, 22(5s):106:1–106:??, October 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3609101.

Cilardo:2015:ECA

[CFGM15] Alessandro Cilardo, Edoardo Fusella, Luca Gallo, and Antonino Mazzeo. Exploiting concurrency for the automated synthesis of MP-SoC interconnects. ACM Transactions on Embedded Computing Systems, 14(3): 57:1–57:??, May 2015. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Choo:2017:EDF

[CFXY17] Kim-Kwang Raymond Choo, Yunsi Fei, Yang Xiang, and Yu Yu. Embedded device forensics and security. ACM Transactions on Embedded Computing Systems, 16(2):50:1–50:??, April 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Chang:2025:CGT

[CGM^{+}25] Liangliang Chang, Serhan Gener, Joshua Mack, Hasan Umut Suluhan, Ali Akoglu, and Chaitali Chakrabarti. Coarse-grained task parallelization by dynamic profiling for heterogeneous SoCbased embedded system. ACM Transactions on Embedded Computing Systems, 24(1):18:1–18:??, January 2025. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3704635.

Castro-Godinez:2019:EBE

[CGSH19] Jorge Castro-Godínez, Muhammad Shafique, and Jörg Henkel. ECAx: Balancing error correction costs in approximate accelerators. ACM Transactions on Embedded Computing Systems, 18(5s):48:1–48:??, October 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/ft_gateway.
cfm?id=3358179.

Courbot:2010:EBD

[CGV10] Alexandre Courbot, Gilles Grimaud, and Jean-Jacques Vandewalle. Efficient offboard deployment and customization of virtual machinebased embedded systems. ACM Transactions on Embedded Computing Systems, 9(3):21:1–21:??, February 2010. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Chen:2018:GEA

[CGZ18] Jiming Chen, Yu (Jason)

Gu, and Gil Zussman. Guest editorial for ACM TECS: Special issue on autonomous battery-free sensing and communication. ACM Transactions on Embedded Computing Systems, 17(1): 2:1–2:??, January 2018. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Choi:2008:SHM

[CH08] Yoonseo Choi and Hwansoo Han. Shared heap management for memory-limited Java virtual machines. ACM Transactions on Embedded Computing Systems, 7(2): 13:1–13:??, February 2008. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Cesana:2010:MBM

[CH10] Ulpian Cesana and Zhen He. Multi-buffer manager: Energy-efficient buffer manager for databases on flash memory. ACM Transactions on Embedded Computing Systems, 9(3):28:1– 28:??, February 2010. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Chang:2013:RDD

[CHCC13] Yu-Ming Chang, Pi-Cheng Hsiu, Yuan-Hao Chang, and Che-Wei Chang. A resourcedriven DVFS scheme for smart handheld devices. ACM Transactions on Embedded Computing Systems, 13(3):53:1–53:??, December 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Chen:2022:SRT

[CHJ22] Cong Chen, Zhong Hong, and Jian-Min Jiang. Scheduling in real-time mobile systems. ACM Transactions on Embedded Computing Systems, 21(3):34:1–34:36, May 2022. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3517747.

Chakraborty:2014:MCH

[CHK^{+}14a] Arup Chakraborty, Houman Homayoun, Amin Khajeh, Nikil Dutt, Ahmed Eltawil, and Fadi Kurdahi. Multicopy cache: a highly energy-efficient cache architecture. ACM Transactions on Embedded Computing Systems, 13(5s):150:1– 150:??, September 2014. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Chen:2014:EOR

[CHK14b] Gang Chen, Kai Huang, and Alois Knoll. Energy optimization for real-time multiprocessor system-on-chip with optimal DVFS and DPM combination. ACM Transactions on Embedded Computing Systems, 13(3s): 111:1–111:??, March 2014. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Cooke:2015:FSM

[CHS15] Patrick Cooke, Lu Hao, and Greg Stitt. Finite-state-machine overlay architectures for fast FPGA compilation and application portability. ACM Transactions on Embedded Computing Systems, 14(3):54:1–54:??, April 2015. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Crenshaw:2007:RIE

[CHTC07] Tanya L. Crenshaw, Spencer Hoke, Ajay Tirumala, and Marco Caccamo. Robust implicit EDF: a wireless MAC protocol for collaborative real-time systems. ACM Transactions on Embedded Computing Systems, 6(4): 28:1–28:??, September 2007. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Ciszewski:2017:EAC

[CI17] Michal Ciszewski and Konrad Iwanicki. Efficient automated code partitioning for microcontrollers with switchable memory banks. ACM Transactions on Embedded Computing Systems, 16(4): 114:1–114:??, August 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Chattopadhyay:2008:PPA

[CIC^{+}08] A. Chattopadhyay, H. Ishebabi, X. Chen, Z. Rakosi, K. Karuri, D. Kammler, R. Leupers, G. Ascheid, and H. Meyr. Prefabrication and postfabrication architecture exploration for partially reconfigurable VLIW processors. ACM Transactions on Embedded Computing Systems, 7(4):40:1–40:??, July 2008. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Chattopadhyay:2009:PPA

[CIC^{+}09] A. Chattopadhyay, H. Ishebabi, X. Chen, Z. Rakosi, K. Karuri, D. Kammler, R. Leupers, G. Ascheid, and H. Meyr. Pre- and postfabrication architecture exploration for partially reconfigurable VLIW processors. ACM Transactions on Embedded Computing Systems, 8(2):12:1–12:??, January 2009. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Cedersjo:2020:TFC

[CJ20] Gustav Cedersjö and Jörn W.

Janneck. Tÿcho: a framework for compiling stream programs. ACM Transactions on Embedded Computing Systems, 18(6):1–25, January 2020. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL

https://dl.acm.org/doi/
abs/10.1145/3362692.

Chen:2017:GEA

[CJL17] Bo-Wei Chen, Wen Ji, and Zhu Li. Guest editorial for ACM TECS special issue on effective divide-and-conquer, incremental, or distributed mechanisms of embedded designs for extremely big data in large-scale devices. ACM Transactions on Embedded Computing Systems, 16(3):72:1–72:??, July 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Chouali:2005:PPM

[CJMB05] S. Chouali, J. Julliand, P.- A. Masson, and F. Bellegarde. PLTL-partitioned model checking for reactive systems under fairness assumptions. ACM Transactions on Embedded Computing Systems, 4(2):267– 301, May 2005. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Chen:2023:TTR

[CK23] Huili Chen and Farinaz Koushanfar. Tutorial: Toward robust deep learning against poisoning attacks. ACM Transactions on Embedded Computing Systems, 22(3):42:1–42:??, May 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3574159.

Crites:2017:DCE

[CKB17] Brian Crites, Karen Kong, and Philip Brisk. Diagonal component expansion for flow-layer placement of flowbased microfluidic biochips. ACM Transactions on Embedded Computing Systems, 16(5s):126:1–126:??, October 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Chang:2014:ISI

[CKGN14] Li-Pin Chang, Tei-Wei Kuo, Chris Gill, and Jin Nakazawa. Introduction to the special issue on realtime, embedded and cyberphysical systems. ACM Transactions on Embedded Computing Systems, 13(5s): 155:1–155:??, September 2014. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Chen:2006:RCS

[CKIR06] G. Chen, M. Kandemir, M. J. Irwin, and J. Ramanujam. Reducing code size through address register assignment. ACM Transactions on Embedded Computing Systems, 5(1):225– 258, February 2006. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Chang:2004:RTG

[CKL04] Li-Pin Chang, Tei-Wei Kuo, and Shi-Wu Lo. Realtime garbage collection for flash-memory storage systems of real-time embedded systems. ACM Transactions on Embedded Computing Systems, 3(4):837– 863, November 2004. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Calderon:2020:GUE

[CKN^{+}20] Alejandro J. Calderón, Leonidas Kosmidis, Carlos F. Nicolás, Francisco J. Cazorla, and Peio Onaindia. GMAI: Understanding and exploiting the internals of GPU resource allocation in critical systems. ACM Transactions on Embedded Computing Systems, 19(5):34:1– 34:23, November 2020. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3391896.

Chen:2013:CMS

[CL13] Yufeng Chen and Gaiyun Liu. Computation of minimal siphons in Petri nets by using binary decision diagrams. ACM Transactions on Embedded Computing Systems, 12(1):3:1– 3:??, January 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Chen:2019:TAF

[CLJ^{+}19] Peng Chen, Weichen Liu, Xu Jiang, Qingqiang He, and Nan Guan. Timing-anomaly free dynamic scheduling of conditional DAG tasks on multi-core systems. ACM Transactions on Embedded Computing Systems, 18(5s): 91:1–91:??, October 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/ft_gateway.
cfm?id=3358236.

Chung:2013:EUE

[CLK13] Yi-Fan Chung, Yin-Tsung Lo, and Chung-Ta King. Enhancing user experiences by exploiting energy and launch delay trade-off of mobile multimedia applications. ACM Transactions on Embedded Computing Systems, 12(1s):37:1–37:??, March 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Chang:2016:SGA

[CLL16] Li-Pin Chang, Yu-Syun Liu, and Wen-Huei Lin. Stable greedy: Adaptive garbage collection for durable pagemapping multichannel SSDs. ACM Transactions on Embedded Computing Systems, 15(1):13:1–13:??, February 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Chen:2018:HEW

[CLL^{+}18] Qi Chen, Ye Liu, Guangchi Liu, Qing Yang, Xianming Shi, Hongwei Gao, Lu Su, and Quanlong Li. Harvest energy from the water: a self-sustained wireless water quality sensing system. ACM Transactions on Embedded Computing Systems, 17(1): 3:1–3:??, January 2018. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Carreon:2021:PET

[CLL21] Nadir A. Carreon, Sixing Lu, and Roman Lysecky. Probabilistic estimation of threat intrusion in embedded systems for runtime detection. ACM Transactions on Embedded Computing Systems, 20(2):14:1– 14:27, March 2021. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3432590.

Chen:2017:RTD

[CLLC17] Xiaowen Chen, Zhonghai Lu, Sheng Liu, and Shuming Chen. Round-trip DRAM access fairness in 3D NoCbased many-core systems. ACM Transactions on Embedded Computing Systems, 16(5s):162:1–162:??, October 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Corliss:2005:IED

[CLR05] Marc L. Corliss, E. Christopher Lewis, and Amir Roth. The implementation and evaluation of dynamic code decompression using DISE. ACM Transactions on Embedded Computing Systems, 4(1):38–72, February 2005. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Chiew:2016:NEI

[CLS16] Wei Ming Chiew, Feng Lin, and Hock Soon Seah. A novel embedded interpolation algorithm with negative squared distance for realtime endomicroscopy. ACM Transactions on Embedded Computing Systems, 15(4): 75:1–75:??, August 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Cheng:2020:DDT

[CLW^{+}20] Yuan Cheng, Guangya Li, Ngai Wong, Hai-Bao Chen, and Hao Yu. DEEPEYE: a deeply tensor-compressed neural network for video comprehension on terminal devices. ACM Transactions on Embedded Computing Systems, 19(3):18:1– 18:25, July 2020. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3381805.

Cachera:2005:VSP

[CMA05] David Cachera and Katell Morin-Allory. Verification of safety properties for parameterized regular systems. ACM Transactions on Embedded Computing Systems, 4(2):228–266, May 2005. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Chielle:2024:CBM

[CMGM24] Eduardo Chielle, Oleg Mazonka, Homer Gamil, and Michail Maniatakos. Coupling bit and modular arithmetic for efficient generalpurpose fully homomorphic encryption. ACM Transactions on Embedded Computing Systems, 23(4):57:1– 57:??, July 2024. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3665280.

Chen:2012:ART

[CMK12] Mingsong Chen, Prabhat Mishra, and Dhrubajyoti Kalita. Automatic RTL test generation from SystemC TLM specifications. ACM Transactions on Embedded Computing Systems, 11(2): 38:1–38:??, July 2012. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Cilardo:2024:ASC

[CMMD24] Alessandro Cilardo, Vincenzo Maisto, Nicola Mazzocca, and Franca Rocco Di Torrepadula. An approach to the systematic characterization of multitask accelerated CNN inference in edge MP-SoCs. ACM Transactions on Embedded Computing Systems, 23(3):48:1–48:??, May 2024. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3611015.

Contreras:2007:XPP

[CMP^{+}07] Gilberto Contreras, Margaret Martonosi, Jinzhang Peng, Guei-Yuan Lueh, and Roy Ju. The XTREM power and performance simulator for the Intel XScale core: Design and experiences. ACM Transactions on Embedded Computing Systems, 6(1): 4:1–4:??, February 2007. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Cunha:2017:DSC

[CMP17] Marcos Aurélio Pinto Cunha, Omayma Matoussi, and Frédéric Pétrot. Detecting software cache coherence violations in MPSoC using traces captured on virtual platforms. ACM Transactions on Embedded Computing Systems, 16(2):30:1– 30:??, April 2017. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Cleaveland:2023:FVN

[CMP23] Rachel Cleaveland, Stefan Mitsch, and André Platzer. Formally verified next-generation airborne collision avoidance games in ACAS X. ACM Transactions on Embedded Computing Systems, 22(1):10:1– 10:??, January 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3544970.

Colaco:2023:CSB

[CMPP23] Jean-Louis Colaço, Michael Mendler, Baptiste Pauget, and Marc Pouzet. A constructive state-based semantics and interpreter for a synchronous data-flow language with state machines. ACM Transactions on Embedded Computing Systems, 22(5s):152:1–152:??, October 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3609131.

Cho:2008:DNP

[CMS08] Young H. Cho and William H. Mangione-Smith. Deep network packet filter design for reconfigurable devices. ACM Transactions on Embedded Computing Systems, 7(2):21:1–21:??, February 2008. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Chen:2017:CRA

[CMS17] Xin Chen, Sergio Mover, and Sriram Sankaranarayanan. Compositional relational abstraction for nonlinear hybrid systems. ACM Transactions on Embedded Computing Systems, 16(5s):187:1– 187:??, October 2017. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Cabodi:2010:BSF

[CMV10] Gianpiero Cabodi, Marco Murciano, and Massimo Violante. Boosting software fault injection for dependability analysis of realtime embedded applications. ACM Transactions on Embedded Computing Systems, 10(2):24:1–24:??, December 2010. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Chi:2013:WNE

[CNC13] Yu M. Chi, Patrick Ng, and Gert Cauwenberghs. Wireless noncontact ECG and EEG biopotential sensors. ACM Transactions on Embedded Computing Systems, 12(4):103:1–103:??, June 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Coron:2004:SSL

[CNK04] Jean-Sebastien Coron, David Naccache, and Paul Kocher. Statistics and secret leakage. ACM Transactions on Embedded Computing Systems, 3(3):492–508, August 2004. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Chen:2022:PRA

[COC22] Xing Chen, Umit Ogras, and Chaitali Chakrabarti. Probabilistic risk-aware scheduling with deadline constraint for heterogeneous SoCs. ACM Transactions on Embedded Computing Systems, 21(2):15:1–15:27, March 2022. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3489409.

Cardo:2013:ISS

[CP13a] José Flich Cardo and Maurizio Palesi. Introduction to the special section on onchip and off-chip network architectures. ACM Transactions on Embedded Computing Systems, 12(4):104:1– 104:??, June 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Chen:2013:ISS

[CP13b] Jian-Jia Chen and Maurizio Palesi. Introduction to the special section on ES-TIMedia’12. ACM Transactions on Embedded Computing Systems, 12(1s):32:1– 32:??, March 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Chatterjee:2017:FTD

[CPC17] Navonil Chatterjee, Suraj Paul, and Santanu Chattopadhyay. Fault-tolerant dynamic task mapping and scheduling for network-on-chip-based multicore platform. ACM Transactions on Embedded Computing Systems, 16(4):108:1–108:??, August 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Castiglione:2017:BFI

[CPP^{+}17] Arcangelo Castiglione, Raffaele Pizzolante, Francesco Palmieri, Barbara Masucci, Bruno Carpentieri, Alfredo De Santis, and Aniello Castiglione. On-board formatindependent security of functional magnetic resonance images. ACM Transactions on Embedded Computing Systems, 16(2):56:1– 56:??, April 2017. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Cotard:2015:SHR

[CQB^{+}15] Sylvain Cotard, Audrey Queudet, Jean-Luc Béchennec, Sébastien Faucou, and Yvon Trinquet. STM–HRT: a robust and wait-free STM for hard real-time multicore embedded systems. ACM Transactions on Embedded Computing Systems, 14(4): 66:1–66:??, December 2015. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Cazorla:2013:PPA

[CQV^{+}13] Francisco J. Cazorla, Eduardo Quiñones, Tullio Vardanega, Liliana Cucu, Benoit Triquet, Guillem Bernat, Emery Berger, Jaume Abella, Franck Wartel, Michael Houston, Luca Santinelli, Leonidas Kosmidis, Code Lo, and Dorin Maxim. PROARTIS: Probabilistically analyzable real-time systems. ACM Transactions on Embedded Computing Systems, 12(2s):94:1– 94:??, May 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Chattopadhyay:2014:CRP

[CR14] Sudipta Chattopadhyay and Abhik Roychoudhury. Cacherelated preemption delay analysis for multilevel noninclusive caches. ACM Transactions on Embedded Computing Systems, 13(5s): 147:1–147:??, September 2014. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Curley:2010:RDT

[CRAJ10] Edward Curley, Binoy Ravindran, Jonathan Anderson, and E. Douglas Jensen. Recovering from distributable thread failures in distributed real-time Java. ACM Transactions on Embedded Computing Systems, 10(1):8:1– 8:??, August 2010. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Chippa:2013:MQV

[CRCR13] Vinay K. Chippa, Kaushik Roy, Srimat T. Chakradhar, and Anand Raghunathan. Managing the quality vs. efficiency trade-off using dynamic effort scaling. ACM Transactions on Embedded Computing Systems, 12(2s):90:1–90:??, May 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Cho:2010:LFS

[CRJ10] Hyeonjoong Cho, Binoy Ravindran, and E. Douglas Jensen. Lock-free synchronization for dynamic embedded real-time systems. ACM Transactions on Embedded Computing Systems, 9(3):23:1–23:??, February 2010. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Clemente:2014:AMR

[CRM14] Juan Antonio Clemente, Javier Resano, and Daniel Mozos. An approach to manage reconfigurations and reduce area cost in hard realtime reconfigurable systems. ACM Transactions on Embedded Computing Systems, 13(4):90:1–90:??, February 2014. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Castrillon:2016:GES

[CS16] Jeronimo Castrillon and Cristina Silvano. Guest editorial: Special issue on Virtual Prototyping of Parallel and Embedded Systems (ViPES). ACM Transactions on Embedded Computing Systems, 16(1):2:1– 2:??, November 2016. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Choi:2022:ECA

[CS22] Kyubaik Choi and Gerald E. Sobelman. An efficient CNN accelerator for low-cost edge systems. ACM Transactions on Embedded Computing Systems, 21(4):44:1– 44:??, July 2022. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3539224.

Chou:2023:RSK

[CSC23] Yi-Quan Chou, Lin-Wei Shen, and Li-Pin Chang. Rectifying skewed kernel page reclamation in mobile devices for improving userperceivable latency. ACM Transactions on Embedded Computing Systems, 22(5s): 140:1–140:??, October 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3607937.

Chang:2017:ESS

[CSCC17] Li-Pin Chang, Po-Han Sung, Po-Tsang Chen, and Po-Hung Chen. Eager synching: a selective logging strategy for fast fsync() on flash-based Android devices. ACM Transactions on Embedded Computing Systems, 16(2):34:1–34:??, April 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Chen:2022:ERD

[CSH^{+}22] Kuan-Hsun Chen, Chiahui Su, Christian Hakert, Sebastian Buschjäger, Chao-Lin Lee, Jenq-Kuen Lee, Katharina Morik, and Jian-Jia Chen. Efficient realization of decision trees for real-time inference. ACM Transactions on Embedded Computing Systems, 21(6): 68:1–68:??, November 2022. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3508019.

Chen:2002:TGC

[CSK^{+}02] G. Chen, R. Shetty, M. Kandemir, N. Vijaykrishnan, M. J. Irwin, and M. Wolczko. Tuning garbage collection for reducing memory system energy in an embedded Java environment. ACM Transactions on Embedded Computing Systems, 1(1):27–55, November 2002. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Chakraborty:2024:TTA

[CSM24] Shounak Chakraborty, Yanshul Sharma, and Sanjay Moulik. TREAFET: Temperature-aware realtime task scheduling for Fin-FET based multicores. ACM Transactions on Embedded Computing Systems, 23(4): 61:1–61:??, July 2024. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3665276.

Caspi:2008:SPM

[CSST08] Paul Caspi, Norman Scaife, Christos Sofronis, and Stavros Tripakis. Semantics-preserving multitask implementation of synchronous programs. ACM Transactions on Embedded Computing Systems, 7(2): 15:1–15:??, February 2008. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Caspi:2005:GGC

[CSVA^{+}05] P. Caspi, A. Sangiovanni-Vincentelli, L. Almeida, A. Benveniste, B. Bouyssounouse, G. Buttazzo, I. Crnkovic, W. Damm, J. Engblom, G. Folher, M. Garcia-Valls, H. Kopetz, Y. Lakhnech, F. Laroussinie, L. Lavagno, G. Lipari, F. Maraninchi, Ph. Peti, J. de la Puente, N. Scaife, J. Sifakis, R. de Simone, M. Torngren, P. Veríssimo, A. J. Wellings, R. Wilhelm, T. Willemse, and W. Yi. Guidelines for a graduate curriculum on embedded software and systems. ACM Transactions on Embedded Computing Systems, 4(3):587–611, August 2005. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Chang:2015:PVL

[CSW15] Li-Pin Chang, Yo-Chuan Su, and I-Chen Wu. Plugging versus logging: Adaptive buffer management for hybrid-mapping SSDs. ACM Transactions on Embedded Computing Systems, 14(2): 29:1–29:??, March 2015. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Chen:2013:ELS

[CTK^{+}13] Yeong-Sheng Chen, Yun-Ju Ting, Chih-Heng Ke, Naveen Chilamkruti, and Jong Hyuk Park. Efficient localization scheme with ring overlapping by utilizing mobile anchors in wireless sensor networks. ACM Transactions on Embedded Computing Systems, 12(2):20:1–20:??, February 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Cullmann:2013:CPA

[Cul13] Christoph Cullmann. Cache persistence analysis: Theory and practice. ACM Transactions on Embedded Computing Systems, 12(1s):40:1– 40:??, March 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Crenne:2013:CMS

[CVG^{+}13] Jérémie Crenne, Romain Vaslin, Guy Gogniat, Jean-Philippe Diguet, Russell Tessier, and Deepak Unnikrishnan. Configurable memory security in embedded systems. ACM Transactions on Embedded Computing Systems, 12(3):71:1– 71:??, March 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Chang:2014:RAC

[CW14] Li-Pin Chang and Chen-Yi Wen. Reducing asynchrony in channel garbagecollection for improving internal parallelism of multichannel solid-state disks. ACM Transactions on Embedded Computing Systems, 13(2s):63:1–63:??, January 2014. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Chen:2016:ICA

[CWH^{+}16] Renhai Chen, Yi Wang, Jingtong Hu, Duo Liu, Zili Shao, and Yong Guan. Image-content-aware I/O optimization for mobile virtualization. ACM Transactions on Embedded Computing Systems, 16(1):12:1– 12:??, November 2016. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Chen:2023:RCR

[CWH^{+}23] Wei-Ju Chen, Peng Wu, Pei-Chi Huang, Aloysius K. Mok, and Song Han. Regular composite resource partitioning and reconfiguration in open systems. ACM Transactions on Embedded Computing Systems, 22(5): 84:1–84:??, September 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3609424.

Chen:2017:LBD

[CWJ17] Xiaogang Chen, Z. Jane Wang, and Xiangyang Ji. A load-balancing divide-and-conquer SVM solver. ACM Transactions on Embedded Computing Systems, 16(3): 74:1–74:??, July 2017. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Chang:2012:AFS

[CWKH12] Yuan-Hao Chang, Po-Liang Wu, Tei-Wei Kuo, and Shih-Hao Hung. An adaptive filesystem-oriented FTL mechanism for flash-memory storage systems. ACM Transactions on Embedded Computing Systems, 11(1):9:1– 9:??, March 2012. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Chen:2023:FNN

[CWX^{+}23] Weiwei Chen, Ying Wang, Ying Xu, Chengsi Gao, Cheng Liu, and Lei Zhang. A framework for neural network architecture and compile co-optimization. ACM Transactions on Embedded Computing Systems, 22(1): 5:1–5:??, January 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3533251.

Chang:2020:DAR

[CWZ^{+}20] Wanli Chang, Ran Wei, Shuai Zhao, Andy Wellings, Jim Woodcock, and Alan Burns. Development automation of real-time Java: Model-driven transformation and synthesis. ACM Transactions on Embedded Computing Systems, 19(5):31:1– 31:26, November 2020. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3391897.

Cai:2023:OOF

[CWZ23] Xuyi Cai, Ying Wang, and Lei Zhang. Optimus: an operator fusion framework for deep neural networks. ACM Transactions on Embedded Computing Systems, 22(1):

1:1–1:??, January 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3520142.

Chen:2017:SBT

[CYH^{+}17] Jiunn-Yeu Chen, Wuu Yang, Wei-Chung Hsu, Bor-Yeh Shen, and Quan-Huei Ou. On static binary translation of ARM/Thumb mixed ISA binaries. ACM Transactions on Embedded Computing Systems, 16(3):81:1– 81:??, July 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Chen:2020:QEO

[CYH20] Fupeng Chen, Heng Yu, and Yajun Ha. Quality estimation and optimization of adaptive stereo matching algorithms for smart vehicles. ACM Transactions on Embedded Computing Systems, 19(2):10:1– 10:24, March 2020. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3372784.

Chang:2013:RED

[CYKH13] Yuan-Hao Chang, Ming-Chang Yang, Tei-Wei Kuo, and Ren-Hung Hwang. A reliability enhancement design under the flash translation layer for MLC-based flash-memory storage systems. ACM Transactions on Embedded Computing Systems, 13(1):10:1–10:??, August 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Calsi:2023:IRA

[CZ23] Davide Li Calsi and Vittorio Zaccaria. Interruptible remote attestation of low-end IoT microcontrollers via performance counters. ACM Transactions on Embedded Computing Systems, 22(5):87:1–87:??, September 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3611674.

Chen:2024:DTO

[CZH^{+}24] Ying Chen, Jie Zhao, Jintao Hu, Shaohua Wan, and Jiwei Huang. Distributed task offloading and resource purchasing in NOMA-Enabled mobile edge computing: Hierarchical game theoretical approaches. ACM Transactions on Embedded Computing Systems, 23(1):2:1– 2:??, January 2024. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3597023.

Chen:2023:ARL

[CZHK23] Huili Chen, Xinqiao Zhang, Ke Huang, and Farinaz Koushanfar. AdaTest: Reinforcement learning and adaptive sampling for onchip hardware Trojan detection. ACM Transactions on Embedded Computing Systems, 22(2):37:1– 37:??, March 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3544015.

Chen:2022:ARF

[CZK^{+}22] Yanfeng Chen, Tianyu Zhang, Fanxin Kong, Lin Zhang, and Qingxu Deng. Attackresilient fusion of sensor data with uncertain delays. ACM Transactions on Embedded Computing Systems, 21(4):39:1–39:??, July 2022. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3532181.

Dupont:2021:EBH

[DAASP21] Guillaume Dupont, Yamine Ait-Ameur, Neeraj Kumar Singh, and Marc Pantel. Event-B hybridation: a proof and refinement-based framework for modelling hybrid systems. ACM Transactions on Embedded Computing Systems, 20(4):35:1– 35:37, June 2021. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3448270.

Das:2016:AHR

[DAHM16] Anup Das, Bashir M. Al-Hashimi, and Geoff V. Merrett. Adaptive and hierarchical runtime manager for energy-aware thermal management of embedded systems. ACM Transactions on Embedded Computing Systems, 15(2):24:1–24:??, May 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

DiBiagio:2012:AOA

[DASS12] Andrea Di Biagio, Giovanni Agosta, Martino Sykora, and Cristina Silvano. Architecture optimization of application-specific implicit instructions. ACM Transactions on Embedded Computing Systems, 11(S2):44:1– 44:??, 2012. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Duggirala:2019:ASR

[DB19] Parasara Sridhar Duggirala and Stanley Bak. Aggregation strategies in reachable set computation of hybrid systems. ACM Transactions on Embedded Computing Systems, 18(5s):99:1– 99:??, October 2019. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

                                        tronic). URL https://dl.
                                        acm.org/ft_gateway.cfm?
                                        id=3358214.

Dua:2014:CSS

[DBFH14] Akshay Dua, Nirupama Bulusu, Wu-Chang Feng, and Wen Hu. Combating software and Sybil attacks to data integrity in crowd-sourced embedded systems. ACM Transactions on Embedded Computing Systems, 13(5s):154:1– 154:??, September 2014. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

DelBarrio:2014:ULP

[DBH14] Alberto A. Del Barrio, Nader Bagherzadeh, and Román Hermida. Ultra-low-power adder stage design for exascale floating point units. ACM Transactions on Embedded Computing Systems, 13(3s):105:1–105:??, March 2014. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Davis:2015:GPM

[DBM^{+}15] Robert I. Davis, Alan Burns, Jose Marinho, Vincent Nelis, Stefan M. Petters, and Marko Bertogna. Global and partitioned multiprocessor fixed priority scheduling with deferred preemption. ACM Transactions on Embedded Computing Systems, 14(3):47:1–47:??, April 2015. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Daghero:2022:HAR

[DBX^{+}22] Francesco Daghero, Alessio Burrello, Chen Xie, Marco Castellano, Luca Gandolfi, Andrea Calimera, Enrico Macii, Massimo Poncino, and Daniele Jahier Pagliari. Human activity recognition on microcontrollers with quantized and adaptive deep neural networks. ACM Transactions on Embedded Computing Systems, 21(4): 46:1–46:??, July 2022. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3542819.

Dai:2019:DMS

[DCZB19] Xiaotian Dai, Wanli Chang, Shuai Zhao, and Alan Burns. A dual-mode strategy for performance-maximisation and resource-efficient CPS design. ACM Transactions on Embedded Computing Systems, 18(5s):85:1– 85:??, October 2019. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3358213.

Deng:2024:SST

[DDC^{+}24] Jianing Deng, Shunjie Dong, Lvcheng Chen, Jingtong Hu, and Cheng Zhuo. STDF: Spatio-temporal deformable fusion for video quality enhancement on embedded platforms. ACM Transactions on Embedded Computing Systems, 23(2):25:1– 25:??, March 2024. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3645113.

Davare:2013:MDE

[DDG^{+}13] Abhijit Davare, Douglas Densmore, Liangpeng Guo, Roberto Passerone, Alberto L. Sangiovanni-Vincentelli, Alena Simalatsar, and Qi Zhu. metroII: a design environment for cyber-physical systems. ACM Transactions on Embedded Computing Systems, 12(1s):49:1–49:??, March 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Dean:2006:STI

[Dea06] Alexander G. Dean. Software thread integration for embedded system display applications. ACM Transactions on Embedded Computing Systems, 5(1):116– 151, February 2006. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Diguet:2011:CLB

[DEG11] Jean-Philippe Diguet, Yvan Eustache, and Guy Gogniat. Closed-loop–based selfadaptive Hardware/Software-Embedded systems: Design methodology and smart CAM case study. ACM Transactions on Embedded Computing Systems, 10(3): 38:1–38:??, April 2011. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Dewan:2014:BAF

[DF14] Farhana Dewan and Nathan Fisher. Bandwidth allocation for fixed-priority-scheduled compositional realtime systems. ACM Transactions on Embedded Computing Systems, 13(4):91:1– 91:??, February 2014. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Silva:2019:RFG

[DFC^{+}19] Lucas Bragança Da Silva, Ricardo Ferreira, Michael Canesche, Marcelo M. Menezes, Maria D. Vieira, Jeronimo Penha, Peter Jamieson, and José Augusto M. Nacif. READY: a fine-grained multithreading overlay framework for modern CPU– FPGA dataflow applications. ACM Transactions on Embedded Computing Systems, 18(5s):56:1–56:??, October 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

                                        tronic). URL https://dl.
                                        acm.org/ft_gateway.cfm?
                                        id=3358187.

Leyva-del-Foyo:2012:ITI

[dFMAdN12] Luis E. Leyva del Foyo, Pedro Mejia-Alvarez, and Dionisio de Niz. Integrated task and interrupt management for real-time systems. ACM Transactions on Embedded Computing Systems, 11(2):32:1–32:??, July 2012. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Das:2020:ALS

[DGC^{+}20] Tuhin Subhra Das, Prasun Ghosal, Navonil Chatterjee, Arnab Nath, Akash Banerjee, and Subhojyoti Khastagir. Application of logical sub-networking in congestion-aware deadlockfree SDmesh routing. ACM Transactions on Embedded Computing Systems, 19(4): 24:1–24:26, July 2020. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3387928.

Deantoni:2024:ISI

[DGG24] Julien Deantoni, Alain Girault, and Daniel Grosse. Introduction to the special issue on specification and design languages (FDL 2021). ACM Transactions on Embedded Computing Systems, 23(5):65:1–65:??, September 2024. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3677316.

Dong:2023:RAS

[DHB^{+}23] Yi Dong, Wei Huang, Vibhav Bharti, Victoria Cox, Alec Banks, Sen Wang, Xingyu Zhao, Sven Schewe, and Xiaowei Huang. Reliability assessment and safety arguments for machine learning components in system assurance. ACM Transactions on Embedded Computing Systems, 22(3):48:1–48:??, May 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3570918.

Dokhanchi:2018:FRD

[DHF18] Adel Dokhanchi, Bardh Hoxha, and Georgios Fainekos. Formal requirement debugging for testing and verification of cyber-physical systems. ACM Transactions on Embedded Computing Systems, 17(2):34:1–34:??, April 2018. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Deshmukh:2017:TCP

[DHJ^{+}17] Jyotirmoy Deshmukh, Marko Horvat, Xiaoqing Jin, Rupak Majumdar, and Vinayak S. Prabhu. Testing cyberphysical systems through Bayesian optimization. ACM Transactions on Embedded Computing Systems, 16(5s): 170:1–170:??, October 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

DeGroote:2015:IAC

[DHKS15] Robert De Groote, Philip K. F. Hölzenspies, Jan Kuper, and Gerard J. M. Smit. Incremental analysis of cyclo-static synchronous dataflow graphs. ACM Transactions on Embedded Computing Systems, 14(4):68:1–68:??, December 2015. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Dietrich:2017:GOF

[DHL17] Christian Dietrich, Martin Hoffmann, and Daniel Lohmann. Global optimization of fixed-priority realtime systems by RTOSaware control-flow analysis. ACM Transactions on Embedded Computing Systems, 16(2):35:1–35:??, April 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Deb:2023:ZDT

[DJ23] Dipika Deb and John Jose.

ZPP: a dynamic technique to eliminate cache pollution in NoC based MPSoCs. ACM Transactions on Embedded Computing Systems, 22(5s):118:1–118:??, October 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3609113.

Deshwal:2019:MMO

[DJJ^{+}19] Aryan Deshwal, Nitthilan Kanappan Jayakodi, Biresh Kumar Joardar, Janardhan Rao Doppa, and Partha Pratim Pande. MOOS: a multi-objective design space exploration and optimization framework for NoC enabled manycore systems. ACM Transactions on Embedded Computing Systems, 18(5s):77:1–77:??, October 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3358206.

Dubach:2012:EPE

[DJO12] Christophe Dubach, Timothy M. Jones, and Michael F. P. O’Boyle. Exploring and predicting the effects of microarchitectural parameters and compiler optimizations on performance and energy. ACM Transactions on Embedded Computing Systems, 11S(1):24:1–24:??, 2012. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Dutt:2016:TSE

[DJS16] Nikil Dutt, Axel Jantsch, and Santanu Sarma. Toward smart embedded systems: a self-aware systemon-chip (SoC) perspective.

ACM Transactions on Embedded Computing Systems, 15(2):22:1–22:??, May 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Ding:2013:DAV

[DJZ13] Zhijun Ding, Changjun Jiang, and Mengchu Zhou. Design, analysis and verification of real-time systems based on time Petri net refinement. ACM Transactions on Embedded Computing Systems, 12(1):4:1– 4:??, January 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Dave:2019:DEP

[DKA^{+}19] Shail Dave, Youngbin Kim, Sasikanth Avancha, Kyoungwoo Lee, and Aviral Shrivastava. dMazeRunner: Executing perfectly nested loops on dataflow accelerators. ACM Transactions on Embedded Computing Systems, 18(5s):70:1–70:??, October 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/ft_gateway.
cfm?id=3358198.

Dhurjati:2005:MSG

[DKAL05] Dinakar Dhurjati, Sumant Kowshik, Vikram Adve, and Chris Lattner. Memory safety without garbage collection for embedded applications. ACM Transactions on Embedded Computing Systems, 4(1):73–111, February 2005. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Das:2014:EAT

[DKV14] Anup Das, Akash Kumar, and Bharadwaj Veeravalli. Energy-aware task mapping and scheduling for reliable embedded computing systems. ACM Transactions on Embedded Computing Systems, 13(2s):72:1–72:??, January 2014. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Dong:2012:UAS

[DL12] Qi Dong and Donggang Liu. Using auxiliary sensors for pairwise key establishment in WSN. ACM Transactions on Embedded Computing Systems, 11(3):59:1– 59:??, September 2012. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Dong:2014:EEE

[DLC^{+}14] Wei Dong, Yunhao Liu, Chun Chen, Lin Gu, and Xiaofan Wu. Elon: Enabling efficient and long-term reprogramming for wireless sensor networks. ACM Transactions on Embedded Computing Systems, 13(4):77:1– 77:??, February 2014. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Dugo:2019:CLC

[DLD^{+}19] Alexy Torres Aurora Dugo, Jean-Baptiste Lefoul, Felipe Gohring De Magalhaes, Dahman Assal, and Gabriela Nicolescu. Cache locking content selection algorithms for ARINC-653 compliant RTOS. ACM Transactions on Embedded Computing Systems, 18(5s):76:1– 76:??, October 2019. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3358196.

Dong:2016:DLD

[DLH16] Wei Dong, Luyao Luo, and Chao Huang. Dynamic logging with dylog in networked embedded systems. ACM Transactions on Embedded Computing Systems, 15(1): 5:1–5:??, February 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Dong:2013:PRS

[DLN13] Qi Dong, Donggang Liu, and Peng Ning. Providing DoS resistance for signature-based broadcast authentication in sensor networks. ACM Transactions on Embedded Computing Systems, 12(3):73:1– 73:??, March 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Duraisamy:2016:HPE

[DLPK16] Karthi Duraisamy, Hao Lu, Partha Pratim Pande, and Ananth Kalyanaraman. High-performance and energy-efficient network-on-chip architectures for graph analytics. ACM Transactions on Embedded Computing Systems, 15(4):66:1– 66:??, August 2016. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Desirena-Lopez:2019:TAR

[DLRTB^{+}19] G. Desirena-López, A. Ramírez-Treviño, J. L. Briz, C. R. Vázquez, and D. Gómez-Gutiérrez. Thermal-aware real-time scheduling using timed continuous Petri nets. ACM Transactions on Embedded Computing Systems, 18(4):36:1–36:??, August 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

                                        /dl.acm.org/ft_gateway.
                                        cfm?id=3322643.

Du:2024:AMC

[DLS^{+}24] Sichun Du, Jun Li, Chen Sun, Pingdan Xiao, Qinghui Hong, and Jiliang Zhang. Analog in-memory circuit design of polynomial multiplication for lattice cipher acceleration application. ACM Transactions on Embedded Computing Systems, 23(6):88:1–88:??, November 2024. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3605891.

DiPietro:2016:CLD

[DLV16] Roberto Di Pietro, Flavio Lombardi, and Antonio Villani. CUDA leaks: a detailed hack for CUDA and a (partial) fix. ACM Transactions on Embedded Computing Systems, 15(1): 15:1–15:??, February 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Das:2023:EFS

[DMPC23] Satyajit Das, Kevin Martin, Thomas Peyret, and Philippe Coussy. An efficient and flexible stochastic CGRA mapping approach. ACM Transactions on Embedded Computing Systems, 22(1):8:1– 8:??, January 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3550071.

Desai:2022:CLR

[DNBL22] Harsh Desai, Matteo Nardello, Davide Brunelli, and Brandon Lucia. Camaroptera: a long-range image sensor with local inference for remote sensing applications. ACM Transactions on Embedded Computing Systems, 21(3):32:1–32:25, May 2022. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3510850.

Dasari:2014:NCA

[DNNP14] Dakshina Dasari, Borislav Nikolić, Vincent Nélis, and Stefan M. Petters. NoC contention analysis using a branch-and-prune algorithm. ACM Transactions on Embedded Computing Systems, 13(3s):113:1–113:??, March 2014. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Dutt:2018:ADA

[DNT18] Sunil Dutt, Sukumar Nandi, and Gaurav Trivedi. Analysis and design of adders for approximate computing. ACM Transactions on Embedded Computing Systems, 17(2):40:1–40:??, April 2018. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

DiNatale:2008:BOM

[DP08] Marco Di Natale and Valerio Pappalardo. Buffer optimization in multitask implementations of Simulink models. ACM Transactions on Embedded Computing Systems, 7(3):23:1–23:??, April 2008. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Durrieu:2019:GAC

[DP19] Guy Durrieu and Claire Pagetti. GRec: Automatic computation of reconfiguration graphs for multi-core platforms. ACM Transactions on Embedded Computing Systems, 18(5):41:1– 41:??, October 2019. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3350533.

Desnos:2016:MRB

[DPNA16] Karol Desnos, Maxime Pelcat, Jean-François Nezan, and Slaheddine Aridhi. On memory reuse between inputs and outputs of dataflow actors. ACM Transactions on Embedded Computing Systems, 15(2):30:1– 30:??, May 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Daneshtalab:2014:ESI

[DPP14] Masoud Daneshtalab, Maurizio Palesi, and Juha Plosila. Editorial: Special issue on design challenges for manycore processors. ACM Transactions on Embedded Computing Systems, 13(3s):

100:1–100:??, March 2014. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Dunbar:2014:DTE

[DQ14] Carson Dunbar and Gang Qu. Designing trusted embedded systems from finite state machines. ACM Transactions on Embedded Computing Systems, 13(5s): 153:1–153:??, September 2014. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Driver:2010:MES

[DRL^{+}10] Cormac Driver, Sean Reilly, Éamonn Linehan, Vinny Cahill, and Siobhán Clarke. Managing embedded systems complexity with aspectoriented model-driven engineering. ACM Transactions on Embedded Computing Systems, 10(2):21:1– 21:??, December 2010. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Dini:2011:LLA

[DS11] Gianluca Dini and Ida M. Savino. LARK: a lightweight authenticated ReKeying scheme for clustered wireless sensor networks. ACM Transactions on Embedded Computing Systems, 10(4):41:1– 41:??, November 2011. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Dixit:2024:PPA

[DS24] Akanksha Dixit and Smruti R. Sarangi. PredATW: Predicting the asynchronous time warp latency for VR systems. ACM Transactions on Embedded Computing Systems, 23(5):80:1– 80:??, September 2024. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3677329.

Devaraj:2017:FTP

[DSB17] Rajesh Devaraj, Arnab Sarkar, and Santosh Biswas. Fault-tolerant preemptive aperiodic RT scheduling by supervisory control of TDES on multiprocessors. ACM Transactions on Embedded Computing Systems, 16(3): 87:1–87:??, July 2017. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Dick:2012:ISS

[DSD12] Robert Dick, Li Shang, and Nikil Dutt. Introduction to special section SCPS’09. ACM Transactions on Embedded Computing Systems, 11(4):74:1–74:??, December 2012. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Derler:2019:GES

[DST19] Patricia Derler, Klaus Schneider, and Jean-Pierre Talpin. Guest editorial: Special issue of ACM TECS on the ACM– IEEE International Conference on Formal Methods and Models for System Design (MEMOCODE 2017). ACM Transactions on Embedded Computing Systems, 18(1): 1:1–1:??, February 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/ft_gateway.
cfm?id=3292422.

DoCarmoLucas:2009:ADF

[DSW^{+}09] Amilcar Do Carmo Lucas, Henning Sahlbach, Sean Whitty, Sven Heithecker, and Rolf Ernst. Application development with the FlexWAFE real-time stream processing architecture for FPGAs. ACM Transactions on Embedded Computing Systems, 9(1):4:1– 4:??, October 2009. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Diamantopoulos:2014:PFS

[DSXS^{+}14]  Dionysios  Diamantopou-

los, Efstathios Sotiriou-Xanthopoulos, Kostas Siozios, George Economakos, and Dimitrios Soudris. Plug&Chip: a framework for supporting rapid prototyping of 3D hybrid virtual SoCs. ACM Transactions on Embedded Computing Systems, 13 (5s):168:1–168:??, November 2014. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Diamantopoulos:2015:GPA

[DSXS15] Dionysios Diamantopoulos, Kostas Siozios, Sotirios Xydis, and Dimitrios Soudris. GENESIS: Parallel application placement onto reconfigurable architectures (invited for the special issue on runtime management). ACM Transactions on Embedded Computing Systems, 14(1): 18:1–18:??, January 2015. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

DeSutter:2013:ISS

[DV13] Bjorn De Sutter and Jan Vitek. Introduction to the special section on LCTES’11. ACM Transactions on Embedded Computing Systems, 12(1s):38:1–38:??, March 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

DeSutter:2007:LTC

[DVC^{+}07] Bjorn De Sutter, Ludo Van Put, Dominique Chanet, Bruno De Bus, and Koen De Bosschere. Link-time compaction and optimization of ARM executables. ACM Transactions on Embedded Computing Systems, 6(1):

5:1–5:??, February 2007. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Dharmaraj:2021:OSP

[DVC21] Celia Dharmaraj, Vinita Vasudevan, and Nitin Chandrachoodan. Optimization of signal processing applications using parameterized error models for approximate adders. ACM Transactions on Embedded Computing Systems, 20(2):12:1– 12:25, March 2021. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3430509.

Durr:2019:EET

[DVCC19] Marco Dürr, Georg Von Der Brüggen, Kuan-Hsun Chen, and Jian-Jia Chen. End-to-end timing analysis of sporadic cause-effect chains in distributed systems. ACM Transactions on Embedded Computing Systems, 18(5s): 58:1–58:??, October 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/ft_gateway.
cfm?id=3358181.

DosSantos:2010:MPB

[DW10] Osmar Marchi Dos Santos and Andy Wellings. Measuring and policing blocking times in real-time systems. ACM Transactions on Embedded Computing Systems, 10(1):2:1–2:??, August 2010. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Natale:2014:ESI

[DWCM14] Marco Di Natale, Rich West, Jian-Jia Chen, and Rahul Mangharam. Editorial: Special issue on real-time and embedded technology and applications. ACM Transactions on Embedded Computing Systems, 13(4s):119:1– 119:??, April 2014. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

DeNiz:2014:UBR

[DWRR14] Dionisio De Niz, Lutz Wrage, Anthony Rowe, and Ragunathan (Raj) Rajkumar. Utility-based resource overbooking for cyber-physical systems. ACM Transactions on Embedded Computing Systems, 13(5s):162:1– 162:??, September 2014. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Dong:2022:EEA

[DZL^{+}22] Jiankuo Dong, Fangyu Zheng, Jingqiang Lin, Zhe Liu, Fu Xiao, and Guang Fan. EC-ECC: Accelerating elliptic curve cryptography for edge computing on embedded GPU TX2. ACM Transactions on Embedded Computing Systems, 21(2): 16:1–16:25, March 2022. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3492734.

Doblander:2009:NSF

[DZR09] Andreas Doblander, Andreas Zoufal, and Bernhard Rinner. A novel software framework for embedded multiprocessor smart cameras. ACM Transactions on Embedded Computing Systems, 8(3):24:1–24:??, April 2009. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Eldstaal-Ahrens:2022:CCL

[EAAS22] Albin Eldstål-Ahrens, Angelos Arelakis, and Ioannis Sourdis. L^{2}C: Combining lossy and lossless compression on memory and I/O. ACM Transactions on Embedded Computing Systems, 21(1):12:1–12:27, January 2022. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3481641.

Editors:2013:ISS

[Edi13] Editors. Introduction to the special section on ES-TIMedia’10. ACM Transactions on Embedded Computing Systems, 13(1s):26:1– 26:??, November 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Editors:2014:MMA

[Edi14] Editors. Monitoring massive appliances by a minimal number of smart meters. ACM Transactions on Embedded Computing Systems, 13(2s):56:1–56:??, January 2014. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Eles:2016:GES

[EE16] Petru Eles and Rolf Ernst.

Guest editorial for special issue of ESWEEK 2015. ACM Transactions on Embedded Computing Systems, 15(4): 63:1–63:??, August 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Eles:2018:GES

[EH18] Petru Eles and Jörg Henkel.

Guest editorial for the special issue of ESWEEK 2016. ACM Transactions on Embedded Computing Systems, 17(1):14:1–14:??, January 2018. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

ElAdoui:2024:CTM

[EHF24] Mohammed El Adoui, Thomas Herpoel, and Benoît Frénay. Constrained tiny machine learning for predicting gas concentration with I4.0 lowcost sensors. ACM Transactions on Embedded Computing Systems, 23(3):51:1– 51:??, May 2024. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3590956.

Eriksson:2012:ICG

[EK12] Mattias Eriksson and Christoph Kessler. Integrated code generation for loops. ACM Transactions on Embedded Computing Systems, 11S(1): 19:1–19:??, 2012. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Elkhatib:2024:CEF

[EKAK24] Rami Elkhatib, Brian Koziel, Reza Azarderakhsh, and Mehran Mozaffari Kermani. Cryptographic engineering a fast and efficient SIKE in FPGA. ACM Transactions on Embedded Computing Systems, 23(2):31:1– 31:??, March 2024. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3584919.

Egger:2008:DSM

[ELS08] Bernhard Egger, Jaejin Lee, and Heonshik Shin. Dynamic scratchpad memory management for code in portable systems with an MMU. ACM Transactions on Embedded Computing Systems, 7(2):

11:1–11:??, February 2008. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

ElYaacoub:2023:SDS

[EMVR23] Ahmed El Yaacoub, Luca Mottola, Thiemo Voigt, and Philipp Rümmer. Scheduling dynamic software updates in mobile robots. ACM Transactions on Embedded Computing Systems, 22(6): 99:1–99:??, November 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3623676.

Elewi:2014:EET

[ESAS14] Abdullah Elewi, Mohamed Shalan, Medhat Awadalla, and Elsayed M. Saad. Energy-efficient task allocation techniques for asymmetric multiprocessor embedded systems. ACM Transactions on Embedded Computing Systems, 13(2s):71:1– 71:??, January 2014. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Ernst:2023:ACN

[ESBK23] Rolf Ernst, Dominik Stöhrmann, Alex Bendrick, and Adam Kostrzewa. Applicationcentric network management — addressing safety and real-time in V2X applications. ACM Transactions on Embedded Computing Systems, 22(2):20:1– 20:??, March 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3528411.

Egilmez:2017:UAF

[ESM^{+}17] Begum Egilmez, Matthew Schuchhardt, Gokhan Memik, Raid Ayoub, Niranjan Soundararajan, and Michael Kishinevsky. User-aware frame rate management in Android Smartphones. ACM Transactions on Embedded Computing Systems, 16(5s):131:1– 131:??, October 2017. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Emeretlis:2016:LBB

[ETAV16] Andreas Emeretlis, George Theodoridis, Panayiotis Alefragis, and Nikolaos Voros. A Logic-Based Benders decomposition approach for mapping applications on heterogeneous multicore platforms. ACM Transactions on Embedded Computing Systems, 15(1):19:1–19:??, February 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Edwards:2019:CDC

[ETBK19] Stephen A. Edwards, Richard Townsend, Martha Barker, and Martha A. Kim. Compositional dataflow circuits.

ACM Transactions on Embedded Computing Systems, 18(1):5:1–5:??, February 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/ft_gateway.
cfm?id=3274280.

Esposito:2017:NMO

[EVS^{+}17] Stefano Esposito, Massimo Violante, Marco Sozzi, Marco Terrone, and Massimo Traversone. A novel method for online detection of faults affecting executiontime in multicore-based systems. ACM Transactions on Embedded Computing Systems, 16(4):94:1–94:??, August 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Erata:2023:EEA

[EYG^{+}23] Ferhat Erata, Eren Yildiz, Arda Goknil, Kasim Sinan Yildirim, Jakub Szefer, Ruzica Piskac, and Gokcin Sezgin. ETAP: Energy-aware timing analysis of intermittent programs. ACM Transactions on Embedded Computing Systems, 22(2):23:1– 23:??, March 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3563216.

Elfar:2017:SER

[EZL^{+}17] Mahmoud Elfar, Zhanwei Zhong, Zipeng Li, Krishnendu Chakrabarty, and Miroslav Pajic. Synthesis of error-recovery protocols for micro-electrode-dot-array digital microfluidic biochips. ACM Transactions on Embedded Computing Systems, 16(5s):127:1– 127:??, October 2017. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Franchino:2016:BOE

[FBM16] Gianluca Franchino, Giorgio Buttazzo, and Mauro Marinoni. Bandwidth optimization and energy management in real-time wireless networks. ACM Transactions on Embedded Computing Systems, 15(3):41:1– 41:??, July 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Furtado:2013:CON

[FC13] Pedro Furtado and José Cecílio. Configuration and operation of networked control systems over heterogeneous WSANs. ACM Transactions on Embedded Computing Systems, 13(1s):34:1– 34:??, November 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Fusella:2016:CAA

[FC16] Edoardo Fusella and Alessandro Cilardo. Crosstalkaware automated mapping for optical networks-on-chip. ACM Transactions on Embedded Computing Systems, 16(1):16:1–16:??, November 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Ferri:2009:RIF

[FF09] B. H. Ferri and A. A.

Ferri. Reconfiguration of IIR filters in response to computer resource availability. ACM Transactions on Embedded Computing Systems, 9(1):2:1–2:??, October 2009. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Fischer:2025:TAC

[FF25] Thilo Leon Fischer and Heiko Falk. Towards analysing cache-related preemption delay in noninclusive cache hierarchies. ACM Transactions on Embedded Computing Systems, 24(1):8:1–8:??, January 2025. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3695768.

Feng:2023:ADS

[FFA^{+}23] Kaijie Feng, Xiaoya Fan, Jianfeng An, Chuxi Li, Kaiyue Di, and Jiangfei Li. ACDSE: a design space exploration method for CNN accelerator based on adaptive compression mechanism.

ACM Transactions on Embedded Computing Systems, 22(6):95:1–95:??, November 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3545177.

Fainekos:2012:ESS

[FGIS12] Georgios Fainekos, Eric Goubault, Franjo Ivancić, and Sriram Sankaranarayanan. Editorial: Special section VCPSS’09. ACM Transactions on Embedded Computing Systems, 11(S2):52:1– 52:??, 2012. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Fradet:2023:RRD

[FGK^{+}23] Pascal Fradet, Alain Girault, Ruby Krishnaswamy, Xavier Nicollin, and Arash Shafiei. RDF: a reconfigurable dataflow model of computation. ACM Transactions on Embedded Computing Systems, 22(1):12:1– 12:??, January 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3544972.

Feng:2019:EUH

[FGL^{+}19] Zhiwei Feng, Nan Guan, Mingsong Lv, Weichen Liu, Qingxu Deng, Xue Liu, and Wang Yi. An efficient UAV hijacking detection method using onboard inertial measurement unit. ACM Transactions on Embedded Computing Systems, 17(6):96:1– 96:??, January 2019. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3289390.

Fleming:2017:CDI

[FHB^{+}17] Tom Fleming, Huang-Ming Huang, Alan Burns, Chris Gill, Sanjoy Baruah, and Chenyang Lu. Corrections to and discussion of “Implementation and Evaluation of Mixed-criticality Scheduling Approaches for Sporadic Tasks”. ACM Transactions on Embedded Computing Systems, 16(3):77:1– 77:??, July 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). See [HGL14].

Fard:2021:APP

[FHK21] Mahdi Mohammadpour Fard, Mahmood Hasanloo, and Mehdi Kargahi. Analytical program power characterization for battery depletiontime estimation. ACM Transactions on Embedded Computing Systems, 20(2): 9:1–9:9, March 2021. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3421511.

Fahmida:2025:RRT

[FJM^{+}25] Sezana Fahmida, Aakriti Jain, Venkata Prashant Modekurthy, Dali Ismail, and Abusayeed Saifullah. RTPL: a real-time communication protocol for LoRa network. ACM Transactions on Embedded Computing Systems, 24(1):17:1– 17:??, January 2025. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3702209.

Fan:2018:SDR

[FKJM18] Chuchu Fan, James Kapinski, Xiaoqing Jin, and Sayan Mitra. Simulationdriven reachability using matrix measures. ACM Transactions on Embedded Computing Systems, 17(1):21:1– 21:??, January 2018. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Fellner:2019:MBM

[FKS^{+}19] Andreas Fellner, Willibald Krenn, Rupert Schlick, Thorsten Tarrach, and Georg Weissenbacher. Modelbased, mutation-driven testcase generation via heuristicguided branching search. ACM Transactions on Embedded Computing Systems, 18(1):4:1–4:??, February 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/ft_gateway.
cfm?id=3289256.

Fezzardi:2017:UEP

[FLF17] Pietro Fezzardi, Marco Lattuada, and Fabrizio Ferrandi. Using efficient path profiling to optimize memory consumption of on-chip debugging for high-level synthesis. ACM Transactions on Embedded Computing Systems, 16(5s):149:1–149:??, October 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Fu:2023:RRB

[FLF^{+}23] Yu Fu, Jingqiang Lin, Dengguo Feng, Wei Wang, Mingyu Wang, and Wenjie Wang. RegKey: a register-based implementation of ECC signature algorithms against one-shot memory disclosure. ACM Transactions on Embedded Computing Systems, 22(6): 97:1–97:??, November 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3604805.

Falk:2012:ISS

[FM12] Heiko Falk and Peter Marwedel. Introduction to the Special Section on SCOPES’09. ACM Transactions on Embedded Computing Systems, 11S(1):17:1– 17:??, 2012. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Farahmand:2023:DAH

[FMHS23] Ebrahim Farahmand, Ali Mahani, Muhammad Abdullah Hanif, and Muhammad Shafique. Design and analysis of high performance heterogeneous block-based approximate adders. ACM Transactions on Embedded Computing Systems, 22 (6):106:1–106:??, November 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3625686.

Furbach:2015:MMA

[FMSS15] Florian Furbach, Roland Meyer, Klaus Schneider, and Maximilian Senftleben. Memory-model-aware testing: a unified complexity analysis. ACM Transactions on Embedded Computing Systems, 14(4):63:1– 63:??, December 2015. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Ferreira:2016:LRF

[FND^{+}16] Ronaldo R. Ferreira, Gabriel L. Nazar, Jean Da Rolt, Álvaro F. Moreira, and Luigi Carro. Live-out register fencing: Interrupt-triggered soft error correction based on the elimination of registerto-register communication.

ACM Transactions on Embedded Computing Systems, 15(3):60:1–60:??, July 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Franke:2003:ARH

[FO03] Björn Franke and Michael O’Boyle. Array recovery and high-level transformations for DSP applications. ACM Transactions on Embedded Computing Systems, 2(2): 132–162, May 2003. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

France-Pillois:2022:SAR

[FPGS22] Maxime France-Pillois, Abdoulaye Gamatié, and Gilles Sassatelli. A segmented adaptive router for near energy-proportional networkson-chip. ACM Transactions on Embedded Computing Systems, 21(4):40:1– 40:??, July 2022. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3529106.

Franke:2012:SPM

[Fra12]  Björn Franke.  Statisti-

cal performance modeling in functional instruction set simulators. ACM Transactions on Embedded Computing Systems, 11S(1):22:1– 22:??, 2012. CODEN ????

ISSN 1539-9087 (print), 1558-3465 (electronic).

Fei:2007:EOS

[FRRJ07] Yunsi Fei, Srivaths Ravi, Anand Raghunathan, and Niraj K. Jha. Energyoptimizing source code transformations for operating system-driven embedded software. ACM Transactions on Embedded Computing Systems, 7(1):2:1– 2:26, December 2007. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Forte:2013:ETA

[FS13] Domenic Forte and Ankur Srivastava. Energy- and thermal-aware video coding via encoder/decoder workload balancing. ACM Transactions on Embedded Computing Systems, 12(2s):96:1– 96:??, May 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Foglia:2014:ERI

[FS14] Pierfrancesco Foglia and Marco Solinas. Exploiting replication to improve performances of NUCAbased CMP systems. ACM Transactions on Embedded Computing Systems, 13(3s): 117:1–117:??, March 2014. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Fatnassi:2024:PNN

[FS24] Wael Fatnassi and Yasser Shoukry. PolyARBerNN: a neural network guided solver and optimizer for bounded polynomial inequalities. ACM Transactions on Embedded Computing Systems, 23(2):22:1–22:??, March 2024. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL

https://dl.acm.org/doi/
10.1145/3632970.

Forsberg:2021:PEM

[FSB^{+}21] Björn Forsberg, Marco Solieri, Marko Bertogna, Luca Benini, and Andrea Marongiu. The predictable execution model in practice: Compiling real applications for COTS hardware. ACM Transactions on Embedded Computing Systems, 20(5):47:1–47:25, July 2021. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3465370.

Filippopoulos:2016:IEM

[FSC^{+}16] Iason Filippopoulos, Namita Sharma, Francky Catthoor, Per Gunnar Kjeldsberg, and Preeti Ranjan Panda. Integrated exploration methodology for data interleaving and data-to-memory mapping on SIMD architectures. ACM Transactions on Embedded Computing Systems, 15(3):59:1–59:??, July 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Feldtkeller:2023:COS

[FSG23] Jakob Feldtkeller, Pascal Sasdrich, and Tim Güneysu. Challenges and opportunities of security-aware EDA. ACM Transactions on Embedded Computing Systems, 22(3):43:1–43:??, May 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3576199.

Fong:2019:ODS

[FSVG19] Daniel D. Fong, Vivek J. Srinivasan, Kourosh Vali, and Soheil Ghiasi. Optode design space exploration for clinically-robust non-invasive fetal oximetry. ACM Transactions on Embedded Computing Systems, 18(5s):63:1–63:??, October 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

                                        /dl.acm.org/ft_gateway.
                                        cfm?id=3358207.

Fischmeister:2017:GES

[FX17] Sebastian Fischmeister and Jason Xue. Guest editorial: Special issue on LCTES 2015. ACM Transactions on Embedded Computing Systems, 16(2):29:1–29:??, April 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Fu:2017:DFA

[FXP^{+}17] Shan Fu, Guoai Xu, Juan Pan, Zongyue Wang, and An Wang. Differential fault attack on ITUbee block cipher. ACM Transactions on Embedded Computing Systems, 16(2):54:1–54:??, April 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Falk:2013:RBQ

[FZHT13] Joachim Falk, Christian Zebelein, Christian Haubelt, and Jürgen Teich. A rulebased quasi-static scheduling approach for static islands in dynamic dataflow graphs. ACM Transactions on Embedded Computing Systems, 12(3):74:1– 74:??, March 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Fei:2008:EAF

[FZJ08] Yunsi Fei, Lin Zhong, and Niraj K. Jha. An energyaware framework for dynamic software management in mobile computing systems. ACM Transactions on Embedded Computing Systems, 7(3):27:1–27:??, April 2008. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Falk:2010:ASA

[FZK^{+}10] Joachim Falk, Christian Zebelein, Joachim Keinert, Christian Haubelt, Juergen Teich, and Shuvra S. Bhattacharyya. Analysis of SystemC actor networks for efficient synthesis. ACM Transactions on Embedded Computing Systems, 10(2):18:1– 18:??, December 2010. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Gaitan:2024:MOC

[GAD^{+}24] Miguel Gutiérrez Gaitán, Luís Almeida, Pedro M. D’orey, Pedro M. Santos, and Thomas Watteyne. Minimal-overlap centrality for multi-gateway designation in real-time TSCH networks. ACM Transactions on Embedded Computing Systems, 23(1):9:1– 9:??, January 2024. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3610583.

Gomony:2015:RTM

[GAG15] Manil Dev Gomony, Benny Akesson, and Kees Goossens. A real-time multichannel memory controller and optimal mapping of memory clients to memory channels. ACM Transactions on Embedded Computing Systems, 14(2):25:1–25:??, March 2015. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Gardner:2005:CCS

[Gar05] William B. Gardner. Converging CSP specifications and C++ programming via selective formalism. ACM Transactions on Embedded Computing Systems, 4(2): 302–330, May 2005. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Gottscho:2017:LCM

[GAS^{+}17] Mark Gottscho, Irina Alam, Clayton Schoeny, Lara Dolecek, and Puneet Gupta. Low-cost memory fault tolerance for IoT devices. ACM Transactions on Embedded Computing Systems, 16(5s): 128:1–128:??, October 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Ghosh:2020:RSD

[GCJD20] Saurav Kumar Ghosh, Jaffer Sheriff R. C., Vibhor Jain, and Soumyajit Dey. Reliable and secure designspace-exploration for cyberphysical systems. ACM Transactions on Embedded Computing Systems, 19(3): 21:1–21:29, July 2020. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3387927.

Gupta:2024:MEE

[GCS24] Saransh Gupta, Rosario Cammarota, and Tajana Simunić. MemFHE: Endto-end computing with fully homomorphic encryption in memory. ACM Transactions on Embedded Computing Systems, 23(2):28:1– 28:??, March 2024. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3569955.

Gunzel:2023:CTA

[GCU^{+}23] Mario Günzel, Kuan-Hsun Chen, Niklas Ueter, Georg von der Brüggen, Marco Dürr, and Jian-Jia Chen. Compositional timing analysis of asynchronized distributed cause-effect chains. ACM Transactions on Embedded Computing Systems, 22(4):63:1–63:??, July 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3587036.

Gong:2014:SBF

[GD14] Lingkan Gong and Oliver Diessel. Simulation-based functional verification of dynamically reconfigurable systems. ACM Transactions on Embedded Computing Systems, 13(4):97:1– 97:??, February 2014. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Ghosh:2019:RRS

[GD19] Bineet Ghosh and Parasara Sridhar Duggirala. Robust reachable set: Accounting for uncertainties in linear dynamical systems. ACM Transactions on Embedded Computing Systems, 18(5s):97:1– 97:??, October 2019. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3358229.

Goli:2024:ESI

[GD24] Mehran Goli and Rolf Drechsler. Early SoCs information flow policies validation using SystemC-based virtual prototypes at the ESL. ACM Transactions on Embedded Computing Systems, 23(5):67:1–67:??, September 2024. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3544780.

Godary-Dejean:2013:FVD

[GDA13] Karen Godary-Dejean and David Andreu. Formal validation of a deterministic MAC protocol. ACM Transactions on Embedded Computing Systems, 12(1):6:1– 6:??, January 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Goyal:2022:HFU

[GDB22] Vidushi Goyal, Reetuparna Das, and Valeria Bertacco. Hardware-friendly user-specific machine learning for edge devices. ACM Transactions on Embedded Computing Systems, 21(5):62:1– 62:??, September 2022. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3524125.

Goncalves:2019:AER

[GDC19] Larissa Rozales Gonçalves, Rafael Fão De Moura, and Luigi Carro. Aggressive energy reduction for video inference with software-only strategies. ACM Transactions on Embedded Computing Systems, 18(5s):46:1– 46:??, October 2019. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

                                        tronic). URL https://dl.
                                        acm.org/ft_gateway.cfm?
                                        id=3358174.

Ghosh:2020:PGI

[GDD20] Sumana Ghosh, Soumyajit Dey, and Pallab Dasgupta. Pattern guided integrated scheduling and routing in multi-hop control networks. ACM Transactions on Embedded Computing Systems, 19(2):9:1–9:28, March 2020. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3372134.

Ghosh:2017:SMP

[GDDD17] Sumana Ghosh, Souradeep Dutta, Soumyajit Dey, and Pallab Dasgupta. A structured methodology for pattern based adaptive scheduling in embedded control. ACM Transactions on Embedded Computing Systems, 16(5s):189:1–189:??, October 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Grun:2003:APB

[GDN03] Peter Grun, Nikil Dutt, and Alex Nicolau. Access pattern-based memory and connectivity architecture exploration. ACM Transactions on Embedded Computing Systems, 2(1):33–73, February 2003. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Gu:2018:EST

[GE18] Xiaozhe Gu and Arvind Easwaran. Efficient schedulability test for dynamicpriority scheduling of mixedcriticality real-time systems. ACM Transactions on Embedded Computing Systems, 17(1):24:1–24:??, January 2018. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Gebotys:2004:DSC

[Geb04] Catherine H. Gebotys. Design of secure cryptography against the threat of powerattacks in DSP-embedded processors. ACM Transactions on Embedded Computing Systems, 3(1):92–113, February 2004. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Gebotys:2006:SMC

[Geb06] Catherine H. Gebotys. A split-mask countermeasure for low-energy secure embedded systems. ACM Transactions on Embedded Computing Systems, 5(3):577– 612, August 2006. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Geilen:2010:SDS

[Gei10] Marc Geilen. Synchronous dataflow scenarios. ACM Transactions on Embedded Computing Systems, 10(2): 16:1–16:??, December 2010. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Geelen:2010:MES

[GFC^{+}10] Bert Geelen, Vissarion Ferentinos, Francky Catthoor, Gauthier Lafruit, Diederik Verkest, Rudy Lauwereins, and Thanos Stouraitis. Modeling and exploiting spatial locality trade-offs in wavelet-based applications under varying resource requirements. ACM Transactions on Embedded Computing Systems, 9(3):17:1– 17:??, February 2010. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Gealy:2024:CPS

[GG24] Calvin B. Gealy and Alan D.

George. Characterizing parameter scaling with quantization for deployment of CNNs on real-time systems. ACM Transactions on Embedded Computing Systems, 23(3):38:1–38:??, May 2024. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3654799.

Get:2008:PFE

[GGGK08] Jerome Hugues Get, Bechir Zalila Get, Laurent Pautet Get, and Fabrice Kordon. From the prototype to the final embedded system using the Ocarina AADL tool suite. ACM Transactions on Embedded Computing Systems, 7(4):42:1–42:??, July 2008. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Gu:2013:DDL

[GGI13] Ji Gu, Hui Guo, and Tohru Ishihara. DLIC: Decoded loop instructions caching for energy-aware embedded processors. ACM Transactions on Embedded Computing Systems, 13(1):6:1– 6:??, August 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Guenterberg:2012:ASR

[GGJ12] Eric Guenterberg, Hassan Ghasemzadeh, and Roozbeh Jafari. Automatic segmentation and recognition in body sensor networks using a hidden Markov model. ACM Transactions on Embedded Computing Systems, 11(S2):46:1–46:??, 2012. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Gretz:2024:SSM

[GGMS24] Friedrich Gretz, Franz-Josef Grosch, Michael Mendler, and Stephan Scheele. Synchronized shared memory and black-box procedural abstraction: Toward a formal semantics of blech. ACM Transactions on Embedded Computing Systems, 23(5): 76:1–76:??, September 2024. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3571585.

Geuns:2013:SST

[GHB13] Stefan J. Geuns, Joost P.

H. M. Hausmans, and Marco J. G. Bekooij. Sequential specification of timeaware stream processing applications. ACM Transactions on Embedded Computing Systems, 12(1s):35:1– 35:??, March 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Garbay:2025:ZCD

[GHD^{+}25] Thomas Garbay, Khalil Hachicha, Petr Dobias, Andrea Pinna, Karim Hocine, Wilfried Dron, Pedro Lusich, Imane Khalis, and Bertrand Granado. ZIP-CNN: Design space exploration for CNN implementation within a MCU. ACM Transactions on Embedded Computing Systems, 24(1):5:1– 5:??, January 2025. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3691343.

Germanos:2015:DUW

[GHKS15] Vasileios Germanos, Stefan Haar, Victor Khomenko, and Stefan Schwoon. Diagnosability under weak fairness. ACM Transactions on Embedded Computing Systems, 14(4):69:1–69:??, December 2015. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Guo:2018:CSP

[GHPP18] Danlu Guo, Mohamed Hassan, Rodolfo Pellizzoni, and Hiren Patel. A comparative study of predictable DRAM controllers. ACM Transactions on Embedded Computing Systems, 17(2): 53:1–53:??, April 2018. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Geeraerts:2015:VCA

[GHR15] Gilles Geeraerts, Alexander Heußner, and Jean-François Raskin. On the verification of concurrent, asynchronous programs with waiting queues. ACM Transactions on Embedded Computing Systems, 14(3):58:1– 58:??, May 2015. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Gu:2014:AES

[GHZH14] Yu Gu, Liang He, Ting Zhu, and Tian He. Achieving energy-synchronized communication in energy-harvesting wireless sensor networks. ACM Transactions on Embedded Computing Systems, 13(2s):68:1–68:??, January 2014. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Gauthier:2024:HRE

[GI24] Lovic Gauthier and Yohei Ishikawa. HDLRuby: a Ruby extension for hardware description and its translation to synthesizable Verilog HDL. ACM Transactions on Embedded Computing Systems, 23(5):73:1– 73:??, September 2024. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3581757.

Gilroy:2011:RHA

[GIA11] Michael Gilroy, James Irvine, and Robert Atkinson. RAID 6 hardware acceleration. ACM Transactions on Embedded Computing Systems, 10(4):43:1–43:??, November 2011. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Girodias:2012:IMO

[GIB^{+}12] Bruno Girodias, Luiza Gheorghe Iugan, Youcef Bouchebaba, Gabriela Nicolescu, El Mostapha Abouhamid, Michel Langevin, and Pierre Paulin. Integrating memory optimization with mapping algorithms for multi-processors system-on-chip. ACM Transactions on Embedded Computing Systems, 11(3):64:1– 64:??, September 2012. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Ghasemzadeh:2013:ULP

[GJ13] Hassan Ghasemzadeh and Roozbeh Jafari. Ultra lowpower signal processing in wearable monitoring systems: a tiered screening architecture with optimal bit resolution. ACM Transactions on Embedded Computing Systems, 13(1):9:1– 9:??, August 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Ganesan:2022:DST

[GK22] Vinod Ganesan and Pratyush Kumar. Design and scaffolded training of an efficient DNN operator for computer vision on the edge. ACM Transactions on Embedded Computing Systems, 21(6):72:1–72:??, November 2022. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3511212.

Gupta:2022:SLC

[GKS^{+}22] Saransh Gupta, Behnam Khaleghi, Sahand Salamat, Justin Morris, Ranganathan Ramkumar, Jeffrey Yu, Aniket Tiwari, Jaeyoung Kang, Mohsen Imani, Baris Aksanli, and Tajana Simunić Rosing. Store-n-learn: Classification and clustering with hyperdimensional computing across flash hierarchy. ACM Transactions on Embedded Computing Systems, 21(3):22:1–22:25, May 2022. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3503541.

Goel:2025:EFE

[GKSB25] Shikha Goel, Rajesh Kedia, Rijurekha Sen, and M. Balakrishnan. EX-PRESS: a framework for execution time prediction of concurrent CNNs on Xilinx DPU accelerator. ACM Transactions on Embedded Computing Systems, 24(1): 11:1–11:??, January 2025. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3697835.

Gurun:2008:NGP

[GKW08] Selim Gurun, Chandra Krintz, and Rich Wolski. NWSLite: a generalpurpose, nonparametric prediction utility for embedded systems. ACM Transactions on Embedded Computing Systems, 7(3):32:1– 32:??, April 2008. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Gay:2007:SDP

[GLC07] David Gay, Philip Levis, and David Culler. Software design patterns for TinyOS. ACM Transactions on Embedded Computing Systems, 6(4):22:1–22:??, September 2007. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Gaglio:2018:DPD

[GLMP18] Salvatore Gaglio, Giuseppe Lo Re, Gloria Martorella, and Daniele Peri. DC4CD: a platform for distributed computing on constrained devices. ACM Transactions on Embedded Computing Systems, 17(1):27:1– 27:??, January 2018. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Gamatie:2011:MDD

[GLP^{+}11] Abdoulaye Gamatié, Sébastien Le Beux, Éric Piel, Rabie Ben Atitallah, Anne Etien, Philippe Marquet, and Jean-Luc Dekeyser. A modeldriven design framework for massively parallel embedded systems. ACM Transactions on Embedded Computing Systems, 10(4):39:1– 39:??, November 2011. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Gubbi:2023:HTD

[GLS^{+}23] Kevin Immanuel Gubbi, Banafsheh Saber Latibari, Anirudh Srikanth, Tyler Sheaves, Sayed Arash Beheshti-Shirazi, Sai Manoj PD, Satareh Rafatirad, Avesta Sasan, Houman Homayoun, and Soheil Salehi. Hardware Trojan detection using machine learning: a tutorial. ACM Transactions on Embedded Computing Systems, 22(3):46:1–46:??, May 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3579823.

Ghiribaldi:2013:CST

[GLT^{+}13] Alberto Ghiribaldi, Daniele Ludovici, Francisco Triviño, Alessandro Strano, José Flich, José Luis Sánchez, Francisco Alfaro, Michele Favalli, and Davide Bertozzi. A complete self-testing and self-configuring NoC infrastructure for cost-effective MPSoCs. ACM Transactions on Embedded Computing Systems, 12(4):106:1–106:??, June 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). See comment [BLG^{+}15].

Guimbretiere:2014:ADP

[GLWM14] François Guimbretiére, Shenwei Liu, Han Wang, and Rajit Manohar. An asymmetric dual-processor architecture for low-power information appliances. ACM Transactions on Embedded Computing Systems, 13(4):98:1– 98:??, February 2014. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Guan:2014:WAM

[GLYY14] Nan Guan, Mingsong Lv, Wang Yi, and Ge Yu. WCET analysis with MRU cache: Challenging LRU for predictability. ACM Transactions on Embedded Computing Systems, 13(4s):123:1– 123:??, April 2014. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Guang:2003:SIC

[GM03] Gao Guang and Trevor Mudge. Special issue on compilers, architecture, and synthesis for embedded systems. ACM Transactions on Embedded Computing Systems, 2(2):131, May 2003. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Garg:2012:IMP

[GM12] Siddharth Garg and Diana Marculescu. On the impact of manufacturing process variations on the lifetime of sensor networks. ACM Transactions on Embedded Computing Systems, 11(2):33:1–33:??, July 2012. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Guo:2018:IWP

[GMCC18] Jie Guo, Chuhan Min, Tao Cai, and Yiran Chen. Improving write performance and extending endurance of object-based NAND flash devices. ACM Transactions on Embedded Computing Systems, 17(1):18:1– 18:??, January 2018. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Garcia:2021:IHG

[GMN21] Andrés Amaya García, David May, and Ed Nutting. Integrated hardware garbage collection. ACM Transactions on Embedded Computing Systems, 20(5):40:1– 40:25, July 2021. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3450147.

Gohringer:2013:RAN

[GMOB13] Diana Göhringer, Lukas Meder, Oliver Oey, and Jürgen Becker. Reliable and adaptive network-on-chip architectures for cyber physical systems. ACM Transactions on Embedded Computing Systems, 12(1s):51:1– 51:??, March 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Gavran:2017:AMR

[GMS17] Ivan Gavran, Rupak Majumdar, and Indranil Saha. Antlab: a multi-robot task server. ACM Transactions on Embedded Computing Systems, 16(5s):190:1– 190:??, October 2017. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Groza:2017:LCL

[GMVV17] Bogdan Groza, Stefan Murvay, Anthony Van Herrewege, and Ingrid Verbauwhede. LiBrA–CAN: Lightweight broadcast authentication for controller area networks. ACM Transactions on Embedded Computing Systems, 16(3):90:1– 90:??, July 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Girault:2006:ARD

[GNP06] Alain Girault, Xavier Nicollin, and Marc Pouzet. Automatic rate desynchronization of embedded reactive programs. ACM Transactions on Embedded Computing Systems, 5(3):687– 717, August 2006. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Guang:2010:HAM

[GNR^{+}10] Liang Guang, Ethiopia Nigussie, Pekka Rantala, Jouni Isoaho, and Hannu Tenhunen. Hierarchical agent monitoring design approach towards self-aware parallel systems-on-chip. ACM Transactions on Embedded Computing Systems, 9(3): 25:1–25:??, February 2010. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Ghiasi:2004:OAM

[GNS04] Soheil Ghiasi, Ani Nahapetian, and Majid Sarrafzadeh. An optimal algorithm for minimizing runtime reconfiguration delay. ACM Transactions on Embedded Computing Systems, 3(2):237–256, May 2004. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Gaujal:2005:SPA

[GNW05] Bruno Gaujal, Nicolas Navet, and Cormac Walsh. Shortestpath algorithms for realtime scheduling of FIFO tasks with minimal energy use. ACM Transactions on Embedded Computing Systems, 4(4):907–933, November 2005. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

George:2022:UPE

[GOC^{+}22] Biji George, Om Ji Omer, Ziaul Choudhury, Anoop V, and Sreenivas Subramoney. A unified programmable edge matrix processor for deep neural networks and matrix algebra. ACM Transactions on Embedded Computing Systems, 21(5):63:1–63:??, September 2022. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3524453.

Goehringer:2014:ISI

[Goe14] Diana Goehringer. Introduction to the Special Issue on Virtual Prototyping of Parallel and Embedded Systems (ViPES). ACM Transactions on Embedded Computing Systems, 13(5s):164:1– 164:??, November 2014. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Gupta:2007:ISL

[GP07] Rajiv Gupta and Yunheung Paek. Introduction to the special LCTES’05 issue. ACM Transactions on Embedded Computing Systems, 6(4):21:1–21:??, September 2007. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Gufran:2023:FHR

[GP23] Danish Gufran and Sudeep Pasricha. FedHIL: Heterogeneity resilient federated learning for robust indoor localization with mobile devices. ACM Transactions on Embedded Computing Systems, 22(5s):125:1– 125:??, October 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3607919.

Gupta:2017:DDP

[GPB^{+}17] Ujjwal Gupta, Chetan Arvind Patil, Ganapati Bhat, Prabhat Mishra, and Umit Y. Ogras. DyPO: Dynamic Pareto-optimal configuration selection for heterogeneous MpSoCs. ACM Transactions on Embedded Computing Systems, 16(5s):123:1– 123:??, October 2017. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Gomez:2024:ODF

[GPH24] Clara Gomez, Davron Patkhullaev, and Alejandra C. Hernandez. OffloaD: Detection failure-based scheduler for offloading object detection. ACM Transactions on Embedded Computing Systems, 23(5):84:1–84:??, September 2024. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3677321.

Gkeka:2023:RSC

[GPT^{+}23] Maria Rafaela Gkeka, Alexandros Patras, Nikolaos Tavoularis, Stylianos Piperakis, Emmanouil Hourdakis, Panos Trahanias, Christos D. Antonopoulos, Spyros Lalis, and Nikolaos Bellas. Reconfigurable system-on-chip architectures for robust visual SLAM on humanoid robots. ACM Transactions on Embedded Computing Systems, 22(2):24:1– 24:??, March 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3570210.

Gai:2017:SES

[GQC^{+}17] Keke Gai, Longfei Qiu, Min Chen, Hui Zhao, and Meikang Qiu. SA–EAST: Security-aware efficient data transmission for ITS in mobile heterogeneous cloud computing. ACM Transactions on Embedded Computing Systems, 16(2):60:1– 60:??, April 2017. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Gordon-Ross:2003:TIC

[GRCV03] Ann Gordon-Ross, Susan Cotterell, and Frank Vahid. Tiny instruction caches for low power embedded systems. ACM Transactions on Embedded Computing Systems, 2(4):449–481, November 2003. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Ghosh:2023:EEA

[GRR23] Soumendu Kumar Ghosh, Arnab Raha, and Vijay Raghunathan. Energyefficient approximate edge inference systems. ACM Transactions on Embedded Computing Systems, 22(4):

77:1–77:??, July 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3589766.

Ghosh:2024:PPA

[GRRR24] Soumendu Kumar Ghosh, Arnab Raha, Vijay Raghunathan, and Anand Raghunathan. PArtNNer: Platformagnostic adaptive edgecloud DNN partitioning for minimizing end-to-end latency. ACM Transactions on Embedded Computing Systems, 23(1):6:1–6:??, January 2024. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL

https://dl.acm.org/doi/
10.1145/3630266.

Gordon-Ross:2012:CCR

[GRVD12] Ann Gordon-Ross, Frank Vahid, and Nikil Dutt. Combining code reordering and cache configuration. ACM Transactions on Embedded Computing Systems, 11(4): 88:1–88:??, December 2012. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Ghasemi:2022:EEE

[GRWV22] Mehdi Ghasemi, Daler Rakhmatov, Carole-Jean Wu, and Sarma Vrudhula. EdgeWise: Energy-efficient CNN computation on edge devices under stochastic communication delays. ACM Transactions on Embedded Computing Systems, 21(5):66:1– 66:??, September 2022. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3530908.

Guha:2019:SBS

[GSC19] Krishnendu Guha, Debasri Saha, and Amlan Chakrabarti. Stigmergybased security for SoC operations from runtime performance degradation of SoC components. ACM Transactions on Embedded Computing Systems, 18(2):14:1– 14:??, April 2019. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

                                        tronic). URL https://dl.
                                        acm.org/ft_gateway.cfm?
                                        id=3301279.

Gressl:2021:DSE

[GSN21] Lukas Gressl, Christian Steger, and Ulrich Neffe. Design space exploration for secure IoT devices and cyberphysical systems. ACM Transactions on Embedded Computing Systems, 20(4): 32:1–32:24, June 2021. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3430372.

Gomez:2018:ELT

[GSS^{+}18] Andres Gomez, Lukas Sigrist, Thomas Schalch, Luca Benini, and Lothar Thiele. Efficient, long-term logging of rich data sensors using transient sensor nodes. ACM Transactions on Embedded Computing Systems, 17(1): 4:1–4:??, January 2018. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Grimm:2024:LGM

[GSSR^{+}24] Lena Grimm, Steven Smyth, Alexander Schulz-Rosengarten, Reinhard von Hanxleden, and Marc Pouzet. From Lustre to graphical models and SCCharts. ACM Transactions on Embedded Computing Systems, 23(5): 66:1–66:??, September 2024. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3544973.

Grimheden:2005:WES

[GT05] Martin Grimheden and Martin Törngren. What is embedded systems and how should it be taught?—results from a didactic analysis. ACM Transactions on Embedded Computing Systems, 4(3):633–651, August 2005. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Gomez:2022:DDP

[GTH^{+}22] Andres Gomez, Andreas Tretter, Pascal Alexander Hager, Praveenth Sanmugarajah, Luca Benini, and Lothar Thiele. Dataflow driven partitioning of machine learning applications for optimal energy use in batteryless systems. ACM Transactions on Embedded Computing Systems, 21(5): 54:1–54:??, September 2022. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3520135.

Gunzel:2025:EEL

[GTvdBC25] Mario Günzel, Harun Teper, Georg von der Brüggen, and Jian-Jia Chen. Endto-end latency of cause– effect chains: a tutorial. ACM Transactions on Embedded Computing Systems, 24(1):22:1–22:??, January 2025. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3703630.

Gunzel:2023:PRT

[GUC^{+}23] Mario Günzel, Niklas Ueter, Kuan-Hsun Chen, Georg von der Brüggen, and Jian-Jia Chen. Probabilistic reaction time analysis. ACM Transactions on Embedded Computing Systems, 22(5s): 143:1–143:??, October 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3609390.

Gupta:2004:GES

[Gup04] Rajesh Gupta. Guest editorial: Special issue on networked embedded systems. ACM Transactions on Embedded Computing Systems, 3(1):1–2, February 2004. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Giraldo:2021:HAE

[GV21a] J. S. P. Giraldo and Marian Verhelst. Hardware acceleration for embedded keyword spotting: Tutorial and survey. ACM Transactions on Embedded Computing Systems, 20(6):111:1– 111:25, November 2021. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3474365.

Girault:2021:ISI

[GV21b] Alain Girault and Reinhard Von Hanxleden. Introduction to the special issue on Specification and Design Languages (FDL 2019). ACM Transactions on Embedded Computing Systems, 20(4):27:1–27:3, June 2021. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3458748.

Ganapathy:2020:DDV

[GVS^{+}20] Sanjay Ganapathy, Swagath Venkataramani, Giridhur Sriraman, Balaraman Ravindran, and Anand Raghunathan. DyVEDeep: Dynamic variable effort deep neural networks. ACM Transactions on Embedded Computing Systems, 19(3): 16:1–16:24, July 2020. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3372882.

Gebotys:2008:EAW

[GW08] Catherine H. Gebotys and Brian A. White. EM analysis of a wireless Java-based PDA. ACM Transactions on Embedded Computing Systems, 7(4):44:1–44:??, July 2008. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Gebotys:2015:SWP

[GW15] Catherine H. Gebotys and Brian A. White. A sliding window phase-only correlation method for side-channel alignment in a Smartphone. ACM Transactions on Embedded Computing Systems, 14(4):80:1–80:??, December 2015. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Gao:2024:SSW

[GWC^{+}24] Xin Gao, Hongyue Wang, Yiyan Chen, Yuhao Zhang, Zhaoyan Shen, and Lei Ju. Static scheduling of weight programming for DNN acceleration with resource constrained PIM. ACM Transactions on Embedded Computing Systems, 23(6):89:1– 89:??, November 2024. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3615657.

Gebotys:2016:PCP

[GWM16] Catherine H. Gebotys, Brian A.

White, and Edgar Mateos. Preaveraging and carry propagate approaches to side-channel analysis of HMAC-SHA256. ACM Transactions on Embedded Computing Systems, 15(1): 4:1–4:??, February 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Gu:2016:CPP

[GWZ16] Zonghua Gu, Chao Wang, and Haibo Zeng. Cachepartitioned preemption threshold scheduling. ACM Transactions on Embedded Computing Systems, 16(1):13:1– 13:??, November 2016. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Girard:2012:VSL

[GZ12] Antoine Girard and Gang Zheng. Verification of safety and liveness properties of metric transition systems. ACM Transactions on Embedded Computing Systems, 11(S2):54:1–54:??, 2012. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Gu:2016:RTF

[GZZ^{+}16] Xiaoqi Gu, Yongxin Zhu, Shengyan Zhou, Chaojun Wang, Meikang Qiu, and Guoxing Wang. A real-time FPGA-Based accelerator for ECG analysis and diagnosis using association-rule mining. ACM Transactions on Embedded Computing Systems, 15(2):25:1–25:??, May 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Hafezan:2024:TFD

[HA24] Mohammad Hassan Hafezan and Ehsan Atoofian. Transient fault detection in tensor cores for modern GPUs. ACM Transactions on Embedded Computing Systems, 23(5):82:1–82:??, September 2024. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3687483.

Huynh:2011:EAR

[HABT11] Johnny Huynh, José Nelson Amaral, Paul Berube, and Sid-Ahmed-Ali Touati. Evaluating address register assignment and offset assignment algorithms. ACM Transactions on Embedded Computing Systems, 10(3): 37:1–37:??, April 2011. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Hilal:2016:CEA

[HB16] Allaa R. Hilal and Otman Basir. A collaborative energy-aware sensor management system using team theory. ACM Transactions on Embedded Computing Systems, 15(3):52:1–52:??, July 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Hussein:2023:CNC

[HB23] Dina Hussein and Ganapati Bhat. CIM: a novel clustering-based energy-efficient data imputation method for human activity recognition. ACM Transactions on Embedded Computing Systems, 22(5s):116:1–116:??, October 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3609111.

Hussein:2024:SND

[HB24] Dina Hussein and Ganapati Bhat. SensorGAN: a novel data recovery approach for wearable human activity recognition. ACM Transactions on Embedded Computing Systems, 23(3):53:1– 53:??, May 2024. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3609425.

He:2004:AAA

[HBSA04] Tian He, Brian M. Blum, John A. Stankovic, and Tarek Abdelzaher. AIDA: Adaptive application-independent data aggregation in wireless sensor networks. ACM Transactions on Embedded Computing Systems, 3(2): 426–457, May 2004. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Huang:2016:EPC

[HC16] Sheng-Min Huang and Li-Pin Chang. Exploiting page correlations for write buffering in page-mapping multichannel SSDs. ACM Transactions on Embedded Computing Systems, 15(1): 12:1–12:??, February 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Ha:2008:IES

[HCK^{+}08] Soonhoi Ha, Kiyoung Choi, Taewhan Kim, Krisztian Flautner, Sanglyul Min, and Wang Yi. Introduction to embedded systems week 2006 special issue. ACM Transactions on Embedded Computing Systems, 7(2): 8:1–8:??, February 2008. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Huang:2017:PSV

[HCL^{+}17] Chao Huang, Xin Chen, Wang Lin, Zhengfeng Yang, and Xuandong Li. Probabilistic safety verification of stochastic hybrid systems using barrier certificates. ACM Transactions on Embedded Computing Systems, 16(5s): 186:1–186:??, October 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Huang:2014:TMP

[HCQ^{+}14] Huang Huang, Vivek Chaturvedi, Gang Quan, Jeffrey Fan, and Meikang Qiu. Throughput maximization for periodic real-time systems under the maximal temperature constraint. ACM Transactions on Embedded Computing Systems, 13(2s):70:1– 70:??, January 2014. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

He:2018:AFI

[HCS18] Zhijian He, Yao Chen, and Zhaoyan Shen. Attitude fusion of inertial and magnetic sensor under different magnetic filed distortions. ACM Transactions on Embedded Computing Systems, 17(2):48:1–48:??, April 2018. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Hakert:2022:SMR

[HCS^{+}22] Christian Hakert, Kuan-Hsun Chen, Horst Schirmeier, Lars Bauer, Paul R. Genssler, Georg von der Brüggen, Hussam Amrouch, Jörg Henkel, and Jian-Jia Chen. Softwaremanaged read and write wear-leveling for non-volatile main memory. ACM Transactions on Embedded Computing Systems, 21(1):5:1– 5:24, January 2022. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3483839.

Hanumaiah:2014:SST

[HDG^{+}14] Vinay Hanumaiah, Digant Desai, Benjamin Gaudette, Carole-Jean Wu, and Sarma Vrudhula. STEAM: a smart temperature and energy aware multicore controller. ACM Transactions on Embedded Computing Systems, 13(5s):151:1– 151:??, September 2014. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Hessel:2006:SRA

[HDR^{+}06] Fabiano Hessel, Vitor M.

Da Rosa, Carlos Eduardo Reif, César Marcon, and Tatiana Gadelha Serra Dos Santos. Scheduling refinement in abstract RTOS models. ACM Transactions on Embedded Computing Systems, 5(2):342–354, May 2006. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

He:2020:BCL

[HDZL20] Wenjian He, Sanjeev Das, Wei Zhang, and Yang Liu. BBB-CFI: Lightweight CFI approach against code-reuse attacks using basic block information. ACM Transactions on Embedded Computing Systems, 19(1):7:1– 7:22, February 2020. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3371151.

Hamers:2012:EMS

[HE12] Juan Hamers and Lieven Eeckhout. Exploiting media stream similarity for energyefficient decoding and resource prediction. ACM Transactions on Embedded Computing Systems, 11(1): 2:1–2:??, March 2012. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Hui:2024:SSM

[HE24] John Hui and Stephen A. Edwards. The sparse synchronous model on real hardware. ACM Transactions on Embedded Computing Systems, 23(5):69:1– 69:??, September 2024. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3572920.

Hettiarachchi:2014:DAF

[HFA^{+}14] Pradeep M. Hettiarachchi, Nathan Fisher, Masud Ahmed, Le Yi Wang, Shinan Wang, and Weisong Shi. A design and analysis framework for thermal-resilient hard realtime systems. ACM Transactions on Embedded Computing Systems, 13(5s):146:1– 146:??, September 2014. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Hashemi:2013:TMF

[HFG13] Matin Hashemi, Mohammad H. Foroozannejad, and Soheil Ghiasi. Throughputmemory footprint trade-off in synthesis of streaming software on embedded multiprocessors. ACM Transactions on Embedded Computing Systems, 13(3):46:1– 46:??, December 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Huang:2019:RRA

[HFL^{+}19] Chao Huang, Jiameng Fan, Wenchao Li, Xin Chen, and Qi Zhu. ReachNN: Reachability analysis of neuralnetwork controlled systems. ACM Transactions on Embedded Computing Systems, 18(5s):106:1–106:??, October 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/ft_gateway.
cfm?id=3358228.

Hashemi:2009:TDS

[HG09] Matin Hashemi and Soheil Ghiasi. Throughputdriven synthesis of embedded software for pipelined execution on multicore architectures. ACM Transactions on Embedded Computing Systems, 8(2):11:1– 11:??, January 2009. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Huang:2014:IEM

[HGL14] Huang-Ming Huang, Christopher Gill, and Chenyang Lu. Implementation and evaluation of mixed-criticality scheduling approaches for sporadic tasks. ACM Transactions on Embedded Computing Systems, 13(4s):

126:1–126:??, April 2014. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). See corrections and comments [FHB^{+}17].

Han:2020:BAP

[HGW^{+}20] Jian-Jun Han, Sunlu Gong, Zhenjiang Wang, Wen Cai, Dakai Zhu, and Laurence T. Yang. Blocking-aware partitioned real-time scheduling for uniform heterogeneous multicore platforms. ACM Transactions on Embedded Computing Systems, 19(1): 1:1–1:25, February 2020. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3366683.

Hsieh:2013:TAM

[HH13] Ang-Chih Hsieh and Tingting Hwang. Thermal-aware memory mapping in 3D designs. ACM Transactions on Embedded Computing Systems, 13(1):4:1–4:??, August 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Hessien:2023:PPS

[HH23] Salah Hessien and Mohamed Hassan. PISCOT: a pipelined split-transaction COTS-Coherent bus for multi-core real-time systems. ACM Transactions on Embedded Computing Systems, 22(1):16:1–16:??, January 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3556975.

He:2005:RFL

[HHB^{+}05] Tian He, Chengdu Huang, Brian M. Blum, John A. Stankovic, and Tarek F. Abdelzaher. Range-free localization and its impact on large scale sensor networks. ACM Transactions on Embedded Computing Systems, 4(4):877– 906, November 2005. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Huang:2012:EFP

[HHB^{+}12] Kai Huang, Wolfgang Haid, Iuliana Bacivarov, Matthias Keller, and Lothar Thiele. Embedding formal performance analysis into the design cycle of MPSoCs for real-time streaming applications. ACM Transactions on Embedded Computing Systems, 11(1):8:1–8:??, March 2012. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Hu:2016:AWM

[HHC^{+}16a] Biao Hu, Kai Huang, Gang Chen, Long Cheng, and Alois Knoll. Adaptive workload management in mixedcriticality systems. ACM Transactions on Embedded Computing Systems, 16(1): 14:1–14:??, November 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Hu:2016:EIR

[HHC^{+}16b] Biao Hu, Kai Huang, Gang Chen, Long Cheng, and Alois Knoll. Evaluation and improvements of runtime monitoring methods for realtime event streams. ACM Transactions on Embedded Computing Systems, 15(3): 56:1–56:??, July 2016. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Hu:2023:HPR

[HHL^{+}23a] Xianghong Hu, Hongmin Huang, Xueming Li, Xin Zheng, Qinyuan Ren, Jingyu He, and Xiaoming Xiong. High-performance reconfigurable DNN accelerator on a bandwidth-limited embedded system. ACM Transactions on Embedded Computing Systems, 22(6):90:1– 90:??, November 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3530818.

Hu:2023:HPI

[HHL^{+}23b] Xinyi Hu, Debiao He, Min Luo, Cong Peng, Qi Feng, and Xinyi Huang. Highperformance implementation of the identity-based signature scheme in IEEE P1363 on GPU. ACM Transactions on Embedded Computing Systems, 22(2):25:1– 25:??, March 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3564784.

Harb:2019:FIE

[HJ19] Salah Harb and Moath Jarrah. FPGA implementation of the ECC over GF(2^{m}) for small embedded applications. ACM Transactions on Embedded Computing Systems, 18(2):17:1–17:??, April 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/ft_gateway.
cfm?id=3310354.

Ha:2025:SOD

[HJ25] Soonhoi Ha and Eunjin Jeong. Software optimization and design methodology for low power computer vision systems. ACM Transactions on Embedded Computing Systems, 24(1):19:1– 19:??, January 2025. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3687310.

Hammari:2018:RPD

[HKC18] Elena Hammari, Per Gunnar Kjeldsberg, and Francky Catthoor. Runtime precomputation of data-dependent parameters in embedded systems. ACM Transactions on Embedded Computing Systems, 17(3):68:1–68:??, June 2018. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Huai:2023:CCR

[HKL^{+}23] Shuo Huai, Hao Kong, Xiangzhong Luo, Shiqing Li, Ravi Subramaniam, Christian Makaya, Qian Lin, and Weichen Liu. CRIMP: Compact & reliable DNN inference on in-memory processing via crossbar-aligned compression and non-ideality adaptation. ACM Transactions on Embedded Computing Systems, 22(5s):123:1– 123:??, October 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3609115.

Huang:2005:ESC

[HKLH05] Tai-Yi Huang, Chung-Ta King, Youn-Long Steve Lin, and Yin-Tsung Hwang. The embedded software consortium of Taiwan. ACM Transactions on Embedded Computing Systems, 4(3):612– 632, August 2005. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Hessell:2008:EES

[HKP08] Fabiano Hessell, Kenneth Kent, and Dionisios Pnevmatikatos. Editorial: Embedded systems — new challenges and future directions. ACM Transactions on Embedded Computing Systems, 7(4):37:1–37:??, July 2008. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Hassan:2018:EID

[HKP18] Mohamed Hassan, Anirudh M.

Kaushik, and Hiren Patel. Exposing implementation details of embedded DRAM memory controllers through latency-based analysis. ACM Transactions on Embedded Computing Systems, 17(5): 90:1–90:??, November 2018. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/ft_gateway.
cfm?id=3274281.

Hu:2005:ADR

[HKVI05] J. Hu, M. Kandemir, N. Vijaykrishnan, and M. J. Irwin. Analyzing data reuse for cache reconfiguration. ACM Transactions on Embedded Computing Systems, 4(4): 851–876, November 2005. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Huang:2014:YES

[HL14] Yu-Jen Huang and Jin-Fu Li.

Yield-enhancement schemes for multicore processor and memory stacked 3D ICs. ACM Transactions on Embedded Computing Systems, 13(3s):106:1–106:??, March 2014. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Hu:2009:CAS

[HLD^{+}09] Jie Hu, Feihui Li, Vijay Degalahal, Mahmut Kandemir, N. Vijaykrishnan, and Mary J. Irwin. Compilerassisted soft error detection under performance and energy constraints in embedded systems. ACM Transactions on Embedded Computing Systems, 8(4):27:1– 27:??, July 2009. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Hong:2018:ISP

[HLF^{+}18] Ding-Yong Hong, Yu-Ping Liu, Sheng-Yu Fu, Jan-Jan Wu, and Wei-Chung Hsu. Improving SIMD parallelism via dynamic binary translation. ACM Transactions on Embedded Computing Systems, 17(3):61:1–61:??, June 2018. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Hung:2023:EEC

[HLL^{+}23] Chen-Tui Hung, Kai Xuan Lee, Yi-Zheng Liu, Ya-Shu Chen, and Zhong-Han Chan.

Energy-efficient communications for improving timely progress of intermittentpowered BLE devices. ACM Transactions on Embedded Computing Systems, 22 (6):102:1–102:??, November 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3626197.

Hsieh:2012:PBP

[HLLL12] Kun-Yuan Hsieh, Chi-Hua Lai, Shang-Hong Lai, and Jenq Kuen Lee. Parallelization of belief propagation on Cell processors for stereo vision. ACM Transactions on Embedded Computing Systems, 11S(1):13:1–13:??, 2012. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Hu:2020:GBT

[HLLL20] Junyan Hu, Kenli Li, Chubo Liu, and Keqin Li. Gamebased task offloading of multiple mobile devices with QoS in mobile edge computing systems of limited computation capacity. ACM Transactions on Embedded Computing Systems, 19(4): 29:1–29:21, July 2020. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3398038.

Hoag:2024:RFM

[HLM^{+}24] Ellis Hoag, Kyungwoo Lee, Julian Mestre, Sergey Pupyrev, and Yongkang Zhu. Reordering functions in mobiles apps for reduced size and faster start-up. ACM Transactions on Embedded Computing Systems, 23(4): 54:1–54:??, July 2024. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3660635.

Haar:2017:MGE

[HM17] Stefan Haar and Roland Meyer. Message from the Guest Editors. ACM Transactions on Embedded Computing Systems, 16(2):40:1– 40:??, April 2017. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

He:2021:GCF

[HMLZ21] Jiaji He, Haocheng Ma, Yanjiang Liu, and Yiqiang Zhao. Golden chip-free Trojan detection leveraging Trojan Trigger’s side-channel fingerprinting. ACM Transactions on Embedded Computing Systems, 20(1):6:1– 6:18, January 2021. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3419105.

Harkin:2004:MOR

[HMM04] J. Harkin, T. M. McGinnity, and L. P. Maguire. Modeling and optimizing run-time reconfiguration using evolutionary computation. ACM Transactions on Embedded Computing Systems, 3(4): 661–685, November 2004. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Huang:2004:DDR

[HMMA04] Zhining Huang, Sharad Malik, Nahri Moreano, and Guido Araujo. The design of dynamically reconfigurable datapath coprocessors. ACM Transactions on Embedded Computing Systems, 3(2):361–384, May 2004. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Halder:2023:OPN

[HMR23] Dipal Halder, Maneesh Merugu, and Sandip Ray. ObNoCs: Protecting networkon-chip fabrics against reverseengineering attacks. ACM Transactions on Embedded Computing Systems, 22(5s): 112:1–112:??, October 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3609107.

Hosseinabady:2018:DEM

[HNY18] Mohammad Hosseinabady and Jose Luis Nunez-Yanez. Dynamic energy management of FPGA accelerators in embedded systems. ACM Transactions on Embedded Computing Systems, 17(3): 63:1–63:??, June 2018. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Hanson:2012:AFE

[HPBL12] Mark A. Hanson, Harry C. Powell, Jr., Adam T. Barth, and John Lach. Applicationfocused energy-fidelity scalability for wireless motionbased health assessment. ACM Transactions on Embedded Computing Systems, 11(S2):50:1–50:??, 2012. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Huang:2009:SFB

[HPLD09] Jian Huang, Matthew Parris, Jooheung Lee, and Ronald F. Demara. Scalable FPGA-based architecture for DCT computation using dynamic partial reconfiguration. ACM Transactions on Embedded Computing Systems, 9(1):9:1–9:??, October 2009. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Howe:2015:PLB

[HPO^{+}15] James Howe, Thomas Pöppelmann, Máire O’Neill, Elizabeth O’Sullivan, and Tim Güneysu. Practical lattice-based digital signature schemes. ACM Transactions on Embedded Computing Systems, 14(3): 41:1–41:??, April 2015. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Hassan:2017:PRA

[HPP17] Mohamed Hassan, Hiren Patel, and Rodolfo Pellizzoni. PMC: a requirementaware DRAM controller for multicore mixed criticality systems. ACM Transactions on Embedded Computing Systems, 16(4):100:1– 100:??, August 2017. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Huang:2013:TBD

[HPS13] Yi-Sheng Huang, Yen-Liang Pan, and Pin-June Su. Transition-based deadlock detection and recovery policy for FMSs using graph technique. ACM Transactions on Embedded Computing Systems, 12(1):11:1– 11:??, January 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Hua:2006:EEE

[HQB06] Shaoxiong Hua, Gang Qu, and Shuvra S. Bhattacharyya. Energy-efficient embedded software implementation on multiprocessor system-on-chip with multiple voltages. ACM Transactions on Embedded Computing Systems, 5(2):321–341, May 2006. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Hua:2007:PDM

[HQB07] Shaoxiong Hua, Gang Qu, and Shuvra S. Bhattacharyya. Probabilistic design of multimedia embedded systems. ACM Transactions on Embedded Computing Systems, 6(3):15:1–15:??, July 2007. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Hammadeh:2020:WHR

[HQE20] Zain A. H. Hammadeh, Sophie Quinton, and Rolf Ernst. Weakly-hard realtime guarantees for earliest deadline first scheduling of independent tasks. ACM Transactions on Embedded Computing Systems, 18(6):1–25, January 2020. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3356865.

Herzog:2022:RDE

[HRH^{+}22] Benedict Herzog, Stefan Reif, Judith Hemp, Timo Hönig, and Wolfgang Schröder-Preikschat. Resourcedemand estimation for edge tensor processing units. ACM Transactions on Embedded Computing Systems, 21(5):58:1–58:??, September 2022. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3520132.

HeydariGorji:2022:LCS

[HRT^{+}22] Ali HeydariGorji, Siavash Rezaei, Mahdi Torabzadehkashi, Hossein Bobarshad, Vladimir Alves, and Pai H. Chou. Leveraging computational storage for power-efficient distributed data analytics. ACM Transactions on Embedded Computing Systems, 21(6):82:1–82:??, November 2022. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3528577.

Hada:2024:DCH

[HS24] Rupendra Pratap Singh Hada and Abhishek Srivastava. Dynamic cluster head selection in WSN. ACM Transactions on Embedded Computing Systems, 23(4): 64:1–64:??, July 2024. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3665867.

Hemmat:2022:CCA

[HSD22] Maedeh Hemmat, Joshua San Miguel, and Azadeh Davoodi. CAP’NN: a classaware framework for personalized neural network inference. ACM Transactions on Embedded Computing Systems, 21(5):59:1– 59:??, September 2022. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3520126.

Hegde:2018:CAC

[HSK18] Gopalakrishna Hegde, Siddhartha, and Nachiket Kapre. CaffePresso: Accelerating convolutional networks on embedded SoCs. ACM Transactions on Embedded Computing Systems, 17(1): 15:1–15:??, January 2018. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Hassan:2016:HSB

[HSMS16] Hadeer A. Hassan, Sameh A. Salem, Ahmed M. Mostafa, and E. M. Saad. Harmonic segment-based semipartitioning scheduling on multi-core real-time systems. ACM Transactions on Embedded Computing Systems, 15(4):73:1–73:??, August 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Harrison:2018:CPR

[HSR18] David C. Harrison, Winston K. G. Seah, and Ramesh Rayudu. Coverage preservation with rapid forwarding in energy-harvesting wireless sensor networks for critical rare events. ACM Transactions on Embedded Computing Systems, 17(2):43:1– 43:??, April 2018. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Higuera-Toledano:2006:HSD

[HT06] M. Teresa Higuera-Toledano.

Hardware support for detecting illegal references in a multiapplication real-time Java environment. ACM Transactions on Embedded Computing Systems, 5(4): 753–772, November 2006. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

He:2025:HHT

[HTB^{+}25] Pengzhou He, Yazheng Tu, Tianyou Bao, Çetin Çetin Koç, and Jiafeng Xie. HSPA: High-throughput sparse polynomial multiplication for code-based post-quantum cryptography. ACM Transactions on Embedded Computing Systems, 24(1):16:1– 16:??, January 2025. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3703837.

Hsiu:2016:UCS

[HTC^{+}16] Pi-Cheng Hsiu, Po-Hsien Tseng, Wei-Ming Chen, Chin-Chiang Pan, and Tei-Wei Kuo. User-centric scheduling and governing on mobile devices with big.LITTLE processors. ACM Transactions on Embedded Computing Systems, 15(1): 17:1–17:??, February 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Higuera-Toledano:2010:ISI

[HTLC10] M. Teresa Higuera-Toledano, Doug Locke, and Angelo Corsaro. Introduction to special issue on Java technologies for real-time and embedded systems. ACM Transactions on Embedded Computing Systems, 10(1):1:1– 1:??, August 2010. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Hester:2016:PCB

[HTR^{+}16] Josiah Hester, Nicole Tobias, Amir Rahmati, Lanny Sitanayah, Daniel Holcomb, Kevin Fu, Wayne P. Burleson, and Jacob Sorber. Persistent clocks for batteryless sensing devices. ACM Transactions on Embedded Computing Systems, 15(4): 77:1–77:??, August 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Hubner:2013:ISS

[Hüb13] Michael Hübner. Introduction to the special section on multiprocessor systemon-chip for cyber-physical systems. ACM Transactions on Embedded Computing Systems, 12(1s):46:1– 46:??, March 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Huang:2013:ASP

[HVG13] Chen Huang, Frank Vahid, and Tony Givargis. Automatic synthesis of physical system differential equation models to a custom network of general processing elements on FPGAs. ACM Transactions on Embedded Computing Systems, 13(2):23:1–23:??, September 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Huang:2017:FBA

[HW17] Guoxian Huang and Lei Wang. An FPGA-based architecture for high-speed compressed signal reconstruction. ACM Transactions on Embedded Computing Systems, 16(3):89:1– 89:??, July 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Hsiao:2020:CHC

[HWC^{+}20] Luke Hsiao, Sen Wu, Nicholas Chiang, Christopher Ré, and Philip Levis. Creating hardware component knowledge bases with training data generation and multi-task learning. ACM Transactions on Embedded Computing Systems, 19(6):42:1– 42:26, November 2020. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3391906.

Huang:2022:HFY

[HWC22] Shihua Huang, Luc Waeijen, and Henk Corporaal. How flexible is your computing system? ACM Transactions on Embedded Computing Systems, 21(4): 37:1–37:??, July 2022. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3524861.

Hsu:2023:GBC

[HWTL23] Yao-Jen Hsu, Chin-Hsien Wu, Yu-Chieh Tsai, and Chia-Cheng Liu. A granularitybased clustering method for reducing write amplification in solid-state drives. ACM Transactions on Embedded Computing Systems, 22(4):

70:1–70:??, July 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3605779.

He:2024:REA

[HXH^{+}24] Zhijian He, Bohuan Xue, Xiangcheng Hu, Zhaoyan Shen, Xiangyue Zeng, and Ming Liu. Robust embedded autonomous driving positioning system fusing Li-DAR and inertial sensors. ACM Transactions on Embedded Computing Systems, 23(1):15:1–15:??, January 2024. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3626098.

Hu:2013:WAR

[HXZ^{+}13] Jingtong Hu, Chun Jason Xue, Qingfeng Zhuge, Wei-Che Tseng, and Edwin H.-M. Sha. Write activity reduction on non-volatile main memories for embedded chip multiprocessors. ACM Transactions on Embedded Computing Systems, 12(3):77:1– 77:??, March 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Hong:2022:EGE

[HY22] Ziyang Hong and C. Patrick Yue. Efficient-grad: Efficient training deep convolutional neural networks on edge devices with gradient optimizations. ACM Transactions on Embedded Computing Systems, 21(2):19:1– 19:24, March 2022. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3504034.

Huang:2015:COM

[HYY^{+}15] Kai Huang, Min Yu, Rongjie Yan, Xiaomeng Zhang, Xiaolang Yan, Lisane Brisolara, Ahmed Amine Jerraya, and Jiong Feng. Communication optimizations for multithreaded code generation from Simulink models. ACM Transactions on Embedded Computing Systems, 14(3):59:1–59:??, May 2015. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Hu:2018:JOS

[HZGW18] Hang Hu, Hang Zhang, Jianxin Guo, and Feng Wang. Joint optimization of sensing and power allocation in energy-harvesting cognitive radio networks. ACM Transactions on Embedded Computing Systems, 17(1): 8:1–8:??, January 2018. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Huang:2018:HPH

[HZH^{+}18] Tian Huang, Yongxin Zhu, Yajun Ha, Xu Wang, and Meikang Qiu. A hardware pipeline with high energy and resource efficiency for FMM acceleration. ACM Transactions on Embedded Computing Systems, 17(2): 51:1–51:??, April 2018. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Huang:2023:DEL

[HZW^{+}23] Wenbo Huang, Lei Zhang, Shuoyuan Wang, Hao Wu, and Aiguo Song. Deep ensemble learning for human activity recognition using wearable sensors via filter activation. ACM Transactions on Embedded Computing Systems, 22(1):15:1– 15:??, January 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3551486.

Hu:2014:MON

[HZX^{+}14] Jingtong Hu, Qingfeng Zhuge, Chun Jason Xue, Wei-Che Tseng, and Edwin H.-M. Sha. Management and optimization for nonvolatile memory-based hybrid scratchpad memory on multicore embedded processors. ACM Transactions on Embedded Computing Systems, 13(4):79:1– 79:??, February 2014. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Huang:2015:JWU

[HZX15] Yazhi Huang, Mengying Zhao, and Chun Jason Xue. Joint WCET and update activity minimization for cyber-physical systems. ACM Transactions on Embedded Computing Systems, 14(1):6:1–6:??, January 2015. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Hassantabar:2022:MMH

[HZYJ22] Shayan Hassantabar, Joe Zhang, Hongxu Yin, and Niraj K. Jha. MHDeep: Mental health disorder detection system based on wearable sensors and artificial neural networks. ACM Transactions on Embedded Computing Systems, 21(6):81:1– 81:??, November 2022. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3527170.

Ikeda:2023:MDD

[IAS23] Sosei Ikeda, Hiromitsu Awano, and Takashi Sato. Modular DFR: Digital delayed feedback reservoir model for enhancing design flexibility. ACM Transactions on Embedded Computing Systems, 22(5s):110:1–110:??, October 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3609105.

Indrusiak:2023:RTG

[IB23] Leandro Soares Indrusiak and Alan Burns. Realtime guarantees in routerless networks-on-chip. ACM Transactions on Embedded Computing Systems, 22(5): 88:1–88:??, September 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3616539.

Irturk:2010:GAG

[IBMK10] Ali Irturk, Bridget Benson, Shahnam Mirzaei, and Ryan Kastner. GUSTO: an automatic generation and optimization tool for matrix inversion architectures. ACM Transactions on Embedded Computing Systems, 9(4):32:1–32:??, March 2010. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Ivanov:2021:VSA

[ICW^{+}21] Radoslav Ivanov, Taylor J.

Carpenter, James Weimer, Rajeev Alur, George J. Pappas, and Insup Lee. Verifying the safety of autonomous systems with neural network controllers. ACM Transactions on Embedded Computing Systems, 20(1):7:1– 7:26, January 2021. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3419742.

Isik:2023:NNC

[ICZ^{+}23] Berivan Isik, Kristy Choi, Xin Zheng, Tsachy Weissman, Stefano Ermon, H.- S. Philip Wong, and Armin Alaghi. Neural network compression for noisy storage devices. ACM Transactions on Embedded Computing Systems, 22(3):58:1–58:??, May 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3588436.

Isuwa:2022:QMQ

[IDO^{+}22] Samuel Isuwa, Somdip Dey, Andre P. Ortega, Amit Kumar Singh, Bashir M. Al-Hashimi, and Geoff V. Merrett. QUAREM: Maximising QoE through adaptive resource management in mobile MPSoC platforms. ACM Transactions on Embedded Computing Systems, 21(4): 38:1–38:??, July 2022. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3526116.

Iida:2016:GET

[IFA^{+}16] Yuki Iida, Yusuke Fujii, Takuya Azumi, Nobuhiko Nishio, and Shinpei Kato.

GPUrpc: Exploring transparent access to remote GPUs. ACM Transactions on Embedded Computing Systems, 16(1):17:1– 17:??, November 2016. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Im:2004:DVS

[IHK04] Chaeseok Im, Soonhoi Ha, and Huiseok Kim. Dynamic voltage scheduling with buffers in low-power multimedia applications. ACM Transactions on Embedded Computing Systems, 3(4): 686–705, November 2004. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Izosimov:2012:SOF

[IPEP12] Viacheslav Izosimov, Paul Pop, Petru Eles, and Zebo Peng. Scheduling and optimization of fault-tolerant embedded systems with transparency/performance trade-offs. ACM Transactions on Embedded Computing Systems, 11(3):61:1– 61:??, September 2012. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Ivanov:2016:ARS

[IPL16] Radoslav Ivanov, Miroslav Pajic, and Insup Lee. Attack-resilient sensor fusion for safety-critical cyberphysical systems. ACM Transactions on Embedded Computing Systems, 15(1): 21:1–21:??, February 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Inoue:2010:RSC

[ISE10] Hiroaki Inoue, Junji Sakai, and Masato Edahiro. A robust seamless communication architecture for nextgeneration mobile terminals on multi-CPU SoCs. ACM Transactions on Embedded Computing Systems, 9(3): 19:1–19:??, February 2010. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Isenberg:2017:IIV

[Ise17] Tobias Isenberg. Incremental inductive verification of parameterized timed systems. ACM Transactions on Embedded Computing Systems, 16(2):47:1–47:??, April 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Irani:2003:OSD

[ISG03] Sandy Irani, Sandeep Shukla, and Rajesh Gupta. Online strategies for dynamic power management in systems with multiple power-saving states. ACM Transactions on Embedded Computing Systems, 2(3):325–346, August 2003. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Ibrahim:2021:MFU

[ISOD21] Omar Adel Ibrahim, Savio Sciancalepore, Gabriele Oligeri, and Roberto Di Pietro. MAGNETO: Fingerprinting USB flash drives via unintentional magnetic emissions. ACM Transactions on Embedded Computing Systems, 20(1):8:1–8:26, January 2021. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL

https://dl.acm.org/doi/
10.1145/3422308.

Inoue:2008:FAC

[ISTE08] Hiroaki Inoue, Junji Sakai, Sunao Torii, and Masato Edahiro. FIDES: an advanced chip multiprocessor platform for secure next generation mobile terminals. ACM Transactions on Embedded Computing Systems, 8(1):1:1–1:??, December 2008. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Ienne:2016:GES

[IT16] Paolo Ienne and Jean-Pierre Talpin. Guest editorial: Special issue on models and methodologies for system design. ACM Transactions on Embedded Computing Systems, 15(2):29:1–29:??, May 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Iyer:2024:HCM

[ITO^{+}24] Vishnuvardhan V. Iyer, Aditya Thimmaiah, Michael Orshansky, Andreas Gerstlauer, and Ali E. Yilmaz. A hierarchical classification method for highaccuracy instruction disassembly with near-field EM measurements. ACM Transactions on Embedded Computing Systems, 23(1):10:1– 10:??, January 2024. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3629167.

Inci:2023:QFQ

[IVJ^{+}23] Ahmet Inci, Siri Virupaksha, Aman Jain, Ting-Wu Chin, Venkata Thallam, Ruizhou Ding, and Diana Marculescu. QUIDAM: a framework for quantization-aware DNN accelerator and model coexploration. ACM Transactions on Embedded Computing Systems, 22(2):33:1– 33:??, March 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3555807.

Inagaki:2023:PSC

[IYL^{+}23] Saya Inagaki, Mingyu Yang, Yang Li, Kazuo Sakiyama, and Yuko Hara-Azumi.

Power side-channel attack resistant circuit designs of ARX ciphers using high-level synthesis. ACM Transactions on Embedded Computing Systems, 22(5):85:1– 85:??, September 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3609507.

Jokic:2022:CKE

[JAB^{+}22] Petar Jokic, Erfan Azarkhish, Andrea Bonetti, Marc Pons, Stephane Emery, and Luca Benini. A construction kit for efficient low power neural network accelerator designs. ACM Transactions on Embedded Computing Systems, 21(5):56:1–56:??, September 2022. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3520127.

Jiang:2019:BSR

[JAD19] Zhe Jiang, Neil Audsley, and Pan Dong. BlueIO: a scalable real-time hardware I/O virtualization system for many-core embedded systems. ACM Transactions on Embedded Computing Systems, 18(3):19:1–19:??, June 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/ft_gateway.
cfm?id=3309765.

Jacob:2002:ITS

[JB02] Bruce Jacob and Shuvra Bhattacharyya. Introduction to the two special issues on memory. ACM Transactions on Embedded Computing Systems, 1(1):2–5, November 2002. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Jacob:2003:ITS

[JB03] Bruce Jacob and Shuvra Bhattacharyya. Introduction to the two special issues on memory. ACM Transactions on Embedded Computing Systems, 2(1):1–4, February 2003. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Johnson:2016:RTR

[JBCS16] Taylor T. Johnson, Stanley Bak, Marco Caccamo, and Lui Sha. Real-time reachability for verified simplex design. ACM Transactions on Embedded Computing Systems, 15(2):26:1–26:??, May 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Jayakodi:2020:DOE

[JBDD20] Nitthilan Kanappan Jayakodi, Syrine Belakaria, Aryan Deshwal, and Janardhan Rao Doppa. Design and optimization of energy-accuracy tradeoff networks for mobile platforms via pretrained deep models. ACM Transactions on Embedded Computing Systems, 19(1):4:1– 4:24, February 2020. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3366636.

Josipovic:2017:OLS

[JBI17] Lana Josipovic, Philip Brisk, and Paolo Ienne. An outof-order load-store queue for spatial computing. ACM Transactions on Embedded Computing Systems, 16(5s): 125:1–125:??, October 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Jia:2013:SLI

[JBN^{+}13] Zai Jian Jia, Tomás Bautista, Antonio Núñez, Andy D. Pimentel, and Mark Thompson. A system-level infrastructure for multidimensional MP-SoC design space co-exploration. ACM Transactions on Embedded Computing Systems, 13(1s):27:1– 27:??, November 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Jacome:2003:SIP

[JC03] Margarida Jacome and Francky Catthoor. Special issue on power-aware embedded computing. ACM Transactions on Embedded Computing Systems, 2(3): 251–254, August 2003. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Jeong:2012:PLT

[JC12] Jaein Jeong and David Culler. Predicting the longterm behavior of a microsolar power system. ACM Transactions on Embedded Computing Systems, 11(2): 35:1–35:??, July 2012. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Ji:2017:LDC

[JCS^{+}17] Cheng Ji, Li-Pin Chang, Liang Shi, Congming Gao, Chao Wu, Yuangang Wang, and Chun Jason Xue. Lightweight data compression for mobile flash storage. ACM Transactions on Embedded Computing Systems, 16(5s):183:1–183:??, October 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Ji:2016:CLO

[JCW^{+}16] Wen Ji, Bo-Wei Chen, Xiangdong Wang, Haiyong Luo, Mucheol Kim, and Yiqiang Chen. Cross-layer opportunistic scheduling for device-to-device video multicast services. ACM Transactions on Embedded Computing Systems, 15(2):37:1– 37:??, May 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Jiang:2024:EMB

[JDW24] Zijing Jiang, Qun Ding, and An Wang. Efficient multibyte power analysis architecture focusing on bitwise linear leakage. ACM Transactions on Embedded Computing Systems, 23(6):102:1– 102:??, November 2024. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3687484.

Jiang:2016:PAD

[JEP16] Ke Jiang, Petru Eles, and Zebo Peng. Power-aware design techniques of secure multimode embedded systems. ACM Transactions on Embedded Computing Systems, 15(1):6:1– 6:??, February 2016. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Jezequel:2015:FPA

[JFK15] Loïg Jezequel, Eric Fabre, and Victor Khomenko. Factored planning: From automata to Petri nets. ACM Transactions on Embedded Computing Systems, 14(2): 26:1–26:??, March 2015. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Javadi:2023:CME

[JFM23] Mohammad Haji Seyed Javadi, Mohsen Faryabi, and Hamid Reza Mahdiani. A comprehensive model for efficient design space exploration of imprecise computational blocks. ACM Transactions on Embedded Computing Systems, 22(6):103:1– 103:??, November 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3625555.

Jati:2024:CCK

[JGCS24] Arpan Jati, Naina Gupta, Anupam Chattopadhyay, and Somitra Kumar Sanadhya. A configurable CRYSTALS– Kyber hardware implementation with side-channel protection. ACM Transactions on Embedded Computing Systems, 23(2):33:1– 33:??, March 2024. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3587037.

Jafari:2009:EPR

[JGD^{+}09] Roozbeh Jafari, Hassan Ghasemzadeh, Foad Dabiri, Ani Nahapetian, and Majid Sarrafzadeh. An efficient placement and routing technique for fault-tolerant distributed embedded computing. ACM Transactions on Embedded Computing Systems, 8(4):28:1–28:??, July 2009. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Jackson:2021:EES

[JGL21] Riley Jackson, Jonathan Gresl, and Ramon Lawrence. Efficient external sorting for memory-constrained embedded devices with flash memory. ACM Transactions on Embedded Computing Systems, 20(4):29:1–29:21, June 2021. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3446976.

Jin:2018:PAR

[JGX^{+}18] Xi Jin, Nan Guan, Changqing Xia, Jintao Wang, and Peng Zeng. Packet aggregation real-time scheduling for large-scale WIA– PA industrial wireless sensor networks. ACM Transactions on Embedded Computing Systems, 17(5):88:1– 88:??, November 2018. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3266228.

Jones:2006:RPW

[JHK^{+}06] Alex K. Jones, Raymond Hoare, Dara Kusic, Gayatri Mehta, Josh Fazekas, and John Foster. Reducing power while increasing performance with SuperCISC. ACM Transactions on Embedded Computing Systems, 5(3):658–686, August 2006. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Jungeblut:2013:SAO

[JHPR13] Thorsten Jungeblut, Boris Hübener, Mario Porrmann, and Ulrich Rückert. A systematic approach for optimized bypass configurations for application-specific embedded processors. ACM Transactions on Embedded Computing Systems, 13(2): 18:1–18:??, September 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Jiang:2025:EAI

[JJS^{+}25] Junqiang Jiang, Shengjie Jin, Zhifang Sun, Jinxue Duan, Lizhi Liu, Li Pan, and Zebo Peng. An efficient approach for improving message acceptance rate and link utilization in timesensitive networking. ACM Transactions on Embedded Computing Systems, 24(1): 1:1–1:??, January 2025. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3690638.

Jeong:2013:RRM

[JKH^{+}13] Jinkyu Jeong, Hwanju Kim, Jeaho Hwang, Joonwon Lee, and Seungryoul Maeng. Rigorous rental memory management for embedded systems. ACM Transactions on Embedded Computing Systems, 12(1s):43:1–43:??, March 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Jeong:2022:TBF

[JKH22] Eunjin Jeong, Jangryul Kim, and Soonhoi Ha. TensorRTbased framework and optimization methodology for deep learning inference on Jetson boards. ACM Transactions on Embedded Computing Systems, 21(5):51:1– 51:??, September 2022. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3508391.

Jung:2010:SFS

[JKJ^{+}10] Dawoon Jung, Jeong-Uk Kang, Heeseung Jo, Jin-Soo Kim, and Joonwon Lee. Superblock FTL: a superblockbased Flash Translation Layer with a hybrid address translation scheme. ACM Transactions on Embedded Computing Systems, 9(4): 40:1–40:??, March 2010. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Jafari:2013:ISS

[JLSK13] Roozbeh Jafari, John Lach, Majid Sarrafzadeh, and William Kaiser. Introduction to the special section on wireless health systems. ACM Transactions on Embedded Computing Systems, 12(4):98:1–98:??, June 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Ji:2018:ACP

[JLSP18] Kecheng Ji, Ming Ling, Longxing Shi, and Jianping Pan. An analytical cache performance evaluation framework for embedded out-of-order processors using software characteristics. ACM Transactions on Embedded Computing Systems, 17(4):79:1– 79:??, August 2018. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Jia:2015:TAD

[JLW^{+}15] Zhiping Jia, Yang Li, Yi Wang, Meng Wang, and Zili Shao. Temperatureaware data allocation for embedded systems with cache and scratchpad memory. ACM Transactions on Embedded Computing Systems, 14(2):30:1–30:??, March 2015. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Jerraya:2006:GEC

[JM06] Ahmed Jerraya and Trevor Mudge. Guest editorial: Concurrent hardware and software design for multiprocessor SoC. ACM Transactions on Embedded Computing Systems, 5(2):259– 262, May 2006. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Jung:2014:HCO

[JMO14] Dong-Heon Jung, Soo-Mook Moon, and Hyeong-Seok Oh. Hybrid compilation and optimization for Java-based digital TV platforms. ACM Transactions on Embedded Computing Systems, 13(2s): 62:1–62:??, January 2014. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Jimenez:2015:LSC

[JNI15] Xavier Jimenez, David Novo, and Paolo Ienne. Libra: Software-controlled cell bitdensity to balance wear in NAND flash. ACM Transactions on Embedded Computing Systems, 14(2):28:1– 28:??, March 2015. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Jin:2014:PPA

[JP14] Yuho Jin and Timothy Mark Pinkston. PAIS: Parallelismaware interconnect scheduling in multicores. ACM Transactions on Embedded Computing Systems, 13(3s): 108:1–108:??, March 2014. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Jain:2020:CHS

[JR20] Shubham Jain and Anand Raghunathan. CxDNN: Hardware-software compensation methods for deep neural networks on resistive crossbar systems. ACM Transactions on Embedded Computing Systems, 18(6): 1–23, January 2020. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3362035.

Jayakumar:2016:SMV

[JRR16] Hrishikesh Jayakumar, Arnab Raha, and Vijay Raghunathan. Sleep-mode voltage scaling: Enabling SRAM data retention at ultra-low power in embedded microcontrollers. ACM Transactions on Embedded Computing Systems, 16(1):10:1– 10:??, November 2016. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Jayakumar:2017:EAM

[JRSR17] Hrishikesh Jayakumar, Arnab Raha, Jacob R. Stevens, and Vijay Raghunathan. Energyaware memory mapping for hybrid FRAM–SRAM MCUs in intermittentlypowered IoT devices. ACM Transactions on Embedded Computing Systems, 16(3): 65:1–65:??, July 2017. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Juang:2023:LCG

[JSD23] Tzung-Han Juang, Christof Schlaak, and Christophe Dubach. Let coarse-grained resources be shared: Mapping entire neural networks on FPGAs. ACM Transactions on Embedded Computing Systems, 22(5s):114:1– 114:??, October 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3609109.

Jiang:2019:ASL

[JSZ^{+}19] Weiwen Jiang, Edwin H.-M.

Sha, Xinyi Zhang, Lei Yang, Qingfeng Zhuge, Yiyu Shi, and Jingtong Hu. Achieving super-linear speedup across multi-FPGA for real-time DNN inference. ACM Transactions on Embedded Computing Systems, 18(5s):67:1– 67:??, October 2019. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3358192.

Jiang:2015:AEB

[JZL^{+}15] Jian-Min Jiang, Huibiao Zhu, Qin Li, Yongxin Zhao, Lin Zhao, Shi Zhang, Ping Gong, and Zhong Hong. Analyzing event-based scheduling in concurrent reactive systems. ACM Transactions on Embedded Computing Systems, 14(4):86:1– 86:??, December 2015. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Kim:2005:DDC

[KAK05] Hyung Seok Kim, Tarek F. Abdelzaher, and Wook Hyun Kwon. Dynamic delayconstrained minimum-energy dissemination in wireless sensor networks. ACM Transactions on Embedded Computing Systems, 4(3): 679–706, August 2005. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Kamal:2015:OHC

[KAKSP15] Mehdi Kamal, Ali Afzali-Kusha, Saeed Safari, and Massoud Pedram. OPLE: a heuristic custom instruction selection algorithm based on partitioning and local exploration of application dataflow graphs. ACM Transactions on Embedded Computing Systems, 14(4): 72:1–72:??, December 2015. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Kadiyala:2020:LLA

[KAS^{+}20] Sai Praveen Kadiyala, Manaar Alam, Yash Shrivastava, Sikhar Patranabis, Muhamed Fauzi Bin Abbas, Arnab Kumar Biswas, Debdeep Mukhopadhyay, and Thambipillai Srikanthan. LAMBDA: Lightweight Assessment of Malware for em-BeddeD architectures. ACM Transactions on Embedded Computing Systems, 19(4): 23:1–23:31, July 2020. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3390855.

Kumar:2007:ESI

[KASD07] Nagendra J. Kumar, Vasanth Asokan, Siddhartha Shivshankar, and Alexander G. Dean. Efficient software implementation of embedded communication protocol controllers using asynchronous software thread integration with timeand space-efficient procedure calls. ACM Transactions on Embedded Computing Systems, 6(1):2:1– 2:??, February 2007. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Kurtin:2017:ART

[KB17] Philip S. Kurtin and Marco J. G. Bekooij. An abstractionrefinement theory for the analysis and design of realtime systems. ACM Transactions on Embedded Computing Systems, 16(5s):173:1– 173:??, October 2017. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Koh:2023:PST

[KB23] Jaime Koh and Bruno Bodin. K-periodic scheduling for throughput-buffering trade-off exploration of CSDF. ACM Transactions on Embedded Computing Systems, 22(1):19:1– 19:??, January 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3559760.

Kim:2013:SIE

[KBCL13] Jongsung Kim, Javier A. Barria, Morris Chang, and Victor C. M. Leung. Special issue on embedded systems for interactive multimedia services (ES-IMS). ACM Transactions on Embedded Computing Systems, 12(2): 19:1–19:??, February 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Kutukcu:2024:SAI

[KBD24] Basar Kutukcu, Sabur Baidya, and Sujit Dey. SLEXNet: Adaptive inference using slimmable early exit neural networks. ACM Transactions on Embedded Computing Systems, 23(6):104:1– 104:??, November 2024. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3689632.

Kim:2008:EAC

[KBDV08] Minyoung Kim, Sudarshan Banerjee, Nikil Dutt, and Nalini Venkatasubramanian. Energy-aware cosynthesis of real-time multimedia applications on MPSoCs using heterogeneous scheduling policies. ACM Transactions on Embedded Computing Systems, 7(2):9:1– 9:??, February 2008. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Kutukcu:2022:CGA

[KBRD22] Basar Kutukcu, Sabur Baidya, Anand Raghunathan, and Sujit Dey. Contention grading and adaptive model selection for machine vision in embedded systems. ACM Transactions on Embedded Computing Systems, 21(5):55:1–55:??, September 2022. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3520134.

Kim:2017:WAF

[KBS17] Yooseong Kim, David Broman, and Aviral Shrivastava. WCET-aware functionlevel dynamic code management on scratchpad memory. ACM Transactions on Embedded Computing Systems, 16(4):112:1–112:??, August 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Ko:2021:LCL

[KCBM21] Yousun Ko, Alex Chadwick, Daniel Bates, and Robert Mullins. Lane compression: a lightweight lossless compression method for machine learning on embedded systems. ACM Transactions on Embedded Computing Systems, 20(2):16:1– 16:26, March 2021. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3431815.

Kuan:2016:SEI

[KCC^{+}16] Yuan-Hung Kuan, Yuan-Hao Chang, Tseng-Yi Chen, Po-Chun Huang, and Kam-Yiu Lam. Space-efficient index scheme for PCM-based multiversion databases in cyberphysical systems. ACM Transactions on Embedded Computing Systems, 16(1):

21:1–21:??, November 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Kung:2017:CPD

[KCCW17] Sun-Yuan Kung, Thee Chanyaswad, J. Morris Chang, and Peiyuan Wu. Collaborative PCA/DCA learning methods for compressive privacy. ACM Transactions on Embedded Computing Systems, 16(3):76:1–76:??, July 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Koopman:2005:UES

[KCG^{+}05] Philip Koopman, Howie Choset, Rajeev Gandhi, Bruce Krogh, Diana Marculescu, Priya Narasimhan, Joann M. Paul, Ragunathan Rajkumar, Daniel Siewiorek, Asim Smailagic, Peter Steenkiste, Donald E. Thomas, and Chenxi Wang. Undergraduate embedded system education at Carnegie Mellon. ACM Transactions on Embedded Computing Systems, 4(3):500– 528, August 2005. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Kim:2016:APA

[KCJ^{+}16] Dongwon Kim, Yohan Chon, Wonwoo Jung, Yungeun Kim, and Hojung Cha. Accurate prediction of available battery time for mobile applications. ACM Transactions on Embedded Computing Systems, 15(3):48:1– 48:??, July 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Kang:2014:HSA

[KCWH14] Chih-Kai Kang, Yu-Jhang Cai, Chin-Hsien Wu, and Pi-Cheng Hsiu. A hybrid storage access framework for high-performance virtual machines. ACM Transactions on Embedded Computing Systems, 13(5s):157:1– 157:??, September 2014. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Kumar:2008:CCP

[KD08] Rajeev Kumar and Dipankar Das. Code compression for performance enhancement of variable-length embedded processors. ACM Transactions on Embedded Computing Systems, 7(3):35:1– 35:??, April 2008. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Kyriakis:2019:SMR

[KDB19] Panagiotis Kyriakis, Jyotirmoy V. Deshmukh, and Paul Bogdan. Specification mining and robust design under uncertainty: a stochastic temporal logic approach.

ACM Transactions on Embedded Computing Systems, 18(5s):96:1–96:??, October 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/ft_gateway.
cfm?id=3358231.

Koushanfar:2007:TMC

[KDN^{+}07] Farinaz Koushanfar, Abhijit Davare, David T. Nguyen, Alberto Sangiovanni-Vincentelli, and Miodrag Potkonjak. Techniques for maintaining connectivity in wireless adhoc networks under energy constraints. ACM Transactions on Embedded Computing Systems, 6(3):16:1– 16:??, July 2007. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Kundu:2023:SRB

[KDR23] Atanu Kundu, Sarthak Das, and Rajarshi Ray. SAT-Reach: a bounded model checker for affine hybrid systems. ACM Transactions on Embedded Computing Systems, 22(2):26:1– 26:??, March 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3567425.

Kerrison:2015:EMS

[KE15] Steve Kerrison and Kerstin Eder. Energy modeling of software for a hardware multithreaded embedded microprocessor. ACM Transactions on Embedded Computing Systems, 14(3): 56:1–56:??, May 2015. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Kundu:2022:TAA

[KFY^{+}22] Souvik Kundu, Yao Fu, Bill Ye, Peter A. Beerel, and Massoud Pedram. Toward adversary-aware noniterative model pruning through dynamic network rewiring of DNNs. ACM Transactions on Embedded Computing Systems, 21(5): 52:1–52:??, September 2022. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3510833.

Krishnaswamy:2005:DCB

[KG05] Arvind Krishnaswamy and Rajiv Gupta. Dynamic coalescing for 16-bit instructions. ACM Transactions on Embedded Computing Systems, 4(1):3–37, February 2005. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Kumar:2012:CMA

[KGR12] T. S. Rajesh Kumar, R. Govindarajan, and C. P. Ravikumar. On-chip memory architecture exploration framework for DSP processorbased embedded system on chip. ACM Transactions on Embedded Computing Systems, 11(1):5:1–5:??, March 2012. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Kloda:2023:LLS

[KGT^{+}23] Tomasz Kloda, Giovani Gracioli, Rohan Tabish, Reza Mirosanlou, Renato Mancuso, Rodolfo Pellizzoni, and Marco Caccamo. Lazy load scheduling for mixed-criticality applications in heterogeneous MP-SoCs. ACM Transactions on Embedded Computing Systems, 22(3):59:1–59:??, May 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3587694.

Kahkonen:2018:TPC

[KH18] Kari Kähkönen and Keijo Heljanko. Testing programs with contextual unfoldings. ACM Transactions on Embedded Computing Systems, 17(1):23:1–23:??, January 2018. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Kohl:2023:MBD

[KH23] Maximilian A. Köhl and Holger Hermanns. Model-based diagnosis of real-time systems: Robustness against varying latency, clock drift, and out-of-order observations. ACM Transactions on Embedded Computing Systems, 22(4):68:1–68:??, July 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3597209.

Khalgui:2013:DRA

[Kha13] Mohamed Khalgui. Distributed reconfigurations of autonomous IEC61499 systems. ACM Transactions on Embedded Computing Systems, 12(1):18:1–18:??, January 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Kohler:2023:RCE

[KHB^{+}23] Leonie Köhler, Phil Hertha, Matthias Beckert, Alex Bendrick, and Rolf Ernst. Robust cause-effect chains with bounded execution time and system-level logical execution time. ACM Transactions on Embedded Computing Systems, 22(3):50:1– 50:??, May 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3573388.

Kim:2014:MBM

[KHHH14] Heeseok Kim, Dong-Guk Han, Seokhie Hong, and Jaecheol Ha. Message blinding method requiring no multiplicative inversion for RSA. ACM Transactions on Embedded Computing Systems, 13(4):80:1–80:??, February 2014. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Kansal:2007:PME

[KHZS07] Aman Kansal, Jason Hsu, Sadaf Zahedi, and Mani B. Srivastava. Power management in energy harvesting sensor networks. ACM Transactions on Embedded Computing Systems, 6(4): 32:1–32:??, September 2007. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Kim:2017:AAS

[KJK17a] Sang-Hoon Kim, Jinkyu Jeong, and Jin-Soo Kim. Application-aware swapping for mobile systems. ACM Transactions on Embedded Computing Systems, 16(5s): 182:1–182:??, October 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Ko:2017:PCS

[KJK^{+}17b] Yohan Ko, Reiley Jeyapaul, Youngbin Kim, Kyoungwoo Lee, and Aviral Shrivastava. Protecting caches from soft errors: a microarchitect’s perspective. ACM Transactions on Embedded Computing Systems, 16(4):93:1– 93:??, August 2017. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Kim:2018:OND

[KJK18] Hyeonggyu Kim, Minho Ju, and Soontae Kim. On-Network+: Network delayaware management for mobile systems. ACM Transactions on Embedded Computing Systems, 17(3):64:1– 64:??, June 2018. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Kim:2016:SMR

[KJKM16] Sang-Hoon Kim, Jinkyu Jeong, Jin-Soo Kim, and Seungryoul Maeng. SmartLMK: a memory reclamation scheme for improving user-perceived app launch time. ACM Transactions on Embedded Computing Systems, 15(3): 47:1–47:??, July 2016. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Kadiyala:2020:HPC

[KJLS20] Sai Praveen Kadiyala, Pranav Jadhav, Siew-Kei Lam, and Thambipillai Srikanthan. Hardware performance counter-based finegrained malware detection. ACM Transactions on Embedded Computing Systems, 19(5):38:1–38:17, November 2020. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3403943.

Kim:2013:DER

[KJRG13] Sunwoo Kim, Won Seob Jeong, Won W. Ro, and Jean-Luc Gaudiot. Design and evaluation of random linear network coding accelerators on FPGAs. ACM Transactions on Embedded Computing Systems, 13(1): 13:1–13:??, August 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Kadayif:2005:DSO

[KK05a] I. Kadayif and M. Kandemir. Data space-oriented tiling for enhancing locality. ACM Transactions on Embedded Computing Systems, 4(2):388–414, May 2005. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Kwon:2005:OVA

[KK05b] Woo-Cheol Kwon and Taewhan Kim. Optimal voltage allocation techniques for dynamically variable voltage processors. ACM Transactions on Embedded Computing Systems, 4(1):211– 230, February 2005. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Kadayif:2005:CDH

[KKC^{+}05] I. Kadayif, M. Kandemir, G. Chen, N. Vijaykrishnan, M. J. Irwin, and A. Sivasubramaniam. Compilerdirected high-level energy estimation and optimization. ACM Transactions on Embedded Computing Systems, 4(4):819–850, November 2005. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Kim:2016:UMA

[KKCS16] Eui-Jik Kim, Jung-Hyok Kwon, Ken Choi, and Taeshik Shon. Unified medium access control architecture for resourceconstrained machine-to-machine devices. ACM Transactions on Embedded Computing Systems, 15(2):40:1– 40:??, May 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Khajeh:2012:EAA

[KKD^{+}12] Amin Khajeh, Minyoung Kim, Nikil Dutt, Ahmed M. Eltawil, and Fadi J. Kurdahi. Error-aware algorithm/architecture coexploration for video over wireless applications. ACM Transactions on Embedded Computing Systems, 11S(1): 15:1–15:??, 2012. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Koutsoukos:2012:PAM

[KKH^{+}12] Xenofon Koutsoukos, Nicholas Kottenstette, Joseph Hall, Emeka Eyisi, Heath Leblanc, Joseph Porter, and Janos Sztipanovits. A passivity approach for model-based compositional design of networked control systems. ACM Transactions on Embedded Computing Systems, 11(4):75:1–75:??, December 2012. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Kim:2011:DPT

[KKK^{+}11] Seungkyun Kim, Kiwon Kwon, Chihun Kim, Choonki Jang, Jaejin Lee, and Sang Lyul Min. Demand paging techniques for flash memory using compiler postpass optimizations. ACM Transactions on Embedded Computing Systems, 10(4): 40:1–40:??, November 2011. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Ko:2016:SBS

[KKL^{+}16] Yohan Ko, Jihoon Kang, Jongwon Lee, Yongjoo Kim, Joonhyun Kim, Hwisoo So, Kyoungwoo Lee, and Yunheung Paek. Software-based selective validation techniques for robust CGRAs against soft errors. ACM Transactions on Embedded Computing Systems, 15(1):20:1– 20:??, February 2016. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Kangas:2006:UBM

[KKO^{+}06] Tero Kangas, Petri Kukkala, Heikki Orsila, Erno Salminen, Marko Hännikäinen, Timo D. Hämäläinen, Jouni Riihimäki, and Kimmo Kuusilinna. UML-based multiprocessor SoC design framework. ACM Transactions on Embedded Computing Systems, 5(2):281–320, May 2006. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Kwon:2023:LRT

[KKS^{+}23] Hyeokdong Kwon, Hyunjun Kim, Minjoo Sim, Wai-Kong Lee, and Hwajeong Seo. Look-up the rainbow: Table-based implementation of rainbow signature on 64-bit ARMv8 processors. ACM Transactions on Embedded Computing Systems, 22(5):80:1–80:??, September 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3607140.

Khalgui:2013:ISI

[KL13] Mohamed Khalgui and Zhiwu Li. Introduction to the Special Issue on Modeling and Verification of Discrete Event Systems. ACM Transactions on Embedded Computing Systems, 12(1): 1:1–1:??, January 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Ko:2010:MME

[KLC^{+}10] Jeonggil Ko, Jong Hyun Lim, Yin Chen, Rvãzvan Musvaloiu-E, Andreas Terzis, Gerald M. Masson, Tia Gao, Walt Destler, Leo Selavo, and Richard P. Dutton. MEDiSN: Medical emergency detection in sensor networks. ACM Transactions on Embedded Computing Systems, 10(1):11:1– 11:??, August 2010. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Kim:2019:AAI

[KLK^{+}19] Jihye Kim, Jiwon Lee, Hankyung Ko, Donghwan Oh, Semin Han, Gwonho Jeong, and Hyunok Oh. AuthCropper: Authenticated image cropper for privacy preserving surveillance systems. ACM Transactions on Embedded Computing Systems, 18(5s):62:1–62:??, October 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/ft_gateway.
cfm?id=3358195.

Kornaros:2024:FUI

[KLK24] George Kornaros, Svoronos Leivadaros, and Filippos Kolimbianakis. Flexible updating of Internet of things computing functions through optimizing dynamic partial reconfiguration. ACM Transactions on Embedded Computing Systems, 23(2): 21:1–21:??, March 2024. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3643825.

Koo:2009:FTG

[KM09] Heon-Mo Koo and Prabhat Mishra. Functional test generation using design and property decomposition techniques. ACM Transactions on Embedded Computing Systems, 8(4): 32:1–32:??, July 2009. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Kirsch:2013:ISS

[KM13] Christoph Kirsch and Vincent Mooney. Introduction to special section on probabilistic embedded computing. ACM Transactions on Embedded Computing Systems, 12(2s):86:1–86:??, May 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Ko:2007:BSA

[KMB07] Ming-Yung Ko, Praveen K. Murthy, and Shuvra S. Bhattacharyya. Beyond singleappearance schedules: Efficient DSP software synthesis using nested procedure calls. ACM Transactions on Embedded Computing Systems, 6(2):14:1–14:??, May 2007. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Khalgui:2013:RRE

[KML13] Mohamed Khalgui, Olfa Mosbahi, and Zhiwu Li. Runtime reconfigurations of embedded controllers. ACM Transactions on Embedded Computing Systems, 12(1): 14:1–14:??, January 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Kang:2022:MLM

[KML^{+}22] Chih-Kai Kang, Hashan Roshantha Mendis, Chun-Han Lin, Ming-Syan Chen, and Pi-Cheng Hsiu. More is less: Model augmentation for intermittent deep inference. ACM Transactions on Embedded Computing Systems, 21(5):49:1–49:??, September 2022. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3506732.

Knapik:2015:ASB

[KMP15] Michal Knapik, Artur Meski, and Wojciech Penczek. Action synthesis for branching time logic: Theory and applications. ACM Transactions on Embedded Computing Systems, 14(4):64:1– 64:??, December 2015. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Koul:2023:AAA

[KMS^{+}23] Kalhan Koul, Jackson Melchert, Kavya Sreedhar, Leonard Truong, Gedeon Nyengele, Keyi Zhang, Qiaoyi Liu, Jeff Setter, Po-Han Chen, Yuchen Mei, Maxwell Strange, Ross Daly, Caleb Donovick, Alex Carsello, Taeyoung Kong, Kathleen Feng, Dillon Huff, Ankita Nayak, Rajsekhar Setaluri, James Thomas, Nikhil Bhagdikar, David Durst, Zachary Myers, Nestan Tsiskaridze, Stephen Richardson, Rick Bahr, Kayvon Fatahalian, Pat Hanrahan, Clark Barrett, Mark Horowitz, Christopher Torng, Fredrik Kjolstad, and Priyanka Raina. AHA: an agile approach to the design of coarse-grained reconfigurable accelerators and compilers. ACM Transactions on Embedded Computing Systems, 22(2):35:1– 35:??, March 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3534933.

Kundu:2025:SES

[KNK^{+}25] Suparna Kundu, Quinten Norga, Angshuman Karmakar, Shreya Gangopadhyay, Jose Maria Bermudo Mera, and Ingrid Verbauwhede. Scabbard: an exploratory study on hardware aware design choices of learning with rounding-based key encapsulation mechanisms. ACM Transactions on Embedded Computing Systems, 24(1):10:1–10:??, January 2025. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3696208.

Kumar:2012:ECI

[KNL12] Karthik Kumar, Yamini Nimmagadda, and Yung-Hsiang Lu. Energy conservation for image retrieval on mobile systems. ACM Transactions on Embedded Computing Systems, 11(3):66:1– 66:??, September 2012. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Kim:2017:PBB

[KNY^{+}17] Hong Seok Kim, Eyee Hyun Nam, Ji Hyuck Yun, Sheayun Lee, and Sang Lyul Min. P-BMS: a bad block management scheme in parallelized flash memory storage devices. ACM Transactions on Embedded Computing Systems, 16(5s):140:1– 140:??, October 2017. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Khan:2022:BIC

[KOL^{+}22] Asif Ali Khan, Sébastien Ollivier, Stephen Longofono, Gerald Hempel, Jeronimo Castrillon, and Alex K. Jones. Brain-inspired cognition in next-generation racetrack memories. ACM Transactions on Embedded Computing Systems, 21(6): 79:1–79:??, November 2022. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3524071.

Krishnakumar:2023:DSA

[KOM^{+}23] Anish Krishnakumar, Umit Ogras, Radu Marculescu, Mike Kishinevsky, and Trevor Mudge. Domainspecific architectures: Research problems and promising approaches. ACM Transactions on Embedded Computing Systems, 22(2):28:1– 28:??, March 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3563946.

Kent:2013:CPS

[KP13] Christopher G. Kent and Joann M. Paul. Contextual partitioning for speech recognition. ACM Transactions on Embedded Computing Systems, 13(1):12:1– 12:??, August 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Khalid:2016:RHL

[KPC^{+}16] Ayesha Khalid, Goutam Paul, Anupam Chattopadhyay, Faezeh Abediostad, Syed Imad Ud Din, Muhammad Hassan, Baishik Biswas, and Prasanna Ravi. Run-Stream: a high-level rapid prototyping framework for stream ciphers. ACM Transactions on Embedded Computing Systems, 15(3):61:1– 61:??, July 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Kim:2019:OBI

[KPK^{+}19] Minsu Kim, Jeong-Keun Park, Sungyeol Kim, Insu Yang, Hyunsoo Jung, and Soo-Mook Moon. Outputbased intermediate representation for translation of test-pattern program. ACM Transactions on Embedded Computing Systems, 18(5s): 55:1–55:??, October 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/ft_gateway.
cfm?id=3358186.

Khan:2023:DDL

[KPS23] Osama Khan, Gwanjong Park, and Euiseong Seo. Da-Capo: an on-device learning scheme for memoryconstrained embedded systems. ACM Transactions on Embedded Computing Systems, 22(5s):142:1–142:??, October 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3609121.

Khan:2014:OLT

[KR14] Umair Ali Khan and Bernhard Rinner. Online learning of timeout policies for dynamic power management. ACM Transactions on Embedded Computing Systems, 13(4):96:1–96:??, February 2014. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Kim:2018:PSC

[KR18] Hyoseung Kim and Ragunathan (Raj) Rajkumar. Predictable shared cache management for multicore real-time virtualization. ACM Transactions on Embedded Computing Systems, 17(1):22:1–22:??, January 2018. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Khan:2020:OTC

[KRHC20] Asif Ali Khan, Norman A. Rink, Fazal Hameed, and Jeronimo Castrillon. Optimizing tensor contractions for embedded devices with racetrack and DRAM memories. ACM Transactions on Embedded Computing Systems, 19(6):44:1–44:26, November 2020. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3396235.

Krishnakumar:2020:APL

[KRR20] Gnanambikai Krishnakumar, Kommuru Alekhya Reddy, and Chester Rebeiro. ALEXIA: a processor with lightweight extensions for memory safety. ACM Transactions on Embedded Computing Systems, 18(6):1–27, January 2020. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3362064.

Kriebel:2016:RAA

[KRS^{+}16] Florian Kriebel, Semeen Rehman, Arun Subramaniyan, Segnon Jean Bruno Ahandagbe, Muhammad Shafique, and Jörg Henkel. Reliabilityaware adaptations for shared last-level caches in multicores. ACM Transactions on Embedded Computing Systems, 15(4):67:1–67:??, August 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Kaiser:2010:ISI

[KS10] William Kaiser and Majid Sarrafzadeh. Introduction to special issue on wireless health. ACM Transactions on Embedded Computing Systems, 10(1):10:1– 10:??, August 2010. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Kim:2013:MPE

[KS13] Yooseong Kim and Aviral Shrivastava. Memory performance estimation of CUDA programs. ACM Transactions on Embedded Computing Systems, 13(2):21:1– 21:??, September 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Kalayappan:2018:PAH

[KS18] Rajshekar Kalayappan and Smruti R. Sarangi. Providing accountability in heterogeneous systems-on-chip. ACM Transactions on Embedded Computing Systems, 17(5):83:1–83:??, November 2018. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/ft_gateway.
cfm?id=3241048.

Krishnan:2022:BCS

[KS22] Archanaa S. Krishnan and Patrick Schaumont. Benchmarking and configuring security levels in intermittent computing. ACM Transactions on Embedded Computing Systems, 21(4):36:1– 36:??, July 2022. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3522748.

Kulkarni:2018:LOC

[KSA^{+}18] Amey Kulkarni, Colin Shea, Tahmid Abtahi, Houman Homayoun, and Tinoosh Mohsenin. Low overhead CSbased heterogeneous framework for big data acceleration. ACM Transactions on Embedded Computing Systems, 17(1):25:1– 25:??, January 2018. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Karl:2024:PQS

[KSFS24] Patrick Karl, Jonas Schupp, Tim Fritzmann, and Georg Sigl. Post-quantum signatures on RISC-V with hardware acceleration. ACM Transactions on Embedded Computing Systems, 23(2): 30:1–30:??, March 2024. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3579092.

Kang:2013:AEC

[KSK13] Hyeong-Ju Kang, Heesuk Seo, and Jin Kwak. Areaefficient convolutional deinterleaver for mobile TV receiver. ACM Transactions on Embedded Computing Systems, 12(2):28:1– 28:??, February 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Kim:2012:FLF

[KSP^{+}12] Jaegeuk Kim, Hyotaek Shim, Seon-Yeong Park, Seungryoul Maeng, and Jin-Soo Kim. FlashLight: a lightweight flash file system for embedded systems. ACM Transactions on Embedded Computing Systems, 11S(1):18:1– 18:??, 2012. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Kartal:2016:MDR

[KSS16] Yusuf Bora Kartal, Ece Güran Schmidt, and Klaus Werner Schmidt. Modeling distributed real-time systems in TIOA and UPPAAL. ACM Transactions on Embedded Computing Systems, 16(1):22:1–22:??, November 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Kashikar:2024:CWA

[KSS24] Prachi Kashikar, Olivier Sentieys, and Sharad Sinha. Combining weight approximation, sharing and retraining for neural network model compression. ACM Transactions on Embedded Computing Systems, 23(6):99:1– 99:??, November 2024. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3687466.

Kim:2012:XFM

[KST^{+}12] Minyoung Kim, Mark-Oliver Stehr, Carolyn Talcott, Nikil Dutt, and Nalini Venkatasubramanian. xTune: a formal methodology for crosslayer tuning of mobile embedded systems. ACM Transactions on Embedded Computing Systems, 11(4): 73:1–73:??, December 2012. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Kang:2017:RLA

[KSY17] Wonkyung Kang, Dongkun Shin, and Sungjoo Yoo. Reinforcement learning-assisted garbage collection to mitigate long-tail latency in SSD. ACM Transactions on Embedded Computing Systems, 16(5s):134:1–134:??, October 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Kumar:2014:WCG

[KT14] Pratyush Kumar and Lothar Thiele. Worst-case guarantees on a processor with temperature-based feedback control of speed. ACM Transactions on Embedded Computing Systems, 13(4s): 122:1–122:??, April 2014. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Kyrkou:2013:HAR

[KTT13] Christos Kyrkou, Christos Ttofis, and Theocharis Theocharides. A hardware architecture for realtime object detection using depth and edge information. ACM Transactions on Embedded Computing Systems, 13(3):54:1–54:??, December 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Kim:2003:PIC

[KVK^{+}03] Soontae Kim, N. Vijaykrishnan, Mahmut Kandemir, Anand Sivasubramaniam, and Mary Jane Irwin. Partitioned instruction cache architecture for energy efficiency. ACM Transactions on Embedded Computing Systems, 2(2):163–185, May 2003. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Kejariwal:2009:ELL

[KVN^{+}09] Arun Kejariwal, Alexander V. Veidenbaum, Alexandru Nicolau, Milind Girkar, Xinmin Tian, and Hideki Saito. On the exploitation of loop-level parallelism in embedded applications. ACM Transactions on Embedded Computing Systems, 8(2): 10:1–10:??, January 2009. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Kim:2010:EAE

[KW10] Minseong Kim and Andy Wellings. Efficient asynchronous event handling in the real-time specification for Java. ACM Transactions on Embedded Computing Systems, 10(1):5:1– 5:??, August 2010. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Klashtorny:2023:PGW

[KWKP23] Artem Klashtorny, Zhuanhao Wu, Anirudh Mohan Kaushik, and Hiren Patel. Predictable GPU wavefront splitting for safety-critical systems. ACM Transactions on Embedded Computing Systems, 22(5s):107:1– 107:??, October 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3609102.

Kwon:2016:CBF

[Kwo16] Se Jin Kwon. A cache-based flash translation layer for TLC-based multimedia storage devices. ACM Transactions on Embedded Computing Systems, 15(1):11:1– 11:??, February 2016. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Klues:2010:LLD

[KXL10] Kevin Klues, Guoliang Xing, and Chenyang Lu. Link layer driver architecture for unified radio power management in wireless sensor networks. ACM Transactions on Embedded Computing Systems, 9(4):41:1– 41:??, March 2010. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Khouzani:2017:DBS

[KY17] Hoda Aghaei Khouzani and Chengmo Yang. A DWMbased stack architecture implementation for energy harvesting systems. ACM Transactions on Embedded Computing Systems, 16(5s): 155:1–155:??, October 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Kindt:2020:EMB

[KYDC20] Philipp H. Kindt, Daniel Yunge, Robert Diemer, and Samarjit Chakraborty. Energy modeling for the Bluetooth low energy protocol. ACM Transactions on Embedded Computing Systems, 19(2):13:1–13:32, March 2020. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL

https://dl.acm.org/doi/
abs/10.1145/3379339.

Koohi:2014:TSL

[KYHY14] Somayyeh Koohi, Yawei Yin, Shaahin Hessabi, and S. J. Ben Yoo. Towards a scalable, low-power all-optical architecture for networks-on-chip. ACM Transactions on Embedded Computing Systems, 13(3s): 101:1–101:??, March 2014. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Kim:2013:NCA

[KYL13] Dongki Kim, Sungjoo Yoo, and Sunggu Lee. A network congestion-aware memory subsystem for manycore. ACM Transactions on Embedded Computing Systems, 12(4):110:1–110:??, June 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Kulkarni:2006:VVI

[KZH^{+}06] Prasad Kulkarni, Wankang Zhao, Stephen Hines, David Whalley, Xin Yuan, Robert van Engelen, Kyle Gallivan, Jason Hiser, Jack Davidson, Baosheng Cai, Mark Bailey, Hwashin Moon, Kyunghwan Cho, and Yunheung Paek. VISTA: VPO interactive system for tuning applications. ACM Transactions on Embedded Computing Systems, 5(4): 819–863, November 2006. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Liu:2011:NBF

[LA11] Xue Liu and Tarek Abdelzaher. Nonutilization bounds and feasible regions for arbitrary fixed-priority policies. ACM Transactions on Embedded Computing Systems, 10(3):31:1–31:??, April 2011. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Lee:2023:CVA

[LAB^{+}23] Edward A. Lee, Ravi Akella, Soroush Bateni, Shaokai Lin, Marten Lohstroh, and Christian Menard. Consistency vs. availability in distributed cyber-physical systems. ACM Transactions on Embedded Computing Systems, 22(5s):138:1– 138:??, October 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3609119.

Luo:2006:EEI

[LAHS06] Liqian Luo, Tarek F. Abdelzaher, Tian He, and John A. Stankovic. EnviroSuite: an environmentally immersive programming framework for sensor networks. ACM Transactions on Embedded Computing Systems, 5(3):543– 576, August 2006. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Liang:2024:REE

[LAMS24] Dehua Liang, Hiromitsu Awano, Noriyuki Miura, and Jun Shiomi. A robust and energy efficient hyperdimensional computing system for voltage-scaled circuits. ACM Transactions on Embedded Computing Systems, 23(6):91:1–91:??, November 2024. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3620671.

Lapalme:2006:NEE

[LAN06] James Lapalme, El Mostapha Aboulhamid, and Gabriela Nicolescu. A new efficient EDA tool design methodology. ACM Transactions on Embedded Computing Systems, 5(2):408–430, May 2006. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Liu:2016:SMA

[LAZ^{+}16] Lu Liu, Nick Antonopoulos, Minghui Zheng, Yongzhao Zhan, and Zhijun Ding. A socioecological model for advanced service discovery in machine-to-machine communication networks. ACM Transactions on Embedded Computing Systems, 15(2): 38:1–38:??, May 2016. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Lach:2004:ESI

[LB04] John Lach and Kia Bazargan. Editorial: Special issue on dynamically adaptable embedded systems. ACM Transactions on Embedded Computing Systems, 3(2): 233–236, May 2004. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Lohstroh:2024:DCA

[LBM^{+}24] Marten Lohstroh, Soroush Bateni, Christian Menard, Alexander Schulz-Rosengarten, Jeronimo Castrillon, and Edward A. Lee. Deterministic coordination across multiple timelines. ACM Transactions on Embedded Computing Systems, 23(5):77:1– 77:??, September 2024. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3615357.

Loghi:2007:PMM

[LBP07] Mirko Loghi, Luca Benini, and Massimo Poncino. Power macromodeling of MPSoC message passing primitives. ACM Transactions on Embedded Computing Systems, 6(4):31:1–31:??, September 2007. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Lu:2015:ECA

[LBS15] Jing Lu, Ke Bai, and Aviral Shrivastava. Efficient code assignment techniques for local memory on software managed multicores. ACM Transactions on Embedded Computing Systems, 14(4):71:1–71:??, December 2015. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Li:2017:TPR

[LC17] Yu Li and Albert M. K.

Cheng. Toward a practical regularity-based model: The impact of evenly distributed temporal resource partitions. ACM Transactions on Embedded Computing Systems, 16(4):111:1–111:??, August 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Liang:2019:ESW

[LCC^{+}19] Yu-Pei Liang, Tseng-Yi Chen, Yuan-Hao Chang, Shuo-Han Chen, Kam-Yiu Lam, Wei-Hsin Li, and Wei-Kuan Shih. Enabling sequential-write-constrained B+-tree index scheme to upgrade shingled magnetic recording storage performance. ACM Transactions on Embedded Computing Systems, 18(5s):66:1– 66:??, October 2019. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3358201.

Lien:2023:FFS

[LCC^{+}23] Yi-Han Lien, Yen-Ting Chen, Yuan-Hao Chang, Yu-Pei Liang, and Wei-Kuan Shih. FSIMR: File-system-aware data management for interlaced magnetic recording. ACM Transactions on Embedded Computing Systems, 22(5s):128:1–128:??, October 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3607922.

Lee:2018:FRT

[LCD18] Ganghee Lee, Ediz Cetin, and Oliver Diessel. Fault recovery time analysis for coarse-grained reconfigurable architectures. ACM Transactions on Embedded Computing Systems, 17(2):42:1– 42:??, April 2018. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Liu:2008:HPP

[LCH^{+}08] Duo Liu, Zheng Chen, Bei Hua, Nenghai Yu, and Xinan Tang. High-performance packet classification algorithm for multithreaded IXP network processor. ACM Transactions on Embedded Computing Systems, 7(2): 16:1–16:??, February 2008. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Lim:2013:DRS

[LCJ13] Seung-Ho Lim, Min Choi, and Young Sik Jeong. Data reorganization for scalable video service with embedded mobile devices. ACM Transactions on Embedded Computing Systems, 12(2): 27:1–27:??, February 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Liu:2019:CDM

[LCL^{+}19] Daibo Liu, Zhichao Cao, Mingyan Liu, Mengshu Hou, and Hongbo Jinag. Contention-detectable mechanism for receiver-initiated MAC. ACM Transactions on Embedded Computing Systems, 18(4):31:1– 31:??, August 2019. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3317683.

Li:2017:RNF

[LCLW17] Yang Li, Mengting Chen, Zhe Liu, and Jian Wang. Reduction in the number of fault injections for blind fault attack on SPN block ciphers. ACM Transactions on Embedded Computing Systems, 16(2):55:1–55:??, April 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Lee:2017:MAA

[LCP^{+}17] Jaewoo Lee, Hoon Sung Chwa, Linh T. X. Phan, Insik Shin, and Insup Lee. MC-ADAPT: Adaptive task dropping in mixed-criticality scheduling. ACM Transactions on Embedded Computing Systems, 16(5s):163:1– 163:??, October 2017. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Lai:2013:RBR

[LCQ^{+}13] Chin-Feng Lai, Min Chen, Meikang Qiu, Athanasios V. Vasilakos, and Jong Hyuk Park. A RF4CE-based remote controller with interactive graphical user interface applied to home automation system. ACM Transactions on Embedded Computing Systems, 12(2): 30:1–30:??, February 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Lin:2003:CMC

[LCS03] Kelvin Lin, Chung-Ping Chung, and Jean Jyh-Jiun Shann. Compressing MIPS code by multiple operand dependencies. ACM Transactions on Embedded Computing Systems, 2(4):482– 508, November 2003. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Lin:2022:DRR

[LCY^{+}22] Wei-Ting Lin, Hsiang-Yun Cheng, Chia-Lin Yang, Meng-Yao Lin, Kai Lien, Han-Wen Hu, Hung-Sheng Chang, Hsiang-Pang Li, Meng-Fan Chang, Yen-Ting Tsou, and Chin-Fu Nien. DL-RSIM: a reliability and deployment strategy simulation framework for ReRAMbased CNN accelerators. ACM Transactions on Embedded Computing Systems, 21(3):24:1–24:29, May 2022. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3507639.

Lakshmanan:2012:OPM

[LDRM12] Karthik Lakshmanan, Dionisio De Niz, Ragunathan (Raj) Rajkumar, and Gabriel Moreno. Overload provisioning in mixed-criticality cyber-physical systems. ACM Transactions on Embedded Computing Systems, 11(4): 83:1–83:??, December 2012. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Lee:2012:EEA

[LDV12] Kyoungwoo Lee, Nikil Dutt, and Nalini Venkatasubramanian. EAVE: Error-Aware Video Encoding supporting extended energy/QoS tradeoffs for mobile embedded systems. ACM Transactions on Embedded Computing Systems, 11(2):37:1–37:??, July 2012. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Lee:2021:D

[Lee21] Edward A. Lee. Determinism. ACM Transactions on Embedded Computing Systems, 20(5):38:1–38:34, July 2021. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3453652.

Leonard:2018:GES

[Leo18] Elizabeth Leonard. Guest editorial: Special issue on formal methods and models for system design. ACM Transactions on Embedded Computing Systems, 17(2): 30:1–30:??, April 2018. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Lingamneni:2013:SPI

[LEPP13] Avinash Lingamneni, Christian Enz, Krishna Palem, and Christian Piguet. Synthesizing parsimonious inexact circuits through probabilistic design techniques. ACM Transactions on Embedded Computing Systems, 12(2s):93:1–93:??, May 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Lee:2014:CFE

[LES14] Jinkyu Lee, Arvind Easwaran, and Insik Shin. Contentionfree executions for real-time multiprocessor scheduling. ACM Transactions on Embedded Computing Systems, 13(2s):69:1–69:??, January 2014. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Llopard:2017:FPA

[LFC17] Ivan Llopard, Christian Fabre, and Albert Cohen. From a formalized parallel action language to its efficient code generation. ACM Transactions on Embedded Computing Systems, 16(2): 37:1–37:??, April 2017. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Li:2018:NOC

[LFHS18] Songyuan Li, Lingkun Fu, Shibo He, and Youxian Sun. Near-optimal co-deployment of chargers and sink stations in rechargeable sensor networks. ACM Transactions on Embedded Computing Systems, 17(1):10:1– 10:??, January 2018. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Langerman:2021:RTH

[LG21] David Langerman and Alan George. Real-time, highresolution depth upsampling on embedded accelerators. ACM Transactions on Embedded Computing Systems, 20(3):18:1–18:22, April 2021. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3436878.

Li:2018:FPS

[LH18] Zheng Li and Shuibing He. Fixed-priority scheduling for two-phase mixed-criticality systems. ACM Transactions on Embedded Computing Systems, 17(2):35:1– 35:??, April 2018. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Liu:2004:FSM

[LHCK04] Rong-Tai Liu, Nen-Fu Huang, Chih-Hao Chen, and Chia-Nan Kao. A fast stringmatching algorithm for network processor-based intrusion detection system. ACM Transactions on Embedded Computing Systems, 3(3): 614–633, August 2004. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Lin:2019:GBM

[LHL^{+}19] Yi-Ting Lin, Hsiang Hsu, Shang-Chien Lin, Chung-Wei Lin, Iris Hui-Ru Jiang, and Changliu Liu. Graphbased modeling, scheduling, and verification for intersection management of intelligent vehicles. ACM Transactions on Embedded Computing Systems, 18(5s):95:1– 95:??, October 2019. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3358221.

Logaras:2014:PAE

[LHM14] Evangelos Logaras, Orsalia G. Hazapis, and Elias S. Manolakos. Python to accelerate embedded SoC design: a case study for systems biology. ACM Transactions on Embedded Computing Systems, 13(4):84:1– 84:??, February 2014. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Lei:2023:FII

[LHP^{+}23] Douwei Lei, Debiao He, Cong Peng, Min Luo, Zhe Liu, and Xinyi Huang. Faster implementation of ideal lattice-based cryptography using AVX512. ACM Transactions on Embedded Computing Systems, 22(5): 83:1–83:??, September 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3609223.

Liu:2014:PRR

[LHX^{+}14] Hengchang Liu, Pan Hui, Zhiheng Xie, Jingyuan Li, David Siu, Gang Zhou, Liusheng Huang, and John A. Stankovic. Providing reliable and real-time delivery in the presence of body shadowing in breadcrumb systems. ACM Transactions on Embedded Computing Systems, 13(4):94:1–94:??, February 2014. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Liang:2018:DFM

[LHYQ18] Xiaoxuan Liang, Zhangqin Huang, Shengqi Yang, and Lanxin Qiu. Device-free motion & trajectory detection via RFID. ACM Transactions on Embedded Computing Systems, 17(4):78:1– 78:??, August 2018. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Li:2021:HCO

[Li21] Keqin Li. Heuristic computation offloading algorithms for mobile users in fog computing. ACM Transactions on Embedded Computing Systems, 20(2):11:1– 11:28, March 2021. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3426852.

Gal:2014:GLC

[LJ14] Bertrand Le Gal and Christophe Jego. GPU-like on-chip system for decoding LDPC codes. ACM Transactions on Embedded Computing Systems, 13(4):95:1–95:??, February 2014. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Li:2024:DMD

[LJ24] Chia-Hao Li and Niraj K.

Jha. DOCTOR: a multidisease detection continual learning framework based on wearable medical sensors. ACM Transactions on Embedded Computing Systems, 23(5):81:1–81:??, September 2024. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3679050.

Liu:2017:CDS

[LJLT17] Qingrui Liu, Changhee Jung, Dongyoon Lee, and Devesh Tiwari. Compiler-directed soft error detection and recovery to avoid DUE and SDC via Tail-DMR. ACM Transactions on Embedded Computing Systems, 16(2): 32:1–32:??, April 2017. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Lindemann:2023:RSS

[LJMP23] Lars Lindemann, Lejun Jiang, Nikolai Matni, and George J. Pappas. Risk of stochastic systems for temporal logic specifications. ACM Transactions on Embedded Computing Systems, 22(3):54:1–54:??, May 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3580490.

Lesi:2017:SAS

[LJP17] Vuk Lesi, Ilija Jovanov, and Miroslav Pajic. Securityaware scheduling of embedded control tasks. ACM Transactions on Embedded Computing Systems, 16(5s): 188:1–188:??, October 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Li:2012:SRS

[LJR12] Chunxiao Li, Niraj K. Jha, and Anand Raghunathan. Secure reconfiguration of software-defined radio. ACM Transactions on Embedded Computing Systems, 11(1): 10:1–10:??, March 2012. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Lohar:2023:SMF

[LJVD23] Debasmita Lohar, Clothilde Jeangoudoux, Anastasia Volkova, and Eva Darulova. Sound mixed fixed-point quantization of neural networks. ACM Transactions on Embedded Computing Systems, 22(5s):136:1– 136:??, October 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3609118.

Lee:2010:CPV

[LK10] Hyung Sun Lee and Byung Kook Kim. Coscheduling of processor voltage and control task period for energyefficient control systems. ACM Transactions on Embedded Computing Systems, 9(3):15:1–15:??, February 2010. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Leech:2018:RPP

[LKA^{+}18] Charles Leech, Charan Kumar, Amit Acharyya, Sheng Yang, Geoff V. Merrett, and Bashir M. Al-Hashimi. Runtime performance and power optimization of parallel disparity estimation on manycore platforms. ACM Transactions on Embedded Computing Systems, 17(2):41:1– 41:??, April 2018. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Lashgar:2014:HHI

[LKB14] Ahmad Lashgar, Ahmad Khonsari, and Amirali Baniasadi. HARP: Harnessing in-Active thReads in many-core Processors. ACM Transactions on Embedded Computing Systems, 13(3s):114:1– 114:??, March 2014. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Lin:2016:CFQ

[LKH16] Chun-Han Lin, Chih-Kai Kang, and Pi-Cheng Hsiu. CURA: a framework for quality-retaining power saving on mobile OLED displays. ACM Transactions on Embedded Computing Systems, 15(4):76:1–76:??, August 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Lee:2002:AAI

[LKW02] Jung-Hoon Lee, Shin-Dug Kim, and Charles Weems. Application-adaptive intelligent cache memory system. ACM Transactions on Embedded Computing Systems, 1(1):56–78, November 2002. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Li:2024:IHA

[LKW^{+}24] Chaojian Li, Kyungmin Kim, Bichen Wu, Peizhao Zhang, Hang Zhang, Xiaoliang Dai, Peter Vajda, and Yingyan (Celine) Lin. An investigation on hardwareaware vision transformer scaling. ACM Transactions on Embedded Computing Systems, 23(3):52:1– 52:??, May 2024. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3611387.

Li:2023:EDS

[LKZ^{+}23] Yueting Li, Wang Kang, Kunyu Zhou, Keni Qiu, and Weisheng Zhao. Experimental demonstration of STT-MRAM-based nonvolatile instantly on/off system for IoT applications: Case studies. ACM Transactions on Embedded Computing Systems, 22(2):29:1– 29:??, March 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3546193.

Lazarescu:2015:ITB

[LL15] Mihai T. Lazarescu and Luciano Lavagno. Interactive trace-based analysis toolset for manual parallelization of C programs. ACM Transactions on Embedded Computing Systems, 14(1): 13:1–13:??, January 2015. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Liang:2017:EKM

[LL17] Yun Liang and Xiuhong Li.

Efficient kernel management on GPUs. ACM Transactions on Embedded Computing Systems, 16(4):115:1– 115:??, August 2017. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Lu:2018:TSI

[LL18] Sixing Lu and Roman Lysecky. Time and sequence integrated runtime anomaly detection for embedded systems. ACM Transactions on Embedded Computing Systems, 17(2):38:1–38:??, April 2018. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Lo:2013:AGH

[LLC^{+}13] Chen-Kang Lo, Mao-Lin Li, Li-Chun Chen, Yi-Shan Lu, Ren-Song Tsay, Hsu-Yao Huang, and Jen-Chieh Yeh. Automatic generation of high-speed accurate TLM models for out-of-order pipelined bus. ACM Transactions on Embedded Computing Systems, 13(1s):37:1– 37:??, November 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Lin:2022:HEI

[LLC^{+}22] Yi-Syuan Lin, Yu-Pei Liang, Tseng-Yi Chen, Yuan-Hao Chang, Shuo-Han Chen, Hsin-Wen Wei, and Wei-Kuan Shih. How to enable index scheme for reducing the writing cost of DNA storage on insertion and deletion. ACM Transactions on Embedded Computing Systems, 21(3):30:1–30:25, May 2022. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3516482.

Li:2020:HSC

[LLG^{+}20] Mengquan Li, Weichen Liu, Nan Guan, Yiyuan Xie, and Yaoyao Ye. Hardwaresoftware collaborative thermal sensing in optical network-on-chip–based manycore systems. ACM Transactions on Embedded Computing Systems, 18(6):1–24, January 2020. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL

https://dl.acm.org/doi/
abs/10.1145/3362099.

Luo:2025:EDL

[LLK^{+}25] Xiangzhong Luo, Di Liu, Hao Kong, Shuo Huai, Hui Chen, Guochu Xiong, and Weichen Liu. Efficient deep learning infrastructures for embedded computing systems: a comprehensive survey and future envision. ACM Transactions on Embedded Computing Systems, 24(1):21:1–21:??, January 2025. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3701728.

Li:2014:BAM

[LLL14] Shuai Li, Yuesheng Lou, and Bo Liu. Bluetooth aided mobile phone localization: a nonlinear neural circuit approach. ACM Transactions on Embedded Computing Systems, 13(4): 78:1–78:??, February 2014. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Lizarraga:2013:DPF

[LLLGR13] Adrian Lizarraga, Roman Lysecky, Susan Lysecky, and Ann Gordon-Ross. Dynamic profiling and fuzzylogic-based optimization of sensor network platforms. ACM Transactions on Embedded Computing Systems, 13(3):51:1–51:??, December 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Lin:2008:MAC

[LLLT08] Yi-Neng Lin, Ying-Dar Lin, Yuan-Cheng Lai, and Kuo-Kun Tseng. Modeling and analysis of core-centric network processors. ACM Transactions on Embedded Computing Systems, 7(4): 41:1–41:??, July 2008. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Lee:2009:CIA

[LLN09] Edward A. Lee, Xiaojun Liu, and Stephen Neuendorffer. Classes and inheritance in actor-oriented design. ACM Transactions on Embedded Computing Systems, 8(4):29:1–29:??, July 2009. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Liu:2014:STD

[LLN^{+}14] Kai Liu, Victor C. S. Lee, Joseph K. Y. Ng, Sang H. Son, and Edwin H.-M. Sha. Scheduling temporal data with dynamic snapshot consistency requirement in vehicular cyber-physical systems. ACM Transactions on Embedded Computing Systems, 13(5s):163:1–163:??, September 2014. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Lu:2017:CDS

[LLP^{+}17] Qining Lu, Guanpeng Li, Karthik Pattabiraman, Meeta S. Gupta, and Jude A. Rivers. Configurable detection of SDC-causing errors in programs. ACM Transactions on Embedded Computing Systems, 16(3):88:1– 88:??, July 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Leon:2021:IPP

[LLP^{+}21] Vasileios Leon, George Lentaris, Evangelos Petrongonas, Dimitrios Soudris, Gianluca Furano, Antonis Tavoularis, and David Moloney. Improving performancepower-programmability in space avionics with edge devices: VBN on myriad2 SoC. ACM Transactions on Embedded Computing Systems, 20(3):22:1–22:23, April 2021. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3440885.

Lee:2007:SCT

[LLPM07] Sheayun Lee, Jaejin Lee, Chang Yun Park, and Sang Lyul Min. Selective code transformation for dual instruction set processors. ACM Transactions on Embedded Computing Systems, 6(2):10:1–10:??, May 2007. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Li:2014:MRT

[LLR14] Zheng Li, Frank Lockom, and Shangping Ren. Maintaining real-time application timing similarity for defecttolerant NoC-based manycore systems. ACM Transactions on Embedded Computing Systems, 13(2s):64:1– 64:??, January 2014. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Liu:2017:PSSa

[LLT^{+}17] Anfeng Liu, Xiao Liu, Zhipeng Tang, Laurence T. Yang, and Zili Shao. Preserving smart sink-location privacy with delay guaranteed routing scheme for WSNs. ACM Transactions on Embedded Computing Systems, 16(3):68:1–68:??, July 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Lin:2009:MAC

[LLTL09] Yi-Neng Lin, Ying-Dar Lin, Kuo-Kun Tseng, and Yuan-Cheng Lai. Modeling and analysis of core-centric network processors. ACM Transactions on Embedded Computing Systems, 8(2): 13:1–13:??, January 2009. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Liu:2017:DMR

[LLW^{+}17] Anfeng Liu, Xiao Liu, Tianyi Wei, Laurence T. Yang, Seungmin (Charlie) Rho, and Anand Paul. Distributed multi-representative re-fusion approach for heterogeneous sensing data collection. ACM Transactions on Embedded Computing Systems, 16(3): 73:1–73:??, July 2017. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Li:2024:ATV

[LLW^{+}24] Fenfang Li, Huizhang Luo, Junqi Wang, Yida Li, Zhuo Tang, and Kenli Li. AMP: Total variation reduction for lossless compression via approximate median-based preconditioning. ACM Transactions on Embedded Computing Systems, 23(6): 86:1–86:??, November 2024. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3605359.

Liu:2017:MCS

[LLZ^{+}17] Jing Liu, Kenli Li, Dakai Zhu, Jianjun Han, and Keqin Li. Minimizing cost of scheduling tasks on heterogeneous multicore embedded systems. ACM Transactions on Embedded Computing Systems, 16(2):36:1– 36:??, April 2017. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Liao:2022:RRS

[LLZ^{+}22] Jianwei Liao, Jun Li, Mingwang Zhao, Zhibing Sha, and Zhigang Cai. Read refresh scheduling and data reallocation against read disturb in SSDs. ACM Transactions on Embedded Computing Systems, 21(2):18:1– 18:27, March 2022. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3495254.

Liang:2013:AAF

[LM13] Yun Liang and Tulika Mitra. An analytical approach for fast and accurate design space exploration of instruction caches. ACM Transactions on Embedded Computing Systems, 13(3):43:1–43:??, December 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Lu:2024:EEE

[LM24] Qianyun Lu and Boris Murmann. Enhancing the energy efficiency and robustness of tinyML computer vision using coarsely-quantized log-gradient input images. ACM Transactions on Embedded Computing Systems, 23(3):41:1–41:??, May 2024. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3591466.

Lee:2019:ESA

[LMA19] Robert P. Lee, Konstantinos Markantonakis, and Raja Naeem Akram. Ensuring secure application execution and platform-specific execution in embedded devices. ACM Transactions on Embedded Computing Systems, 18(3):26:1–26:??, June 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/ft_gateway.
cfm?id=3284361.

Lemaire:2022:SAH

[LMB^{+}22] Edgar Lemaire, Benoît Miramond, Sébastien Bilavarn, Hadi Saoud, and Nassim Abderrahmane. Synaptic activity and hardware footprint of spiking neural networks in digital neuromorphic systems. ACM Transactions on Embedded Computing Systems, 21(6):75:1– 75:??, November 2022. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3520133.

Lohstroh:2021:TLF

[LMBL21] Marten Lohstroh, Christian Menard, Soroush Bateni, and Edward A. Lee. Toward a lingua franca for deterministic concurrent systems. ACM Transactions on Embedded Computing Systems, 20(4):36:1–36:27, June 2021. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3448128.

Liang:2018:EVL

[LMK^{+}18] Lihao Liang, Tom Melham, Daniel Kroening, Peter Schrammel, and Michael Tautschnig. Effective verification for low-level software with competing interrupts. ACM Transactions on Embedded Computing Systems, 17(2):36:1–36:??, April 2018. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Lukyanov:2020:FVS

[LML20] Georgy Lukyanov, Andrey Mokhov, and Jakob Lechner. Formal verification of spacecraft control programs. ACM Transactions on Embedded Computing Systems, 19(5):37:1–37:18, November 2020. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3391900.

Lin:2023:TBV

[LML^{+}23] Shaokai Lin, Yatin A. Manerkar, Marten Lohstroh, Elizabeth Polgreen, Sheng-Jung Yu, Chadlia Jerad, Edward A. Lee, and Sanjit A. Seshia. Towards building verifiable CPS using Lingua Franca. ACM Transactions on Embedded Computing Systems, 22(5s):155:1– 155:??, October 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3609134.

Lentaris:2019:SMF

[LMS^{+}19] George Lentaris, Konstantinos Maragos, Dimitrios Soudris, Xenophon Zabulis, and Manolis Lourakis. Single- and multi-FPGA acceleration of dense stereo vision for planetary rovers.

ACM Transactions on Embedded Computing Systems, 18(2):16:1–16:??, April 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

                                        /dl.acm.org/ft_gateway.
                                        cfm?id=3312743.

Li:2022:FDN

[LMS^{+}22] Shuwei Li, Changhai Man, Ao Shen, Ziyi Guan, Wei Mao, Shaobo Luo, Rumin Zhang, and Hao Yu. A fall detection network by 2D/ 3D spatio-temporal joint models with tensor compression on edge. ACM Transactions on Embedded Computing Systems, 21(6):83:1– 83:??, November 2022. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3531004.

Lanotte:2004:IFH

[LMST04] Ruggero Lanotte, Andrea Maggiolo-Schettini, and Simone Tini. Information flow in hybrid systems. ACM Transactions on Embedded Computing Systems, 3(4): 760–799, November 2004. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Liu:2017:PMH

[LMW^{+}17] Jianghua Liu, Jinhua Ma, Wei Wu, Xiaofeng Chen, Xinyi Huang, and Li Xu. Protecting mobile health records in cloud computing: a secure, efficient, and anonymous design. ACM Transactions on Embedded Computing Systems, 16(2): 57:1–57:??, April 2017. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Liu:2004:MBA

[LN04] Donggang Liu and Peng Ning. Multilevel μTESLA: Broadcast authentication for distributed sensor networks. ACM Transactions on Embedded Computing Systems, 3(4):800– 836, November 2004. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Leipnitz:2019:HLS

[LN19] Marcos T. Leipnitz and Gabriel L. Nazar. Highlevel synthesis of approximate designs under real-time constraints. ACM Transactions on Embedded Computing Systems, 18(5s):59:1– 59:??, October 2019. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3358182.

Loke:2015:MCS

[LNA^{+}15] Seng W. Loke, Keegan Napier, Abdulaziz Alali, Niroshinie Fernando, and Wenny Rahayu. Mobile computations with surrounding devices: Proximity sensing and MultiLayered work stealing. ACM Transactions on Embedded Computing Systems, 14(2):22:1– 22:??, March 2015. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Lee:2013:LAB

[LO13] Daeyoung Lee and Hyunok Oh. A lifetime aware buffer assignment method for streaming applications on DRAM/PRAM hybrid memory. ACM Transactions on Embedded Computing Systems, 12(1s):36:1– 36:??, March 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Li:2018:ECB

[LOD18] He Li, Kaoru Ota, and Mianxiong Dong. Energy cooperation in battery-free wireless communications with radio frequency energy harvesting. ACM Transactions on Embedded Computing Systems, 17(2):44:1– 44:??, April 2018. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Luppold:2020:CWC

[LOF20] Arno Luppold, Dominic Oehlert, and Heiko Falk. Compiling for the worst case: Memory allocation for multitask and multi-core hard real-time systems. ACM Transactions on Embedded Computing Systems, 19(2): 14:1–14:26, March 2020. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3381752.

Lhuillier:2014:HHA

[LOG^{+}14] Yves Lhuillier, Maroun Ojail, Alexandre Guerre, Jean-Marc Philippe, Karim Ben Chehida, Farhat Thabet, Caaliph Andriamisaina, Chafic Jaber, and Raphaël David. HARS: a hardware-assisted runtime software for embedded many-core architectures. ACM Transactions on Embedded Computing Systems, 13(3s):102:1–102:??, March 2014. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Liu:2013:RAE

[LOXL13] Tiantian Liu, Alex Orailoglu, Chun Jason Xue, and Minming Li. Register allocation for embedded systems to simultaneously reduce energy and temperature on registers. ACM Transactions on Embedded Computing Systems, 13(3):50:1– 50:??, December 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Li:2009:ELC

[LP09a] Zhiyuan Li and Santosh Pande. Editorial: Languages, compilers, and tools for embedded systems. ACM Transactions on Embedded Computing Systems, 8(4): 25:1–25:??, July 2009. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Lubbers:2009:RMP

[LP09b] Enno Lübbers and Marco Platzner. ReconOS: Multithreaded programming for reconfigurable computers. ACM Transactions on Embedded Computing Systems, 9(1):8:1–8:??, October 2009. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Lee:2010:IHM

[LP10] Jupyung Lee and Kyu Ho Park. Interrupt handler migration and direct interrupt scheduling for rapid scheduling of interrupt-driven tasks. ACM Transactions on Embedded Computing Systems, 9(4):42:1–42:??, March 2010. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Lal:2019:CGA

[LP19] Ratan Lal and Pavithra Prabhakar. Counterexample guided abstraction refinement for polyhedral probabilistic hybrid systems. ACM Transactions on Embedded Computing Systems, 18(5s): 98:1–98:??, October 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/ft_gateway.
cfm?id=3358217.

Loghi:2006:CCT

[LPB06] Mirko Loghi, Massimo Poncino, and Luca Benini. Cache coherence tradeoffs in shared-memory MPSoCs. ACM Transactions on Embedded Computing Systems, 5(2):383–407, May 2006. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Lumpp:2024:DFB

[LPBF24] Francesco Lumpp, Marco Panato, Nicola Bombieri, and Franco Fummi. A design flow based on Docker and Kubernetes for ROSbased robotic software applications. ACM Transactions on Embedded Computing Systems, 23(5):74:1– 74:??, September 2024. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3594539.

Lee:2007:LBB

[LPC^{+}07] Sang-Won Lee, Dong-Joo Park, Tae-Sun Chung, Dong-Ho Lee, Sangwon Park, and Ha-Joo Song. A log bufferbased flash translation layer using fully-associative sector translation. ACM Transactions on Embedded Computing Systems, 6(3):18:1– 18:??, July 2007. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Latifis:2020:RMC

[LPD^{+}20] Ioannis Latifis, Karthick Parashar, Grigoris Dimitroulakos, Hans Cappelle, Christakis Lezos, Konstantinos Masselos, and Francky Catthoor. A retargetable MATLAB-to-C compiler exploiting custom instructions and data parallelism. ACM Transactions on Embedded Computing Systems, 19(6): 50:1–50:27, November 2020. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3391898.

Loveless:2023:CML

[LPE^{+}23] Andrew Loveless, Linh Thi Xuan Phan, Lisa Erickson, Ronald Dreslinski, and Baris Kasikci. CrossTalk: Making low-latency fault tolerance cheap by exploiting redundant networks. ACM Transactions on Embedded Computing Systems, 22(5s): 147:1–147:??, October 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3609436.

Liu:2013:AEE

[LPFG13] Shaoshan Liu, Richard Neil Pittman, Alessandro Forin, and Jean-Luc Gaudiot. Achieving energy efficiency through runtime partial reconfiguration on reconfigurable systems. ACM Transactions on Embedded Computing Systems, 12(3):72:1– 72:??, March 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Le:2016:CBR

[LPFL16] Thi Thieu Hoa Le, Roberto Passerone, Uli Fahrenberg, and Axel Legay. Contractbased requirement modularization via synthesis of correct decompositions. ACM Transactions on Embedded Computing Systems, 15(2): 33:1–33:??, May 2016. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Liu:2017:HPI

[LPO^{+}17] Zhe Liu, Thomas Pöppelmann, Tobias Oder, Hwajeong Seo, Sujoy Sinha Roy, Tim Güneysu, Johann Großschädl, Howon Kim, and Ingrid Verbauwhede. High-performance ideal lattice-based cryptography on 8-bit AVR microcontrollers. ACM Transactions on Embedded Computing Systems, 16(4):117:1– 117:??, August 2017. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Leon:2021:IPD

[LPP^{+}21] Vasileios Leon, Theodora Paparouni, Evangelos Petrongonas, Dimitrios Soudris, and Kiamal Pekmestzi. Improving power of DSP and CNN hardware accelerators using approximate floating-point multipliers. ACM Transactions on Embedded Computing Systems, 20(5):39:1– 39:21, July 2021. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3448980.

Li:2013:TAT

[LQN^{+}13] Jiayin Li, Meikang Qiu, Jian-Wei Niu, Laurence T. Yang, Yongxin Zhu, and Zhong Ming. Thermal-aware task scheduling in 3D chip multiprocessor with real-time constrained workloads. ACM Transactions on Embedded Computing Systems, 12(2): 24:1–24:??, February 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Lu:2016:VCV

[LRZ16] Yaojie Lu, Seyedamin Rooholamin, and Sotirios G. Ziavras. Vector coprocessor virtualization for simultaneous multithreading. ACM Transactions on Embedded Computing Systems, 15(3): 57:1–57:??, July 2016. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Lloyd:2009:PSN

[LS09] Scott Lloyd and Quinn Snell.

A packet-switched network architecture for reconfigurable computing. ACM Transactions on Embedded Computing Systems, 9(1): 7:1–7:??, October 2009. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Lee:2012:PPI

[LS12] Jongeun Lee and Aviral Shrivastava. PICA: Processor Idle Cycle Aggregation for energy-efficient embedded systems. ACM Transactions on Embedded Computing Systems, 11(2):26:1– 26:??, July 2012. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Lee:2013:SBR

[LS13] Jongeun Lee and Aviral Shrivastava. Software-based register file vulnerability reduction for embedded processors. ACM Transactions on Embedded Computing Systems, 13(1s):38:1– 38:??, November 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Landy:2017:SAS

[LS17] Aaron Landy and Greg Stitt. Serial arithmetic strategies for improving FPGA throughput. ACM Transactions on Embedded Computing Systems, 16(3):84:1– 84:??, July 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Lang:2020:DIE

[LS20] Clemens Lang and Isabella Stilkerich. Design and implementation of an escape analysis in the context of safety-critical embedded systems. ACM Transactions on Embedded Computing Systems, 19(1):6:1–6:20, February 2020. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL

                                        https://dl.acm.org/doi/
                                        abs/10.1145/3372133.

Lam:2014:REC

[LSC14] Siew-Kei Lam, Thambipillai Srikanthan, and Christopher T. Clarke. Rapid evaluation of custom instruction selection approaches with FPGA estimation. ACM Transactions on Embedded Computing Systems, 13(4): 75:1–75:??, February 2014. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Lee:2019:TAS

[LSC19] Youngmoon Lee, Kang G.

Shin, and Hoon Sung Chwa. Thermal-aware scheduling for integrated CPUs–GPU platforms. ACM Transactions on Embedded Computing Systems, 18(5s):90:1– 90:??, October 2019. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3358235.

Lee:2008:DFR

[LSK^{+}08] Sheayun Lee, Insik Shin, Woonseok Kim, Insup Lee, and Sang Lyul Min. A design framework for real-time embedded systems with code size and energy constraints. ACM Transactions on Embedded Computing Systems, 7(2):18:1–18:??, February 2008. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Lizarraga:2020:AMB

[LSL20] Adrian Lizarraga, Jonathan Sprinkle, and Roman Lysecky. Automated modelbased optimization of dataadaptable embedded systems. ACM Transactions on Embedded Computing Systems, 19(1):8:1–8:22, February 2020. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL

https://dl.acm.org/doi/
abs/10.1145/3372142.

Li:2023:IIE

[LSL^{+}23] Wentong Li, Liang Shi, Hang Li, Changlong Li, and Edwin Hsing-Mean Sha. IOSR: Improving I/O efficiency for memory swapping on mobile devices via scheduling and reshaping. ACM Transactions on Embedded Computing Systems, 22(5s):129:1– 129:??, October 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3607923.

Li:2024:CIS

[LTQ^{+}24] Lu Li, Qi Tian, Guofeng Qin, Shuaiyu Chen, and Weijia Wang. Compact instruction set extensions for Dilithium. ACM Transactions on Embedded Computing Systems, 23(2):23:1– 23:??, March 2024. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3643826.

Lysecky:2009:DIM

[LV09] Roman Lysecky and Frank Vahid. Design and implementation of a MicroBlazebased warp processor. ACM Transactions on Embedded Computing Systems, 8(3): 22:1–22:??, April 2009. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Leon-Vega:2023:AGR

[LVSVRFCG23] Luis G. León-Vega, Eduardo Salazar-Villalobos, Alejandro Rodriguez-Figueroa, and Jorge Castro-Godínez. Automatic generation of resource and accuracy configurable processing elements. ACM Transactions on Embedded Computing Systems, 22(4):75:1–75:??, July 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3594540.

Lee:2013:HPL

[LWB13] Dongwon Lee, Marilyn Wolf, and Shuvra S. Bhattacharyya. High-performance and low-energy buffer mapping method for multiprocessor DSP systems. ACM Transactions on Embedded Computing Systems, 12(3): 82:1–82:??, March 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Lin:2018:MCV

[LWB18] Shuoxin Lin, Jiahao Wu, and Shuvra S. Bhattacharyya. Memory-constrained vectorization and scheduling of dataflow graphs for hybrid CPU–GPU platforms. ACM Transactions on Embedded Computing Systems, 17(2): 50:1–50:??, April 2018. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Lin:2025:IRI

[LWCL25] Gao-Yu Lin, Po-Yuan Wang, Shin-Ming Cheng, and Hahn-Ming Lee. Improving robustness in IoT malware detection through execution order analysis. ACM Transactions on Embedded Computing Systems, 24(1):4:1– 4:??, January 2025. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3689427.

Liu:2017:EEC

[LWHS17] Zhe Liu, Jian Weng, Zhi Hu, and Hwajeong Seo. Efficient elliptic curve cryptography for embedded devices. ACM Transactions on Embedded Computing Systems, 16(2):53:1–53:??, April 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Lin:2010:SSA

[LWK^{+}10] Chang Hong Lin, Marilyn Wolf, Xenefon Koutsoukos, Sandeep Neema, and Janos Sztipanovits. System and software architectures of distributed smart cameras. ACM Transactions on Embedded Computing Systems, 9(4):38:1–38:??, March 2010. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Liu:2017:BEB

[LWK^{+}17] Qingrui Liu, Xiaolong Wu, Larry Kittinger, Markus Levy, and Changhee Jung. BenchPrime: Effective building of a hybrid benchmark suite. ACM Transactions on Embedded Computing Systems, 16(5s):179:1–179:??, October 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Liang:2023:TRS

[LWS^{+}23] Tailin Liang, Lei Wang, Shaobo Shi, John Glossner, and Xiaotong Zhang. TCX: a RISC style tensor computing extension and a programmable tensor processor. ACM Transactions on Embedded Computing Systems, 22(3):47:1–47:??, May 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3568310.

Li:2016:USS

[LWZ^{+}16] Fang Li, Jiafu Wan, Ping Zhang, Di Li, Daqiang Zhang, and Keliang Zhou. Usage-specific semantic integration for cyber-physical robot systems. ACM Transactions on Embedded Computing Systems, 15(3):50:1– 50:??, July 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Li:2024:TEE

[LWZ^{+}24] Yueting Li, Xueyan Wang, He Zhang, Biao Pan, Keni Qiu, Wang Kang, Jun Wang, and Weisheng Zhao. Toward energy-efficient STT-MRAM-based near memory computing architecture for embedded systems. ACM Transactions on Embedded Computing Systems, 23(3): 37:1–37:??, May 2024. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3650729.

Lee:2012:IMR

[LX12] Jaehwan John Lee and Xiang Xiao. Instant multiunit resource hardware deadlock detection scheme for systemon-chips. ACM Transactions on Embedded Computing Systems, 11(3):67:1– 67:??, September 2012. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Liu:2016:ETA

[LX16] Weichen Liu and Chunhua Xiao. An efficient technique of application mapping and scheduling on realtime multiprocessor systems for throughput optimization. ACM Transactions on Embedded Computing Systems, 15(4):65:1–65:??, August 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Leite:2022:REC

[LX22] Clayton Frederick Souza Leite and Yu Xiao. Resourceefficient continual learning for sensor-based human activity recognition. ACM Transactions on Embedded Computing Systems, 21(6): 85:1–85:??, November 2022. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3530910.

Li:2010:SMA

[LXK10] Lian Li, Jingling Xue, and Jens Knoop. Scratchpad memory allocation for data aggregates via interval coloring in superperfect graphs. ACM Transactions on Embedded Computing Systems, 10(2):28:1–28:??, December 2010. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Liu:2013:JVP

[LXL13] Tiantian Liu, Chun Jason Xue, and Minming Li. Joint variable partitioning and bank selection instruction optimization for partitioned memory architectures. ACM Transactions on Embedded Computing Systems, 12(3): 76:1–76:??, March 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Li:2018:TES

[LYC^{+}18] Feng Li, Yanbing Yang, Zicheng Chi, Liya Zhao, Yaowen Yang, and Jun Luo. Trinity: Enabling selfsustaining WSNs indoors with energy-free sensing and networking. ACM Transactions on Embedded Computing Systems, 17(2):57:1– 57:??, April 2018. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Lin:2015:SLP

[LYH^{+}15] Ye-Jyun Lin, Chia-Lin Yang, Jiao-We Huang, Tay-Jyi Lin, Chih-Wen Hsueh, and Naehyuck Chang. System-level performance and power optimization for MPSoC: a memory access-aware approach. ACM Transactions on Embedded Computing Systems, 14(1):8:1– 8:??, January 2015. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Li:2013:SDM

[LYL13] Mo Li, Zheng Yang, and Yunhao Liu. Sea depth measurement with restricted floating sensors. ACM Transactions on Embedded Computing Systems, 13(1): 1:1–1:??, August 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Liu:2017:PSSb

[LYY^{+}17] Bo Liu, Xiao-Tong Yuan, Yang Yu, Qingshan Liu, and Dimitris N. Metaxas. Parallel sparse subspace clustering via joint sample and parameter blockwise partition. ACM Transactions on Embedded Computing Systems, 16(3):75:1–75:??, July 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Liu:2017:PNM

[LZJ17] Guanjun Liu, Mengchu Zhou, and Changjun Jiang. Petri net models and collaborativeness for parallel processes with resource sharing and message passing. ACM Transactions on Embedded Computing Systems, 16(4):113:1–113:??, August 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Luo:2019:PFC

[LZJ^{+}19] Zhengxiong Luo, Feilong Zuo, Yu Jiang, Jian Gao, Xun Jiao, and Jiaguang Sun. Polar: Function code aware fuzz testing of ICS protocol. ACM Transactions on Embedded Computing Systems, 18(5s):93:1–93:??, October 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/ft_gateway.
cfm?id=3358227.

Liu:2020:AML

[LZJ^{+}20] Ke Liu, Mengying Zhao, Lei Ju, Zhiping Jia, Jingtong Hu, and Chun Jason Xue. Applying multiple level cell to non-volatile FP-GAs. ACM Transactions on Embedded Computing Systems, 19(4):27:1–27:22, July 2020. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3400885.

Li:2015:CIB

[LZL15] Guohui Li, Yi Zhang, and Jianjun Li. Crenel-interval-based dynamic power management for periodic realtime systems. ACM Transactions on Embedded Computing Systems, 14(4):74:1– 74:??, December 2015. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Liang:2023:SIA

[LZNL23] Yun (Eric) Liang, Wei Zhang, Stephen Neuendorffer, and Wayne Luk. Special issue: “AI Acceleration on FPGAs”. ACM Transactions on Embedded Computing Systems, 22(6):89:1– 89:??, November 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3626323.

Lin:2018:OOP

[LZS^{+}18] Chi Lin, Yanhong Zhou, Houbing Song, Chang Wu Yu, and Guowei Wu. OPPC: an optimal path planning charging scheme based on schedulability evaluation for WRSNs. ACM Transactions on Embedded Computing Systems, 17(1):7:1– 7:??, January 2018. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Li:2020:DIA

[LZS20] Xinyi Li, Lei Zhang, and Xipeng Shen. DIAC: an inter-app conflicts detector for open IoT systems. ACM Transactions on Embedded Computing Systems, 19(6):46:1–46:25, November 2020. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3391895.

Liu:2019:XBL

[LZZ^{+}19] Weiqiang Liu, Lei Zhang, Zhengran Zhang, Chongyan Gu, Chenghua Wang, Maire O’neill, and Fabrizio Lombardi. XOR-based low-cost reconfigurable PUFs for IoT security. ACM Transactions on Embedded Computing Systems, 18(3):25:1– 25:??, June 2019. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

tronic). URL https://dl.

                                        acm.org/ft_gateway.cfm?
                                        id=3274666.

Luo:2023:SEF

[LZZ^{+}23] Yuling Luo, Shiqi Zhang, Shunsheng Zhang, Junxiu Liu, Yanhu Wang, and Su Yang. A secure and efficient framework for outsourcing large-scale matrix determinant and linear equations. ACM Transactions on Embedded Computing Systems, 22(5):86:1–86:??, September 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3611014.

Lam:2025:MMR

[LZZK25] Kam-Yiu Lam, Xiaofei Zhao, Chunjiang Zhu, and Tei-Wei Kuo. MVLevelDB+: Meeting relative consistency requirements of temporal queries in sensor stream databases. ACM Transactions on Embedded Computing Systems, 24(1):7:1– 7:??, January 2025. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3694787.

Mirzoyan:2014:PVA

[MAG14] Davit Mirzoyan, Benny Akesson, and Kees Goossens. Process-variation-aware mapping of best-effort and realtime streaming applications to MPSoCs. ACM Transactions on Embedded Computing Systems, 13(2s):61:1– 61:??, January 2014. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Munir:2015:MAF

[MAGR15] Arslan Munir, Joseph Antoon, and Ann Gordon-Ross. Modeling and analysis of fault detection and fault tolerance in wireless sensor networks. ACM Transactions on Embedded Computing Systems, 14(1):3:1– 3:??, January 2015. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Mandal:2019:APM

[MAKO19] Sumit K. Mandal, Raid Ayoub, Michael Kishinevsky, and Umit Y. Ogras. Analytical performance models for NoCs with multiple priority traffic classes. ACM Transactions on Embedded Computing Systems, 18(5s):52:1–52:??, October 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/ft_gateway.
cfm?id=3358176.

Mejia-Alvarez:2004:ASS

[MALM04] Pedro Mejia-Alvarez, Eugene Levner, and Daniel Mossé. Adaptive scheduling server for power-aware real-time tasks. ACM Transactions on Embedded Computing Systems, 3(2):284– 306, May 2004. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Mamish:2024:NSP

[MAS^{+}24] John Mamish, Rawan Alharbi, Sougata Sen, Shashank Holla, Panchami Kamath, Yaman Sangar, Nabil Alshurafa, and Josiah Hester. NIR-sighted: a programmable streaming architecture for low-energy human-centric vision applications. ACM Transactions on Embedded Computing Systems, 23(6):101:1– 101:??, November 2024. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3672076.

Mirzoyan:2015:MNG

[MASG15] Davit Mirzoyan, Benny Akesson, Sander Stuijk, and Kees Goossens. Maximizing the number of good dies for streaming applications in NoC-based0 MPSoCs under process variation. ACM Transactions on Embedded Computing Systems, 14(4): 83:1–83:??, December 2015. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Missimer:2022:TRT

[MAW22] Katherine Missimer, Manos Athanassoulis, and Richard West. Telomere: Realtime NAND flash storage. ACM Transactions on Embedded Computing Systems, 21(1):10:1–10:24, January 2022. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3479157.

McLoughlin:2010:RTR

[MB10] Ian Vince McLoughlin and Timo Rolf Bretschneider. Reliability through redundant parallelism for microsatellite computing. ACM Transactions on Embedded Computing Systems, 9(3): 26:1–26:??, February 2010. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Matthews:2015:PTS

[MBB^{+}15] Adam Matthews, Stanislav Bobovych, Nilanjan Banerjee, James P. Parkerson, Ryan Robucci, and Chintan Patel. Perpetuu: a tiered solar-powered GIS microserver. ACM Transactions on Embedded Computing Systems, 14(4):78:1– 78:??, December 2015. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Mendez:2022:EIC

[MBCM22] Javier Mendez, Kay Bierzynski, M. P. Cuéllar, and Diego P. Morales. Edge intelligence: Concepts, architectures, applications, and future directions. ACM Transactions on Embedded Computing Systems, 21(5):48:1– 48:??, September 2022. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3486674.

Mangeruca:2007:USU

[MBFSV07] Leonardo Mangeruca, Massimo Baleani, Alberto Ferrari, and Alberto Sangiovanni-Vincentelli. Uniprocessor scheduling under precedence constraints for embedded systems design. ACM Transactions on Embedded Computing Systems, 7(1):6:1– 6:30, December 2007. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Murray:2009:CTI

[MBFT09] Alastair C. Murray, Richard V. Bennett, Björn Franke, and Nigel Topham. Code transformation and instruction set extension. ACM Transactions on Embedded Computing Systems, 8(4):26:1– 26:??, July 2009. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Modi:2023:CRR

[MBJ^{+}23] Garima Modi, Aritra Bagchi, Neetu Jindal, Ayan Mandal, and Preeti Ranjan Panda. CABARRE: Request response arbitration for shared cache management. ACM Transactions on Embedded Computing Systems, 22(5s): 130:1–130:??, October 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3608096.

Medhat:2015:RMC

[MBKF15] Ramy Medhat, Borzoo Bonakdarpour, Deepak Kumar, and Sebastian Fischmeister. Runtime monitoring of cyber-physical systems under timing and memory constraints. ACM Transactions on Embedded Computing Systems, 14(4):79:1– 79:??, December 2015. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Murillo:2016:MSD

[MBLA16] Luis Gabriel Murillo, Ròbert Lajos Bücs, Rainer Leupers, and Gerd Ascheid. MPSoC software debugging on virtual platforms via execution control with event graphs. ACM Transactions on Embedded Computing Systems, 16(1):7:1–7:??, November 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Mancuso:2014:OPA

[MBP14] Giulio M. Mancuso, Enrico Bini, and Gabriele Pannocchia. Optimal priority assignment to control tasks. ACM Transactions on Embedded Computing Systems, 13(5s):161:1–161:??, September 2014. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Massolino:2015:OSC

[MBR15] Pedro Maat C. Massolino, Paulo S. L. M. Barreto, and Wilson V. Ruggiero. Optimized and scalable coprocessor for McEliece with binary Goppa codes. ACM Transactions on Embedded Computing Systems, 14(3): 45:1–45:??, April 2015. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Mishra:2022:SCF

[MCG22] Tanmaya Mishra, Thidapat Chantem, and Ryan Gerdes. Survey of controlflow integrity techniques for real-time embedded systems. ACM Transactions on Embedded Computing Systems, 21(4):41:1–41:??, July 2022. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3538275.

McInnes:2013:MAT

[McI13] Allan I. McInnes. Modeling and analysis of TinyOS sensor node firmware: a CSP approach. ACM Transactions on Embedded Computing Systems, 12(1):5:1– 5:??, January 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Mera:2017:ATP

[MCM^{+}17] Maria Isabel Mera, Jonah Caplan, Seyyed Hasan Mozafari, Brett H. Meyer, and Peter Milder. Area, throughput, and power trade-offs for FPGA- and ASIC-based execution stream compression. ACM Transactions on Embedded Computing Systems, 16(4):96:1–96:??, August 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Mathew:2017:GES

[MCP17] Jimson Mathew, Rajat Subhra Chakraborty, and Dhiraj K. Pradhan. Guest editorial: Special issue on “Secure and Fault-Tolerant Embedded Computing”. ACM Transactions on Embedded Computing Systems, 16(4):92:1– 92:??, August 2017. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Mathew:2015:NMB

[MCS^{+}15] Jimson Mathew, Rajat Subhra Chakraborty, Durga Prasad Sahoo, Yuanfan Yang, and Dhiraj K. Pradhan. A novel memristor-based hardware security primitive. ACM Transactions on Embedded Computing Systems, 14(3): 60:1–60:??, April 2015. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Mark:2012:HBC

[MCSW12] Cindy Mark, Scott Y. L. Chin, Lesley Shannon, and Steven J. E. Wilton. Hierarchical benchmark circuit generation for FPGA architecture evaluation. ACM Transactions on Embedded Computing Systems, 11(S2): 42:1–42:??, 2012. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Mishra:2004:MVP

[MD04] Prabhat Mishra and Nikil Dutt. Modeling and validation of pipeline specifications. ACM Transactions on Embedded Computing Systems, 3(1):114– 139, February 2004. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Maity:2021:SSO

[MDS^{+}21] Biswadip Maity, Bryan Donyanavard, Anmol Surhonne, Amir Rahmani, Andreas Herkersdorf, and Nikil Dutt. SEAMS: Self-optimizing runtime manager for approximate memory hierarchies. ACM Transactions on Embedded Computing Systems, 20(5):48:1–48:26, July 2021. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3466875.

Ma:2023:ICD

[MDWL23] Zhengzheng Ma, Tuo Dai, Xuechao Wei, and Guojie Luo. An intermediatecentric dataflow for transposed convolution acceleration on FPGA. ACM Transactions on Embedded Computing Systems, 22(6):92:1– 92:??, November 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3561053.

Morris:2022:HUH

[MEK^{+}22] Justin Morris, Kazim Ergun, Behnam Khaleghi, Mohen Imani, Baris Aksanli, and Tajana Simunic. Hy-DREA: Utilizing hyperdimensional computing for a more robust and efficient machine learning system. ACM Transactions on Embedded Computing Systems, 21(6):78:1–78:??, November 2022. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3524067.

Manolache:2004:SAA

[MEP04] Sorin Manolache, Petru Eles, and Zebo Peng. Schedulability analysis of applications with stochastic task execution times. ACM Transactions on Embedded Computing Systems, 3(4):706– 735, November 2004. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Manolache:2008:TMP

[MEP08] Sorin Manolache, Petru Eles, and Zebo Peng. Task mapping and priority assignment for soft real-time applications under deadline miss ratio constraints. ACM Transactions on Embedded Computing Systems, 7(2): 19:1–19:??, February 2008. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Murray:2012:ASL

[MF12] Alastair Murray and Björn Franke. Adaptive sourcelevel data assignment to dual memory banks. ACM Transactions on Embedded Computing Systems, 11S(1): 20:1–20:??, 2012. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

MartinezSantos:2013:LSA

[MF13] Juan Carlos Martinez Santos and Yunsi Fei. Leveraging speculative architectures for runtime program validation. ACM Transactions on Embedded Computing Systems, 13(1):3:1–3:??, August 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Mandal:2016:DIW

[MFG16] Kalikinkar Mandal, Xinxin Fan, and Guang Gong. Design and implementation of Warbler family of lightweight pseudorandom number generators for smart devices. ACM Transactions on Embedded Computing Systems, 15(1):1:1– 1:??, February 2016. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Motamedi:2017:MIR

[MFG17] Mohammad Motamedi, Daniel Fong, and Soheil Ghiasi. Machine intelligence on resource-constrained IoT devices: The case of thread granularity optimization for CNN inference. ACM Transactions on Embedded Computing Systems, 16(5s): 151:1–151:??, October 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Manilov:2017:FRS

[MFMA17] Stanislav Manilov, Björn Franke, Anthony Magrath, and Cedric Andrieu. Free Rider: a source-level transformation tool for retargeting platform-specific intrinsic functions. ACM Transactions on Embedded Computing Systems, 16(2):38:1– 38:??, April 2017. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Muresan:2005:ICM

[MG05] Radu Muresan and Catherine Gebotys. Instantaneous current modeling in a complex VLIW processor core. ACM Transactions on Embedded Computing Systems, 4(2):415–451, May 2005. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Malik:2015:HRT

[MG15] Avinash Malik and David Gregg. Heuristics on reachability trees for bicriteria scheduling of stream graphs on heterogeneous multiprocessor architectures. ACM Transactions on Embedded Computing Systems, 14(2): 23:1–23:??, March 2015. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Marshall:2021:PCP

[MGB^{+}21] James Marshall, Robert Gifford, Gedare Bloom, Gabriel Parmer, and Rahul Simha. Precise cache profiling for studying radiation effects. ACM Transactions on Embedded Computing Systems, 20(3):25:1–25:25, April 2021. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3442339.

Mondal:2023:PPR

[MGC^{+}23] Anupam Mondal, Shreya Gangopadhyay, Durba Chatterjee, Harishma Boyapally, and Debdeep Mukhopadhyay. PReFeR: Physically related function based remote attestation protocol. ACM Transactions on Embedded Computing Systems, 22(5s):109:1–109:??, October 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3609104.

Mohanty:2019:SPE

[MGLP19] Ram Prasad Mohanty, Hasindu Gamaarachchi, Andrew Lambert, and Sri Parameswaran. SWARAM: Portable energy and cost efficient embedded system for genomic processing. ACM Transactions on Embedded Computing Systems, 18(5s):61:1– 61:??, October 2019. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

                                        tronic). URL https://dl.
                                        acm.org/ft_gateway.cfm?
                                        id=3358211.

Mohammad:2025:CAE

[MGP25] Rafiuzzaman Mohammad, Sathish Gopalakrishnan, and Karthik Pattabiraman. Coapproximator: Enabling performance prediction in colocated applications. ACM Transactions on Embedded Computing Systems, 24(1): 6:1–6:??, January 2025. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3677180.

Mendis:2019:ADU

[MH19] Hashan Roshantha Mendis and Pi-Cheng Hsiu. Accumulative display updating for intermittent systems. ACM Transactions on Embedded Computing Systems, 18(5s):72:1–72:??, October 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

                                        /dl.acm.org/ft_gateway.
                                        cfm?id=3358190.

Mack:2023:CCI

[MHK^{+}23] Joshua Mack, Sahil Hassan, Nirmal Kumbhare, Miguel Castro Gonzalez, and Ali Akoglu. CEDR: a compiler-integrated, extensible DSSoC runtime. ACM Transactions on Embedded Computing Systems, 22(2): 36:1–36:??, March 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3529257.

Mhamdi:2013:FMT

[MHT13] Tarek Mhamdi, Osman Hasan, and Sofiène Tahar. Formalization of measure theory and Lebesgue integration for probabilistic analysis in HOL. ACM Transactions on Embedded Computing Systems, 12(1):13:1– 13:??, January 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Mitra:2021:ERA

[Mit21]  Tulika Mitra.  Editorial:

Reimagining ACM Transactions on Embedded Computing Systems (TECS). ACM Transactions on Embedded Computing Systems, 20(3): 18e:1–18e:3, April 2021. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3450438.

Mozaffari-Kermani:2017:FDA

[MKAA17] Mehran Mozaffari-Kermani, Reza Azarderakhsh, and Anita Aghaie. Fault detection architectures for post-quantum cryptographic stateless hash-based secure signatures benchmarked on ASIC. ACM Transactions on Embedded Computing Systems, 16(2):59:1–59:??, April 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Mozaffari-Kermani:2018:ERE

[MKASJ18] Mehran Mozaffari-Kermani, Reza Azarderakhsh, Ausmita Sarker, and Amir Jalali. Efficient and reliable error detection architectures of hash-counter-hash tweakable enciphering schemes. ACM Transactions on Embedded Computing Systems, 17(2): 54:1–54:??, April 2018. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Mohaqeqi:2013:ASR

[MKD13] Morteza Mohaqeqi, Mehdi Kargahi, and Maryam Dehghan. Adaptive scheduling of real-time systems cosupplied by renewable and nonrenewable energy sources. ACM Transactions on Embedded Computing Systems, 13(1s):36:1–36:??, November 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Millo:2015:MAD

[MKD15] Jean-Vivien Millo, Emilien Kofman, and Robert De Simone. Modeling and analyzing dataflow applications on NoC-based many-core architectures. ACM Transactions on Embedded Computing Systems, 14(3):46:1– 46:??, April 2015. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Morse:2018:LAW

[MKE18] Jeremy Morse, Steve Kerrison, and Kerstin Eder. On the limitations of analyzing worst-case dynamic energy of processing. ACM Transactions on Embedded Computing Systems, 17(3): 59:1–59:??, June 2018. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Min:2023:MTK

[MKM^{+}23a] Donghyun Min, Kihyun Kim, Chaewon Moon, Awais Khan, Seungjin Lee, Changhwan Yun, Woosuk Chung, and Youngjae Kim. A multitenant key-value SSD with secondary index for search query processing and analysis. ACM Transactions on Embedded Computing Systems, 22(4):65:1–65:??, July 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3590153.

Mondal:2023:HTD

[MKM^{+}23b] Anindan Mondal, Shubrojyoti Karmakar, Mahabub Hasan Mahalat, Suchismita Roy, Bibhash Sen, and Anupam Chattopadhyay. Hardware Trojan detection using transition probability with minimal test vectors. ACM Transactions on Embedded Computing Systems, 22(1): 11:1–11:??, January 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3545000.

Maier:2018:FIT

[MKMGS18] Petra R. Maier, Veit B. Kleeberger, Daniel Mueller-Gritschneder, and Ulf Schlichtmann. Fault injection for test-driven development of robust SoC firmware. ACM Transactions on Embedded Computing Systems, 17(1): 19:1–19:??, January 2018. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Mack:2025:TNR

[MKOA25] Joshua Mack, Anish Krishnakumar, Umit Ogras, and Ali Akoglu. Tutorial: a novel runtime environment for accelerator-rich heterogeneous architectures. ACM Transactions on Embedded Computing Systems, 24(1):20:1–20:??, January 2025. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3687463.

Misailovic:2013:PSP

[MKR13] Sasa Misailovic, Deokhwan Kim, and Martin Rinard. Parallelizing sequential programs with statistical accuracy tests. ACM Transactions on Embedded Computing Systems, 12(2s):88:1– 88:??, May 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Mehrabian:2017:TTL

[MKS^{+}17] Mohammadreza Mehrabian, Mohammad Khayatian, Aviral Shrivastava, John C. Eidson, Patricia Derler, Hugo A. Andrade, Ya-Shian Li-Baboud, Edward Griffor, Marc Weiss, and Kevin Stanton. Timestamp temporal logic (TTL) for testing the timing of cyber-physical systems. ACM Transactions on Embedded Computing Systems, 16(5s):169:1– 169:??, October 2017. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Metz:2023:BBS

[MKS23] David Metz, Vineet Kumar, and Magnus Själander. BISDU: a bit-serial dotproduct unit for microcontrollers. ACM Transactions on Embedded Computing Systems, 22(5):79:1– 79:??, September 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3608447.

Mousavi:2023:DDA

[MLAD23] Hamid Mousavi, Mohammad Loni, Mina Alibeigi, and Masoud Daneshtalab. DASS: Differentiable architecture search for sparse neural networks. ACM Transactions on Embedded Computing Systems, 22(5s):105:1– 105:??, October 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3609385.

Mitra:2008:VAD

[MLL08] Sayan Mitra, Daniel Liberzon, and Nancy Lynch. Verifying average dwell time of hybrid systems. ACM Transactions on Embedded Computing Systems, 8(1): 3:1–3:??, December 2008. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Ma:2017:DPE

[MLL^{+}17] Kaisheng Ma, Xueqing Li, Huichu Liu, Xiao Sheng, Yiqun Wang, Karthik Swaminathan, Yongpan Liu, Yuan Xie, John Sampson, and Vijaykrishnan Narayanan. Dynamic power and energy management for energy harvesting nonvolatile processor systems. ACM Transactions on Embedded Computing Systems, 16(4):107:1– 107:??, August 2017. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Medhat:2017:MPE

[MLR^{+}17] Ramy Medhat, Michael O.

Lam, Barry L. Rountree, Borzoo Bonakdarpour, and Sebastian Fischmeister. Managing the performance/ error tradeoff of floatingpoint intensive applications. ACM Transactions on Embedded Computing Systems, 16(5s):184:1–184:??, October 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Mozumdar:2009:CSP

[MLV09] Mohammad Mostafizur Rahman Mozumdar, Luciano Lavagno, and Laura Vanzago. A comparison of software platforms for wireless sensor networks: MANTIS, TinyOS, and ZigBee. ACM Transactions on Embedded Computing Systems, 8(2): 17:1–17:??, January 2009. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Majmudar:2016:AOR

[MM16] Charvi A. Majmudar and Bashir I. Morshed. Autonomous OA removal in real-time from single channel EEG data on a wearable device using a hybrid algebraic-wavelet algorithm. ACM Transactions on Embedded Computing Systems, 16(1):20:1–20:??, November 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Min:2023:SBM

[MMA^{+}23] Chulhong Min, Akhil Mathur, Utku Günay Acer, Alessandro Montanari, and Fahim Kawsar. SensiX++: Bringing MLOps and multi-tenant model serving to sensory edge devices. ACM Transactions on Embedded Computing Systems, 22(6):98:1– 98:??, November 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3617507.

Mishra:2004:PMC

[MMD04] Prabhat Mishra, Mahesh Mamidipaka, and Nikil Dutt. Processor-memory coexploration using an architecture description language. ACM Transactions on Embedded Computing Systems, 3(1): 140–162, February 2004. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Mukherjee:2022:AFD

[MMD22] Arijit Mukherjee, Jayeeta Mondal, and Swarnava Dey. Accelerated fire detection and localization at edge. ACM Transactions on Embedded Computing Systems, 21(6):70:1–70:??, November 2022. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3510027.

Mishra:2023:VVA

[MMH^{+}23] Vishesh Mishra, Sparsh Mittal, Neelofar Hassan, Rekha Singhal, and Urbi Chatterjee. VADF: Versatile approximate data formats for energy-efficient computing. ACM Transactions on Embedded Computing Systems, 22(5s):111:1–111:??, October 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3609106.

Manohar:2022:CUC

[MMK22] Sheel Sindhu Manohar, Sparsh Mittal, and Hemangee K. Kapoor. CORI-DOR: Using COherence and TempoRal LocalIty to mitigate read disurbance ErrOR in STT–RAM caches. ACM Transactions on Embedded Computing Systems, 21(1): 2:1–2:24, January 2022. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3484493.

Mohan:2010:PTA

[MMR^{+}10] Sibin Mohan, Frank Mueller, Michael Root, William Hawkins, Christopher Healy, David Whalley, and Emilio Vivancos. Parametric timing analysis and its application to dynamic voltage scaling. ACM Transactions on Embedded Computing Systems, 10(2):25:1–25:??, December 2010. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Memik:2006:ENP

[MMS06] Gokhan Memik and William H. Mangione-Smith. Evaluating Network Processors using NetBench. ACM Transactions on Embedded Computing Systems, 5(2):453– 471, May 2006. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Mahdavikhah:2014:MFP

[MMSN14] Behzad Mahdavikhah, Ramin Mafi, Shahin Sirouspour, and Nicola Nicolici. A multiple-FPGA parallel computing architecture for realtime simulation of softobject deformation. ACM Transactions on Embedded Computing Systems, 13(4): 81:1–81:??, February 2014. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Moazzemi:2019:HFL

[MMY^{+}19] Kasra Moazzemi, Biswadip Maity, Saehanseul Yi, Amir M. Rahmani, and Nikil Dutt. HESSLE–FREE: Heterogeneous systems leveraging fuzzy control for runtime resource management. ACM Transactions on Embedded Computing Systems, 18(5s): 74:1–74:??, October 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/ft_gateway.
cfm?id=3358203.

Mosbahi:2013:CFM

[Mos13] Olfa Mosbahi. Combining formal methods for the development of reactive systems. ACM Transactions on Embedded Computing Systems, 12(1):16:1–16:??, January 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Motamedi:2019:DNA

[MPFG19] Mohammad Motamedi, Felix A. Portillo, Daniel Fong, and Soheil Ghiasi. Distill-Net: Application-specific distillation of deep convolutional neural networks for resource-constrained IoT platforms. ACM Transactions on Embedded Computing Systems, 18(5):44:1– 44:??, October 2019. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

tronic). URL https://dl.

                                        acm.org/ft_gateway.cfm?
                                        id=3360512.

Marinelli:2022:MES

[MPT^{+}22] Tommaso Marinelli, José Ignacio Gómez Pérez, Christian Tenllado, Manu Komalan, Mohit Gupta, and Francky Catthoor. Microarchitectural exploration of STT–MRAM last-level cache parameters for energyefficient devices. ACM Transactions on Embedded Computing Systems, 21(1): 3:1–3:20, January 2022. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3490391.

Mariani:2013:DSE

[MPZS13] Giovanni Mariani, Gianluca Palermo, Vittorio Zaccaria, and Cristina Silvano. Design-space exploration and runtime resource management for multicores. ACM Transactions on Embedded Computing Systems, 13(2):20:1–20:??, September 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Malik:2017:MCH

[MRA^{+}17] Avinash Malik, Partha S. Roop, Sidharta Andalam, Mark Trew, and Michael Mendler. Modular compilation of hybrid systems for emulation and large scale simulation. ACM Transactions on Embedded Computing Systems, 16(5s):118:1– 118:??, October 2017. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Majumdar:2013:TRO

[MRT13] Rupak Majumdar, Elaine Render, and Paulo Tabuada. A theory of robust omegaregular software synthesis. ACM Transactions on Embedded Computing Systems, 13(3):48:1–48:??, December 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Manzanares:2010:CER

[MRY^{+}10] Adam Manzanares, Xiaojun Ruan, Shu Yin, Xiao Qin, Adam Roth, and Mais Najim. Conserving energy in real-time storage systems with I/O burstiness. ACM Transactions on Embedded Computing Systems, 9(3): 20:1–20:??, February 2010. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Mueller:2005:ISI

[MS05] Frank Mueller and Per Stenström. Introduction to the special issue. ACM Transactions on Embedded Computing Systems, 4(1):1–2, February 2005. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Marinescu:2013:FSJ

[MS13a] Maria-Cristina Marinescu and César Sánchez. Fusing statecharts and Java. ACM Transactions on Embedded Computing Systems, 12(1s): 45:1–45:??, March 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Mery:2013:FSM

[MS13b] Dominique Méry and Neeraj Kumar Singh. Formal specification of medical systems by proof-based refinement. ACM Transactions on Embedded Computing Systems, 12(1):15:1–15:??, January 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Ma:2021:CSA

[MS21] Mingze Ma and Rizos Sakellariou. Code-size-aware scheduling of synchronous dataflow graphs on multicore systems. ACM Transactions on Embedded Computing Systems, 20(3):20:1– 20:24, April 2021. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3440034.

Monniaux:2023:FVL

[MS23] David Monniaux and Cyril Six. Formally verified loop-invariant code motion and assorted optimizations. ACM Transactions on Embedded Computing Systems, 22(1):3:1–3:??, January 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL

https://dl.acm.org/doi/
10.1145/3529507.

Middha:2008:MMS

[MSB08] Bhuvan Middha, Matthew Simpson, and Rajeev Barua. MTSS: Multitask stack sharing for embedded systems. ACM Transactions on Embedded Computing Systems, 7(4):46:1–46:??, July 2008. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Malik:2012:SLA

[MSCJ12] Avinash Malik, Zoran Salcic, Christopher Chong, and Salman Javed. System-level approach to the design of a smart distributed surveillance system using SystemJ. ACM Transactions on Embedded Computing Systems, 11(4):77:1–77:??, December 2012. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Manna:2016:ITS

[MSCS16] Kanchan Manna, Shivam Swami, Santanu Chattopadhyay, and Indranil Sengupta. Integrated through-silicon via placement and application mapping for 3D meshbased NoC design. ACM Transactions on Embedded Computing Systems, 16(1): 24:1–24:??, November 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Micolet:2017:SDP

[MSD17] Paul-Jules Micolet, Aaron Smith, and Christophe Dubach. A study of dynamic phase adaptation using a dynamic multicore processor. ACM Transactions on Embedded Computing Systems, 16(5s):121:1–121:??, October 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Moussalli:2014:SPX

[MSH^{+}14] Roger Moussalli, Mariam Salloum, Robert Halstead, Walid Najjar, and Vassilis J. Tsotras. A study on parallelizing XML path filtering using accelerators. ACM Transactions on Embedded Computing Systems, 13(4):93:1–93:??, February 2014. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Ma:2019:RFD

[MSHS19] Chenlin Ma, Zhaoyan Shen, Lei Han, and Zili Shao. RMW-F: a design of RMW-Free cache using built-in NAND-Flash for SMR storage. ACM Transactions on Embedded Computing Systems, 18(5s):65:1–65:??, October 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3358210.

Mu:2013:POS

[MSL13] Jingqing Mu, Karthik Shankar, and Roman Lysecky. Profiling and online systemlevel performance and power estimation for dynamically adaptable embedded systems. ACM Transactions on Embedded Computing Systems, 12(3):85:1–85:??, March 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Modekurthy:2021:DRT

[MSM21] Venkata P. Modekurthy, Abusayeed Saifullah, and Sanjay Madria. A distributed real-time scheduling system for industrial wireless networks. ACM Transactions on Embedded Computing Systems, 20(5): 46:1–46:28, July 2021. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3464429.

McIntire:2012:EES

[MSR^{+}12] Dustin McIntire, Thanos Stathopoulos, Sasank Reddy, Thomas Schmidt, and William J.

Kaiser. Energy-efficient sensing with the Low Power, Energy Aware Processing (LEAP) architecture. ACM Transactions on Embedded Computing Systems, 11(2): 27:1–27:??, July 2012. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Migliore:2017:HSA

[MSR^{+}17] Vincent Migliore, Cédric Seguin, Maria Méndez Real, Vianney Lapotre, Arnaud Tisserand, Caroline Fontaine, Guy Gogniat, and Russell Tessier. A high-speed accelerator for homomorphic encryption using the Karatsuba algorithm. ACM Transactions on Embedded Computing Systems, 16(5s): 138:1–138:??, October 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Mazumder:2024:RTR

[MSRM24] Arnab Neelim Mazumder, Farshad Safavi, Maryam Rahnemoonfar, and Tinoosh Mohsenin. Reg-Tune: a regression-focused finetuning approach for profiling low energy consumption and latency. ACM Transactions on Embedded Computing Systems, 23(3): 44:1–44:??, May 2024. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3623380.

Martin:2003:CSS

[MSS^{+}03] Thomas L. Martin, Daniel P.

Siewiorek, Asim Smailagic, Matthew Bosworth, Matthew Ettus, and Jolin Warren. A case study of a systemlevel approach to poweraware computing. ACM Transactions on Embedded Computing Systems, 2(3): 255–276, August 2003. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Majumdar:2023:NAB

[MSS23] Rupak Majumdar, Mahmoud Salamati, and Sadegh Soudjani. Neural abstractionbased controller synthesis and deployment. ACM Transactions on Embedded Computing Systems, 22(5s): 141:1–141:??, October 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3608104.

Minakova:2022:SBR

[MSSP22] Svetlana Minakova, Dolly Sapra, Todor Stefanov, and Andy D. Pimentel. Scenario based run-time switching for adaptive CNN-based applications at the edge. ACM Transactions on Embedded Computing Systems, 21(2): 14:1–14:33, March 2022. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3488718.

Malek:2016:RRQ

[MST^{+}16] Alirad Malek, Ioannis Sourdis, Stavros Tzilis, Yifan He, and Gerard Rauwerda. RQNoC: a resilient quality-of-service networkon-chip with service redirection. ACM Transactions on Embedded Computing Systems, 15(2):28:1–28:??, May 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Maggio:2014:TSC

[MTL14] Martina Maggio, Federico Terraneo, and Alberto Leva. Task scheduling: a controltheoretical viewpoint for a general and flexible solution. ACM Transactions on Embedded Computing Systems, 13(4):76:1–76:??, February 2014. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Marco:2020:ODL

[MTWE20] Vicent Sanz Marco, Ben Taylor, Zheng Wang, and Yehia Elkhatib. Optimizing deep learning inference on embedded systems through adaptive model selection. ACM Transactions on Embedded Computing Systems, 19(1):2:1– 2:28, February 2020. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3371154.

Musoll:2003:SRU

[Mus03]  Enric Musoll.  Speculat-

ing to reduce unnecessary power consumption. ACM Transactions on Embedded Computing Systems, 2(4): 509–536, November 2003. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Musoll:2010:CEL

[Mus10]  Enric Musoll.  A cost-

effective load-balancing policy for tile-based, massive multi-core packet processors. ACM Transactions on Embedded Computing Systems, 9(3):24:1–24:??, February 2010. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Marz:2016:RPC

[MV16] Stephen Marz and Brad Vander Zanden. Reducing power consumption and latency in mobile devices using an event stream model. ACM Transactions on Embedded Computing Systems, 16(1):11:1–11:??, November 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Masse:2013:MWE

[MVS^{+}13] Fabien Massé, Martien Van Bussel, Aline Serteyn, Johan Arends, and Julien Penders. Miniaturized wireless ECG monitor for realtime detection of epileptic seizures. ACM Transactions on Embedded Computing Systems, 12(4):102:1– 102:??, June 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Meyer:2016:SSC

[MWF^{+}16] Rolf Meyer, Jan Wagner, Bastian Farkas, Sven Horsinka, Patrick Siegl, Rainer Buchty, and Mladen Berekovic. A scriptable standard-compliant reporting and logging framework for SystemC. ACM Transactions on Embedded Computing Systems, 16(1):6:1– 6:??, November 2016. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Mujdei:2024:SCA

[MWK^{+}24] Catinca Mujdei, Lennert Wouters, Angshuman Karmakar, Arthur Beckers, Jose Maria Bermudo Mera, and Ingrid Verbauwhede. Sidechannel analysis of latticebased post-quantum cryptography: Exploiting polynomial multiplication. ACM Transactions on Embedded Computing Systems, 23(2): 27:1–27:??, March 2024. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3569420.

Martin:2015:ROS

[MWS15] Paul Martin, Lucas Wanner, and Mani Srivastava. Runtime optimization of system utility with variable hardware. ACM Transactions on Embedded Computing Systems, 14(2):24:1– 24:??, March 2015. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Mao:2022:TEA

[MYL^{+}22] Jiachen Mao, Qing Yang, Ang Li, Kent W. Nixon, Hai Li, and Yiran Chen. Toward efficient and adaptive design of video detection system with deep neural networks. ACM Transactions on Embedded Computing Systems, 21(3):33:1–33:21, May 2022. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3484946.

Mihajlovic:2014:DIQ

[MZG14] Bojan Mihajlović, Zeljko Zilić, and Warren J. Gross. Dynamically instrumenting the QEMU emulator for Linux process trace generation with the GDB debugger. ACM Transactions on Embedded Computing Systems, 13(5s):167:1– 167:??, November 2014. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Mihajlovic:2015:AAR

[MZG15] Bojan Mihajlović, Zeljko Zilić, and Warren J. Gross. Architecture-aware real-time compression of execution traces. ACM Transactions on Embedded Computing Systems, 14(4):75:1– 75:??, December 2015. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Namazi:2018:MBR

[NASM18] Alireza Namazi, Meisam Abdollahi, Saeed Safari, and Siamak Mohammadi. A majority-based reliabilityaware task mapping in highperformance homogeneous NoC architectures. ACM Transactions on Embedded Computing Systems, 17(1): 28:1–28:??, January 2018. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Noguera:2004:MRA

[NB04] Juanjo Noguera and Rosa M. Badia. Multitasking on reconfigurable architectures: microarchitecture support and dynamic scheduling. ACM Transactions on Embedded Computing Systems, 3(2):385–406, May 2004. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Nodeh:2018:EAM

[NBE18] Mohammad Taghi Teimoori Nodeh, Mostafa Bazzaz, and Alireza Ejlali. Exploiting approximate MLC-PCM in low-power embedded systems. ACM Transactions on Embedded Computing Systems, 17(1):26:1– 26:??, January 2018. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Nahapetian:2009:AAS

[NBGS09] Ani Nahapetian, Philip Brisk, Soheil Ghiasi, and Majid Sarrafzadeh. An approximation algorithm for scheduling on heterogeneous reconfigurable resources. ACM Transactions on Embedded Computing Systems, 9(1):5:1– 5:??, October 2009. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Nassar:2023:APM

[NBH23] Hassan Nassar, Lars Bauer, and Jörg Henkel. ANV-PUF: Machine-learning-resilient NVMbased arbiter PUF. ACM Transactions on Embedded Computing Systems, 22(5s): 133:1–133:??, October 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3609388.

Nouri:2016:ARA

[NBM^{+}16] Ayoub Nouri, Marius Bozga, Anca Molnos, Axel Legay, and Saddek Bensalem. AS-TROLABE: a rigorous approach for system-level performance modeling and analysis. ACM Transactions on Embedded Computing Systems, 15(2):31:1–31:??, May 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Narayan:2018:MTR

[NCJF18] Apurva Narayan, Greta Cutulenco, Yogi Joshi, and Sebastian Fischmeister. Mining timed regular specifications from system traces. ACM Transactions on Embedded Computing Systems, 17(2):46:1–46:??, April 2018. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Nguyen:2009:MAE

[NDB09] Nghi Nguyen, Angel Dominguez, and Rajeev Barua. Memory allocation for embedded systems with a compiletime-unknown scratch-pad size. ACM Transactions on Embedded Computing Systems, 8(3):21:1–21:??, April 2009. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Nazemzadeh:2013:FMD

[NDZ13] Payam Nazemzadeh, Abbas Dideban, and Meisam Zareiee. Fault modeling in discrete event systems using Petri nets. ACM Transactions on Embedded Computing Systems, 12(1):12:1– 12:??, January 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Niknafs:2023:RRM

[NEP23] Mina Niknafs, Petru Eles, and Zebo Peng. Runtime resource management with multiple-step-ahead workload prediction. ACM Transactions on Embedded Computing Systems, 22(4):71:1– 71:??, July 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3605213.

Nie:2022:HRA

[NFL^{+}22] Lanshun Nie, Chenghao Fan, Shuang Lin, Li Zhang, Yajuan Li, and Jing Li. Holistic resource allocation under federated scheduling for parallel real-time tasks. ACM Transactions on Embedded Computing Systems, 21(1): 13:1–13:29, January 2022. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3489467.

Naresh:2017:CCC

[NGL17] Vignyan Reddy Kothinti Naresh, Dibakar Gope, and Mikko H. Lipasti. The CURE: Cluster communication using registers. ACM Transactions on Embedded Computing Systems, 16(5s): 124:1–124:??, October 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Neshatpour:2020:IIC

[NHS20] Katayoun Neshatpour, Houman Homayoun, and Avesta Sasan. ICNN: The iterative convolutional neural network. ACM Transactions on Embedded Computing Systems, 18(6):1–27, January 2020. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3355553.

Nam:2012:MTI

[NKP^{+}12] Min-Young Nam, Kyungtae Kang, Rodolfo Pellizzoni, Kyung-Joon Park, Jung-Eun Kim, and Lui Sha. Modeling towards incremental early analyzability of networked avionics systems using virtual integration. ACM Transactions on Embedded Computing Systems, 11(4): 81:1–81:??, December 2012. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Napapetian:2012:ESS

[NKS12] Ani Napapetian, William Kaiser, and Majid Sarrafzadeh. Editorial: Special section on WHS’09. ACM Transactions on Embedded Computing Systems, 11(S2): 45:1–45:??, 2012. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Nuzzo:2019:SAG

[NLSV^{+}19] Pierluigi Nuzzo, Jiwei Li, Alberto L. Sangiovanni-Vincentelli, Yugeng Xi, and Dewei Li. Stochastic assume–guarantee contracts for cyber-physical system design. ACM Transactions on Embedded Computing Systems, 18(1):2:1– 2:??, February 2019. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3243216.

Nirjon:2014:MSR

[NNH^{+}14] Shahriar Nirjon, Angela Nicoara, Cheng-Hsin Hsu, Jatinder Pal Singh, and John A. Stankovic. Multi-Nets: a system for real-time switching between multiple network interfaces on mobile devices. ACM Transactions on Embedded Computing Systems, 13(4s):121:1– 121:??, April 2014. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Nadezhkin:2013:AGP

[NNS13] Dmitry Nadezhkin, Hristo Nikolov, and Todor Stefanov. Automated generation of polyhedral process networks from affine nested-loop programs with dynamic loop bounds. ACM Transactions on Embedded Computing Systems, 13(1s): 28:1–28:??, November 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Narang:2024:TTE

[NODP24] Gaurav Narang, Chukwufumnanya Ogbogu, Janardhan Rao Doppa, and Partha Pratim Pande. TEFLON: Thermally efficient dataflowaware 3D NoC for accelerating CNN inferencing on manycore PIM architectures. ACM Transactions on Embedded Computing Systems, 23(5):78:1–78:??, September 2024. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3665279.

Naik:2004:CCS

[NP04] Mayur Naik and Jens Palsberg. Compiling with codesize constraints. ACM Transactions on Embedded Computing Systems, 3(1): 163–181, February 2004.

CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Nghiem:2012:TTI

[NPAG12] Truong Nghiem, George J.

Pappas, Rajeev Alur, and Antoine Girard. Timetriggered implementations of dynamic controllers. ACM Transactions on Embedded Computing Systems, 11(S2): 58:1–58:??, 2012. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Nikitakis:2013:NLP

[NPP13] Antonis Nikitakis, Savvas Papaioannou, and Ioannis Papaefstathiou. A novel low-power embedded object recognition system working at multi-frames per second. ACM Transactions on Embedded Computing Systems, 12(1s):33:1–33:??, March 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Nam:2013:PAR

[NRL13] Yunyoung Nam, Seungmin Rho, and Chulung Lee. Physical activity recognition using multiple sensors embedded in a wearable device. ACM Transactions on Embedded Computing Systems, 12(2):26:1–26:??, February 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Niu:2024:EMF

[NRZ^{+}24] Linwei Niu, Danda B. Rawat, Dakai Zhu, Jonathan Musselwhite, Zonghua Gu, and Qingxu Deng. Energy management for faulttolerant (m, k)-constrained real-time systems that use standby-sparing. ACM Transactions on Embedded Computing Systems, 23(3): 36:1–36:??, May 2024. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3648365.

Nagar:2016:FPW

[NS16] Kartik Nagar and Y. N. Srikant. Fast and precise worst-case interference placement for shared cache analysis. ACM Transactions on Embedded Computing Systems, 15(3):45:1– 45:??, July 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Nagar:2017:RCB

[NS17] Kartik Nagar and Y. N. Srikant. Refining cache behavior prediction using cache miss paths. ACM Transactions on Embedded Computing Systems, 16(4):103:1– 103:??, August 2017. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Nair:2011:EHB

[NSL11] Ajay Nair, Karthik Shankar, and Roman Lysecky. Efficient hardware-based nonintrusive dynamic application profiling. ACM Transactions on Embedded Computing Systems, 10(3):32:1– 32:??, April 2011. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Nejatollahi:2020:SFA

[NVB^{+}20] Hamid Nejatollahi, Felipe Valencia, Subhadeep Banik, Francesco Regazzoni, Rosario Cammarota, and Nikil Dutt. Synthesis of flexible accelerators for early adoption of ring-LWE post-quantum cryptography. ACM Transactions on Embedded Computing Systems, 19(2):11:1– 11:17, March 2020. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3378164.

Namin:2012:EFF

[NWA12] Ashkan Hosseinzadeh Namin, Huapeng Wu, and Majid Ahmadi. An efficient finite field multiplier using redundant representation. ACM Transactions on Embedded Computing Systems, 11(2): 31:1–31:??, July 2012. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Navarro:2020:MLM

[NYH^{+}20] Osvaldo Navarro, Jones Yudi, Javier Hoffmann, Hector Gerardo Muñoz Hernandez, and Michael Hübner. A machine learning methodology for cache memory design based on dynamic instructions. ACM Transactions on Embedded Computing Systems, 19(2):12:1– 12:20, March 2020. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3376920.

Noltsis:2019:CLC

[NZCS19] Michail Noltsis, Nikolaos Zambelis, Francky Catthoor, and Dimitrios Soudris. A closed-loop controller to ensure performance and temperature constraints for dynamic applications. ACM Transactions on Embedded Computing Systems, 18(5): 40:1–40:??, October 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

                                        /dl.acm.org/ft_gateway.
                                        cfm?id=3343030.

Oneal:2017:GPE

[OBA^{+}17] Kenneth O’neal, Philip Brisk, Ahmed Abousamra, Zack Waters, and Emily Shriver. GPU performance estimation using software rasterization and machine learning. ACM Transactions on Embedded Computing Systems, 16(5s):148:1– 148:??, October 2017. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Odema:2023:MMA

[OBO^{+}23] Mohanad Odema, Halima Bouzidi, Hamza Ouarnoughi, Smail Niar, and Mohammad Abdullah Al Faruque. MaGNAS: a mapping-aware graph neural architecture search framework for heterogeneous MPSoC deployment. ACM Transactions on Embedded Computing Systems, 22(5s):108:1–108:??, October 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3609386.

Olivier:2016:MEP

[OBSO16] Pierre Olivier, Jalil Boukhobza, Eric Senn, and Hamza Ouarnoughi. A methodology for estimating performance and power consumption of embedded flash file systems. ACM Transactions on Embedded Computing Systems, 15(4):79:1– 79:??, August 2016. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Owaida:2015:EDS

[OFA^{+}15] Muhsen Owaida, Gabriel Falcao, Joao Andrade, Christos Antonopoulos, Nikolaos Bellas, Madhura Purnaprajna, David Novo, Georgios Karakonstantis, Andreas Burg, and Paolo Ienne. Enhancing design space exploration by extending CPU/GPU specifications onto FPGAs. ACM Transactions on Embedded Computing Systems, 14(2):33:1– 33:??, March 2015. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Oza:2024:DAT

[OHCK24] Pratham Oza, Nathaniel Hudson, Thidapat Chantem, and Hana Khamfroush. Deadline-aware task offloading for vehicular edge computing networks using traffic light data. ACM Transactions on Embedded Computing Systems, 23(1):8:1– 8:??, January 2024. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3594541.

Ost:2013:PAD

[OMA^{+}13] Luciano Ost, Marcelo Mandelli, Gabriel Marchesan Almeida, Leandro Moller, Leandro Soares Indrusiak, Gilles Sassatelli, Pascal Benoit, Manfred Glesner, Michel Robert, and Fernando Moraes. Power-aware dynamic mapping heuristics for NoC-based MPSoCs using a unified model-based approach. ACM Transactions on Embedded Computing Systems, 12(3):75:1– 75:??, March 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Oh:2023:MFL

[OMH^{+}23] Deok-Jae Oh, Yaebin Moon, Do Kyu Ham, Tae Jun Ham, Yongjun Park, Jae W. Lee, Jung Ho Ahn, and Eojin Lee. MaPHeA: a framework for lightweight memory hierarchy-aware profile-guided heap allocation. ACM Transactions on Embedded Computing Systems, 22(1):2:1–2:??, January 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL

https://dl.acm.org/doi/
10.1145/3527853.

Ogras:2023:ISI

[OMMK23] Umit Y. Ogras, Radu Marculescu, Trevor N. Mudge, and Michael Kishinevsky. Introduction to the special issue on domain-specific system-on-chip architectures and run-time management techniques. ACM Transactions on Embedded Computing Systems, 22(2):27:1– 27:??, March 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3567834.

Ozer:2008:SBE

[ÖNG08] Emre Özer, Andy P. Nisbet, and David Gregg. A stochastic bitwidth estimation technique for compact and lowpower custom processors. ACM Transactions on Embedded Computing Systems, 7(3):34:1–34:??, April 2008. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Ottoni:2006:OAU

[OOAL06] Desiree Ottoni, Guilherme Ottoni, Guido Araujo, and Rainer Leupers. Offset assignment using simultaneous variable coalescing. ACM Transactions on Embedded Computing Systems, 5(4): 864–883, November 2006. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Ou:2006:DSE

[OP06] Jingzhao Ou and Viktor K. Prasanna. Design space exploration using arithmetic-level hardware– software cosimulation for configurable multiprocessor platforms. ACM Transactions on Embedded Computing Systems, 5(2):355– 382, May 2006. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Oneto:2016:LHF

[ORA16] Luca Oneto, Sandro Ridella, and Davide Anguita. Learning hardware-friendly classifiers through algorithmic stability. ACM Transactions on Embedded Computing Systems, 15(2):23:1–23:??, May 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Omar:2018:DRH

[OSA^{+}18] Hamza Omar, Qingchuan Shi, Masab Ahmad, Halit Dogan, and Omer Khan. Declarative resilience: a holistic soft-error resilient multicore architecture that trades off program accuracy for efficiency. ACM Transactions on Embedded Computing Systems, 17(4):76:1– 76:??, August 2018. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Oehlert:2019:CIT

[OSF19] Dominic Oehlert, Selma Saidi, and Heiko Falk. Codeinherent traffic shaping for hard real-time systems. ACM Transactions on Embedded Computing Systems, 18(5s):108:1–108:??, October 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/ft_gateway.
cfm?id=3358215.

Ouyang:2022:WWF

[OZ22] Xiangzhen Ouyang and Yian Zhu. wfspan: Wait-free dynamic memory management. ACM Transactions on Embedded Computing Systems, 21(4):43:1–43:??, July 2022. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3533724.

Paissan:2022:PSB

[PAF22] Francesco Paissan, Alberto Ancilotto, and Elisabetta Farella. PhiNets: a scalable backbone for low-power AI at the edge. ACM Transactions on Embedded Computing Systems, 21(5):53:1– 53:??, September 2022. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3510832.

Paterna:2012:VTW

[PAP^{+}12] Francesco Paterna, Andrea Acquaviva, Francesco Papariello, Giuseppe Desoli, and Luca Benini. Variabilitytolerant workload allocation for MPSoC energy minimization under real-time constraints. ACM Transactions on Embedded Computing Systems, 11(4):71:1– 71:??, December 2012. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Patterson:2009:SMB

[PAS^{+}09] C. Patterson, P. Athanas, M. Shelburne, J. Bowen, J. Surís, T. Dunham, and J. Rice. Slotless modulebased reconfiguration of embedded FPGAs. ACM Transactions on Embedded Computing Systems, 9(1): 6:1–6:??, October 2009. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Paul:2014:RTP

[Pau14]  Anand Paul.  Real-time

power management for embedded M2M using intelligent learning methods. ACM Transactions on Embedded Computing Systems, 13(5s): 148:1–148:??, September 2014. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Parikh:2014:FCF

[PB14] Ritesh Parikh and Valeria Bertacco. ForEVeR: a complementary formal and runtime verification approach to correct NoC functionality. ACM Transactions on Embedded Computing Systems, 13(3s):104:1–104:??, March 2014. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Pan:2022:BWH

[PBC22] Hongyi Pan, Diaa Badawi, and Ahmet Enis Cetin. Block Walsh-Hadamard transformbased binary layers in deep neural networks. ACM Transactions on Embedded Computing Systems, 21(6): 69:1–69:??, November 2022. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3510026.

Plaks:2009:GECa

[PBP09a] Toomas P. Plaks, Neil Bergmann, and Bernard Pottier. Guest editorial CAPA’08 configurable computing: Configuring algorithms, processes, and architecture issue I: Configuring algorithms and processes. ACM Transactions on Embedded Computing Systems, 9(1):1:1–1:??, October 2009. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Plaks:2009:GECb

[PBP09b] Toomas P. Plaks, Neil Bergmann, and Bernard Pottier. Guest editorial CAPA’08 Configurable computing: Configuring algorithms, processes, and architecture Issue II: Configuring hardware architecture. ACM Transactions on Embedded Computing Systems, 9(2): 10:1–10:??, October 2009.

CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Panainte:2007:MCR

[PBV07] Elena Moscu Panainte, Koen Bertels, and Stamatis Vassiliadis. The Molen compiler for reconfigurable processors. ACM Transactions on Embedded Computing Systems, 6(1):6:1– 6:??, February 2007. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Pagani:2014:EEA

[PC14] Santiago Pagani and Jian-Jia Chen. Energy efficiency analysis for the single frequency approximation (SFA) scheme. ACM Transactions on Embedded Computing Systems, 13(5s):158:1– 158:??, September 2014. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Pimentel:2012:ISS

[PCB12] Andy D. Pimentel, Naehyuck Chang, and Mladen Berekovic. Introduction to special section ESTI-Media’09. ACM Transactions on Embedded Computing Systems, 11(4):70:1– 70:??, December 2012. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Paul:2013:VSI

[PCBW13] Anand Paul, Bo-Wei Chen, Karunanithi Bharanitharan, and Jhing-Fa Wang. Video search and indexing with reinforcement agent for interactive multimedia services. ACM Transactions on Embedded Computing Systems, 12(2):25:1–25:??, February 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Pagliari:2017:ABC

[PCC17] Daniele Jahier Pagliari, Mario R. Casu, and Luca P. Carloni. Accelerators for breast cancer detection. ACM Transactions on Embedded Computing Systems, 16(3):80:1–80:??, July 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Paul:2021:ATA

[PCGD21] Suraj Paul, Navonil Chatterjee, Prasun Ghosal, and Jean-Philippe Diguet. Adaptive task allocation and scheduling on NoC-based multicore platforms with multitasking processors. ACM Transactions on Embedded Computing Systems, 20(1): 4:1–4:26, January 2021. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3408324.

Park:2008:RFF

[PCK^{+}08] Chanik Park, Wonmoon Cheon, Jeonguk Kang, Kangho Roh, Wonhee Cho, and Jin-Soo Kim. A reconfigurable FTL (flash translation layer) architecture for NAND flash-based applications. ACM Transactions on Embedded Computing Systems, 7(4):38:1–38:??, July 2008. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Pajic:2012:RAE

[PCM12] Miroslav Pajic, Alexander Chernoguzov, and Rahul Mangharam. Robust architectures for embedded wireless network control and actuation. ACM Transactions on Embedded Computing Systems, 11(4):82:1– 82:??, December 2012. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Papagiannopoulou:2015:EEH

[PCM^{+}15] Dimitra Papagiannopoulou, Giuseppe Capodanno, Tali Moreshet, Maurice Herlihy, and R. Iris Bahar. Energy-efficient and highperformance lock speculation hardware for embedded multicore systems. ACM Transactions on Embedded Computing Systems, 14(3): 51:1–51:??, May 2015. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Pasricha:2008:FEB

[PDBR08] Sudeep Pasricha, Nikil Dutt, and Mohamed Ben-Romdhane. Fast exploration of busbased communication architectures at the CCATB abstraction. ACM Transactions on Embedded Computing Systems, 7(2):22:1– 22:??, February 2008. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Paul:2023:ANI

[PDHC23] Sibendu Paul, Utsav Drolia, Y. Charlie Hu, and Srimat Chakradhar. AQuA: a new image quality metric for optimizing video analytics systems. ACM Transactions on Embedded Computing Systems, 22(4):74:1–74:??, July 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3568423.

Park:2021:IML

[PDL21] Jurn-Gyu Park, Nikil Dutt, and Sung-Soo Lim. An interpretable machine learning model enhanced integrated CPU–GPU DVFS governor. ACM Transactions on Embedded Computing Systems, 20(6):108:1– 108:28, November 2021. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3470974.

Parra:2023:TMV

[PDL^{+}23] Pablo Parra, Antonio Da Silva, Borja Losa, J. Ignacio García, Óscar R. Polo, Agustín Martínez, and Sebastián Sánchez. Tailormade virtualization monitor design for CPU virtualization on LEON processors. ACM Transactions on Embedded Computing Systems, 22(4):60:1–60:??, July 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3584702.

Peeck:2023:IWC

[PE23] Jonas Peeck and Rolf Ernst.

Improving worst-case TSN communication times of large sensor data samples by exploiting synchronization. ACM Transactions on Embedded Computing Systems, 22(5s):139:1–139:??, October 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3609120.

Pop:2005:SDF

[PEP05] Paul Pop, Petru Eles, and Zebo Peng. Schedulabilitydriven frame packing for multicluster distributed embedded systems. ACM Transactions on Embedded Computing Systems, 4(1): 112–140, February 2005. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Peres:2024:PTU

[PG24] Florent Peres and Mohamed Ghazel. A proven translation from a UML state machine subset to timed automata. ACM Transactions on Embedded Computing Systems, 23(5):72:1–72:??, September 2024. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3581771.

Popovici:2008:PBS

[PGR^{+}08] Katalin Popovici, Xavier Guerin, Frederic Rousseau, Pier Stanislao Paolucci, and Ahmed Amine Jerraya. Platform-based software design flow for heterogeneous MPSoC. ACM Transactions on Embedded Computing Systems, 7(4):39:1– 39:??, July 2008. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Poddar:2016:DHP

[PGR16] Soumyajit Poddar, Prasun Ghosal, and Hafizur Rahaman. Design of a high-performance CDMAbased broadcast-free photonic multi-core network on chip. ACM Transactions on Embedded Computing Systems, 15(1):2:1– 2:??, February 2016. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Papakonstantinou:2013:ECC

[PGS^{+}13] Alexandros Papakonstantinou, Karthik Gururaj, John A. Stratton, Deming Chen, Jason Cong, and Wen-Mei W. Hwu. Efficient compilation of CUDA kernels for high-performance computing on FPGAs. ACM Transactions on Embedded Computing Systems, 13(2):25:1– 25:??, September 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Park:2018:SCG

[PHDL18] Jurn-Gyu Park, Chen-Ying Hsieh, Nikil Dutt, and Sung-Soo Lim. Synergistic CPU–GPU frequency capping for energy-efficient mobile games. ACM Transactions on Embedded Computing Systems, 17(2):45:1– 45:??, April 2018. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Procter:2017:PAS

[PHG^{+}17] Adam Procter, William L.

Harrison, Ian Graves, Michela Becchi, and Gerard Allwein. A principled approach to secure multi-core processor design with ReWire. ACM Transactions on Embedded Computing Systems, 16(2): 33:1–33:??, April 2017. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Pajic:2014:SCM

[PJL^{+}14] Miroslav Pajic, Zhihao Jiang, Insup Lee, Oleg Sokolsky, and Rahul Mangharam. Safety-critical medical device development using the UPP2SF model translation tool. ACM Transactions on Embedded Computing Systems, 13(4s):127:1– 127:??, April 2014. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Park:2017:FPC

[PJL^{+}17] Jaehyun Park, Hitesh Joshi, Hyung Gyu Lee, Sayfe Kiaei, and Umit Y. Ogras. Flexible PV-cell modeling for energy harvesting in wearable IoT applications. ACM Transactions on Embedded Computing Systems, 16(5s): 156:1–156:??, October 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Pager:2015:SSM

[PJS15] Jared Pager, Reiley Jeyapaul, and Aviral Shrivastava. A software scheme for multithreading on CGRAs. ACM Transactions on Embedded Computing Systems, 14(1): 19:1–19:??, January 2015. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Pujol:2023:VEC

[PJT^{+}23] Roger Pujol, Josep Jorba, Hamid Tabani, Leonidas Kosmidis, Enrico Mezzetti, Jaume Abella, and Francisco Cazorla. Vector extensions in COTS processors to increase guaranteed performance in real-time systems. ACM Transactions on Embedded Computing Systems, 22(2): 21:1–21:??, March 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3561054.

Paul:2012:PRC

[PJWY12] Anand Paul, Yung-Chuan Jiang, Jhing-Fa Wang, and Jar-Ferr Yang. Parallel reconfigurable computingbased mapping algorithm for motion estimation in advanced video coding. ACM Transactions on Embedded Computing Systems, 11(S2): 40:1–40:??, 2012. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Park:2013:EEN

[PK13] Sang Oh Park and Sung Jo Kim. ENFFiS: an enhanced NAND flash memory file system for mobile embedded multimedia system. ACM Transactions on Embedded Computing Systems, 12(2): 23:1–23:??, February 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Papaioannou:2023:ULP

[PKIT23] Alexios Papaioannou, Charalampos S. Kouzinopoulos, Dimosthenis Ioannidis, and Dimitrios Tzovaras. An ultra-low-power embedded AI fire detection and crowd counting system for indoor areas. ACM Transactions on Embedded Computing Systems, 22(4):61:1–61:??, July 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3582433.

Park:2022:QST

[PKL22] Jun-Hyung Park, Kang-Min Kim, and Sangkeun Lee. Quantized sparse training: a unified trainable framework for joint pruning and quantization in DNNs. ACM Transactions on Embedded Computing Systems, 21(5): 60:1–60:??, September 2022. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3524066.

Pearce:2023:HLA

[PKT23] Hammond Pearce, Ramesh Karri, and Benjamin Tan. High-level approaches to hardware security: a tutorial. ACM Transactions on Embedded Computing Systems, 22(3):45:1–45:??, May 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3577200.

Peng:2010:OWZ

[PL10] Huan-Kai Peng and Youn-Long Lin. An optimal warning-zone-length assignment algorithm for real-time and multiple-QoS on-chip bus arbitration. ACM Transactions on Embedded Computing Systems, 9(4):35:1– 35:??, March 2010. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Palem:2013:TYB

[PL13] Krishna Palem and Avinash Lingamneni. Ten years of building broken chips: The physics and engineering of inexact computing. ACM Transactions on Embedded Computing Systems, 12(2s):87:1–87:??, May 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Plaks:2012:ESS

[Pla12] Toomas P. Plaks. Editorial:

Special section on CAPA’09. ACM Transactions on Embedded Computing Systems, 11(S2):39:1–39:??, 2012. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Park:2008:QSL

[PLKH08] Jiyong Park, Jaesoo Lee, Saehwa Kim, and Seongsoo Hong. Quasistatic shared libraries and XIP for memory footprint reduction in MMU-less embedded systems. ACM Transactions on Embedded Computing Systems, 8(1):6:1–6:??, December 2008. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Petrucci:2015:EET

[PLM^{+}15] Vinicius Petrucci, Orlando Loques, Daniel Mossé, Rami Melhem, Neven Abou Gazala, and Sameh Gobriel. Energyefficient thread assignment optimization for heterogeneous multicore systems. ACM Transactions on Embedded Computing Systems, 14(1):15:1–15:??, January 2015. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Pradhan:2023:ETB

[PLT23] Chetana Pradhan, Martin Letras, and Jürgen Teich. Efficient table-based function approximation on FP-GAs using interval splitting and BRAM instantiation. ACM Transactions on Embedded Computing Systems, 22(4):73:1–73:??, July 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3580737.

Ponugoti:2019:EFH

[PM19] Mounika Ponugoti and Aleksandar Milenkovic. Enabling on-the-fly hardware tracing of data reads in multicores. ACM Transactions on Embedded Computing Systems, 18(4):34:1– 34:??, August 2019. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3322642.

Plassan:2019:MMA

[PMAB19] Guillaume Plassan, Katell Morin-Allory, and Dominique Borrione. Mining missing assumptions from counterexamples. ACM Transactions on Embedded Computing Systems, 18(1):3:1– 3:??, February 2019. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3288759.

Piccolboni:2017:CCH

[PMDC17] Luca Piccolboni, Paolo Mantovani, Giuseppe Di Guglielmo, and Luca P. Carloni. COSMOS: Coordination of high-level synthesis and memory optimization for hardware accelerators. ACM Transactions on Embedded Computing Systems, 16(5s):150:1–150:??, October 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Paolieri:2013:HRT

[PMM^{+}13] Marco Paolieri, Jörg Mische, Stefan Metzlaff, Mike Gerdes, Eduardo Quiñones, Sascha Uhrig, Theo Ungerer, and Francisco J. Cazorla. A hard real-time capable multicore SMT processor. ACM Transactions on Embedded Computing Systems, 12(3): 79:1–79:??, March 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Papagiannopoulou:2017:ETE

[PMM^{+}17] Dimitra Papagiannopoulou, Andrea Marongiu, Tali Moreshet, Maurice Herlihy, and R. Iris Bahar. Edge-TM: Exploiting transactional memory for error tolerance and energy efficiency. ACM Transactions on Embedded Computing Systems, 16(5s):153:1–153:??, October 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Piccolboni:2017:ECF

[PMP17] Luca Piccolboni, Alessandro Menon, and Graziano Pravadelli. Efficient controlflow subgraph matching for detecting hardware Trojans in RTL models. ACM Transactions on Embedded Computing Systems, 16(5s): 137:1–137:??, October 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Park:2014:AWL

[PMPP14] Sung Kyu Park, Min Kyu Maeng, Ki-Woong Park, and Kyu Ho Park. Adaptive wear-leveling algorithm for PRAM main memory with a DRAM buffer. ACM Transactions on Embedded Computing Systems, 13(4): 88:1–88:??, February 2014. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Pan:2024:MTR

[PMS^{+}24] Yungang Pan, Rouhollah Mahfouzi, Soheil Samii, Petru Eles, and Zebo Peng. Multi-traffic resource optimization for real-time applications with 5G configured Grant scheduling. ACM Transactions on Embedded Computing Systems, 23(4): 63:1–63:??, July 2024. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3664621.

Puthal:2017:DDK

[PNRC17] Deepak Puthal, Surya Nepal, Rajiv Ranjan, and Jinjun Chen. DLSeF: a dynamic key-length-based efficient real-time security verification model for big data stream. ACM Transactions on Embedded Computing Systems, 16(2):51:1– 51:??, April 2017. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Petrov:2005:RCF

[PO05] Peter Petrov and Alex Orailoglu. A reprogrammable customization framework for efficient branch resolution in embedded processors. ACM Transactions on Embedded Computing Systems, 4(2): 452–468, May 2005. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Pavan:2024:TAD

[POCR24] Massimo Pavan, Eugeniu Ostrovan, Armando Caltabiano, and Manuel Roveri. TyBox: an automatic design and code generation toolbox for TinyML incremental on-device learning. ACM Transactions on Embedded Computing Systems, 23(3):42:1–42:??, May 2024. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3604566.

Paulin:2013:PPP

[PÖG^{+}13] Pierre G. Paulin, Ali Erdem Özcan, Vincent Gagné, Bruno Lavigueur, and Olivier Benny. Parallel programming patterns for multiprocessor SoC: Application to video processing. ACM Transactions on Embedded Computing Systems, 12(1s): 47:1–47:??, March 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Pradhan:2012:AVJ

[PP12] Gaurav N. Pradhan and B. Prabhakaran. Analyzing and visualizing jump performance using wireless body sensors. ACM Transactions on Embedded Computing Systems, 11(S2):47:1–47:??, 2012. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Pan:2019:MTP

[PP19] Runyu Pan and Gabriel Parmer. MxU: Towards predictable, flexible, and efficient memory access control for the secure IoT. ACM Transactions on Embedded Computing Systems, 18(5s):103:1–103:??, October 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/ft_gateway.
cfm?id=3358224.

Pandey:2024:NTM

[PP24] Shailja Pandey and Preeti Ranjan Panda. NeuroTAP: Thermal and memory access pattern-aware data mapping on 3D DRAM for maximizing DNN performance. ACM Transactions on Embedded Computing Systems, 23(6):96:1–96:??, November 2024. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3677178.

Pederson:2019:BCL

[PQA^{+}19] Daniel J. Pederson, Christopher J. Quinkert, Muhammad A. Arafat, Jesse P. Somann, Jack D. Williams, Rebecca A. Bercich, Zhi Wang, Gabriel O. Albors, John G. R. Jefferys, and Pedro P. Irazoqui. The Bionode: a closed-loop neuromodulation implant. ACM Transactions on Embedded Computing Systems, 18(1):9:1– 9:??, February 2019. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3301310.

Peon-quiros:2015:PLD

[PqBM^{+}15] Miguel Peón-quirós, Alexandros Bartzas, Stylianos Mamagkakis, Francky Catthoor, José Manuel Mendías, and Dimitrios Soudris. Placement of linked dynamic data structures over heterogeneous memories in embedded systems. ACM Transactions on Embedded Computing Systems, 14(2):37:1– 37:??, March 2015. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Palossi:2015:CDP

[PRB15] Daniele Palossi, Martino Ruggiero, and Luca Benini. 3D CV descriptor on parallel heterogeneous platforms. ACM Transactions on Embedded Computing Systems, 14(4):73:1–73:??, December 2015. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Pan:2015:HFY

[PRK15] Abhisek Pan, Rance Rodrigues, and Sandip Kundu. A hardware framework for yield and reliability enhancement in chip multiprocessors. ACM Transactions on Embedded Computing Systems, 14(1):12:1– 12:??, January 2015. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Ponzina:2023:OFC

[PRL^{+}23] Flavio Ponzina, Marco Rios, Alexandre Levisse, Giovanni Ansaloni, and David Atienza. Overflow-free compute memories for edge AI acceleration. ACM Transactions on Embedded Computing Systems, 22(5s):121:1– 121:??, October 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3609387.

Poudel:2021:MFU

[PRM21] Prawar Poudel, Biswajit Ray, and Aleksandar Milenkovic. Microcontroller fingerprinting using partially erased NOR flash memory cells. ACM Transactions on Embedded Computing Systems, 20(3):26:1–26:23, April 2021. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3448271.

Pinisetty:2017:REC

[PRS^{+}17] Srinivas Pinisetty, Partha S. Roop, Steven Smyth, Nathan Allen, Stavros Tripakis, and Reinhard Von Hanxleden. Runtime enforcement of cyber-physical systems. ACM Transactions on Embedded Computing Systems, 16(5s):178:1–178:??, October 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Passerone:2019:CEC

[PRSV19] Roberto Passerone, Íñigo Íncer Romeo, and Alberto L.

Sangiovanni-Vincentelli. Coherent extension, composition, and merging operators in contract models for system design. ACM Transactions on Embedded Computing Systems, 18(5s):86:1– 86:??, October 2019. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3358216.

Park:2004:LLS

[PS04] Taejoon Park and Kang G.

Shin. LiSP: a lightweight security protocol for wireless sensor networks. ACM Transactions on Embedded Computing Systems, 3(3): 634–660, August 2004. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Park:2008:ATL

[PS08a] Taejoon Park and Kang G.

Shin. Attack-tolerant localization via iterative verification of locations in sensor networks. ACM Transactions on Embedded Computing Systems, 8(1):2:1– 2:??, December 2008. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Park:2008:SRB

[PS08b] Taejoon Park and Kang G.

Shin. Secure routing based on distributed key sharing in large-scale sensor networks. ACM Transactions on Embedded Computing Systems, 7(2):20:1–20:??, February 2008. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Pitter:2010:RTJ

[PS10] Christof Pitter and Martin Schoeberl. A real-time Java chip-multiprocessor. ACM Transactions on Embedded Computing Systems, 10(1): 9:1–9:??, August 2010. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Palesi:2014:ESS

[PS14] Maurizio Palesi and Todor Stefanov. Editorial: Special section on ESTIMedia’13. ACM Transactions on Embedded Computing Systems, 13(3s):110:1–110:??, March 2014. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Park:2019:ERR

[PS19] Taeju Park and Kang G. Shin. EACAN: Reliable and resource-efficient CAN communications. ACM Transactions on Embedded Computing Systems, 18(1):8:1– 8:??, February 2019. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

                                        tronic). URL https://dl.
                                        acm.org/ft_gateway.cfm?
                                        id=3301309.

Pazzaglia:2021:GWH

[PSD21] Paolo Pazzaglia, Youcheng Sun, and Marco Di Natale. Generalized weakly hard schedulability analysis for real-time periodic tasks. ACM Transactions on Embedded Computing Systems, 20(1):3:1–3:26, January 2021. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL

https://dl.acm.org/doi/
10.1145/3404888.

Palermo:2012:VAR

[PSZ12a] Gianluca Palermo, Cristina Silvano, and Vittorio Zaccaria. A variability-aware robust design space exploration methodology for onchip multiprocessors subject to application-specific constraints. ACM Transactions on Embedded Computing Systems, 11(2):29:1– 29:??, July 2012. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Peng:2012:BHA

[PSZ12b] Chunyi Peng, Guobin Shen, and Yongguang Zhang. BeepBeep: a high-accuracy acoustic-based system for ranging and localization using COTS devices. ACM Transactions on Embedded Computing Systems, 11(1): 4:1–4:??, March 2012. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Pal:2022:OEI

[PVSG22] Subhankar Pal, Swagath Venkataramani, Viji Srinivasan, and Kailash Gopalakrishnan. OnSRAM: Efficient inter-node on-chip scratchpad management in deep learning accelerators. ACM Transactions on Embedded Computing Systems, 21(6): 86:1–86:??, November 2022. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3530909.

Panopoulos:2024:CCA

[PVV24] Ioannis Panopoulos, Stylianos Venieris, and Iakovos Venieris. CARIn: Constraintaware and responsive inference on heterogeneous devices for single- and multi-DNN workloads. ACM Transactions on Embedded Computing Systems, 23(4): 60:1–60:??, July 2024. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3665868.

Parmer:2013:PCC

[PW13] Gabriel Parmer and Richard West. Predictable and configurable component-based scheduling in the Composite OS. ACM Transactions on Embedded Computing Systems, 13(1s):32:1– 32:??, November 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Park:2019:MOE

[PWL^{+}19] Sunghyun Park, Youfeng Wu, Janghaeng Lee, Amir Aupov, and Scott Mahlke. Multi-objective exploration for practical optimization decisions in binary translation. ACM Transactions on Embedded Computing Systems, 18(5s):57:1–57:??, October 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/ft_gateway.
cfm?id=3358185.

Pan:2018:MAC

[PX18] Wen Pan and Tao Xie. A mirroring-assisted channel-RAID5 SSD for mobile applications. ACM Transactions on Embedded Computing Systems, 17(4):75:1– 75:??, August 2018. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Pan:2017:EMW

[PXY^{+}17] Chen Pan, Mimi Xie, Chengmo Yang, Yiran Chen, and Jingtong Hu. Exploiting multiple write modes of nonvolatile main memory in embedded systems. ACM Transactions on Embedded Computing Systems, 16(4):110:1–110:??, August 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Pan:2015:SGP

[PYJL15] Gung-Yu Pan, Jed Yang, Jing-Yang Jou, and Bo-Cheng Charles Lai. Scalable global power management policy based on combinatorial optimization for multiprocessors. ACM Transactions on Embedded Computing Systems, 14(4):70:1– 70:??, December 2015. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Pan:2023:BBS

[PYL^{+}23] Yunjie Pan, Jiecao Yu, Andrew Lukefahr, Reetuparna Das, and Scott Mahlke. Bit-SET: Bit-serial early termination for computation reduction in convolutional neural networks. ACM Transactions on Embedded Computing Systems, 22(5s):98:1– 98:??, October 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3609093.

Pande:2012:PDP

[PZ12] Amit Pande and Joseph Zambreno. Poly-DWT: Polymorphic wavelet hardware support for dynamic image compression. ACM Transactions on Embedded Computing Systems, 11(1):6:1– 6:??, March 2012. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Que:2024:LGL

[QFL^{+}24] Zhiqiang Que, Hongxiang Fan, Marcus Loo, He Li, Michaela Blott, Maurizio Pierini, Alexander Tapper, and Wayne Luk. LL-GNN: Low latency graph neural networks on FPGAs for high energy physics. ACM Transactions on Embedded Computing Systems, 23(2):17:1– 17:??, March 2024. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3640464.

Quan:2007:EED

[QH07] Gang Quan and Xiaobo Sharon Hu. Energy efficient DVS schedule for fixed-priority real-time systems. ACM Transactions on Embedded Computing Systems, 6(4): 29:1–29:??, September 2007. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Qu:2003:SSS

[QP03] Gang Qu and Miodrag Potkonjak. System synthesis of synchronous multimedia applications. ACM Transactions on Embedded Computing Systems, 2(1): 74–97, February 2003. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Quan:2015:HTM

[QP15] Wei Quan and Andy D. Pimentel. A hybrid task mapping algorithm for heterogeneous MPSoCs. ACM Transactions on Embedded Computing Systems, 14(1): 14:1–14:??, January 2015. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Quwaider:2010:TPA

[QRB10] Muhannad Quwaider, Jayanthi Rao, and Subir Biswas. Transmission power assignment with postural position inference for on-body wireless communication links. ACM Transactions on Embedded Computing Systems, 10(1):14:1–14:??, August 2010. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Qi:2024:MCS

[QRW^{+}24] Huamei Qi, Fang Ren, Leilei Wang, Ping Jiang, Shaohua Wan, and Xiaoheng Deng. Multi-compression scale DNN inference acceleration based on cloud-edge-end collaboration. ACM Transactions on Embedded Computing Systems, 23(1): 16:1–16:??, January 2024.

CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3634704.

Qian:2018:ECD

[QWY^{+}18] Kun Qian, Chenshu Wu, Zheng Yang, Yunhao Liu, Fugui He, and Tianzhang Xing. Enabling contactless detection of moving humans with dynamic speeds using CSI. ACM Transactions on Embedded Computing Systems, 17(2):52:1–52:??, April 2018. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Qiu:2014:BPD

[QZXO14] Keni Qiu, Mengying Zhao, Chun Jason Xue, and Alex Orailoglu. Branch prediction-directed dynamic instruction cache locking for embedded systems. ACM Transactions on Embedded Computing Systems, 13(5s): 156:1–156:??, September 2014. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Rodrigues:2014:LPI

[RAK14] Rance Rodrigues, Arunachalam Annamalai, and Sandip Kundu. A low-power instruction replay mechanism for design of resilient microprocessors. ACM Transactions on Embedded Computing Systems, 13(4):85:1– 85:??, February 2014. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Ren:2024:DOL

[RALR24] Haoyu Ren, Darko Anicic, Xue Li, and Thomas Runkler. On-device online learning and semantic management of TinyML systems. ACM Transactions on Embedded Computing Systems, 23(4):55:1–55:??, July 2024. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3665278.

Rathor:2025:AFE

[Rat25] Mahendra Rathor. ALOHA-FP2I: Efficient algorithms and hardware for multimode rounding of floating point to integer. ACM Transactions on Embedded Computing Systems, 24(1): 12:1–12:26, January 2025. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3701560.

Ray:2021:HAS

[RB21] Kaustabha Ray and Ansuman Banerjee. Horizontal auto-scaling for multiaccess edge computing using safe reinforcement learning. ACM Transactions on Embedded Computing Systems, 20(6):109:1–109:33, November 2021. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL

https://dl.acm.org/doi/
10.1145/3475991.

Roy:2019:CPR

[RBNM19] Debapriya Basu Roy, Shivam Bhasin, Ivica Nikolić, and Debdeep Mukhopadhyay. Combining PUF with RLUTs: a two-party pay-per-device IP licensing scheme on FP-GAs. ACM Transactions on Embedded Computing Systems, 18(2):12:1–12:??, April 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/ft_gateway.
cfm?id=3301307.

Razavi:2010:SAB

[RBS^{+}10] Niloofar Razavi, Razieh Behjati, Hamideh Sabouri, Ehsan Khamespanah, Amin Shali, and Marjan Sirjani. Sysfier: Actor-based formal verification of SystemC. ACM Transactions on Embedded Computing Systems, 10(2):19:1–19:??, December 2010. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Raman:2008:ASW

[RC08] Balaji Raman and Samarjit Chakraborty. Applicationspecific workload shaping in multimedia-enabled personal mobile devices. ACM Transactions on Embedded Computing Systems, 7(2):10:1– 10:??, February 2008. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Rai:2017:UCG

[RC17] Siddharth Rai and Mainak Chaudhuri. Using criticality of GPU accesses in memory management for CPU–GPU heterogeneous multi-core processors. ACM Transactions on Embedded Computing Systems, 16(5s): 133:1–133:??, October 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Ravi:2024:SCF

[RCDB24] Prasanna Ravi, Anupam Chattopadhyay, Jan Pieter D’Anvers, and Anubhab Baksi. Side-channel and fault-injection attacks over lattice-based post-quantum schemes (Kyber, Dilithium): Survey and new results. ACM Transactions on Embedded Computing Systems, 23(2):35:1–35:??, March 2024. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3603170.

Ravi:2023:VLV

[RCS23] Akshara Ravi, Vivek Chaturvedi, and Muhammad Shafique. ViT4Mal: Lightweight vision transformer for malware detection on edge devices. ACM Transactions on Embedded Computing Systems, 22(5s):117:1–117:??, October 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3609112.

Reshadi:2006:RFI

[RDM06] Mehrdad Reshadi, Nikil Dutt, and Prabhat Mishra. A retargetable framework for instruction-set architecture simulation. ACM Transactions on Embedded Computing Systems, 5(2):431– 452, May 2006. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Rouxel:2017:TCD

[RDP17] Benjamin Rouxel, Steven Derrien, and Isabelle Puaut. Tightening contention delays while scheduling parallel applications on multicore architectures. ACM Transactions on Embedded Computing Systems, 16(5s): 164:1–164:??, October 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Roy:2021:SQL

[RDSS21] Sanjit Kumar Roy, Rajesh Devaraj, Arnab Sarkar, and Debabrata Senapati. SLAQA: Quality-level aware scheduling of task graphs on heterogeneous distributed systems. ACM Transactions on Embedded Computing Systems, 20(5):45:1– 45:31, July 2021. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3462776.

Rafiliu:2015:SOR

[REPL15] Sergiu Rafiliu, Petru Eles, Zebo Peng, and Michael Lemmon. Stability of online resource managers for distributed systems under execution time variations. ACM Transactions on Embedded Computing Systems, 14(2): 21:1–21:??, March 2015. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Reineke:2013:SCR

[RG13] Jan Reineke and Daniel Grund. Sensitivity of cache replacement policies. ACM Transactions on Embedded Computing Systems, 12(1s): 42:1–42:??, March 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Razaghi:2014:HCM

[RG14] Parisa Razaghi and Andreas Gerstlauer. Host-compiled multicore system simulation for early real-time performance evaluation. ACM Transactions on Embedded Computing Systems, 13 (5s):166:1–166:??, November 2014. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Riemens:2014:TSA

[RGdZS14] Danny P. Riemens, Georgi N.

Gaydadjiev, Chris I. de Zeeuw, and Christos Strydis. Towards scalable arithmetic units with graceful degradation. ACM Transactions on Embedded Computing Systems, 13(4):87:1– 87:??, February 2014. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Raghunathan:2004:EEW

[RGSS04] Vijay Raghunathan, Saurabh Ganeriwal, Mani Srivastava, and Curt Schurgers. Energy efficient wireless packet scheduling and fair queuing. ACM Transactions on Embedded Computing Systems, 3(1):3–23, February 2004. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Rautakoura:2023:DSH

[RH23] Antti Rautakoura and Timo Hämäläinen. Does SoC hardware development become agile by saying so: a literature review and mapping study. ACM Transactions on Embedded Computing Systems, 22(3):44:1–44:??, May 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3578554.

Reardon:2012:REE

[RHG^{+}12] Casey Reardon, Brian Holland, Alan D. George, Greg Stitt, and Herman Lam. RCML: An environment for estimation modeling of reconfigurable computing systems. ACM Transactions on Embedded Computing Systems, 11(S2):43:1–43:??, 2012. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Rossebo:2014:ISI

[RHG^{+}14] Judith E. Y. Rossebø, Siv Hilde Houmb, Geri Georg, Virginia N. L. Franqueira, and Dimitrios Serpanos. Introduction to special issue on risk and trust in embedded critical systems. ACM Transactions on Embedded Computing Systems, 13(5s):152:1–152:??, September 2014. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Robertson:2004:DFP

[RI04] Ian Robertson and James Irvine. A design flow for partially reconfigurable hardware. ACM Transactions on Embedded Computing Systems, 3(2):257–283, May 2004. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Rahman:2021:LTW

[RIMS21] Mahbubur Rahman, Dali Ismail, Venkata P. Modekurthy, and Abusayeed Saifullah. LPWAN in the TV white spaces: a practical implementation and deployment experiences. ACM Transactions on Embedded Computing Systems, 20(4): 30:1–30:26, June 2021. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3447877.

Ruaro:2019:SAQ

[RJM19] Marcelo Ruaro, Axel Jantsch, and Fernando Gehm Moraes. Self-adaptive QoS management of computation and communication resources in many-core SoCs. ACM Transactions on Embedded Computing Systems, 18(4): 37:1–37:??, August 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/ft_gateway.
cfm?id=3328755.

Rhisheekesan:2019:CFC

[RJS19] Abhishek Rhisheekesan, Reiley Jeyapaul, and Aviral Shrivastava. Control flow checking or not? (for soft errors). ACM Transactions on Embedded Computing Systems, 18(1):11:1– 11:??, February 2019. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

                                        tronic). URL https://dl.
                                        acm.org/ft_gateway.cfm?
                                        id=3301311.

Resch:2022:EER

[RKC^{+}22] Salonik Resch, S. Karen Khatamifard, Zamshed I. Chowdhury, Masoud Zabihi, Zhengyang Zhao, Husrev Cilasun, Jian-Ping Wang, Sachin S. Sapatnekar, and Ulya R. Karpuzcu. Energyefficient and reliable inference in nonvolatile memory under extreme operating conditions. ACM Transactions on Embedded Computing Systems, 21(5):57:1– 57:??, September 2022. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3520130.

Rodrigues:2015:DSE

[RKK15] Rance Rodrigues, Israel Koren, and Sandip Kundu. Does the sharing of execution units improve performance/power of multicores? ACM Transactions on Embedded Computing Systems, 14(1):17:1–17:??, January 2015. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Rashid:2024:TNV

[RKM24] Hasib-Al Rashid, Utteja Kallakuri, and Tinoosh Mohsenin. TinyM 2 Net-V2: a compact low-power software hardware architecture for multimodal deep neural networks. ACM Transactions on Embedded Computing Systems, 23(3):47:1– 47:??, May 2024. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3595633.

Romaszkan:2020:PPP

[RLG20] Wojciech Romaszkan, Tianmu Li, and Puneet Gupta. 3PXNet: Pruned-permuted-packed XNOR networks for edge machine learning. ACM Transactions on Embedded Computing Systems, 19(1): 5:1–5:23, February 2020. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3371157.

Ren:2023:PWB

[RLL^{+}23] Jiankang Ren, Chunxiao Liu, Chi Lin, Ran Bi, Simeng Li, Zheng Wang, Yicheng Qian, Zhichao Zhao, and Guozhen Tan. Protection window based security-aware scheduling against schedulebased attacks. ACM Transactions on Embedded Computing Systems, 22(5s): 103:1–103:??, October 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3609098.

Rodionova:2023:TRT

[RLMP23] Alena Rodionova, Lars Lindemann, Manfred Morari, and George Pappas. Temporal robustness of temporal logic specifications: Analysis and control design. ACM Transactions on Embedded Computing Systems, 22(1):13:1–13:??, January 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3550072.

RibeiroDaSilva:2021:MCH

[RLP^{+}21] Junio Cezar Ribeiro Da Silva, Lorena Leão, Vinicius Petrucci, Abdoulaye Gamatié, and Fernando Magno Quintão Pereira. Mapping computations in heterogeneous multicore systems with statistical regression on program inputs. ACM Transactions on Embedded Computing Systems, 20 (6):112:1–112:35, November 2021. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3478288.

Ramaprasad:2010:TBF

[RM10] Harini Ramaprasad and Frank Mueller. Tightening the bounds on feasible preemptions. ACM Transactions on Embedded Computing Systems, 10(2):27:1– 27:??, December 2010. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Reissmann:2020:RIR

[RMBS20] Nico Reissmann, Jan Christian Meyer, Helge Bahmann, and Magnus Själander. RVSDG: an intermediate representation for optimizing compilers. ACM Transactions on Embedded Computing Systems, 19(6):49:1– 49:28, November 2020. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3391902.

Reshadi:2009:HCS

[RMD09] Mehrdad Reshadi, Prabhat Mishra, and Nikil Dutt. Hybrid-compiled simulation: an efficient technique for instruction-set architecture simulation. ACM Transactions on Embedded Computing Systems, 8(3):20:1– 20:??, April 2009. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Reyhani-Masoleh:2004:EDS

[RMH04a] Arash Reyhani-Masoleh and M. Anwar Hasan. Efficient digit-serial normal basis multipliers over binary extension fields. ACM Transactions on Embedded Computing Systems, 3(3):575–592, August 2004. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Reyhani-Masoleh:2004:TFT

[RMH04b] Arash Reyhani-Masoleh and M. Anwar Hasan. Towards fault-tolerant cryptographic computations over finite fields. ACM Transactions on Embedded Computing Systems, 3(3):593–613, August 2004. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Rouhani:2017:RAF

[RMK17] Bita Darvish Rouhani, Azalia Mirhoseini, and Farinaz Koushanfar. RISE: an automated framework for realtime intelligent video surveillance on FPGA. ACM Transactions on Embedded Computing Systems, 16(5s): 158:1–158:??, October 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Rusu:2003:MRR

[RMM03] Cosmin Rusu, Rami Melhem, and Daniel Mossé. Maximizing rewards for realtime applications with energy constraints. ACM Transactions on Embedded Computing Systems, 2(4): 537–559, November 2003. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Raravi:2014:TAA

[RN14] Gurulingesh Raravi and Vincent Nélis. Task assignment algorithms for heterogeneous multiprocessors. ACM Transactions on Embedded Computing Systems, 13(5s): 159:1–159:??, September 2014. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Rajib:2018:PRI

[RN18] MD. Majharul Islam Rajib and Asis Nasipuri. Predictive retransmissions for intermittently connected sensor networks with transmission diversity. ACM Transactions on Embedded Computing Systems, 17(1): 12:1–12:??, January 2018. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Rabbah:2003:DRD

[RP03] Rodric M. Rabbah and Krishna V. Palem. Data remapping for design space optimization of embedded memory systems. ACM Transactions on Embedded Computing Systems, 2(2):186– 218, May 2003. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Reddy:2010:CPE

[RP10] Rakesh Reddy and Peter Petrov. Cache partitioning for energy-efficient and interference-free embedded multitasking. ACM Transactions on Embedded Computing Systems, 9(3):16:1– 16:??, February 2010. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Ragel:2011:HHS

[RP11] Roshan G. Ragel and Sri Parameswaran. A hybrid hardware–software technique to improve reliability in embedded processors. ACM Transactions on Embedded Computing Systems, 10(3):36:1–36:??, April 2011. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Restuccia:2019:YBA

[RPB^{+}19] Francesco Restuccia, Marco Pagani, Alessandro Biondi, Mauro Marinoni, and Giorgio Buttazzo. Is your bus arbiter really fair? Restoring fairness in AXI interconnects for FPGA SoCs. ACM Transactions on Embedded Computing Systems, 18(5s):51:1–51:??, October 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/ft_gateway.
cfm?id=3358183.

Reynolds:2019:MME

[RPHA19] Thomas N. Reynolds, Adam Procter, William L. Harrison, and Gerard Allwein.

The mechanized marriage of effects and monads with applications to high-assurance hardware. ACM Transactions on Embedded Computing Systems, 18(1):6:1– 6:??, February 2019. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3274282.

Raha:2017:QIA

[RR17] Arnab Raha and Vijay Raghunathan. qLUT: Inputaware quantized table lookup for energy-efficient approximate accelerators. ACM Transactions on Embedded Computing Systems, 16(5s): 130:1–130:??, October 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Raj:2022:RMV

[RRC22] Pani Prithvi Raj, Pakala Akhil Reddy, and Nitin Chandrachoodan. Reduced memory Viterbi decoding for hardware-accelerated speech recognition. ACM Transactions on Embedded Computing Systems, 21(3):31:1– 31:18, May 2022. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3510028.

Ravi:2004:SES

[RRKH04] Srivaths Ravi, Anand Raghunathan, Paul Kocher, and Sunil Hattangady. Security in embedded systems: Design challenges. ACM Transactions on Embedded Computing Systems, 3(3):461– 491, August 2004. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Rho:2016:GEC

[RRM16] Seungmin Rho, Wenny Rahayu, and Geyong Min. Guest editorial: Challenges of embedded systems as they evolve into M2M, Internet of Things. ACM Transactions on Embedded Computing Systems, 15(2):34:1– 34:??, May 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Rrushi:2022:PDP

[Rru22] Julian L. Rrushi. Physicsdriven page fault handling for customized deception against CPS malware. ACM Transactions on Embedded Computing Systems, 21(3): 23:1–23:36, May 2022. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3502742.

Regehr:2005:ESO

[RRW05] John Regehr, Alastair Reid, and Kirk Webb. Eliminating stack overflow by abstract interpretation. ACM Transactions on Embedded Computing Systems, 4(4): 751–778, November 2005. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Ratschan:2007:SVH

[RS07] Stefan Ratschan and Zhikun She. Safety verification of hybrid systems by constraint propagation-based abstraction refinement. ACM Transactions on Embedded Computing Systems, 6(1):8:1– 8:??, February 2007. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Riccobene:2009:SCB

[RSB^{+}09] Elvinia Riccobene, Patrizia Scandurra, Sara Bocchio, Alberto Rosti, Luigi Lavazza, and Luigi Mantellini. SystemC/C-based model-driven design for embedded systems. ACM Transactions on Embedded Computing Systems, 8(4):30:1–30:??, July 2009. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Reghenzani:2020:DUP

[RSF20] Federico Reghenzani, Luca Santinelli, and William Fornaciari. Dealing with uncertainty in pWCET estimations. ACM Transactions on Embedded Computing Systems, 19(5):33:1– 33:23, November 2020. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3396234.

Riazi:2017:CSC

[RSK17] M. Sadegh Riazi, Mohammad Samragh, and Farinaz Koushanfar. CAMsure: Secure content-addressable memory for approximate search. ACM Transactions on Embedded Computing Systems, 16(5s):136:1– 136:??, October 2017. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Rottleuthner:2021:SYP

[RSW21] Michel Rottleuthner, Thomas C. Schmidt, and Matthias Wählisch. Sense your power: The ECO approach to energy awareness for IoT devices. ACM Transactions on Embedded Computing Systems, 20(3):24:1–24:25, April 2021. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3441643.

Rakhmatov:2003:EMB

[RV03] Daler Rakhmatov and Sarma Vrudhula. Energy management for battery-powered embedded systems. ACM Transactions on Embedded Computing Systems, 2(3): 277–324, August 2003. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Rao:2007:EOS

[RV07] Ravishankar Rao and Sarma Vrudhula. Energy optimal speed control of a producer– consumer device pair. ACM Transactions on Embedded Computing Systems, 6(4): 30:1–30:??, September 2007. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Real:2018:ADS

[RWL^{+}18] Maria Méndez Real, Philipp Wehner, Vianney Lapotre, Diana Göhringer, and Guy Gogniat. Application deployment strategies for spatial isolation on many-core accelerators. ACM Transactions on Embedded Computing Systems, 17(2):55:1– 55:??, April 2018. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Sababha:2018:RBF

[SA18] Belal H. Sababha and Yazan A. Alqudah. A reconfiguration-based faulttolerant anti-lock brake-by-wire system. ACM Transactions on Embedded Computing Systems, 17(5):87:1– 87:??, November 2018. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

tronic). URL https://dl.

acm.org/ft_gateway.cfm?
id=3242178.

Servais:2021:ACR

[SA21] Jason Servais and Ehsan Atoofian. Adaptive computation reuse for energyefficient training of deep neural networks. ACM Transactions on Embedded Computing Systems, 20(6):114:1– 114:24, November 2021. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3487025.

Schmitz:2004:ISO

[SAHE04] Marcus T. Schmitz, Bashir M. Al-Hashimi, and Petru Eles. Iterative schedule optimization for voltage scalable distributed embedded systems. ACM Transactions on Embedded Computing Systems, 3(1):182–217, February 2004. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Seo:2020:MMP

[SAKH20] Hwajeong Seo, Kyuhwang An, Hyeokdong Kwon, and Zhi Hu. Montgomery multiplication for public key cryptography on MSP430X. ACM Transactions on Embedded Computing Systems, 19(3):20:1–20:15, July 2020. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3387919.

Seth:2006:FFA

[SAMR06] Kiran Seth, Aravindh Anantaraman, Frank Mueller, and Eric Rotenberg. FAST: Frequency-Aware Static Timing analysis. ACM Transactions on Embedded Computing Systems, 5(1):200– 224, February 2006. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Szeto:2023:BAB

[SAS^{+}23] Matthew Szeto, Edward Andert, Aviral Shrivastava, Martin Reisslein, Chung-Wei Lin, and Christ Richmond. B-AWARE: Blockage aware RSU scheduling for 5G enabled autonomous vehicles. ACM Transactions on Embedded Computing Systems, 22(5s):154:1–154:??, October 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3609133.

Suresh:2009:EEE

[SAYN09] Dinesh C. Suresh, Banit Agrawal, Jun Yang, and Walid Najjar. Energyefficient encoding techniques for off-chip data buses. ACM Transactions on Embedded Computing Systems, 8(2): 9:1–9:??, January 2009. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

So:2008:UHS

[SB08] Hayden Kwok-Hay So and Robert Brodersen. A unified hardware/software runtime environment for FPGAbased reconfigurable computers using BORPH. ACM Transactions on Embedded Computing Systems, 7(2): 14:1–14:??, February 2008. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Schneider:2023:CCM

[SB23] Klaus Schneider and Anoop Bhagyanath. Consistency constraints for mapping dataflow graphs to hybrid dataflow/von Neumann architectures. ACM Transactions on Embedded Computing Systems, 22(5):81:1– 81:??, September 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3607869.

Seyoum:2019:FFO

[SBB19] Biruk B. Seyoum, Alessandro Biondi, and Giorgio C. Buttazzo. FLORA: FLoorplan optimizer for reconfigurable areas in FPGAs. ACM Transactions on Embedded Computing Systems, 18(5s):73:1–73:??, October 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/ft_gateway.
cfm?id=3358202.

Song:2022:DTC

[SBDK22] Shihao Song, Adarsha Balaji, Anup Das, and Nagarajan Kandasamy. Designtechnology co-optimization for NVM-Based neuromorphic processing elements. ACM Transactions on Embedded Computing Systems, 21(6):77:1–77:??, November 2022. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3524068.

Sztipanovits:2005:IES

[SBF^{+}05] Janos Sztipanovits, Gautam Biswas, Ken Frampton, Aniruddha Gokhale, Larry Howard, Gabor Karsai, T. John Koo, Xenofon Koutsoukos, and Douglas C. Schmidt. Introducing embedded software and systems education and advanced learning technology in an engineering curriculum. ACM Transactions on Embedded Computing Systems, 4(3):549–568, August 2005. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Siddhu:2023:DTM

[SBK^{+}23] Lokesh Siddhu, Aritra Bagchi, Rajesh Kedia, Isaar Ahmad, Shailja Pandey, and Preeti Ranjan Panda. Dynamic thermal management of 3D memory through rotating low power states and partial channel closure. ACM Transactions on Embedded Computing Systems, 22 (6):104:1–104:??, November 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3624581.

Singh:2013:MCN

[SBLM13] Anshul Singh, Arindam Basu, Keck-Voon Ling, and Vincent J. Mooney III. Models for characterizing noise based PCMOS circuits. ACM Transactions on Embedded Computing Systems, 13(1s):39:1–39:??, November 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Scheir:2015:ASC

[SBR^{+}15] Marijn Scheir, Josep Balasch, Alfredo Rial, Bart Preneel, and Ingrid Verbauwhede. Anonymous split E-cash-toward mobile anonymous payments. ACM Transactions on Embedded Computing Systems, 14(4): 85:1–85:??, December 2015. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Scholz:2008:MPB

[SBX08] Bernhard Scholz, Bernd Burgstaller, and Jingling Xue. Minimal placement of bank selection instructions for partitioned memory architectures. ACM Transactions on Embedded Computing Systems, 7(2):12:1– 12:??, February 2008. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Swaminathan:2005:PBE

[SC05] Vishnu Swaminathan and Krishnendu Chakrabarty. Pruning-based, energy-optimal, deterministic I/O device scheduling for hard realtime systems. ACM Transactions on Embedded Computing Systems, 4(1):141– 167, February 2005. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Schulze:2017:IIM

[SC17] Christoph Schulze and Rance Cleaveland. Improving invariant mining via static analysis. ACM Transactions on Embedded Computing Systems, 16(5s):167:1– 167:??, October 2017. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Stitt:2020:PAI

[SC20] Greg Stitt and David Campbell. PANDORA: an architecture-independent parallelizing approximationdiscovery framework. ACM Transactions on Embedded Computing Systems, 19(5): 39:1–39:17, November 2020. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3391899.

Struhar:2024:HRO

[SCA^{+}24] Václav Struhár, Silviu S. Craciunas, Mohammad Ashjaei, Moris Behnam, and Alessandro V. Papadopoulos. Hierarchical resource orchestration framework for real-time containers. ACM Transactions on Embedded Computing Systems, 23(1): 4:1–4:??, January 2024. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3592856.

Song:2022:DDB

[SCB^{+}22] Shihao Song, Harry Chong, Adarsha Balaji, Anup Das, James Shackleford, and Nagarajan Kandasamy. DF-Synthesizer: Dataflow-based synthesis of spiking neural networks to neuromorphic hardware. ACM Transactions on Embedded Computing Systems, 21(3):27:1– 27:35, May 2022. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3479156.

Scharfenberger:2012:RIP

[SCF12] Christian Scharfenberger, Samarajiit Chakraborty, and Georg Färber. Robust image processing for an omnidirectional camera-based smart car door. ACM Transactions on Embedded Computing Systems, 11(4):87:1– 87:??, December 2012. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Santinelli:2015:PCP

[SCG15] Luca Santinelli and Liliana Cucu-Grosjean. A probabilistic calculus for probabilistic real-time systems. ACM Transactions on Embedded Computing Systems, 14(3):52:1–52:??, May 2015. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Schepers:2007:GEI

[Sch07] Henk Schepers. Guest editorial: Introduction to the special issue on software and compilers for embedded systems. ACM Transactions on Embedded Computing Systems, 6(2):9:1–9:??, May 2007. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Schlich:2010:MCS

[Sch10]  Bastian Schlich.  Model

checking of software for microcontrollers. ACM Transactions on Embedded Computing Systems, 9(4):36:1– 36:??, March 2010. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Shrivastava:2023:ATS

[SCKD23] Aviral Shrivastava, Jian-Jia Chen, Akash Kumar, and Anup Das. ACM TECS special issue on embedded system security tutorials. ACM Transactions on Embedded Computing Systems, 22(3):41:1–41:??, May 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3594872.

Smeets:2020:ARS

[SCM20] Hugues Smeets, Matteo Ceriotti, and Pedro José Marrón. Adapting recursive sinusoidal software oscillators for low-power fixedpoint processors. ACM Transactions on Embedded Computing Systems, 19(3): 17:1–17:26, July 2020. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3378559.

Shen:2016:UAS

[SCRY16] Jie Shen, Yingjue Cai, Yang Ren, and Xiao Yang. A universal application storage system based on smart card. ACM Transactions on Embedded Computing Systems, 15(4):72:1–72:??, August 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Starke:2016:EDV

[SCS16] Renan Augusto Starke, Andreu Carminati, and Rômulo Silva De Oliveira. Evaluating the design of a VLIW processor for real-time systems. ACM Transactions on Embedded Computing Systems, 15(3):46:1–46:??, July 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Shrivastava:2020:ISIa

[SCZ20a] Aviral Shrivastava, Jian-Jia Chen, and Youtao Zhang. Introduction to the special issue on languages, compilers, tools, and theory of embedded systems: Part 1. ACM Transactions on Embedded Computing Systems, 19(5):30:1–30:3, November 2020. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3417732.

Shrivastava:2020:ISIb

[SCZ20b] Aviral Shrivastava, Jian-Jia Chen, and Youtao Zhang. Introduction to the special issue on languages, compilers, tools, and theory of embedded systems: Part 2. ACM Transactions on Embedded Computing Systems, 19(6):41:1–41:2, November 2020. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3417734.

Schirner:2008:QAS

[SD08] Gunar Schirner and Rainer Dömer. Quantitative analysis of the speed/accuracy trade-off in transaction level modeling. ACM Transactions on Embedded Computing Systems, 8(1):4:1– 4:??, December 2008. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

So:2013:STI

[SD13] Won So and Alexander G. Dean. Software thread integration for instruction-level parallelism. ACM Transactions on Embedded Computing Systems, 13(1):8:1– 8:??, August 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Sun:2017:WHS

[SD17] Youcheng Sun and Marco Di Natale. Weakly hard schedulability analysis for fixed priority scheduling of periodic real-time tasks. ACM Transactions on Embedded Computing Systems, 16(5s): 171:1–171:??, October 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Shoushtari:2018:SIS

[SDBD18] Majid Shoushtari, Bryan Donyanavard, Luis Angel D. Bathen, and Nikil Dutt. ShaVe-ICE: Sharing distributed virtualized SPMs in many-core embedded systems. ACM Transactions on Embedded Computing Systems, 17(2):47:1–47:??, April 2018. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Song:2019:EEP

[SDMK19] Shihao Song, Anup Das, Onur Mutlu, and Nagarajan Kandasamy. Enabling and exploiting partition-level parallelism (PALP) in phase change memories. ACM Transactions on Embedded Computing Systems, 18(5s): 53:1–53:??, October 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/ft_gateway.
cfm?id=3358180.

Staschulat:2007:SPC

[SE07] Jan Staschulat and Rolf Ernst. Scalable precision cache analysis for real-time software. ACM Transactions on Embedded Computing Systems, 6(4):25:1– 25:??, September 2007. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Schliecker:2010:RTP

[SE10] Simon Schliecker and Rolf Ernst. Real-time performance analysis of multiprocessor systems with shared memory. ACM Transactions on Embedded Computing Systems, 10(2):22:1– 22:??, December 2010. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Schlatow:2017:RTA

[SE17] Johannes Schlatow and Rolf Ernst. Response-time analysis for task chains with complex precedence and blocking relations. ACM Transactions on Embedded Computing Systems, 16(5s):172:1– 172:??, October 2017. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Samaddar:2023:ODS

[SE23] Ankita Samaddar and Arvind Easwaran. Online distributed schedule randomization to mitigate timing attacks in industrial control systems. ACM Transactions on Embedded Computing Systems, 22(6):100:1– 100:??, November 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3624584.

Shokry:2012:HSS

[SEB12] Hesham Shokry and Hatem M.

El-Boghdadi. On heuristic solutions to the simple offset assignment problem in address-code optimization. ACM Transactions on Embedded Computing Systems, 11(3):63:1–63:??, September 2012. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Seo:2018:CSI

[Seo18] Hwajeong Seo. Compact software implementation of public-key cryptography on MSP430X. ACM Transactions on Embedded Computing Systems, 17(3):66:1– 66:??, June 2018. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Seviora:2005:CES

[Sev05] Rudolph E. Seviora. A curriculum for embedded system engineering. ACM Transactions on Embedded Computing Systems, 4(3): 569–586, August 2005. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Saberi:2023:POT

[SFB23] Iman Saberi, Fathiyeh Faghih, and Farzad Sobhi Bavil. A passive online technique for learning hybrid automata from Input/ output traces. ACM Transactions on Embedded Computing Systems, 22(1):9:1– 9:??, January 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3556543.

Shu:2023:TID

[SFCW23] Jiwu Shu, Kedong Fang, Youmin Chen, and Shuo Wang. TH-iSSD: Design and implementation of a generic and reconfigurable near-data processing framework. ACM Transactions on Embedded Computing Systems, 22(6): 96:1–96:??, November 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3563456.

Sui:2018:LOP

[SFZX18] Yulei Sui, Xiaokang Fan, Hao Zhou, and Jingling Xue. Loop-oriented pointer analysis for automatic SIMD vectorization. ACM Transactions on Embedded Computing Systems, 17(2):56:1– 56:??, April 2018. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Sah:2024:ATE

[SG24] Ramesh Kumar Sah and Hassan Ghasemzadeh. Adversarial transferability in embedded sensor systems: an activity recognition perspective. ACM Transactions on Embedded Computing Systems, 23(2):20:1– 20:??, March 2024. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3641861.

Subramanian:2012:GOP

[SGDP12] Varun Subramanian, Michael Gilberti, Alex Doboli, and Dan Pescaru. A goaloriented programming framework for grid sensor networks with reconfigurable embedded nodes. ACM Transactions on Embedded Computing Systems, 11(4):79:1– 79:??, December 2012. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Sogokon:2017:OMP

[SGJ17] Andrew Sogokon, Khalil Ghorbal, and Taylor T. Johnson. Operational models for piecewise-smooth systems. ACM Transactions on Embedded Computing Systems, 16(5s):185:1–185:??, October 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Sanz:2013:SLM

[SGT^{+}13] Concepción Sanz, José Ignacio Gómez, Christian Tenllado, Manuel Prieto, and Francky Catthoor. Systemlevel memory management based on statistical variability compensation for framebased applications. ACM Transactions on Embedded Computing Systems, 13(1s): 35:1–35:??, November 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Sun:2016:FFJ

[SGW^{+}16] Jinghao Sun, Nan Guan, Yang Wang, Qingxu Deng, Peng Zeng, and Wang Yi. Feasibility of fork-join realtime task graph models: Hardness and algorithms. ACM Transactions on Embedded Computing Systems, 15(1):14:1–14:??, February 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Sepulveda:2021:BCA

[SGZS21] Johanna Sepúlveda, Mathieu Gross, Andreas Zankl, and Georg Sigl. Beyond cache attacks: Exploiting the bus-based communication structure for powerful on-chip microarchitectural attacks. ACM Transactions on Embedded Computing Systems, 20(2):17:1– 17:23, March 2021. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3433653.

Siirtola:2015:PMI

[SH15] Antti Siirtola and Keijo Heljanko. Parametrised modal interface automata. ACM Transactions on Embedded Computing Systems, 14(4): 65:1–65:??, December 2015. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Sifakis:2023:TAS

[SH23] Joseph Sifakis and David Harel. Trustworthy autonomous system development. ACM Transactions on Embedded Computing Systems, 22(3):40:1–40:??, May 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3545178.

Sun:2019:STC

[SHK^{+}19] Youcheng Sun, Xiaowei Huang, Daniel Kroening, James Sharp, Matthew Hill, and Rob Ashmore. Structural test coverage criteria for deep neural networks. ACM Transactions on Embedded Computing Systems, 18(5s):94:1–94:??, October 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/ft_gateway.
cfm?id=3358233.

Shen:2017:MQC

[SHL^{+}17] Zhaoyan Shen, Zhijian He, Shuai Li, Qixin Wang, and Zili Shao. A multiquadcopter cooperative cyberphysical system for timely air pollution localization. ACM Transactions on Embedded Computing Systems, 16(3): 70:1–70:??, July 2017. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Schmoll:2013:IFR

[SHME13] Florian Schmoll, Andreas Heinig, Peter Marwedel, and Michael Engel. Improving the fault resilience of an H.264 decoder using static analysis methods. ACM Transactions on Embedded Computing Systems, 13(1s): 31:1–31:??, November 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Sun:2019:DOS

[SHQX19] Hui Sun, Jianzhong Huang, Xiao Qin, and Changsheng Xie. DLSpace: Optimizing SSD lifetime via an efficient distributed log space allocation strategy. ACM Transactions on Embedded Computing Systems, 17(6):92:1– 92:??, January 2019. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3284749.

Shukla:2014:EEE

[Shu14a] Sandeep K. Shukla. Editorial: Embedded everywhere for everyone. ACM Transactions on Embedded Computing Systems, 13(4): 74:1–74:??, February 2014. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Shukla:2014:EES

[Shu14b] Sandeep K. Shukla. Editorial: Embedded systems — more than methodology. ACM Transactions on Embedded Computing Systems, 13(3s):99:1–99:??, March 2014. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Shukla:2015:EBD

[Shu15a] Sandeep K. Shukla. Editorial: Big data, Internet of Things, cybersecurity — a new trinity of embedded systems research. ACM Transactions on Embedded Computing Systems, 14(4):61:1–61:??, December 2015. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Shukla:2015:EOS

[Shu15b] Sandeep K. Shukla. Editorial: Oh security — where art thou? ACM Transactions on Embedded Computing Systems, 14(2):20:1– 20:??, March 2015. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Shukla:2015:ERS

[Shu15c] Sandeep K. Shukla. Editorial: Regular, special, and related issues. ACM Transactions on Embedded Computing Systems, 14(1):1:1– 1:??, January 2015. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Shukla:2015:ESD

[Shu15d] Sandeep K. Shukla. Editorial: Schizoid design for critical embedded systems. ACM Transactions on Embedded Computing Systems, 14(3): 40e:1–40e:??, May 2015. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Shukla:2016:EDP

[Shu16a] Sandeep K. Shukla. Editorial: Distributed public ledgers and block chains — what good are they for embedded systems? ACM Transactions on Embedded Computing Systems, 16(1): 1:1–1:2, November 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Shukla:2016:EFI

[Shu16b] Sandeep K. Shukla. Editorial: Fence itself grazing the field — security from the sentries. ACM Transactions on Embedded Computing Systems, 15(3):41:1–41:??, July 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Shukla:2016:ESB

[Shu16c] Sandeep K. Shukla. Editorial: Science of the big and small and embedded computing systems. ACM Transactions on Embedded Computing Systems, 15(2): 21:1–21:??, May 2016. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Shukla:2016:ESE

[Shu16d] Sandeep K. Shukla. Editorial: Security of embedded systems and cyber irons — embedded systems for security. ACM Transactions on Embedded Computing Systems, 15(4):62:1– 62:??, August 2016. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Shukla:2017:ECC

[Shu17a] Sandeep K. Shukla. Editorial: Continuing the course. ACM Transactions on Embedded Computing Systems, 16(2):28:1–28:??, April 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Shukla:2017:ECS

[Shu17b] Sandeep K. Shukla. Editorial: Cyber security, IoT, block chains-risks and opportunities. ACM Transactions on Embedded Computing Systems, 16(3):62:1– 62:??, July 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Shukla:2017:ESM

[Shu17c] Sandeep K. Shukla. Editorial: Security of mobile devices. ACM Transactions on Embedded Computing Systems, 16(4):91:1– 91:??, September 2017. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Shukla:2018:EEC

[Shu18a] Sandeep K. Shukla. Editorial: Early career researchers in embedded computing. ACM Transactions on Embedded Computing Systems, 17(4):73:1– 73:??, August 2018. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Shukla:2018:EIC

[Shu18b] Sandeep K. Shukla. Editorial: Industry 4.0 — a confluence of embedded artificial intelligence, machine learning, robotics and security. ACM Transactions on Embedded Computing Systems, 17(2):29:1–29:??, April 2018. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Shukla:2018:ENA

[Shu18c] Sandeep K. Shukla. Editorial: Need for artifact verified articles in ACM Transactions. ACM Transactions on Embedded Computing Systems, 17(5):82:1– 82:??, November 2018. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3282437.

Shukla:2018:EUE

[Shu18d] Sandeep K. Shukla. Editorial: To use or not to? Embedded systems for voting. ACM Transactions on Embedded Computing Systems, 17(3):58:1–58:??, June 2018. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Shukla:2018:ETS

[Shu18e] Sandeep K. Shukla. Editorial: Trust and security must become a primary design concern in embedded computing. ACM Transactions on Embedded Computing Systems, 17(1):1:1– 1:??, January 2018. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Shukla:2019:EAR

[Shu19a] Sandeep K. Shukla. Editorial: Adversaries and robustness. ACM Transactions on Embedded Computing Systems, 18(4):30:1– 30:??, August 2019. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3345556.

Shukla:2019:EES

[Shu19b] Sandeep K. Shukla. Editorial: Embedded security challenge: Cyber security contests in the embedded computing domain. ACM Transactions on Embedded Computing Systems, 17(6): 91:1–91:??, January 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Shukla:2019:EHF

[Shu19c] Sandeep K. Shukla. Editorial: Human factors in embedded computing. ACM Transactions on Embedded Computing Systems, 18(1): 1:1–1:??, February 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/ft_gateway.
cfm?id=3302888.

Shukla:2019:ERH

[Shu19d] Sandeep K. Shukla. Editorial: Reflections on the history of cyber-physical versus embedded systems. ACM Transactions on Embedded Computing Systems, 18(3):

19:1–19:??, June 2019. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3325115.

Shukla:2020:EEC

[Shu20a] Sandeep K. Shukla. Editorial: Embedded computing and society. ACM Transactions on Embedded Computing Systems, 18(6):1–3, January 2020. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL

https://dl.acm.org/doi/
abs/10.1145/3368250.

Shukla:2020:TER

[Shu20b] Sandeep K. Shukla. TECS editorial: Rethinking and reevaluating in the time of crisis. ACM Transactions on Embedded Computing Systems, 19(3):16e:1–16e:3, July 2020. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3395923.

Sridhar:2019:SEC

[SIC19] Aditya Sridhar, Mohamed Ibrahim, and Krishnendu Chakrabarty. Synterface: Efficient chip-to-world interfacing for flow-based microfluidic biochips using pin-count minimization. ACM Transactions on Embedded Computing Systems, 18(5s):54:1– 54:??, October 2019. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

                                        tronic). URL https://dl.
                                        acm.org/ft_gateway.cfm?
                                        id=3358188.

Santanna:2017:DIS

[SIR^{+}17] Francisco Sant’anna, Roberto Ierusalimschy, Noemi Rodriguez, Silvana Rossetto, and Adriano Branco. The design and implementation of the synchronous language CéU. ACM Transactions on Embedded Computing Systems, 16(4):98:1– 98:26, August 2017. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Singh:2024:AEF

[SISS24] Richa Singh, Saad Islam, Berk Sunar, and Patrick Schaumont. Analysis of EM fault injection on bit-sliced number theoretic transform software in Dilithium. ACM Transactions on Embedded Computing Systems, 23(2): 32:1–32:??, March 2024. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3583757.

Shim:2003:LEC

[SJC^{+}03] Hojun Shim, Yongsoo Joo, Yongseok Choi, Hyung Gyu Lee, and Naehyuck Chang. Low-energy off-chip SDRAM memory systems for embedded applications. ACM Transactions on Embedded Computing Systems, 2(1): 98–130, February 2003. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Samragh:2020:ERB

[SJK20] Mohammad Samragh, Mojan Javaheripi, and Farinaz Koushanfar. EncoDeep: Realizing bit-flexible encoding for deep neural networks. ACM Transactions on Embedded Computing Systems, 19(6):43:1–43:29, November 2020. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3391901.

Seo:2018:CIA

[SJLK18] Hwajeong Seo, Ilwoong Jeong, Jungkeun Lee, and Woo-Hwan Kim. Compact implementations of ARXbased block ciphers on IoT processors. ACM Transactions on Embedded Computing Systems, 17(3):60:1– 60:??, June 2018. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Shin:2022:EED

[SJOL22] Dongsuk Shin, Hakbeom Jang, Kiseok Oh, and Jae W. Lee. An energyefficient DRAM cache architecture for mobile platforms with PCM-based main memory. ACM Transactions on Embedded Computing Systems, 21(1):7:1– 7:22, January 2022. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3451995.

Sem-Jacobsen:2013:ELC

[SJRS^{+}13a] Frank Olaf Sem-Jacobsen, Samuel Rodrigo, Tor Skeie, Alessandro Strano, and Davide Bertozzi. An efficient, low-cost routing framework for convex mesh partitions to support virtualization. ACM Transactions on Embedded Computing Systems, 12(4): 107:1–107:??, June 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Sem-Jacobsen:2013:EPE

[SJRS^{+}13b] Frank Olaf Sem-Jacobsen, Samuel Rodrigo, Alessandro Strano, Tor Skeie, Davide Bertozzi, and Francisco Gilabert. Enabling power efficiency through dynamic rerouting on-chip. ACM Transactions on Embedded Computing Systems, 12(4): 111:1–111:??, June 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Sartori:2013:ETE

[SK13] John Sartori and Rakesh Kumar. Exploiting timing error resilience in processor architecture. ACM Transactions on Embedded Computing Systems, 12(2s): 89:1–89:??, May 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Seo:2019:ETM

[SK19] Minjun Seo and Fadi Kurdahi. Efficient tracing methodology using automata processor. ACM Transactions on Embedded Computing Systems, 18(5s):80:1– 80:??, October 2019. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3358200.

Seo:2012:RGV

[SKH^{+}12] Suk-Hyun Seo, Jin-Ho Kim, Sung-Ho Hwang, Key Ho Kwon, and Jae Wook Jeon. A reliable gateway for invehicle networks based on LIN, CAN, and FlexRay. ACM Transactions on Embedded Computing Systems, 11(1):7:1–7:??, March 2012. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Song:2014:POA

[SKK^{+}14] Wook Song, Yeseong Kim, Hakbong Kim, Jehun Lim, and Jihong Kim. Personalized optimization for Android smartphones. ACM Transactions on Embedded Computing Systems, 13(2s): 60:1–60:??, January 2014. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Schoeberl:2011:HAL

[SKKR11] Martin Schoeberl, Stephan Korsholm, Tomas Kalibera, and Anders P. Ravn. A hardware abstraction layer in Java. ACM Transactions on Embedded Computing Systems, 10(4):42:1– 42:??, November 2011. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Semenova:2025:CSS

[SKL^{+}25] Sofiya Semenova, Steven Ko, Yu David Liu, Lukasz Ziarek, and Karthik Dantu. A comprehensive study of systems challenges in visual simultaneous localization and mapping systems. ACM Transactions on Embedded Computing Systems, 24(1):2:1– 2:??, January 2025. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3677317.

Shresthamali:2017:APM

[SKN17] Shaswot Shresthamali, Masaaki Kondo, and Hiroshi Nakamura. Adaptive power management in solar energy harvesting sensor node using reinforcement learning.

ACM Transactions on Embedded Computing Systems, 16(5s):181:1–181:??, October 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Seo:2010:DAS

[SKPL10] Euiseong Seo, Sangwon Kim, Seonyeong Park, and Joonwon Lee. Dynamic alteration schemes of realtime schedules for I/O device energy efficiency. ACM Transactions on Embedded Computing Systems, 10(2): 23:1–23:??, December 2010. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Saini:2021:IFC

[SKS21] Kanika Saini, Sheetal Kalra, and Sandeep K. Sood. IoTfog-cloud centric earthquake monitoring and prediction. ACM Transactions on Embedded Computing Systems, 20(6):115:1–115:26, November 2021. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL

https://dl.acm.org/doi/
10.1145/3487942.

Scharwaechter:2007:AAE

[SKW^{+}07] Hanno Scharwaechter, David Kammler, Andreas Wieferink, Manuel Hohenauer, Kingshuk Karuri, Jianjiang Ceng, Rainer Leupers, Gerd Ascheid, and Heinrich Meyr. ASIP architecture exploration for efficient IPSec encryption: a case study. ACM Transactions on Embedded Computing Systems, 6(2): 12:1–12:??, May 2007. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Serpanos:2004:GES

[SL04] Dimitrios N. Serpanos and Haris Lekatsas. Guest editorial: Special issue on embedded systems and security. ACM Transactions on Embedded Computing Systems, 3(3):459–460, August 2004. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Shin:2008:CRT

[SL08] Insik Shin and Insup Lee. Compositional real-time scheduling framework with periodic model. ACM Transactions on Embedded Computing Systems, 7(3):30:1– 30:??, April 2008. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Sauer:2016:LFD

[SL16] Christian Sauer and Hans-Peter Loeb. A lightweight framework for the dynamic creation and configuration of virtual platforms in SystemC. ACM Transactions on Embedded Computing Systems, 16(1):5:1–5:??, November 2016. CODEN ????

ISSN 1539-9087 (print), 1558-3465 (electronic).

Sayyah:2015:VPB

[SLB^{+}15] Parinaz Sayyah, Mihai T.

Lazarescu, Sara Bocchio, Emad Ebeid, Gianluca Palermo, Davide Quaglia, Alberto Rosti, and Luciano Lavagno. Virtual platformbased design space exploration of power-efficient distributed embedded applications. ACM Transactions on Embedded Computing Systems, 14(3):49:1–49:??, May 2015. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Spasic:2016:IHR

[SLCS16] Jelena Spasic, Di Liu, Emanuele Cannella, and Todor Stefanov. On the improved hard real-time scheduling of cyclo-static dataflow. ACM Transactions on Embedded Computing Systems, 15(4):68:1– 68:??, August 2016. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Stilkerich:2017:PGU

[SLE^{+}17] Isabella Stilkerich, Clemens Lang, Christoph Erhardt, Christian Bay, and Michael Stilkerich. The perfect getaway: Using escape analysis in embedded real-time systems. ACM Transactions on Embedded Computing Systems, 16(4):99:1– 99:30, August 2017. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Spellini:2019:CDM

[SLFC19] Stefano Spellini, Michele Lora, Franco Fummi, and Sudipta Chattopadhyay. Compositional design of multirobot systems control software on ROS. ACM Transactions on Embedded Computing Systems, 18(5s):71:1– 71:??, October 2019. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3358197.

Soliman:2022:FFF

[SLK^{+}22] Taha Soliman, Nellie Laleni, Tobias Kirchner, Franz Müller, Ashish Shrivastava, Thomas Kämpfe, Andre Guntoro, and Norbert Wehn. FELIX: a ferroelectric FET based low power mixedsignal in-memory architecture for DNN acceleration. ACM Transactions on Embedded Computing Systems, 21(6):84:1–84:??, November 2022. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3529760.

Seo:2016:HMR

[SLN^{+}16] Hwajeong Seo, Zhe Liu, Yasuyuki Nogami, Jongseok Choi, and Howon Kim. Hybrid Montgomery reduction. ACM Transactions on Embedded Computing Systems, 15(3):58:1–58:??, July 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Su:2019:TFR

[SLS^{+}19] Fang Su, Yongpan Liu, Xiao Sheng, Hyung Gyu Lee, Naehyuck Chang, and Huazhong Yang. A task failure rate aware dual-channel solar power system for nonvolatile sensor nodes. ACM Transactions on Embedded Computing Systems, 18(4):33:1– 33:??, August 2019. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3320270.

Sun:2024:ACA

[SLZ^{+}24] Hui Sun, Bendong Lou, Chao Zhao, Deyan Kong, Chaowei Zhang, Jianzhong Huang, Yinliang Yue, and Xiao Qin. Asynchronous compaction acceleration scheme for near-data processingenabled LSM-tree-based KV stores. ACM Transactions on Embedded Computing Systems, 23(6):93:1– 93:??, November 2024. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3626097.

Stamoulias:2013:PAK

[SM13a] Ioannis Stamoulias and Elias S. Manolakos. Parallel architectures for the kNN classifier — design of soft IP cores and FPGA implementations. ACM Transactions on Embedded Computing Systems, 13(2):22:1– 22:??, September 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Stanley-Marbell:2013:LPP

[SM13b] Phillip Stanley-Marbell. L24: Parallelism, performance, energy efficiency, and cost trade-offs in future sensor platforms. ACM Transactions on Embedded Computing Systems, 13(1):7:1– 7:??, August 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Serpanos:2004:EHS

[SMG04] Dimitrios N. Serpanos, Poluxeni Mountrouidou, and Maria Gamvrili. Evaluation of hardware and software schedulers for embedded switches. ACM Transactions on Embedded Computing Systems, 3(4):736– 759, November 2004. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Sarkar:2015:STP

[SMR15] Abhik Sarkar, Frank Mueller, and Harini Ramaprasad. Static task partitioning for locked caches in multicore real-time systems. ACM Transactions on Embedded Computing Systems, 14(1): 4:1–4:??, January 2015. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Smyth:2018:SSC

[SMR^{+}18] Steven Smyth, Christian Motika, Karsten Rathlev, Reinhard Von Hanxleden, and Michael Mendler. SCEst: Sequentially constructive Esterel. ACM Transactions on Embedded Computing Systems, 17(2): 33:1–33:??, April 2018. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Sood:2020:RDV

[SMR20] Surinder Sood, Avinash Malik, and Partha Roop. Robust design and validation of cyber-physical systems. ACM Transactions on Embedded Computing Systems, 18(6):1–21, January 2020. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3362098.

Sandoval:2017:TTS

[SMW^{+}17] Nathan Sandoval, Casey Mackin, Sean Whitsitt, Vijay Shankar Gopinath, Sachidanand Mahadevan, Andrew Milakovich, Kyle Merry, Jonathan Sprinkle, and Roman Lysecky. Task transition scheduling for data-adaptable systems. ACM Transactions on Embedded Computing Systems, 16(4):105:1–105:??, August 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Shi:2021:TGH

[SMZ^{+}21] Zhendong Shi, Haocheng Ma, Qizhi Zhang, Yanjiang Liu, Yiqiang Zhao, and Jiaji He. Test generation for hardware Trojan detection using correlation analysis and genetic algorithm. ACM Transactions on Embedded Computing Systems, 20(4):28:1–28:20, June 2021. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3446837.

Singh:2010:CPD

[SN10] Montek Singh and Steven M. Nowick. Call for papers: Deadline: March 15, 2011. ACM Transactions on Embedded Computing Systems, 10(2):29:1–29:??, December 2010. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Sahin:2025:HAE

[SNTÖ25] Suhap Sahin, Oguz Narli, Muhammet Bahadir Türkoglu, and Hikmetcan Özcan. Hardware area efficient and realtime FPGA implementation of PHMMRGB. ACM Transactions on Embedded Computing Systems, 24(1):15:1– 15:??, January 2025. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3701727.

Schaumont:2015:IEP

[SOG15] Patrick Schaumont, Maire O’Neill, and Tim Güneysu. Introduction for embedded platforms for cryptography in the coming decade. ACM Transactions on Embedded Computing Systems, 14(3): 40:1–40:??, April 2015. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Schurmans:2016:FAE

[SOL^{+}16] Stefan Schürmans, Gereon Onnebrink, Rainer Leupers, Gerd Ascheid, and Xiaotao Chen. Frequency-aware ESL power estimation for ARM Cortex-A9 using a black box processor model. ACM Transactions on Embedded Computing Systems, 16(1):26:1–26:??, November 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Schoeberl:2010:NRT

[SP10] Martin Schoeberl and Wolfgang Puffitsch. Nonblocking real-time garbage collection. ACM Transactions on Embedded Computing Systems, 10(1):6:1–6:??, August 2010. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Smith:2012:OSH

[SP12] Melissa C. Smith and Gregory D. Peterson. Optimization of shared highperformance reconfigurable computing resources. ACM Transactions on Embedded Computing Systems, 11(2): 36:1–36:??, July 2012. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Sheikh:2019:EEM

[SP19a] Saad Zia Sheikh and Muhammad Adeel Pasha. Energyefficient multicore scheduling for hard real-time systems: a survey. ACM Transactions on Embedded Computing Systems, 17(6):94:1– 94:??, January 2019. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Siddhu:2019:PLA

[SP19b] Lokesh Siddhu and Preeti Ranjan Panda. PredictNcool: Leakage aware thermal management for 3D memories using a lightweight temperature predictor. ACM Transactions on Embedded Computing Systems, 18(5s):64:1– 64:??, October 2019. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3358208.

Sheikh:2020:EER

[SP20] Saad Zia Sheikh and Muhammad Adeel Pasha. Energyefficient real-time scheduling on multicores: a novel approach to model cache contention. ACM Transactions on Embedded Computing Systems, 19(4):28:1– 28:25, July 2020. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3399413.

Singh:2017:EER

[SPB^{+}17] Amit Kumar Singh, Alok Prakash, Karunakar Reddy Basireddy, Geoff V. Merrett, and Bashir M. Al-Hashimi. Energy-efficient run-time mapping and thread partitioning of concurrent OpenCL applications on CPU–GPU MPSoCs. ACM Transactions on Embedded Computing Systems, 16(5s):147:1– 147:??, October 2017. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Sharma:2016:DFT

[SPC^{+}16] Namita Sharma, Preeti Ranjan Panda, Francky Catthoor, Min Li, and Prashant Agrawal. Data flow transformation for energy-efficient implementation of Givens rotation-based QRD. ACM Transactions on Embedded Computing Systems, 15(1): 18:1–18:??, February 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Saarikivi:2017:MTS

[SPDLK^{+}17] Olli Saarikivi, Hernán Ponce-De-León, Kari Kähkönen, Keijo Heljanko, and Javier Esparza. Minimizing test suites with unfoldings of multithreaded programs. ACM Transactions on Embedded Computing Systems, 16(2):45:1–45:??, April 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Salmon:2024:NCA

[SPGB24] Loic Salmon, Pierre-Yves Pillain, Goulven Guillou, and Jean-Philippe Babau. NAVIDRO, a CARES architectural style for configuring drone co-simulation. ACM Transactions on Embedded Computing Systems, 23(3):39:1–39:??, May 2024. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3651889.

Smirnov:2019:IGM

[SPGT19] Fedor Smirnov, Behnaz Pourmohseni, Michael Glaß, and Jürgen Teich. IGOR, get me the optimum! Prioritizing important design decisions during the DSE of embedded systems. ACM Transactions on Embedded Computing Systems, 18(5s): 78:1–78:??, October 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/ft_gateway.
cfm?id=3358204.

Shin:2012:CTC

[SPK^{+}12] Donghwa Shin, Jaehyun Park, Younghyun Kim, Jaeam Seo, and Naehyuck Chang. Control-theoretic cyber-physical system modeling and synthesis: a case study of an active direct methanol fuel cell. ACM Transactions on Embedded Computing Systems, 11(4): 76:1–76:??, December 2012. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Spring:2010:RAI

[SPP^{+}10] Jesper Honig Spring, Filip Pizlo, Jean Privat, Rachid Guerraoui, and Jan Vitek. Reflexes: Abstractions for integrating highly responsive tasks into Java applications. ACM Transactions on Embedded Computing Systems, 10(1):4:1–4:??, August 2010. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Shamsa:2021:UUB

[SPT^{+}21] Elham Shamsa, Alma Pröbstl, Nima TaheriNejad, Anil Kanduri, Samarjit Chakraborty, Amir M. Rahmani, and Pasi Liljeberg. UBAR: User- and battery-aware resource management for smartphones. ACM Transactions on Embedded Computing Systems, 20(3):23:1–23:25, April 2021. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3441644.

Sharma:2023:FCD

[SPT^{+}23] Harsh Sharma, Lukas Pfromm, Rasit Onur Topaloglu, Janardhan Rao Doppa, Umit Y. Ogras, Ananth Kalyanraman, and Partha Pratim Pande. Florets for chiplets: Data flow-aware high-performance and energyefficient network-on-interposer for CNN inference tasks. ACM Transactions on Embedded Computing Systems, 22(5s):132:1–132:??, October 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3608098.

Salamy:2012:SOT

[SR12a] Hassan Salamy and J. Ramanujam. Storage optimization through offset assignment with variable coalescing. ACM Transactions on Embedded Computing Systems, 11S(1):16:1–16:??, 2012. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Seshia:2012:QAS

[SR12b] Sanjit A. Seshia and Alexander Rakhlin. Quantitative analysis of systems using game-theoretic learning. ACM Transactions on Embedded Computing Systems, 11(S2):55:1–55:??, 2012. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Strobel:2019:PMA

[SR19] Manuel Strobel and Martin Radetzki. Power-mode-aware memory subsystem optimization for low-power system-on-chip design. ACM Transactions on Embedded Computing Systems, 18(5): 43:1–43:??, October 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/ft_gateway.
cfm?id=3356583.

Suris:2012:RSC

[SRA12] Jorge A. Surís, Adolfo Recio, and Peter Athanas. RapidRadio: Signal classification and radio deployment framework. ACM Transactions on Embedded Computing Systems, 11(S2): 41:1–41:??, 2012. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Sarwar:2023:CPE

[SRB23] Mir Sarwar, Rajarshi Ray, and Ansuman Banerjee. A contrastive plan explanation framework for hybrid system models. ACM Transactions on Embedded Computing Systems, 22(2):22:1– 22:??, March 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3561532.

Segarra:2015:ASP

[SRG^{+}15] Juan Segarra, Clemente Rodríguez, Rubén Gran, Luis C. Aparicio, and Víctor Viñals. ACDC: Small, predictable and highperformance data cache. ACM Transactions on Embedded Computing Systems, 14(2):38:1–38:??, March 2015. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Sutar:2018:DPI

[SRK^{+}18] Soubhagya Sutar, Arnab Raha, Devadatta Kulkarni, Rajeev Shorey, Jeffrey Tew, and Vijay Raghunathan. D-PUF: an intrinsically reconfigurable DRAM PUF for device authentication and random number generation. ACM Transactions on Embedded Computing Systems, 17(1):17:1–17:??, January 2018. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Seiculescu:2013:DBE

[SRM^{+}13] Ciprian Seiculescu, Dara Rahmati, Srinivasan Murali, Hamid Sarbazi-Azad, Luca Benini, and Giovanni De Micheli. Designing best effort networks-on-chip to meet hard latency constraints. ACM Transactions on Embedded Computing Systems, 12(4):108:1– 108:??, June 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Santini:2016:BCS

[SRNW16] Thiago Santini, Paolo Rech, Gabriel Luca Nazar, and Flávio Rech Wagner. Beyond cross-section: Spatiotemporal reliability analysis. ACM Transactions on Embedded Computing Systems, 15(1):3:1–3:??, February 2016. CODEN ????

ISSN 1539-9087 (print), 1558-3465 (electronic).

Singh:2023:KWC

[SRR^{+}23] Nikhilesh Singh, Karthikeyan Renganathan, Chester Rebeiro, Jithin Jose, and Ralph Mader. Kryptonite: Worstcase program interference estimation on multi-core embedded systems. ACM Transactions on Embedded Computing Systems, 22(5s): 149:1–149:??, October 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3609128.

Schurgers:2003:PME

[SRS03] Curt Schurgers, Vijay Raghunathan, and Mani B. Srivastava. Power management for energy-aware communication systems. ACM Transactions on Embedded Computing Systems, 2(3): 431–447, August 2003. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Schulz-Rosengarten:2021:TOO

[SRSM21] Alexander Schulz-Rosengarten, Steven Smyth, and Michael Mendler. Toward objectoriented modeling in SCCharts. ACM Transactions on Embedded Computing Systems, 20(4):37:1–37:26, June 2021. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3453482.

Seo:2013:AIG

[SRY13] Sanghyun Seo, Seungtaek Ryoo, and Kyunghyun Yoon. Artistic image generation for emerging multimedia services by impressionist manner. ACM Transactions on Embedded Computing Systems, 12(2):22:1–22:??, February 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Seo:2021:SBA

[SSA21] Hwajeong Seo, Pakize Sanal, and Reza Azarderakhsh. SIKE in 32-bit ARM processors based on redundant number system for NIST level-II. ACM Transactions on Embedded Computing Systems, 20(3):19:1– 19:23, April 2021. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3439733.

Saha:2024:TPA

[SSA^{+}24a] Swapnil Sayan Saha, Sandeep Singh Sandha, Mohit Aggarwal, Brian Wang, Liying Han, Julian De Gortari Briseno, and Mani Srivastava. TinyNS: Platform-aware neurosymbolic auto tiny machine learning. ACM Transactions on Embedded Computing Systems, 23(3):43:1– 43:??, May 2024. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3603171.

Shi:2024:ISI

[SSA^{+}24b] Liang Shi, Jingtong Shi, Hussam Amrouch, Kuan-Hsun Chen, Mengying Zhao, and Weichen Liu. Introduction to special issue on in/near memory and storage computing for embedded systems. ACM Transactions on Embedded Computing Systems, 23(6):85:1– 85:??, November 2024. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3677018.

Shin:2024:VEM

[SSB24] Yong-Jun Shin, Donghwan Shin, and Doo-Hwan Bae. Virtual environment model generation for CPS goal verification using imitation learning. ACM Transactions on Embedded Computing Systems, 23(1):13:1– 13:??, January 2024. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3633804.

Salamati:2019:MEM

[SSD^{+}19] Mahmoud Salamati, Rocco Salvia, Eva Darulova, Sadegh Soudjani, and Rupak Majumdar. Memory-efficient mixed-precision implementations for robust explicit model predictive control. ACM Transactions on Embedded Computing Systems, 18(5s):100:1–100:??, October 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/ft_gateway.
cfm?id=3358223.

Siozios:2014:FSA

[SSH14] Kostas Siozios, Dimitrios Soudris, and Michael Hübner. A framework for supporting adaptive fault-tolerant solutions. ACM Transactions on Embedded Computing Systems, 13(5s):169:1– 169:??, November 2014. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Saha:2021:MWR

[SSK21] Debasri Saha and Susmita Sur-Kolay. Minimization of WCRT with recovery assurance from hardware Trojans for tasks on FPGAbased cloud. ACM Transactions on Embedded Computing Systems, 20(1):1:1– 1:25, January 2021. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3409479.

Shahhosseini:2022:OLO

[SSK^{+}22] Sina Shahhosseini, Dongjoo Seo, Anil Kanduri, Tianyi Hu, Sung-Soo Lim, Bryan Donyanavard, Amir M. Rahmani, and Nikil Dutt. Online learning for orchestration of inference in multi-user endedge-cloud networks. ACM Transactions on Embedded Computing Systems, 21(6): 73:1–73:??, November 2022. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3520129.

S:2023:PQA

[SSK23] Aswathy N. S., Arnab Sarkar, and Hemangee Kapoor. A predictable QoS-aware memory request scheduler for soft real-time systems. ACM Transactions on Embedded Computing Systems, 22(2):39:1–39:??, March 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3561052.

Shen:2023:TMS

[SSPP23] Yixian Shen, Leo Schreuders, Anuj Pathania, and Andy D. Pimentel. Thermal management for 3D-stacked systems via unified corememory power regulation. ACM Transactions on Embedded Computing Systems, 22(5s):120:1–120:??, October 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3608040.

Stuart:2011:RRN

[SSS11] Matthias Bo Stuart, Mikkel Bystrup Stensgaard, and Jens Sparsø. The ReNoC reconfigurable Network-on-Chip: Architecture, configuration algorithms, and evaluation. ACM Transactions on Embedded Computing Systems, 10(4):45:1–45:??, November 2011. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Shukla:2005:GES

[ST05] Sandeep K. Shukla and Jean-Pierre Talpin. Guest editorial: Special issue on models and methodologies for co-design of embedded systems. ACM Transactions on Embedded Computing Systems, 4(2):225–227, May 2005. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Sridhar:2024:SSR

[STB^{+}24] Upasana Sridhar, Nicholai Tukanov, Elliott Binder, Tze Meng Low, Scott McMillan, and Martin D. Schatz. SMaLL: Software for rapidly instantiating machine learning libraries. ACM Transactions on Embedded Computing Systems, 23(3):46:1– 46:??, May 2024. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3607870.

Siirtola:2017:WDW

[STH17] Antti Siirtola, Stavros Tripakis, and Keijo Heljanko. When do we not need complex assume-guarantee rules? ACM Transactions on Embedded Computing Systems, 16(2):48:1–48:??, April 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Shafique:2022:ISIa

[STLX22a] Muhammad Shafique, Theocharis Theocharides, Hai Li, and Chun Jason Xue. Introduction to the special issue on accelerating AI on the edge — Part 1. ACM Transactions on Embedded Computing Systems, 21(5): 47:1–47:??, September 2022. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3558078.

Shafique:2022:ISIb

[STLX22b] Muhammad Shafique, Theocharis Theocharides, Hai (Helen) Li, and Chun Jason Xue. Introduction to the special issue on accelerating AI on the edge — Part 2. ACM Transactions on Embedded Computing Systems, 21(6):67:1–67:??, November 2022. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3563127.

Sifakis:2013:ISS

[STW13] Joseph Sifakis, Lothar Thiele, and Reinhard Wilhelm. Introduction to the special section on rigorous embedded systems design. ACM Transactions on Embedded Computing Systems, 12(1s):41:1– 41:??, March 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Su:2014:RVP

[STY^{+}14] Tzu-Hsiang Su, Hsiang-Jen Tsai, Keng-Hao Yang, Po-Chun Chang, Tien-Fu Chen, and Yi-Ting Zhao. Reconfigurable vertical profiling framework for the Android runtime system. ACM Transactions on Embedded Computing Systems, 13(2s): 59:1–59:??, January 2014. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Sahoo:2023:ATS

[SUK23] Siva Satyendra Sahoo, Salim Ullah, and Akash Kumar. AxOTreeS: a tree search approach to synthesizing FPGA-based approximate operators. ACM Transactions on Embedded Computing Systems, 22(5s):101:1– 101:??, October 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3609096.

Schuster:2017:DSE

[SUS^{+}17] Simon Schuster, Peter Ulbrich, Isabella Stilkerich, Christian Dietrich, and Wolfgang SchröDer-Preikschat. Demystifying soft-error mitigation by control-flow checking — a new perspective on its effectiveness. ACM Transactions on Embedded Computing Systems, 16(5s): 180:1–180:??, October 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Susu:2020:VLA

[Sus20]  Alexandru E. Susu.  A

vector-length agnostic compiler for the Connex-S accelerator with scratchpad memory. ACM Transactions on Embedded Computing Systems, 19(6):51:1– 51:30, November 2020. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3406536.

Sorrentino:2023:HCA

[SVC^{+}23] Giuseppe Sorrentino, Marco Venere, Davide Conficconi, Eleonora D’Arnese, and Marco Domenico Santambrogio. Hephaestus: Codesigning and automating 3D image registration on reconfigurable architectures. ACM Transactions on Embedded Computing Systems, 22(5s):134:1–134:??, October 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3607928.

Stitt:2004:ESS

[SVN04] Greg Stitt, Frank Vahid, and Shawn Nematbakhsh. Energy savings and speedups from partitioning critical software loops to hardware in embedded systems. ACM Transactions on Embedded Computing Systems, 3(1): 218–232, February 2004. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Sangiovanni-Vincentelli:2005:OES

[SVP05] Alberto L. Sangiovanni-Vincentelli and Alessandro Pinto. An overview of embedded system design education at Berkeley. ACM Transactions on Embedded Computing Systems, 4(3): 472–499, August 2005. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Srinivasavarma:2021:TBC

[SVS21] Vegesna S. M. Srinivasavarma, Shiv Vidhyut, and Noor Mahammad S. A TCAMbased caching architecture framework for packet classification. ACM Transactions on Embedded Computing Systems, 20(1):2:1– 2:19, January 2021. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3409109.

Sunder:2013:FVD

[SVZ13] Christoph Sünder, Valeriy Vyatkin, and Alois Zoitl. Formal verification of downtimeless system evolution in embedded automation controllers. ACM Transactions on Embedded Computing Systems, 12(1):17:1– 17:??, January 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Salajegheh:2013:HWS

[SWJ^{+}13] Mastooreh Salajegheh, Yue Wang, Anxiao (Andrew) Jiang, Erik Learned-Miller, and Kevin Fu. Half-wits: Software techniques for lowvoltage probabilistic storage on microcontrollers with NOR flash memory. ACM Transactions on Embedded Computing Systems, 12(2s): 91:1–91:??, May 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Saeed:2019:LDB

[SWK19] Samah Mohamed Saeed, Robert Wille, and Ramesh Karri. Locking the design of building blocks for quantum circuits. ACM Transactions on Embedded Computing Systems, 18(5s):60:1– 60:??, October 2019. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3358184.

Sassone:2007:SSS

[SWL07] Peter G. Sassone, D. Scott Wills, and Gabriel H. Loh. Static strands: Safely exposing dependence chains for increasing embedded power efficiency. ACM Transactions on Embedded Computing Systems, 6(4):24:1– 24:??, September 2007. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Schumacher:2014:LLS

[SWL^{+}14] Christoph Schumacher, Jan Henrik Weinstock, Rainer Leupers, Gerd Ascheid, Laura Tosoratto, Alessandro Lonardo, Dietmar Petras, and Andreas Hoffmann. legaSCi: Legacy SystemC model integration into parallel simulators. ACM Transactions on Embedded Computing Systems, 13(5s):165:1– 165:??, November 2014. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Sha:2023:PSR

[SWL^{+}23] Zhibing Sha, Jiaojiao Wu, Jun Li, Balazs Gerofi, Zhigang Cai, and Jianwei Liao. Proactive stripe reconstruction to improve cache use efficiency of SSD-based RAID systems. ACM Transactions on Embedded Computing Systems, 22(5s):104:1– 104:??, October 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3609099.

Shijubo:2023:PBB

[SWS23] Junya Shijubo, Masaki Waga, and Kohei Suenaga. Probabilistic blackbox checking via active MDP learning. ACM Transactions on Embedded Computing Systems, 22(5s):148:1– 148:??, October 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3609127.

Saifullah:2014:NOR

[SWT^{+}14] Abusayeed Saifullah, Chengjie Wu, Paras Babu Tiwari, You Xu, Yong Fu, Chenyang Lu, and Yixin Chen. Near optimal rate selection for wireless control systems. ACM Transactions on Embedded Computing Systems, 13(4s): 128:1–128:??, April 2014. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Sun:2017:ESD

[SWWW17] Yuliang Sun, Lanjun Wang, Chen Wang, and Yu Wang. Exploiting stable data dependency in stream processing acceleration on FPGAs. ACM Transactions on Embedded Computing Systems, 16(4):116:1–116:??, August 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Sun:2013:DEI

[SWWY13] Hung-Min Sun, Chi-Yao Weng, Shiuh-Jeng Wang, and Cheng-Hsing Yang. Data embedding in imagemedia using weight-function on modulo operations. ACM Transactions on Embedded Computing Systems, 12(2): 21:1–21:??, February 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Su:2017:EWA

[SWX17] Xuesong Su, Hui Wu, and Jingling Xue. An efficient WCET-aware instruction scheduling and register allocation approach for clustered VLIW processors. ACM Transactions on Embedded Computing Systems, 16(5s):120:1–120:??, October 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Sha:2019:CED

[SXH^{+}19] Le-Tian Sha, Fu Xiao, Hai-Ping Huang, Yu Chen, and Ru-Chuan Wang. Catching escapers: a detection method for advanced persistent escapers in industry Internet of Things based on identity-based broadcast encryption (IBBE). ACM Transactions on Embedded Computing Systems, 18(3): 29:1–29:??, June 2019. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3319615.

Sotiriou-Xanthopoulos:2018:OBV

[SXMX^{+}18] Efstathios Sotiriou-Xanthopoulos, Leonard Masing, Sotirios Xydis, Kostas Siozios, Jürgen Becker, and Dimitrios Soudris. OpenCL-based virtual prototyping and simulation of many-accelerator architectures. ACM Transactions on Embedded Computing Systems, 17(5):86:1–86:??, November 2018. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3242179.

Sotiriou-Xanthopoulos:2016:FIA

[SXXS^{+}16a] Efstathios Sotiriou-Xanthopoulos, Sotirios Xydis, Kostas Siozios, George Economakos, and Dimitrios Soudris. A framework for interconnectionaware domain-specific manyaccelerator synthesis. ACM Transactions on Embedded Computing Systems, 16(1): 8:1–8:??, November 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Sotiriou-Xanthopoulos:2016:IEV

[SXXS^{+}16b] Efstathios Sotiriou-Xanthopoulos, Sotirios Xydis, Kostas Siozios, George Economakos, and Dimitrios Soudris. An integrated exploration and virtual platform framework for many-accelerator heterogeneous systems. ACM Transactions on Embedded Computing Systems, 15(3):43:1– 43:??, July 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Song:2017:SSI

[SYC^{+}17] Jun Song, Fan Yang, Kim-Kwang Raymond Choo, Zhijian Zhuang, and Lizhe Wang. SIPF: a secure installment payment framework for drive-thru Internet. ACM Transactions on Embedded Computing Systems, 16(2):52:1–52:??, April 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Shadab:2023:HHC

[SZG^{+}23] Rakin Muhammad Shadab, Yu Zou, Sanjay Gandham, Amro Awad, and Mingjie Lin. HMT: a hardwarecentric hybrid bonsai Merkle tree algorithm for highperformance authentication. ACM Transactions on Embedded Computing Systems, 22(4):64:1–64:??, July 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3595179.

Shu:2017:WDD

[SZL^{+}17] Junliang Shu, Yuanyuan Zhang, Juanru Li, Bodong Li, and Dawu Gu. Why data deletion fails? A study on deletion flaws and data remanence in Android systems. ACM Transactions on Embedded Computing Systems, 16(2):61:1–61:??, April 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Tsoutsouras:2018:HDR

[TAMS18] Vasileios Tsoutsouras, Iraklis Anagnostopoulos, Dimosthenis Masouros, and Dimitrios Soudris. A hierarchical distributed runtime resource management scheme for NoC-based many-cores. ACM Transactions on Embedded Computing Systems, 17(3):65:1–65:??, June 2018. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Tsounis:2023:MFT

[TAP23] Ioannis Tsounis, Dimitris Agiakatsikas, and Mihalis Psarakis. A methodology for fault-tolerant Paretooptimal approximate designs of FPGA-based accelerators. ACM Transactions on Embedded Computing Systems, 22(4):78:1–78:??, July 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3568021.

Thilakasiri:2023:MRP

[TB23] Thilanka Thilakasiri and Matthias Becker. Methods to realize preemption in phased execution models. ACM Transactions on Embedded Computing Systems, 22(5s):153:1–153:??, October 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3609132.

Tan:2017:ASA

[TBAS17] Benjamin Tan, Morteza Biglari-Abhari, and Zoran Salcic. An automated security-aware approach for design of embedded systems on MPSoC. ACM Transactions on Embedded Computing Systems, 16(5s):143:1– 143:??, October 2017. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Tilli:2015:GCR

[TBCB15] Andrea Tilli, Andrea Bartolini, Matteo Cacciari, and Luca Benini. Guaranteed computational resprinting via model-predictive control. ACM Transactions on Embedded Computing Systems, 14(3):48:1–48:??, April 2015. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Tuncel:2023:SSC

[TBCGO23] Yigit Tuncel, Toygun Basaklar, Dina Carpenter-Graffy, and Umit Ogras. A selfsustained CPS design for reliable wildfire monitoring. ACM Transactions on Embedded Computing Systems, 22(5s):135:1–135:??, October 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3608100.

Touati:2011:ESR

[TBDdD11] Sid-Ahmed-Ali Touati, Frederic Brault, Karine Deschinkel, and Benoît Dupont de Dinechin. Efficient spilling reduction for software pipelined loops in presence of multiple register types in embedded VLIW processors. ACM Transactions on Embedded Computing Systems, 10(4):47:1– 47:??, November 2011. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Tanasa:2016:CAP

[TBEP16] Bogdan Tanasa, Unmesh D.

Bordoloi, Petru Eles, and Zebo Peng. Correlationaware probabilistic timing analysis for the dynamic segment of FlexRay. ACM Transactions on Embedded Computing Systems, 15(3): 54:1–54:??, July 2016. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Tigori:2017:FMB

[TBFR17] Kabland Toussaint Gautier Tigori, Jean-Luc Béchennec, Sébastien Faucou, and Olivier Henri Roux. Formal model-based synthesis of applicationspecific static RTOS. ACM Transactions on Embedded Computing Systems, 16(4): 97:1–97:??, August 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Tripakis:2013:CSD

[TBG^{+}13] Stavros Tripakis, Dai Bui, Marc Geilen, Bert Rodiers, and Edward A. Lee. Compositionality in synchronous data flow: Modular code generation from hierarchical SDF graphs. ACM Transactions on Embedded Computing Systems, 12(3):83:1– 83:??, March 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Tang:2017:TFC

[TBG^{+}17] Qi Tang, Twan Basten, Marc Geilen, Sander Stuijk, and Ji-Bo Wei. Task-FIFO coscheduling of streaming applications on MPSoCs with predictable memory hierarchy. ACM Transactions on Embedded Computing Systems, 16(2):49:1–49:??, April 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Tran:2019:SVC

[TCD^{+}19] Hoang-Dung Tran, Feiyang Cai, Manzanas Lopez Diego, Patrick Musau, Taylor T. Johnson, and Xenofon Koutsoukos. Safety verification of cyber-physical systems with reinforcement learning control. ACM Transactions on Embedded Computing Systems, 18(5s):105:1– 105:??, October 2019. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3358230.

Tajik:2016:SRS

[TDD^{+}16] Hossein Tajik, Bryan Donyanavard, Nikil Dutt, Janmartin Jahn, and Jörg Henkel. SPMPool: Runtime SPM management for memory-intensive applications in embedded manycores. ACM Transactions on Embedded Computing Systems, 16(1):25:1–25:??, November 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Tardieu:2005:LE

[TdS05] Olivier Tardieu and Robert de Simone. Loops in ES-TEREL. ACM Transactions on Embedded Computing Systems, 4(4):708– 750, November 2005. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Staff:2012:APA

[TEC12] TECS Staff. Abstracts of papers to appear in Special Supplemental Issue of TECS (v11, iSupplemental1). ACM Transactions on Embedded Computing Systems, 11(2):25:1–25:??, July 2012. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Theocharides:2024:ISI

[TFC24] Theocharis Theocharides, Charlotte Frenkel, and Lukas Cavigelli. Introduction to the special issue on tinyML. ACM Transactions on Embedded Computing Systems, 23(3):40:1–40:??, May 2024. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3658375.

Tuncali:2016:APM

[TFL16] Cumhur Erkan Tuncali, Georgios Fainekos, and Yann-Hang Lee. Automatic parallelization of multirate block diagrams of control systems on multicore platforms. ACM Transactions on Embedded Computing Systems, 16(1):15:1– 15:??, November 2016. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Tretter:2017:MAC

[TGBT17] Andreas Tretter, Georgia Giannopoulou, Matthias Baer, and Lothar Thiele. Minimising access conflicts on shared multi-bank memory. ACM Transactions on Embedded Computing Systems, 16(5s):135:1–135:??, October 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Tijero:2025:ALE

[TGP25] Hector Perez Tijero, J. Javier Gutiérrez García, and Diego García Prieto. Application-level evaluation of IEEE 802.1AS synchronized time and Linux for distributed real-time systems. ACM Transactions on Embedded Computing Systems, 24(1):13:1–13:??, January 2025. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL

https://dl.acm.org/doi/
10.1145/3701300.

Trub:2017:IPM

[TGTT17] Roman Trüb, Georgia Giannopoulou, Andreas Tretter, and Lothar Thiele. Implementation of partitioned mixed-criticality scheduling on a multi-core platform. ACM Transactions on Embedded Computing Systems, 16(5s):122:1–122:??, October 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Tang:2012:UMS

[TGV12] Qinghui Tang, Sandeep K. S. Gupta, and Georgios Varsamopoulos. A unified methodology for scheduling in distributed cyber-physical systems. ACM Transactions on Embedded Computing Systems, 11(S2):57:1–57:??, 2012. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Terechko:2012:BPS

[THA^{+}12] Andrei Terechko, Jan Hoogerbrugge, Ghiath Alkadi, Surendra Guntur, Anirban Lahiri, Marc Duranton, Clemens Wüst, Phillip Christie, Axel Nackaerts, and Aatish Kumar. Balancing programmability and silicon efficiency of heterogeneous multicore architectures. ACM Transactions on Embedded Computing Systems, 11S(1):14:1–14:??, 2012. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Thoma:2024:AAS

[THG24] Jan Philipp Thoma, Darius Hartlief, and Tim Güneysu. Agile acceleration of stateful hash-based signatures in hardware. ACM Transactions on Embedded Computing Systems, 23(2):29:1– 29:??, March 2024. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3567426.

Tsutsui:2012:HTP

[THON12] Hiroshi Tsutsui, Koichi Hattori, Hiroyuki Ochi, and Yukihiro Nakamura. A highthroughput pipelined parallel architecture for JPEG XR encoding. ACM Transactions on Embedded Computing Systems, 11(4):72:1– 72:??, December 2012. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Tahaee:2010:PAP

[TJ10] Seyed-Abdoreza Tahaee and Amir Hossein Jahangir. A polynomial algorithm for partitioning problems. ACM Transactions on Embedded Computing Systems, 9(4): 34:1–34:??, March 2010. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Terway:2023:RGM

[TJ23] Prerit Terway and Niraj K.

Jha. REPAIRS: Gaussian mixture model-based completion and optimization of partially specified systems. ACM Transactions on Embedded Computing Systems, 22(4):69:1–69:??, July 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3605147.

Turjan:2007:CIC

[TKD07] Alexandru Turjan, Bart Kienhuis, and Ed Deprettere. Classifying interprocess communication in process network representation of nested-loop programs. ACM Transactions on Embedded Computing Systems, 6(2): 13:1–13:??, May 2007. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Theodoropoulos:2013:CAM

[TKG13] Dimitris Theodoropoulos, Georgi Kuzmanov, and Georgi Gaydadjiev. Custom architecture for multicore audio beamforming systems. ACM Transactions on Embedded Computing Systems, 13(2):19:1–19:??, September 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Trajkovic:2022:PMA

[TKHZ22] Jelena Trajkovic, Sara Karimi, Samantha Hangsan, and Wenlu Zhang. Prediction modeling for applicationspecific communication architecture design of optical NoC. ACM Transactions on Embedded Computing Systems, 21(4):35:1–35:??, July 2022. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3520241.

Tsai:2015:JPI

[TKL^{+}15] Chun-Jen Tsai, Han-Wen Kuo, Zigang Lin, Zi-Jing Guo, and Jun-Fu Wang. A Java processor IP design for embedded SoC. ACM Transactions on Embedded Computing Systems, 14(2):35:1– 35:??, March 2015. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Ttofis:2015:HEA

[TKT15] Christos Ttofis, Christos Kyrkou, and Theocharis Theocharides. A hardwareefficient architecture for accurate real-time disparity map estimation. ACM Transactions on Embedded Computing Systems, 14(2): 36:1–36:??, March 2015. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Tan:2018:LLP

[TKV^{+}18] Cheng Tan, Aditi Kulkarni, Vanchinathan Venkataramani, Manupa Karunaratne, Tulika Mitra, and Li-Shiuan Peh. LOCUS: Low-power customizable many-core architecture for wearables. ACM Transactions on Embedded Computing Systems, 17(1):16:1–16:??, January 2018. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Tillenius:2015:RAT

[TLBM15] Martin Tillenius, Elisabeth Larsson, Rosa M. Badia, and Xavier Martorell. Resourceaware task scheduling. ACM Transactions on Embedded Computing Systems, 14(1): 5:1–5:??, January 2015. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Thatte:2012:KEE

[TLL^{+}12] Gautam Thatte, Ming Li, Sangwon Lee, Adar Emken, Shrikanth Narayanan, Urbashi Mitra, Donna Spruijt-Metz, and Murali Annavaram. KNOWME: An energy-efficient multimodal body area network for physical activity monitoring. ACM Transactions on Embedded Computing Systems, 11(S2):48:1–48:??, 2012. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Tseng:2009:FSA

[TLLL09] Kuo-Kun Tseng, Yuan-Cheng Lai, Ying-Dar Lin, and Tsern-Huei Lee. A fast scalable automatonmatching accelerator for embedded content processors. ACM Transactions on Embedded Computing Systems, 8(3):19:1–19:??, April 2009. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Tuli:2023:CNA

[TLSJ23] Shikhar Tuli, Chia-Hao Li, Ritvik Sharma, and Niraj K. Jha. CODEBench: a neural architecture and hardware accelerator co-design framework. ACM Transactions on Embedded Computing Systems, 22(3):51:1–51:??, May 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3575798.

Tan:2007:TAP

[TM07] Yudong Tan and Vincent Mooney. Timing analysis for preemptive multitasking real-time systems with caches. ACM Transactions on Embedded Computing Systems, 6(1):7:1– 7:??, February 2007. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Taniuchi:2015:AUI

[TM15] Daisuke Taniuchi and Takuya Maekawa. Automatic update of indoor location fingerprints with pedestrian dead reckoning. ACM Transactions on Embedded Computing Systems, 14(2):27:1– 27:??, March 2015. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Tsoutsouras:2017:SSO

[TMXS17] Vasileios Tsoutsouras, Dimosthenis Masouros, Sotirios Xydis, and Dimitrios Soudris. SoftRM: Self-organized faulttolerant resource management for failure detection and recovery in NoC based many-cores. ACM Transactions on Embedded Computing Systems, 16(5s):144:1– 144:??, October 2017. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Tiloca:2017:ADB

[TNR17] Marco Tiloca, Kirill Nikitin, and Shahid Raza. Axiom: DTLS-based secure IoT group communication. ACM Transactions on Embedded Computing Systems, 16(3):66:1–66:??, July 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Thomas:2016:EDP

[TP16] Anna Thomas and Karthik Pattabiraman. Error detector placement for soft computing applications. ACM Transactions on Embedded Computing Systems, 15(1): 8:1–8:??, February 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Tabrizi:2019:DLC

[TP19] Farid Molazem Tabrizi and Karthik Pattabiraman. Designlevel and code-level security analysis of IoT devices. ACM Transactions on Embedded Computing Systems, 18(3):20:1–20:??, June 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/ft_gateway.
cfm?id=3310353.

Tiku:2020:OSV

[TP20] Saideep Tiku and Sudeep Pasricha. Overcoming security vulnerabilities in deep learning-based indoor localization frameworks on mobile devices. ACM Transactions on Embedded Computing Systems, 18(6):1–24, January 2020. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL

https://dl.acm.org/doi/
abs/10.1145/3362036.

Thuijsman:2024:SCD

[TR24] Sander Thuijsman and Michel Reniers. Supervisory control for dynamic feature configuration in product lines. ACM Transactions on Embedded Computing Systems, 23(5):71:1–71:??, September 2024. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3579644.

Tan:2005:EME

[TRJ05] T. K. Tan, A. Raghunathan, and N. K. Jha. Energy macromodeling of embedded operating systems. ACM Transactions on Embedded Computing Systems, 4(1): 231–254, February 2005. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Thiele:2013:PTT

[TSBY13] Lothar Thiele, Lars Schor, Iuliana Bacivarov, and Hoeseok Yang. Predictability for timing and temperature in multiprocessor system-on-chip platforms. ACM Transactions on Embedded Computing Systems, 12(1s):48:1– 48:??, March 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Tripakis:2005:TDT

[TSCC05] Stavros Tripakis, Christos Sofronis, Paul Caspi, and Adrian Curic. Translating discrete-time Simulink to Lustre. ACM Transactions on Embedded Computing Systems, 4(4):779– 818, November 2005. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Tichy:2010:GAF

[TSG10] Milan Tichy, Jan Schier, and David Gregg. GS-FAP adaptive filtering using log arithmetic for resourceconstrained embedded systems. ACM Transactions on Embedded Computing Systems, 9(3):29:1–29:??, February 2010. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Tsouvalas:2022:FST

[TSO22] Vasileios Tsouvalas, Aaqib Saeed, and Tanir Ozcelebi. Federated self-training for semi-supervised audio recognition. ACM Transactions on Embedded Computing Systems, 21(6):74:1– 74:??, November 2022. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3520128.

Tamas-Selicean:2015:DOM

[TSP15] Domitian Tamas-Selicean and Paul Pop. Design optimization of mixed-criticality real-time embedded systems. ACM Transactions on Embedded Computing Systems, 14(3):50:1–50:??, May 2015. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Tan:2017:ITM

[TSW^{+}17] Wilson M. Tan, Paul Sullivan, Hamish Watson, Joanna Slota-Newson, and Stephen A. Jarvis. An indoor test methodology for solar-powered wireless sensor networks. ACM Transactions on Embedded Computing Systems, 16(3):82:1– 82:??, July 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Tan:2010:MSE

[TSWL10] Chiu C. Tan, Bo Sheng, Haodong Wang, and Qun Li. Microsearch: a search engine for embedded devices used in pervasive computing. ACM Transactions on Embedded Computing Systems, 9(4):43:1–43:??, March 2010. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Tan:2016:SSH

[TSY^{+}16] Song Tan, Wen-Zhan Song, Steve Yothment, Junjie Yang, and Lang Tong. Score-Plus: a software-hardware hybrid and federated experiment environment for smart grid. ACM Transactions on Embedded Computing Systems, 16(1):19:1– 19:??, November 2016. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Tiwari:2020:RRA

[TTA^{+}20] Sakshi Tiwari, Shreshth Tuli, Isaar Ahmad, Ayushi Agarwal, Preeti Ranjan Panda, and Sreenivas Subramoney. REAL: REquest arbitration in last level caches. ACM Transactions on Embedded Computing Systems, 18(6): 1–24, January 2020. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3362100.

Tavana:2014:SHT

[TTAG14] Mohammad Khavari Tavana, Nasibeh Teimouri, Meisam Abdollahi, and Maziar Goudarzi. Simultaneous hardware and time redundancy with online task scheduling for low energy highly reliable standbysparing system. ACM Transactions on Embedded Computing Systems, 13(4):86:1– 86:??, February 2014. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Tabanelli:2023:DAY

[TTB23] Enrico Tabanelli, Giuseppe Tagliavini, and Luca Benini. DNN is not all you need: Parallelizing non-neural ML algorithms on ultra-low-power IoT processors. ACM Transactions on Embedded Computing Systems, 22(3): 56:1–56:??, May 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3571133.

Turan:2019:CFF

[TV19] Furkan Turan and Ingrid Verbauwhede. Compact and flexible FPGA implementation of Ed25519 and X25519. ACM Transactions on Embedded Computing Systems, 18(3):24:1–24:??, June 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/ft_gateway.
cfm?id=3312742.

Trajkovic:2008:ISA

[TVK08] Jelena Trajkovic, Alexander V. Veidenbaum, and Arun Kejariwal. Improving SDRAM access energy efficiency for low-power embedded systems. ACM Transactions on Embedded Computing Systems, 7(3):24:1– 24:??, April 2008. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Tanase:2018:SML

[TWTH18] Alexandru Tanase, Michael Witterauf, Jürgen Teich, and Frank Hannig. Symbolic multi-level loop mapping of loop programs for massively parallel processor arrays. ACM Transactions on Embedded Computing Systems, 17(2):31:1–31:??, April 2018. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Tan:2012:ACF

[TXL^{+}12] Rui Tan, Guoliang Xing, Xue Liu, Jianguo Yao, and Zhaohui Yuan. Adaptive calibration for fusionbased cyber-physical systems. ACM Transactions on Embedded Computing Systems, 11(4):80:1–80:??, December 2012. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Unsal:2003:CCC

[UAK^{+}03] Osman S. Unsal, Raksit Ashok, Israel Koren, C. Mani Krishna, and Csaba Andras Moritz. Cool-Cache: a compiler-enabled energy efficient data caching framework for embedded/multimedia processors. ACM Transactions on Embedded Computing Systems, 2(3):373– 392, August 2003. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Ungerer:2016:PIH

[UBF^{+}16] Theo Ungerer, Christian Bradatsch, Martin Frieb, Florian Kluge, Jörg Mische, Alexander Stegmeier, Ralf Jahr, Mike Gerdes, Pavel Zaykov, Lucie Matusova, Zai Jian Jia Li, Zlatko Petrov, Bert Böddeker, Sebastian Kehr, Hans Regler, Andreas Hugl, Christine Rochange, Haluk Ozaktas, Hugues Cassé, Armelle Bonenfant, Pascal Sainrat, Nick Lay, David George, Ian Broster, Eduardo Quiñones, Milos Panic, Jaume Abella, Carles Hernandez, Francisco Cazorla, Sascha Uhrig, Mathias Rohde, and Arthur Pyka. Parallelizing industrial hard real-time applications for the parMERASA multicore. ACM Transactions on Embedded Computing Systems, 15(3):53:1–53:??, July 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Unnikrishnan:2009:RMR

[UCK^{+}09] P. Unnikrishnan, G. Chen, M. Kandemir, M. Karakoy, and I. Kolcu. Reducing memory requirements of resource-constrained applications. ACM Transactions on Embedded Computing Systems, 8(3):17:1– 17:??, April 2009. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Udayakumaran:2006:DAS

[UDB06] Sumesh Udayakumaran, Angel Dominguez, and Rajeev Barua. Dynamic allocation for scratch-pad memory using compile-time decisions. ACM Transactions on Embedded Computing Systems, 5(2):472–511, May 2006. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Ulus:2024:ETP

[UFA^{+}24] Dogan Ulus, Thomas Ferrère, Eugene Asarin, Dejan Nickovic, and Oded Maler. Elements of timed pattern matching. ACM Transactions on Embedded Computing Systems, 23(4):59:1– 59:??, July 2024. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3645114.

Ungureanu:2021:FAT

[UGS^{+}21] George Ungureanu, José Edil Guimarães De Medeiros, Timmy Sundström, Ingemar Söderquist, Anders Åhlander, and Ingo Sander. ForSyDe-Atom: Taming complexity in cyber physical system design with layers. ACM Transactions on Embedded Computing Systems, 20(2):10:1–10:27, March 2021. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3424667.

Uzelac:2013:HBL

[UM13] Vladimir Uzelac and Aleksandar Milenković. Hardwarebased load value trace filtering for on-the-fly debugging. ACM Transactions on Embedded Computing Systems, 12(2s):97:1–97:??, May 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Ullah:2022:ADA

[USA^{+}22] Salim Ullah, Siva Satyendra Sahoo, Nemath Ahmed, Debabrata Chaudhury, and Akash Kumar. AppAxO: Designing application-specific approximate operators for FPGA-based embedded systems. ACM Transactions on Embedded Computing Systems, 21(3):29:1–29:31, May 2022. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3513262.

Vatanparvar:2018:DAB

[VA18] Korosh Vatanparvar and Mohammad Abdullah Al Faruque. Design and analysis of battery-aware automotive climate control for electric vehicles. ACM Transactions on Embedded Computing Systems, 17(4):74:1– 74:??, August 2018. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Varea:2006:DFN

[VAHC^{+}06] Mauricio Varea, Bashir M.

Al-Hashimi, Luis A. CortéS, Petru Eles, and Zebo Peng. Dual Flow Nets: Modeling the control/data-flow relation in embedded systems. ACM Transactions on Embedded Computing Systems, 5(1):54–81, February 2006. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Valmari:2017:SIS

[Val17] Antti Valmari. Stop it, and be stubborn! ACM Transactions on Embedded Computing Systems, 16(2):46:1– 46:??, April 2017. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Valtazanos:2013:LSS

[VAR13] Aris Valtazanos, D. K. Arvind, and Subramanian Ramamoorthy. Latent space segmentation for mobile gait analysis. ACM Transactions on Embedded Computing Systems, 12(4):101:1– 101:??, June 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Venkataramani:2019:SMM

[VCM19] Vanchinathan Venkataramani, Mun Choon Chan, and Tulika Mitra. Scratchpadmemory management for multi-threaded applications on many-core architectures. ACM Transactions on Embedded Computing Systems, 18(1):10:1–10:??, February 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

                                        /dl.acm.org/ft_gateway.
                                        cfm?id=3301308.

Varma:2008:AFS

[VDK^{+}08] Ankush Varma, Eric Debes, Igor Kozintsev, Paul Klein, and Bruce Jacob. Accurate and fast system-level power modeling: an XScale-based case study. ACM Transactions on Embedded Computing Systems, 7(3):25:1– 25:??, April 2008. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Vijzelaar:2017:MVS

[VF17] Stefan Vijzelaar and Wan Fokkink. Multi-valued simulation and abstraction using lattice operations. ACM Transactions on Embedded Computing Systems, 16(2): 42:1–42:??, April 2017. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Valente:2021:CMS

[VFS^{+}21] Giacomo Valente, Tiziana Fanni, Carlo Sau, Tania Di Mascio, Luigi Pomante, and Francesca Palumbo. A composable monitoring system for heterogeneous embedded platforms. ACM Transactions on Embedded Computing Systems, 20(5): 42:1–42:34, July 2021. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3461647.

VanPinxten:2019:PSC

[VGB19] Joost Van Pinxten, Marc Geilen, and Twan Basten. Parametric scheduler characterization. ACM Transactions on Embedded Computing Systems, 18(5s):110:1– 110:??, October 2019. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

                                        tronic). URL https://dl.
                                        acm.org/ft_gateway.cfm?
                                        id=3358226.

Vyas:2013:HAS

[VGG^{+}13] Sudhanshu Vyas, Adwait Gupte, Christopher D. Gill, Ron K. Cytron, Joseph Zambreno, and Phillip H. Jones. Hardware architectural support for control systems and sensor processing. ACM Transactions on Embedded Computing Systems, 13(2):16:1–16:??, September 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Vasilios:2018:CSC

[VGN18] Kelefouras Vasilios, Keramidas Georgios, and Voros Nikolaos. Combining software cache partitioning and loop tiling for effective shared cache management. ACM Transactions on Embedded Computing Systems, 17(3):72:1–72:??, June 2018. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Voros:2013:MHD

[VHB^{+}13] Nikolaos S. Voros, Michael Hübner, Jürgen Becker, Matthias Kühnle, Florian Thomaitiv, Arnaud Grasset, Paul Brelet, Philippe Bonnot, Fabio Campi, Eberhard Schüler, Henning Sahlbach, Sean Whitty, Rolf Ernst, Enrico Billich, Claudia Tischendorf, Ulrich Heinkel, Frank Ieromnimon, Dimitrios Kritharidis, Axel Schneider, Joachim Knaeblein, and Wolfram Putzke-Röming. MORPHEUS: a heterogeneous dynamically reconfigurable platform for designing highly complex embedded systems. ACM Transactions on Embedded Computing Systems, 12(3): 70:1–70:??, March 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Varma:2007:AFS

[VJD^{+}07] Ankush Varma, Bruce Jacob, Eric Debes, Igor Kozintsev, and Paul Klein. Accurate and fast system-level power modeling: an XScalebased case study. ACM Transactions on Embedded Computing Systems, 6(4): 26:1–26:??, September 2007. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Vashist:2019:UTS

[VKDG19] Abhishek Vashist, Andrew Keats, Sai Manoj Pudukotai Dinakarrao, and Amlan Ganguly. Unified testing and security framework for wireless network-on-chip enabled multi-core chips. ACM Transactions on Embedded Computing Systems, 18(5s):75:1–75:??, October 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/ft_gateway.
cfm?id=3358212.

Venkataramani:2020:SSD

[VKMP20] Vanchinathan Venkataramani, Aditi Kulkarni, Tulika Mitra, and Li-Shiuan Peh. SPECTRUM: a softwaredefined predictable manycore architecture for LTE/ 5G baseband processing. ACM Transactions on Embedded Computing Systems, 19(5):32:1–32:28, November 2020. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3400032.

Vogel:2017:EVM

[VKW^{+}17] Pirmin Vogel, Andreas Kurth, Johannes Weinbuch, Andrea Marongiu, and Luca Benini. Efficient virtual memory sharing via onaccelerator page table walking in heterogeneous embedded SoCs. ACM Transactions on Embedded Computing Systems, 16(5s):154:1– 154:??, October 2017. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Vera:2007:DCL

[VLX07] Xavier Vera, Björn Lisper, and Jingling Xue. Data cache locking for tight timing calculations. ACM Transactions on Embedded Computing Systems, 7(1):4:1– 4:38, December 2007. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Vreman:2023:SAC

[VM23] Nils Vreman and Martina Maggio. Stochastic analysis of control systems subject to communication and computation faults. ACM Transactions on Embedded Computing Systems, 22(5s): 144:1–144:??, October 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3609123.

Venkataramani:2003:ACC

[VNK^{+}03] Girish Venkataramani, Walid Najjar, Fadi Kurdahi, Nader Bagherzadeh, Wim Bohm, and Jeff Hammes. Automatic compilation to a coarse-grained reconfigurable system-on-a-chip. ACM Transactions on Embedded Computing Systems, 2(4): 560–589, November 2003. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

VonMaurich:2015:IQM

[VOG15] Ingo Von Maurich, Tobias Oder, and Tim Güneysu. Implementing QC–MDPC McEliece encryption. ACM Transactions on Embedded Computing Systems, 14(3): 44:1–44:??, April 2015. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Vinco:2016:ESI

[VP16] Sara Vinco and Christian Pilato. Editorial: Special issue on innovative design methods for smart embedded systems. ACM Transactions on Embedded Computing Systems, 15(2):22:1–22:??, May 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

VanHulst:2015:MSH

[VRF15] A. C. Van Hulst, M. A. Reniers, and W. J. Fokkink. Maximal synthesis for Hennessy– Milner logic. ACM Transactions on Embedded Computing Systems, 14(1):10:1– 10:??, January 2015. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Verbauwhede:2005:SES

[VS05] Ingrid Verbauwhede and Patrick Schaumont. Skiing the embedded systems mountain. ACM Transactions on Embedded Computing Systems, 4(3):529– 548, August 2005. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Voyiatzis:2008:SFS

[VS08] Artemios G. Voyiatzis and Dimitrios N. Serpanos. The security of the Fiat–Shamir scheme in the presence of transient hardware faults. ACM Transactions on Embedded Computing Systems, 7(3):31:1–31:??, April 2008. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Vougioukas:2017:NFS

[VSD^{+}17] Ilias Vougioukas, Andreas Sandberg, Stephan Diestelhorst, Bashir M. Al-Hashimi, and Geoff V. Merrett. Nucleus: Finding the sharing limit of heterogeneous cores. ACM Transactions on Embedded Computing Systems, 16(5s):152:1–152:??, October 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Vasilikos:2013:HSA

[VSSS13] Vasileios Vasilikos, Georgios Smaragdos, Christos Strydis, and Ioannis Sourdis. Heuristic search for adaptive, defect-tolerant multiprocessor arrays. ACM Transactions on Embedded Computing Systems, 12(1s):44:1– 44:??, March 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Vali:2023:BSD

[VVKG23] Kourosh Vali, Ata Vafi, Begum Kasap, and Soheil Ghiasi. BASS: Safe deep tissue optical sensing for wearable embedded systems. ACM Transactions on Embedded Computing Systems, 22(5s):122:1–122:??, October 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3607916.

VanPinxten:2017:OSR

[VWG^{+}17] Joost Van Pinxten, Umar Waqas, Marc Geilen, Twan Basten, and Lou Somers. Online scheduling of 2-re-entrant flexible manufacturing systems. ACM Transactions on Embedded Computing Systems, 16(5s):160:1– 160:??, October 2017. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Whitham:2014:ERC

[WAD14] Jack Whitham, Neil C. Audsley, and Robert I. Davis. Explicit reservation of cache memory in a predictable, preemptive multitasking real-time system. ACM Transactions on Embedded Computing Systems, 13(4s):120:1–120:??, April 2014. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Winter:2006:TPC

[WBF^{+}06] Victor L. Winter, Jason Beranek, Fares Fraij, Steve Roach, and Greg Wickstrom. A transformational perspective into the core of an abstract class loader for the SSP. ACM Transactions on Embedded Computing Systems, 5(4):773– 818, November 2006. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Wiggers:2010:BCC

[WBS10] Maarten H. Wiggers, Marco J. G. Bekooij, and Gerard J. M. Smit. Buffer capacity computation for throughputconstrained modal task graphs. ACM Transactions on Embedded Computing Systems, 10(2):17:1– 17:??, December 2010. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Wu:2016:JJO

[WC16] Chin-Hsien Wu and Syuan-An Chen. JOM: a joint operation mechanism for NAND flash memory. ACM Transactions on Embedded Computing Systems, 15(4):74:1– 74:??, August 2016. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Wang:2020:CTC

[WCB20] Chundong Wang, Sudipta Chattopadhyay, and Gunavaran Brihadiswarn. Crabtree: a crash recoverable B+-tree variant for persistent memory with ARMv8 architecture. ACM Transactions on Embedded Computing Systems, 19(5):35:1– 35:26, November 2020. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3396236.

Wu:2023:WTM

[WCH^{+}23] Shin-Ting Wu, Liang-Chi Chen, Po-Chun Huang, Yuan-Hao Chang, Chien-Chung Ho, and Wei-Kuan Shih. WARM-tree: Making quadtrees write-efficient and space-economic on persistent memories. ACM Transactions on Embedded Computing Systems, 22(5s):119:1– 119:??, October 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3608033.

Wagner:2007:HSI

[WCJ07] Flávio R. Wagner, Wander Cesário, and Ahmed A. Jerraya. Hardware/software IP integration using the ROSES design environment. ACM Transactions on Embedded Computing Systems, 6(3): 17:1–17:??, July 2007. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Wang:2019:ALA

[WCK^{+}19] Wei-Chen Wang, Yuan-Hao Chang, Tei-Wei Kuo, Chien-Chung Ho, Yu-Ming Chang, and Hung-Sheng Chang. Achieving lossless accuracy with lossy programming for efficient neuralnetwork training on NVMbased systems. ACM Transactions on Embedded Computing Systems, 18(5s):68:1– 68:??, October 2019. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3358191.

Wu:2016:SAR

[WCM^{+}16] Xueguang Wu, Liqian Chen, Antoine Miné, Wei Dong, and Ji Wang. Static analysis of runtime errors in interrupt-driven programs via sequentialization. ACM Transactions on Embedded Computing Systems, 15(4): 70:1–70:??, August 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Wagemann:2018:OEN

[WDJ^{+}18] Peter Wägemann, Tobias Distler, Heiko Janker, Phillip Raffeck, Volkmar Sieh, and Wolfgang Schröder-Preikschat. Operating energy-neutral real-time systems. ACM Transactions on Embedded Computing Systems, 17(1):11:1–11:??, January 2018. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Wang:2017:PAP

[WDM17] Yi Wang, Lisha Dong, and Rui Mao. P-Alloc: Processvariation tolerant reliability management for 3D chargetrapping flash memory. ACM Transactions on Embedded Computing Systems, 16(5s): 142:1–142:??, October 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Wang:2023:EBN

[WDM^{+}23] Erwei Wang, James J. Davis, Daniele Moro, Piotr Zielinski, Jia Jie Lim, Claudionor Coelho, Satrajit Chatterjee, Peter Y. K. Cheung, and George A. Constantinides. Enabling binary neural network training on the edge. ACM Transactions on Embedded Computing Systems, 22(6):105:1– 105:??, November 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3626100.

Wang:2016:GTB

[WDY^{+}16] Kun Wang, Miao Du, Dejun Yang, Chunsheng Zhu, Jian Shen, and Yan Zhang. Game-theory-based active defense for intrusion detection in cyber-physical embedded systems. ACM Transactions on Embedded Computing Systems, 16(1): 18:1–18:??, November 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Wilhelm:2008:WCE

[WEE^{+}08] Reinhard Wilhelm, Jakob Engblom, Andreas Ermedahl, Niklas Holsti, Stephan Thesing, David Whalley, Guillem Bernat, Christian Ferdinand, Reinhold Heckmann, Tulika Mitra, Frank Mueller, Isabelle Puaut, Peter Puschner, Jan Staschulat, and Per Stenström. The worst-case execution-time problem— overview of methods and survey of tools. ACM Transactions on Embedded Computing Systems, 7(3):36:1– 36:??, April 2008. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

West:2023:RTU

[WGN23] Richard West, Ahmad Golchin, and Anton Njavro. Real-time USB networking and device I/O. ACM Transactions on Embedded Computing Systems, 22(4):67:1–67:??, July 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3604429.

Wollinger:2004:SFS

[WGP04] Thomas Wollinger, Jorge Guajardo, and Christof Paar. Security on FPGAs: State-of-the-art implementations and attacks. ACM Transactions on Embedded Computing Systems, 3(3): 534–574, August 2004. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Wijsman:2013:TME

[WGPH13] Jacqueline Wijsman, Bernard Grundlehner, Julien Penders, and Hermie Hermens. Trapezius muscle EMG as predictor of mental stress. ACM Transactions on Embedded Computing Systems, 12(4):99:1–99:??, June 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Wang:2017:DRM

[WH17] Yi Wang and Yajun Ha. A DFA-resistant and masked PRESENT with area optimization for RFID applications. ACM Transactions on Embedded Computing Systems, 16(4):102:1– 102:??, August 2017. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Whalley:2007:GE

[Wha07] David Whalley. Guest Editorial. ACM Transactions on Embedded Computing Systems, 6(1):1:1–1:??, February 2007. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Wu:2023:SFA

[WHL23] Jun-Shen Wu, Tsen-Wei Hsu, and Ren-Shuo Liu. SG-Float: Achieving memory access and computing power reduction using self-gating float in CNNs. ACM Transactions on Embedded Computing Systems, 22(6):101:1– 101:??, November 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3624582.

Wang:2017:HAT

[WHN^{+}17] Tianyi Wang, Soamar Homsi, Linwei Niu, Shaolei Ren, Ou Bai, Gang Quan, and Meikang Qiu. Harmonicityaware task partitioning for fixed priority scheduling of probabilistic real-time tasks on multi-core platforms. ACM Transactions on Embedded Computing Systems, 16(4):101:1–101:??, August 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Wu:2017:SVA

[WJ17] Jian Wu and Roozbeh Jafari. Seamless vision-assisted placement calibration for wearable inertial sensors. ACM Transactions on Embedded Computing Systems, 16(3):71:1–71:??, July 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Wu:2007:EBT

[WKC07] Chin-Hsien Wu, Tei-Wei Kuo, and Li Ping Chang. An efficient B-tree layer implementation for flash-memory storage systems. ACM Transactions on Embedded Computing Systems, 6(3): 19:1–19:??, July 2007. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Wade:2020:EIP

[WKJ20] April W. Wade, Prasad A. Kulkarni, and Michael R. Jantz. Exploring impact of profile data on code quality in the HotSpot JVM. ACM Transactions on Embedded Computing Systems, 19(6):48:1–48:26, November 2020. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3391894.

Wu:2024:HPP

[WKP24] Zhuanhao Wu, Anirudh Kaushik, and Hiren Patel. High performance and predictable shared last-level cache for safety-critical systems. ACM Transactions on Embedded Computing Systems, 23(6):97:1–97:??, November 2024. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3687308.

Wang:2009:TTD

[WL09] Xiaojun Wang and Miriam Leeser. A truly twodimensional systolic array FPGA implementation of QR decomposition. ACM Transactions on Embedded Computing Systems, 9(1): 3:1–3:??, October 2009. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Wei:2018:SAE

[WLC^{+}18] Yi-Hung Wei, Quan Leng, Wei-Ju Chen, Aloysius K. Mok, and Song Han. Schedule adaptation for ensuring reliability in RT-WiFi-based networked embedded systems. ACM Transactions on Embedded Computing Systems, 17(5):85:1– 85:??, November 2018. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3236011.

Wei:2022:SBD

[WLC^{+}22] Qian Wei, Bingzhe Li, Wanli Chang, Zhiping Jia, Zhaoyan Shen, and Zili Shao. A survey of blockchain data management systems. ACM Transactions on Embedded Computing Systems, 21(3): 25:1–25:28, May 2022. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3502741.

Wang:2016:UUP

[WLH16] Yichuan Wang, Xin Liu, and Cheng-Hsin Hsu. UPDATE: User-Profile-Driven Adaptive TransfEr for mobile devices. ACM Transactions on Embedded Computing Systems, 15(3):42:1–42:??, July 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Wu:2018:RRA

[WLH^{+}18] Die Wu, Li Lu, Muhammad Jawad Hussain, Songfan Li, Mo Li, and Fengli Zhang. R^{3}: Reliable over-the-air reprogramming on computational RFIDs. ACM Transactions on Embedded Computing Systems, 17(1):9:1– 9:??, January 2018. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Wang:2018:TAB

[WLHC18] Zejue Wang, Hongjia Li, Dan Hu, and Song Ci. Transmission adaptation for batteryfree relaying. ACM Transactions on Embedded Computing Systems, 17(1):5:1– 5:??, January 2018. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Wijerathne:2019:CHT

[WLK^{+}19]  Dhananjaya  Wijerathne,

Zhaoying Li, Manupa Karunarathne, Anuj Pathania, and Tulika Mitra. CASCADE: High throughput data streaming via decoupled access-execute CGRA. ACM Transactions on Embedded Computing Systems, 18(5s):50:1– 50:??, October 2019. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3358177.

Woehrle:2012:CTC

[WLT12] Matthias Woehrle, Kai Lampka, and Lothar Thiele. Conformance testing for cyber-physical systems. ACM Transactions on Embedded Computing Systems, 11(4): 84:1–84:??, December 2012. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Wu:2024:SGC

[WLTW24] Chin-Hsien Wu, Cheng-Tze Lee, Yi-Ren Tsai, and Cheng-Yen Wu. A spacegrained cleaning method to reduce long-tail latency of DM-SMR disks. ACM Transactions on Embedded Computing Systems, 23(2): 24:1–24:??, March 2024. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3643827.

Wang:2024:LBR

[WLW^{+}24] Meng Wang, Yiqin Lu, Haihan Wang, Zhuoxing Chen, and Jiancheng Qin. Loadbalanced routing heuristics for bandwidth allocation of AVB flow in TSN. ACM Transactions on Embedded Computing Systems, 23(5): 83:1–83:??, September 2024. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3687307.

Wang:2015:TWA

[WLWS15] Tianzheng Wang, Duo Liu, Yi Wang, and Zili Shao. Towards write-activity-aware page table management for non-volatile main memories. ACM Transactions on Embedded Computing Systems, 14(2):34:1–34:??, March 2015. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Wang:2023:EEM

[WLZ^{+}23] Yitu Wang, Shiyu Li, Qilin Zheng, Andrew Chang, Hai Li, and Yiran Chen. EMSi: an efficient memory system design with specialized caching mechanism for recommendation inference. ACM Transactions on Embedded Computing Systems, 22(5s):100:1–100:??, October 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3609384.

Weerasena:2024:RCA

[WM24] Hansika Weerasena and Prabhat Mishra. Revealing CNN architectures via sidechannel analysis in dataflowbased inference accelerators. ACM Transactions on Embedded Computing Systems, 23(6):100:1–100:??, November 2024. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL

https://dl.acm.org/doi/
10.1145/3688001.

Wang:2012:DCR

[WMGR12] Weixun Wang, Prabhat Mishra, and Ann Gordon-Ross. Dynamic cache reconfiguration for soft realtime systems. ACM Transactions on Embedded Computing Systems, 11(2):28:1– 28:??, July 2012. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Weinstock:2016:PSS

[WMLA16] Jan Henrik Weinstock, Luis Gabriel Murillo, Rainer Leupers, and Gerd Ascheid. Parallel SystemC simulation for ESL design. ACM Transactions on Embedded Computing Systems, 16(1):27:1– 27:??, November 2016. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Wongpiromsarn:2012:VPC

[WMLM12] Tichakorn Wongpiromsarn, Sayan Mitra, Andrew Lamperski, and Richard M. Murray. Verification of periodically controlled hybrid systems: Application to an autonomous vehicle. ACM Transactions on Embedded Computing Systems, 11(S2): 53:1–53:??, 2012. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Wang:2017:TAS

[WMRB17] Jiajie Wang, Michael Mendler, Partha Roop, and Bruno Bodin. Timing analysis of synchronous programs using WCRT algebra: Scalability through abstraction. ACM Transactions on Embedded Computing Systems, 16(5s):177:1–177:??, October 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Wandeler:2012:UGS

[WMT12] Ernesto Wandeler, Alexander Maxiaguine, and Lothar Thiele. On the use of greedy shapers in real-time embedded systems. ACM Transactions on Embedded Computing Systems, 11(1):1:1– 1:??, March 2012. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Wu:2013:AMC

[WMZY13] Yulei Wu, Geyong Min, Dakai Zhu, and Laurence T. Yang. An analytical model for on-chip interconnects in multimedia embedded systems. ACM Transactions on Embedded Computing Systems, 13(1s):29:1–29:??, November 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Wolf:2002:III

[Wol02] Wayne Wolf. Introduction to the inaugural issue. ACM Transactions on Embedded Computing Systems, 1(1): 1, November 2002. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

West:2011:ASS

[WP11] Richard West and Gabriel Parmer. Application-specific service technologies for commodity operating systems in real-time environments. ACM Transactions on Embedded Computing Systems, 10(3):30:1–30:??, April 2011. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Wollinger:2004:EHC

[WPW^{+}04] Thomas Wollinger, Jan Pelzl, Volker Wittelsberger, Christof Paar, Gökay Saldamli, and Çetin K. Koç. Elliptic and hyperelliptic curves on embedded μP. ACM Transactions on Embedded Computing Systems, 3(3):509–533, August 2004. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Wen:2022:SHD

[WQGR22] Fei Wen, Mian Qin, Paul Gratz, and Narasimha Reddy. Software hint-driven data management for hybrid memory in mobile systems. ACM Transactions on Embedded Computing Systems, 21(1):8:1–8:18, January 2022. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3494536.

Watkins:2015:UNT

[WRB15] Lanier Watkins, William H. Robinson, and Raheem Beyah. Using network traffic to infer hardware state: a kernel-level investigation. ACM Transactions on Embedded Computing Systems, 14(3):55:1–55:??, May 2015. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Wu:2006:EEU

[WRJL06] Haisang Wu, Binoy Ravindran, E. Douglas Jensen, and Peng Li. Energy-efficient, utility accrual scheduling under resource constraints for mobile embedded systems. ACM Transactions on Embedded Computing Systems, 5(3):513–542, August 2006. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Wehner:2016:SRM

[WRKG16] Philipp Wehner, Jens Rettkowski, Tobias Kalb, and Diana Göhringer. Simulating reconfigurable multiprocessor systems-on-chip with MPSoCSim. ACM Transactions on Embedded Computing Systems, 16(1): 4:1–4:??, November 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Wang:2021:VSH

[WRW^{+}21] Yu Wang, Nima Roohi, Matthew West, Mahesh Viswanathan, and Geir E. Dullerud. Verifying stochastic hybrid systems with temporal logic specifications via model reduction. ACM Transactions on Embedded Computing Systems, 20 (6):113:1–113:27, November 2021. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3483380.

Wu:2014:EIE

[WSHC14] I-Wei Wu, Jean Jyh-Jiun Shann, Wei-Chung Hsu, and Chung-Ping Chung. Extended instruction exploration for multiple-issue architectures. ACM Transactions on Embedded Computing Systems, 13(4):92:1– 92:??, February 2014. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Wu:2014:EDF

[WSK14] Chenye Wu, Yiyu Shi, and Soummya Kar. Exploring demand flexibility in heterogeneous aggregators: an LMPbased pricing scheme. ACM Transactions on Embedded Computing Systems, 13(2s): 57:1–57:??, January 2014. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Wittig:2022:AES

[WSMF22] Robert Wittig, Philipp Schulz, Emil Matus, and Gerhard P. Fettweis. Accurate estimation of service rates in interleaved scratchpad memory systems. ACM Transactions on Embedded Computing Systems, 21(1): 4:1–4:15, January 2022. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3457171.

Wu:2015:SDE

[WT15] Ming-Ju Wu and Chun-Jen Tsai. A storage device emulator for system performance evaluation. ACM Transactions on Embedded Computing Systems, 14(4): 82:1–82:??, December 2015. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Waluyo:2013:MQS

[WTSR13] Agustinus Borgy Waluyo, David Taniar, Bala Srinivasan, and Wenny Rahayu. Mobile query services in a participatory embedded sensing environment. ACM Transactions on Embedded Computing Systems, 12(2): 31:1–31:??, February 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Wu:2010:SAF

[Wu10] Chin-Hsien Wu. A selfadjusting flash translation layer for resource-limited embedded systems. ACM Transactions on Embedded Computing Systems, 9(4): 31:1–31:??, March 2010. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Weng:2009:AMN

[WW09] Ning Weng and Tilman Wolf.

Analytic modeling of network processors for parallel workload mapping. ACM Transactions on Embedded Computing Systems, 8(3): 18:1–18:??, April 2009. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Weichslgartner:2018:DTR

[WWG^{+}18]  Andreas  Weichslgartner,

Stefan Wildermann, Deepak Gangadharan, Michael Glaß, and Jürgen Teich. A design– time/run-time application mapping methodology for predictable execution time in MPSoCs. ACM Transactions on Embedded Computing Systems, 17(5):89:1– 89:??, November 2018. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3274665.

Witterauf:2021:SLC

[WWHT21] Michael Witterauf, Dominik Walter, Frank Hannig, and Jürgen Teich. Symbolic loop compilation for tightly coupled processor arrays. ACM Transactions on Embedded Computing Systems, 20(5): 49:1–49:31, July 2021. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3466897.

Wen:2023:WCP

[WWN23] Elliott Wen, Gerald Weber, and Suranga Nanayakkara. WasmAndroid: a crossplatform runtime for native programming languages on Android. ACM Transactions on Embedded Computing Systems, 22(1):4:1– 4:??, January 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3530286.

Wang:2022:RID

[WWT^{+}22] Tse-Yuan Wang, Chun-Feng Wu, Che-Wei Tsao, Yuan-Hao Chang, Tei-Wei Kuo, and Xue Liu. Rethinking the interactivity of OS and device layers in memory management. ACM Transactions on Embedded Computing Systems, 21(4): 42:1–42:??, July 2022. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3530876.

Wang:2019:DES

[WWTSM19] Youchao Wang, Sam Willis, Vasileios Tsoutsouras, and Phillip Stanley-Marbell. Deriving equations from sensor data using dimensional function synthesis. ACM Transactions on Embedded Computing Systems, 18(5s):84:1– 84:??, October 2019. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

tronic). URL https://dl.

                                        acm.org/ft_gateway.cfm?
                                        id=3358218.

Wang:2013:DLE

[WWY13] Shouguang Wang, Chengying Wang, and Yanping Yu. Design of livenessenforcing supervisors for S3PR based on complementary places. ACM Transactions on Embedded Computing Systems, 12(1):2:1– 2:??, January 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Wolf:2017:GES

[WX17] Marilyn Wolf and Jason Xue. Guest editorial: Special issue on embedded computing for IoT. ACM Transactions on Embedded Computing Systems, 16(3):63:1–63:??, July 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Wang:2017:SRS

[WXY^{+}17] Wei Wang, Peng Xu, Laurence Tianruo Yang, Willy Susilo, and Jinjun Chen. Securely reinforcing synchronization for embedded online contests. ACM Transactions on Embedded Computing Systems, 16(2):58:1– 58:??, April 2017. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Wang:2018:STW

[WXY^{+}18] Zhongqin Wang, Fu Xiao, Ning Ye, Ruchuan Wang, and Panlong Yang. A see-through-wall system for device-free human motion sensing based on batteryfree RFID. ACM Transactions on Embedded Computing Systems, 17(1):6:1– 6:??, January 2018. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Wang:2014:STN

[WYJ^{+}14] Xiaohang Wang, Mei Yang, Yingtao Jiang, Peng Liu, Masoud Daneshtalab, Maurizio Palesi, and Terrence Mak. On self-tuning networks-on-chip for dynamic network-flow dominance adaptation. ACM Transactions on Embedded Computing Systems, 13(2s): 73:1–73:??, January 2014. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Wang:2019:NNA

[WYL^{+}19] Xiaokang Wang, Laurence T.

Yang, Hongguo Li, Man Lin, Jianjun Han, and Bernady O. Apduhan. NQA: a nested anti-collision algorithm for RFID systems. ACM Transactions on Embedded Computing Systems, 18(4):32:1–32:??, August 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

                                        /dl.acm.org/ft_gateway.
                                        cfm?id=3330139.

Waluyo:2010:MMB

[WYP^{+}10] Agustinus Borgy Waluyo, Wee-Soon Yeoh, Isaac Pek, Yihan Yong, and Xiang Chen. MobiSense: Mobile body sensor network for ambulatory monitoring. ACM Transactions on Embedded Computing Systems, 10(1): 13:1–13:??, August 2010. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Wouhaybi:2013:ECM

[WYS^{+}13] Rita H. Wouhaybi, Mark D. Yarvis, Sangita Sharma, Philip Muse, Chieh-Yih Wan, Sai Prasad, Lenitra Durham, Ritu Sahni, Robert Norton, Merlin Curry, Holly Jimison, Richard Harper, and Robert A. Lowe. Experiences with context management in emergency medicine. ACM Transactions on Embedded Computing Systems, 12(4):100:1–100:??, June 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Wu:2012:MCB

[WZ12] Lan Wu and Wei Zhang. A model checking based approach to bounding worstcase execution time for multicore processors. ACM Transactions on Embedded Computing Systems, 11(S2): 56:1–56:??, 2012. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Wang:2019:SVH

[WZBP19] Yu Wang, Mojtaba Zarei, Borzoo Bonakdarpour, and Miroslav Pajic. Statistical verification of hyperproperties for cyber-physical systems. ACM Transactions on Embedded Computing Systems, 18(5s):92:1– 92:??, October 2019. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3358232.

Wu:2017:PIE

[WZD^{+}17] Fei Wu, Meng Zhang, Yajuan Du, Xubin He, Ping Huang, Changsheng Xie, and Jiguang Wan. A program interference error aware LDPC scheme for improving NAND flash decoding performance. ACM Transactions on Embedded Computing Systems, 16(5s):141:1–141:??, October 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Wu:2023:EAD

[WZG^{+}23] Yirui Wu, Lilai Zhang, Zonghua Gu, Hu Lu, and Shaohua Wan. Edge-AI-driven framework with efficient mobile network design for facial expression recognition. ACM Transactions on Embedded Computing Systems, 22(3):57:1–57:??, May 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3587038.

Wu:2013:OSL

[WZH13] Naiqi Wu, Mengchu Zhou, and Gang Hu. One-step look-ahead maximally permissive deadlock control of AMS by using Petri nets. ACM Transactions on Embedded Computing Systems, 12(1):10:1–10:??, January 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Wang:2018:SLL

[WZJ^{+}18] Guan Wang, Chuanqi Zang, Lei Ju, Mengying Zhao, Xiaojun Cai, and Zhiping Jia. Shared last-level cache management and memory scheduling for GPGPUs with hybrid main memory. ACM Transactions on Embedded Computing Systems, 17(4): 77:1–77:??, August 2018. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Wang:2017:CEG

[WZM17] Siqi Wang, Guanwen Zhong, and Tulika Mitra. CG-Predict: Embedded GPU performance estimation from single-threaded applications. ACM Transactions on Embedded Computing Systems, 16(5s):146:1–146:??, October 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Wang:2024:HSD

[WZS^{+}24] Xingbin Wang, Boyan Zhao, Yulan Su, Sisi Zhang, Fengkai Yuan, Jun Zhang, Dan Meng, and Rui Hou. A hybrid sparse-dense defensive DNN accelerator architecture against adversarial example attacks. ACM Transactions on Embedded Computing Systems, 23(5): 79:1–79:??, September 2024. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3677318.

Wang:2023:FCF

[WZY^{+}23] Xiaoyang Wang, Zhe Zhou, Zhihang Yuan, Jingchen Zhu, Yulong Cao, Yao Zhang, Kangrui Sun, and Guangyu Sun. FD-CNN: a frequency-domain FPGA acceleration scheme for CNNbased image-processing applications. ACM Transactions on Embedded Computing Systems, 22(6):91:1– 91:??, November 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3559105.

Xia:2023:SPS

[XCZW23] Chengpeng Xia, Yawen Chen, Haibo Zhang, and Jigang Wu. STADIA: Photonic stochastic gradient descent for neural network accelerators. ACM Transactions on Embedded Computing Systems, 22(5s):126:1– 126:??, October 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3607920.

Xie:2018:ESA

[XDL^{+}18] Xinfeng Xie, Dayou Du, Qian Li, Yun Liang, Wai Teng Tang, Zhong Liang Ong, Mian Lu, Huynh Phung Huynh, and Rick Siow Mong Goh. Exploiting sparsity to accelerate fully connected layers of CNN-based applications on mobile SoCs. ACM Transactions on Embedded Computing Systems, 17(2): 37:1–37:??, April 2018. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Xi:2016:FSS

[XHK16] Kai Xi, Jiankun Hu, and B. V. K. Vijaya Kumar. FE-SViT: a SViT-based fuzzy extractor framework. ACM Transactions on Embedded Computing Systems, 15(4): 78:1–78:??, August 2016.

CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Xu:2024:ODI

[XHL^{+}24] Runqing Xu, Debiao He, Min Luo, Cong Peng, and Xiangyong Zeng. Optimizing Dilithium implementation with AVX2/-512. ACM Transactions on Embedded Computing Systems, 23(6): 98:1–98:??, November 2024. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3687309.

Xue:2010:IRP

[XHSS10] Chun Jason Xue, Jingtong Hu, Zili Shao, and Edwin Sha. Iterational retiming with partitioning: Loop scheduling with complete memory latency hiding. ACM Transactions on Embedded Computing Systems, 9(3):22:1–22:??, February 2010. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Xu:2017:AFA

[XKK17] Ye Xu, Israel Koren, and C. Mani Krishna. AdaFT: a framework for adaptive fault tolerance for cyberphysical systems. ACM Transactions on Embedded Computing Systems, 16(3): 79:1–79:??, July 2017. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Xu:2018:GEP

[XLY18] Chi Xu, Wei Liang, and Haibin Yu. Green-energy-powered cognitive radio networks: Joint time and power allocation. ACM Transactions on Embedded Computing Systems, 17(1):13:1– 13:??, January 2018. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Xie:2007:ISP

[XQ07] Tao Xie and Xiao Qin. Improving security for periodic tasks in embedded systems through scheduling. ACM Transactions on Embedded Computing Systems, 6(3):20:1–20:??, July 2007. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Xu:2024:SLB

[XQQ^{+}24] Haitao Xu, Saiyu Qi, Yong Qi, Wei Wei, and Naixue Xiong. Secure and lightweight blockchain-based truthful data trading for real-time vehicular crowdsensing. ACM Transactions on Embedded Computing Systems, 23(1):7:1– 7:??, January 2024. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3582008.

Xiao:2022:CIA

[XSP22] Jun Xiao, Yixian Shen, and Andy D. Pimentel. Cache interference-aware task partitioning for non-preemptive real-time multi-core systems. ACM Transactions on Embedded Computing Systems, 21(3):28:1–28:28, May 2022. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3487581.

Xu:2006:DMA

[XWHC06] Jiang Xu, Wayne Wolf, Joerg Henkel, and Srimat Chakradhar. A design methodology for applicationspecific networks-on-chip. ACM Transactions on Embedded Computing Systems, 5(2):263–280, May 2006. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Xia:2025:TBA

[XWL25] Zhuoqun Xia, Ziyu Wang, and Xiao Liu. Trust based active game data collection scheme in smart cities. ACM Transactions on Embedded Computing Systems, 24(1): 3:1–3:??, January 2025. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3677319.

Xu:2023:LVL

[XYLC23] Zirui Xu, Fuxun Yu, Chenchen Liu, and Xiang Chen. LanCeX: a versatile and lightweight defense method against condensed adversarial attacks in image and audio recognition. ACM Transactions on Embedded Computing Systems, 22(1):14:1– 14:??, January 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3555375.

Xie:2019:EWA

[XZK^{+}19] Guoqi Xie, Gang Zeng, Ryo Kurachi, Hiroaki Takada, Renfa Li, and Keqin Li. Exact WCRT analysis for message-processing tasks on gateway-integrated invehicle CAN clusters. ACM Transactions on Embedded Computing Systems, 17(6): 95:1–95:??, January 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Yang:2012:UEP

[YC12] Yoon Seok Yang and Gwan Choi. Unequal error protection based on DVFS for JSCD in low-power portable multimedia systems. ACM Transactions on Embedded Computing Systems, 11(2): 30:1–30:??, July 2012. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

You:2016:VVA

[YC16] Yi-Ping You and Szu-Chien Chen. VecRA: a vectoraware register allocator for GPU shader processors. ACM Transactions on Embedded Computing Systems, 15(4):64:1–64:??, August 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Chiou:2005:SAS

[yCBR05] Lih yih Chiou, Swarup Bhunia, and Kaushik Roy. Synthesis of application-specific highly efficient multi-mode cores for embedded systems. ACM Transactions on Embedded Computing Systems, 4(1):168–188, February 2005. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Ye:2024:HEA

[YCJW24] Chongnan Ye, Meng Chen, Qisheng Jiang, and Chundong Wang. Hercules: Enabling atomic durability for persistent memory with transient persistence domain. ACM Transactions on Embedded Computing Systems, 23(6):87:1– 87:??, November 2024. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3607473.

Yassin:2018:AAC

[YCK^{+}18] Yahya H. Yassin, Francky Catthoor, Fabian Kloosterman, Jyh-Jang Sun, JoãO Couto, Per Gunnar Kjeldsberg, and Nick Van Helleputte. Algorithm/architecture cooptimisation technique for automatic data reduction of wireless read-out in highdensity electrode arrays. ACM Transactions on Embedded Computing Systems, 17(3):67:1–67:??, June 2018. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Ykman-Couvreur:2002:SLE

[YCLV^{+}02] Ch. Ykman-Couvreur, J. Lambrecht, A. Van Der Togt, F. Catthoor, and H. De Man. System-level exploration of association table implementations in telecom network applications. ACM Transactions on Embedded Computing Systems, 1(1):106– 140, November 2002. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Ykman-Couvreur:2011:FMM

[YCNCC11] Ch. Ykman-Couvreur, V. Nollet, F. Catthoor, and H. Corporaal. Fast multidimension multichoice knapsack heuristic for MP-SoC runtime management. ACM Transactions on Embedded Computing Systems, 10(3): 35:1–35:??, April 2011. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Yang:2016:BAU

[YCT16] Ming-Chang Yang, Yuan-Hao Chang, and Che-Wei Tsao. Byte-addressable update scheme to minimize the energy consumption of PCM-based storage systems. ACM Transactions on Embedded Computing Systems, 15(3):55:1–55:??, July 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Yang:2010:HPO

[YDLC10a] Lei Yang, Robert P. Dick, Haris Lekatsas, and Srimat Chakradhar. Highperformance operating system controlled online memory compression. ACM Transactions on Embedded Computing Systems, 9(4): 30:1–30:??, March 2010. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Yang:2010:OMC

[YDLC10b] Lei Yang, Robert P. Dick, Haris Lekatsas, and Srimat Chakradhar. Online memory compression for embedded systems. ACM Transactions on Embedded Computing Systems, 9(3):27:1– 27:??, February 2010. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Yuan:2022:MFC

[YDS^{+}22] Geng Yuan, Peiyan Dong, Mengshu Sun, Wei Niu, Zhengang Li, Yuxuan Cai, Yanyu Li, Jun Liu, Weiwen Jiang, Xue Lin, Bin Ren, Xulong Tang, and Yanzhi Wang. Mobile or FPGA? A comprehensive evaluation on energy efficiency and a unified optimization framework. ACM Transactions on Embedded Computing Systems, 21(5):65:1–65:??, September 2022. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3528578.

Yantir:2017:AMM

[YEK17] Hasan Erdem Yantir, Ahmed M. Eltawil, and Fadi J. Kurdahi. Approximate memristive in-memory computing. ACM Transactions on Embedded Computing Systems, 16(5s):129:1–129:??, October 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Yaghoubi:2019:WCS

[YF19] Shakiba Yaghoubi and Georgios Fainekos. Worst-case satisfaction of STL specifications using feedforward neural network controllers: a Lagrange multipliers approach.

ACM Transactions on Embedded Computing Systems, 18(5s):107:1–107:??, October 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/ft_gateway.
cfm?id=3358239.

Yang:2014:CTR

[YFPJ14] Qiang Yang, Jian Fu, Raphael Poss, and Chris Jesshope. On-chip traffic regulation to reduce coherence protocol cost on a microthreaded many-core architecture with distributed caches. ACM Transactions on Embedded Computing Systems, 13(3s):103:1– 103:??, March 2014. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Yang:2002:FVL

[YG02] Jun Yang and Rajiv Gupta.

Frequent value locality and its applications. ACM Transactions on Embedded Computing Systems, 1(1):79–105, November 2002. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Yuce:2017:AFI

[YGD^{+}17] Bilgiday Yuce, Nahid Farhady Ghalaty, Chinmay Deshpande, Harika Santapuri, Conor Patrick, Leyla Nazhandali, and Patrick Schaumont. Analyzing the fault injection sensitivity of secure embedded software. ACM Transactions on Embedded Computing Systems, 16(4):95:1– 95:??, August 2017. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Yan:2019:CAR

[YGD^{+}19] Yin Yan, Girish Gokul, Karthik Dantu, Steven Y. Ko, Lukasz Ziarek, and Jan Vitek. Can Android run on time? Extending and measuring the Android platform’s timeliness. ACM Transactions on Embedded Computing Systems, 17(6): 97:1–97:??, January 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Yan:2008:DOD

[YGHS08] Ting Yan, Yu Gu, Tian He, and John A. Stankovic. Design and optimization of distributed sensing coverage in wireless sensor networks. ACM Transactions on Embedded Computing Systems, 7(3):33:1–33:??, April 2008. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Yang:2012:PAA

[YGW^{+}12] Shengqi Yang, Pallav Gupta, Marilyn Wolf, Dimitrios Serpanos, Vijaykrishnan Narayanan, and Yuan Xie. Power analysis attack resistance engineering by dynamic voltage and frequency scaling. ACM Transactions on Embedded Computing Systems, 11(3):62:1– 62:??, September 2012. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Young:2023:CAD

[YHL23] May Young, Alan J. Hu, and Guy G. F. Lemieux. Cache abstraction for data race detection in heterogeneous systems with non-coherent accelerators. ACM Transactions on Embedded Computing Systems, 22(1):6:1– 6:??, January 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3535457.

Yan:2017:FFI

[YJD^{+}17] Hao Yan, Lei Jiang, Lide Duan, Wei-Ming Lin, and Eugene John. FlowPaP and FlowReR: Improving energy efficiency and performance for STT-MRAM-based handheld devices under read disturbance. ACM Transactions on Embedded Computing Systems, 16(5s): 132:1–132:??, October 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Yun:2003:EOV

[YK03] Han-Saem Yun and Jihong Kim. On energy-optimal voltage scheduling for fixedpriority hard real-time systems. ACM Transactions on Embedded Computing Systems, 2(3):393–430, August 2003. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Yu:2024:MAE

[YKD^{+}24] Wangyang Yu, Jinming Kong, Zhijun Ding, Xiaojun Zhai, Zhiqiang Li, and Qi Guo. Modeling and analysis of ETC control system with colored Petri net and dynamic slicing. ACM Transactions on Embedded Computing Systems, 23(1): 14:1–14:??, January 2024. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3633450.

Yun:2013:DHS

[YKK^{+}13] Jeong-Han Yun, Chul-Joo Kim, Seonggun Kim, Kwang-Moo Choe, and Taisook Han. Detection of harmful schizophrenic statements in Esterel. ACM Transactions on Embedded Computing Systems, 12(3):80:1– 80:??, March 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Yi:2023:EEE

[YKKD23] Saehanseul Yi, Tae-Wook Kim, Jong-Chan Kim, and Nikil Dutt. EASYR: Energyefficient adaptive system reconfiguration for dynamic deadlines in autonomous driving on multicore processors. ACM Transactions on Embedded Computing Systems, 22(3):52:1–52:??, May 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3570503.

Yu:2019:TND

[YLDM19] Jiecao Yu, Andrew Lukefahr, Reetuparna Das, and Scott Mahlke. TF-Net: Deploying sub-byte deep neural networks on microcontrollers. ACM Transactions on Embedded Computing Systems, 18(5s):45:1–45:??, October 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/ft_gateway.
cfm?id=3358189.

Yuan:2021:CCB

[YLTY21] Bo Yuan, Xiaofen Lu, Ke Tang, and Xin Yao. Cooperative coevolution-based design space exploration for multi-mode dataflow mapping. ACM Transactions on Embedded Computing Systems, 20(3):21:1–21:25, April 2021. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3440246.

Yang:2015:ESV

[YLW15] Zhengfeng Yang, Wang Lin, and Min Wu. Exact safety verification of hybrid systems based on bilinear SOS representation. ACM Transactions on Embedded Computing Systems, 14(1):16:1– 16:??, January 2015. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Yim:2019:TFS

[YMHB19] Keun Soo Yim, Iliyan Malchev, Andrew Hsieh, and Dave Burke. Treble: Fast software updates by creating an equilibrium in an active software ecosystem of globally distributed stakeholders. ACM Transactions on Embedded Computing Systems, 18(5s):104:1–104:??, October 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

                                        /dl.acm.org/ft_gateway.
                                        cfm?id=3358237.

Yen:2023:KBR

[YMKH23] Chih-Hsuan Yen, Hashan Roshantha Mendis, Tei-Wei Kuo, and Pi-Cheng Hsiu. Keep in balance: Runtime-reconfigurable intermittent deep inference. ACM Transactions on Embedded Computing Systems, 22(5s):124:1–124:??, October 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3607918.

Yu:2010:FSB

[YRF10] Yue Yu, Shangping Ren, and Ophir Frieder. Feasibility of semiring-based timing constraints. ACM Transactions on Embedded Computing Systems, 9(4):33:1– 33:??, March 2010. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Yoong:2012:ICC

[YRS12] Li Hsien Yoong, Partha S.

Roop, and Zoran Salcic. Implementing constrained cyber-physical systems with IEC 61499. ACM Transactions on Embedded Computing Systems, 11(4):78:1– 78:??, December 2012. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Yang:2023:EEP

[YS23] Zhao Yang and Qingshuang Sun. Energy-efficient personalized federated search with graph for edge computing. ACM Transactions on Embedded Computing Systems, 22(5s):99:1–99:??, October 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3609435.

Yang:2022:DDC

[YSC22] Yi Yang, Murugan Sankaradas, and Srimat Chakradhar. DyCo: Dynamic, contextualized AI models. ACM Transactions on Embedded Computing Systems, 21(6): 76:1–76:??, November 2022. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3520131.

Yan:2020:TCH

[YTL^{+}20] Kaige Yan, Jingweijia Tan, Longjun Liu, Xingyao Zhang, Stanko R. Brankovic, Jinghong Chen, and Xin Fu. Toward customized hybrid fuelcell and battery-powered mobile device for individual users. ACM Transactions on Embedded Computing Systems, 18(6):1–20, January 2020. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3362033.

Yin:2023:CBR

[YV23] Jun Yin and Marian Verhelst. CNN-based robust sound source localization with SRP-PHAT for the extreme edge. ACM Transactions on Embedded Computing Systems, 22(3):55:1– 55:??, May 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3586996.

You:2013:EAC

[YW13] Yi-Ping You and Shen-Hong Wang. Energy-aware code motion for GPU shader processors. ACM Transactions on Embedded Computing Systems, 13(3):49:1– 49:??, December 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Yeh:2023:WRR

[YWLW23] Po-Chen Yeh, Chin-Hsien Wu, Yung-Hsiang Lin, and Ming-Yan Wu. A writerelated and read-related DRAM allocation strategy inside solid-state drives (SSDs). ACM Transactions on Embedded Computing Systems, 22(1):17:1– 17:??, January 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3561301.

Yoon:2018:SAF

[YYKK18] Su-Kyung Yoon, Jitae Yun, Jung-Geun Kim, and Shin-Dug Kim. Self-adaptive filtering algorithm with PCMbased memory storage system. ACM Transactions on Embedded Computing Systems, 17(3):69:1–69:??, June 2018. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Yan:2008:AWC

[YZ08] Jun Yan and Wei Zhang. Analyzing the worst-case execution time for instruction caches with prefetching. ACM Transactions on Embedded Computing Systems, 8(1):7:1–7:??, December 2008. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Yu:2013:ANC

[YZA13] Qiaoyan Yu, Meilin Zhang, and Paul Ampadu. Addressing network-on-chip router transient errors with inherent information redundancy. ACM Transactions on Embedded Computing Systems, 12(4):105:1– 105:??, June 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Ye:2023:AAM

[YZZ^{+}23] Wenhua Ye, Xu Zhou, Joey Zhou, Cen Chen, and Kenli Li. Accelerating attention mechanism on FP-GAs based on efficient reconfigurable systolic array. ACM Transactions on Embedded Computing Systems, 22(6):93:1–93:??, November 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3549937.

Yan:2023:MBD

[YZZG23] Zujia Yan, Yi Zhuang, Weining Zheng, and Jingjing Gu. Multi-bit data flow error detection method based on SDC vulnerability analysis. ACM Transactions on Embedded Computing Systems, 22(3):49:1–49:??, May 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3572838.

Zhang:2007:RBP

[ZA07] Wei Zhang and Bramha Allu. Reducing branch predictor leakage energy by exploiting loops. ACM Transactions on Embedded Computing Systems, 6(2):11:1– 11:??, May 2007. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Zou:2022:DHA

[ZAL22] Yu Zou, Amro Awad, and Mingjie Lin. DirectNVM: Hardware-accelerated NVMe SSDs for high-performance embedded computing. ACM Transactions on Embedded Computing Systems, 21(1): 9:1–9:24, January 2022. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3463911.

Zhang:2013:SAE

[ZB13] Fengxiang Zhang and Alan Burns. Schedulability analysis of EDF-scheduled embedded real-time systems with resource sharing. ACM Transactions on Embedded Computing Systems, 12(3): 67:1–67:??, March 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Zhao:2009:STT

[ZBCM09] Weisheng Zhao, Eric Belhaire, Claude Chappert, and Pascale Mazoyer. Spin transfer torque (STT)-MRAM– based runtime reconfiguration FPGA circuit. ACM Transactions on Embedded Computing Systems, 9(2): 14:1–14:??, October 2009. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Zhao:2020:NLD

[ZBG20] Zhuoran Zhao, Kamyar Mirzazad Barijough, and Andreas Gerstlauer. Network-level design space exploration of resource-constrained networksof-systems. ACM Transactions on Embedded Computing Systems, 19(4):22:1– 22:26, July 2020. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3387918.

Zhang:2004:BAP

[ZC04a] Fan Zhang and Samuel T.

Chanson. Blocking-aware processor voltage scheduling for real-time tasks. ACM Transactions on Embedded Computing Systems, 3(2): 307–335, May 2004. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Zhang:2004:DAF

[ZC04b] Ying Zhang and Krishnendu Chakrabarty. Dynamic adaptation for fault tolerance and power management in embedded real-time systems. ACM Transactions on Embedded Computing Systems, 3(2):336–360, May 2004. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Zou:2004:SDT

[ZC04c] Yi Zou and Krishnendu Chakrabarty. Sensor deployment and target localization in distributed sensor networks. ACM Transactions on Embedded Computing Systems, 3(1):61–91, February 2004. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Zhuo:2008:EED

[ZC08] Jianli Zhuo and Chaitali Chakrabarti. Energyefficient dynamic task scheduling algorithms for DVS systems. ACM Transactions on Embedded Computing Systems, 7(2):17:1–17:??, February 2008. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Zhao:2022:CBI

[ZCG^{+}22] Qingling Zhao, Mingqiang Chen, Zonghua Gu, Siyu Luan, Haibo Zeng, and Samarjit Chakrabory. CAN bus intrusion detection based on auxiliary classifier GAN and out-of-distribution detection. ACM Transactions on Embedded Computing Systems, 21(4):45:1– 45:??, July 2022. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3540198.

Zahavi:2013:GNL

[ZCK13] Eitan Zahavi, Israel Cidon, and Avinoam Kolodny. Gana: a novel low-cost conflict-free NoC architecture. ACM Transactions on Embedded Computing Systems, 12(4):109:1–109:??, June 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Zambreno:2005:SOA

[ZCS^{+}05] Joseph Zambreno, Alok Choudhary, Rahul Simha, Bhagi Narahari, and Nasir Memon. SAFE-OPS: an approach to embedded software security. ACM Transactions on Embedded Computing Systems, 4(1):189– 210, February 2005. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Zhu:2022:TUO

[ZDL22] Shien Zhu, Luan H. K.

Duong, and Weichen Liu. TAB: Unified and optimized ternary, binary, and mixed-precision neural network inference on the edge. ACM Transactions on Embedded Computing Systems, 21(5):50:1–50:??, September 2022. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3508390.

Zhong:2019:SHS

[ZDTM19] Guanwen Zhong, Akshat Dubey, Cheng Tan, and Tulika Mitra. Synergy: an HW/SW framework for high throughput CNNs on embedded heterogeneous SoC. ACM Transactions on Embedded Computing Systems, 18(2):13:1–13:??, April 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/ft_gateway.
cfm?id=3301278.

Zeng:2014:MSC

[ZDZ14] Haibo Zeng, Marco Di Natale, and Qi Zhu. Minimizing stack and communication memory usage in real-time embedded applications. ACM Transactions on Embedded Computing Systems, 13(5s):149:1– 149:??, September 2014. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Zacharopoulos:2023:TEH

[ZEJ^{+}23]  Georgios  Zacharopoulos,

Adel Ejjeh, Ying Jing, En-Yu Yang, Tianyu Jia, Iulian Brumar, Jeremy Intan, Muhammad Huzaifa, Sarita Adve, Vikram Adve, Gu-Yeon Wei, and David Brooks. Trireme: Exploration of hierarchical multilevel parallelism for hardware acceleration. ACM Transactions on Embedded Computing Systems, 22(3): 53:1–53:??, May 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3580394.

Ziegler:2019:HSE

[ZGH^{+}19] Andreas Ziegler, Julian Geus, Bernhard Heinloth, Timo Hönig, and Daniel Lohmann. Honey, I shrunk the ELFs: Lightweight binary tailoring of shared libraries. ACM Transactions on Embedded Computing Systems, 18(5s):102:1– 102:??, October 2019. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3358222.

Zhao:2015:RSP

[ZGZ15] Qingling Zhao, Zonghua Gu, and Haibo Zeng. Resource synchronization and preemption thresholds within mixed-criticality scheduling. ACM Transactions on Embedded Computing Systems, 14(4):81:1–81:??, December 2015. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Zhong:2012:SNL

[ZH12a] Ziguo Zhong and Tian He.

Sensor node localization with uncontrolled events. ACM Transactions on Embedded Computing Systems, 11(3):65:1–65:??, September 2012. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Zhong:2012:WSN

[ZH12b] Ziguo Zhong and Tian He.

Wireless sensor node localization by multisequence processing. ACM Transactions on Embedded Computing Systems, 11(1):3:1– 3:??, March 2012. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Zhou:2013:ARD

[ZHCY13] Bo Zhou, Xiaobo Sharon Hu, Danny Z. Chen, and Cedric X. Yu. Accelerating radiation dose calculation: a multi-FPGA solution. ACM Transactions on Embedded Computing Systems, 13(1s): 33:1–33:??, November 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Zhu:2014:CCL

[ZHM^{+}14] Xiuming Zhu, Pei-Chi Huang, Jianyong Meng, Song Han, Aloysius K. Mok, Deji Chen, and Mark Nixon. ColLoc: a collaborative location and tracking system on WirelessHART. ACM Transactions on Embedded Computing Systems, 13(4s):125:1– 125:??, April 2014. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Zhu:2010:RAD

[Zhu10]  Dakai Zhu.  Reliability-

aware dynamic energy management in dependable embedded real-time systems. ACM Transactions on Embedded Computing Systems, 10(2):26:1–26:??, December 2010. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Zheng:2017:RTS

[ZJC^{+}17] Xi Zheng, Christine Julien, Hongxu Chen, Rodion Podorozhny, and Franck Cassez. Real-time simulation support for runtime verification of cyber-physical systems. ACM Transactions on Embedded Computing Systems, 16(4):106:1–106:??, August 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Zhou:2020:BBT

[ZJZL20] Jia Zhou, Prachi Joshi, Haibo Zeng, and Renfa Li. BTMonitor: Bit-time-based intrusion detection and attacker identification in controller area network. ACM Transactions on Embedded Computing Systems, 18(6): 1–23, January 2020. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/doi/abs/10.
1145/3362034.

Zyarah:2025:TSF

[ZK25] Abdullah Zyarah and Dhireesha Kudithipudi. Time-series forecasting and sequence learning using memristorbased reservoir system. ACM Transactions on Embedded Computing Systems, 24(1): 14:1–14:??, January 2025. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3703446.

Zhang:2005:RDC

[ZKKC05] Wei Zhang, Mahmut Kandemir, Mustafa Karakoy, and Guangyu Chen. Reducing data cache leakage energy using a compiler-based approach. ACM Transactions on Embedded Computing Systems, 4(3):652– 678, August 2005. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Zhou:2008:CIA

[ZL08] Ye Zhou and Edward A. Lee. Causality interfaces for actor networks. ACM Transactions on Embedded Computing Systems, 7(3):29:1– 29:??, April 2008. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Zimmerman:2013:MBR

[ZLF13] Andrew T. Zimmerman, Jerome P. Lynch, and Frank T. Ferrese. Marketbased resource allocation for distributed data processing in wireless sensor networks. ACM Transactions on Embedded Computing Systems, 12(3):84:1– 84:??, March 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Zhou:2011:ARA

[ZLL^{+}11] Gang Zhou, Qiang Li, Jingyuan Li, Yafeng Wu, Shan Lin, Jian Lu, Chieh-Yih Wan, Mark D. Yarvis, and John A. Stankovic. Adaptive and radio-agnostic QoS for body sensor networks. ACM Transactions on Embedded Computing Systems, 10(4):48:1– 48:??, November 2011. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Zhang:2018:PEP

[ZLL^{+}18] Jiutian Zhang, Yuhang Liu, Haifeng Li, Xiaojing Zhu, and Mingyu Chen. PTAT: an efficient and precise tool for tracing and profiling detailed TLB misses. ACM Transactions on Embedded Computing Systems, 17(3): 62:1–62:??, June 2018. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Zhou:2019:RTA

[ZLL^{+}19] Quan Zhou, Guohui Li, Jianjun Li, Chenggang Deng, and Ling Yuan. Response time analysis for tasks with fixed preemption points under global scheduling. ACM Transactions on Embedded Computing Systems, 18(5): 111:1–111:??, October 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

/dl.acm.org/ft_gateway.
cfm?id=3360513.

Zhang:2015:MPA

[ZLLC15] Weihua Zhang, Jiaxin Li, Yi Li, and Haibo Chen. Multilevel phase analysis. ACM Transactions on Embedded Computing Systems, 14(2): 31:1–31:??, March 2015. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Zhang:2017:FAK

[ZLSQ17] Jiacheng Zhang, Youyou Lu, Jiwu Shu, and Xiongjun Qin. FlashKV: Accelerating KV performance with open-channel SSDs. ACM Transactions on Embedded Computing Systems, 16(5s): 139:1–139:??, October 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Zheng:2023:IIA

[ZLX^{+}23] Yongchun Zheng, Changlong Li, Yi Xiong, Weihong Liu, Cheng Ji, Zongwei Zhu, and Lichen Yu. iAware: Interaction aware task scheduling for reducing resource contention in mobile systems. ACM Transactions on Embedded Computing Systems, 22(5s):145:1–145:??, October 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3609391.

Zhao:2024:ICV

[ZLZ^{+}24] Liang Zhao, Hongxuan Li, Enchao Zhang, Ammar Hawbani, Mingwei Lin, Shaohua Wan, and Mohsen Guizani. Intelligent caching for vehicular dew computing in poor network connectivity environments. ACM Transactions on Embedded Computing Systems, 23(2):19:1– 19:??, March 2024. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3643038.

Zhu:2007:ESA

[ZM07] Yifan Zhu and Frank Mueller.

Exploiting synchronous and asynchronous DVS for feedback EDF scheduling on an embedded platform. ACM Transactions on Embedded Computing Systems, 7(1): 3:1–3:26, December 2007. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Zhao:2003:SRM

[ZMB03] Qin Zhao, Bart Mesman, and Twan Basten. Static resource models for code-size efficient embedded processors. ACM Transactions on Embedded Computing Systems, 2(2):219–250, May 2003. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Zhai:2013:MSA

[ZNS13] Jiali Teddy Zhai, Hristo Nikolov, and Todor Stefanov. Mapping of streaming applications considering alternative application specifications. ACM Transactions on Embedded Computing Systems, 12(1s):34:1– 34:??, March 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Zhu:2016:SDW

[ZO16] Zhenhuan Zhu and S. Olutunde Oyadiji. Structure design of wireless sensor nodes with energy and cost awareness for multichannel signal measurement. ACM Transactions on Embedded Computing Systems, 15(1): 16:1–16:??, February 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Zhuang:2006:PLS

[ZP06] Xiaotong Zhuang and Santosh Pande. Parallelizing load/stores on dual-bank memory embedded processors. ACM Transactions on Embedded Computing Systems, 5(3):613–657, August 2006. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Zhuang:2007:PEP

[ZP07] Xiaotong Zhuang and Santosh Pande. Power-efficient prefetching for embedded processors. ACM Transactions on Embedded Computing Systems, 6(1):3:1– 3:??, February 2007. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Zhou:2008:DAT

[ZP08] Xiangrong Zhou and Peter Petrov. Direct address translation for virtual memory in energy-efficient embedded systems. ACM Transactions on Embedded Computing Systems, 8(1): 5:1–5:??, December 2008. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Zhou:2009:CLC

[ZP09] Xiangrong Zhou and Peter Petrov. Cross-layer customization for rapid and low-cost task preemption in multitasked embedded systems. ACM Transactions on Embedded Computing Systems, 8(2):14:1– 14:??, January 2009. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Zhuang:2011:CST

[ZP11] Xiaotong Zhuang and Santosh Pande. Compilersupported thread management for multithreaded network processors. ACM Transactions on Embedded Computing Systems, 10(4): 44:1–44:??, November 2011. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Zhao:2017:ORT

[ZPZG17] Yecheng Zhao, Chao Peng, Haibo Zeng, and Zonghua Gu. Optimization of realtime software implementing multi-rate synchronous finite state machines. ACM Transactions on Embedded Computing Systems, 16(5s): 175:1–175:??, October 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Zhou:2024:RRC

[ZQ24] Kunyu Zhou and Keni Qiu. REC: REtime convolutional layers to fully exploit harvested energy for ReRAMbased CNN accelerators. ACM Transactions on Embedded Computing Systems, 23(6):95:1–95:??, November 2024. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3652593.

Zhu:2016:GES

[ZQC16] Dakai Zhu, Meikang Qiu, and Samarjit Chakraborty. Guest editorial: Special issue on emerging technologies in embedded software and systems. ACM Transactions on Embedded Computing Systems, 16(1):9:1– 9:??, November 2016. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Zhao:2023:FSN

[ZQD^{+}23] Hanrui Zhao, Niuniu Qi, Lydia Dehbi, Xia Zeng, and Zhengfeng Yang. Formal synthesis of neural barrier certificates for continuous systems via counterexample guided learning. ACM Transactions on Embedded Computing Systems, 22(5s): 146:1–146:??, October 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3609125.

Zhao:2022:MSM

[ZQGZ22] Qingling Zhao, Mengfei Qu, Zonghua Gu, and Haibo Zeng. Minimizing stack memory for partitioned mixed-criticality scheduling on multiprocessor platforms. ACM Transactions on Embedded Computing Systems, 21(2):20:1– 20:30, March 2022. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3506703.

Zappi:2012:NLP

[ZRF^{+}12] Piero Zappi, Daniel Roggen, Elisabetta Farella, Gerhard Tröster, and Luca Benini. Network-level powerperformance trade-off in wearable activity recognition: a dynamic sensor selection approach. ACM Transactions on Embedded Computing Systems, 11(3):68:1– 68:??, September 2012. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Zhang:2019:CCL

[ZRZ^{+}19] Jeff (Jun) Zhang, Parul Raj, Shuayb Zarar, Amol Ambardekar, and Siddharth Garg. CompAct: On-chip compression of activations for low power systolic array based CNN acceleration. ACM Transactions on Embedded Computing Systems, 18(5s):47:1–47:??, October 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

                                        /dl.acm.org/ft_gateway.
                                        cfm?id=3358178.

Ziller:2005:CSS

[ZS05] Roberto Ziller and Klaus Schneider. Combining supervisor synthesis and model checking. ACM Transactions on Embedded Computing Systems, 4(2):331–362, May 2005. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Zhou:2021:RAS

[ZSEP21] Yuanbin Zhou, Soheil Samii, Petru Eles, and Zebo Peng. Reliability-aware scheduling and routing for messages in time-sensitive networking. ACM Transactions on Embedded Computing Systems, 20(5):41:1–41:24, July 2021. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3458768.

Zhou:2019:LIN

[ZSH^{+}19] Lu Zhou, Chunhua Su, Zhi Hu, Sokjoon Lee, and Hwajeong Seo. Lightweight implementations of NIST P-256 and SM2 ECC on 8bit resource-constraint embedded device. ACM Transactions on Embedded Computing Systems, 18(3):23:1– 23:??, June 2019. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3236010.

Zhang:2024:SBN

[ZSH^{+}24] Xinqiao Zhang, Mohammad Samragh, Siam Hussain, Ke Huang, and Farinaz Koushanfar. Scalable binary neural network applications in oblivious inference. ACM Transactions on Embedded Computing Systems, 23(3):45:1–45:??, May 2024. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3607192.

Zhu:2012:PAR

[ZSJ12] Jun Zhu, Ingo Sander, and Axel Jantsch. Performance analysis of reconfigurations in adaptive real-time streaming applications. ACM Transactions on Embedded Computing Systems, 11S(1): 12:1–12:??, 2012. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Zhang:2013:RAB

[ZSM13] Fumin Zhang, Zhenwu Shi, and Shayok Mukhopadhyay. Robustness analysis for battery-supported cyberphysical systems. ACM Transactions on Embedded Computing Systems, 12(3): 69:1–69:??, March 2013. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Zhou:2019:LCP

[ZSY19] Lu Zhou, Chunhua Su, and Kuo-Hui Yeh. A lightweight cryptographic protocol with certificateless signature for the Internet of Things. ACM Transactions on Embedded Computing Systems, 18(3): 28:1–28:??, June 2019. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

                                        tronic). URL https://dl.
                                        acm.org/ft_gateway.cfm?
                                        id=3301306.

Zhang:2006:RDL

[ZTD^{+}06] W. Zhang, Y.-F. Tsai, D. Duarte, N. Vijaykrishnan, M. Kandemir, and M. J. Irwin. Reducing dynamic and leakage energy in VLIW architectures. ACM Transactions on Embedded Computing Systems, 5(1):1– 28, February 2006. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Zhou:2003:AMC

[ZTRC03] Huiyang Zhou, Mark C. Toburen, Eric Rotenberg, and Thomas M. Conte. Adaptive mode control: a staticpower-efficient cache design. ACM Transactions on Embedded Computing Systems, 2(3):347–372, August 2003. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Zhu:2019:SEA

[ZTZ^{+}19] Siwen Zhu, Yi Tang, Junxiang Zheng, Yongzhi Cao, Hanpin Wang, Yu Huang, and Marian Margraf. Sample essentiality and its application to modeling attacks on arbiter PUFs. ACM Transactions on Embedded Computing Systems, 18(5):42:1– 42:??, October 2019. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3344148.

Zhang:2004:STC

[ZVL04] Chuanjun Zhang, Frank Vahid, and Roman Lysecky. A self-tuning cache architecture for embedded systems. ACM Transactions on Embedded Computing Systems, 3(2):407–425, May 2004. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Zhang:2005:HCC

[ZVN05] Chuanjun Zhang, Frank Vahid, and Walid Najjar. A highly configurable cache for low energy embedded systems. ACM Transactions on Embedded Computing Systems, 4(2):363–387, May 2005. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Zerzelidis:2010:FFS

[ZW10] Alexandros Zerzelidis and Andy Wellings. A framework for flexible scheduling in the RTSJ. ACM Transactions on Embedded Computing Systems, 10(1):3:1–3:??, August 2010. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Zhang:2013:SCE

[ZW13] Zhiming Zhang and Weimin Wu. Sequence control of essential siphons for deadlock prevention in Petri nets. ACM Transactions on Embedded Computing Systems, 12(1):8:1–8:??, January 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Zheng:2017:DDC

[ZW17] Wenguang Zheng and Hui Wu. Dynamic data-cache locking for minimizing the WCET of a single task. ACM Transactions on Embedded Computing Systems, 16(2):31:1–31:??, April 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Zhang:2016:IRW

[ZWH^{+}16] Daqiang Zhang, Jiafu Wan, Zongjian He, Shengjie Zhao, Ke Fan, Sang Oh Park, and Zhibin Jiang. Identifying region-wide functions using urban taxicab trajectories. ACM Transactions on Embedded Computing Systems, 15(2):36:1–36:??, May 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Zhang:2023:OCS

[ZWK23] Lin Zhang, Zifan Wang, and Fanxin Kong. Optimal checkpointing strategy for real-time systems with both logical and timing correctness. ACM Transactions on Embedded Computing Systems, 22(4):66:1–66:??, July 2023. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3603172.

Zhou:2010:MMS

[ZWY^{+}10] Gang Zhou, Yafeng Wu, Ting Yan, Tian He, Chengdu Huang, John A. Stankovic, and Tarek F. Abdelzaher. A multifrequency MAC specially designed for wireless sensor network applications. ACM Transactions on Embedded Computing Systems, 9(4):39:1–39:??, March 2010. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Zhong:2008:SWE

[ZX08] Xiliang Zhong and Cheng-Zhong Xu. System-wide energy minimization for realtime tasks: Lower bound and approximation. ACM Transactions on Embedded Computing Systems, 7(3): 28:1–28:??, April 2008. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Zhou:2013:GOV

[ZXCH13] Bo Zhou, Kai Xiao, Danny Z. Chen, and X. Sharon Hu. GPU-optimized volume ray tracing for massive numbers of rays in radiotherapy. ACM Transactions on Embedded Computing Systems, 13(3):42:1–42:??, December 2013. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Zhuge:2003:CSR

[ZXS03] Qingfeng Zhuge, Bin Xiao, and Edwin H.-M. Sha. Code size reduction technique and implementation for software-pipelined DSP applications. ACM Transactions on Embedded Computing Systems, 2(4):590– 613, November 2003. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Zeng:2017:SLD

[ZYL^{+}17] Jing Zeng, Laurence T.

Yang, Man Lin, Zili Shao, and Dakai Zhu. Systemlevel design optimization for security-critical cyberphysical-social systems. ACM Transactions on Embedded Computing Systems, 16(2): 39:1–39:??, April 2017. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Zeng:2016:SLM

[ZYM16] Jing Zeng, Laurence T.

Yang, and Jianhua Ma. A system-level modeling and design for cyber-physical-social systems. ACM Transactions on Embedded Computing Systems, 15(2):35:1– 35:??, May 2016. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Zou:2022:APS

[ZZA^{+}22] Yu Zou, Kazi Abu Zubair, Mazen Alwadi, Rakin Muhammad Shadab, Sanjay Gandham, Amro Awad, and Mingjie Lin. ARES: Persistently secure non-volatile memory with processortransparent and hardwarefriendly integrity verification and metadata recovery. ACM Transactions on Embedded Computing Systems, 21(1):11:1–11:32, January 2022. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3492735.

Zhang:2024:EAA

[ZZG24] Yi-Wen Zhang, Hui Zheng, and Zonghua Gu. Energyaware adaptive mixed-criticality scheduling with semi-clairvoyance and graceful degradation. ACM Transactions on Embedded Computing Systems, 23(1):11:1–11:??, January 2024. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://

dl.acm.org/doi/10.1145/
3632749.

Zhang:2015:CDR

[ZZX^{+}15] Shiwen Zhang, Qingquan Zhang, Sheng Xiao, Ting Zhu, Yu Gu, and Yaping Lin. Cooperative data reduction in wireless sensor network. ACM Transactions on Embedded Computing Systems, 14(4):84:1–84:??, December 2015. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Zhu:2012:OTA

[ZZZ^{+}12] Qi Zhu, Haibo Zeng, Wei Zheng, Marco Di Natale, and Alberto Sangiovanni-Vincentelli. Optimization of task allocation and priority assignment in hard realtime distributed systems. ACM Transactions on Embedded Computing Systems, 11(4):85:1–85:??, December 2012. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).